A Complete Bibliography of Publications in the Journal of the History of Ideas

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 April 2025 Version 1.03

Title word cross-reference

$3.00 [Bar42].

1602= [Sha02]. 1602=-1652 [Sha02]. 1628 [Poc78]. 1634= [Sol00]. 1634=-1706 [Sol00]. 1770s [Lau98]. 1794 [Rap21]. 1848 [Ger51]. 1873 [Wie73]. 18th [Ste58]. 1900 [Boa56]. 1904 [Coa61]. 1915 [Whi45a]. 1930s [Fef05, HL19]. 1950 [Wei69]. 1950s [Gal17]. 1960s [Sar14]. 1970s [dS23]. 1976 [Lar78].

57 [Fog12].

6 [Dym21, Ede22].

Ōdō [Nol84].

A. [Boa48a, Ede63, GB87, Hef78, Nic48, Rif53]. A.-L [And02]. Aaron [Ste18]. Abad [Chr23]. Abbé [Per57, And02, Hal74, The23]. Abbot [O’C64]. Abbt [Red97]. ’Abd [Lea80]. Abelard [Hel21a]. Abolition [Hef73, Kea64, Mae73]. Abortion [Dom88]. Abraham [Kja02, Pop06]. Absence [Gor08]. Absent [All75]. Absolute [Gou61, Mir77a, Oak68, Oak98b, Oak98a, Pow70, Rab62]. Absolutism [Keo74]. Absolutist [Mil87]. Abstract [Dow40, McK82b, Yol79b]. Abstraction [Mer53, Mor92, Mor96, Pax93]. Abundance [Mai77]. Abuse [Bea82]. Aby [Gui08, San20, Wen24]. ac [Lee11]. Academic [Mai97, Nob65, Sch02, Val22, Voz25]. Academics [Smi24b]. Academy [And55, vI11, Ano47k]. Acceptance [Sch51]. Accident [Spi61]. Accidents [Koe77]. Accolti [Bla82]. Accommodation [Ben84, Gar13]. according [Doe22]. Account [Bui02, Gil69, Gri12, GS91, Har85a, McK87]. Accountability [Bro06]. Accountants [Ben23]. Accounting [Kel12]. Accounts [Gas00, Sil78]. Achievement [Hun74, Nor47, Mit71a]. Achievements [Str22]. Acknowledgements [Ano57a, Ano58a, Ano59a, Ano70a, Ano71a]. Acknowledgments [Ano62a, Ano63a, Ano65a, Ano67a, Ano68a, Ano69a, Ano73a, Ano75a]. Acosta [Ver22]. Acquaintance [Yol79a]. Acquired [Blo79]. Acrasia [Hug43]. Across [Ver22, BJ19, Ram14, TW21]. Act [Fal88, Ila12]. Acta [Mah15]. Action [Bis52, HTW21, Nak22, Sup54, Vir50, Wil68, Wil73]. Activists [Smi24b]. Acton [Dea72, Lan02, Nur61, Ton59]. Acts [Kem00]. Actuality [Tay98]. Ad [Tur66, Smi67]. Adage [Ede98]. Adam [Tie11, Arr94, Ber74b, Bro92, BS21, Cam20, CR81a, DB77, Har11, Het83, Hil01, Lam74, Lan77, Lev78, Lom02, Mac49, McN73, Meg75, Mos60, Mye75, Nie86, NSR80, Pac02, Pri17a, Qua19, Rap69, Ski76, SRN84, Tun14, Was84, Whi76a]. Adams [Bur87, Gre16, Min65, O’N07, Par88b, Sch83]. Adaptation [Bea82]. Adapting [Hes12]. Addams [Cur61, Sch93]. Addiction [Afr61]. Addison [BB51, BB60, Tae50, Wol49]. Addition [Hat75]. Additional [Bow75b]. Address [Joh42, Kee92]. ’Adi [RS64]. Adler [Lac10]. Admiration [Kau48]. Adolescence [Sum17]. Adorno [Wil23]. Adriatic [De 03]. Advantages [Blu63]. Advent [Doj05]. Adventures [Chi84]. Adventus [Whi71]. Adversary [CJ60]. Advertisements [Ano59b, Ano66a, Ano68b, Ano69b]. Advice [CR81a]. Adviser [LS23]. Advocacy [Sus12]. Advocate [Eki50]. Advocates [Ful94]. Aeneas [Wat72]. Aesthetic [Ber62, Bre61, Del89, Del07, Ger13, Hil66, Kiv67, Kor75, Man88, Man91, Ogd74, Pek99, Spe98, Tow87, Bar88]. Aesthetician [Tho44]. Aesthetics [Bla93, Blo84, Cas88, For03, Jac90, Jaf80, Koc21, Kri51b, Kri52, Ogd49, Wil08, Bea76]. Affairs [Hol21]. Affinity [Mag16]. Affirming [Wil05]. Africa [Mil06, Rob20c]. African [Syp41]. After [Nad91, Rie52, Wal19, Ada49, Han95, Lan12b, O’N12, Pan21, Urb12, Vel19, Wil87]. Afterli [HG21]. Afterlife [Fog12]. Aftermath [Dus81, Nis21]. Again [Pat86, Rab62, Smi67]. Against [Bur74, Dud57, Flo49, Fub96, Gla01, Har99, Mor89, MF06, Sch43b, Whi47, Bed21, D’E07, Gro96, Hia09, Pet15, Sca99]. Agamben [Gre23b]. Agamemnons [Bro47]. Agassiz [Nar05]. Age [Bro61, Coa61, Cro53a, Cru51, Dav83, Deg23, Gor19b, Gri58, Jon40, Kel05b, Mar23, Mul09, Pro72, Ran41, Rig86, Tri56, Twi20, Vie49, Wei19a, Bri00, Sin00, Wen24, Gor08, Gor19a, Mar23, Sty19]. Ageing [Dug15]. Agency [Hol21, Tea62]. Agenda [Mah15]. Ages [Bal56, Boa63, Gri56, Joh60, Lyn00, Bar59b, Cad95, Chr46, Dal80, Deu45, Lac47, San44, Zuc75]. Agostino [Ano67g, Del96, Sch66]. Agrarian [Orr14, Ste53]. Agricola [Han92]. Agricultural [Lov67]. Agrippa [Nau57]. Agro [Dra90]. Ahmad [Har85a]. Aids [Ger72, Hip04, Sha70]. Akenside [Ald44]. Al [Alj19, Fak65, Lea80, Res63, Sta89]. Al-Attas [Alj19]. Al-Farabi [Fak65, Res63]. Al-Jabbar [Lea80]. Al-Kindi [Sta89]. Albert [Was73]. Alberti [Sch23b, Wes69]. Albertus [Cun67]. Albion [Sch53]. Albrecht [Son74]. Alchemical [Jan08]. Alchemies [Sch80b]. Alchemist [Mor41]. Alchemy [Dem72, Lin74]. Alciato [Mon97]. Aldrovandi [Mar23]. Aleramo [Dra90]. Alexander [Ada89, Fis54, Kel80, Low49, dM49]. Alexandria [HP18, Run00, Thi79]. Alexis [Thu76, Wac46]. Alfred [Bar42, Ger15]. Algernon [Con82]. Algorisms [Sch47]. Alice [Pet85]. Alienation [Ham72, Phi83]. Allanic [Ger25]. Allegiance [Ecc76]. Allegory [Elu92]. Allen [Mor86]. Alliance [Ken90]. Allusion [McC42]. Almagest [Sch23b]. Almost [Ros18]. Alois [Gub05]. Altars [She06a, Gre15]. Alternate [Ker24]. Alternative [Kra18, Sny11]. Alternatives [Sch93]. am [Str22]. Amazing [Ned99]. Ambassador [Cha60]. Ambiguities [MH04]. Ambiguity [Par04]. Ambition [Har02a]. Ambivalence [Fer71]. Ambrose [Lov42]. Amelot [Sol00]. America [Gra75, Ard63, BJ19, Cha58b, Cog07, Con60, Dea74, Goo90b, Gua82, Haa48, Hol40, Kla71, Kol91, Leb86, McG13, Mil63, Nad93, Neu77, Ort18, Pal06, Pal09, Stu07, Wel55, Wer72, Wol53, Zea59]. ’American [Wie56, Ber67, Bid54, Coh40, Fis47, Gab56, Pal18b, Sch44, SC40, Whi45a, And21, Ans09, Bok17, Byn84, Cas78, Chi40, Coa61, Coh78, CH15, Cot94, Cre79, Cur42, Cur66, Dav95, Dew41, Dud61, Ent18, Feu59, Fir71, Fri73, Ful94, Gre54, Har93, Hob80, Joh04, Kam70, Kap56, Liu04, Mai89, Man08, Mil74, Mor18, Mor54, Pop52, Que63, Ras82, Rie68, Rik57, Ros12, Rub91, Sch45b, Shu76, Ski76, Sko64, Spi76, Ste51, Sto62, Tuc03, Wal96, Whi47, Zub18]. Americana [Poc87]. Americas [Ham96]. Amerindians [Tun14]. Ames [Gib72]. Amicus [Gue78]. Amin [Har85a]. Among [Sav60, For84, Ioz22, Moy23, Vid95]. amongst [Jos06]. Amoralities [O’B08]. Amos [Moy03b]. Analogical [Car03]. Analogies [Gue83, Gui13, Kli24]. Analogue [Wil66b]. Analogy [Ber87, Kar69, Maz54, Mor57, Osm73]. Analysis [Bec69, Bou03, Gro11, Man42, Moo76, Oak55, Pre79, Tim99, dBJN^{+}20]. Analytic [Bev12, Pie09]. Analytical [Clo75, Sko65, Sir18]. Anarchism [Avr66, Chr23, Mad45, Mil76, Pre54]. Anarchist [Ada16]. Anarchocapitalism [Jen22]. Anatole [Por72]. Anatomy [Cha81, Gar77, Gre69a, Mor16, Rod07]. Anaximander [She00, Nad98]. Ancient [Boo00, Boy02, Cav16, Ces05, Chr46, Doo09, Ede52, Gas55, Gil85, Had52, Joh67a, KM94, Kap08, San44, Ste58, Tur66, Wal11, Wil05, Ber78b, Cal97, Chr54, Cog07, Cra45, Lan12b, McK81a, Ram14, Str16, Urq14, Woo24, Nor74]. Ancients [Bar59a, Bla82, Fon00, Lev91, Ski12, Tod76]. and/as [Wil97]. Andes [Mac06]. Andrea [Mon97]. Andres [Kil61]. Andrew [Bra57]. Anew [Gol88]. Angel [All75]. Angellick [Pat62]. Angelological [Wes53]. Angels [Fix89, Ste75]. Anglican [Ten82]. Anglicanism [Tum96]. Anglicus [Gre80]. Anglo [Hor76, Kli55, Van58]. Anglo-Saxon [Kli55]. Anglo-Saxonism [Hor76]. Anglo-Saxons [Van58]. Anian [Owe75]. Animae [Nan21]. Animal [Gue89, HL65, Kap74, Ser01]. Animals [Har98, Pas75, Wol00]. Animate [Jay70]. Animated [Rob80]. Animation [Dal80]. Animism [Hen01]. Annihilation [Sme20]. Anniversary [Gor19a]. Annotates [Sol00]. annotations [Gar13]. Announcement [Ano41a, Ano42a, Ano46a, Ano46b, Ano56a, Ano57b, Ano68c, Ano70b, Ano80a, Ano80b, Ano83a, Ano92a]. Announcements [Ano45a, Ano47a, Ano52a, Ano58b, Ano59c, Ano67b, Hen44]. Anomy [Orr86]. Anonymous [Dav94, Rap97]. Anquetil [Stu07]. Anselm [Nis21]. Answer [Bar58a, Dus81, Fau78, Hor82, Mos51]. Antal [Mom50]. Ante [Nor47]. Ante-Nicene [Nor47]. Antecedents [Ada42, Cum10]. Anthologizing [Wil20]. Anthony [CH95]. Anthropocene [Cha17]. Anthropologia [Ros18]. Anthropological [Ada91, Von81]. Anthropology [Bat74, CH15, Doj05, Kel84, Sta04, Wil22a]. Anti [Avr66, Can73, Der10, Elu92, Fat05, Gre53, Han18, Hou52, Lie98, Nis21, Pan21, Rob15, Sch03, Sut03, Zub18, vE01, Deu45]. Anti-Catholic [Elu92, Fat05]. Anti-English [Pan21]. Anti-intellectual [Nis21]. Anti-Intellectualism [Avr66, Hou52]. Anti-Jacobins [Sch03]. Anti-Newtonianism [Gre53]. Anti-Puritan [Can73]. Anti-racist [Zub18]. Anti-Religious [Sut03]. Anti-Semitic [Deu45]. Anti-Skeptic [vE01]. Anti-Socialism [Rob15]. Anti-System [Lie98]. Anti-totalitarian [Han18]. Anti-Utopianism [Der10]. Anticipation [Len97]. Anticipations [Hil01]. Anticlericalism [Mac56]. Antidemocratic [Pan21]. Antidote [Ela15]. Antifascism [Cha23a]. Antihumanism [Bia24]. Antinomy [Goo83]. Antioch [Dow40]. Antiqua} [Obe87]. Antiquarianism [Jan08, Kau01, Lat24, Phi96, Sir03]. Antiquarianization [Mil01]. Antiquarians [Moy03a]. Antiqui [Cou87]. Antiquitas [Tri87]. Antiquity [Cal12, Cel01, Cla11, Clo69, Clo71, Cop91, Eld24, Gil69, Lac47, Mac01, Mor98, Str21, Tat63a, Von58, Dod68]. Antirepublicanism [Her92]. Antisemitism [Cla23]. Antislavery [Mor22a]. Antithesis [Mar76, Mue58]. Antoine [Wol07]. Antoine-Yves [Wol07]. Antonello [Boa56]. Antonio [Wil60]. Anxieties [ME21, Pag88]. Apart [Web79]. Apeirotheism [Ano88e, DeA87]. Apes [Gou10]. Aphorism [Rin41]. Aphorisms [Rom61]. Apocalypse [Alm93]. Apocalyptic [Sta08b]. Apologetics [Col76, Har99]. Aporias [Coh62]. Apostle [New95]. Apostles [Ful94]. Apotheoses [Ham96]. Apparent [Bar19]. Apparitions [Jen20]. Appearance [Bro77]. Apperception [Bel15]. Application [Coh42]. Applied [Bou03, Ski76, Spi44]. Appointment [Yin22]. Appreciation [Boc51, Mer89, Mor94]. Approach [Ada79, Bri42, Cro85, Rat41, Sil86]. Approaches [Aer19, Gla90, Gug80]. Appropriation [Del89, Oli74, Sin00]. Approved [Ros75]. Aquinas [Bly86, McK75, Tay98, Wau20]. Arabic [AU19, Ada01, Koc99, Kuk00, Res60, Tru12]. Arabica [Tay98]. Arbitrariness [Lif12]. Arc [Phi71]. Archaeology [Ger25, Goo11, Gri56, Low74, Ger25]. Archetype [Koc21]. Archipelago [Tsa13]. Architectural [Bea83, Mar19]. Architecture [Eig09]. Archive [Bev17, Dun21a]. Archived [Rei17]. Archives [Mon19]. Areas [De 15]. Arendt [Kan13, Sim16, Bri00, Kee12, Moy23]. Areopagite [Wil91]. Areopagus [Bat24]. Arezzo [Wei42]. Argentina [Gri21]. Argument [Dom84, Har99, Iva24, Lan95, Lie56, Nad93, Teg47]. Arguments [Ruk18]. Arian [Pfi97]. Arianism [Pat64]. Aristocracy [Kel65, Sie52]. Aristocratic [Car91]. Aristoteles [Chr45]. Aristotelian [HV02, Hut78, Kap23, Mac98, Pes14]. Aristotelianism [Ned96, Ada01, Lar62, Mal09, Ned91, Sga15]. Aristotelians [Jos06, Wal88]. Aristotelicum [Bos89]. Aristotelis [Del17]. Aristotle [Ano73e, Bir62, Bla22, Elt97, Eva55, Fak65, Gra86, Gre72b, Gue68, Ham59, Joh67a, Joh85, Kin56, Lac65, Lan95, Liz22, McK47, Mul74, Plo53, Rif53, Shu62, Sku58, Smi83, Sol58, Sol78, Sta89, Tod76, VR05, Von81, Wis49]. Arithmetics [Dav60b]. Arminianism [Som67]. Arms [Cha66, Par04]. Army [Ha20]. Arnaldo [Phi96]. Arnall [Hor80]. Arnauld [Nad88, Nad94]. Arnold [Ala67, Ran23]. Aron [Cam86, Dro19]. Ars [Van93]. Art [Ada89, Alt07, Bar75, Bar45, Bri85, Clo69, Clo71, Duf95, Goc99, Gui23, Hai50, Hut87, Jon73, Jor85, Kau01, Kin00, Loe61a, Mad57, Man94, Mat57, Mie73, Mil99a, Mil99b, Mon44, Mor92, Mor96, Mor94, RS64, Rog53, Sar08, Sch56a, Sch47, Sir91, Vid95, Wil23, Wit61, Wol44, HP18, Kri70, Lar78]. Art-Historical [Sch56a]. Artaud [Dem72]. Arthur [Ano48a, Col16, Dig06, Duf80, Fog12, Got75, Man48, Ned13, Str16, Wie73, Wil80]. Article [Ano51a, Car78, Koh69, Mur54]. Articles [Ano44e, Ano44f, Ano45k, Ano45l, Ano45m, Ano45n, Ano49d, Ano52f, Ano80g, Ano80h, Ano80i, Hai50, Kla93]. Articulation [Sta09b]. Artificial [Cot79, Rus75]. Artificialis [Ban14]. Artisans [Kel50]. Artist [Nah47]. Artistic [Bro81, Jen44, Mit71a, Nor47]. Artists [Wit61]. Arts [Bil11, Hia09, HP18, Kri51b, Kri52, Sto46, Tat63a, Tat63b, Wes69, Kiv70]. Arundelliana [Wal11]. Aryan [Coo14]. Asceticism [Bla01]. Ascham [Sta08b]. Ashcraft [Sch89b]. Ashish [Ger25]. Ashmole [Jan08]. Asia [Nak69]. Asian [Hel21b, Jen13, Wes07]. Asiatique [McG03]. Aspect [Che59, Fie76, Hen69]. Aspects [Fre73, Had50, Bin67]. Assault [Spe55, Tay01]. Assent [Moo76]. Assimilation [Ard63]. Assmann [Deg23]. Association [Beh08, Obe76]. Astell [Ell03, Tay01]. Astrological [Wen24]. Astrology [Bro05, Nel76, Pop06]. Astronomical [GS23]. Astronomy [Rid56]. Asylum [OR13]. Asymmetry [NS16]. Athanasius [Bra82]. Atheism [Cri86, Goo02, Roe85, For89]. Athena [Blo96, Kri95]. Athenee [Sta95]. Athenian [KM94, Bat24, Bur81, Sta67a]. Athens [Gut41, Lan12b, McK81a, Vis84]. Atlantic [Jur01, Pir25]. Atlantis [Ada49, For23]. Atomism [Fle63, Goo02, Kar65, Kar66, Sai64, Sea68]. Atomists [Fen86]. Atonal [Dav95]. Attack [Goo18, Nas46, Not11]. Attas [Alj19]. Attempt [Wil51]. Attitude [Ban55, Hel19, Nor47, Rat61, Ros60, Ros71a, Sie50, Wol00, Woo60]. Attitudes [Att83, Ber74a, DM88, Hay59, Nad82, Rie68, Ros43b, Tor84]. Attributed [Rap97, Whi76b, Zah12]. Auctoritates [Mor16]. Audience [Wil66a]. Audiences [Ros92]. Audier [SJB18]. Auerbach [Bre61]. August [Mah15, Kit73, Lid66, Rap21]. Augustan [Joh58a]. Augustans [Hun86]. Auguste [Cas78, Sim65, Ver84]. Augustine [Bur93, Dom88, Har66, Mar72, Maz62, Mom51, Qui92, Ree91]. Aurelius [Afr61]. Aurillac [Tru12]. Austin [Smi62a]. Australia [Bou90]. Australian [Goo64, Tre21]. Authentic [Ros43a]. Authenticity [Cla11]. Author [Col07, Llo69, Vie42]. Authoritarianism [Spe02]. Authorities [CM05]. Authority [Aer19, Axi71, Bou00, Cla11, Lac10, Loc07, McA99, Pek99, Sho79, Von58, Wat72, Yar86]. Authors [Lom02]. Authorship [Dun64, Gro67]. Autobiographical [Lat57]. Autobiography [Bee20]. Automation [Sim16]. Automaton [Cam70, Cap54]. Autonomous [Sou89]. Autonomy [Ces04, Kri73, Roe03, Sei87, Van15, Vic02]. Autrecourt [Wei42]. Avant [Tan98]. Avempace [Fra03, Moo51a, Moo51b]. Averroism [Spi65]. Avikunthak [Ger25]. Avoiding [Sei90]. Awakening [Shu76]. Awarded [Ano47k]. aware [Str22]. Awareness [Cas63]. Away [Blo46, Ver84, Wal98]. Axial [Deg23]. Aztecs [Hil72].

B [Daw93]. B.C. [Bur81]. Bach [Sch46a]. Bachofen [Mal99, Val09]. Back [Mer89, Pet94, Feu59]. Background [Ber67, Cla85, Cou44, Fre64, Gli88, Man57a, Mom51, Nah47, Osb02, Pes14, Pri69, Rou67, Sil86, Von48, Wol44, Mom50]. Backward [Jum12]. Bacon [Bar81, Boa51, Bow56, Hat78, Hor82, Hou41, Kel80, Koc57, Lar62, Lin74, McR57, Mer08, Mil97, Mor77b, Par08, Pes14, Pri54, Ren76, Ser17, Vic92, Vic08, Whi89]. Baconian [Ras85, Coh80, Nad93]. Badiano [Gim08]. Baedekers [Lar10]. Bagehot [Hal58]. Bagya [Goo83]. Bain [Fis54]. Baird [Mil56]. Baker [Cha58a]. Bakhtin [Bra02]. Balazs [Sch65]. Balthasar [Fix89]. Bancroft [Lew65]. Banishing [Bel08]. Baptism [Cam22a]. Baptismal [Col16]. Barbados [Smi17]. Barbarian [Spe55]. Barbarism [Blu63, Isr16a, Ros16a, Val04, Wol16b]. Barbeyrac [Bis15]. Barcan [Str22]. Bare [Wac96]. Barlow [Bla49]. Barnum [Bet59]. Baron [Han95, Tri56, Fer58, Tri56]. Baronio [Maz14]. Baroque [GCM10, Ste40, Wes01]. Barre [Wil19, Sei74, Stu97]. Barrow [Mal97, Pyc87]. Bartas [Gar13]. Barth [Cre95]. Barthelemy [Rin61]. Barthes [Tag86]. Bartholomaeus [Gre80]. Bartolus [Ces04]. Bartram [Cla85, Sil78, Sil86]. Based [Coh94]. Basileus [Che59]. Basis [Aro87, Lov67, Psa60]. Basnage [Elu92]. Battista [Sch23b]. Battle [Atk74, Tin88]. Battles [Bed21]. Baudouin [Ano64j, Kel64, Lyo03]. Bax [Bev93]. Baxter [Bra57]. Bayle [Die61, Bra64, Hic10, Hic18, Isr19, Len97, Len02, Mor04a, O’C84, Sol03a, Tin96, vdL17]. Bayreuth [Lar78]. Be [Ang00, Cas81, Gue86, Str22]. Beach [Col16]. Beard [Dei54, Gol52, Mar53, Phi53, Wie56]. Beast [Shu68]. Beast-Machine [Shu68]. Beatle [Dav69]. Beattie [Wol60]. Beautiful [Gre15]. Beauty [Bro74, Sto61]. Beauvais [Tob74, Zah12]. Beauvoir [Sim90]. Bebel [Lid66]. Beccaria [Mae73]. Beck [Guy02]. Becker [Nel87]. Become [Kim14]. Becomes [Mil06]. Bee [Joh61]. Beethoven [Cas10, SF72]. before [AG95, Bot17, Mil87, Eyc57, Fol17, Gue86, Hor76, Jor59, Kau01, Lev50, Mar10, Pal40, Smi62b, Tig70]. Begging [Spa66]. Beginner [Ioz22]. Beginning [Tri56, dC11]. Beginnings [Bat24, Lum41, MB81]. Begriffsgeschichte [Ano98e, Ric87]. Behavior [Ric79]. Behemoth [Mac70]. behind [Ina20, Mah15]. Behmenistic [Wor49]. Being [And76, Cha17, Cog07, Lov46a, Lyn45, Mah87, Tay98, Wil80, Wil87, Ebe75]. Beings [Goo02]. Bekker [Fix89]. Belfort [Bev93]. Belgium [Tol98]. Belief [Cha58b, Doo99, O’C84, Tut22]. believe [dC11]. Bell [And21]. Belli [Lee11]. Bello [Kil61]. Belloc [Tre21]. Bellum [Nad93]. ben [Nad19]. Beneath [Ger25]. Benedetto [Sch41, Sga15]. Benefits [Jac76a]. Benevolence [Rad93]. Benjamin [Ben82b, D’E69, Fin72, Mae75, Mal99, Pet55, Sin00, Was72]. Bentham [Tay82, Bau45, Coo81, Cri86, Cri94, Cri02, Din75, Ela15, Gol72, Itz78, Par69b, Rod68, Sch07b, Smi70, Wil75]. Benthamism [Coa50a]. Benthamite [Can84]. Bentley [GJ69, Lev89, Smi61, Tin88]. Berdyaev [RG70]. Berengarian [Nis21]. Bergson [Boa59, Gun71, Gun82, Hab93, Low49, McG13, Sin16]. Berkeley [Arm56, Ber71, Ber86b, Bra56, Bra57, Can77, Car07, Con60, Con61, Fri54, Gel52, Gre53, Her70, Lea77, McK82b, McK87, Mir77a, Par69a, Pat86, Pea14, Pyc87, Sil73, Sil76, Sil93, Thr77, Tip69, Tur55, Wen76, Wen82]. Berkeleyan [De 06]. Berlin [Nor08, Ska05]. Bernard [Coh40, Whi60, Ano69l, Car19, Coo99, Gol88, Hun95, Mar59a, Vir86, Wei42]. Bernardin [Men18]. Bernhard [Sta67b]. Bertrand [Boa47, Woo94]. Bertrando [Gri41]. Best [Dev70, Sch83, Yeo03]. Between [Bul05, Col74, Kla93, Mor93, Poc87, Sco98, Sil01, Syp65, Tay17, Wes15, Bec85, Bly86, Bri23, Mil87, CR81b, Hes18, Hun72a, Ken90, Koh56, Lev91, Moo45, Mos60, Nan21, Nau02, Nel00, Pic89, Sac18, Sar14, Sch46a, Sga15, Sho99, Som83, Tou13, Van95, dC11, Ede98, Wel84]. Betwixt [Lyn00]. Beuve [Nel00]. Bevir [Tho12]. Beyond [Gei91, Gra06, Mon44, Pal06, Pal09, Pal18a, Spe98]. Bias [Pre71, Boa64]. Bible [Alb17, Fra13, Kil11, Mil01]. Biblical [Har13, Kov76, Lan66, Mil01, She03, Twi20]. Bibliomancy [Wei19b]. Bibliothéque [Yol79b]. Bicentennial [Pal09, Pal18b]. Big [Coe16]. Bigamy [Mai21]. Bilbao [Spi80]. Bilderdijk [vE01]. Bildung [Sor83]. Bill [San60]. Bills [Gro67]. Biographia [Cor80, Mil99b, Roy07]. Biography [Wit43]. Biological [Gra86, Gue83]. Biology [Ano73e, Gre72b, Jor59, Man57a, Plo53, Rig86, Söl16]. Biran [Hal57]. Birth [AK23, Car74, Fri12, Gro50, Hau11, Jen22, McL76b, Mul06, Spi61, Swe99]. Birthday [Wie73]. Bishop [Cau94, Tip69]. Bishop-Eaters [Cau94]. Bismarck [Hol60]. bist [Nor08]. Black [BS21, Cha23a, Dun21b, Hun78, Gre91c, vG20, Blo96, Kri95]. Blackloism [Sou92, Sou00]. Blackstone [Jez71]. Blanc [Lou56]. Blending [Gou10]. Blökende [Nor08]. Blondel [Eng08]. Bloomer [Mon74]. Bloomsbury [Lan64]. Blount [Red74]. Blue [Rol71]. Blumenberg [Bou87, Mon19, Ada91, Bri00, Gor19a, Gor19b, Hel19, Laz03, Pal97, Sty19]. Blut [Lov67]. Board [Cla15, San20]. Boas [Ano81e, CH15, Gom81, Wie80a, Eli67]. Boccaccio [Sti98]. Boccaccio-Petrarch [Sti98]. Bocchi [Bar75]. Boden [Lov67]. Bodies [Bro18, Woo94, Woo24]. Bodin [Mal17, McR55, McR63, Nic19, Pop88, Smi64, Sol03a]. Body [Bar98, Beh08, Bla93, Gor94, Gud96, Mar76, Mic03, Osm73, Rei96, Sch02, Wil22a, Wil22b, vdL17]. Boétie [Keo77]. Bohemian [Von48]. Boileau [Del07]. Bolingbroke [Hed59, Nad62]. Bonald [KCR46, Nis44, Ree83]. Bonaventura [McK75]. Bongars [Wei16]. Bonnet [And76]. Book [Ack73, Ano55a, Ano64a, Bar47, Bar50, Bar60a, Bar62, Bar42, Bea71, Bea76, Bid54, Bin67, Bla92, Boa47, Boa48b, Boa56, Boa58, Boa63, Boa64, Boa67, BB70, Boa70, Bri42, Bro61, Bur57a, Cas45, Cas48, Cen89, Chi84, Coa50b, Coh40, Coh49, Cri66, Cul88, Del80, Die61, Dod68, Dup74, Dup82, Dyn81, Eck65, Ede52, Ede22, EE70, Eli67, EST66, Fac75, Fel77, Fer68, Fie55, Fis47, Fis70, For89, Fri40, Gab56, Gas55, Ger72, Ger73, Gey48, Gey55, Gil53, Gra75, Gre69b, Gug80, Had52, Had58a, HE63, Ham58, Han91, Har90, HH57, Hea87, Hes12, Hig91, Hil72, Hor78, Jac76a, Jac82, Joh59a, Kah61, KM94, Kas77, KK66, Kel79, Kel95, Kni64, Koh64b, Koh70, Kri70, Kri47, Kri51a, Kri58, LaC88, La 93, Lar78, Lev71, Lev72a, Lev87a, Lin91, Lun77]. Book [Mat43, Met63, Mie73, Mil41, Mil56, Min51, Mit71a, Mit94, Mod94, Mom50, Moo68, Mor57, Mor78b, Mos98, Mos66, Nag68, Nak69, Neg79, Nor74, Pag88, Par62, Phi52, Poc78, Pri58, Pri63, Pri74, Ran41, Ran49, Rie61, Ros57a, Ros74, Ros85a, San59, Sap79, Sar82, Sav44, Sch41, Sch65, Sch40, Sch44, Sch50, Sch89b, Smi65b, Spe48, Spi66, Sta73, Sta74, Tor84, Tri56, Tri90, Tur71, Var73, Vie42, Wal63, War67, Web79, Wei69, Wei19b, Wei72, Wel84, Wes76, Whi45a, Whi76a, Wie43, WW50, Wie51, Wie68, Wie70, Wie75, Wri49, Yeo03, Her05]. Bookburning [Cra45]. Books [Ano40a, Ano40b, Ano40c, Ano41b, Ano41c, Ano41d, Ano41e, Ano42b, Ano42c, Ano42d, Ano42e, Ano43a, Ano43b, Ano43c, Ano43d, Ano44a, Ano44b, Ano44c, Ano44d, Ano45b, Ano45c, Ano45d, Ano45e, Ano46c, Ano46d, Ano46e, Ano46f, Ano47b, Ano47c, Ano47d, Ano47e, Ano48b, Ano48c, Ano48d, Ano48e, Ano49a, Ano49b, Ano49c, Ano50a, Ano50b, Ano52b, Ano52c, Ano52d, Ano53a, Ano54a, Ano54b, Ano55b, Ano56b, Ano56c, Ano56d, Ano57c, Ano57d, Ano57e, Ano57f, Ano58c, Ano58d, Ano58e, Ano58f, Ano59d, Ano59e, Ano59f, Ano59g, Ano60a, Ano61a, Ano61b, Ano61c, Ano62b, Ano62c, Ano62d, Ano62e, Ano63b, Ano63c, Ano63d, Ano63e, Ano64b, Ano64c, Ano64d, Ano65b, Ano65c, Ano65d, Ano65e, Ano66b, Ano66c, Ano66d, Ano66e, Ano67c, Ano67d, Ano67e]. Books [Ano67f, Ano68d, Ano68e, Ano68f, Ano68g, Ano69c, Ano69d, Ano69e, Ano69f, Ano70c, Ano70d, Ano70e, Ano70f, Ano71b, Ano71c, Ano71d, Ano71e, Ano72a, Ano72b, Ano72c, Ano72d, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano74c, Ano74d, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83b, Ano83c, Ano83d, Ano83e, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b]. Books [Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11a, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b]. Books [Ano14c, Ano15a, Ano15b, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, Ano18a, Ano18b, Ano18c, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano22a, Ano22b, Ano22c, Ano23a, Ano23b, Ano23c, Ano23d, Boa58, Coh54, Dac04, KM94, Lac10, Lun77, McG63, Mer11, Ogi03, Ric12, Tin88, Wie75, Yeo03, Zag99]. Boom [Fog10]. Borderlands [Hel21b]. Boreel [Qua19]. Borkenau [Jon92b]. Bos [Dav10]. Boscovichean [Kar65]. Boswell [Sil93]. Both [Sch22b]. Bottom [Nan21]. Boulainvilliers [Bur57b]. Boundaries [Sta09b, Sta08a]. Bourbon [Hol82]. Bourgeois [Coh68, Cor77, Til45]. Bourne [Ber57]. Boutroux [Nye79]. Bovelles [Lun21]. Boyle [Ans99, App64, Dav94, Gre62, GJ69, Hou41, Jac77, Loe55, McG72, Mor41, Mul94, MF06, Pri95, Ren76, Rog66]. Boys [Col16]. Bradlaugh [Arn57]. Bradley [Ran66a]. Bradwardine [Wal71]. Brahe [Bla90]. Brain [Bas13]. Bramble [Rou67]. Brandes [Sei42]. Brazil [Pim18]. Brazilian [Bur64]. Breakdown [Lev50]. Breaking [Mai77, Val24, Cha17]. Brentano [Mer45, Mer49a, Sch07a]. Brethren [Som98]. Brian [Sea12, Par08]. Brief [Atw71, KK66]. Briefwechsel [Ham72]. Bringing [Ber86a, Ano86g]. Brinton [Mum81]. Britain [All74, Bok17, Cai07, Cha24, Dac04, Goo02, Han17, Her48, Hor76, Koh56, Lin48, Phi96, Tow19, Yol79a]. Britannica [Lac10]. British [Bar62, Alb17, Ano47k, Ben23, Bou94, Bra24, Coo14, Fie96, Flo46, May04, Sch03, Sch52, Sim55, Smi50, Sul83, Tsa13, Vin07, Zeb94].

Brobdingnag [Ben57]. Broken [Hat78, Hor82]. Brooks [Mil41]. Brougham [Pea94]. Brown [Col10, Neu67]. Browne [Gru73, Hun53, Wil48]. Brucioli [Cum10]. Bruegel [Zag03]. Bruni [Han95, Ian00, Ian08, Qui10]. Bruno [Cat10, Mai77, Mas77, Kri47]. Brush [Ell03, Fri12]. Brutus [Sac72]. Buber [Mur68]. Budapest [Glu13]. Buddhism [Ano84e, Dum81, Elm83, Wri51]. Buddhist [Liu12, Sha88]. Bude [Mon97]. Building [Sho24]. Bultmann [Cre95]. Buranelli [HH57]. Burckhardt [Sav44, Wil59, Gil86, Gro99, Hin92, How99]. Burdon [Vin07]. Bureaucracy [Jou23]. Bureaucratic [Ger25]. Buridan [Wal66]. Burke [Arm00, Bou00, Con86, Con99, Dea72, Dwa11, Kir53, Mel75, Rot58, Rya01, Sar08, Wal19, Wes58]. Burlamaqui [Del62, Dou11]. Burnt [Thi79]. Burton [Bar73, Can73, Gar77, Mor16]. Bush [Min51]. Business [Dav60b]. Busting [Dud57]. Butler [New01]. Butterfield [Sew03, Hob71, Wie80c]. Buzzword [Jou23]. Byron [Rus40]. Byzantine [Wil20]. Byzantium [Rob23].

C [Wil08, Win70, Bac07, Dod08, Ned13, Rob75]. C. [Eis57]. ca [Bla03, Bra24, De 15, Sch22a, Sch23b, Zwi20]. Cabalism [Bou54]. Caceres [Mot01]. Cádiz [Wes15]. Caesar [Mar04, Pri17b, Bli57, Sac72]. Cairo [Har85a]. Cajetan [vL97]. Calais [Jen07b]. Calculans [Wes99]. Calculation [Cal22]. Calf [Iva24]. Calhoun [Har85b]. California [Ger25, Rob59]. Calixtus [Cal12]. Callahan [Nor04]. Called [Mye08]. Calvin [Rat61, Bat48, Mue43, Ros60]. Calvinism [Tea62]. Calvinist [Geo57, Gue95]. Cambridge [Ger25, Sav59, Sta74, Arm69, Cou75, Rog79, Sav60, Str23, Wal07]. Camerata [Kat84]. Campaign [Cau94, O’B85]. Campanella [Hea88]. Camporeale [Bul05, Reg05]. Can [Cha17, Fri12, Gue86]. Canaanites [Ila12]. canales [Fel04]. Candidus [Cou04]. Candle [Gre91b]. Cannons [HV02]. Cannot [Eng08, Fri12]. Canon [Gre64, MP04, Sho99, Sim90]. Canonists [Tie54]. Canonization [Han18]. Cantor [Dau77]. Capital [Dun21b, Mae73, Smi17, Fay78]. Capitalism [Ben23, Col76]. Capitals [Dun21b]. Capitulate [Hel81]. Cardano [Cor18, Pri17b, Sir91]. Cardinal [Tho07]. Care [Cla23]. Career [Boa63, Dra77a, Kla93, WB93]. Caricature [Ela15]. Carl [Kel04, Mag16, Nel87, Sme20]. Carlyle [Gol72, Rif53]. Carneades [Dot86]. Carnegie [Isa10]. Carolina [Hsu02]. Caroline [Bri42, Mel99]. Carolingian [Tro73]. Carrie [Ami07]. Carroll [Sch15]. Cartesian [Aqu88, Bou01, Don55, Gew41, Goo18, Hey79, Ken90, McC79, Mie69, Nad88, Nas46, Pet15, Shu68, Ste67b, Wac96, vR00]. Cartesianism [Ham08, Sch99, Stu97, Tho24]. Cartography [Dun21b]. Carus [Bis74, Hay56, Jac68]. Carus-James [Bis74]. Casaubon [Cal12]. Case [Aer19, Arn57, Ber77, Blo96, Cha66, Chr89, Con82, Con99, Cor18, Cra40, Had79, Hay56, HH66, Hil01, Hil72, Jac68, Jac82, Kat84, Kri64, Mac88, Pal72, Por81, Pri48, Sca99, Sch43b, Sir03, Syp63, Tie63, TA06, Urb12, Wes15, Woo19, Bla22]. Case-History [Cra40]. Case-Study [Had79, Hay56]. Cases [Dom84]. Cassiodorus [Sto85]. Cassirer [MT13, Pri57, Ver69, Wis97, Wie51]. Cast [Lyn00]. Castiglionchio [Wat92]. Casualty [Lar10]. Casuistry [Sta67a]. Catastrophic [LS23]. Catchers [Hes12]. Categories [God08]. Cathedral [Mor20]. Catholic [Cam16, Dav10, Elu92, Ent18, Fat05, Gra44, Liu12, Rob20b, Ros96a, Spe02]. Catholicism [Bia24, Hel73, Mil76, Sch07a, Tut22]. Catholics [Gre18, Sar14]. Cato [CW13]. Caton [Ken73]. Catt [Ami07]. Causal [Len58, Man42, Ste60]. Causation [Bar63, Coh42, Sch43c, Teg42]. Cause [Duc06, GS91, Wal63]. Causes [Guy02, Tod76]. Cauterising [Sou92]. Cavendish [Boy06]. Caverns [Lea78]. Caxton [Ant68]. CE [Fog12]. Celebration [Mit71b]. Celestial [Moh04, Odo66, Pat59, Ros85b, Tru12]. Censorial [Bau45]. Censors [Bro24]. Censorship [Cra45, Jen44, McC79, Syp63, Tho55]. Centenary [Fle59, Wie73]. Center [NS16, Sma83]. Central [Hel21b]. Centuries [Afr93, Ger72, Lyo03, Ble58, Gre14, Kib46, Kno65, Mar22, Oak98a, Phi70, Sch23a, dC11]. Century [Bur81, Cal97, Cha24, Dur43, Eme84, Gar84, God08, Han17, Jac76a, Jac82, Jen20, Kle99, McC79, cN93, O’B85, Whi71, ZMM10, All49, And04, Ant24, Ari92, BB89, Bar43, Bas13, Bea82, Beh08, Bil11, BB51, Bri53, Bro05, Bur03, Cam16, Can77, Cat10, Cha58b, Cha60, Cla82, Coc62, Cog07, Coh76, Cro52, Dac04, Dal88, Dan81, Dav66, Dav60b, DeM74, Dod08, Dum81, Eas62, Ell57, Fie76, Fie96, Fin47a, Fix89, Fle63, For74, Frä75, Goo90b, Goo02, Gra65, Gre67, Gre53, Gue89, Gun68, Gur00, Hac83, Han67, Har98, Hei78, Her92, Hil01, Hol21, Hou41, Hou42a, Hou42b, How01, Hyd56, Ila12, Jac76b, Jaf80, Jay70, Jen07a, Jun53, Kel81, Koc21, Kri60, Kri64, Lan74, Lan76, Leg25, Lev56, Lev89, Mar19, Mas00, Mas77, McL76b, McR65, Mer90]. Century [MM89, Mil43, Mor78a, Mor77a, Mor20, Moy03a, Mul06, Neh77, Nor77, Not11, Oak98b, Ogd49, Ogd74, Oti02, Pac79, Pau81, Pax93, Pey49, Phi96, Rad93, Rau88, RH79, Res60, Ric12, Rid56, Ros96a, Rou67, Row56a, Rud74, Sch04, Sch80b, Sch22b, Scl20, Sel13, Sho99, Shy14, Sig62, Sil78, Sim65, Sin89, Sko64, Smi54, Smi78, Smi24a, Smi24b, Söl16, Som98, Sou89, Ste82, Ste58, Str51a, Swa62, Tan98, Tat65, The23, Tho51, Tho55, Tow19, Tri91, Tru12, Van16, Var50, Wal86, Wei96, Wei75, Wes11, Whi47, Wol76b, Yol55, Yol79a, YY84, Zeb94, dG80, Bin67, Sch40]. Centuryi [Ned91]. Certain [Aqu88, For84]. Certainty [Sho85, Sou00]. Cervantes [Cer05, Sti98]. Cesare [Mae73, Maz14]. Ch’ên [Sch51]. Chaadaev [McN66]. Chain [And76, Cha17, Lov46a, Lyn45, Wil80, Wil87]. Chair [Wal07]. Challenge [Coa50b, Dig06, Mag16, Phi89, Sta89]. Challenges [Mor22a]. Chamberlin [Win70]. Chambers [Bra56, Bal50, Rau88, Yeo96, Yeo03]. Change [Coe16, Dil74, Lac65, Moo45, Nau02]. Changing [Bla91a, Bow75a, Fra42, Hav53, Hui43, Jaf80, Mor78a, Osl70, Ran61, Ran66b, Rie68, SB83, Wes76, Whi71, Woo60]. Channeling [Kos24]. Chant [Zon14]. Ch’Ao [Lev50]. Chaos [Cha63]. Chapman [Ami07]. Chapoutot [Ger25]. Chapter [Cra45, Dyk52, Eas67, Haa48, Nad01, Par44, Wri49]. Chapters [Ros16a]. Character [Bar75, Cla82, Dil74, Dym21, Goc99, Hab95, Har02a, Haw01, Koh40, Wol68].

Characteristics [Che51, Dav69]. Characterization [Gre64]. Characters [Sin89]. Charcot [Gel89]. Chardin [Sho24]. Chariot [Mul58]. Charisma [Zha25]. Charity [Win81]. Charkha [Men20]. Charles [Alt08, And55, And76, Bel15, Bui02, CH95, Col74, Dav10, Day08, Dei54, Fal54, Gor08, Hun72b, Lun21, Mad57, Man57b, Nad93, Phi53, Pri17a, Red74, Ros84, Rus75, Skj22, Ter21]. Charleton [Bro18, Lew01, Osl79]. Charm [Ros71b]. Charts [Doa17]. Chaucer [Gre64]. Chauncey [Ano45f, Ano45o, Gri12, Mad54, Wie45]. Chemical [Mil70]. Chemistry [Fle63, HH66, Rob23]. Chernov [Cro71]. Chesse [Ant68]. Chester [Ash78, Ash79, Roy79]. Chew [Gal74]. Cheyne [Ebe75]. Ch’I [Lev50]. Ch’I-Ch’Ao [Lev50]. Chickens [Naj99]. Child [Kiv67, SR81, You01, Sch15]. Childhood [Ber74a, Hab90, Eli67]. Children [Arr94, Col16]. Chile [Wol76b]. Chilean [Spi80]. Chimera [Rob60]. China [EST66, Dav83, GS23, Lac45, Nee51a, Nee51b, Ram01, Web22, Yin22, Zwi20, Sch65]. Chinese [Chi84, Bal50, CR81b, Hig61, Liu04, Loe61a, Mum81, Mun79, Mun80a, Mun80b, Nee51b, cN93, Per02, Pet88, Sha88, TC47, Wri60, Wym49]. Ch’ing [cN93, Oxt01]. Chivalric [Str76a]. Choosing [Nad94]. Chouet [Hey79]. Chrestomathia [Tay82, Itz78]. Christ [Tur66]. Christendom [Le 45]. Christiaan [Wol00]. Christian [Att83, Bar71, Bar59b, Bla61, Cas47, Cer05, Cor75b, Ede18, Fin14, Fis82, Frä75, Fry54, Gre14, Hun00, Koh17, Lac53, Lid66, McK82b, Mom51, Mue43, Mul09, Ras77, Sho18, Tun14, Vov12, Wel73, Wen76, Whi42, Wun03, vdZ98]. Christianity [Cum73, GSJ18, Gre81, Mor22b, Poj83, Pre03, Wal84a]. Christianization [Kla76]. Christians [Ioz22, Shy14]. Christine [Gab55, Mar86, Wil66a]. Christmas [Not11, Sol70]. Christoph [Mah15]. Christopher [Car08, She57]. Chronicler [Ban55]. Chronicles [Gre67]. Chronological [Not11]. Chronology [GS23, Gra03, Not16, Tri56]. Church [Mor86, All85, All89, Fal20, Nau02, Reg05, Wat72, Wil20, Zuc75]. Churches [Lid66]. Cicero [Gou10, Gre75, Hat68a, Ian00, Kli24, Pri64a, Van95, vdZ98]. Ciceronian [Ned88]. Cinquecento [Hau11]. Circe [Hug43]. Circle [Bir96, Don55, Hun53, Jac77, Mai77, Bra02, Hac98, Kar66, Lev13, Nye79]. Circles [Pet88]. Circuit [Oti02]. Circulation [Low74, Mer11]. Circumference [Sma83]. Circumspect [Kel87]. Citation [EMR13]. Cities [Mar72]. Citizen [Ron12]. Citizens [HM79, Mer90]. Citizenship [Hil11, Mil96, Mor18, Pla10]. City [Afr62, Bur93, Cas63, Eme67, Joh42, Mom51, Run00, Spe13, Wei75, Whe55, Lat24, Kah61]. Civic [Har93, Tri56]. Civil [Mac70, Ned13, Per57, Rib52, Bon56, Nad82, Rik54]. Civilian [Glo57]. Civility [Lee95]. Civilization [Gar96, McA10, Nee51b, Pau81, Wri60, Bar62]. Civilizational [Tsa13]. Civilizations [Dan52, Deu45, Kro53, Gey48]. Clairvaux [Whi60]. Clandestine [Sut03]. Clarendon [Hoe15, Mal15]. Clarification [Cre49, Pri49]. Clarissa [Ste88, Moy75]. Clark [Ano75f]. Clarke [Gay63, Guy20, Mel99, Per69, Sil86, Ste81, Wig09]. Class [Ano84f, Clo84, Kos04, Mer48, Per18, Ran23, Wal86]. Classical [Afr93, CM05, Clo69, Clo71, Cum10, Dan52, Dil74, Dod68, Fit97, Ger25, Joh61, Jon40, Kit66, Mil74, Mor98, Ols91, Par68b, Whi86, Wun03, Cuf66, Vis84]. Classicism [Del89, SF72, Tri56]. Classics [Hol53]. Classification [Ree52, Spi55a, Tat63a]. Claude [Bat00a, Boo00, Vir86]. Clausewitz [Par68a, Par88a, Par04]. Clausum [Vie03]. Clayton [Ber71]. Clear [Gau92]. Cleghorn [Nob65]. Cleo [Tri91]. Clerical [De 53]. Clerk [Mos60]. Clever [Coo99]. Cliché [Rod07]. Climacteric [Bal44]. Climate [Coe16, Gat67]. Climes [Joh60]. Clinker [Rou67]. Clio [Mul58, Tri91]. Cloche [Fal20]. Clocks [Sco97]. Clouds [Ans78]. Club [Pea15, Wie46c]. Cobbett [Aus52, Keg58]. Coburn [Pri58]. Cocceius [Yof04]. Code [Shu63, Wat92]. Codex [Gim08]. Coercion [Lub24]. Coffee [Glu13]. Cognitive [Str16]. Cohen [Ran49, Ano47f, Bro53, Cah17, Coh48, Lar59, Nag57, Ros57b]. Coherence [For09, McG63]. Coherent [Ril73]. Coincidence [Eme84]. Coit [Cop51]. Cold [Moy23]. Coleridge [All85, All89, Bos70, Coa77, Hav59, Hed98, Hes12, Hip04, Jac83, Koo99, Lin48, Met60, Mil64, Mil99b, Mor86, Pip59, Roy07, Sha70, Tho44, Val95, Pri58]. Colet [Duh53, Loc07]. Colish [Ned24]. Collection [Rom64]. Collective [Kat84, Str56, Öza14]. Collectivism [Coa50a, Guy20]. College [Feu58, Joh40]. Collegiant [Qua19]. Collegiants [Fix89]. Collingwood [Bou97, Ska05]. Collini [Cai07]. Colloquium [Nad13b]. Colombia [Mor22a]. Colón [Kim21]. Colonial [Bur64, Kla76, Leg25, Ort18, Ros24, Sum17, Tsa13]. Colonialism [Arn94]. Colonization [Kam70, Urq14]. Color [Gue86, Lau90, Mor18]. Colors [Gue86]. Combe [Gra65]. Comedies [Fle66]. Comes [Coa61, McG13]. Cometography [Joh04]. Comets [Ari92]. Comfort [Kra18]. Coming [Con86]. Comma [Lev97]. Commandino [Cla15]. Commandment [Ila12, Nic90]. Commemorating [Wie73]. Comment [Gil87, Kri95, Pet12, Sch49, Toy55, BB70, WW50]. Commentaires [Gar13]. Commentary [Atw71, Car08, Gre74, Kal95, Poc16]. Commentators [Wal81, Yof04]. Comments [Dyk71, Rat61, Shu63, Wie70]. Commerce [Jur01, Pla10, Tho07, Ven76, Boe83]. Commission [And21]. Commitment [Ver73]. Commits [Zhu02]. Common [Ede98, Jac12, Ram14, Ros92, Ros71a, Woo24]. Commonplace [Bla92, Clo69, Dac04, Hes12, Mos98]. Commonplace-Book [Hes12, Mos98]. Commonplaces [Kap23, Yeo96]. Commons [Poc78]. Commonwealth [Fer55]. Commonwealthwoman [BB89]. Communal [Ned13, Ste51]. Commune [Bas13]. Communication [Hof54, Oti02, Rus93]. Communism [Ham53, Ham72, Lov42, Rob20b]. Communist [Fef05, Kla71, Kos24, Moo45]. Communists [Sar14]. Community [Bok17, May04, Rob20c, Scl20, Ski76, Woo88]. Commynes [Arc67]. Comparative [Chi84, JN12, Lau44, Rub91, Sho16, Smi64]. Comparison [Ham59, Sho52, Ska05]. Compassion [Fie76]. Compatibility [Pet84, Ruk18]. Compendium [Aug65]. Competing [Col09, Urq14]. Compilation [Wei19b]. Complete [Fer68, Mos48, Zak24]. Completion [Gre23b]. Complex [Van17]. Complexity [Dil74, Jum08]. Composing [Har89]. Composition [Gil53]. Compromise [Smi76a]. Compulsory [McQ14]. Comte [Cas78, Lew72, Pic89, Sim65, Ver84]. Conceit [Gan56, Wat55]. Conceitos [TT24]. Concept [Åsb21, Ano84f, Arr94, Atk14, Bar50, Bej97, Bow76, Cas92, Clo84, Coh76, Cuf66, Dav66, De 79, Ell62, Gou61, Gre15, Gre80, Hac83, HM79, Har86, Hav53, Jaf80, Kal87, Kra18, Lea80, Low74, McN73, Mil64, Mul63, Mul68b, Mun80b, Nag68, Nis44, Pod10, Ran66b, Rou44, Sac76, Sch14a, Smi63, Sol83, Tae42, Tow87, Tow19, Tro73, Wil60, Zil45, dBJN^{+}20]. Conception [Alj19, Ano64j, Bre01a, Bur49, Coh80, Die41, Eck44, Fal88, Gou74, Joh85, Kel64, Lat94, Le 45, Maz57, McG72, McK47, Nel72, Phi52, Rap81, Shu62, Smi80a, Tim99]. Conceptions [Fat05, Fra42, Hei78, Str21]. Conceptos [TT24]. Concepts [AU19, Bay49, Deg12, Fac21, Hol49, Jer13, Mor92, Pim18, Was84, Wen82]. Conceptual [Cla86, Ede22, Gou10, Mor22a, Ort18]. Conceptualization [Rap21]. Concerning [Dor40, Goc99, Lan77, Man55]. Conches [Cad95]. Conciliar [Fas97, Tie54, Rue64]. Conciliarism [Sig62]. Conciliarist [May88]. Conciliarists [vL97]. Concordia [For42]. Concreteness [Bra02]. Condillac [Sch10]. Condition [Cog07]. Conditional [Noo72]. Condorcet [Ber18, Koy48]. Cone [Ebe75]. Conférence [Ros84]. Confessional [Val22, Ver22, Twi20]. Confessions [Smi54]. Configurations [BJ19]. Conflict [Bre52a, Mil87, Deu45, Don82, Fef90, Hes18, Hil01, Ran23, Sch46a]. Conflicting [Sil76]. Conflicts [Lam56]. Confluence [Mun79]. Confraternity [Scl20]. Confrontation [God08]. Confronting [Ger25]. Confucian [Chi74, Dav83, Kim14, Liu12, Pri48, Wan05, Wil51]. Confucianism [Hos24, Lev50, Lun83]. Confucius [Ham59, Lun83]. Confused [Bar92]. Conjectures [Wol07]. Conjugal [Ach20]. Connection [All74, Ces04, Mit93, Nad94]. Connections [McR63]. Connective [Dun21b]. Connoisseur [Fri12]. Conquest [Lam56, Loc04, Tho77]. Conscience [Atk14, Gre91a, Hic18, Kli06, Nur61, Sku64, Öza14, Hip04]. Conscious [Han17]. Consciousness [Aqu88, Cor10, Gri12, Hip04, KM94, Kre00, Neu25, Tho75]. Consensus [Cal12]. Consent [HM79, Lub24, Wat72]. Consequences [Mac67]. Conservation [Ano42f, Cro53b, De 15]. Conservatism [Fis58]. Conservative [Bla52, Bra71, Bro24, Hud56]. Considerations [Ber74b, Gil20, Lan77, Smi80b]. Conspiracy [D’E07, Rat86]. Const [Mie73]. Constance [Oak66]. Constant [Ben82b, And02, Fin72]. Constantine [Cam96, Con96, Del96, Fub96]. Constitution [Bly86, Gol52, Chi40, Wes15, Woo24]. Constitutional [Bro24]. Constitutions [Hsu02]. Constructing [Rob20c]. Construction [AU19, Wan05, Wun03]. Constructions [And08]. Contacts [Col74]. Contemplative [Lun21]. Contemporaries [Der10, McK81b, RMG81]. Contemporary [Cop05, Sea11, Sli53, Tie11, Wit77]. Content [Gut41, Mer47, Sei87]. Contents [Ano03e, Ano04e, Ano05c, Ano06e, Ano07e, Ano08e, Ano09e, Ano10c, Ano11b, Ano12e, Ano13d, Ano14d, Ano15c, Ano16e, Ano17e, Ano18d, Ano19d, Ano20e, Ano21d, Ano22d, Ano23e, Ano24a, Kem00]. Contestation [Rob15]. Contested [OH88]. Contesting [Kre00]. Context [Ano14e, Bar16, Byr06, Cel14, For82, Got80, Hoe15, Lea77, Mer89, Par68b, Per18, Pre79, Ros16a, Sch14a, Sho99]. Contexts [Col09]. Contextual [Rot90]. Contingency [Ter21]. Contingent [Mag16]. Continua [Thi92]. Continuing [Bla91a]. Continuity [Aga73, Bar63, Gra81a, Par88a]. Continuous [Eva55]. Contra [AA85, GM80, Vie15, Zwi20]. Contract [Lev67, McK82a, Mul79, Ril73, Shu62, Ber77]. Contractility [Sar08]. Contractual [Lev78]. Contradiction [Bos73]. ContrariesHappy [Pac79]. Contrast [Ogd49]. Contrat [Ech72]. contre [Tan98]. Contribution [Chr45, Dew18, Fer58, Men61, Reg05]. Contributions [Bre01a, Bul05]. Control [Cha66, McL76b, Sta87]. Controversies [Bla91a, Tho77]. Controversy [Bis74, Bla19, Blo79, Bur74, Cal69, Cog07, Dam79, Gor94, Gre69a, Gre72a, Lat24, Mai89, Nis21, Rap69, Ros70, Rou67, Str55, Tho24, Wag54]. Convention [Che56]. Conventional [Pea54]. Conventionalism [Nye79]. Conventions [Ste65]. Converging [Ban14]. Conversation [Cot08a, Wal93]. Conversible [Gur00]. Conversion [Maz14]. Conway [Bro18]. Cooking [Tri91]. Cooley [Pal72]. Cooperation [Don82]. Copernican [Bla90, Han61, Hen01, McC42, Rid62]. Copernicana [Ros74, Sta73]. Copernicus [Rat61, Bla90, Gra62, Gue68, Han59, Mos73, Neh77, RH79, Ros43a, Ros60, Ros71a, Ros75, Ros83, Tay55, Zil40]. Coping [Bla03]. Cornewall [Boc51]. Cornhill [Blo84]. Corporation [Bec69]. Corporations [Haw94]. Corporative [Ada42]. Corporeal [MM89]. Corpus [Bos89]. Correction [Ano46g, Car78, Lam91, Smi61]. Correctness [TW21]. Correspondence [Ano47f, Coh48, Cop51, Isr19, Maz53, Mul06, Ros40, Spe96, Tur71]. Corrigenda [Ano51a]. Corrupt [Hor80]. Corruption [Jac76a, Mit87]. Cortes [Spe02]. Cosmic [Bar73, Tur75]. Cosmogonies [Fei68]. Cosmography [Kim21]. Cosmological [Gui13, Nad98]. Cosmologies [Ver82]. Cosmology [Gun71]. Cosmopolitan [Stu07]. Cosmopolitanism [Had43, Kle99]. Cosmos [Boa67, Egg99, Gue68, HV02, Kub67, Sma83]. Cotkin [Jum08, Liv08, O’B08, Per08]. Cotton [Jes86]. Coulanges [Wil73]. Council [Bro24]. Count [For42]. Counter [Kos24, Nor07, Ree83, Kri51a]. Counter-Enlightenment [Nor07]. Counter-Reformation [Kos24]. Counter-Renaissance [Kri51a]. Counter-Revolution [Ree83]. Counterblast [Mac98]. Countering [Les07]. Country [Chr23, Fon99]. Course [Bur87]. Court [May88]. Courtly [Moo79]. Courts [Bro04b]. Cousin [Sim65]. Couturat [Rem79]. Covenant [Joh71a]. Cowper [Kro60]. Cracks [Hac83]. Craft [Pau72]. Craftsman [Sol63]. Crane [Mum81]. Crashing [Mor20]. Craving [Tho51]. Crawshay [Ano47k]. Creation [Car72, Ent18, Mar19, McD05, Nar05, Poo05, Sta89]. Creative [Mye93]. Creativity [Rol71, Kri83]. Creator [Nah47]. Credibility [Tut22].

Credulity [Dor40, Tut22]. Creel [Pri48, Pri49]. Cressida [Elt97]. Cretan [Kah61]. Creuzerstreit [Ger19]. Crime [Fre64, MP04, Zhu02]. Crimean [Ada79]. Criminal [Shu63]. Crise [Len77]. Crisis [And41, Ban14, Bre01b, Fol17, How99, KR06, Kri65, Lin48, Rig98, Sti98, Tri56, Zee46, Bro61]. Critic [Bec72, Bra57, Got75, Mye75, Pop52, Shu85, Tho44]. Critical [Cla23, Gar55, Mul09, Ruk18, Smi76b, Soc06, Wal96, Wea80, Wil23]. Criticism [Cap54, Chr59, Coh40, Daw93, Dil74, Doo90, Duf95, Jon40, Kov76, Lar41, Lev89, Mor25, Ran67, Rey53, Rif53, Sch10, Sir18, SC40, Sol03a, Sol70, Tri56, Twi20, Wei69, Wel69, Ano69l, Cul88, Mat43]. Criticisms [DB77]. Critics [Con99, Fal20, Fei77, Lee79, Lyn77, Ski12, Sli53]. Critique [Alj19, Aro87, Bid76, Bla90, Cri66, Dun64, Gre75, Hol50, Imm09, Kar69, MB81, MS61, PB73, Pri48, Rya01, Sme20, Sny11, Wal81, Wat71, Zuc75, Cro05]. Critiques [Cho15, O’C84]. Croce [Sch41, Cas83, Jac75, Poi70]. Crocker [Bro61]. Cromwell [Maz60]. Cronin [Ger25]. Cross [Ada79, Chi84, PD63, Sho16, Tan99]. Cross-Cultural [Ada79, Chi84, Sho16, Tan99]. Crosses [Tut22]. Crossings [dG84]. Crown [Bat19]. Crucial [Tou57]. Crudelis [Dom88]. Cruelty [Bar98, Dem72]. Crusade [Rem79]. Cruz [Gim08]. Cudende [RH79]. Cudworth [Sai62, Sai88]. Cul [CH95]. Cul-de-Sac [CH95]. Cult [Cor75a, Eli67, McN45]. Cultivating [Koc99]. Cultural [Ada79, And41, Bou03, Bri42, Cha81, Chi84, Dav51, Ebe69, Gil86, Kel98, Mer11, Mit71b, NS16, Per02, Ros96a, Sch56b, Sho16, Sin00, Sta91, Ste78b, Vic02, Wer13, Tan99, Ker24]. Culture [Ano84f, Bar83, Bou03, Byn84, Clo84, CH15, Cot94, D’E07, Doo99, Gat67, Gra88, HE63, Kal87, LS17, Low54, Maz54, Mil96, ME21, Pau81, San20, Sch22a, Sha70, Sim65, Spe98, The23, Tho07, Vuc60, Wri49, Mit94]. Cultures [Joh06, Mac88, Ree83, Val22, Wei19b, Web79]. ’Cunning [Ben84, Cat73b]. Curia [Wat92]. Curiosities [Har85a, Mar10]. Curious [Ros71b, Wes15]. Currency [Tay55]. Current [Mar11, Pag88, Sap79]. Currents [MB81, Rey53]. Curti [Sch44]. Curtius [Vic02]. Curvature [Egg99]. Cusa [Ned05]. Cusanus [Wat72]. Custodiet [Ell67]. CXLI [Goo90a]. Cybernetic [Sim16]. Cycle [Ash78, Roy79, Ash79]. Cyclical [Kub67]. Cyclopaedia [Yeo96, Yeo03]. Cynicism [Say45]. Cyrenaica [Ros24].

D [Ano03h, Ano09g, Ano15e]. D.A.P. [Von48]. d’ėtat [Bak91]. D.N.S.A.P [Von48]. D’Alembert [Chr89, Shk81]. Dalibray [Rid56]. Damaris [Bro06]. Dana [Gue95, San65]. Dance [Wil22a]. Dangers [Vic90]. Daniel [And21, Cop51, Fer71, OH88, She02, Woo88]. Daniels [Dew59]. Danish [Lau98]. Dante [Egg99, Maz57, Sin49, Ste49, Her21]. Danton [Mit87]. Dares [Cla11]. Dark [Gue86]. Darwin [CF79, Alt08, Ano45f, Ber78a, Bow76, Cau94, Col74, Day08, Fay78, Gar55, Har71, Hun72b, Man58, Pri17a, Pri64b, Ran61, Ros71b, Rus75, Sch45b, Vor69, Wie45, Wil65]. Darwin-Marx [CF79]. Darwinian [Ell57, Eva65]. Darwinism [Whi45a, Bur65, Cla00, Ell58, Hje11, Mik70, Pre03, Rog72, San65, Wei93, Wei02].

Dating [Bro92]. d’Aubignac [The23]. Daughter [Koh17]. David [Hey79, Hun72b, Hur56, Lau98, Les72, Mas00, Mos60, Mos78, Obe76, Pop53a, Rap69, Sho52, Thi92, Tsu61, Vid95, Wer72]. Davidson [Ano88e, Bla57, DeA87, Lat57, Les57]. Davies [Orr14]. Davila [Gre16]. Dawson [Car08]. Daybreak [LS91]. Days [Teg47]. De-Animation [Dal80]. Dead [Von84]. Deadlock [Val24]. Deafness [Bau23]. Deane [Fra91]. Death [CW13, Mil99a, Wei02]. Debate [Bra24, Bro24, Cal22, Dah15, Dic80, Goo02, Gor19b, Gre54, Hef81, Jen07a, Jur01, Kuk00, Lau98, McK82a, Mol97, Pri17b, Rau88, Sel13, Ste82, Twi20, Val24, Vie03, Pre03, Web22]. Debates [Lan80, Mer08, Poc78]. Debating [Bla22]. Debt [Coa77, Dou11, O’C64, Sai88, Tha83, Wol60]. Debunkery [Eng07]. Decadência [Mar20]. Decade [O’N12]. Decadence [For93, dH45]. Decades [But15, McG03]. Decay [Hei73]. Deceit [Fle66]. Decision [LS23]. Decision-Making [LS23]. Declarations [NL20]. Decline [Bro81, Har85b, Hei73, Moh04, Pat59]. Decolonization [Ant24, Rob20a]. Decolonizing [Ger25]. Decomposing [EG84]. Decorum [Sch52]. Decretists [Wau20]. Dedicate [Fay78]. Dee [Tra64]. Deeper [Bra24]. Defender [Wol49]. Defenders [McC79]. Defending [Wal19]. Defense [Ano45f, Arn94, Hor80, Jon73, Len58, Mar59a, Osb02, Sul83, Wie45, Wil91]. Defenses [Alb55]. Deference [Sch14b]. Defining [Duh53, Par07, Rob15]. Definition [Lan80, Lev79, McL86, Sta14, Sta91]. Defoe [Bis52, She44]. Deformations [Alj19]. Deformities [Han17]. Degeneration [Mai89]. Degradation [Col76]. Deism [Bid76, Kro60, Red74]. Deists [Col59]. del [Boa56]. Deleuze [Bia24]. Delimitation [Kro53]. Dell [Dur43]. Della [Bre52a, Bre52b, Cas42a, Cas42b, For42]. Delusion [Mac67]. Demagogue [Lan12b]. Demarcation [Lee95]. Democracy [AU19, Bak76, Bat77, BJ19, Dea74, Fef05, FR07, Lac10, Las47, Lid66, Mal17, Mos85, Moy07, Nol84, O’N07, Par24, Rei17, Sim04, Wol58, Ste51, KM94]. Democratic [Las47, Sch43a, Wol16a]. Democraties [Dre64]. Democritus [Deg12, Joh67a, Mor16, Tig70]. Demography [Ege66, You75]. Demon [Cat73b, Nad97, Sta08a]. Demonology [Tri69]. Demons [Mac06]. Demonstratio [She02]. Demostration [Con61]. Demotion [Wal19]. Demoule [Ger25]. Denial [Gri21]. Denis [Keo74]. Denmark [Hje11, Lau00]. Dennis [Del07]. Denudation [Dav66]. Denying [Rei96]. Deo [Bat00b]. Departures [La 93]. Depiction [Gai00]. Depopulation [You75]. Depravity [Ban16]. Depression [Ano70g, Bor71, Kes68]. Derangement [Was72]. Derodon [Hey79]. Derrida [Her05]. Desacralization [Hun00]. Descartes [AG95, Aug65, Bar92, Cat73a, CJ60, Gab98, GCM10, Gau92, Gor94, Kee92, Ken71, Lo17, Lyn82, McR57, Nad88, Nad97, Nad22, Osl79, Osl85, Pet15, Rei96, Rub90, Sai62, Sav59, Sav60, Sch80a, Sha63, Sup54, Tho24, Tim99, Wil97]. Descendit [Smi67, Tur66]. Descensus [For68]. Descent [Tur66]. Described [Sch52]. Descriptio [Sch23b]. Desert [Cla57]. Desiderata [Kel12]. Design [Alt08]. Desire [Pac89, Pet01]. d’Espagnet [Lin91].

Despair [Spi66]. Despotism [Ven63]. Destruction [Kel80]. Destutt [Ken79]. Detected [Eig09]. Determination [Pri13a, Ran70]. Determinism [Ban16, Hac83, Wel58]. Deus [Tie63]. Deussen [Rol78]. deux [Ans09]. Devaluing [Wei02]. Development [AK23, Ada91, Bea91, Bis96, Bos73, Bow85, Gol72, Gre15, Hal74, Hin92, Hof85, Hos24, Jer13, Kel12, Lea77, McK47, McK66, McK87, Mul68b, Par44, Pop54, Raj52, Ran40, Ran66a, Ran70, Sim81, Smi80a, Spi55a, Tow87, Tow19, Tsa13, Ver22, Zag93, Zan79]. Developments [Man91]. Devil [Rea72]. Devils [Fix89]. Devotional [Rub90]. Dewey [Ano70g, Ano88e, Bor71, Bot17, BJ19, Bux84, CH15, Cyw69, DeA87, Dew59, Dyk52, Dyk59, Dyk61, Dyk62, Feu58, Feu59, Nol84, Phi71, Rat92, Sch50, Wil56, Wil23]. Deweyan [CH15, Fef05]. D’Holbach [Lad62]. D’Huisseau [Som67]. Diagnosis [Pim18, Sch07a]. Diagram [Wan05]. Diagrams [Dun21a]. Dialectic [Ber78b, Got80, Ina20, Kuh49]. Dialectical [MS61, Mue45]. Dialectics [McK83, Sea80]. Dialogi [Cum10]. Dialogue [Bla82, Mur68, Sar14, Sho16, Sta16]. Dialogues [Coh62, Den99, Pop88, Qui10, Sil76]. Dichotomy [Vic90]. Dick [Sto51, Tri69]. Dickens [Gol72, Man57b]. Dictionaries [Ger72, Mul06]. Dictionary [Ano03f, Gab98, Har85a, Kor74, McL86, Sch05b, Sch03]. Dictionnaire [Sho85]. Did [Fay78, Gri58, Hel81, Len02]. Diderot [Ans09, Die41, Isr16b, Sch96, Var73]. Diego [Chr23, Spi65]. Dietzgen [Eas58]. Difference [DeV08, Pax93, Sta14]. Differentially [Cha23a]. Differing [Joh60, Öza14, Sav60]. Diffusion [Flo95, Haa48, McC20]. Dig [Ger25]. Digger [Jur75]. Dignity [Con79, Dal77, For42, Her21, Saa10, Ter18]. Dilemma [Dra90, Sil01]. Dilemmas [Rei96]. Dilthey [Cro05, Fro95, Hol50, Mas52, Ran64, Fri40]. Dionysian [Del89]. Dionysius [Loc07, Wil91]. Diplomacy [Bri23, Hel21b, HTW21]. Diplomatic [Hia09, Hol21]. Direction [Nau02]. Disability [Bau23]. Disappearance [GCM10, Gud96]. Discernment [Whi09]. Disciple [Spi80]. Discipline [Dod08]. Disciplines [Wal19]. Disclaimer [Lov49]. Disclosure [Hun86]. Discontents [Gar96]. Discontinuities [Mon19]. Discontinuity [Aga73, Par88a]. Discord [Pac79]. Discorsi [Gil53]. Discourse [Bea83, Ina20, Kre00, Loe47, Mor93, Mos60, Pan21, Per18, Sac18, Sea11, Tie11]. Discourses [Gre16, Mar20, Tuc03]. Discoveries [CF79, Ros57a]. Discovery [Boo43, Com71, Dav13, Had79, Har46, Koh17, Mue45, Tim99, Efr97]. discussed [Her05]. Discussion [Arm56, Chr45, Coa77, Cro52, RGE45]. Discussions [Dup82, Pac79, dC11]. Disenchanted [Phi83]. Disenchantment [Kei04, vR00, Wen24]. Disenfranchisement [For03]. Disharmony [Har85b]. Disinterestedness [Van01]. Dismissal [JW98]. Disobedience [Ste60]. Dispersion [Pop88]. Disputa [Boa56]. Disputation [Bla22]. Dispute [Wal93]. Disputed [Mac71a]. Disraeli [Rat86]. Dissent [Cam96, Con96, Hef78]. Dissertation [Rub91]. dissimulare [Bak91]. Dissolution [Ros85b, Hor97]. Distance [Ogd74, Sup54, Dyn81]. Distinct [Gau92, Hob80]. Distinction [Bly86, Dyn81, Gud96, Lan12b, Som83, Was84, BS21]. Distortion [JW98].

Distortions [Kri64]. Distributional [dBJN^{+}20]. Disturbance [Han61]. Diversity [Sea12]. Divide [Ned05, Ver22]. Divination [Str16, Tru12]. Divine [Ash78, Ben84, Fal88, Ila12, Ins11, Osl83, Ram14, Roy79, Tea62]. Divinization [Kro60]. Division [Sta14]. Divorce [Haw01]. Divus [Loc07]. Dixon [Wis49]. Do [Cha23a, Sch99]. Doctoral [Fen86]. Doctrine [Arm65, Atl52, Ban70, Bat48, Bla61, Bon56, Gau92, Löw45, Nas46, Pre54, Sam43, Tae53, Wal59, Yar42]. Documentary [Bac07]. Does [Ban14]. Dogma [Wei19a]. Dogs [Day08]. Doing [Fis66, Lan12a, Puc05]. Dollinger [Ton59]. Domestic [Cyw69]. Dominion [Par08, Vic08, Vie03, Whi89]. Domodossola [Bla57]. Don [Ham98, Nad97]. Donald [Car23]. Donatello [Bar75]. Donation [Cam96, Con96, Del96, Fub96]. Doni [Gre65]. Donne [Llo69, Low61, Nic40, Ruo66]. Donoso [Spe02]. Doom [Kue42]. Doors [Low24]. Dostoevskii [LS91]. Dostoevsky [Dah15, Koh45, Sca99]. Dostoyevsky [Bea42]. Double [Ber57, Con14, Men18, Pin68, Tho59]. Doubt [GCM10, Sou00, Wal59]. Doubts [Hal02]. Douglas [Min51]. Down [Ano86g, Ber86a, Tre21]. Downfall [Kra18, Sto46]. Dr [HH57]. Dr. [Ano51a, D’E69, Mid60, Neu67, Ros61, Ste78c]. Draft [Wal04]. Drake [Ros57a, Fra77, Ham96]. Drama [Bla22]. Dramatic [Nau02]. Dramaturgy [Cam20]. Drawing [Cla15, Hel21b]. Dream [Loe70, Sch15, Ste65, Zak24]. Dream-child [Sch15]. Dreamers [Cur66]. Dreams [Cur66, Gab98, Kee92, Ruk18]. Drieu [Sou70]. Drop [Nic68]. Droysen [Gil83]. Drucker [Imm09]. Dryden [Ben82a]. Dualism [Gor94]. Dublin [Ras77]. Duc [Her05]. Dugald [Nor01, Ras85]. Duhem [Pau72]. Dühring [Ada74]. Dumézil [Eld24]. Duns [Rum85]. Dunyi [Wan05]. Duperron [Stu07]. Duran [Bla93]. Durant [Fas97]. During [Ari92, Bak91, Cla23, Han67, Rik54, Sel58, Wei45, Ent18, Ros12, Sar14, Tho94]. Durkheim [Cla92, Haw94, Öza14, Sei87]. Dusk [Afr93]. Dutch [Rik57, Stu89, Wes11, vI11]. Duties [McQ14]. Dwight [Har93]. Dynamically [Kra18]. Dynamics [Moo51a, Moo51b, NS16, Val22].

E. [Cas10, Daw93, Haa48, Har85a, Van93, Mos66]. Earl [Bin67]. Earliest [Sen58]. Early [AK23, Ach20, And64, Ano85e, Bed21, But15, But77, Bux84, Cal22, Cam22b, Cap54, Coh62, Con82, Doo90, Doo99, Dyk52, Eng63, Fel77, For23, Ful94, Gau92, Gol82a, Gol82b, Gra03, Gra78, Har99, HTW21, Hin92, Igg59, Joh67b, Kar66, Kla93, Lee95, Lew01, Mai97, McC42, McG03, Min98, Mos48, Mul68b, Nev09, Nor02, Obe87, Pal09, Phi70, Pip59, Poj83, Ric79, Ros85a, Ros03, Ros18, She06a, Sta09a, Sut03, Tri56, Twi20, Wei19b, Wil88, Woo24, Zag93, But06, Efr97, Arr94, Bea83, Bla22, Bri53, Cam22a, Cam96, Can64, Cha24, Cla11, Col59, Con96, Cor75a, Cre95, Fac21, Fer55, Gom43, Gou06a, Har02b, Hyd56, Jen20, Joh04, Kap74, Kas06, Kil11, Kim21, Lat24, LS17, Mal17, Mil96, ME21, Mul97, Neu67, Par01, Poo05]. Early [Pop52, Pop06, Sch84, Sha02, She03, She06b, Sho99, Sho10, Smi19, Söl16, Sol03b, Sol70, Tri50, Wes01, Wit43, Zah12, San59]. Early-Modern [Lee95]. Earth [Dun69, Sho24]. Earthquake [Van82]. East [Cas48, HE63, Bri23, Fei68, Goo90b, Jac68, Sil01, Vov12, Gas55]. Eastern [EST66, Sin00, AU19, Rob20c]. Eaters [Cau94]. Eats [Che63]. Ecclesiastica [Spr94]. Ecclesiastical [Cra44, Ha20, Loc07, Lyn77, Pre71]. Ecclesiology [Kov76]. Echard [Zah12]. Echoes [Zet82]. Eckstein [McC15]. Eclecticism [Ald44, For09, Kel01, Lew01, Mul97, Sch98, Whi09, Met63]. Eclipses [Gra03]. Eclipsing [Mil99b]. Ecological [Rog94]. Economic [Arn94, Bri41, Chr23, Col76, Gol52, Imm09, Jac12, Lev54, Lin48, Low74, Mac78, Mul23, Pau79, Ski76, Sos73, Spi55a, Web22, Guy20]. Economics [Chr73, Het83, Kit66, MR20, Ski76]. Economy [Cha24, Hun74, Lyo70, Mor98, Pac02, Ras77, Ras85, She77]. Ecstasy [Fin47a, Fin47b, Ste67b]. Edel [Mie73]. Edelstein [Dod68]. Edenic [Lew65]. Edens [Ker24]. Edger [Smi80a]. Edicts [Yin22]. Edinburgh [Lom02]. Edith [Han18]. Edition [Gil83, Hoe15, Mal15]. Editor [Kel05a]. Editorial [Ano63f, Ano64e]. Edmund [Arm00, Con86, Con99, Dea72, Dwa11, Kee47]. Eduard [Bre01b]. Education [Bla91a, BJ19, Cha60, Con60, Cou04, Gug80, Itz78, Kil61, Lun87, Par68a, Par88b, Pos86, Rom61, Smi62b, Som83, Tay82, Tob74]. Educational [Gab55]. Edward [Fer68, Ran41, Wie70, Car19, Dra78, Wie86]. Edwards [And64, Ban16, Dav63, Gue95, Rea72, Wie68, Yar86]. Edwin [Feu79]. Effective [Duh49]. Effects [Che51, Kap56, Men20]. Egalitarianism [Stu07]. ’Egemonia [Wil60]. Egoism [Ger67, Sca99]. Egyptian [Tas69]. Egyptomania [Reh14]. Eighteenth [Bin67, Cal97, For74, Gar84, Ger72, Han17, Jac76a, Kle99, Lan76, cN93, O’B85, Sch23a, And04, BB89, Bas13, Bil11, BB51, Can77, Coc62, Coh76, Cro52, Dac04, Dod08, Fie76, Fie96, Frä75, Gre53, Gur00, Han67, Hei78, Hil01, Hol21, Ila12, Jaf80, Jen07a, Kel81, Koc21, Lan74, Lev89, Mer90, Mor78a, Mor77a, Neh77, Ogd74, Pac79, Pax93, Pey49, Phi96, Rad93, Row56a, Rud74, Scl20, Sel13, Sil78, Smi78, Som98, Ste82, Str51a, Tow19, Var50, Wal86, Wei75, Yol79a, YY84, Zeb94, dC11]. Eighteenth- [Jen07a]. Eighteenth-Century [Cal97, Gar84, Han17, Kle99, cN93, O’B85, BB89, Bas13, BB51, Coc62, Coh76, Dac04, Dod08, Fie76, Fie96, Gre53, Hei78, Hil01, Hol21, Ila12, Jaf80, Kel81, Koc21, Lev89, Mor78a, Mor77a, Pac79, Pax93, Phi96, Row56a, Scl20, Sil78, Ste82, Tow19, Var50, Wal86, Wei75, Yol79a, YY84]. Eilhard [Val22]. Einbildung [Koc21]. Eklektik. [Ano98e]. Elder [Eig09]. Eleatics [Lac65]. Élémens [Chr89]. Element [Wal70]. Elements [Lyn82, Nau02, Rus85, Wal84b]. Eleusinian [Pri64b]. Eleventh [Nic90]. Elia [Jes54]. Elias [Jan08]. Elie [Fro07]. Eliot [Mad57, Bre72, Hab90, Hab93, Rya09]. Elise [CW13]. Elizabeth [Coo81, Ebe69]. Elizabethan [Dor40, Gar77, Jor52, Koc50, Tea62, Tra64]. Eloquence [Gra63b, Sei65, Sta77, Tri91]. Emancipation [Avi64]. Embodiment [Hab95]. Embodying [Mat06]. Emergence [Cro66, Duf80, Ran53, Sil20, Tho24]. Emerson [Ber67, Bro74, Cor10, Goo90b, Gre15, Kli55, Neu77, Rie67, Rob80, Ste53]. Émile [Öza14]. Emotion [New01, Ros12]. Emotions [Dwa11, Fei70, Hey82].

Emotivism [Sat82]. Empathy [Mor96]. Empire [Bou00, Eig09, Mar22, Met14, Ned05, Ros96a, Ven76, Bat19, Kla76, Men16, Sul83]. Empires [Ade18]. Empirical [Kam52, Sol03a]. Empiricism [Eas58, Hen01, Ran65, Rat92, Ren76, Van16]. Empiricus [Pop93, Flo95, Pop53b]. Enchantment [Mor96, Zha25]. Enclosure [Rot90]. Encyclopedia [Doo09, Lær14, She03, Wie68, Mar23]. Encyclopedias [Mar10]. Encyclopedie [Cas63, EMR13]. Encyclopedists [Loe61b]. End [Ste62, Syp65, dC11, Ano00f, Hol00]. Endgame [Bro04b]. Enduring [Jur01]. Engagement [Tay17]. Engdahl [Åsb21]. Engels [Ham72, Ada74, Kri53, Mos85]. Engine [Wes77]. Engineers [Kap56]. England [Coa50b, Mil41, Sch40, Alb78, Bat19, Cog07, Dav66, DeM74, Dod08, Fai41, Fer55, Fin47a, Gal94, Gar55, Gar84, Gra61, Gue89, Jon73, Jor52, Kar66, Kas06, Kil11, Le 45, Lev89, Mal17, Mas77, Mil43, Not16, Pop88, Ric12, Sal89, Sch43a, Sch80b, Sei69, Som83, Spu88, Ste81, Tea62, Tra64, Tru12, Wis68, Jac82]. English [Min51, Sch03, Tay82, All49, Ang00, Bal44, Ban55, Ber74a, Bri79, Col76, Col59, Coo14, Coo81, Duh53, Eyc57, Fat05, Geo57, Gre91a, Hat68b, Hou42a, Hou42b, Itz78, Jon40, Kit66, Kli81, Koh40, Lev03, Loc04, Lum41, Lyn00, Mac70, Mal40b, McC58, Orr86, Pan21, Par73, Phi53, Pop05, Pre71, Rac95, Ran66a, Red74, Ros92, Sch43b, Sch43c, Shu68, Sie52, Sin89, Smi54, Sta91, Tas69, Tin88, Voz25, Wal86, Wal81, Wei50, Whi71, Woo92, Wri40, Zal93]. Englishwoman [Nad82]. Enigma [Men18]. Enkuklios [Doo09]. Enkyklioi [Bos89]. Enkyklios [Bos89]. Enlightened [Bat00a, Bla49, Gre16, Hud56, Moy75, Red97]. Enlightenment [Hig91, Nor07, Sch14a, Bak76, Bra79, Bur64, Cha66, Cro53a, Cro85, Dav83, Dem58, Ell03, Fis75, Gar84, GS23, Gol88, Gro14, Hal74, Hoo68, Hun95, Isr06, Isr16a, Isr16b, Kne90, Lef76, Les07, Lif12, Mor22a, Mor92, Mul97, Nel00, Pri57, Pri13a, Rig86, Sch89a, Sch03, Sei69, Shu76, Sin00, Ste75, SR81, Stu97, Stu07, Sut03, Wes07, Wol16a, Zah12, vdZ10, Boa63, Koh70]. Enormity [Lev72a]. Enough [Bra24]. Entangled [Cha24]. Entanglement [Rei21]. Entertainment [Tat65]. Enthusiasm [Bur74, Kei04, Ham58]. Enthusiastic [Dav43]. Entretiens [Hic10]. Environmental [Col76]. Environmentalism [Tsa13, Woo92]. Ephesus [Fis62b]. Ephraim [Yeo96, Yeo03]. Epic [Bil11, RS68, Ste62]. Epicurean [Alb78]. Epicureanism [Hol17, Joy92, Mer49b]. Epigram [Rus93]. Epileptic [Dah15]. Episode [Feu79, Shu76, Smi80a]. Episteme [Cla82, Mac98]. Epistemic [Aer19]. Epistemological [Bed21]. Epistemology [Bau70, Car03, Hip04, Joh04, Ode65, Pea14]. Equalitarianism [Zee46]. Equality [Con79, Fin72, Man56, Rif53, Saa10, Stu04]. Equity [Kli06, McC58]. Equivocacy [Bou01]. Era [Col76, Ros12, Zub18]. Erasmian [Ede98]. Erasmus [Cas47, Chr71, Cou04, Fer54, Fer73, Gar55, Gou10, Har71, Kos24, Lev97, Mar76, Pri64b, Ric50, Ros71b]. Erastianism [Mar09]. Erasure [Wil22a]. Eric [Ger15, Por02]. Ernest [Bev93]. Ernst [Hol71, MT13, Pri57, Ton64, Wis97]. Errata [Ano45f, Ano61d, Ano69g, Ano75g, Ano81e, Ano81f, Ano84f, Ano84e, Ano86e, Ano86f, Ano88e, Ano00e, Ano00f]. Erratum [Ano56e, Ano66f, Ano67g, Ano70g, Ano73e, Ano77e, Ano83f, Ano85e, Ano86g, Ano87e, Ano88f]. Error [Wil97]. Erudite [Tho07]. Eschatology [Lev72b, Sty15]. Esoteric [Rus93, Ter18]. Esotericism [Bag92]. Espana [Fol58]. Esprit [Mil76]. Essais [Lo17]. Essay [Ano45o, Ber71, Bos70, Bro92, Bro99, Cla82, Duh53, Gro14, Jum12, Kel84, Nak22, Puz59, Ram01, Rog78, Sei87, Ski76, Sny86, Str23, Vog93, Wal93, Wal04, Wel55, You75]. Essay-Review [Str23]. Essays [Gos60, Gru73, Spe48]. Essence [Kan13]. Essenism [Elu98]. Est [HL65, Tie63]. Established [CR81b, Las47]. Establishing [San60]. Establishment [All85, Tay55]. Establishments [Sus12]. Esteem [Saa10]. Estimation [Mos60]. Estonia [Vel19]. Estrangement [Odo66]. etate [Not16]. Eternal [CJ60, Löw45, Osl85, Pra65]. Eternity [Dal88]. Ethic [Ada89, Con79, Gal94]. Ethical [Bok17, Eck44, Mul63, Nak22, Sat82, Thr11]. Ethics [Bar98, Boy06, Dav10, Eas58, Elt97, Gue89, O’B58, Pet84, Pri17a, Sch91a, Ste62, Von81, War11, Was84, And11]. Ethnography [Day08, Nev09, Rub06, Sta04]. Ethnohistory [Bid54]. Ethnology [Mil06]. Ethnopluralism [Åsb21]. Ethology [Lea82]. Étienne [Keo77]. Etymologic [Jac83]. Etymology [Bor68, Jac83, Ste67a]. Eucharist [Tas69]. Eudaimonism [Gas00]. Eugenics [Pau84]. Euhemerism [Sch23a]. Euhemerus [Sch23a]. Euripides [Atk14]. Eurocentrism [Sho16]. Europe [Ger25, Bur03, Cla11, DeV08, Efr97, Gil20, Goo11, Gou06a, Har02b, Kri63, LS17, Mer11, Mil96, Pop06, Sch84, She03, She06b, Sin00, Smi19, Sol03b, Tho94, TA06, Fel77, Nak69]. European [BBGM09, Car19, Dav83, Fal54, Fro07, Mas00, Mor89, Ned05, Ned13, NS16, Spi76, Wil70a]. Europeans [Ger25]. Eusèbe [Zwi20]. Evaluates [Gra53]. Evaluation [Bar43, Lov49, Mor45]. evangelica [She02]. Evangelicalism [WB93]. Evangelism [Jun53]. Evangelizing [Lev71]. Eve [Cla57, Fin14, Koh51, Ste65, Tun14]. Evelyn [Hou41]. Events [Teg42, Yos51]. Ever [Alb17]. Everything [Cla23]. Everywhere [Sma83]. Evidence [Bur03, Dia82, Fay78, Mil06, Pic89, Rap97, Sol03b]. Evil [Cat73a, Cat73b, Fix89, Her70, Ken71, Lat94, Men18, Sty15]. Evils [Tay55]. Evolution [Bak77, Bes48, Bow75a, Bow75b, Bro91, Feu79, Fin14, Goo64, Har71, Ker24, Pri64b, Ric79, Sei69, Sti09, Wel67, Mor40, Sch50]. Evolutional [Ste78b]. Evolutionary [Ami07, Man57a, Phi07]. Evolutionism [Ano46g, Lov46b, Wie46a]. Evolutionist [Blo84]. Evremond [Len02, Mor04a]. Ex [Bat00b]. Exact [Smi17, Swe93]. Examination [All66, Hab95, Joh58b, Tan99]. Examiner [Hai50]. Example [Dah15, Ros96a, Sal89]. Examples [Ham98]. Exasperating [For16]. Exceptionalism [Fef05]. Excerpts [Rom64]. Exchange [Mar10, Mar09, Per02]. Exclusion [Ber18]. Executive [Kel05a]. Exegesis [Haw94, Sac18]. Exemplarities [Cor98]. Exemplarity [JW98, Rig98, Sti98]. Exercise [For05, Bui02]. Exercitium [Gro79]. Exhibition [Wie40]. Exile [Ben23, Lev50, Mul09]. Exist [Ban14]. Existence [Cre83, Res60, Val95]. Existential [Til44]. Exoterikoi [Bos89]. Expansion [Deu45, Tra64]. Expediency [Ben82b]. Experience [Doo99, Gew41, McA10, Meg75, Ode65, Pal09, Tow87, Val04, Wil23, vdZ98]. Experiment [Moo51a, Moo51b, Suł89, Tri91]. Experimental [Dra70, Eas67, HH66, Hen01, Joh59a]. Experimentalism [Rat92]. Experimentation [Gue89]. Experiments [Tou57]. Experts [Sch14b]. Explained [Wan05]. Explaining [Elu02, Wal98]. Explains [Tou98]. Explanation [McL86, Pat01]. Explications [Str57]. Exploitation [Wea80]. Exploration [Cur66]. Exposition [Coa61, Tod76]. Expressing [Boe83]. Expression [Bar75, Ber78b, Duh49, Pit86]. Expressionism [Mor96, Nor08]. extant [Guc00]. Extension [Nas46]. Extermination [Ila12]. External [Bro92]. Extradeical [Wol61]. Extraterrestrial [Dic80, Hef81]. Extremism [LaC70, Por02]. Eye [Bro85, Ros71b]. Eyes [Cor18, Der10].

F [Bel15, Bri42, Nor04]. F. [Gre15, O’C64, Ran66a]. Faber [Hun72a]. Fabians [Irv47, Mac55]. Fable [Ber67, Man08]. Fables [Wes11]. Faces [Mos66]. Facing [NS16]. Fact [Hoo73, Lev03, Loe46, Suł89]. Factions [Ste81]. Facts [Nad19, Web81, Cha24, Mos41, Wol07]. Factum [Mor78c, Min98]. Faerie [Fei68, Gre74, Kas06]. Failed [Pau83]. Failure [Pal72]. Fair [BS21, Gal17]. Faire [Alb55, Coa50a, Fra78, Jen44, Kit66, Ros63]. Fairies [Web81]. Faith [Bar81, Cro10, Dei54, Der10, Liu12, McC79, Ned13, Osb02, Phi07, Rob75, Sga15, Sny86, Str54, Wal84a]. Fall [Bro81, Elu02, Goo84, Joh59b, Lau00, Rum85, Ulr71, Ste60]. Fallacy [Smi50]. Fallen [Wei75]. Fallibility [Sch07b]. Falsehood [Cro53a]. Falsification [Ano77e, McL76a]. Fame [Boy06]. Family [Ber18, Cam22a]. Fanon [Rob20a]. Farabi [Fak65, Res63]. Farewell [Joh22]. Fascination [Deg12]. Fascism [Coa50b, Åsb21, Ada80, Kit73, Rei21, Sou70, Mos66]. Fascisms [HG21]. Fascist [Yar42]. Fate [Dan60, Gou74, Gre15, Lar10, Lea82, Men16, Val09]. Father [Zhu02]. Fathers [Ano69g, Nor47, Smi69]. Faustus [Mor09]. Favorite [Lev47a]. Fay [Car78]. faysoms [Tan98]. Fear [Cha23b, Kap08, Kja02]. Feasting [Mac71a]. Federalism [Rik57]. Federalist [Pol67]. Federico [Cla15]. Feeling [Ros12]. Feet [Mar53]. Felden [Bla13]. Felicity [Ste62]. Felix [Par91]. Female [Cou04, Kim14]. Feminine [Val09]. Femininity [And02]. Feminism [Kim14, Wil19]. Feminist [Gol82b, Ham08, Nad82, Par08, Vic08]. Ferenczi [Spe96]. Ferguson [Bar58a, BS21, Hil01, Mos60, Phi52]. Festschrift [Pri63]. Feudal [Skj22]. Feuerbach [Ber69, Che63, Cro71, Ors69, Ste78a]. Fichte [Wel84, Bre01a, Cro09, Koh49]. Ficino [All75, Bur44, DJ69, Jac85, Kri40, Kri44, Sny11, Sta68, Cas45]. Fiction [Hed59, Rya09, Sta67a, Suł89, Mat43]. Fictions [Nad19]. Fideism [Mor04a]. Field [Dur43, Hui17, Sac76]. Fifteenth [Kib46, Bar43, Dur43]. Fifteenth-Century [Bar43]. Fifth [Wei42, Bur81]. Fiftieth [Gor19a]. Fifty [Wil87]. Fight [Mai21]. Figurative [Ben82a]. Figure [Bro99, Zal93]. Figuring [Lo17]. Figurists [Row56a]. Filangieri [Mae83]. Filia [Doo99]. fill’d [Kas06]. Fin [Spe98, Fal54, Glu13]. Fin-de-Siècle [Spe98, Fal54]. Final [Ban14, Blo79, Pas90]. Fine [Kiv70]. Finitude [Ken71]. Fire [Smi66]. First [Bar50, Ber74a, Cra44, Fin14, Fro07, Goo02, Goo18, Lan77, Mos51, New95, Rap21, Rem79, Rid62, Row56b, Ruk18, Ste60, Tho55, Wil67, Pan21, Ioz22]. Fiscal [Liz22]. Fisher [Ano90m, Rik54]. Fishes [Cop91]. ’Fit [Han60, Rin41]. Fittest [Ban70, Cla00]. Five [Ano48a, Bal56, Gri56]. FL. [Gre80]. Flacius [Lyo03]. Flaubert [Neu47, Lev72a]. Flight [Bro74]. Florence [Tri56, Moy03a, Qui10, Tri91]. Florentine [Bar47, Mom50, Cel01, Gre67, Ian08, Jur01, Jur02, Phi83]. Florists [McG03]. Fly [Hes12]. Fly-Catchers [Hes12]. Flying [Mer66, Wil71]. Fog [Vog93]. Foigny [Kno63]. folio [Mar10]. Folk [She00]. Folktales [Web81]. Follow [Suł89]. Following [Mye08]. Folly [Chr71]. Fontenelle [Adk00, Mar59a]. Food [Ros24]. Fool [Sch96]. Force [Bar41, Ell62, Mir77a, Ver22, Wil08, vE01, Sav44]. Forced [Cam22a]. Forces [Coa50b, SC40, Coh40]. Foreign [God08, Hol21]. Forerunner [Dot86]. Foreshadowing [Joh67b]. Forest [Smi76b]. Forged [Ros77]. Forgetfulness [Ada92]. Forging [Yar42, Han91]. Forgiveness [New01]. Forgotten [Bur60, Ros18, Sou89]. Forkosch [Ano03h, Ano09g, Ano15e]. Form [Boo43, Cra44, Hes12, Hui43, Kat44, Mer47, Pat58]. Formal [dS23]. Formalism [Erl73, Whi47]. Formalist [Gar70]. Formation [Cla86, Dym21, Lan77, Red97, Sor83]. Former [Liu12]. Forms [Ver69, vR00]. Formulating [Ada16]. Forth [Mac71a]. Forthcoming [Ano80g, Ano80h, Ano80i]. Fortuna [Bel08]. Fortunate [Ulr71]. Fortune [Ned99]. Fortunes [Spe55]. Fortunio [Del17]. Forty [Han95]. Forum [GSJ18, Spe13]. Forward [Jum12]. Fossils [She03]. Foucault [Eld24, Mac98, Meg87, Mon19]. Foucaultian [Sei90]. Found [Kel12]. Foundation [Bis15, Sel63]. Foundations [Mie73, Mil97, Osl85, Kel79]. Founders [Sch50]. Founding [San62]. Four [Cam86, Hor74, RS64, Tod76]. Fourier [Bru76, Gol82b]. Fourierism [Gua82]. Fourteenth [Sho99, Gre67, Kib46, Tan98]. Fourteenth-Century [Gre67, Tan98]. Fourth [McM66]. Fragment [Eng63, She00]. Framework [Ruo66]. Francaise [Vir50, Wil68, Wil73]. France [Coa50b, Arc67, Bar41, Doe22, Eme84, Fal20, For93, Her92, Kel81, Keo74, Koh56, Lan80, Leg25, McC15, McC79, Mor40, O’B85, Pal40, Par01, Pek99, Por72, Rid62, Rin61, Sei69, Sim65, Sol03a, Som67, Syp63]. Francesco [Del80, Blu00, Jur02]. Francis [Skj22, Bis96, Bow74, Kel80, Koc57, Lin74, Mil97, Par08, Pes14, Poo05, Ser17, Vic92, Vic08, Whi89]. Francisco [Spi80, Wei19b, Wes15]. Franco [Ter21]. Francois [Hig91, Ano64j, Kel64, Stu97, Wil19]. Franklin [Dra77b, Mae75, Sin00]. Frankness [Wes11]. Frantz [Rob20a]. Franz [CH15, Con14, Jon92b, Sch07a]. Frazer [Ack75]. Frederic [Hor97]. Frederick [Eki50, Mom50]. Free [Bos73, Dam79, Eki50, Goo83, Gra65, Men16, Mie73, Mol97, Mor18, Ned99, Ram19, Red74, Rem79, Sta08a, Tri49, Van58, Ver22]. Free-Will [Dam79]. Freedom [Ber78b, Cra45, Eck44, Fal88, Fat05, Gay63, Gue95, Har72, Jac12, Kne90, Lac99, Lau98, NL20, Pet12, Pit86, Qua19, Rob75, Roe03, San60, Sav44, Sea80, Smi80a, Ter21, Voz25, Woo24, Gre15]. Freethought [Rus93]. Freitas [Vie03]. French [Cen89, Ano84f, Atk74, Axi71, Bea82, Beh08, BI60, Ber87, Bos73, Bre10, Bro24, Cha66, Cla57, Clo84, Coc62, Dav60b, Ede18, Eyc57, Gar73, Ger51, Gol82b, Got44, Hel73, Hut76, Jaf80, Jen07a, Kar69, Kel65, Kim18, Koy98, Man88, Man91, Mat06, McL76b, McN45, Mil76, Mor92, Mos85, Mur58, O’N07, Oos13, Pey49, Pla13, Rid56, Ros98, Sch02, Sta87, Sta04, Sut03, Tho94, Urb12, Var50, Vir50, Wat71, Zam98]. Freud [Dia82, Eng80, Ful83, Gel89, Gup75, Loe70, Mer45, Mer49a, Rat86, Rie56, Sch56a, Spe96, Tri04, Wie56]. Friedrich [Don82, Eck45, Les57, Pfa90, Pur15, TA06, Wei52, Wel73, Wol56]. Friend [Mon48, Pol67]. Friends [Ede98, Gue78]. Friendship [Feu79, TW21]. Frightful [Mac67]. Frobenius [San20]. Front [Ano40d, Ano40e, Ano40f, Ano41f, Ano45g, Ano45h, Ano45i, Ano45j, Ano46h, Ano46i, Ano46j, Ano46k, Ano47g, Ano47h, Ano47i, Ano47j, Ano48f, Ano48g, Ano48h, Ano48i, Ano52e, Ano53b, Ano54c, Ano55c, Ano56f, Ano57h, Ano57g, Ano57i, Ano57j, Ano57k, Ano58g, Ano58h, Ano58i, Ano58j, Ano59h, Ano59i, Ano59j, Ano59k, Ano60b, Ano61e, Ano62f, Ano62g, Ano62h, Ano62i, Ano63g, Ano63h, Ano63i, Ano63j, Ano64f, Ano64g, Ano64h, Ano64i, Ano65f, Ano65g, Ano65h, Ano65i, Ano66g, Ano66h, Ano66i, Ano66j, Ano67h, Ano67i, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano69h, Ano69i, Ano69j, Ano69k, Ano70h, Ano70i, Ano70j, Ano70k, Ano71f, Ano71g, Ano71h, Ano71i, Ano72e, Ano72f, Ano72g, Ano72h, Ano73f, Ano73g, Ano73h, Ano73i]. Front [Ano74e, Ano74f, Ano74g, Ano74h, Ano75h, Ano75i, Ano75j, Ano75k, Ano76e, Ano76g, Ano76f, Ano76h, Ano76i, Ano77f, Ano77g, Ano77h, Ano77i, Ano78e, Ano78f, Ano78g, Ano78h, Ano79c, Ano79d, Ano79e, Ano79f, Ano80j, Ano80k, Ano80l, Ano80m, Ano81g, Ano81h, Ano81i, Ano81j, Ano82e, Ano82f, Ano82g, Ano82h, Ano83g, Ano83h, Ano83i, Ano83j, Ano84g, Ano84h, Ano84i, Ano84j, Ano85f, Ano85g, Ano85h, Ano85i, Ano86h, Ano86i, Ano86j, Ano86k, Ano87f, Ano87g, Ano87h, Ano87i, Ano88g, Ano88h, Ano88i, Ano88j, Ano89e, Ano89f, Ano89g, Ano89h, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h, Ano92f, Ano92g, Ano92h, Ano92i, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f, Ano95g]. Front [Ano95h, Bed21, Cha23a]. Frontier [Hay56, Kes68]. Frontinus [Woo67]. Frustration [Sca97]. Fry [Coo81]. Frye [Mur54]. Fu [Wri51]. Fuchs [Kus97]. full [Kas06]. Fuller [Kea64]. Function [Chr46, Duh49, Ede49, Kat44, Lie98]. Functional [Pre79, Wol50]. Functionalist [Kar84]. Functions [Cas47, Dav51]. Fundamenta [Ver94]. Fundamental [Hsu02, Kla93]. Fundus [Nan21]. Funkenstein [Moy03b, Soc06]. Furet [Hig91]. Furniture [Nor75]. Furor [Tig70]. Further [Bak77, Cla86, How67, Igg59]. Fuseli [All74]. Fusion [Had58b]. Fustel [Wil73]. Future [And21, Ger25, Kim24a, MB95, Smi24b, Son74].

G [All74, Bet09, Bou97, Gut79, Har72, Hyd56, Isr16a, Koh64a, Men56, Mer48, Not55, Pri48, Ski25, Sti09, Str54, Wol16b]. G. [Bre52b, Por78, Wie56]. Gabriel [Pop05]. Gabrielle [Wil19]. Gadamer [Lam91]. Gaetano [Mae83, Puz59]. Gainsborough [WA97]. Galiani [Mer50b]. Galilean [Tuv66]. Galileo [Bro85, Dra77a, Duc06, Gir70, Gro11, Koe77, Koy43, Moo51a, Moo51b, Rap16, Ree91, Ros78, Rud01, Smi76a, Sta68, Wal71, Wal88, Ros57a]. Gambling [Som98]. Game [Ant68, Gal17]. Games [Col16, Hal02]. Gandhi [Men20, Mur68]. Gans [Bre01b]. Garde [Tan98]. Garden [Bal50, Hed98, Neu47, Spe98, Ste67b]. Gardens [Com71]. Garve [vdZ98]. Gases [HH66]. Gassendi [Gli88, Kro84, MM90, Osl83, Sak09, Sar85]. Gaudin [Hal74]. Gay [Eme67, Sch54]. Gefühls [Cro09]. Geistesgeschichte [Spi44]. Gelais [Rid62]. Gender [Ber18, Hil11, Vel19]. Genealogical [Lan12a]. Genealogy [Bla01, Fis54, Gor19b, Liz22, Man94, Van17]. General [Nad01, Wil05]. Generation [Goo02, Roe85]. Genesis [Koh40, Wie46c, Wil48, Zil45]. Genetics [Mik70, VR05]. Genius [Ala77, Cat73a, Dah15, Die41, Jaf80, Ken71, Tow19]. Genotype [Gud96]. Genre [Lef76, Low24]. Gentile [Gre63]. gentium [Wau20]. Geoffrey [Mar19]. Geography [Cav16, Fac21, Wit09, May04]. Geologist [Por78]. Geology [Swi75]. Geometrical [Mal97]. Geometrically [She02]. Geometry [Bir96]. Georg [Ale66, Cal12, Dau77, Lam91, Nak22, Sei42]. George [Rik54, Ano81e, Bar75, Boc51, Con61, Ebe75, Fef90, Gom81, Hep55, Kam52, Kar84, Lew65, Nor75, Pop52, Rya09, Smi41, Sol70, Sta91, Wes77, Wie80a]. Georgian [Wit43]. Georgius [Han92]. Gerard [Spr80]. Gerbert [Tru12]. Gerbi [Boa56]. Germaine [Hil11]. German [Fri40, Neu67, Alt07, And41, Dum81, Eyc57, Fir71, Fis75, Gus50, Hos24, Joh22, Kne90, Koc21, Koh51, Kon97, Kue42, Lev87b, Lid66, Lin10, Mac88, MB81, Mit71b, Mor96, Par88b, Pic89, Rah22, Ram19, Red97, Sak09, Sch07a, Van95, Vin07, Von48, Whi86, vdZ10]. Germans [Ger25, La 93]. Germany [Ant21, Bau23, Bra71, Bro05, Cre95, De 53, Esk04, Hen95, Kle99, Koh56, Las47, Lee79, Lut70, Mul06, Pan21, Pop88, Pur15, Vic02, Wei93, Wei02, Wer13]. Gerrard [Van65, Jur75]. Gerson [Rue64]. Gersonides [Nad01]. Gervase [Byr23]. Geschichte [All74]. Geschichtliche [RR06]. Gesture [Kno65]. Ghazali [Hal02]. Ghost [Tin96]. Giambattista [DH71, Lev91, Min98]. Gibbon [Bro77, Bro81, For16, Ros16a]. Giddens [CH95]. Gilbert [Boa51, Hen01, Par91, Zil41]. Gildon [And55]. Giles [McA99]. Gilman [Cop51]. Giordano [Mai77, Mas77]. Giorgio [Ger72, Gre23b]. Giovanni [Nau11, Bre52a, Cas42a, Cas42b, For42, Gre63]. Girardin [MR20]. Girolamo [Cor18, Pri17b, Sir91, Sir03]. Giulio [Fri12]. Giving [Hsu02]. Glanvill [Pop54, Pop53a]. Glass [Yol79a]. Glicksman [Str22]. Global [Ach20, HTW21, Kel05b, Tad22, Tun14]. Globalization [Meg05, Sch05a]. Globalized [Tad22]. Glorious [Sch90]. Glory [Kuh61]. Gnosticism [Laz03, Wei19a]. Gnostics [Kat54]. God [Ben84, Nad93, Afr62, Bro99, Bur93, Col62, Ell03, Fon99, Gor08, Hed98, Koc50, Lac99, Mom51, Nad91, Ned99, Oak98b, Oak98a, Osl79, Smi61, Som98, Str52, Tra64, Von84]. Godefroy [Tho07]. Godless [Day08]. Gods [Mac06, Wal98, Mul09]. Godwin [Lam09, McC20, Mon74, Pie10, Pol64, Pol67, Pre54, Ros70, Scr78]. Goethe [Kau49, Lar67, Lau90, Nau52, Neu47, Sch96, Wel67, Wel71]. Gog [Poc87]. Goguet [Wol07]. Going [Mae80]. Gold [Alb17, Fog12]. Golden [Gri58]. Goldsmith [Lov46a, Lyn45]. Good [Jac12, Ran67, Sei74, Ven76, Wal63]. Goodness [Mae83]. goods [Woo24]. Got [Cas10]. Gothicism [Nev09]. Gottfried [ZMM10]. Goulart [Gar13]. Government [Boy06, Fac21, HM79, Leb86, Lee11, Mor54, Nic19, Was72, Wat72, Zuc75]. Grace [Col16, Mon44, Ned99]. Gradual [Lin91]. Graetz [Elu98]. Grammar [Nau11]. Grammaticus [Mal58, Del96]. Gramsci [Fon00, Ada80, Bat75, Tod74, Wil60]. Grand [Dem72, Pec96]. Grant [Bur11]. Gratian [Tie01b]. Gratitudo [Gal94]. Gravitation [Tho58]. Gravity [Coh87, Kuh61]. Gray [Man94]. Great [Mor57, Shu76, Thi79, Was73, Wil80, Wil87, dM49, Ano70g, Bor71, HP18, Hor76, Kes68, Lar10, Lew58, Lin48, Pre03, Ser17, Sha02, Lac10]. Greatest [Alb17, Har46]. Greaves [Sha02]. Greco [Had43]. Greco-Roman [Had43]. Greece [Ger25, KM94, Chr54, Str16, Ces05]. Greek [Atk14, Ber78b, Cuf66, Cum73, Fen86, Fur81, Gom43, Hal92, Hol49, Jer13, Lar41, Mul79, Nor47, Pea54, Say45, SF72, Smi73, Sol63, Str23, Tig70, Urq14, Vic02, Wal98, Whe55]. Greeks [Ger25]. Green [And11, Ran66a, Wie48]. Greenberg [Ger25]. Gregorian [Whi60]. Gregory [Dun64, HP18, Wie63]. Grendler [Bla91b]. Grene [Str22]. Gresham [Joh40]. Grimké [Eki50]. Grondt [Mie73]. Grosseteste [Eas67]. Grotius [Ach20, Jon13, Lee11, Mau05, Rub91, Vie03, vI11]. Grouchy [Ber18, Kim24b]. Ground [Cot94, Hip04, Mar53]. Group [Nis44]. Grows [Boc65]. Growth [Sch60, Sch44]. Grundbegriffe [RR06]. Grundtvig [Mal40a]. Guerre [Goo90a]. Guido [Kel12]. Guillaume [Mon97]. Gulliver [Gou45, McM66]. Gunpowder [Wol70]. Gustav [Gil83]. Guyau [Fid94]. Gypsies [vG20].

H [And11, Fal54, Fer68, Kel97, Men56, Por78, Pri63, Ran66a, Sta09b, Sta57]. H. [Hyd56, Lal54, Pri48, Rot90, Sti09, Wil08]. Habermas [Deg23]. Habet [KS24]. Habit [Wol50]. Habits [BB60]. Haeckel [Hol71]. Häfner [Mul09]. Hagiography [Wei19b]. Hajime [EST66]. Hakewill [Hep55]. Haldane [Vin07]. Halévy [Fro07, Rem79]. Half [Tho55]. Halifax [Sta41]. Haller [Son74]. Halley [Alb78]. Hamann [Ale66, Bet09, Mer48, Mer49c, Mer50b]. Hamanniana [Mer49c, Mer48, Mer50b]. Hamburg [Lev13]. Hamilakis [Ger25]. Hamilton [Kar65]. Hamlet [Fei46]. Hammer [Puc05]. Hampton [Cor98]. Hand [Har11, Mac71b]. Hanging [Kil11]. Hannah [Kan13, Sim16, Bri00, Kee12, Moy23]. Hans [Boa58, Tri56, Ada91, Bri00, Fer58, Gor19a, Hel19, Lam91, Laz03, Pal97, Sty19, Wol76a]. Hans-Georg [Lam91]. Happening [Kel90]. Happiness [Car07, Psa60]. Harclay [Dal84]. Hard [Web81]. Hariot [Kar66]. Harlow [Gar54]. Harmony [CR81b, Gri74, Sco98, Was72]. Harrington [Coa77, Coh94, Cot81]. Harriot [Ham96]. Harry [Eng95, Mer89]. Harsdorffer [Wes99]. Hartley [Hat75, Hav59, Les72, Mar59b, Obe76]. Hartz [Gab56]. Harvard [Ken90, Nar05, Ros12]. Harvey [Bat00b, Coh94, Gor94, Pop05]. Hastings [Sav44]. Hattaway [Hor82]. Haunted [Jen20]. Havelock [Hal92]. Having [De 06]. Hawthorne [Fis58]. Haydn [Kri51a]. Hayek [Jac12]. Hayim [Scl20]. Hazard [Har72]. Hazards [Eng07]. Hazlitt [Lar41, MB95, Par69b]. Head [Tru12]. Health [Rom61]. Heart [Boy02]. Hearts [Maz14]. Heathens [Eig09]. Heaven [Hea88]. Heavenly [Eme67]. Heavens [Ano86g, Ber86a, Dal80, Met96, Ree91, Sch23b]. Hebraic [Neh77]. Hebrew [Sap79, Cul70, Sim47, Twi20]. Hegel [Avi61, Ber69, Ber77, Chr64, De 79, Fle47, Ger19, Hof85, Hun86, Lau90, Mue58, Ors69, Par77, Roy07, Sch96, Sil01, Smi65a, Tei77, Ver69, Coh49, Cri66]. Hegelian [Got80, Ste85a, Wil56]. Hegelianism [Bou90, Bre01b, Eas62]. Hegeliano [And11]. Hegelians [Sch03]. Hegemony [Bat75, Fon00]. Heidegger [Gra53, Kol91, Lam91, Von84]. Heinrich [Cro10, Elu98]. Heir [Sak09]. Heirs [Acw77]. Heisenberg [Bru76, O’F92]. Helena [Isr16b]. Hell [Tur66]. Hellenic [Dem58]. Hellenistic [Gli88, Had50, Had58b]. Helvetius [Lad62, McC63, Blo79]. Hempel [Dew18]. Henry [Boc76, May88, All74, Alm93, Ber62, Bou90, Bur87, Dal84, Gre62, Jac85, Kam52, Min65, Par88b, Pow70, Smi76b, Sta68]. Heptameron [Ham98]. Her [Ber18]. Heraclitean [Hel64]. Heraclitus [Fis62b, Smi66]. Herakleitos [Sin63]. Heralds [Joh04]. Herbart [Bel15]. Herbert [Fef90, Kar84, Ada16, Bak77, Bow75b, Fra78, Fra91, Hob71, Pau83, Sew03, Sim60, Wes77, Wie80c]. Herborn [Loe61b]. Hercules [Sak09]. Herder [All74, Bar63, Bar83, Blu63, Ges65, Sch56b, Spi55b, Wel58, Wel60, ZMM10]. Heresies [Hun60, Wes53]. Heresy [Min68, Mye08, Spr94, Wil70b]. Heretical [Mye08]. Heretics [Rud01]. heretofore [Guc00]. Heritage [Bri53, Clo75, Eva65]. Herman [Boa63, KNS^{+}81]. Hermann [Gui08]. Hermaphrodite [DeV08]. Hermeneutical [Wal96]. Hermeneutics [Mar90, Rud01, Ste85b]. Hermetic [Tin54]. Hernando [Kim21]. Hero [Cam84, Lev56]. Herodotus [LS23]. Heroes [Rud01]. Heroic [Doj05]. Heroine [Moy75]. Herschel [Can61, Ros68, Wil74b]. Herzen [Kel80]. Hesiod [Bal56, Bea91, Gri56, Gri58, Teg47]. Heterodoxies [Sar08]. Heumann [Mah15]. Hexter [Fer68, Kel97]. Heydenreich [Cro09]. Hickes [Hia09]. Hidden [Loe70]. Hierarchy [Con79, Mah87, Moh04, Pat59]. Hierarchy2 [Loc07]. Hieroglyphs [Sin89]. High [Mit94]. Higher [Chr59, Eng07]. Highgate [Ban16]. Hilaire [Tre21]. Hind [Ben82a]. Hindu [Pra55]. Hinduism [Goo90b]. Hintikka [GM80]. Hinweisen [Ano98e]. Hiram [Kri51a]. Hispanic [Hob80]. Histoire [Lef76, Ans09, Syp63]. Histoires [Gar13]. Historia [Ano64j, Kel64, Spr94]. Historian [BI60, Boa48a, Lov41, Mar86, Mos41, New95, Ron12, Wis97, Wol76a, Pau72, Rat41].

Historians [Bea93, Bon56, Boo00, Bur49, Cha23b, Gar75, Ger51, Glo57, Joh71b, Kri53, Mac88, Meg87, Mos85, Moy03a, Pla13, Pre71, Sch05b, dS23, Fri40]. Historical [Aer19, Atl52, Bac07, Bar50, Bat74, Ber87, Bow85, Bro04a, Bro81, Bur57b, Car03, Cle69, Coh80, Con14, Cou44, Cro05, De 03, Dei54, Dun21a, Elu92, Fer71, For82, Fre64, Gra44, Gre67, Gug80, Had79, Hea87, Hol50, Hut92, Kra97, Lev87b, McC94, Men61, Mon19, Mor20, Nak22, Nev09, Pre77, Pri48, Ram01, Reh14, Rey53, Sch56a, Sch43c, Sim81, Sol03b, Sol83, Str05, Swi75, Syp63, Syp65, Teg42, Tol98, Ton59, Ton64, Tor84, Van58, Ven76, Vir86, Whi45b, Wom86, Woo88, Car08]. Historicism [Bev12, Bre61, Gre16, Igg95, Mor78c, Moy03b, cN93, Poi70, Ran64, Soc06]. Historicisms [Bou03]. Historicist [Pau83]. Historicity [Ant21, Mit94]. Historicization [Rub06]. Historicizing [Zak24]. Histories [Ant24, Bac07, Bou03, Cha24, Gou06a, Gou06b, Ian08, Kim24a, Pop06, Rot90, Sko64, Tsa13, Van16]. Historik [Gil83]. Historiographical [Cel01, Puz59, Sou00]. Historiography [Bar41, Bec85, Cal97, How99, Liu04, Lov49, Man48, Man55, May90, Ned05, Par08, Phi83, Phi89, Phi96, Pim18, Sol82, Spi52, Tri50, Urq14, Vic08, Wer93, Wol07]. History [Aga69, Aga73, Ala67, Alt08, Ano56e, Ano60c, Ano64j, Ano03f, Ban70, Bar16, Bar63, Bar50, Bar59b, Bar45, Bat24, Bea76, Bet59, Bet09, Bla82, Blu00, Boa40, Boa47, Boa58, Boa64, Boc51, Bre01a, Bre10, Bro72, Bro77, BS21, Bur87, Byr06, Car86, Cas42a, Cas42b, Cat10, Cen89, Chr64, Coe16, Coh40, Coh42, Coh54, CH95, Con82, Coo14, Cop91, Cot08a, Cot08b, Cra40, Cra45, Cro66, Cum64, Dan60, De 79, Dew18, Dig06, Doj05, Doo09, Doo90, Eas67, Ede46, Ede22, Ege66, Elu92, Erl02, Far91, FR07, Fir71, Fis47, Fis70, Fri40, Fro07, Fro95, Gar70, Ger72, Ger19, Gil86, GS23, Got75, Gra75, Gra06, Gro99, Gub05, Guy20, Har90, Har11, Hat68a, Hau98, Hau11, Hed59, HH66, Hel19, HTW21, Hig54, Hil72, Hob71, Hol48, Hol49]. History [Hol53, Hol40, Hou41, Hui43, Hun72a, Hut76, Igg95, Isr19, Jan08, Joh46, Joh59a, Jon67, JN12, Jum08, Jur02, Kan13, Kat44, Kat84, Kau49, Kau01, Kel64, Kel87, Kel90, Kel01, Kel05b, Kel68, Ker24, Kir50, Kir94, Koh64b, Kon97, Koo99, Kri60, Kri70, Kri73, Kri46, Kri51b, Kri52, Kuh49, Kva77, Lac10, Lan95, Lan66, Lev05, Lew65, Lou56, Lov40, Mac67, Mac70, Mac88, Mac01, Mal40a, Man42, Man13, Mar12, Mar20, Mas52, Mat57, Maz14, Maz72, MM22, McK75, Meg05, Mil06, Mon19, Mor77b, Moy07, Mud68, Mul23, Mun79, Nad62, Nad01, Nak22, Nel87, New95, O’N12, Oak55, Ort18, Osm73, Pal09, Pal18b, PB73, Par44, PD63, Par07, Pea48, Pea54, Pet88, Phi52, Phi96, Pim18, Pod10, Pra55, Ran66b, Reg05, Rei96, Ric87]. History [Rin41, Rom48, Ros92, Ros43b, Ros55, Ros18, Rou44, Rut14, San44, San20, Sch05a, Sea12, Shk81, Sho16, Sim56, Sir03, Smi95, Smi17, Smi24a, Smi24b, Sol03a, Spe48, Spi55a, Spi55b, Spi44, Ste78b, Ste85b, Ste18, Ste58, Sto61, Str49, Str51a, Str54, Str16, Sty15, Swe93, Tae42, Teg40, Ter21, Thi79, Tho12, Tru12, Val24, Van17, Von58, Wei19a, Wei69, Wei45, Wel60, Wer75, Wes07, Whi42, Wie46b, Wie61, Wil22a, Wil68, Wil85, Wit09, Wol76b, Zea59, Zwi20, dS23, vdZ10, GB87, Pal06, Pal18a, Bri42, Sav44, Ano69l, Bou97, Cas45, Ces05, Had58a, Har90, Kni64, Pag88, Par62, Ran49, Ran23, Rie61, Sch41]. Histrionics [Ban70]. Hitherto [O’F92, Rap69]. Hitler [Bar42]. Hitlerism [Spi44]. Hobbes [Bir96, Lun87, AA85, App64, Bar81, But06, Col09, Cot81, Cri02, Dam79, Gan56, Gei91, Ger67, Gra78, Har86, Her87, Hey82, Hoe15, Kee12, Lev54, Lub24, Mac70, Mal15, Mar09, Min68, Min70, Mis77, Mis83, Nau02, Par15, Pas90, Pat01, Pec96, Pet12, Pyc87, Rus85, Sar85, Sch45a, Sel74, Spr94, Ver82, Wat55, Wil70b, Zag90, Zag93, Zag99]. Hobbesian [Joh85]. Hobhouse [Gri74, Wei96]. Hobson [Fre73, Kru55, Mit65]. Hodgson [Sta09b]. Hoffmann [Cas10]. Hofstadter [MM22, Whi45a]. Hogarth [Asf99]. Holbein [Cha60]. Hold [Fri12]. Holland [Pop88]. Holmes [Ano47f, Coh48, Rum65]. Holocaust [Fac75, Ano75g, Gri21]. Holt [Neu25]. Holy [Cam16, Cam22a, Llo69, Ruo66]. Homage [Mal99]. Home [Bed21]. Homer [Had79, The23]. Homeric [RS68]. Homiletic [Ong53]. Homme [Mor78a]. Homo [Giu85, Gre23b, Hun72a]. Homonoia [dM49]. Homosexuality [Rut14, Wat92]. Hood [Joh42]. Hooker [Alm78, Ecc76, McG63]. Hoornbeeck [Ver22]. Hope [Cas81, Kue42, Tay81]. Hopkins [Dyk61, Spr80]. Hoplite [KE23]. Horace [Mer49b]. Horatian [Bar60b, Maz60]. Horizons [Byr23, Kel87]. House [Glu13, Zet82]. Household [vI11]. Houssaye [Sol00]. Houten [Stu89]. Houyhnhnm [HL65]. Howard [Wie80b, Smi65b]. Hsiu [Sch51]. Hudibras [Nel76]. Huet [She02]. Hugh [Sol58]. Hughes [Rot90]. Hughlings [Ful83]. Hugo [Ach20, Jon13, Lee11, Rub91, vI11]. Huguenot [Her92]. Huizinga [Kat44]. Hulme [Ano86e, Shu85]. Human [Arm07, Arr94, Ber18, Bre01a, Cur61, Dal77, Day08, Dym21, Ede18, Eze00, Goo02, GSJ18, Gre15, Gri12, Her21, Lee09, McK81b, Mor09, Nee51a, Nee51b, Ran49, Ree52, Saa10, Sea11, Sho18, Tea62, Tie11, Ver82, Wei02, Wer75, Zub18]. ’Humanism [Giu85, Bar60a, BG92, Cam96, Cas47, Cer05, Con96, Cou44, Cou04, Fer58, Fub96, Gra63b, Hau11, Keo77, Lev72b, Log75, Mar90, MH04, Mul09, Phi07, Qui92, Sie52, Tin88, Tri56, Wei72]. Humanist [Bla92, Byr06, Gra88, Han92, Par68b, Sie50, Tri87, Wei16]. Humanistic [Gun62, Tri56]. Humanitarianism [Fie76]. Humanité [And02]. Humanities [BBGM09, Hol48]. Humanity [Arm07, Whi89]. Humans [Goo02]. Humanus [Giu85]. Humboldt [Sor83, Swe73]. Hume [Ber77, But59, Cot79, Den99, Des67, Dun64, Eze00, Fie96, For82, For84, Gos60, Gre43, GS91, Hal57, Har99, Hen69, Hun74, Hun72b, Hur56, Imm92, Kem00, Lau98, Len58, Lyo70, Mas00, Mat55, McR51, Mos41, Mos48, Mos60, Mos78, Nob65, Ols91, Pea80, Pop52, Pop53a, Pop93, Pri64a, Rap69, Rol71, Rus85, Rus88, Rus93, Sho52, Sie75, Sie84, Sus12, Thi92, Tsu61, Ven76, Wal84b, Wer72, Wer75, Wer93, Whi76b, Wol60, Yol79b]. Humes [Mos51]. Humphry [Rou67]. Hundred [Rob59]. Hupsous [Hil66]. Hurst [Par62]. Hurtado [Spi65]. Hutcheson [Bar92, Bis96, Fra55, Kor75]. Hutchinson [Kuh61]. Huxley [Bli82, Phi07, Sta57]. Huygens [Wol00]. Hylomorphism [Man13]. Hyperspheres [Egg99]. Hypotheses [Far87, Lau67, Mac78, Yos51]. Hypothesis [Goc99, Lal54]. Hysteria [Wri80].

Iberia [Kim21]. Iberian [LS17, Mar22, TT24]. Ibn [RS64, Gat67]. Id [Tie63]. Idea [AK23, All85, Ano56e, Ano66f, Bak70, Bat19, Bau23, Bid54, Bos89, Bou00, Byn84, Can61, Cel14, Col71, De 06, Dod68, Dod08, Gri74, Gui23, HL19, Hol40, Hor74, Joh71a, Kel50, Kem00, Kno65, Koc50, Koo99, Kri63, Las47, Les72, Lev56, Lev66, Mac92, Mar20, May92, Mil70, Mom51, Pal40, Rob20c, Run00, Sch53, Sim56, Sim47, Sin89, Smi61, Smi62a, Smi73, Sto61, Str56, Stu04, Tae42, Tho77, Tri90, Wag67, Wel55, Wol70, dBJN^{+}20, McR65]. Ideal [Dav83, Dou11, Kno63, Leg25, Osl70, Par04, Ram01, Ran67, Whi60]. Idealism [Bou94, Hos24, Ins11, Ste78b, Wil56]. Idealist [Bev93, Poo99]. Ideals [Gug80, Sei65, Vic02]. Ideas [Adk00, Ano70g, Ano83f, Ano03f, Ano14e, AG95, Arm69, Bak83, Bar16, Bet09, Bor71, Cas42a, Cas42b, Cat10, Cel14, Cha24, Cha23b, Coe16, Cre83, Cro53b, Cum64, Deu45, Doj05, Dow40, Ede46, Fis82, Gab55, Gat67, Gau92, Gli88, Gra75, Gra06, Gre72a, Gre14, Isr19, JN12, Kau49, Kel90, Kel01, Kva77, Lan95, Lov40, Lov41, Mar12, Mas77, Mas52, Maz72, McD05, McK81b, McK82b, Meg05, MM89, MM90, Mol78, Moo45, Mye75, Nob65, NL20, O’N12, Oak53, Obe76, Pal06, PB73, Par07, Pat58, Pea48, Pey49, Ric87, Rog79, Sak09, San60, Sch22a, Sch90, Sea12, Sko64, Smi62b, Spe13, Spi44, Teg40, Tho94, Tho12, Tri90, Tun14, Van17, Wei45, Wel55, Wie61, Wol61, Yol55, Zea59, dS23, Ano60c, GB87, Pal18a, Ger72, Boa64, Fis70]. Ideas [Koh64b, Moo68, Pri74, Spe48]. Ideen [All74]. Identification [Mar10]. Identity [All66, Coo14, Kel98, Lat24, Loc04, Sea12, Shu22, Ten82, Tsu61, TA06]. Ideologia [Poc87]. Ideological [Cre79, Lov67, Tod74, Pal18a]. Ideology [And41, Bal44, Bur11, Bur49, Cha23a, Chi74, Hun72a, Joh22, KM94, Lar78, Lie98, McC94, Pau71, Que63, Rou44, Van59, Ken79]. Idolatrous [Joh06]. Idolatry [Gar13, MF06, Rub06, Sie84]. Idols [Kel80, Mac06, She06a, She06b]. Idolum [Cor18]. If [Ras82, Fri12, Zhu02]. Ifa [San20]. Ifa-Board [San20]. Ignored [Gas59]. II [Kri52, Mer49c, Moo51b, Nee51b, Ano51a, Aro87, Cla23, Hat68a, Hou42b, Kos24, Par68b, She44, Wil22b]. iii [Nad91, Nad91, Rie61]. IIII [Cor18]. Illustration [Mue43]. Im [Kim14]. Image [But06, Lun83, Mor78a, Van95, Wei72]. Imagery [Tuv42]. Images [Bur03, Cal97, Jon73, Tur55]. Imaginaries [Van17]. Imagination [Ald44, Cla15, Koc21, Koc99, Mac49, Pre89, Boa67]. Imaginative [Kne90, Lea78]. Imagined [Ham96, May04]. Immanence [Hei78]. Immanuel [vG20]. Immaterial [Ruk18]. Immaterialism [And64, De 06, McK87, Sta74]. Immeasurability [Hel21a]. Immeasurable [Mai77]. Immediacy [Pfa90]. Immigration [Shu22]. Immoraliste [Fid94]. Immortality [Kri40, Ste78c]. Impact [Bar58b, Cer05, Fol58, Frä75, Hor78, Ioz22, Que63, Ran61]. Impartial [Cam20]. Imperative [Mye08]. Imperial [Bat19, Ben23, Ces04, Day08, Rei21, Woo18, Wun03]. Imperialism [Bur65, Had50, Kru55, Rem79, Sul83, Var53]. Impetus [Dra75, Fra03, Fra77]. Implements [Goo11]. Implications [Dam79, Got80, Kap08, RG70]. Implicit [Wis68]. Importance [Kus97, NL20, Sta67a, Gra75]. Importing [Gua82]. Impossibility [Wen82]. Impossible [Rut14]. Improvement [Smi17]. Inaccessible [Ram14]. Inanimate [Fog12]. Inaugural [Joh42]. Inca [Kla76]. Incarnation [Hun60, Sie84]. Incident [Som67]. Incommensurability [Sma91]. Incorrectly [Whi76b]. Independence [Ha20, Mor18, Ort18, Pal09, Pal18b, Pim18, Ras82, Pal18a]. Independent [CF79]. Indes [Ans09]. Indeterminism [Bru76]. Index [Ano02e, Ano02f, Ano03g, Ano04f, Ano05d, Ano06f, Ano07f, Ano08f, Ano09f, Ano10d, Ano11c, Ano12f, Ano13e, Ano14f, Ano15d, BBGM09]. Indexed [Kim21]. Indexing [AK23]. India [Ger25, Stu07, TA06, EST66]. India-China-Tibet-Japan [EST66]. Indian [Lun77, Mil41, Pea52, Raj52, Rie67, Tsa13]. Indigeneity [Pri48]. Indigenous [De 15]. Individual [Cla92, Öza14, Ran66b, Sta87, Van15]. Individualism [But61, Hig61, Loe61a, Nol61, Wit61, Cla86, Luk71, Swa62]. Individualization [MP04]. Indivisibles [Smi76a]. Indo [Ger25]. Indo-Europeans [Ger25]. Indoctrination [Som83]. Industrial [Nol84]. Industry [Wes76]. Inequality [Wea80]. Inertia [Sil73]. Infallibility [Sch07a, Sch07b]. Infecting [Ger13]. Infection [Rap69]. Inference [Len58]. Inferiority [Bak70]. Inferos [Smi67, Tur66]. Infinite [Atl52, Bar73, Dal84, Dau77, Fur81]. Inflation [Tay55]. Influence [Ano88e, Bro06, Bux84, Cri02, DeA87, Ech72, Ger15, Han67, Har59, Hay59, Het83, Hof85, Koh17, KS24, Kra40, Lea82, Les57, Low49, LS91, Man88, Man91, McL86, Pey49, Pre54, Ric79, Sav59, Sch45b, She00, Sig62, Spi76]. Influences [Eyc57, Sha88]. Information [Ano40i, Ano40j, Ano41k, Ano42i, Ano42j, Ano43f, Ano44g, Ano44h, Ano45p, Ano45q, Ano46p, Ano46q, Ano47p, Ano48n, Ano48o, Ano49e, Ano49f, Ano50d, Ano51b, Ano52g, Ano53c, Ano53d, Ano54d, Ano55d, Ano56g, Ano57l, Ano57m, Ano58k, Ano58l, Ano59l, Ano59m, Ano59n, Ano60d, Ano60e, Ano61f, Ano61g, Ano62j, Ano62k, Ano62l, Ano62m, Ano63k, Ano63l, Ano63m, Ano63n, Ano64k, Ano64l, Ano64m, Ano64n, Ano65j, Ano65k, Ano65l, Ano65m, Ano66k, Ano66l, Ano66m, Ano66n, Ano67l, Ano67m, Ano67n, Ano67o, Ano68l, Ano68m, Ano68n, Ano68o, Ano69m, Ano69n, Ano69o, Ano69p, Ano70l, Ano70m, Ano70n, Ano70o, Ano71j, Ano71k, Ano71l, Ano71m, Ano72i, Ano72j, Ano72k, Ano72l, Ano73j, Ano73k, Ano73l, Ano73m, Ano74i, Ano74j, Ano74k, Ano74l]. Information [Ano75l, Ano75m, Ano75n, Ano75o, Ano76j, Ano76k, Ano76l, Ano76m, Ano77j, Ano77k, Ano77l, Ano77m, Ano78i, Ano79g, Ano80n, Ano81k, Ano82i, Ano83k, Ano84k, Ano85j, Ano86o, Ano87l, Ano88o, Ano89l, Ano90n, Ano91m, Ano92n, Ano93m, Ano94m, Ano95l, Bla03, Dug15, Ogi03, Ros03]. Ingenuity [Gre81]. Inhabited [Hef78]. Inhuman [Bia24]. Innate [Arm69, Bid76, Blo79, Gre72a, Rog79, Smi61]. Innocence [McA10].

Innocent [Har66]. Innovation [Cam96, Con96, Dur43, For05]. Inquiries [Gre15]. Inquisition [Kei04, Mot01]. Insanity [Lan02, OR13]. Insights [Wet05]. Inspiration [Tig70]. Instauratio [Whi89]. Instauration [Ser17]. Instinct [Gre97]. Institute [Sta87]. Institutes [Mue43]. Institutional [Rob20a]. Institutions [Eki50, KM94]. Instruction [CM05]. Instrument [Bat00b]. Instrumentalism [Wol49]. Insurrection [Rib52]. Integra [Ano64j, Kel64]. Integral [Kan45]. Intellect [Pag96]. Intellectual [All85, Bar16, Byr23, Car19, Cas45, Cen89, Dav10, Dig06, FR07, Feu79, Gra81a, Hed98, HTW21, Hig54, Hil72, Hoe15, Kel87, Kel05b, Kib46, Kri73, Lev05, Moy07, Mul23, Mun79, Öza14, Pag88, Pal09, Rut14, Sch73b, Sch05b, Sch22a, Sel13, Sho16, Smi24a, Smi24b, Ste85b, Val22, Wal19, Nis21, Kri70]. Intellectualism [Avr66, Hou52]. Intellectuals [Avi67, Cai07, NS16]. Intelligible [Ran69]. Intensification [Pan21]. Intentionality [Sch07a]. Interactionism [Fef90]. Interest [Lam74, Mun80b, Gun68]. Interests [Kib46, Pau81]. Intérieurs [Hal57]. Interior [Wil20]. Interloper [Het81]. Internal [Cla98, Hum72, Tou13]. International [Bou94, De 15, Deu45, Flo46, Mil61, Mor20, Per18, Sch05b, Sch46b, Ano60c]. Internationalisms [Dun21b]. Interpretation [Ada80, Bar58a, Bis52, Ces05, Ede52, Fer51, Fer58, Fis75, Gol52, Gre67, Haw94, Lev67, MB81, Mar09, McK81a, McM66, Pfa90, Sim47, Ste18, Tan99, Wei50, Wil60, Bea71]. Interpretations [Fen86, Gar13, Maz72, Mor78b, Öza14, Par88a, Wei72, Wol61]. Interpreters [Mon97]. Interpreting [Col09, Cro85, Mar09, Naj99]. Interrelations [Swi75]. Intertextuality [Joh04]. Intertwining [CH15]. Interventionist [Duc06]. Interview [FR07, Moy07]. Interwar [Chr23, Dav10, Per18]. Intimacy [Nie86]. Intimate [Bro18]. Intradeical [Wol61]. Intrinsic [Ces04]. Introducing [Fis70]. Introduction [Ano14e, Bev17, Bur12, Cam22b, Dom83, Gor19a, Gou06a, GSJ18, HG21, Lat57, Moy07, Nad13b, OS22, Pal18b, RR06, She06b, Sol03b, Spe13, Sta09a, Wol16a, Wol16b, Zam98]. Introductory [W.51]. Intuition [Gra81a, Hed98, Lan64]. Invasion [Bok17]. Invective [Can73]. Inveniendi [Van93]. Invent [Gri58]. Invented [Qui92]. Inventing [Sch03, Sch47]. Invention [Chr23, Con14, Hau98, Kel12]. Inventions [Wie40]. Invisible [Har11, Mac71b]. Iran [Mas00]. Ireland [Ber55, Con86, Lee95, Loc04, Orr14]. Irish [Cam16]. Ironic [Tay01, Wes58]. Irony [Ano81f, Lac10, Nel87, Rei81, Wis68, Nag68]. Irrelevant [Hil72]. Irresistible [Fie76]. Irreversibility [Bru76]. Isaac [Aus70, Cal12, Het83, Mac92, Mal97, Pfi97, Pow70]. Isaiah [Nor08]. Ishmael [Mil56]. Islam [Shu22, Sta89, Wal98, vdL17]. Islamic [Goc99, Rob23]. Island [Mer66]. Isolationism [Cra40]. Israel [Ger25, Nad19, Sta16, Wol16a]. Israeli [Sap79]. Issue [Cri02, Pal18a]. Italia [Dur43]. Italian [Bar62, Tri56, Bla91a, Coc62, Fer58, Jac75, Kal95, Ros24, Tri87]. Italy [Cer05, Dur43, Gre65, Gul62, Jun53, Jur01, Ren76, Sar14, Tou13, Urq14, Zan79]. Item [Mos51]. Itinerary [Sei90]. Itzkin [Tay82]. ius [Wau20]. IV [Lan95].

J [Coa50b, Fer68, Gre15, Rif53]. J. [All74, Bee20, Bet09, Bur59, Fre73, Hai50, Ina20, Isr16a, Kat44, Kel97, Men56, Mer48, Ran66a, Rin61, Rut14, Shi45, Ski25, Smi80a, Sul83, Wol16b]. J.-J. [Rin61]. Jabbar [Lea80]. Jackson [Ful83, Hut78]. Jacob [Sav44, Gil86, Gro99, Hin92, How99, Sty15]. Jacobean [Oak68, Sal89]. Jacobin [Pla10]. Jacobins [Sch03]. Jacobitism [Ber86b]. Jacques [Car08, Elu92, Gra44, Her05, Sim04, Wei16, Zah12]. Jakobson [Bri23]. James [Sav44, Wie43, Ano88e, Ber62, Ber80, Bis74, Bra79, Bux84, Cap54, Cop51, Cot81, DeA87, Dew41, Dot86, Fei70, Fun94, Gue95, Jez71, Kau63, Ket58, Kra40, Kre00, Lev47a, Lev47b, Low49, Phi71, Rya09, Sch93, Sho52, Thr11, Tin54, Wal07, Wei52, Wie63, Wil56]. Jamieson [Con14]. Jandun [Tho58]. Jane [Cur61, Sch93]. Jansenism [Len77, Sch99]. Jansenist [Ham58, O’B85]. Japan [God08, How01, Nol84, Sho10, EST66]. Japanese [San59, Fac21, Per18, Sel13]. Jaspers [Deg23]. Jaures [Nol61]. Jay [Col82]. Jean [Ger25, Bis15, Doe22, Fid94, Hey79, Kla93, Lin91, McR55, Nic19, Nol61, Shk81, Sim04, Smi64, Tho58]. Jean-Jacques [Sim04]. Jean-Marie [Fid94]. Jean-Paul [Ger25, Doe22]. Jean-Robert [Hey79]. Jeans [Wie43]. Jefferson [Coo73, Pol67, San60, Val04, Whi76b]. Jeffersonian [Hof41, Pos86]. Jena [Wer13]. Jeremy [Coo81, Sch07b]. Jerusalem [McQ14]. Jesuit [GS23, Ren76, Row56a]. Jesus [DeV08, Lin10]. jettatura [dC11]. Jevons [Pau79]. Jewish [Fac75, Glu13, Ano75g, Avi64, Ber14, Bla93, Cam22a, Efr97, Koh17, Nad01, Nad09, Rat86, Sin00, Ste18, Yof04]. Jewishness [Dia82]. Jews [Ben23, DM88, Elu92, Lan66, Shy14, vG20, La 93]. Jhering [Ang00]. JHI [Kel00]. Joachim [Pri13a]. Joanna [Isr16b]. Joel [Bla49]. Johann [Ger25, Ale66, Bel15, Gil83, Mor25, Pri13a, ZMM10]. Johannes [Bla13, Goo18, Par04, Ver22]. Johannine [Lev97]. John [Boa63, Orr14, Ros68, Sch50, Wie48, Ada92, Ano70g, Ano88e, Ano90m, Arm65, Arn94, Bat74, Bev17, Bla22, Bor71, Bro06, Bux84, Cam16, Can61, Car62, Cla15, Cla82, Col82, CH15, Con14, Cyw69, DeA87, Del07, Dew59, Duh53, Dun21a, Dyk52, Dyk59, Dyk61, Dyk62, Eig09, Feu49, Feu58, Feu59, Fis75, Ful83, Gal17, Gre16, Hai46, Haw01, Hes18, Hun72a, Joh58b, Jon92a, Kel96, KNS^{+}81, Lea82, Les57, Loc07, Low61, Mic03, Mil61, Nic40, Nol84, Nor04, Nor01, O’N07, Osl70, Pla10, Ram01, Ran65, Rap97, Ras07, Rat92, Rog06, Sav59, Sch43a, Sch83, Sha02, Smi62a, Spr76, Str55, Tay01, Thi92, Tra64, Wal07, Wal84a, Wal11, Wil23, Yol51]. Johns [Dyk61]. Johnson [Ebe75, Fis58, For82, Kor74, Lyn00, McL86, Mid60, Pat86, Sie75, Sil93, Ste78c]. Jonas [Laz03]. Jonathan [And64, Can73, Gue95, Man94, Rea72, Sta16, Sta59, Wol16a, Yar86]. Jones [Smi65b, Bou90, Cha24, Wie80b]. José [Wol76b, Ver22]. Joseph [Ano51a, Arm07, BB51, Can84, Fau78, Gar96, Gar73, Hat75, Kel08, New01, Pop53a, Pop54, Sch64, Ste78b]. Journal [Bro72, Sta09a]. Journalism [Doo90]. Journals [Ano42f]. Journey [BJ19]. Joyce [Tin54]. Jr. [Boa63, KNS^{+}81]. Juan [Nau15]. Judaeus [Jon13]. Judaism [Koh17, Sut03, Tay17]. Judging [Ger13]. Judgment [Eck45, Lun04, Nak22, Pek99]. Judgments [Loe46]. Judicial [Got44]. Julia [Oxt01]. Julian [Phi07]. Juliet [Kei57]. Julius [Mar04, Pri17b, Sac72]. Jung [Gra81b, Gun82]. Junior [Mor16]. Jupiter [Mac71b, Wen24]. Jure [Lee11]. Jurisprudence [Koc57, Rum65]. Juristic [Tie63]. Jury [Bro24]. Just [Fer73, Gal17, Sch22b, Tur66]. Justi [Men56]. Justice [Arc67, Bla13, Bla19, Bro53, Chr46, Cot79, Ha20, Lac47, McC58, Mul68b, Pie10, Rum65, Sch46b, Wil05, Win81]. Justifiability [Sho79]. Justification [Per66, Wal84a]. Justifications [De 03]. Justinian [For68]. Justus [Mos98].

K. [Blo96]. Kabbalah [Ste18, Ter18]. Kabbalist [Cou75]. Kaleidoscope [Kim24a]. Kames [Tsu61]. Kant [Wel84, Ale66, Axi71, Bec67, Bec71, Bec72, Ben82b, Bre01a, Car84, Deg12, Des67, Fal90, Fol17, Gre23a, Hab90, Han59, Het73, Hum72, Ins11, Kel68, Kra18, Kri65, Mat55, McQ14, Pac89, Poo99, Ree52, Rei56, Ruk18, Sch89a, Sei83, Sha70, Tie01a, Tie01b, Tow87, Van95, Van01, Ver69, Wal81, Wol60, vG20]. Kantian [Fen86]. Kantianism [MB81]. Kantorowicz [Her21]. Kapital [Ber78a]. Karel [Mie73]. Karl [Ran41, Wel84, Col74, Cre83, Feu79, Gre63, Gui08, Hac98, Har03, Imm09, Joh67b, Smi41, Sty15, Wen24]. Karlheinz [Cor98]. Kathleen [Pri58]. Kaufmann [Cri66]. Keeping [Sta09b]. Keill [Thi92]. Kelley [Car23]. Kennington [Cat73a]. Kepler [Ano85e, CM05, GCM10, Gra92, Nic40, Wil74a, Ros85a]. Keplerian [Han61, Sak09]. Keywords [Rod88]. Khaldûn [Gat67]. Khan [Che59, Hed98]. Kicking [Pat86]. Kierkegaard [Nag68, Gri65, Kja02, Poj83, Sar83]. Killing [Har66, Lac99]. Kind [Gui23]. Kinde [Cog07]. Kindi [Sta89]. King [Oak68, Ben57, Oak98a, Sol58, Ste73]. Kingdoms [MS15]. Kings [Kil11]. Kipling [Var53]. Kircher [Bra82]. Kirchliche [Ram19]. Kloyz [Wei16]. Knickerbocker [Hed59]. Knopf [Bar42]. Know [Pet01]. Knowing [Bul05, Lo17, Mar23, Nau15]. Knowledge [Bis52, De 41, Har02b, Hat78, Hor82, Jan08, McK81b, Min70, Osl70, Osl83, Pir25, Pop93, Ram14, Ros71b, She57, Smi17, Str55, Tou13, Vic92, Wil22a, Wil22b, vI11, Wie51]. Knox [Ham58]. Kohn [Wol76a]. Kołakowski [Kos24]. Koselleck [RR06]. Kratos [Fon00]. Krieger [Sch91b]. Krise [RR06]. Kristeller [Hor78, Ano00e, Mah99]. Kropotkin [Ada16]. Kubla [Hed98]. Kuhn [GR86]. Kurt [RGE45]. Kyndenesse [Gal94].

L [And02, Haa48, Ned24]. L. [Gri74, Wei96]. L’Être [Aro87]. Labor [Car19, Cla23, Con79, McN73]. Labriola [Jac75]. Lack [Bau46]. Lactantius [Fis82]. Lacuna [Kae96]. Ladies [Bro04b]. Lafayette [Mae78]. Laisser [Jen44]. Laisser-Faire [Jen44]. Laissez [Alb55, Coa50a, Fra78, Ros63, Kit66]. Laissez-Faire [Alb55, Fra78, Ros63]. Lake [Tuv66]. Lamarckian [Ste78b]. Lamarckianism [Sto62]. Lamennais [Spi76, Spi80]. Land [Cav16, Par73, SF93, Kas06]. Landscape [Gai00, Neu47, Pac79, Sha88]. Landscapes [MR20]. Lane [Har85a]. Langevin [BV88]. Language [Ben82a, Ber74b, Car19, Cor80, Eng63, For74, Gel52, Ger25, Giv62, Gra88, Kno65, Mon89, Moo76, Nau11, Nau15, Pax93, Qui92, Rau88, Ree83, Rob06, Sch04, Sei69, Ser01, Sin89, Smi19, Spr80, Tan99, Vel19, Was73, dG80, Isr16a]. Languages [Kno63, Lan77, dS23]. Lankester [Feu79]. Laozegetics [Tad22]. Lapo [Wat92]. Lapouge [Hec00]. Lasaulx [Ton64]. Lassalle [Fis62b]. Last [Feu79, Hic10, Lal54, ZMM10]. Lastarria [Wol76b]. Late [Ash78, Bla19, Cel01, Cou83, Cou87, Gra62, Mar90, McN66, Obe87, Sel74, Str21, Str76a, Sum17, Cal97, Fle63, For74, Gal94, Gar84, God08, Han17, Jac76b, Kle99, Lan76, MH04, O’B85, Phi70, Pop05]. Late-Medieval [Gal94]. Later [Man91, Ros77, Van65, Jon73]. Latest [Bok17]. Latin [Gra75, Ard63, BJ19, Dea74, Gra88, Kla71, Pal06, Pal09, Pal18b, Que63, Spi76, Zea59]. Latin-American [Que63]. Laughter [Hey82]. Lavater [Gra61]. Lavoisier [Tou57]. Lavoisiers [Vid95]. Law [Bat24, Bis15, Bur87, Chr46, Coh78, Coh87, Cun67, Eas67, Gra63a, Gre97, Hor74, Jac12, Jac90, Jon13, Kri60, Kri65, Lac47, Lee09, Lee11, Lev87b, Lev78, MP04, Min68, Min70, Mon97, Mor22b, Nau02, Noo72, Oak98a, Orr14, Pal72, Par62, Rub86, Rus85, Sac18, Sch91a, Sel63, Sho99, Sin63, Sku64, Sme20, Smi64, Sos73, Spi55b, Tho07, Tie01a, Tie01b, Wil70b, Woo88, Wor49, Woo24]. Lawrence [Mud68]. Laws [Cra44, Elu02, Har95, Het83, KS24, Mae75, Mae78, Mae83, Nee51a, Nee51b, Osl85, Per66, Ver82, Wil74a, Shu63]. Lazarus [Kal87]. Leadership [Bal42]. Leading [Cor75a]. League [Pol43, Str72]. Leaning [Moo51a, Moo51b]. Learn [Jum12]. Learned [Bro72, Pir25]. Learning [Alt07, Hat68b]. Leaves [Bet09]. Leavis [Woo19]. LeBrun [Ros84]. LeClerc [Kla93]. Lecons [Val24]. Leconte [Ste78b]. Lecture [Bou90, Wag54, Col16, Fog12, Ned13, Str16]. Lecturer [Joh42]. Lectures [BS21, Duh53, Fro07, GJ69, Bot17, Kiv70, Phi53]. Left [Esk04, Hel73, Kla71, Lut70, McL76b, O’S14, Pau84]. Legacy [Deg23, Itz78, Jer13, Mar19, Spe02, Swe93, Tay82]. Legal [Bat24, Kli24, Loc04, Rum65, Wal59, Wal11, Wie48, Coh49]. Legem [KS24]. Legend [Mue58]. Legendary [Bau46, Kah47]. Legislation [Mor22a]. Legitimacy [Bar83, Gor19a, Gor19b, Ha20, Laz03, Sty19]. Legitimate [Sho79]. Legitimism [Hol82]. Leibniz [Aga69, Ano67g, Bal60, Bar81, Bro04b, Car19, Car07, Car72, CR81b, Cor75b, Fel04, Fis72, Gri65, Gro96, Gro11, Har46, Jol78, Lac45, Lær14, Lat94, Loe46, Loe47, Loe55, Loe61b, McR57, Mel99, Mul68b, Nad94, Nan21, Nas46, Nor46, Per02, Per69, Ril94, Ril99, Sch66, Sch46b, Sch47, Sco97, She46, Spi52, Wes99, Wie40, Wil67, Tur71]. Leisure [Kin00]. L’Encyclopédie [Var73]. Lenin [Kru55]. Leninism [Mik70]. Lennon [Mor04a]. Leo [Dau77, Alt07, Hol17, Kee12, San20, Tay17]. Leon [Ber87, Sch23b]. Leonard [Sch91b]. Leonardo [Han95, Ian00, Ian08, Qui10, Ran53, Sch56a]. Leonhard [Wel84]. Leonhart [Kus97]. Leonor [Mot01]. Leopold [Bit17]. Leroux [And02, Bak76]. Leroy [Gun62]. Lèse [Kel81]. Lèse-Majesté [Kel81]. Lèse-Nation [Kel81]. L’Esprit [Lyn77]. Lessons [Rei17]. Lester [Nel72]. l’Estoile [Mon97]. Leszek [Kos24]. Letelier [Wol76b]. Letter [CF79, Goo90a, Pol67, Shi45, Wei42, Str22, Lom02]. Letters [Dew59, Kau63, Lun21, Nad62, Nic68, O’F92, Pol64, Ski25, Car19, Elu92, Esk04, Hia09, May04, Pir25, Twi20, Kim24b]. Lettres [O’C84, You75]. Leveller [McD05]. Levels [Ede46]. Lévi [Gra81b]. Lévi-Strauss [Gra81b]. Leviathan [Hoe15, Mal15, Mis77, MS15, Nau02, Par15, Ste67a, Vie15, Hoe15]. Levin [Eng95, Mer89]. Levine [Kel08]. Lewes [Kam52]. Lewis [Bla52, Boc51, Cas92, Guy02, Sch15]. Lexicography [Con14]. Lexicon [Str23]. l’Expression [Ros84]. Liang [Ang00, Lev50]. Liberal [Bar45, Bec69, Fal54, Fat05, Flo46, Fre73, Gab56, Hol82, Wes69, Wol76a]. Liberalism [Bla73, BB51, Boe83, Bro99, Dro19, Eyc57, Fol58, Gar12, Guy20, Kat89, Kee12, Kel65, Met14, Mue45, Por72, Sch43a, Stu89, Sul83, Wei96, Wes15, Coa50b]. Liberalisms [SJB18]. Liberality [Liz22]. Liberals [Kap56, Mor90, Moy23]. Liberating [Lac99]. Libertarian [You75]. Libertas [Sut53, Voz25]. liberties [Woo24]. Liberty [Bou00, Col71, Coo73, CW13, Ehr52, Ela15, Gre18, How01, Ina20, Lee09, Lee11, Mor09, Pap79, Pes14, Pod10, Pos86, Ros68, Ski12, Sta77, Ten69, Tri56, dBJN^{+}20, Sch41, Eck65]. Liberum [Vie03]. Libido [Dom88]. Libraries [Bur64, Kib46]. Library [Ano75f, Bro97, Thi79]. Libri [Cor18]. Licence [Tan77]. Licencing [Tan77]. Liceti [Del17]. l’Idée [Mor40]. L’Idéologie [MR20]. Lie [Gun68]. Lieberman [Ste18]. Life [Dav60b, Dic80, Dyk52, Hef81, Hel19, Ian00, Lin10, Mad54, Mar90, Par68a, Ran67, Roe85, Ski25, Van65, Wei02, vB17, Ano88f, Joh88]. Lifts [Wil20]. Light [Bar47, Bev17, Gra68, Nad62, Rap69]. Lighthouse [Woo94]. Ligon [Smi17]. Likeness [Wei72]. Liminal [Tou13]. Limits [Bro99, Elu02, Hat78, Hig95, Kuk00, SF93]. Line [Wis49]. Lines [Hel21b]. Linguistic [Bri23, Byn84, CH95, Doe22, Hen69, Lif12, Pag88, Sko65]. Linguistics [Gui08, Mie69]. Link [Mor93, Pic89]. Links [Han91]. Linnæus [Lar67, Mar10]. Lippmann [Jac12]. Lipsius [Cor75a, Mos98]. Lisa [Boa40]. List [Ger72]. Literacy [Bur81, Hal92]. Literaria [Cor80, Mil99b, Roy07]. Literary [Atk74, Bou03, Bri79, Cas88, Coh40, Doo90, Fra42, Gar70, Hau11, Jon67, Lan80, Lev72a, Moy75, Phi89, Pri95, Sir18, Tan99, Tho75, Van15, War11, Wil41, Wol68]. Literature [Bea83, Coc62, Cre83, Dor40, Dup74, Fer55, Fir71, Hef73, Hig61, Jon40, Neh77, Nor77, Sch52, She77, She57, Shu68, Ste51, Sti09, Str76a, Tig70, Tri56, Vir86, Boa70, Neg79, Sap79]. litterarum [Oos13]. Little [Shu76]. Lives [Nad22]. Living [Plo53]. Livre [Wil66a]. Livres [Nic19]. Livy [Kos04]. Local [Lat24, NL20]. Localizing [BJ19]. Locating [OS22]. Location [Jez71]. Locke [All66, AJ81, Arm65, Arm69, Arn94, Bat74, Bid76, Bro06, Bro99, Cla98, Dav63, Dus81, Eng80, Far87, Fis75, Giv62, Gla90, Gre72a, HM79, Hel81, Hun72a, Joh58b, Jol78, Kro84, Lau67, Leb86, Mal09, McL86, MM90, Moo76, Nor01, Ode65, Oli74, Osl70, Par69a, Pos86, Rap81, RMG81, Rog66, Rog78, Rog79, Rom61, Rom64, Row56b, Sch80a, Sch90, Sel63, SF93, Smi62a, Smi62b, Sny86, Tay01, Ten82, Vog93, Wal84a, Wal93, Wal04, Wil67, Win81, Yol51, Yol55, YY84, Yol01, Yos51, Sch89b]. Lockean [Tie11, Pea14, San60, Sch90, Sea11]. Lockeanism [Wat71]. Locked [Wat71]. Lodwick [Poo05]. Logic [Bel15, Fis62a, Har59, Ina20, Jer13, JN12, Ken90, Mar12, O’N12, Puc05, RS64, Ron12, Smi17, Tho12, Tre67, Tuv42, Wal07, Was73, Yol55]. Logical [Bir62, Puc05, Res63, Wie46b, Zan79]. Logician [Dun21a, Res60]. Logico [Roz02, War11]. Logico-Philosophicus [Roz02, War11]. Logik [Lev47b]. Logoi [Bos89]. Logos [Fon00]. Lois [Lyn77]. Lollards [Jon73]. London [Joh42, Mil01, Pec96, Pop05, Syp41, Wei75]. longer [Her05]. Longevity [Ege66]. Look [Bur57a, Cal62]. Looking [Yol79a, Zag03]. Looking-Glass [Yol79a]. Loosened [Ros77]. Lord [Dea72, Dus81, Gre91b, Kli06, Lan76, Lan02, Nur61, Pea94, Ton59, Tsu61]. Lorenz [Men61]. Lorenzo [Bul05, Cam96, Cel05, Con96, Mon89, Tri96]. Lost [Kel12, Sou89, Cha63, Den99, Lin91, McD05, Par68b, Ulr71]. Lot [Som98]. Lotze [Kra40]. Louis [Bak91, Bat00a, Gun62, Keo74, KCR46, Lou56, Nar05]. Louis-Claude [Bat00a]. Louisiana [Coa61]. Love [DJ69, Ell03, Fon99, Lev79, Maz57, Moo79, Rut14, Sam43, Sca97, Wil97]. Lovejovian [Lac10]. Lovejoy [Col16, Fog12, Ned13, Str16, Ano48a, Ano87e, Bec72, Boa48a, Cat10, Dig06, Duf80, Ede63, GB87, GM80, Mah87, Man48, Mon48, Nic48, Oak87, Wie73, Wil80, Wil87, Spe48]. Lovely [Dus81]. Lowell [Hef81, Het81]. Löwith [Gor19b, Sty15]. Lubbock [Tha83]. Lubin [Val22]. Luce [McK87]. Lucretian [Sny11]. Lucretius [Mer50a, Mul58, Pal12]. Ludwig [Dod68, Cal22, OS22, Rah22, Roz02, Ste78a, Web22]. Luis [Nau15]. Lukács [Nak22, Shu73]. Lunatic [OR13]. Luther [Ste73, Osb02, Por81, Sho79]. Lutheran [Mor25]. Lutheranism [Gus50]. Luxury [Jen07a]. Luzzatto [Scl20]. Lyceum [Wag54]. Lycophron [Mul79]. Lying [Ben82b]. Lynn [Joh59a]. Lyotard [Van01]. Lyres [Hil11]. Lyric [Bee20]. Lysenko [Jor59].

M [Boa58, Kel08, Mor04a, Mur54, Wie56]. M. [Par24]. Mab [Kei57]. Mabillon [Hia09]. Macaulay [Phi89]. Mach [Sea68]. Machiavelli [Gil53, Bar60b, Bea82, Ble58, Col71, Col99, Dym21, Fle66, Fon99, Gee76, Gee99, Har89, Jur02, Kap23, Liz22, McK82a, Naj99, Ned99, Pre79, Sta14, Wal70, Woo67, dG84]. Machiavellian [Dro19, Ano83f, Bak83, Maz60]. Machiavellism [Hea88]. Machina [Bat00b]. Machine [Mor78a, Shu68, Vie49]. MacIntyre [Woo19]. Macrocosm [Nor77]. Mad [Guc00, Sti09]. Madagascar [Wri43]. Made [Liv08]. Madison [Bra79, Bur11, Ket58]. Madness [Nad97]. Magazine [Blo84]. Magdeburg [Lyo03]. Magic [Dav13, Joh59a, Kas06, Nau57]. Magical [Cop91]. Magisterial [Whi86]. Magnetism [Kap74]. Magnitude [Eva55]. Magnus [Cun67]. Magog [Poc87]. Maharal [Wei16]. Mahomet [vdL17]. Maimon [Atl52, Bre01a]. Maimonides [Bla93, Elu02, Sac76]. Main [MB81]. Maine [Boc76, Smi63]. Maistre [Arm07, Gar96, Spe02, Wat71]. Majesté [Kel81].

Major [Ros43a]. Majority [Lan66]. Makers [Tuv50]. Making [Åsb21, Ant21, Coo14, Gal94, Glu13, Hun72a, Jou23, LS23, Lun21, Nak69]. Malebranche [Nad94, Acw77, Bar92, Fri54, Ham08, Han67, Joh58b, Mun80a, Nad91]. Malthus [Ano66f, Bow76, LeM79, Lev66, Nic90, San82, Sim55, Vor69]. Malthusian [Ros70]. Malthusianism [Aus52, Keg58]. Man [Bat48, Boc51, Che63, Col62, For42, Gel52, Hab95, Hor74, Ket58, Kli81, Kri44, Liv08, Mor93, Mor78a, Mud68, Pri13a, Pri54, Sei74, Ter18, Val04, Wri80, Yol01, You01, Pat58, Scl20, Ste60, Kat89, Ran41]. Mancini [Fri12]. Mander [Mie73]. Mandeville [Jac76a, Alt08, Coo99, Dan86, Gol88, Hun95, Jac78, Ros63]. Mania [Dav69, Kap74]. Mankind [Ada91, Gol88, Pea52, Ker24]. Manley [Spr80]. Manners [Kon97]. Mannheim [Ran41]. Manuscript [AA85, Bos70, Hab93, Kiv70, Min68, Rap69, Wil66a]. Manuscripts [Hab90]. Many [Hui17, Man08, Mun51, Nad22, Ogi03, SJB18]. Mao [Tod74]. Mapping [Del07, For23, May04, Pet88, Pri74]. Maps [For23]. Marcia [Ned24]. Marcus [Afr61, Str22]. Mare [Vie03]. Margaret [Boy06, Car78, Kea64]. Marginal [Dom84, Ros68, Yol51]. Margins [Spe98]. Marie [Doa17, Fid94]. Marina [Wes15]. Maritain [Car08, Cou44, Gra44]. Maritime [LS17]. Marjorie [Pri63, Ano47k, Str22, Tay81]. Mark [Ald44, Tho12]. Market [Web22]. Marmora [Wal11]. Marquard [Sty19]. Marriage [Ach20, Cla92, Joh71a, Nad82, Osm73, Tun14]. Mars [Gou45]. Marsh [Dew41]. Marsiglio [Ned05]. Marsilio [Bur44, Jac85, Kri40, Sny11]. Marsilius [Lee09, Sig62, Tie91]. Martin [Bat00a, Goo90a, Ste73]. Martínez [Wes15]. Martyr [Zhu21]. Martyrdom [Ber57]. Marvell [Bar60b, Maz60, Rob62, Sch60, Ste67b]. Marvelous [Dor40]. Marx [CF79, Dup74, Ada74, Avi64, Avi67, Ber78a, Col74, Cro71, Eat80, Fay78, Fen86, Feu79, Gre63, Ham70, Ham72, Joh67b, Kel84, Ken79, Kri53, Mos85, Ols68, Pri69, Sim81, Smi41, Ste18]. Marxian [Dan60, Feu49]. Marxism [Cas83, Cro71, Hoo68, Jac75, Jor59, Koh64a, Mar68, Mik70, Sko65, Dup74]. Marxist [Bev93, Cho15, Ell67, Hoo73, Kov76, Pau81, Sch73a]. Mary [BB89, Con99, Ell03, Jan78, O’N07, Sir18, Tay01]. Masaryk [Koh64a]. Masculinity [Hen95]. Masham [Bro06]. Mason [Wis49]. Mass [Mer11]. Mass-circulation [Mer11]. Masses [Gre74]. Master [Cum78]. Masters [Mai21, Kel95]. Materia [Kin56]. Material [Mad54]. Materialism [Dom83, Lev87b, Ram19, Roe85, Var50, Wat92]. Materialist [Ste18]. Materialists [Gro96]. Mathematical [Frä75, Gew41, Gla01, Joh42, Str51b]. Mathematician [Pet55]. Mathematics [Boc65, Byr06, Cas88, Dau77, Gar87, Gos60, Gou06b, Har59, Ols71, Oos13, Pag96, Pop06, Pyc87, Rat92, Vuc60, Whi45b, Bro91, Pop06]. Mathematization [Can77, Mal97]. Mather [Jes86]. mathesis [Wes99]. Matt [Rou67]. Matter [Adk00, Ano40d, Ano40e, Ano40f, Ano41f, Ano45g, Ano45h, Ano45i, Ano45j, Ano46h, Ano46i, Ano46j, Ano46k, Ano47g, Ano47h, Ano47i, Ano47j, Ano48f, Ano48g, Ano48h, Ano48i, Ano52e, Ano53b, Ano54c, Ano55c, Ano56f, Ano57h, Ano57g, Ano57i, Ano57j, Ano57k, Ano58g, Ano58h, Ano58i, Ano58j, Ano59h, Ano59i, Ano59j, Ano59k, Ano60b, Ano61e, Ano62f, Ano62g, Ano62h, Ano62i, Ano63g, Ano63h, Ano63i, Ano63j, Ano64f, Ano64g, Ano64h, Ano64i, Ano65f, Ano65g, Ano65h, Ano65i, Ano66g, Ano66h, Ano66i, Ano66j, Ano67h, Ano67i, Ano67j, Ano67k, Ano68h, Ano68i, Ano68j, Ano68k, Ano69h, Ano69i, Ano69j, Ano69k, Ano70h, Ano70i, Ano70j, Ano70k, Ano71f, Ano71g, Ano71h, Ano71i, Ano72e, Ano72f, Ano72g, Ano72h, Ano73f, Ano73g, Ano73h]. Matter [Ano73i, Ano74e, Ano74f, Ano74g, Ano74h, Ano75h, Ano75i, Ano75j, Ano75k, Ano76e, Ano76g, Ano76f, Ano76h, Ano76i, Ano77f, Ano77g, Ano77h, Ano77i, Ano78e, Ano78f, Ano78g, Ano78h, Ano79c, Ano79d, Ano79e, Ano79f, Ano80j, Ano80k, Ano80l, Ano80m, Ano81g, Ano81h, Ano81i, Ano81j, Ano82e, Ano82f, Ano82g, Ano82h, Ano83g, Ano83h, Ano83i, Ano83j, Ano84g, Ano84h, Ano84i, Ano84j, Ano85f, Ano85g, Ano85h, Ano85i, Ano86h, Ano86i, Ano86j, Ano86k, Ano87f, Ano87g, Ano87h, Ano87i, Ano88g, Ano88h, Ano88i, Ano88j, Ano89e, Ano89f, Ano89g, Ano89h, Ano90e, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano91h, Ano92f, Ano92g, Ano92h, Ano92i, Ano93e, Ano93f, Ano93g, Ano93h, Ano94e, Ano94f, Ano94g, Ano94h, Ano95e, Ano95f]. Matter [Ano95g, Ano95h, De 06, Gay63, Gun71, Hun52, Joh67a, Lev03, Sch64, Sol58, Tay01, YY84]. Maurice [Eng08, Her05]. Maurras [Wil68]. Max [Ste74, Bur11, Cal22, Der10, KE23, Mag16, Ste78a, Ste85a, Zha25]. Maxim [Mac71a, Gun68]. Maxime [Hic10]. Maxims [Hun95]. Maxwell [Kar69, Sta08a]. may [Str22]. Mazzeo [War67]. McKeon [Cal86]. Mckim [Wen82]. Mead [Fef90, Kar84]. Meadow [Wil20]. Meaning [Bat00b, Bow75a, Cra92, Ede46, Igg95, Joh58a, Kam70, Kra97, Lov41, Mos73, Mul94, Nau11, Ned96, Rac95, Ran49, Ros16a, Von84]. Meanings [Bit17, Giu85, Joh67a, Kem00, Lie98, Luk71, Ran64, Wil41]. means [McC20]. Measure [Sku64]. Measurer [Sha02]. Mechanical [BH59, Kub67, Sar85, Wri80]. Mechanics [Han67, Odo66, Tho07, Wal71, Zil40]. Media [McC20]. Mediaeval [Che59]. Medical [Dah15, LaC70, Mor25, Rou67, Sir91, Tou13]. Medici [Ian08]. Medicine [Bro87, Fri12, Gim08, Koc50, Mar23, Sar08, Sir03, Sir04, Tho51, Wen24, Zan87]. Medicines [Sch14b]. Medico [Rom64]. Medico-Philosophical [Rom64]. Medieval [Ash78, Bar98, Bej97, Bla93, Chr45, Col16, Cou83, Cou87, Cum10, Dal77, Gra62, Gut41, Jon73, Kae96, KS24, Lee95, Liz22, MP04, Mar90, Mol78, Ned88, Ned96, Ned05, Obe87, Ong53, Rei81, Rob60, Ros98, Sap79, Str76a, Tur66, Zal85, Ano81f, Egg99, Gal94, Koc99, Koh17, Kuk00, Not16, Yin22]. Medieval/Modern [Ned05]. Medievalist [Mar20]. Medievalists [Coc58]. Méditatifs [Hal57]. Meditation [Jor85]. Meditations [Rub90]. Mediterranean [AU19]. Meeting [Cas48, Jac68]. Meets [Wen24]. Mein [Bui02]. Meinecke [Poi70, Ran64, Wol56]. Melancholy [Bar19, Cha81, Gar77, Mor16, Mor25, Sch04]. Melanchthon [Bat24, Met96].

Melancthon [Bre52b]. Melchioris [Goo18]. Mellin [Rid62]. Mellon [Rom64]. Melody [Sol82]. Melon [Gab98]. Melville [Tri69]. Members [For84]. Memoires [Arc67]. Memoranda [Mos48]. Memoriam [KNS^{+}81, Nor04, Pal97, Rog06, Tay81, Wie80a, Wie80b]. Memory [Ano81e, Bar97, Ede63, Gom81, Hut87, Wil22b, Rei96]. Men [Bla82, Dud57, Hil11, Nad82, Nic68]. Menasseh [Nad19]. Mendel [Gas59]. Mendelssohn [Erl02, McQ14, Sch89a]. Mendoza [Spi65]. Menschheit [All74]. Mental [Wal07]. Menzel [Kan45]. Mercantilism [Hay59, Mid60]. Mercuriale [Sir03]. Mercurius [Ant66]. Mercury [Sta67a]. Mère [And02]. Merle [Sch44]. Merleau [Cho15, Her05]. Merleau-Ponty [Cho15, Her05]. Mesmeric [Kap74]. Message [Hic10, She46]. Metacritic [Ale66]. Metamorphoses [Chr71]. Metaphor [Alb17, Bea42, Cla98, Hun53, Lew65, Min65, Mul94, Pet94, Vog93]. Metaphoric [Oti02]. Metaphorology [Ada91]. Metaphors [Ada91, Kre00, Lev47a, Pan21, SB83, Sum17, Vic02]. Metaphysical [Gan56, Lat94, Maz53, Ned13, Pea15, Sel74, Tuv42, Wat55, Wie46c]. Metaphysics [Arm65, Cah17, Gue95, Hel64, Kam52, Kuk00, Mer53, Mil99b, Nad88, Per69, Que63, Rie60, Ruk18, Sha70, Tay98, Wan05]. Metapolitics [Vie42]. Metapolitics. [Bar42]. Metempsychotic [Cor10]. Meteorology [She00]. Meteors [Pet15]. Method [Bau45, DH71, Dia82, Far87, Frä75, Gew41, Gou06b, Hat78, Hen01, Hen69, Lev67, McK47, Mil99b, Mis77, Mor16, Pea48, Ran40, Ras07, Sei87, Wel76, Wel55, Woo67, Zil41]. Methodism [Ong53]. Methodist [Rob62]. Methodology [Coe16, Gir70, Jac90, Kva77, PB73, Ras85]. Methods [Bla92, Bro04a, Gom43, Lew72, McK66, OH88, Rub91, Sch91a, Wie61]. Metonymy [Pet94]. Metric [Mae80]. Metrology [Sha02]. Mettrie [Cam70, Var50]. Mexican [Flo49, Mot01]. Mexico [Lam56, Rom60, Gra75]. Meyer [Dyn81, Ros96b]. Michael [Hor82, O’S14, Pod02, Tho12, Wel94]. Michelangelo [Duf95]. Michelet [Kim24a, Mit87, Mit96]. Michigan [Dyk62]. Microcosm [Nor77]. Microscope [Wil88, vdZ98]. Microscopic [Sil76, Yos51]. Mid [Can77, Cog07, McC79, Swa62]. Mid-Eighteenth-Century [Can77]. Mid-Nineteenth [Swa62]. Mid-Seventeenth [McC79]. Mid-Seventeenth-Century [Cog07]. Midcentury [Imm09]. Middle [Boa63, Cot94, Low61, Mer48, O’N12, Bar59b, Cad95, Chr46, Dal80, Deu45, Lac47, San44, Zuc75]. Middle-Ground [Cot94]. Militant [Pag96]. Militarism [Lau44, Ros82]. Military [Cre79]. Militia [Cre79]. Mill [Ada92, Bee20, Bur59, Car62, Cum64, Feu49, Hai46, Hai50, Ina20, Jon92a, Lea82, Lew72, Mil61, Pap64, Ran65, Ran66a, Ros68, Sch43a, Shi45, Smi80a, Spi61, Str55, Str76b, Sul83, Ten69, Whi45b]. Millenarian [Jur75, vE01]. Millenarianism [Por81, Jac76b]. Millennium [Gar73, Jac77]. Miller [Sch40, Wis68]. Million [Gri21]. Milton [Ant66, Fal88, Fis56, Fis62a, Gra63a, Haw01, Hun52, Hun60, Lev89, Lyn00, Mye08, Ogd49, Pat64, Rum85, Sie50, Wes53, Woo24]. Miltonic [Fol17]. Mime [Col62]. Mimicry [Eva65]. Mind [Ano75g, BB60, Cor10, Dac04, Del07, Gor94, Han17, Hel21a, Jen20, Jes54, Kir50, Kir94, Lan09, Mat06, Mic03, Nor75, Pea52, Rob06, Rya09, Sch40, Sta09a, Sta09b, Sti09, Str54, TC47, War71, Was72, Wis68, Zag90, Bre01a, Rol71, Fac75]. Minds [vR00]. Ministerium [Sac18]. Minoda [Per18]. Miracles [For82, Fra13, Har95, Har99, Har13, Sie75]. Mirandola [For42, Bre52a, Bre52b, Cas42a, Cas42b]. Mirarchi [Sil77]. Mirror [Ste40]. Mises [Cal22, Jen22, OS22, Rah22, Web22]. Misnomer [Tri04]. Misplaced [Pal06]. Misprisions [Lun04]. Misreading [Was89]. Missing [Mor93]. Missiology [Ver22]. Mission [BS83, GS23]. Mistake [Pea80]. Misunderstood [Les72]. Mittwochsgesellschaft [Sch89a]. Mixed [Alb17, Bly86, Bro91]. Mixing [vG20]. Mobilized [Ros98]. Mobilizing [Cha23a]. Moby [Sto51, Tri69]. Moby-Dick [Sto51, Tri69]. Modal [Bou01]. Mode [Cre83]. Model [Ha20, Het73, Kar69, Nad98, Rap16, Suł89, Wer13]. Models [Mor78a, Pet94, Rie60]. Moderate [Mar12]. Modern [Ach20, Ano69l, Ano73e, Bla22, But61, But77, Cam22a, Cam22b, Cas88, Cla11, Cor75a, Cra92, Cul88, Doo90, Doo99, Fac21, Fei44, Fel77, For89, For23, Gor19a, Gor19b, Gou06a, Gra44, Gra03, Gre72b, Gun71, Har99, Har02b, HH66, HTW21, Hol49, Jon13, Kal87, Kap08, Kas06, Kel79, Kil11, Kim21, Kri51b, Kri52, Lat24, Lee95, LS17, Lew01, Mag16, Mai97, Mar19, McK83, Mil70, Mil96, Mor53, ME21, Ned05, Nev09, O’B58, Pag96, Par44, Par01, Pop06, Ran41, Ran53, Ros03, Ros43b, Ros18, Rus40, Sch84, Sea11, She03, She06a, She06b, Smi19, Sol03b, Sty15, Sty19, Tat63b, Tie11, Twi20, VR05, Von81, Wag67, Wei19a, Wei69, Wei19b, Wer13, Wes01, Wil88, Efr97, Arr94, Bri00, Sch51, Sha02]. Moderna [Mar90, Obe87]. Moderni [Cou87]. Modernism [Erl02, Mor92]. Modernitas [Tri87]. Modernity [Atk74, Cav16, Deg23, Fro95, Glu13, Har89, Hol17, Hun95, Mar20, Sty19, Wei19a, Wil22b]. Modernizing [Vov12]. Moderns [Bar59a, Bla82, Lev91, Ski12, Vid95]. Modes [Bea71]. Modest [Wit43]. Moivre [Bel08]. Möllendorf [Gro50]. Molyneux [Dav60a, Hel81, Par69a, Tur55]. Moment [And04, Col99, Eng98, Gre16, Jur01, Kja02]. Moments [Sti98]. Momigliano [Phi96]. Mona [Boa40]. Monadology [Gal74]. Monastic [Hel21a, Mar90]. Monboddo [Lan76]. Monde [Lyn82]. Mondo [Boa56]. Monete [RH79]. Money [Par01, RH79]. Monism [Woo94]. Monistic [Dea74, Hol71]. Monopoly [Dud57]. Montague [Neu25]. Montaigne [For05, Ber04, Bon85, For09, Gre75, Gru73, Log75, Sch84, Sti98]. Montaignian [Eng98]. Montalvo [Spi76]. Montesquieu [Ans09, Car86, Car91, Gil20, Lyn77, Oak53, Oak55, You75]. Montesquieuean [Gre16]. Montgomery [Kee47]. Montmort [Bel08]. Montucla [Swe93]. Moons [Gou45]. Moore [Tae50, Wol49]. Moot [Bar58a]. Moral [Adk00, Alb55, Arm07, Bis96, BB60, Boc76, Bro06, BS21, Cam20, Chr73, Con61, Cot08b, DB77, Fra55, Gol88, Gre97, Hey17, Hip04, Joy92, Jum08, Kar84, Kra18, Mos60, Ned96, Pac89, Pri17a, Rad93, Ril99, Sch04, Smi83, Val04, Wil80]. morale [Zah12]. Moralisation [Lad62].

Moralisés [Mor77a]. Moralist [Hen77, Jon92a, Not55]. Moraliste [Col66]. Moralists [Smi50]. Morality [Car84, Kar84, McC58, Mor94, Per02, Sar85, Smi62a, Was84, Win70]. Moralizing [Van82]. Morals [Ano83f, Bak83, Cot08a, Mad57, NSR80, SRN84]. Moravian [Som98]. More-Vaughan [Bur74]. Mores [Smi78]. Mori [Len04]. Morphological [Kim24a, San20]. Morphology [Ton64]. Morris [Ano03h, Ano09g, Ano15e, Bro53, Cah17, Kin00, Lar59, Nag57, Ros57b, Ran49]. Morrow [All89]. Mortality [Gro67]. Mosaic [Elu02]. Mosaics [Dow40]. Moscow [Poo99]. Moses [Erl02, Gee99, Sai64, Scl20]. Mosque [Ger25]. Motif [Spi61]. Motifs [Zal85]. Motion [Car72, Fra03, Jay70, Mer66, Mir77a, Per66, Pes14, Sar85, Wil22b]. Motionless [Mer66]. Motivation [Bis96, Gra86]. Motive [Ell62, Hun74, Pri69]. Motives [Was89]. Mount [Tut22]. Mountain [Kir50, Kir94]. Move [Bar16]. Movement [Bed21, Bri79, Cha23a, Gul62, Her48, Kea64, Mat55, Feu59, Ram19, Rue64]. Mozart [Kiv67]. Mr. [Ces05, Mou49, Mur54, Rum65]. Mrs. [Mon74]. MS [Mos60]. Mulcaster [DeM74]. Mulgan [Gut79]. Muller [Blo96, Par04]. Multidimensional [Koh70]. Multilingual [Hol21]. Multiple [Lo17]. Multitude [Yeo03]. Mumford [Smi65b, Wie80b, Cas92]. mundi [Not16]. Muneki [Per18]. Municipal [Rob15]. Münsterite [Sta67b]. Munz [McC90]. Murray [Jen22]. Muse [Mur58, Sim16]. Music [Aug65, Bil11, But15, Chr89, Dav51, Dav95, Dra70, Fin47a, Hig86, Kat84, Low54, Pau71, Reh14, Sch52, Sim04, Sol82]. Music-Historical [Reh14]. Musical [Cum10, Gui13, Pri17b, Rin61, Sco98, Tan98]. Musick [Fin47b]. Musicology [Zon14]. Musique [Chr89]. Muskets [Hil11]. Muslim [Wun03]. Muting [Fal20]. My [Ano69l, Mon48, Ski12, Loe70]. Myops [Loe70]. Mysteries [Pri64b]. Mysteriousness [Van82]. Mystery [Bat00a, Tut22]. Mystic [Eng80]. Mystical [Beh08, Dav43, Mos61, Gar84]. Mysticism [Bat00a, Elu98, Eme84, Les72, Wym49]. Mystics [Hav59]. Myth [Ack75, Afr60, Bea91, But15, Cou04, Dan86, Deg23, Ede49, Fra78, Gra81b, Hoo73, Lew65, Mal99, Nor07, Owe75, Pas90, Sch54, Sin49, Ste75, Tag86, Van58, Whi42, Ack73]. Mythic [Gas55]. Mythology [Rie60, Wat92, Wil22b]. Myths [Bra71, Gra86].

N [Par24]. N. [Wie48]. N.S.D.A.P [Von48]. Naboth [Pog63]. Nachlass [Wil96]. Naïveté [Atk74]. Nakamura [EST66]. Naming [Ong53]. Naples [dC11]. Napoleonic [Tho94]. Naquib [Alj19]. Narrative [Ces05, Har02a, Rot90, Sir91]. Narratives [Kra97, Ter21]. Nascitur [Rin41]. Nation [Cav16, Ger25, Kel81, Mar20, Mat06, Mil74, Per18, Tol98, Tun14]. Nation-State [Tol98]. National [Bal42, Bar83, Bri40, Byn84, Cav16, Chr23, Kim24a, Mor86, Mos61, Pal40, Ran70, Von48]. Nationalism [Blu63, Ebe69, Had43, Kan45, Koh40, Koh45, Koh49, Koh51, Kra97, Lar10, Lum41, Mal40b, Mar68, Sch56b, Wil70a, Wol76a]. Nationalist [Had50].

Nationality [Gri41, Hil11, Lan02]. Nationalization [Par73]. Nations [Ade18, Hei73, Pac02, Pol43, Sho24, Str72, Whi76a]. Nativist [Fal90]. Natura [Tie63]. Naturae [Sac18]. Natural [Alt08, Bat24, Bet59, Bis15, Bla92, Bow74, BS21, Cad95, Coh78, Cop91, Cun67, Doa17, Doo09, Gar87, Gar13, Goo02, Gra63a, Gre97, Gro67, Hal02, Hes18, Het83, Hey17, Hor74, Jac90, Jon13, Joy92, Kim24a, Kri65, Lær14, Laz03, Mol97, Mor93, Mud68, Noo72, Poo05, Rei69, Rob80, Rot58, Rus88, Saa10, Sac18, Sch91a, Sel63, Sin89, Spi55b, Str57, Tae53, Tie01b, Tun14, Van82, Vor69, Wal93, Wal04, Wil19, Wor49, Lev78]. Natural-Philosophical [Gar13]. Naturalist [Gri12]. Naturalistic [Jen20]. Naturalists [Ogi03, Rob80]. Naturalization [De 15]. ’Nature [Sta57, Ada92, Arm07, Att83, Bea40, Cam22b, Cla85, Clo69, Clo71, Col62, CH15, Cur61, Dou11, Eze00, Gre54, Gre62, Gre97, Gre15, Gre80, Han17, Har95, Hei78, Hep55, Hof85, Jac78, Jez71, Jon67, Jun53, Ket58, Kro60, Kuk00, Lau67, Lun87, Mal97, McG72, McK47, McK81b, Mer08, Mil64, Mit96, Mor09, Mor22b, MF06, Ned88, Ned13, Nee51a, Nee51b, Nel72, Ogi03, Oli74, Osl79, Osl85, Par08, Plo53, Pri64b, Ram19, Rob80, Sch80a, Sch64, Sil78, Sil86, Sin63, Sol63, Sta04, Tan98, Val95, Val04, Ver82, Vic08, Wer75, Wil41, You01, Zet82, vR00, Pri17b]. Nazi [Ano75g, Fac75, Hec00, Lov67, O’F92, Por02]. Nazis [Ger25]. Nazism [Mai21]. Néant [Aro87]. Neapolitan [Cal97, Mor22a]. Near [Gas55, Fei68]. Necessitas [KS24]. Necessity [Gre15, Naj99, Ven76]. Needham [Ano51a, Roe85]. Needless [Bon56]. Negating [Les07]. Neglected [Bar88, Clo75, Nad93, Shu85]. Negotiating [Guy20]. Negro [Bak70, D’E69, Gre54]. Neighbors [Hig54]. Neill [Mou49]. Nelson [Chr64]. Neo [Bra71, Dil74, Fro95, Hos24, Joh61, Jon40, Lun83, MB81, McK82b, Pea94, Poo99, Shu22, Wen76, Dem58]. Neo-Classical [Dil74, Joh61, Jon40]. Neo-Confucianism [Hos24, Lun83]. Neo-Conservative [Bra71]. Neo-Hellenic [Dem58]. Neo-Idealist [Poo99]. Neo-Kantianism [MB81]. Neo-Paganism [Pea94]. Neo-Platonism [McK82b, Wen76]. Neo-reactionary [Shu22]. Neo-Stoicism [Fro95]. Neoclassical [Cam20]. Neohumanist [Vic02]. Neoliberal [Mul23, Web22]. Neoliberalism [Cal22, Rah22]. Neoliberals [Bed21]. Neoplatonism [Aer19, Kri58]. Neoplatonist [Ros83]. Neorealism [Neu25]. Nerves [Ste88]. Nervii [Lat24]. Nervous [Bas13]. Nervousnes [Wie56]. nescit [Bak91]. Netherlands [Lat24]. Network [Goo18]. Networks [Åsb21]. Neumann [Ras07]. Neutral [Woo94]. Neutrality [Ano45f, Wie45]. New-Right [Shu22]. Newness [Tho51]. News [Ben07]. Newspaper [Boe83]. Newton [Alb78, Aus70, BH59, Mil87, Coh87, Ell62, EG84, Eva55, Gar87, Gro11, GJ69, Gue83, Gue86, Gui13, Han60, Het83, Ins11, Jac90, Kub67, Kuh61, Mur58, Per66, Per69, Pfi97, Pow70, Rog78, Rog79, Sai88, Str51b, Str52, Wes87, Wil74a, Wor49]. Newtonian [Wes76, Asf99, Cal69, Har95, Str57, Wig09]. Newtonianism [Gre53, Ste81, Wil74b]. Nicene [Nor47]. Nicholas [D’E07, Ned05, RH79, RG70, Wei42]. Nichols [Sav44]. Nicolas [Del07]. Nicole [Keo74]. Nicolson [Ano47k, Tay81, Pri63]. Nicomachean [Elt97].

Niebuhr [Cra92]. Nietzsche [Lac99, Ano84e, AP92, Ber04, Bri40, Bro97, Bro04a, Car84, Col66, Del89, Don82, DM88, Eck45, Elm83, Eng63, For93, Gra53, Gro50, Hig86, Kau48, Les57, Löw45, LS91, McG80, Mor45, Nor02, Rie52, Rol78, Sma83, Tha83, Til45, Tou98, Vir50, Von84, Whi86, Wil96, dH45]. Nightmare [Cou75]. Nihilism [Alt07]. Nineteenth [Bri53, Cha24, God08, Jen20, Phi70, Beh08, Cat10, Dum81, Eas62, Ell57, Fle63, Goo90b, Hac83, How01, Jen07a, Leg25, Lev56, Mas00, McL76b, Oti02, Pau81, Rau88, Sim65, Swa62, Van16, Wei96, Whi71, Wol76b, dC11]. Nineteenth-Century [Cha24, God08, Jen20, Beh08, Cat10, Dum81, Eas62, Ell57, Fle63, Goo90b, How01, Jen07a, Leg25, Mas00, Rau88, Sim65, Van16, Wei96, Wol76b]. Nineveh [Ben23]. Ninth [Res60]. Ninth-Century [Res60]. Nizan [Sch73b]. Nizolius [Fel04]. no [Her05, Jur75, BS21]. Noah [Byn84]. Nobiliaire [MR20]. Nobility [Arr94]. Noble [Arr94, HH66, Mie73, Wri43]. Noetic [Rie60]. Nolte [Mos66]. Nominalism [Cou83, Roy79]. Nominalist [Ash79, Osb02]. Non [Ano75g, Eas67, Fac75, Gew41, Goo90a, Han17, Kla71, KS24, Rin41, Spi61, Ste73, Wil70a]. Non-Communist [Kla71]. Non-Conscious [Han17]. Non-European [Wil70a]. Non-Experimental [Eas67]. Non-Jewish [Fac75, Ano75g]. Non-Mathematical [Gew41]. Non-Persian [Goo90a]. Non-Utilitarian [Spi61]. Non-Violence [Ste73]. Nonconformist [Keo74]. Nonhuman [Cha17]. nor [Gri21]. Nordau [Bur11, Mai89]. Nordic [Fie77, KS24, NS16]. Norfolk [Spr76]. Norm [Del89]. Normal [Arm07]. Normalizing [Doa17]. Normative [Can64]. Norris [Joh58b]. North [Bid54, Ger15, Fir71]. Northern [NL20, Bor68]. Northrop [Wri49, Cas48]. Northumberland [Kar66]. Norton [Les07, Mad57]. Norwich [Chr59]. Nos [Tan98]. Nostalgia [Dud61]. Note [Ano63f, Ano64e, Bak77, Bow75b, Cum78, Dra77b, Dra78, Gra65, Gre53, HM79, Kah47, Lev47a, Loe47, Lov46b, Met60, Moo45, Mul68a, Nad91, Par65, Pea48, Rot58, Soc06, Tat63b, Tay82, W.51, Wea80, Whi45b, Wol50]. Notebooks [Pri58]. Notes [And64, Cla86, Kar84, Kil61, Lyo70, Mac67, Rin41, Ros68, Sca99, Shu73, Wil59, Wil60]. Notice [Ano55a, Ano06g, Ano08g, Ano08h, Ano09h, Ano09i, Ano10e]. Notices [Ano86l, Ano86m, Ano86n, Ano87j, Ano87k, Ano88k, Ano88l, Ano88m, Ano88n, Ano89i, Ano89j, Ano89k, Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Ano91j, Ano91k, Ano91l, Ano92j, Ano92k, Ano92l, Ano92m, Ano93i, Ano93j, Ano93k, Ano93l, Ano94i, Ano94j, Ano94k, Ano94l, Ano95i, Ano95j, Ano95k, Ano98f, Ano98g, Ano98h, Ano03i, Ano04g, Ano04h, Ano06h, Ano06i, Ano07g, Ano07h, Ano07i, Ano08i, Ano10f, Ano10g, Ano10h, Ano11d, Ano11e, Ano11f, Ano12g, Ano12h, Ano12i, Ano13f, Ano13g, Ano14g, Ano14h, Ano14i, Ano15f, Ano15g, Ano16f, Ano16g, Ano17f, Ano17g, Ano17h, Ano17i, Ano18e, Ano18f, Ano18g, Ano18h, Ano19e, Ano19f, Ano19g, Ano19h, Ano20f, Ano20g, Ano21e, Ano21f, Ano21g, Ano22e, Ano22f, Ano22g, Ano23f, Ano23g, Ano23h, Ano23i, Ano24b, Ano24c, Ano24d]. Notices [Ano24e, Ano25, Pri12, Pri13b, Pri14, Pri15, Pri16a, Pri16b, Pri20, PP20, Pri21, Pri22]. Noting [Dac04]. Notion [Aqu88, Arc67, Duc06, Str76b, Van01]. Notions [BJ19]. Nottingham [Kli06]. Nouvelles [O’C84]. Novel [Bec85, Lev65]. Novels [Bis52]. Novelty [Duf80, Tho51]. Novissima [Ril99]. Novum [Lar62]. Nowhere [Ben07, Sma83]. Nueva [Fol58]. number [Mar10]. Nuovo [Boa56]. Nurture [Sta04]. Nybakken [Ger25].

O [Ano48a, Blo96, Boa48a, Col16, Dig06, Duf80, Ede63, Fog12, GB87, Man48, Ned13, Nic48, Smi65b, Str16, Wie73, Wil80]. Oakeshott [Mon74, O’S14, Pod02, Tho12, Wel94]. Oaths [McQ14]. Obedience [Ber86b]. Obituary [Ano90m, Axi93, Eng95, Fra91, Par91, Ros96b, Sch91b, Wie80c]. Object [DJ69, Fog12, Pre79, Ram14, Thr77]. Objection [Ber69, Ors69]. Objective [Cah17]. Objectivity [Syp63]. Objects [Cop91, Kau01, Sir18]. Obscurities [Str72]. Observations [Gro67, Hat75]. Observer [GCM10]. Occasion [Bar19, Sch60]. Occasionalism [Ans99]. Occult [She57, Wil88]. Occultism [Spi66]. Occupational [Haw94]. Occupy [Kap23]. Occurrence [How67]. Ockham [Bay49, Bir62, Tie54]. October [Wie73]. Ode [Alb78, Bar60b, Maz60]. Odington [Not16]. Odo [Sty19]. Oeuvre [Dem72]. offensive [Str22]. Offer [Fay78]. Office [De 53]. Officiis [Des67, vdZ98]. Ogilvie [Pol67]. Ohio [Eas62]. Old [Gra68, Kel84, La 93, Mar23, Rap69, Sea11, Tie11, Pek99, Sos73, Nor74]. Olof [For23]. Omnibus [Gre23b]. Onanism [Mac67]. Once [GJ69]. One [Chr23, Fri12, Jac78, Lev47a, Mun51, Phi07, Pod10, Rob59, Sar82, Smi24a, Scl20]. Onions [Tas69]. Only [Sar82]. Ontological [Lin91]. Ontology [Aro87, Pet12]. Open [Shi45]. Opening [Hel73]. Ophites [Ano69g, Smi69]. Opinion [Ell58, Geo57]. Opinions [Ros57a]. Opium [Afr61]. Opposites [Eme84, Gra81b]. Opposition [McL76b]. Optical [Gui13]. Optics [Can77, CM05, GCM10, Mal97, Pet94, Wel71]. Optimism [Gun62]. Option [Ano87e, Oak87]. Oral [Hut92, RS68]. Orality [Hal92]. Oratio [Cam96, Con96]. Oration [Byr06, For42]. Oratory [Sir04, Sto46]. Orchestrated [Goo18]. Ordained [Oak98b, Oak98a]. Order [Lær14, Nau15, Ned13, Pog63, Ram01, Had58a]. Orders [Hsu02, Pat62]. Ordinary [Mon89, Oak68]. Oresme [Sma91]. Orestes [Atk14]. Organic [Lew65, Oti02, Vic02]. Organical [Fin47b]. Organicism [Kee47, Man56, Phi70]. Organicist [Cas92]. Organism [Sim60]. Organized [Pea15]. Organs [Tou13]. Organum [Lar62]. Orient [Val09, Zwi20, Wri49]. Oriental [Bar58b, Ven63]. Orientalism [McC15]. Orientation [San20, Wil05]. Origen [Ter18, Tza91]. Origin [Bar88, Bos89, Bow85, BS83, Cau94, Fle59, Fri54, Goo02, Kan13, Loe47, Nad98, Oli74, Rau88, Rub91, Spr80, Was84]. Original [Har02b, Val04]. Originality [CJK^{+}43, Shu73, Tim99, Wil20]. Origins [Arm07, Car19, Cla00, Coh76, Dic80, Fei44, For89, Ger25, Gol52, Hal50, Ham08, HL19, Hen01, Her21, Hor76, Ker24, Lan12b, Lev87b, McK83, Mer11, Moo68, Mos61, Moy03b, Ned88, Ned91, Öza14, Pal18a, Par65, Par73, Pie09, Rie56, Rub86, Sil01, Skj22, Smi24a, Soc06, Ste82, Stu97, Stu04, Ver69, Vic02, Wag67, Wal59, Wei93, Wei50, Wel73, Zil41]. Orphans [Hig91]. Orsini [Ber69]. Ørsted [Wil08]. Ortega [Rog94, Rom60]. Orthodoxy [Kla93, Smi80b, Vov12]. Oskar [Ano00e, Hor78, Mah99]. Ossian [Hau98]. Osthoff [Gui08]. Oswald [Pla10]. Oswego [Man57b]. Other [Cog07, Cor77, Gue78, Lev13, Sil78, Swi75]. Otherworld [Zal85]. Otto [Har03]. Ottoman [Men16]. Our [Cop05, Lan12a, Sto46, Ban70, Hef78, Mar53, Woo24, Zag90, Val04]. Ourselves [Lan12a]. Overcoming [Laz03, Wei19a]. Overload [Bla03, Ogi03, Ros03]. Overton [McD05]. Owen [Mil54]. Owes [Pri17a]. Owl [Afr93]. Own [Bla82, Cla98, Lan12a]. owne [Woo24]. Ownership [SF93]. Oxford [Ger25, And11, Bla22, Bri79, Duh53, Hut78, Ras77, Sch03]. Oxidation [Sch64].

P [Axi93, Boa72, EST66, Sch50, Shi72]. P. [Bet59]. Pacis [Lee11]. Pad [Eng80]. Padua [Byr06, Lee09, Ned05, Ran40, Scl20, Sig62]. Paduan [Wal88]. Paganism [McC15, Pea94]. Paideia [Doo09, Bos89]. Pailleur [Rid56]. Pain [Sku64]. Paine [Chr73, Kat89, Pro72]. Painting [Fri12, Mie73, Mom50, Rog53, Ros84, Sch02, Sha88, Wes69]. Paleontology [Ant21]. Palgrave [Bow74]. Palingénésie [McC94]. Pamphlet [Wit43]. Pamphlets [Pan21]. Panacea [Tip69]. Pancirolli [Kel12]. Panormitanus [Wat72]. Pansemioticism [Wes01]. Pantheism [Par77, Tie63]. Pantheistic [Pip59]. Pantheists [Rus93]. Panther [Ben82a]. Papal [D’E07]. Paper [Dra77b]. Paperback [Mer11]. Papers [Gos60, Poj83, Rei17, Rom64, Bev17]. Papirius [Naj99]. Papist [Dav94]. Paradigm [Van15, Woo92]. Paradise [Ber14, Com71, Dun69, Ulr71, Wei75, Cha63, Den99, Lin91, McD05, Par68b, Ulr71]. Paradox [All60, Gor00, Hei73, Koh49, Lac10, Rod68, Von58, Wil48, Mul68a]. Paradoxes [Hat68b, Mul74]. Paragraph [Wes07]. Parallels [Cel01]. Paranoid [Gui23, MM22]. Pareto [Jon92b]. Paretos [Cam86]. Paris [Ari92, Dav10, Hor97, Mer90, Pec96, Rob80, Sac18, Sta95, The23]. Parlement [Mer90]. Parlements [Got44]. Parliament [Che56, Jez71]. Parliamentary [Che56, Par24]. Parmenides [Bec47]. Parody [Bec85]. Parrington [Hof41]. Parsons [Isa10]. Part [Mar59b, Smi76b, Smi24a, Smi24b, Ano51a, Hou42a, Hou42b, Kri51b, Sch23b, Wil22a, Wil22b]. Parthenian [Lea78]. Participation [Har93, Phi83]. Participatory [Chr23]. Particles [Han60]. Particular [Dug15]. Particulate [Dug15]. Partisans [Lau90]. Parva [Mer48, Mer49c, Mer50b]. Pascal [For05, Mie69, Mit93, Mue45, Ste78c]. Passage [Kei57, Whi76b]. Passion [Bar92, Jac82, Tou13]. Passions [Bos70, Haa16, Rog53, Ros84, Ser01]. Passive [Ber86b]. Passow [Con14]. Past [Ach20, Ant21, Ber55, Coe16, Ger25, Kas06, Kli55, Qui10, Ger25, Han91, Kel95]. Pastoral [Ano61d, Hen61, Pog63]. Pastors [Cla57]. Patch [Rol71]. Pater [And11, Bli82]. Path [Gre18]. Pathology [Dah15]. Patot [Kno63].

Patriarchalism [HM79]. Patriarchs [Ege66, Mah15]. Patriarchy [Hen95]. Patrick [Mor89, Zal85]. Patriotism [Chr54, De 53, Sho10]. Patrizi [Blu00]. Patronage [And04]. patrum [Wei19b]. Pattern [Fis58, Mum81, Ram14]. Patterns [Lee95]. Paul [Ger25, Wie68, Ano00e, BV88, Bla91b, Doe22, Don82, Hay56, Jac68, Mah99, Men18, Sch73b, Smi73, Hor78]. Paulhan [Doe22]. Pauline [Cum73]. Paysages [Mor77a]. Peace [Llo69]. Peak [Kja02]. Pearse [Mor89]. Pearson [Fal54]. Peasant [Por81]. Peasants [Dra90]. Pedagogy [Sch10, Whi09]. Pediatricians [Ber74a]. Peirce [Ano46g, Bar88, Bel15, CH95, Eis57, Fis72, Lov46b, Nad93, O’C64, Pet55, Tur75, Wie46a, Wie46c]. Peircean [De 06]. Peiresc [Mil06]. Penal [Mae75, Mae78, Rod68]. Penetration [Mac78]. Pennsylvania [Bro24, Cle69]. People [De 15, De 79, Feu59, Jez71, Mor18, Par04, Ros71a, RMG81]. Peoples [EST66, Rub91]. Peopling [Cog07]. Perception [Bar92, Gub05]. Perceptions [Shu22]. Perceptual [Yol79a]. Percival [Hef81, Het81]. Perfecting [Scl20]. Perfidious [Sch53]. Performativity [Fel04]. Peri [Hil66]. Peril [McC20]. Perils [Ber87, GM80]. Periodic [Phi52]. Periodical [Ano44e, Ano44f, Ano45k, Ano45l, Ano45m, Ano45n, Ano49d, Ano52f]. Periodicals [Ano40g, Ano40h, Ano41g, Ano41h, Ano41i, Ano41j, Ano42g, Ano42h, Ano43e, Ano46l, Ano46m, Ano46n, Ano46o, Ano47l, Ano47m, Ano47n, Ano47o, Ano48j, Ano48k, Ano48l, Ano48m, Ano50c, Pir25]. Periphery [NS16]. Periwigged [Joh04]. Permissive [Bis15, Tie01a, Tie01b]. Perrenial [Ano67g, Sch66]. Perry [Sch40, Neu25, Wis68]. Persanes [You75]. Persian [Goo90a]. Persisting [Ano75g, Fac75]. Personal [All66, Ten82, Tsu61]. Personalism [Ste73]. Personalists [Hel73]. Personalities [Pea54]. Personality [Sei87, Wes87]. Perspectiva [Ban14]. Perspective [Ban14, Bia24, Gug80]. Perspectives [Bla91a, Elk92, Sim81, Sol83, Wei16]. Persuasion [Blo96]. Peru [Kla76]. Perversion [Ano61d, Hen61]. Pessimism [Mit71b]. Peter [Bar42, Cal86, Ada16, Eme67, Gou06b, Hel21a, Imm09, Lew58, Ong53]. Petrarch [Bla01, Qui92, Sei65, Sti98]. Petrarchan [Sca97]. Petty [AA85, Hou41]. Phalansteries [Cro59]. Phenomena [Tou13]. Phenomenalism [Car07]. Phenomenology [Her87, Mon19, Rob20b]. Phenomenon [Bau70, Rob20b]. Phenotype [Gud96]. Phenotype/Genotype [Gud96]. Philip [EST66, Rib52, Sch50, Shi72, Axi93, Bat24, Boa72, Met96]. Philippe [Arc67]. Philo [HP18, Jon13, Rob06, Run00, Boa48b]. Philological [Fra42]. Philology [Bre61, Cop05, Hui17, Lev89, Mul09, Nau11, She03, Whi86]. Philoponus [Cla15]. Philosophandi [Sut53, Voz25]. Philosopher [Del17, Die61, Eis57, Lun21, McR51, Mon89, Pet55, Shu85, Spr76]. Philosophers [Ros78, Wil74a]. Philosophes [Hun78, Bur57a]. Philosophic [Lev56]. Philosophical [Ada92, Alt08, And64, Ber67, Bur74, Clo71, Dav95, Dup82, For03, Gar13, Gar75, Gra78, Hof85, Koy98, Kri46, Lan74, McR65, NL20, Pop54, Ram14, Rom64, Sch45b, Sch10, Sim90, Tow19, Van17, Wel94, Zag93, Zam98, Car08, Gre15]. Philosophicus [Roz02, War11]. Philosophie [All74, Ano98e]. Philosophie- [Ano98e]. Philosophies [Ell57, KR41, Pek99, Wel60]. Philosophique [Sho85, McC94]. Philosophorum [Mah15]. Philosophus [Lun83]. Philosophy [Ada89, Adk00, All75, And04, Ano67g, Ano86g, Atk14, Bar59b, Bay49, BV88, Ber86a, Bir96, Bis96, Bla92, Blu00, Boa48a, Boa63, Boc76, Bre01a, Bre52a, Bre61, Bri41, BS21, Bul05, Cad95, Car86, Cat10, Cel05, Cel14, Che51, Chr64, Coh49, Cop05, Cot94, Cro05, Cro09, Cum73, Dan60, DJ69, Dew41, Dew18, Dra77a, Dum81, Ede49, Erl02, Far91, Fir71, Fis66, Fro95, Ful94, Gar87, Gel52, Ger19, Giv62, Goo90b, Gre63, Gri41, Hal02, Har02a, Hay56, Hun00, Jer13, Joh46, Joh67b, Joh71b, Joy92, Kee92, Kil61, Kir53, Koc99, Kra40, Kub67, Lac10, Lou56, Lyo70, Mah15, Mai97, Mal40a, Mar72, Mat06, McK66, McK75, Mer47, Mir77a, Mis83, Mor53, Mor77b, Mue45, Mun80a, Nau11, O’B58, Oak98a, Ols71, Osb02, Osl85, Pag96, Par88b, Pic89]. Philosophy [Pie09, Poj83, Poo05, Por78, Pra55, Pyc87, Rad93, Ran61, Rat92, Rei17, Rei69, RG70, Rie67, Rie61, Rie63, Ril99, Rob59, Rog94, Ros55, Ruk18, Sac72, Sch04, Sch66, Sch05a, Sco98, Scr78, Sin16, Söl16, Spi61, Sta14, Sta09b, Str49, Str57, Tad22, Tay17, Ter21, Til44, Van95, Van16, Ver69, Vin07, Wac46, Wal07, Wal93, Wal04, Wie48, Wol44, Wor49, Bar71, Ger73, God08, Boa47, Boa58, Wie43, Wie68, Chi84, Fis47, HE63, Har90, Kni64, Lun77, Mit71a]. Phrase [Spe55, Sut53]. Phrenologists [Sta59]. Phrenology [Gra65, Sta95]. Phrygius [Cla11]. Phylarchus [Afr60]. Physical [Fle63, Gue83, PB43, Ver82, Gue86]. Physices [Ver94]. Physicist [BV88]. Physicists [Kar69]. Physico [Jes86]. Physico-Theologian [Jes86]. Physics [Aga69, Gla01, Hen01, Mir77a, Per69, Sac76, Wil08, Wit77, Lan95, Wie43]. Physiocracy [Mou49, Nei48]. Physiocrat [Alb55]. Physiognomy [Bar75, Gra61, Sta95]. Physiological [Rya01]. Physiology [Bas13, Coh94, Cru51, Jen20, Pac02, Ree52, Ste88]. Pico [Bre52a, Bre52b, Cas42a, Cas42b, For42, Ter18]. Picture [Lo17]. Pictures [Kus97]. Piedmont [Cav16]. Pierre [And02, Bak76, FR07, Hic18, Isr19, Keo74, Men18, Mon97, Moy07, O’C84, Pau72, Sak09, Sar85, She02, Sho24, Tin96, vdL17, Per57, Die61]. Pierre-Daniel [She02]. pietate [Del17]. Pieter [Zag03]. Pietist [Whi09]. Pietistic [Scl20]. Pilgrimage [Zal85]. Pill [Nic68]. Pilot [Tuv66]. Pioneer [Kan45, Kee47, Mae73, Sch43a]. Pious [Oos13]. Pirate [Pie10]. Pisan [Gab55]. Pitfalls [MM22]. Pity [Car84, Sir18]. Pivotal [Ano03f, Bay49]. Pizan [Mar86, Wil66a]. Place [Aga69, Ber14, Cas45, Cla82, Gor08, Gre54, Hey82, Kri44, Ran53, Wit09]. Plagiarism [Ces05, Eig09, Sel13, Zah12]. Plan [Lyo03]. Planck [GR86]. Planning [And21, Chr23, Jac12]. Plantations [Arn94]. Planter [Pop06]. Platner [Bre01a]. Plato [Coh49, And11, Bec47, Boy02, Coh62, Ede49, Fak65, Fis66, Gro96, Gue78, Had58b, Joh67a, Koy43, Lev79, Mer47, Ran67, Sha63, Shu63, Sol83, Tha75, Tig70, Wis49, Woo60, Kah61]. Platonic [Bur44, O’B58, Sol78, Wol61, Zan87]. Platonica [Jac85]. Platonism [Kri58, And11, Bat48, McK82b, Ram14, Wen76, Sta74]. Platonist [Sma83, Zeb94, Cou75, Hut78, Nar05, Sav59]. Platonistic [Gir70]. Platonists [Arm69, Sav60, Rog79, Sta74]. Play [McC58]. Playe [Ant68]. Pleasure [Pac89]. Pleasures [Ald44, Sca97]. Plekhanov [Bar58b]. Plenitude [GM80, Kuk00]. Pliant [Wes11]. Pliny [Doo09]. Plot [Mac01]. Plotiniana [Tay98]. Plotinos [Ran69]. Plotinus [Ada01, Bro74, Ioz22, Kat54]. Pluralism [Sta91]. Pocock [For16, Isr16a, Ros16a, Ski25, Wol16b]. Poem [Roz02]. Poet [Shu85, Vie49, vE01, Fie55]. Poeta [Rin41, Wes99, Kal95]. Poetess [Mar86]. Poetic [Aqu88, Maz53, McC42, Tig70]. Poetics [Kel96, Mai77, Mar04, Sku58, Smi83, Tuv42]. Poeticus [Tig70]. Poetry [Ada79, Bee20, Bri85, Mar23, Maz53, Ogd49, Pip59, Pod02, Rid56, Ros71b, Sel74, Ste58, Tha75, Wil08, Min51]. Poets [Gre53, Tas69]. Poincaré [Lal54, Nye79]. Point [Koh69]. Points [Twi20]. Pointsman [Sta08a]. Poland [Kos24, Lew58, Sko65]. Polanyi [Imm09]. Polarity [Lar59]. Polemic [Alm78, Elu92, Ila12, Ros96a, Tri87]. Polemica [Boa56]. Polemics [Puc05, Sto85]. Poles [Poi70]. Police [Dod08]. Policies [Cro53b]. Policy [CR81a, San82, Vin07]. Polis [KE23]. Polish [Rie63]. Polite [Oos13]. Politica [Mos98]. Political [Ach20, Bar83, Bar47, Bar60a, Bar41, Bay49, Beh08, Ben82b, Bly86, Bur57b, Bur60, Cam84, Car08, Cha24, Che59, Cri94, Cro85, Din75, Eck44, Ede22, Fac21, Fat05, Fon99, Gol72, Gol88, Gro67, Gun68, Ha20, Har85b, Har86, Hin92, Hob80, Hun74, Ila12, Ina20, Jen07a, JN12, KM94, Kap08, Kel79, Kel98, Kel04, Kil11, Kim24b, Kli24, Koc99, Kri64, Lin48, Lyo70, Mac78, Mar72, McA99, McK81a, McR51, Mis83, Mor98, Nak22, Ned88, Ned96, Ned05, Ned13, Nob65, Nur61, Oak68, Pac02, PB73, Pau79, Pod10, Por81, Por02, Ras77, Ras85, RG70, RMG81, Rie56, Ril94, Ril99, Ros70, SB83, Sch22a, Sho99, Sho16, Skj22, Smi80b, Sol03a, Spi76, Sta14, Sta91, Sta87, Str49, Tre21, TW21, Tri56, Wal19, Wea80, Woo18]. Political [Yin22, Ned91, Smi19, Koh64b]. Politician [Coo99]. Politics [Ami07, Ano83f, Bak83, Boy06, But06, Cal22, Cro53b, Fri73, Gar12, Gee99, Gui23, Haw01, Hor80, Ian00, Jac77, Joh85, Lar78, Lin10, Mag16, Mal09, MM22, Mil54, NL20, Oak98a, Par68a, Pet84, Por78, Rob15, Sch73b, Sch89b, Sel63, Sho18, Sim04, Spe98, Ste89, Sum17, Tag86, dM49, Bla22, Had52]. Politicus [Ant66, Bat77, Nad13a]. Politique [Col66, Lad62]. Politiques [Bea93]. Polity [Cra44, McG63, She06a]. Pollution [Mul96, Vis84]. Polybius [Chi40, Dym21, Ede22]. Polydaedali [Pop05]. Polyglot [Mil01]. Polyhistor [Wes01]. Polyhistors [Mul06]. Polyp [Var50]. Polytheism [Fun94]. Pomponazzi [Kri44, Pin68]. Pontano [Nau11]. Ponty [Cho15, Her05]. Pope [D’E07, Dau77, Ros75]. Popeliniere [Syp63]. Popery [Sch60]. Popper [Ano77e, Cre83, Hac98, Har03, Len97, McL76a, Wet05]. Popular [Ala67, Ces04, Cha23a, Lee11, Mac56]. Popularization [Bet59, Cot94]. Population [Bla73, Ros70]. Porcari [D’E07]. Porphyry [Dom84]. Port [Mie69]. Port-Royal [Mie69]. Portable [Byr23]. Portuguese [Mar20, YY84, vG20]. Positive [Bar43]. Positivism [Ard63, Dav95, Flo49, Gra75, Mil63, Pea15, Puc05, Wol76b]. Positivist [Bev93]. Possession [OR13]. Possibility [Kuk00, Wac96]. Possible [Sch83, Woo67]. Post [Asf99, Bev12, Fen86, Gui85, Lac99, Lan66, Nad93, Ste81, Tut22]. Post-Analytic [Bev12]. Post-Bellum [Nad93]. Post-Biblical [Lan66]. Post-God [Lac99]. Post-Kantian [Fen86]. Post-Newtonian [Asf99]. Post-Reformation [Tut22]. Post-Revolutionary [Ste81]. Post-Socratic [Gui85]. Postcolonial [Ger25]. Postel [Bou54]. Posterity [Eck45]. Posthumous [Kla76]. Postmodern [Hab95]. Postmodernisms [Ano00f, Hol00]. Postscript [McR63]. Potential [Ham08]. Poulain [Stu97, Wil19, Sei74]. Poverty [Tei77]. Power [Ash78, Bly86, Bro24, But15, Ces04, Cha23a, De 03, Hun52, McG63, Neu77, Oak98b, Oak98a, Roy79, Sch02, Tea62, Wes11, Wie63, Wil96]. Powers [Oak68, Rik54]. pp [Bar42, Ger25]. pp. [Ger25]. Practical [Bou90, Kra40, Min98, Bar71]. Practice [Bag92, Gra06, Hol17, Hsu02, Joh06, McK81a, Rap16, Sor83, Vin07, Wal19]. Practices [Car19, Mar10, Mul06, Wal11]. Pragmatic [Mat55, Wie48, Bre01a]. Pragmaticism [Ano46g, Wie46a]. Pragmatism [Bar88, Bra24, Bun95, Dot86, Fis54, Kra40, Wie46c, Zan79, Sch50, Gre69b, Wie75]. Pragmatist [Gul62, Hor97, Kau63]. Pragmatists [Cot94]. Prague [Wei16]. Praise [Per02]. Prayers [Hab90]. Pre [CR81b, Dan52, Ech72, Fal20, Lef76, Pri48, Ruk18]. Pre-Classical [Dan52]. Pre-Confucian [Pri48]. pre-Critical [Ruk18]. Pre-Enlightenment [Lef76]. Pre-Established [CR81b]. Pre-Revolutionary [Ech72, Fal20]. Preacher [Spr76]. Precedent [Cul70]. Precept [Gra06]. Préciosité [Mor40]. Preconceptions [Giv62]. Precursor [Joh40, Pol43, Pop53a]. Predecessor [For16]. Predecessors [Boa59, Van17]. Predestination [Goo83, Men16]. Predicament [Nis21]. Predicate [Res60]. Preeminence [Bla82]. Preface [Fel04, LS91, Ros92, Pri69]. Pregnant [Kja02]. Preliminary [Pim18]. Prelude [Dud57]. Premodern [DeV08]. Preparation [Mil43]. Prescription [Kir53]. Prescriptions [Shu22]. Presence [Bro18, Fef05, Nad88, Osb02]. Present [Koy98, Zam98]. Presentness [Ber55]. Preserved [Gar54]. President [Ban16]. Presocratic [Che51]. Press [Ell58, Ger25, Lau98, NL20]. Presse [Tan77]. Presumption [Bla19]. Presupposed [New01]. Previous [Zuc75]. Previously [AA85, Bla22]. Price [Bla19, Mol97, Pet84, Zeb94]. Priestley [Can84, Fau78, Gar73, Hat75, Sch64, Tou57]. Priests [Sus12]. Prima [Kin56]. Primates [Mor93]. Primato [Dur43]. Prime [Sol58]. Primeval [Goo11]. Primitive [Gui23]. Primitives [Mor93, De 15]. Primitivism [Mal58]. Primitivist [Mer50a, Wal70]. Prince [Har89, Liz22, Maz60, Syp41]. Princely [Lee11]. Princes [Smi80b]. Principes [Lyn82]. Principia [Alb78, Bro87, Rog78]. Principle [Bal42, BH59, Cho15, KS24, Lar59, Mor78c, Mos60, Rap81, Ros70, Sil73]. Principled [SR81]. Principles [Bid76, Con14, Hil66, Ogd49, Sch46b, Bou97]. print [Guc00, ME21]. Prior [Ano84f, Clo84]. Priorities [Str05]. Prisca [MH04]. Prism [Sel13]. Prison [Coo81]. Prisoner [Lub24]. Private [Bro97, Jac76a]. Privatum [Gro79]. Prize [Ano47k, Ano03h, Ano09g, Ano15e]. Probabilism [Sch14b]. Probability [Coh80, Ran66b, Wal93]. Problem [Avi61, Bar59a, Bec47, Bur93, Cro53a, Dav60a, Day08, De 41, Ecc76, Fle47, Hal50, Har02b, Har66, Her70, Hey17, Hut92, Jay70, Kan13, Kat84, Koe77, Lan74, Lev97, Liz22, Loe47, McM66, Moo79, Mor78c, Mul97, Nad13a, Nie86, Pal06, Par24, Par69a, Pax93, Pin68, Pre71, Ran23, Sch43c, Sup54, Teg40, Tie01a, Tri49, Tri96, Wag54, Wie51, Wil51, Wit61, Sew03]. Problem-Solving [Kat84]. Problems [Bar58a, Gom43, Lac65, Von81, Wie61, Ell67]. Process [Cla86, Par62, Söl16]. Processualization [Rei21]. Producing [Mat06]. Production [Val95, Wal86, vI11]. Productive [Met63]. Profession [DeM74]. Professional [Het81]. Professor [Mai21, Wie70, Dra77b, Lic73, Lov44, Mac80, Mir77b, Shi72, Sil77, Boa48b, Fis47, Wri49]. Professors [Sch73b]. Professorship [Ano75f]. Profiat [Bla93]. Program [Lov67]. Programmers [dS23]. Progress [Ano56e, Ano66f, Ber18, Cha58b, Dod68, Fal54, Ger73, Hei73, Kel50, Les72, Lev66, Mol78, Mom51, Mor90, Nel87, Pri64b, Ran23, Sim56, Smi63, Vic92, Wag67, Wel55, Wol70, Zil45, Har02a, Jac76a]. Progressive [Cro66, Rob15]. Progressivist [Mer50a]. Project [Bul05, Isa10, Wie40]. Projectile [Fra03]. Prolegomena [Obe87, And11]. Proletarians [Coh68]. Prologue [Spr68]. Prometheus [Bea91]. Promises [McQ14]. Promotion [Fit97]. Proof [Blo96]. Propaganda [Chr89, Pol67]. Proper [Thr77]. Property [Arn94, Bis15, Bow85, Lee09, Oli74, Pie10, Pos86, Rap81, Tei77, Tie01a, Tie01b, Ven76, Win81, RMG81]. Property-Rights [Pos86]. Prophecies [Vir86]. Prophecy [For82, Har99]. Prophet [Pau83, Wol76a, Gey55]. Prophets [Kue42]. Proposal [Ras82, Wit43]. Proposed [Lea82]. Prose [Woo24]. Proslogion [Nis21]. Prospect [Kel87]. Prostitution [Rie68]. Protected [De 15]. Protectorate [Jac77, Woo18]. Protestant [Cre95, Dav94, Hje11, Kla93, Pri13a, Sch22a, Sta08b, Tas69, Gug80]. Protestantism [Ros16a]. Protestants [Gre18, O’B85, Zub18]. Protoplasm [Bli82, Hal50]. Prototypical [Cas81]. Proud [Kee12]. Proudhon [Nol67]. Proverb [Iva24]. Providence [Col76, GJ69, Hic10, Kub67, Nad01, Osl83]. Proximity [Ant21]. Prussian [Ter21]. Psammetichus [Suł89]. Pseudo [Cam96, Con96]. Pseudo-Donation [Cam96, Con96]. Pseudonymity [Mul06]. Psychathanasia [Sta68]. Psyche [Sol83]. Psychiatry [Ant24, Gui23, Rob20a]. Psychoanalysis [Hut87]. Psychodia [Jac85]. Psychological [Dav69, Ger67, Gus50, Mat57, Poo99, Wei52]. Psychological-Romantic [Mat57]. Psychologism [Smi50]. Psychologist [Gol82a]. Psychologization [Nan21]. Psychologizing [MM22]. Psychology [Bel15, Blo84, Cru51, Gar77, Gar54, MM89, Rat41, Rie56, Ros12, RS68, Rya09]. Psychopathology [Sch15]. Psychotherapy [Gil85, Rob20a]. Ptolemy [Sch23b]. Puberty [McA10]. Public [Ell58, Jac76a, Lin10, Mac92, Red97, Tho24, Vin07, Wie40]. Publica [Hat68a, Oos13]. Publication [Ber18, Lin10]. Publicity [Cau94]. Published [Wil96]. Publishers [Sto85]. Pufendorf [Haa16, Kri60, Saa10]. Punch [Cam16]. Punic [Sal96]. Punishment [Fle47, Mae73]. Pupils [McC63]. Purchase [Coa61]. Pure [Bou03, Boy02, MB81, Wal81]. Purgatory [Zal85]. Puritan [Ada89, Can73, Con79, Joh71a, Mac56, Mal17, Pea52, Spr68, Spr76, Van82, Wol68]. Puritanism [Gre69a, Ken90]. Puritans [Nel76]. Purity [Mul96]. Purpose [Alm78, Wes58]. Pursuit [Dac04, Gra63b, Not16]. Pyramid [Sha02]. Pyrrhic [AJ81]. Pyrrhonian [Ber04]. Pyrrhonienne [Len77]. Pyrrhonism [Sou92]. Pythagoras [Oos13].

Q. [PB73]. Qichao [Ang00]. Quaker [Cla85, Sil86]. Quakers [Sta59]. Qualitative [Wel71]. Quantitative [Eas67]. Quarrel [Lev91, Tha75, Tho24]. Quarrels [La 93]. Quarter [ZMM10]. Quattrocento [Tri56, Qui10]. Queen [Kei57]. Queene [Fei68, Gre74]. Quentin [Lan12a, Pet12]. Querelle [Bar59a, Bul05]. Quesnay [Nei48]. Quest [Pal97, Phi07, Web22]. Question [Cai07, CJK^{+}43, Fin14, Gra88, Kee12, Kro84, Loe61a, Ort18, Sch89a, Spa66, TW21]. Questions [Dor40, Rap97, RS64, Tip69]. Qui [Bak91]. Quiet [Pur15]. Quis [Ell67]. Quixote [Ham98, Nad97]. Quotations [Bau46, Kah47]. Quote [EMR13].

R [Bin67, Cah17, Car23, Ger25, Hef78, Nag57, Ran49, Ros57b, Sim55, Sol58]. R. [Bou97, Gut79]. Rôle [Che56, Wac46]. Rabbinic [ME21]. Rabble [Smi80b]. Race [Åsb21, Ami07, BS21, Dun21b, Eze00, Ger25, Mor18, Pau81, Sel58, vG20]. Race-Thinking [Sel58]. Races [Fal54]. Racial [Hor76]. Racialized [Bur11]. Racism [Åsb21, Fie77, Imm92, Puz64]. racist [Zub18]. Radical [Cre79, Dav13, Ell03, Keo77, Mal17, Spr76, Wes07, Isr16b, Sch89b]. Radicalism [Cri94, Din75, Kat89]. Rainbow [EG84]. Raison [Bak91]. Raisonnée [Yol79b, Lef76]. Ralegh [Ham96, Sal96]. Raleigh [Sta41]. Ralph [Gar54, Gre15, Mul09]. Rameau [Hun86, Pau71]. Ramism [McR63]. Ramist [McR55, Ong53, Spr76]. Ramos [Rom60]. Ramus [Tuv42, Cum78, Gou06b, Ong53, Ros40]. Randall [Boa63, KNS^{+}81, Wal88]. Randolph [Ber57]. Rank [Wal86]. Ranters [Guc00]. Raphael [Ger25, Gut41]. Rasa [Eng80, Woo92]. Rasselas [Fis58]. Ratio [RH79]. Rational [Sca99, Web22, Spu88]. Rationale [HL65]. Rationalism [Kla93, Ros96a, Van16, Ver73]. Rationalist [Hey79]. Rationalists [Ros43b]. Rationality [Bui02, Cal22, Dan86]. Rationalization [Kae96, Sta67b]. Ratner [Ros61]. Rauh [Hor97]. Rawls [Bev17, Bev17, Bok17, Bot17, Gal17, Kel96, Rei17]. Ray [Feu79, Cha58a]. Raymond [Cam86, Dro19]. Raynal [Ans09]. Re [All66, Hat68a, Zha25]. Re-Enchantment [Zha25]. Re-Examination [All66]. Reach [Mon44]. Reaction [Ble58, Fel77]. reactionary [Shu22]. Reactions [Ell58, Sav60, Sou89, YY84]. Read [Len02, Pop05]. Reader [Gra92, Gro99, Mor16, Ros92]. Readers [Kee12, Mal17, Mor09]. Reading [Ada16, Bet09, Bla03, Bro97, Feu58, Gim08, Gro99, Loc07, Pal12, Qui10, Rap16, Ree91, Ric12, Rya09, Sch10, Sim90]. Readings [Cat10]. Readjustments [Moo45]. Reads [Sty19, Zah12]. Real [Cla82, Nad88, Osb02, Pea54, Sin89]. Realism [Cah17, Fei44, Fer55, Rum65, Sou70, Del80]. Reality [Bro77, Nad93, Run00, Sch54, Web81]. Really [Pri17a]. Realpolitik [Hol60, Nak22]. Reappraisal [Arn57, Sch84]. Reappraised [Dra75, Fra77]. Rearming [Hea88]. Reason [Ano48a, Arm00, Atl52, Bar81, Bea40, Con82, FS80, Gri12, Hol50, Kuk00, MB81, Mor89, MS61, Orr86, Per02, Pro72, Rap81, Rog66, Rya01, Sga15, Sny86, Spi66, Ste75, Wal81, Bui02, Cro05, Mul94, Tum96]. Reasonable [Wal59]. Reasonableness [Gre81, Wal84a]. Reasoning [Car03]. Reasons [Dav94, Guy02]. Reassessed [Mac98]. Reassessing [BG92]. Reassessment [Fay78]. Reassessments [Mor20]. Rebellion [Gel89, Rei56]. Received [Ano40a, Ano40b, Ano40c, Ano40g, Ano40h, Ano41b, Ano41c, Ano41d, Ano41e, Ano41g, Ano41h, Ano41i, Ano41j, Ano42b, Ano42c, Ano42d, Ano42e, Ano42g, Ano42h, Ano43a, Ano43b, Ano43c, Ano43d, Ano43e, Ano44a, Ano44b, Ano44c, Ano44d, Ano44e, Ano44f, Ano45b, Ano45c, Ano45d, Ano45e, Ano45k, Ano45l, Ano45m, Ano45n, Ano46c, Ano46d, Ano46e, Ano46f, Ano46l, Ano46m, Ano46n, Ano46o, Ano47b, Ano47c, Ano47d, Ano47e, Ano47l, Ano47m, Ano47n, Ano47o, Ano48b, Ano48c, Ano48d, Ano48e, Ano48j, Ano48k, Ano48l, Ano48m, Ano49a, Ano49b, Ano49c, Ano49d, Ano50a, Ano50b, Ano50c, Ano52b, Ano52c, Ano52d, Ano53a, Ano54a, Ano54b, Ano55b, Ano56b, Ano56c, Ano56d, Ano57c, Ano57d, Ano57e, Ano57f, Ano58c, Ano58d, Ano58e, Ano58f]. Received [Ano59d, Ano59e, Ano59f, Ano59g, Ano60a, Ano61a, Ano61b, Ano61c, Ano62b, Ano62c, Ano62d, Ano62e, Ano63b, Ano63c, Ano63d, Ano63e, Ano64b, Ano64c, Ano64d, Ano64a, Ano65b, Ano65c, Ano65d, Ano65e, Ano66b, Ano66c, Ano66d, Ano66e, Ano67c, Ano67d, Ano67e, Ano67f, Ano68d, Ano68e, Ano68f, Ano68g, Ano69c, Ano69d, Ano69e, Ano69f, Ano70c, Ano70d, Ano70e, Ano70f, Ano71b, Ano71c, Ano71d, Ano71e, Ano72a, Ano72b, Ano72c, Ano72d, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano74c, Ano74d, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c]. Received [Ano82d, Ano83b, Ano83c, Ano83d, Ano83e, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c]. Received [Ano04d, Ano05a, Ano05b, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano11a, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, Ano18a, Ano18b, Ano18c, Ano19a, Ano19b, Ano19c, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano22a, Ano22b, Ano22c, Ano23a, Ano23b, Ano23c, Ano23d]. Reception [Dav13, Gre23a, Jan78, Man08, Meg87, Par15, Poo99, Sak09]. Reciprocity [Yin22]. Reclaiming [Mit93]. Reconceived [Hut87]. Reconciliation [Fak65, Gra81b, Mal99]. Reconciling [Nad88]. Reconsideration [App64, Ham08]. Reconsiderations [Afr62, Mos41, Phi96]. Reconsidered [Cla98, Ham70, Liu04, Sch14b, Swe73]. Reconsidering [Dym21, Van15]. Reconstruction [Cyw69, Had79, Jur02, Pal72, Ran41]. Recovering [Goo11]. Rectification [Che63]. Recurrence [Löw45, Sma91]. Red [PD63, Hac98]. Redefining [Yin22]. Rediscovered [Sch98]. Redivivus [Wal88]. Rée [Don82]. Reenvisioning [Wei96]. Reexamined [Cri94]. Refer [Mar10]. Reference [BBGM09, Gro11, Rid62]. References [Bau46, Cum10]. Reflected [Kib46]. Reflection [Hip04, Sha70]. Reflections [But61, Con86, Gab98, Gor19a, Lov40, Sar83, Sav44, Stu04, Wol68]. Reflex [Phi71]. Reflexive [Jer13]. Reform [Cha58a, Coo81, Her48, Rod68, Ruk18, Web22, Ram19]. Reformation [Dam79, Dod08, Hea87, Kos24, Lin74, Obe87, Tri49, Voz25, Hen95, Tut22]. Reformed [Cas78, Ver22]. Reformer [Mae78]. Reforming [Bee20]. Refraction [Eas67]. Refutation [De 06, Pat86, Sil93]. Regal [Bly86]. Regarding [RS64]. Regeneration [For93, Wer13]. Regenerative [Men20]. Regime [Dev70, Las47, Pek99]. Regimens [Ric12]. Regimes [Bur93]. Reginald [Dav13]. Regiomontanus [Byr06]. Regius [Ver94]. regnare [Bak91]. Regulating [Gol88]. Regulation [Bla19]. Regulative [Las47]. Rehabilitation [Mul96, Shk81]. Reich [Bra71]. Reid [Cal62, Can77, DB77, Fau78, Fei44, Kiv70, NSR80, SRN84]. Reign [Bak91, Ebe69]. Reimarus [CW13]. Reinhart [RR06]. Reinhold [Cra92, Wel84]. Reintegrating [Jur01]. Reinvention [Kos24]. Rejection [Sea68, Wri51, Yos51]. Rejoinder [Cat73b, Len04, Mir77b]. Relate [Lun77]. Relating [Fis82]. Relation [Ano83f, Bak83, Dic80, For84, Gar87, Hou41, Kro84, Lid66, Log75, Mat55, McK81b, Moo45, Vir86]. Relational [Mil87, Van15]. Relationalism [Cas88]. Relations [Atl52, Bed21, Bou94, Flo46, Mil61, Mor20, Sel13]. Relationship [Jac85, Osm73, Sho99, Zuc75]. Relative [Mae83, NL20]. Relativism [Cah17, Hig95, Kor75, Lan74, Mar53]. Relevance [Von81]. Religion [Aus70, Beh08, Bos73, But61, Col99, Cri86, Cro05, Cro09, Cro10, Dav13, Fon99, For84, Gar55, God08, Gro14, Hes18, Hey17, Hob71, Hol71, How99, Isr16a, Joh06, Kae96, Mac92, Mar09, Mil06, Mit96, Mos78, Mue43, Naj99, Nau52, Nau02, Ned13, New95, Odo66, Par44, Phi07, Pre79, Pre89, Psa60, Rob62, Ros16a, Rus88, She06a, Spu88, Tum96, Wac46, WB93, Wol16b, Dup82]. Religionis [Gro79]. Religionist [Bla49]. Religiosity [Liu12]. Religious [Car62, Fai41, Fer54, Gre18, Jon73, Kei04, Lam56, Man58, McM66, Nur61, Oak53, Pea14, Ram19, Ram14, Ric50, Rob20b, Row56a, San60, Sus12, Sut03, Syp63, Tae53, Tou13, Wri40]. Reluctant [Car62, Erl02, Pie10]. Remains [Wun03]. Remapping [Lea78]. Remarks [CJK^{+}43, Coh80, Gor19b, Igg59]. Rembrandt [Sli53]. Remembering [Shu85, Wie86, Ano86e]. Remembrance [Hig95]. Remnants [Gre23b]. Remonstrances [Mer90]. Removed [Åsb21]. Renaissance [Bar62, Hor78, Spi66, Tri56, Wei72, Ano61d, Atk74, Bar75, Bar43, Bar58a, Bar59a, Bat48, Bau70, Bau23, Ber14, Bla82, Bla91a, Bor68, Bou54, Boy02, Bul05, Cam96, Cas42a, Cas42b, CJK^{+}43, Cel01, Cel14, Cer05, Clo69, Com71, Con96, Dan52, Dra70, Duh53, Elk92, Fer51, Fer54, Fer55, Fer71, Flo95, Gee99, Gra63b, Gre91a, Gre65, Hat68b, Hen61, Hug43, Joy92, Kal95, Kel50, Kli81, Kor74, KR41, Le 45, Liz22, Loc04, Log75, Low54, Lyn00, Mac01, Mac98, Mar23, Maz54, Moh04, Mon97, MH04, Nel00, Ogi03, Oos13, Orr86, Pal12, Pat58, Pat59, Pat62, Rig98, Shy14, Sir04, Ste62, Ste67a, Syp65, Tou13, Tri49, Tro73, Von58, Wei45, Wei50, Wil41, Wit61, Woo18, Wri40, Zan87, Kri51a, Bar47, Bar50, Bar60a, Phi52]. Renaissance [Tri90, War67]. Renan [Ala67]. Renaudot [Zwi20]. Rendered [Sir18]. René [Wei69, Cre83, MR20, Nad22]. Renouvier [Ter21]. Replies [Yol51]. Reply [Ano69l, Ber69, Ber80, Bla91b, Bre52b, Dra78, Gan56, Gre91c, Gut79, HH57, Isr16b, Kat44, Ken73, Lic73, Lov44, Mac80, McC90, Min70, Mir77b, Mor04a, Mor04b, Pri49, Rod88, Ros61, Sea12, Sil77, Ski12, Sol58, Sol73, Vie42, Wen82, EE70, KK66, Wie70]. Report [Hui17]. Representation [Gro11, Kel04, Sch02, Tol98]. Representations [Dow40]. Reprints [Ano40g, Ano40h, Ano41g, Ano41h, Ano41i, Ano41j, Ano42g, Ano42h, Ano43e, Ano44e, Ano44f, Ano45k, Ano45l, Ano45m, Ano45n, Ano46l, Ano46m, Ano46n, Ano46o, Ano47l, Ano47m, Ano47n, Ano47o, Ano48j, Ano48k, Ano48l, Ano48m]. Reproduction [Vel19]. Republic [AU19, Pir25, Thr11, Twi20, Car19, Elu92, Esk04, Har93, Her92, Hia09, Lie98, May04, Öza14, Ran23, vI11]. Republican [D’E07, Fat05, Gue95, Pal09, Poc87, Tri56, Jur01]. Republicanism [Car91, Col99, Ela15, Wes11, Ent18, Urb12]. Republics [Car91, Ent18]. République [Nic19]. Repurposing [Jen22]. Reputation [Pop54]. Rereading [Bro99, Ros92]. Res [Oos13]. Resartus [Hab93]. Rescue [Bun95]. Research [And21, Bar19, Fra42, HL19, Zhu21]. Resignation [Por72]. Resistance [Ros98, Rue64, Sho79, Tuc03]. Resisting [Mul96]. Resolution [Sew03]. Resonance [Tri87]. Responsa [ME21]. Response [AHJ95, Col07, Cor98, Gar55, Gel89, Jum08, Les07, Liv08, Mal15, McQ14, Mor86, O’B08, Par08, Per08, Poc16, Ren76, Tay82, Ten82, Tie11]. Responses [Fin14, Hje11, Mor25]. Responsibilities [Ada49]. Responsibility [Str76b]. Restatement [Hig95]. Restoration [Kel65, Sel58, Lut70, McC15, Spu88]. Resurrection [Ols68]. Rethinking [Cam22b, Smi24a, Pag88]. Rethinks [Aro87]. Retinal [Tur55]. Retracing [HP18]. Retrospect [Kel87, Nag57]. Return [Aro87, Liv08]. Returns [Hui17]. Revaluation [Scr78]. Revelation [Har93, Jac77]. Review [Ack73, Ano69l, Bar47, Bar50, Bar60a, Bar62, Bar42, Bea71, Bea76, Bid54, Bin67, Boa47, Boa48b, Boa56, Boa58, Boa63, Boa64, Boa67, BB70, Boa70, Bri42, Bro61, Bur57a, Cas45, Cas48, Cen89, Chi84, Coa50b, Coh40, Coh49, Cri66, Cul88, Del80, Die61, Dod68, Dup74, Dup82, Dyn81, Eck65, Ede52, EE70, Eli67, EST66, Fac75, Fel77, Fer68, Fie55, Fir71, Fis47, Fis70, For89, Fri40, Gab56, Gas55, Ger72, Ger73, Ger25, Gey48, Gey55, Gil53, Got44, Gra75, Gre69b, Gro14, Gug80, Had52, Had58a, HE63, Ham58, Han91, Har90, HH57, Hea87, Hig91, Hil72, Hor78, Jac76a, Jac82, Joh59a, Jum12, Kah61, KM94, Kas77, KK66, Kel79, Kel95, Kni64, Koh64b, Koh69, Koh70, Kri70, Kri47, Kri51a, Kri58, LaC88, La 93, Lar78, Lev71, Lev72a, Lev87a]. Review [Lun77, Mat43, Met63, Mie73, Mil41, Mil56, Min51, Mit71a, Mit94, Mod94, Mom50, Moo68, Mor57, Mor78b, Mos66, Nag68, Nak69, Neg79, Nor74, Pag88, Par62, Phi52, Poc78, Pri58, Pri63, Pri74, Ran41, Ran49, Rie61, Ros57a, Ros74, Ros85a, San59, Sap79, Sar82, Sav44, Sch41, Sch65, Sch40, Sch44, Sch50, Sch89b, Smi65b, Spe48, Spi66, Sta73, Sta74, Str23, Tor84, Tri90, Tur71, Var73, Vie42, Wal63, War67, Web79, Wei69, Wei72, Wel84, Wes76, Whi45a, Whi76a, Wie43, WW50, Wie51, Wie68, Wie70, Wie75, Wri49, Ano98e, EE70, LaC88, Lom02, Sir18]. Review-Article [Koh69]. Review-Essay [Gro14, Jum12]. Reviews [Fie96, She44, Tri56]. Réville [Smi80a]. Revised [Imm92]. Revising [Pal18b]. Revision [Koh64a]. Revisionism [Ell67, Pal09, Gri21]. Revisions [Pos86]. Revisited [Afr62, Bou03, Cou83, Kir94, Mer08, Mit65, Nel87, Oak66, Pal06, Pal18a, Pro72, Sku58, Woo88, ZMM10]. Revisiting [Dew18, Jen13, O’N12, Sil93]. Revitalizing [Cen89]. Revival [Alb78, Ber04, Fai41, Ste74]. Revolt [Bur74, Flo49, Mor89, Whi47]. Revolution [Ano84f, Bec71, Clo84, Coh76, Col76, Con86, Cop91, Dic80, Ede22, Ent18, Gra62, Hal74, Ham72, Han59, Har86, Hoo73, Mel75, Mer11, Mit96, O’F92, Pal40, Pre77, Rac95, Ree83, Sch73a, Sch90, Skj22, Smi76a, Tho77, Wes15, ZMM10, Ans09, Axi71, BI60, Ber87, Bro24, Cla57, Coh78, Ede18, Gar73, Ger51, Guy20, Han61, Kim18, KE23, Lev03, Mac56, McN45, Mum81, O’N07, Pey49, Pla13, Rac95, Sch43c, Tho94, Tuc03, Urb12, Voz25, Cen89, Fel77]. Revolutionaries [Mos85]. Revolutionary [Ech72, Fal20, Leg25, Mit87, Pag96, Pau79, Rad93, Val24, Cam16, Cre79, Hut76, Kam70, Nak22, Rin61, Ste81, Zal93]. Revolutionizing [Ha20, Jur75]. Revolutions [Ade18, Ano86f, GR86, Kah85, Pal18a, Syp65, Pal18b, Ros75]. Rheticus [Ros40]. Rhetoric [Bac07, Bla82, Bou01, Bre52a, Bul05, But06, Cad95, Duh49, Fit97, Hut87, Maz62, Mun90, Nis21, Nor74, Sir04, Sta08b, Wal93, Wol70]. Rhetorical [Blu00, Gar12, Yar86]. Rhoda [Ger25]. Rhythm [Doa17]. Richard [Ger25, Whi45a, Alm78, Cal86, Cha24, DeM74, Ecc76, McD05, MM22, McK81b, Mol97, Pet84, Ras77, Smi61, Smi17, Ste88, Vin07, Zeb94]. Richardson [Ada89, Man94]. Richelieu [Tho07]. Rickert [Cro10]. Ricoeur [Cho15]. Riegl [Gub05]. Riezler [RGE45]. Right [Bec71, Cas10, Per18, Rei56, Rue64, Shu22, Wil19, Woo24, Mul94]. Right-Wing [Per18]. Rightist [Vir50]. Rights [Ang00, Bis15, Ede18, GSJ18, Jan78, Lah97, Lee09, Mol97, Pos86, Sea11, Sho18, Sir18, Sta87, Tae53, Tie91, Tie11, Wil75, Zub18, Kat89]. Rigorism [Not55]. Rise [Ano00f, Cas88, Elu02, Goo84, Hec00, Hol00, Joh59b, KE23, Mal97, Mal40b, Mye93, Ros96a, Sat82, Sch84]. Rites [Boy02]. Ritual [Ack75]. Rival [Nob65]. Road [Gil86]. Robert [Hes18, Bar73, Bas13, Can73, Dav94, Dew59, Gre62, Gre91c, Hey79, Hor80, Les07, Mil54, Mor16, Mos60, Mul94, Pri95, Rau88, Smi78]. Roberton [Ces05]. Robertson [WB93, Har02a, Kon97, Wom86]. Robot [Cam70]. Rochefoucauld [Ano83f, Bak83]. Rochelle [Sou70]. Rodgers [OH88]. Roederer [Rap21]. Roger [Sta08b]. Roland [Mur54]. Role [Can64, Cre79, Gui13, Han59, Hel73, Jaf80, Met96, Ven76, Ver22]. Romae [Sch23b]. Roman [Cam22a, Bri23, Had43, Had50, Lee09, Lee11, Mon97, Oak55, Sal89]. Romance [Pri95, Ran23]. Romanian [Kel98]. Romano [Dra90]. Romans [Ger25]. Romantic [And02, Bal44, Bar41, Duf95, Mat57, Pri95, Ros82, Wol53, Cas10, Tol98]. Romanticism [And41, Bri41, Bri79, Fai41, Koo99, Lan80, Lov41, Mil63, Mit96, Mor96, New95, Nor02, Pou54, Sou70, Tho75, Neu67]. Romantics [Bar42]. Rome [Cal97, Cer05, Cra45, D’E07, McA99, Pra65, Sch23b, Val09, Ran23]. Romeo [Kei57]. Roosevelt [Bur65]. Roosevelts [Cro53b]. Roots [Ber95, Cre79, Fat05, Hel64, Kae96, Lac10, Ols91, Sea11, Tie11]. Rosanvallon [Moy07, FR07]. Rosen [Dra78, Rat61, Wie86]. Rosenblatt [Isr16b]. Rosmini [Bla57]. Rothbard [Jen22]. Rothmann [Sta67b]. Rotterdam’ [Die61]. Rousseau [Arm07, Bar83, Blo79, Cam84, Del62, Dou11, Ech72, Eck44, Gei91, Hol17, Isr16b, Jac78, Kel68, Lev67, McA10, McK82a, McN45, Mit93, Mor93, Neu47, Nol67, Noo72, Pau71, Ros16b, Sco98, Sim04, Wil05, You01, Eck65]. Routes [Bra02]. Rowan [Kar65]. Rowland [Har72]. Roy [Par24]. Royal [Poo05, Bat19, But15, For84, Joh40, Mie69]. Rubens [Wil91]. Rudbeck [For23]. Rüdiger [MF06]. Rudolf [Ang00]. Rudyard [Var53]. Ruines [Pea52]. Rule [Jac12, Smi70]. Rulership [Fac21]. Rural [Joh44, MR20]. Rush [Cha23b, D’E69, Was72]. Ruskin [She77]. Russell [Boa47, Rem79, Woo94]. Russia [Avr66, Bar58b, Ges65, Las47, Lew58, Mar68, Thu76, Vov12]. Russian [Che59, Erl73, Gar70, McC63, Sil01, Vuc60, Wil70a]. Russians [Eat80].

Rustic [Iva24]. Ruth [Str22].

S [Bee20, Bel15, Bre72, Bur59, Eis57, Hab90, Hab93, Hai50, Ina20, Par62, Ran66a, Shi45, Smi80a, Sul83]. S. [All85, Tie11]. Sabians [Elu02]. Sac [CH95]. Sacer [Gre23b]. Sack [Blu00]. Sacra [Ell67]. Sacred [Cla92, Gor08, O’B85, Sch46a]. Sacrifice [She06a]. Sage [Ban16, Kim14]. Sailor [Lub24]. Saint [Ano56e, Han18, Bat00a, Car74, Hai46, Har66, Len02, Lov42, Men18, Mor04a, Pol43, Rid62, Shi45, Sim56, Smi73, Whi60, Per57]. Saint-Evremond [Len02, Mor04a]. Saint-Gelais [Rid62]. Saint-Martin [Bat00a]. Saint-Pierre [Men18, Per57]. Saint-Simon [Ano56e, Pol43, Sim56]. Saint-Simonian [Shi45]. Saint-Simonians [Car74]. Sainte [Nel00]. Sainte-Beuve [Nel00]. Salmasius [Gra63a]. Salomon [Zet82]. Salonnières [Pek99]. Salt [Nor74]. Salutati [Wit77]. Salvation [Mil43, Pat67, Tur66]. Salvatore [Bul05, Reg05]. Salvatus [Chr45]. Salvemini [Puz59]. Salwyn [Coa50b]. Samuel [Pri58, Ebe75, Fer71, Gay63, Hud56, Lin48, Pop53b, Rom60, Ste81, Stu89, Wig09, Woo88]. Sanctis [Del80]. Sanders [CH95]. Santayana [Lev87a]. Santillán [Chr23]. Sapientia [MH04]. Sartre [Lev72a, Aro87, Bia24, Doe22, Ful94, MS61]. Sartreanism [Ful94]. Satan [Bli57, Pat67, Rum85]. Satire [Gui85, Hun86]. Saturn [Wen24]. Saturnalia [Not11]. Satyagraha [Mur68]. Savage [Bid54, SR81, Wri43, You01]. Savages [Day08]. Saved [Åsb21]. Saving [Del17]. Saviors [Kap56]. Savonarolan [Col99]. Saw [Ras82]. Saxo [Mal58]. Saxoferrato [Ces04]. Saxon [Kli55]. Saxonism [Hor76]. Saxons [Van58]. Saxonum [Whi71]. Scabies [Kei57]. Scale [Lin91]. Scaliger [Mar04, Pri17b]. Scandal [Ban14]. Scarcity [LeM79]. Sceptic [Bra64, Gre43]. Scepticism [But59, Car62]. Sceptics [Pri64a]. Schapiro [Dyn81, Coa50b, Ros96b]. Schelling [Gre15, Kra18, Pur15]. Schilder [Mie73]. Schilder-Const [Mie73]. Schiller [Bar88, Gai00, Hof85, Roe03, Wei52]. Schlegel [TA06]. Schleiermacher [Pfa90, Pur15]. Schmalkaldic [Bat24]. Schmitt [Kel04, Mag16, Sme20]. Schmohl [Mer49c]. Schneider [Fis47]. Scholar [Ber67, Cur42]. Scholarly [Ano42f]. Scholars [Ste40, Sto85]. Scholarship [Afr93, Bar59a, Coa61, Elu02, For23, Gra44, Hut92, Mar11, Mil01, Not11, Sap79, Twi20, Whi71, Zah12, ZMM10, Mod94]. Scholastic [Bla19, Dan81, Iva24, Pet15]. Scholastica [Voz25]. Scholasticism [Bre72, Byr23, Cam22b, Koh17]. School [Gut41, Lev87b, Pet15, Ran40, Lev13, San62, Tre67]. Schoolmaster [Sta08b]. Schoolmen [Ste49]. Schopenhauer [Car84, Got75, Gup75, Hab95, Lau90, Van01]. Schumpeter [Kru55]. Science [Ada49, Aga73, Ala77, Ano45f, Aus70, Bay65, Bed21, Ber86a, BG92, Boc51, Bow85, Bur59, But15, Cad95, Can61, Can64, Car07, Car74, Cha24, Coh54, Cor80, Cre83, DH71, DeV08, Dra70, Dur43, Eas67, Ell57, Eng08, For84, Fri73, Gar55, Ger25, Gom43, Gou06a, Gre69a, Haa48, Han67, Har95, Hec00, HH66, Hel19, Hes18, HP18, Igg59, Jac76b, Jon40, Kar69, Kel84, Koc57, Lal54, Lea82, LS17, Lew72, Mar59a, McK75, Min65, Min51, Mul63, MF06, Nad93, Nel72, Neu77, Oak98a, Ong53, Par44, Pau72, Pri54, Ram19, Ran53, Ren76, Ros12, Sch80a, Sen58, Sho85, Sho24, Skj22, Sol78, Son74, Sou89, Sta09b, Sta08a, Sta87, Ste75, Sto62, Tat65, Tho51, Tho55, Tru12, Tum96, Van82, Vid95, Vir86, Wer93, Wie45, Wil88, Win70, Zak24]. Science [Zon14, dS23, Ano00f, Cla86, Eng08, HV02, Hol00, Ned91, Pea15, Smi19, Ede52, Jac82, Joh59a]. Sciences [Eng08, Gel52, Gou06b, McR57, Rap16, Ree52, Swe93, Swi75, HP18]. Scientific [All49, Chr89, Coh76, CF79, Cop91, Dia82, Dic80, GR86, Giv62, Gra62, Hal50, Hat78, Hol40, Hur56, Ker24, Man57a, McC94, McK47, McK66, Mol78, Not16, Osl70, Osl83, Pet94, Ran40, Ras07, RS64, Rub86, Str55, Syp65, Wes87, Wie40, Zil41, Zil45, Efr97, Pla13]. Scientism [Ree83]. Scientist [Eis57, Het81, Sti09]. Scientists [Gre53]. Scope [Bro81]. Scot [Dav13]. Scotism [Cam22a, Sch22b]. Scotist [Cam16]. Scotland [Jen20]. Scott [Gar75, Mar19, Phi89]. Scottish [Bow85, Bra79, Gol88, Man88, Man91, Ols71, Pla10, Pla13, Sch22a, SR81]. Scotus [Rum85]. Scripture [Bla93, Moo76, Nad13b, Nad13a]. Scrutiny [Woo19]. Sculpture [Lar41]. Scythian [Joh59b]. Seagrave [Tie11]. Seal [Fog12, Mos73]. Search [Bro53, Cha58a, Ger25, Ger13, Kar84, Nad91]. Searcher [Maz14]. Searching [Har03, San20]. Searchlight [Har03]. Seas [Vie03]. Seascape [Vog93]. Seaton [Ber80]. Second [Bur87, Her48, Leb86, Mar59b, Mor93, Row56b, Sim90, Smi70, Smi80b, Wei19a]. Secrets [Ben57, Mer08, Pet01, Pri17b]. Sect [Sho85]. Secular [Alj19, Cri86, Gor08, O’B85, Sch46a]. Secularism [Alj19, Gra65]. Secularization [Gor19b, Hic18, dG80, Bar60a]. Security [Str56]. Seduction [Tri04]. Seeing [Kim24a]. Seeley [Hes18]. Seen [Dia82, Hou41]. Seer [Ruk18]. Selden [Vie03, Wal11]. Selected [Ano03f]. Selecting [Car07]. Selection [Alt08, Rus75, Vor69]. Self [Dac04, Fri54, Hun86, Kar84, Lam74, Liv08, Loc07, MB95, Mit93, Ran70, Saa10, Sam43, Sor83, Whe55]. Self-Determination [Ran70]. Self-Development [Ran70]. Self-Disclosure [Hun86]. Self-Esteem [Saa10]. Self-Formation [Sor83]. Self-Interest [Lam74]. Self-Love [Sam43]. Self-Made [Liv08]. Self-Sufficiency [Whe55]. Selma [Ano03h]. Selz [Har03]. Semantics [Kva77, Pet12]. Semitic [Deu45]. Seneca [Bar98, Sal89, Wal66]. Senecan [And04]. Sensationalist [Ric79]. Sense [Fra55, Gre97]. Senses [Bro92]. Sensibilities [Chi84]. Sensibility [Dav43, Ham08, Ste88]. Sensible [Mor78a]. Sensitive [Ham08]. Sensorium [Ins11, Bas13]. Sensus [Lie56]. Sentiments [Cam20, DB77, Gol88, Pri17a]. Separation [Bar81, Rah22, Rik54]. Sepmaine [Gar13]. Sequel [Mer49a]. Serge [SJB18]. Sergeant [Yol51]. Seriously [Bra24]. Sermons [Van82]. Service [Bur93, MR20, Tho07]. Servile [Tre21]. Servitude [Woo18]. Settecento [Coc58]. Settler [Leg25]. Seven [Bro47]. Seventeenth [Ble58, Cog07, Eme84, Ger72, Gre14, Jac76b, Jac82, Mar22, Oak98a, Sch40, All49, Ari92, Bur03, Cam16, Cla82, Dan81, Dav66, Fin47a, Fix89, Goo02, Gue89, Gun68, Har98, Her92, Hou41, Hou42a, Hou42b, Jay70, Kri60, Mas77, McC79, McR65, MM89, Mil43, Mul06, cN93, Nor77, Oak98b, Ogd49, Rid56, Ros96a, Sch23a, Sch04, Sch80b, Sch22b, Sin89, Sou89, The23, Tho51, Tho55, Tri91, Wes11, Yol55, dG80]. Seventeenth- [cN93]. Seventeenth-Century [Eme84, Jac82, Bur03, Cla82, Dan81, Dav66, Fin47a, Fix89, Goo02, Gue89, Gun68, Har98, Her92, Hou41, Mas77, MM89, Mil43, Mul06, Nor77, Oak98b, Rid56, Ros96a, Sch04, Sch80b, Sch22b, Sin89, The23, Tho51, Tri91, Wes11, Yol55]. Seventy [Ano48a]. Seventy-Five [Ano48a]. Seville [Wun03]. Sex [Ami07, DeV08, Fre64, Mai21, McL76b, Nic90, vdL17, Sim90]. Sex-Crime [Fre64]. Sexism [Sim90]. Sextus [Pop93, Flo95, Pop53b]. Sexual [Nad82, Smi78]. Seyssel [Boo00]. Shōwa [Sho10]. Shackles [Ros77]. Shadow [Bat24, Bro87, Ter21]. Shadows [CM05]. Shadworth [Sta09b]. Shaftesbury [Ans78, Dus81, Mor94, Tow87]. Shakespeare [Cru51, Elt97, Qui65, Smi76b, Smi80b, Lev71]. Shall [Cas81]. Shanghai [Fog10]. Shaping [Pri95]. Shaw [Irv47]. Shedd [Str54]. Shelley [Woo60]. Shemuel [Ste18]. Sheridan [Skj22]. Sheriff [Tur75]. Shifting [Rey53]. Shils [Ran41]. Ship [Alt08]. Shipwreck [Leg25]. Shook [HV02]. Shores [Leg25]. Short [Ger72, Mor04b]. Short-Title [Ger72]. Should [Ang00, Dav94]. Shue [Lic73]. shuo [Wan05]. Shut [Cor18]. Sibilla [Dra90]. Sibylline [Bet09]. Sides [Sch22b]. Sidgwick [Mul63]. Sidney [Con82, Rib52, Rik54]. Siècle [Mor40, Spe98, Fal54, Glu13]. Sieyès [Bro24]. Sigmund [Dia82, Ful83]. Sign [Lif12]. Significance [Bou97, Bou54, Goo11, Gre15, Isr19, Kri46, Spi52, Vir50, Wie46b, Wri40]. Signification [Mor40]. Significs [Pie09]. Signs [Arm65, Rud74, Sei69, Wes01]. Sigwart [Lev47b]. Silence [Maz62, Sei65]. Silver [Mir77b]. Similarities [Syp65]. Simile [Bow56]. Similitude [Lun04]. Simmel [Sma91]. Simon [Ano56e, Gar13, Pol43, Sim56]. Simonian [Shi45]. Simonians [Car74, Gol82b, Hai46]. Simonides [Tha75]. Simplicity [Hav53, Zon14]. Simulating [Ger13]. Sin [Fry54, Har02b, Ned88]. Sinai [Bat24]. Sinarum [Lun83]. Since [MB81, Gee76]. Singing [Sim04]. Singularity [Hef78]. Sinica [Ril99]. sinicus [Wes07]. Sinism [Cre49, Pri48]. Sino [Sel13]. Sinophilism [Lac53, Men56]. Sir [Bal50, Boc76, Bow74, Gru73, Hun53, Orr14, Rib52, Ros68, Wil48]. Sismonde [Urb12]. Sismondi [Pap79, Urb12]. Sit [Chr45]. Six [Gri21, Kle99, Nic19]. Sixteenth [Ger72, Bea82, Ble58, Bro05, Cha60, Dav60b, DeM74, Eme84, Gre14, Jun53, Mar22, Moy03a, Neh77, Not11, Oak98b, Oak98a, RH79, Ric12, Smi54]. Sixteenth- [Eme84, Oak98b]. Sixteenth-Century [Bea82, Bro05, Cha60, Dav60b, DeM74, Jun53, Moy03a, Not11, Ric12]. Skeptic [vE01]. Skeptical [Hal02, Tay17]. Skepticism [Dei54, Der10, Gru73, Mai97, Mis83, Mul97, Nau57, O’C84, Ols91, Rus88, Sch84, Sch91a, Sou92, Wac96]. Skeptics [Mau05]. Sketch [Ber18, Lat57]. Sketches [Jur02]. Skinner [Lan12a, PB73, Pet12]. Sky [Bro05]. Slave [Gla90, Hef73]. Slavery [Cuf66, Hun78]. Sleepwalking [Han17]. Slogan [Sch53]. Slovenia [Rob20b]. Smart [Gre53, She57]. Smith [Sav59, Whi76a, Ber74b, Bro92, Cam20, CR81a, Coh40, DB77, Gar54, Har11, Het83, Hey17, Hud56, Lam74, Lan77, Lev78, Lom02, Mac49, McN73, Meg75, Mos60, Mye75, Nie86, NSR80, Pac02, Pri17a, Rap69, Ski76, SRN84, Was84, WB93]. Smithsonian [Eis57]. Smyth [BI60]. Sociability [Fef90, Haa16]. Social [Ada49, Ada89, Bed21, Bos73, Bur59, Bur65, Cas92, Cas47, Cla00, Coa50b, Daw93, Ech72, Fef90, Fre73, Gal94, Goo64, Got80, Gre74, Ha20, Hay59, Hel64, Igg59, Jac82, Kap56, Lea77, Lev67, Lew72, Lid66, McC20, McK82a, Mit96, Mom50, Moo45, Mor90, Mos85, Mul79, Nis44, Par44, Par62, Pog63, Ran41, Ril73, Rog72, San82, Sen58, Sim60, Sta91, Sta87, Sto62, Str76b, Stu97, Van17, Ver82, Wac46, Wal84b, Wal86, Wei93, Whi47, Wol44, Woo92, Zee46, Ber77, Cla86, Whi45a]. Sociale} [McC94]. Socialism [And02, Bak76, Bal42, Feu49, Fro07, Gol82b, Hyd56, McL76b, Mil63, Mos61, Rob20b, Rob15, Web22, Cla86, Von48]. Socialist [Bes48, Bri40, Cal22, New95, Nol61, Ste51]. Société [McG03]. Society [Ano60c, Bar58b, Beh08, Bla13, Car74, Cla92, Cri86, Fis82, For74, Gal17, Hor80, Imm09, Lee79, Lev78, Loc07, Mai21, Mit71b, Ned88, Ran41, Rap69, Rig86, Rot58, Sim55, Sos73, Ven76, But15, For84, Joh40, Poo99, Poo05, Shi45]. Socinian [Tin96]. Socinianism [Jol78, Wal84a]. Socinus [Mor09]. Socio [Fre64]. Socio-Historical [Fre64]. Sociological [Rum65, Wei50]. Sociology [Ada16, De 41, Hil01, Lev65, Men61, Par44, WB93]. Socrates [Ans78, Ber86a, BS83, Gui85, Joh44, Kau48, Lev56, Sar83]. Socratic [Mul68a, All60, Ans78, Gui85, Hab95, Hat68a, Ina20, Mul74, Wel76]. Soldier [Lub24]. Solid [Ros85b]. Solitary [You01]. Solmsen [Wel73]. Solomon [Atl52, Hat68b]. Solution [Yeo03, Car08]. Solving [Kat84]. Some [Ada42, CJK^{+}43, Chi84, Coh54, Coh80, Dor40, Fre73, Gra03, Gra86, GS91, Han95, Hun60, KM94, Kap56, Lac65, Mac78, Mad54, Maz72, Mil06, Moo45, Mun79, Pac79, Par88a, Rat61, RMG81, Rin41, Rom64, Ros43b, Sch43c, Sch80b, Smi62b, Sol83, Sou89, Sto61, Wie61, Wil59, Wil60, Wil41, Wil74a, YY84, Cog07, Dup82, Kas77, Tor84]. Something [Bri53]. Somniorum [Cor18]. Somnium [Nic40, CM05]. Son [Loe70]. Sonnet [Llo69, Ruo66]. Sophie [Ber18, Kim24b]. Sophismata [Tan98]. Sorbière [Pop53b]. Sorbonne [Tho55]. Sorel [Tag86, Ver73]. Søren [Kja02, Nag68]. Soul [Ada01, Bas13, Bla61, Del17, Mar76, Nan21, Osm73, Sga15, Sol83, Wig09, Wil70a, Wil20]. Souls [Bro18]. Sound [Tay55]. Source [Ans99, Cor75a, Goo90a, Gro99, Woo67]. Source-Reader [Gro99]. Sources [Ada89, Bac07, Bar97, GS91, Het73, Lau67, Pip59, Pop93, San59, Von84, Wal84a]. South [Hay59]. Southern [AU19]. Sovereignty [Jez71, Joh85, Nic19, Par01, Tun14, Vel19, Zhu21]. Soviet [Bea83, Jor59, Mik70]. Sozzini [Tin96]. Space [Bal60, Gou61, Her87, Nor46, NS16, PB43, Pow70, Rab62, TW21]. Space-Again [Rab62]. Spain [FS80]. Spalding [Pri13a]. Spanish [Ble58, Cor75a, Ent18, Kei04, Mil63, Mor18, Mor54, Ort18, Wes15]. Spark [Gre91a]. Spartan [Afr60, Ran67]. Spatial [Ogd74, Wit09]. Spaventa [Gri41]. Speak [Cha17, Fri12]. Speaking [Tri96]. Special [GSJ18]. Species [Cau94, Fle59]. Specific [LS91]. Spectator [Cam20]. Specter [Roy07]. Spectre [Ham53, Ras82]. Speculum [Zah12]. Speech [Ber78b, Cra45, Ful83, Low24, Qua19, Sea80]. Spence [Par73]. Spencer [Ada16, Bak77, Bow75b, Fra78, Pau83, Sch45b, Sim60]. Spengler [Far91]. Spenser [Gre74, Hug43]. Sphere [Lin10]. Spheres [Nie86, Ros85b]. Spider [Bow56, Wil71]. Spinning [Men20]. Spinoza [Kas77, Bat77, Col59, Eck44, Ell03, Fra13, Gei91, Goo18, Har13, Lau00, Met60, Mor80, Nad09, Nad13b, Nad13a, Nad19, Par77, Pit86, Pre89, Rud01, Sac76, Wes07, Wol53, vB17]. Spirit [Bec69, Ber71, Gre62, Nic90, Pag96, Ruk18, Til45]. Spirit-Seer [Ruk18]. Spirits [Fix89, Ruk18, vR00]. Spiritual [For05, Mor25, Sch80b]. Spirituality [Mar23, Sho24]. Spitzer [Lov44]. Splitting [Tin88]. Spokesman [Mer48]. Spontaneous [Goo02]. Spread [Gat67]. Sprigge [Sch41]. Spunge [Tan77]. Spy [Lub24]. Squares [Pet88]. Squaring [Bir96]. St [Fer68, Wie48]. St. [Bar75, Dom88, Gol82b, Maz62, Mer53, Mom51, Smi41, Zal85, vG20]. St.-Simonians [Gol82b]. Staatstheorie [Fri73]. Staël [Hil11]. Stages [Blo79, Rei21, Sto61, Wel69]. Staging [CW13]. Stalnaker [Isr16b]. Standard [Ger13, Iva24, Tho59]. Stands [Kja02]. Stanhope [Hud56]. Stanley [Pau79]. Stannard [Cha58a]. Starkey [May88]. Started [Mae80]. Startled [Ras82]. State [All89, Arm00, Ben57, Bou94, Bra71, Coh94, De 79, Dym21, FS80, Ger25, Gil20, Hui17, Jac78, Kri63, Lun87, Mac71a, Mai21, Mar11, McA10, Mor86, Oli74, Tol98, Ver84, Woo88, All85, Blo46, Gri21, Tre21]. Statecraft [Cor75a, FS80]. States [Bla13, Han17, Imm09, Kri63, Lew65, Mad45]. Statesman [Lan12b]. Stating [Mul68a]. Statistics [Arm07]. Statistik [vdZ10]. Status [Cum73, MM22, Wal84b]. Stefan [Cai07]. Stefano [D’E07]. Stein [Han18, Men61]. Steinthal [Kal87]. Stellar [Het73]. Stereotype [Cla23]. Sterne [Mac49]. Steuco [Ano67g, Del96, Sch66]. Stewart [Nor01, Ras85, Wal19]. Stierle [Cor98]. Stillman [Fra77, Ros57a]. Stirner [Ste74, Ste78a, Ste85a]. Stockholm [Hol21]. Stoic [And08, Gas00, Gou74, Hor74, Tza91, Was84]. Stoicism [Dev70, Fro95, Sal89]. Stoics [Sei83]. Stone [Goo11, Pat86, Sil93]. Stones [Tou13]. Stop [Eng08]. Stopes [Doa17]. Storia [Boa56]. Stories [Ham98]. Story [Goo90a, Nar05, Sch41, Wei19a]. Strands [CH15]. Strange [Cor18, Smi65b]. Strategic [Tho24]. Strategies [Bla03, EMR13]. Strategy [Bec67]. Stratigraphers [Cha23b]. Strauss [Gra81b, Alt07, Hol17, Kee12, Lin10, Tay17]. Street [The23]. Stress [Pet85]. Strife [Cho15]. Strong [Pac79]. Structure [Gil53, Hum72, Ste60, Thi92, Ran41]. Structures [SB83]. Struensee [Lau00]. Struggle [Bow76]. Stuart [Ada92, Car62, Feu49, Hai46, Jon92a, Lea82, Mil61, Ran65, Rot90, Sch43a, Str55]. Student [Gil86]. Studia [Ros74]. Studies [All49, Ano03f, Bro04a, Chi84, Fog10, Gee76, Han95, Jan08, Liu04, Mun79, Ran41, Sch43c, Tri56, Hor78]. Study [Boa70, Car62, Cas42a, Cas42b, Coh40, Cum73, DH71, Fer71, Gus50, Had79, Hay56, Hol53, JN12, KR41, Kri51b, Kri52, Lau44, Les57, Mac70, Pax93, Rus93, San44, San20, Sch56a, Tad22, Wil70a, Wri60]. Stuff [Wes11]. Style [Gui23, Gui08, MM22, Pfa90, Pri95]. Suárez [Mor53]. Sub [Yos51]. Sub-Microscopic [Yos51]. Subject [Ger72, Lac99, Sei90]. Subjection [Woo18]. Subjects [Ham98, Mer90]. Sublim’d [Lin91]. Sublime [Duf95, Kra18, Rya01, Sar08]. Sublimity [Fol17]. Substantialism [Mon19]. Substratum [TC47]. Success [Ioz22]. Suchon [Wil19]. Sudhoff [Wen24]. Sufficiency [Whe55]. Sufficient [Rap81]. Suggestion [YY84]. Suggestions [Fer51]. Suicide [Cro52, Noo78, Sac72, Sei83]. Sukkot [Not11]. Sully [Rya09]. Sulphur [Rou67]. Sulzer [Nan21]. Summa [Mer53]. Summer [Boc76, Mil41]. Sumner [Not55]. Superhuman [Her21]. Superintendence [Dod08]. Superior [Fal54]. Supernatural [Ban16, Tun14]. Supernature [Cam22b]. Superstition [Gui23, Tho55]. Superstitious [Day08]. Superstructure [Tod74]. Supplementary [Gil69]. Supposed [Ber69, Ors69]. Supranationalism [Vel19]. Supremacy [Che56, Van59]. Supreme [Wan05]. Surface [Wil88]. Surtz [Fer68]. Survival [Ban70, Cla00, Had50]. Survivals [Daw93]. Sweden [Skj22]. Swedenborg [Gar84]. Swift [Can73, Ehr52, Fry54, LaC70, Mer66, Mur54, Pre54, Sta59, Tin88, Tuv50, Wit43]. Sylvia [Sch41]. Sylvius [Wat72]. Symbol [Kiv67, Neu47, Thi79, Wil88]. Symbolic [Mos73, Pag96, Ver69]. Symbolism [Sto51]. Symbols [Cas81, Fry54, Min65]. Symonds [Rut14]. Sympathy [Fie76, Lam74, Lan09, Mac49, Kim24b]. Symposium [Bar62, Ede46, Lev79, Pal18b, Wol16a, Wol16b, Bev17]. Synagogue [Cul70]. Syncretism [Wil51]. Syncretist [Joh71b]. Synderesis [Gre91a, Gre91b, Gre97]. Synesiorum [Cor18]. Syntax [Lan76]. Synthesis [Ecc76, Fer51, Hor74, Mue58]. Synthetic [Ste89]. systêmes [Sch10]. System [Ada42, Bla90, Dus81, Gey48, Gra78, Het73, Kri51b, Kri52, Lam74, Lie98, Man57b, Mar59b, Pap79, Rid62, Sch10, Tat63b, Rap21]. Systems [Doo90].

T [All85, Bet59, Cas10, Gri74, OH88, Str54, Wei96]. T. [And11, Bre72, Hab90, Hab93, Koh64a, Ran66a, Sim55, Sta57]. Tables [Ste82]. Tabula [Eng80, Woo92]. Tacit [HM79]. Tacitus [Sal89, Sol00]. Taijitu [Wan05]. Taine [Wol44]. Takes [Cla23]. Taking [Bra24]. Talcott [Isa10]. Tale [Cop91]. Tales [Fis58]. Talking [Tru12]. Tallien [Rap21]. Taming [Sus12]. Tanaka [Nol84]. Tao [Chi74]. Tao-T’Ung [Chi74]. Taoist [Sha88, TC47]. Tarkiewiez [Rie61]. Task [Hor97]. Taste [Boa40, Gur00, Ham08]. Tatarkiewicz’ [Bea76]. Taubes [Sty15]. Tautology [Rob60]. Taxonomy [And76]. Tayler [Bla52]. Taylor [Lin48, Pri58, Chr59, Deg23, Gor08, Har85b, Van17]. Teacher [Nic48, Ros57b]. Teachers [Fef05]. Teaching [DeM74, Rob59]. Technocracy [Car74]. Technological [Oti02]. Technologies [Whi09]. Technology [McG80, Neu77, Sch10, Sme20]. Technometria [Spr68]. Technometry [Gib72]. Teilhard [Sho24]. Teleology [May92, Par65]. Telescope [Bro85]. Television [Mat06]. Temple [Cul70, Pri64b]. Temporis [Doo99]. Tendencies [Bal44, Fer54, McR55]. Tennyson [Ada79]. Tension [cN93]. Term [Bro91, Bur59, Igg59, Igg95, Kem00, McR65, Sen58]. Terminology [Moo79, Str76a]. Terminus [Mil97, Ser17]. Terms [Des67, Fac21, Wal86]. Terror [Doe22, Rap21]. Terrorism [Rap21, Gri21]. Tertullian [Lie56]. Testament [Sol70, Sos73, Wel69]. Tetralogy [Smi80b]. Text [Bar75, Mos48, Pfa90, Rap69]. Textbook [Rei69]. Texts [Bou03]. Textual [Hau98]. Texture [Sil20]. Thalheimer [Kit73]. Thao [Her05]. Theater [Dem72, Zhu21]. Theatre [Cam20]. Theatrica [How67, Tat65]. Theatricality [Mel75]. Theft [KS24]. Their [Cum10, Eck44, Gar13, HG21, Kee92, Mat55, Ste82, Str55]. Theism [Hur56]. Theismus [Cro09]. Theistic [Alt08]. Theme [Lif12, Loe70, Ran67]. Themes [Gol82b, Hor74]. Thémiste [Hic10]. Theodicy [Bar81, Fis56, Men18, Nad94, San82, Sty19]. Theodor [Wil23]. Theodore [Bur65, Tho07]. Theologia [Jac85]. Theologian [Dam79, Hal02, Jes86]. Theologians [Cas78]. Theological [Gue83, Mal97, Mil97, Moy03b, Nah47, Osl85, Sac18, Soc06, Sou00]. Theologico [Nad13a]. Theologico-Politicus [Nad13a]. Theologus [Kal95]. Theology [Bow74, Bur44, Cre95, Dau77, Gas00, Hey79, LeM79, McK75, Oak68, Oak98b, Per57, Poo05, Ram19, Reg05, Rob80, Rub06, Sch45b, Sho99, Sho18, Spr68, Sta08b, Ton59, Zon14]. Theophrastus [Mod94]. Theoretical [Gri12, Hel19, Jor52, Par44, Rom48]. Theories [Bur11, Clo69, Clo71, Eck44, Flo46, Jen20, Mar22, Mul79, Rat86, Ric79, Spi55a, Wal11, Wal86, Zuc75]. Theorist [Arr94, Fre73, Isa10]. Theorists [Bok17]. Theory [Ada42, All66, Ari92, Arm56, Bal60, Bat75, Bis15, Bre10, Cap54, Mil87, Car91, Che59, Chr89, Cla23, Coh78, Coh94, CH95, Cro10, Dav95, Daw93, Doo90, Dra75, DB77, Eck65, Ell57, Eva55, Fas97, Fei46, Fra03, Fra55, Fra77, Fur81, Gai00, Gal17, Gli88, Goo64, Gra88, Gra89, Gub05, Gue95, Gue86, Gun71, Har03, Har85b, Hey82, Hob80, Hof85, Hol17, Hoo73, Hsu02, Hut76, Isa10, Jen44, Jon92b, JN12, Kar84, Kel04, Kit73, Koh64b, KE23, Kor75, Kri64, Kri40, Lan76, Lau90, Lea77, Low74, Lyn82, Mac78, Man48, Man55, Man57a, McC42, McK81a, Meg75, MM90, Mil61, Moo45, Mor92, Mor54, Nah47, New01, Noo72, Nor46, NSR80, Pfa90, Pri17a, Rif53, Rob80, Ros70, Rud74, Sat82, Sch73a, Sch73b, Sch64, Sco98, Sho16, Smi76a]. Theory [Sor83, Spi55b, SRN84, Tay01, Ten82, Tie54, Vin07, Von58, Vor69, Wal96, Wal19, Wea80, Wil41, Win81, Wol50, dH45, Tri04]. Theosophy [Sta74]. Therapeutic [Sch04]. Therapy [Coo99]. There [Sar82, Ger73, Gue86, Pre77]. Theriophily [Gil69]. Thermodynamics [Bur87]. Thersites [Stu04]. Theses [Cro71]. Thesis [Cum78, Fen86, Gra81a, Han95, Kes68, LS17, Mue58, Poc87]. Thesis-Antithesis-Synthesis [Mue58]. Thing [Coe16, Plo53]. Things [Cha23a, Dou11, Kel12, Sha02]. Thinking [Ach20, Ben23, Dun21a, EST66, Lær14, Lan12a, Lo17, Mue43, Rob23, Sel58, She06b, She02, Tay01, Web22, YY84]. Third [Bra71, Coh87, vE01, Dav10]. Thirteenth [Dal88]. Thirty [Gri21]. Thomas [Fer68, Gru73, Hun53, Wil48, Ano77e, Ano88e, Bla57, Bly86, But06, Cal62, Col09, Cot81, DeA87, DB77, Hut78, Joh42, Jor85, Kar66, Kiv70, Lev54, Mac70, Man94, Mar09, May88, McK75, McL76a, Mer53, Mor04a, NSR80, Pal72, Par73, Pro72, Ran23, Red97, San60, Sar85, Sch45a, SRN84, Val04, Win70, Zag99, vG20]. Thomasius [Bar71, Hun00]. Thomist [Zah12]. Thomistic [Bre72]. Thoreau [Bay65]. Thorndike [Joh59a]. Thought [Ach20, Ash78, Ash79, Avi61, Bar98, Bar92, Beh08, Ber14, Ber78b, Bos73, Bur57b, Byn84, Cam84, Can64, Car62, Cas92, Col16, Con82, CR81b, Cou87, Cra92, CJ60, Ede22, Fat05, Fer54, Fer71, Feu59, Fix89, Fre73, Gas55, Gay63, Goo64, Gra62, Gue83, Guy20, Har98, Hei78, Hey17, Hin92, Hob71, Hou41, Hum72, Jen07a, Kam70, Kim24b, Kli24, Koh17, Koy98, Kri46, Liz22, Man88, Man91, Mar19, Mas00, McK75, McR55, Met96, Moh04, Mun80b, Nak22, Nau57, Nau15, Ned88, Ned96, Ned99, Ned05, Ned13, cN93, Nol61, Nur61, Obe87, Pat59, Pet88, Pod02, Raj52, Ran66a, Red74, Row56a, Row56b, Run00, San82, Sch04, Sea80, Sho99, Sim04, Sim47, Sin89, Sol63, Spi76, Sta91, Sut03, Tre21, Vir50, Wal70, Whi45a, Whi47, Wie46b, Wil60]. Thought [Wil70a, Woo18, Woo88, Yin22, Zal93, Zam98, Efr97, Coh40, Sch44, Hea87, Kas77, Kel79, Spi66]. Thoughts [Lal54, Bee20]. Thousand [Gri21, Mar17]. Thread [Mil63]. Three [Bri53, Cur42, Ger51, Gla90, Joh67a, MS15, Mos66, O’F92, Sil76, Sti98, Mor78b, Wie75]. Thresholds [Mon19]. throughout [Mar19]. Thucydides [Arc67, LS23, Sch45a]. Thy [Ros71b]. Tibet [EST66]. Tierney [Sea12]. Tilbury [Byr23]. Tillich [Cre95]. Time [Bal60, Bla82, Blu00, Mil87, Dal88, Dan81, Gro11, Gub05, Lun87, Pet85, Pop93, Qui65, TW21, Tza91, AP92]. Time-Sack [Blu00]. Timelessness [Pou54]. Times [Bal44, Bre10, But77, Mil70]. Timothy [Cor98, Har93]. Timpanaro [Dom83]. Tingyun [Liu12]. Title [Cum78, Ger72, Hic10, Ros43a]. Tocqueville [Ano86f, Boe83, Dre64, Kah85, Man08, Pap64, Par88a, Thu76, Wac46]. Tocquevilles [Man08]. Today [Rie63, Her05]. Together [Rap69]. Toland [Rap97, Bid76]. Tolerance [Str21]. Tolerantia [Bej97]. Toleration [But77, Gro79, Hal58, Hic10, O’B85, Sto85, Tin96]. Tolstoy [Ger13]. Tom [Chr73, Kat89]. Tommaso [Hea88]. Tonelli [Ger72]. Tönnies [Mit71b]. Too [Man08, Wal63]. Toothache [Hel19]. Topic [Ege66]. Topics [Bir62, Kap23, Nau15]. Topography [Lea78]. Topos [Gor00]. Tortured [Gre18]. Tory [Gre43, Mos41]. Total [And55]. totalitarian [Han18]. Totalitarianism [Jon92b]. Touching [Smi80b]. Toulmin [Por78]. Tour [Bou90, Pec96]. Tower [Moo51a, Moo51b]. Toynbee [Afr60, Afr62, Fie55, Gey48, Gey55]. Tract [Rap97]. Tractatus [Bat77, Nad13a, Puc05, Roz02, War11]. Tracts [Haw01]. Tracy [Ken79]. Trade [Arn94, Gla90, Hef73, Rem79]. Trades [Hou41]. Tradition [Ans78, Ber62, Bir62, Bor68, Cam16, Can73, Coc62, Dea74, Dur43, Ede98, Hob80, Hof41, Hut76, Hut92, Kap23, Kri83, Las47, Moh04, Ned88, Pat59, Puz64, Qui92, Rei69, Res63, Ric50, Ric79, Ril73, RS68, Sma83, Sou00, Ste53, Tin54, Wes15, Whi86, Wil66a, Woo19, Yeo96, Sar82, Gab56, San59].

Traditional [Wri49]. Traditionalism [Wil68]. Traditionalist [McC15, Ree83, Ros16b]. Traditions [Cel05, Lat24, Ram14, Sko64, Urq14]. Tragedy [Dav10, Gro50, Hau11, Lun87, Swe99]. Tragic [Bil11, Kra18]. Traherne [Jor85]. Traité [Sch10]. Traits [Blo79]. Tran [Her05]. trans [Ger25]. Transcendental [Cro09, Ins11, Ram19]. Transcendentalism [Cle69, Wil66b]. Transcendentalists [Cro59, Wil65]. Transcontinental [Rob20c]. Transcultural [Ant24]. Transformation [Ard63, Far91, Sol03a, Wal86]. Transformed [Cat10]. Transition [Din75, Fro95, How99, Mul23]. Transitional [Mil63]. Translating [How01, TW21]. Translation [Ebe69, Eig09, Fac21, Hel21b, HTW21, Hol21, Lan74, Pop53b, RR06, vdZ98]. Translations [Boo00, Ede22]. Transmission [Chi74, Cum78, Jan08, ME21]. Transmissions [Cel05]. Transnational [Bar16, Liu04]. Transposing [Les07]. Transubstantiation [Ben82a, Bou01, Nad88]. Transvaluation [dH45]. Trap [LS23]. Trauma [Ano75g, Fac75]. Travel [Hef73, Lar10]. Travellers [HM79]. Travels [Pir25]. Tread [Cha23b]. Treason [Chr54, Kel81, Smi54]. Treatise [Cot79, Kem00, Leb86, Mos51, RS64, Row56b, Rus85, Rus88, Rus93, Sch89b, Wal84b, Ioz22]. Treatises [Bro99, HM79]. Treatment [Bar98, Bar63, Pas75, Ran67, Sta57]. Treatments [Dan81]. Tree [Iva24, Lin91]. Tree-Trunk [Iva24]. Trembley [Var50]. Trembling [Kja02]. Trends [Ede52, Koy98, Man55, Zam98, Zan87]. Trial [Kri47]. Trials [Mor25, Smi54]. Tribute [Rom60]. Trinitarian [Ruo66]. Trinitatis [Kli81]. Trinity [Tri96]. Tripolitania [Ros24]. Troeltsch [Ran64]. Troilus [Elt97]. Trois [Wil66a]. Tropes [Fel04]. Troporum [Fel04]. Trotsky [Ber87]. Troubadour [Ros98]. Trouble [Vel19]. True [Bla52, Had79, Psa60, dC11, Smi17]. Trunk [Iva24]. Trust [Bou00, Dud57, Fle66]. Trust-Busting [Dud57]. Truth [AP92, Bat00a, Chi74, Cro53a, De 41, Fis82, Fra13, Fub96, Gra89, McC20, Nad91, Nad13a, Pin68, Sch96, Smi95, Tan99]. Truths [CJ60, OH88, Osl85, Wil67]. Tschirnhaus [Van93]. Tse [Tod74]. Tse-Tung [Tod74]. Tu [Sch51]. Tu-Hsiu [Sch51]. Tucker [Pop52, Ras82]. Tudor [Fer55, Pop05, Ber74a, Sie52, Tri50, Zee46]. Tumour [Sou92]. T’Ung [Chi74, Tod74]. Turbulent [McG03]. Turkey [Öza14]. Turn [Cot08b, Dav94, Lan12a, Pag88, Per08, Wal96, Wit09]. Turnbull [Nor75]. Turning [Koh69, O’B08]. Turns [BV88]. Tuscany [Coc62]. Tuzet [Boa67]. Twelfth [Ned91, Shy14, Tru12]. Twelfth-Century [Shy14, Tru12]. Twelve [Ste82]. Twentieth [Hyd56, Phi70, Söl16, Whi71, Ant24, Cha58b, Mar19, Mor20, Sel13, Smi24a, Smi24b, Whi47]. Twentieth-Century [Cha58b, Mor20, Sel13]. Twisse [Hut78]. ’Two [Sim65, Afr93, Ano86f, Bra71, Bra02, Bro99, Bur64, Cop91, Cro53b, Dre64, Fen86, Gos60, Gru73, Gug80, HM79, Kah85, Kem00, Lyn00, Mac88, Mar72, Poi70, Rah22, Ran64, Rap16, Ree83, Sco97, Sko64, Smi24b, Tip69, Was84, Wel60, Wen82, Zag99, OR13]. Two-Cultures [Mac88]. Tycho [Bla90]. Tylor [Daw93]. Type [Kim24a]. Types [Rah22, Sch14b, Wei52]. Tyranny [Sus12, Tri56]. Tyrant [Liz22, Zal93, Zhu21]. Tyre [Kos04]. Tyssot [Kno63].

U [Par62]. Ubermensch [AP92]. Uchronic [Ter21]. Uganda [Sum17]. Ulisse [Mar23]. Ulterior [Pri69]. Ultimate [Wan05]. Un-Benthamite [Can84]. Unapologetic [Str22]. Unbelief [Ber95]. Uncertainties [Str72]. Uncertainty [Sch14b]. Uncivil [Tun14]. Unconscious [Bas13, Gol82a]. Undemocratic [Sto46]. Underground [Bea42, Sca99]. Understand [Jum12]. Understanding [Bra24, Pat01, Ree52]. Unended [Pal97]. UNESCO [Ker24]. Unexamined [Ans99]. Unexplored [Ano87e, Oak87]. Unfinished [Bat77]. Unification [Val95]. Union [Bil11, Cro59]. Uniqueness [Day08]. United [Imm09, Kri63, Lew65, Mad45, Sho24]. Unity [McR57, Phi07, Wil08]. Universal [Cho15, Coh87, Gel52, Kno65, Maz54, Rad93]. universalis [Wes99]. Universalism [Cho15, Lan74]. Universals [Zuc75]. Universe [Asf99, Fur81, Kar84, Kri44, Mun51, Ran69, Yeo03]. Universities [All49, Sch22a, Voz25]. University [Ger25, Mai21, Bla22, Dyk62, Wal07]. Unjust [Bur93]. Unknown [Bla22, May88]. Unmasking [Mul06]. Unnatural [Cro59]. Unnoticed [Mos51]. Unorthodoxy [Val22]. Unpublished [Ano45o, Gos60, Kiv70, Met60, Mos60, O’F92, Rap69, Yol51]. Unremarked [All74]. Unruly [Wes77]. Unseen [Dia82]. Until [Fre73]. Untranslated [AA85]. Unusual [Boe83]. Unwin [Sta91]. Unworldly [Bri00]. Upelekwa [Rob20c]. Upon [Bar58b, Gro67, Spi76, Gre23b]. Uppsala [San62, Tre67]. Upright [Pat58]. Urban [Lat24, Lee79, Wei75]. Urbild [Koc21]. urbis [Sch23b]. Usable [Kli55]. Usages [RMG81]. Use [Bea82, Bri40, Cro71, Dah15, Dor40, For23, Rub91, Sch52, Sen58, Som98, Tri69]. Useful [Ric12, Smi17]. Uselessness [Lea80]. Uses [Ala67, Fon99, Gra03, Hol82, Igg59, Kap08, Kim24a, Leb86, Lev67, Min65, Qui10, Sol03b]. Usucaption [Bla13]. Usurped [Ger25]. Usury [Geo57, Tae42]. Utilitarian [Itz78, Lam09, Pau79, Spi61, Tay82]. Utilitarianism [Bee20, CR81a, Can84, Mac55, Wei96]. Utilitarians [Irv47]. Utopia [Ano56e, Gre65, Rin61, Ros24, Sim56, Ada49, Fer68]. Utopian [Gol82b, Neg79, Pol43, Sar82, Mil63]. Utopianism [Der10, Sme20]. Utrecht [Goo18, Tho24]. Utrum [Chr45].

V [Ano03h, Dew59, Smi76b, HG21, D’E07, Lin91]. Vacher [Hec00]. Vagaries [JW98]. Valentín [Wol76b]. Valerius [Mil97, Ser17]. Valla [Fub96, Bul05, Cam96, Cel05, Con96, Cop05, Del96, Mon89, Tri96]. Valori [Jur02]. Value [Bla19, Bro04a, Cro10, Ode65, RH79, Sat82, Val24]. Values [Jen13, Rut14, Sol78, dH45]. Vampire [Wil85]. Varchi [Sga15]. Variae [Sto85]. Variations [Ans09, Lif12]. Varieties [Kle99, Bee20]. Variety [Ogd49]. Vatican [Kos24]. Vaughan [Bur74, Lev67]. Veblen [Daw93]. Vedānta [Rol78]. Vegetarianism [Dom84]. Veils [Kel96]. Veiras [Keo74, Kno63]. Venery [Smi78]. Venetian [De 03]. Vengeance [Vis84]. Venice [Tri56, Car91, Pec96]. Venn [Dun21a, Wal07]. Verax [Pol64].

Veritas [Doo99]. Vermont [Dyk59]. Vernacular [Atk74, Gra88, Ric12, Rid62, Sga15]. Verse [Joh67b]. Version [Bro74, Wil74b]. Versus [Dav43, McN66, Roe85, Tri87, Wal86, Win81, Car08, Kra18, O’N07, Sou92, Sum17]. Vertical [Ban14]. Verticals [Ban14]. Vertus [Wil66a]. Verum [Min98, Mor78c]. Verum-factum [Min98]. Very [Kel84]. Vestibule [Ros71b]. Vestige [vE01]. Vestiges [Goo11]. Vestigia [Kli81]. Vesuvius [Tut22]. Vexed [Lan95]. Via [Mar90, Obe87, Wol60]. Vicar [And08]. Vices [Jac76a]. Vickers [Par08]. Vico [DH71, Had79, Hut76, Hut92, Lev91, Mar11, Min98, Mor78c, Mor80, Pre89, Sim81, Mor78b]. Victor [Sim65]. Victorian [Ala77, Can64, Car03, Dun21a, Hes18, Hou52, Kar69, OR13, Rut14, Sch15, Sta08a, WB93, Zon14]. Victorians [Kap74]. Victorino [Wol76b]. Victory [AJ81]. Vida [Kal95]. Vienna [Hac98, Joh71b, Sch54, Spe98]. Viereck [Bar42]. View [Bay65, Bro04a, Cla85, Dal77, Gra68, Har71, Hat68a, Jen07b, Joh71a, Mat57, Pog63, Sar85, Sho16, Smi83, Ste78b, Tae53, Tas69, Tza91, Wat92, Wes69]. Viewed [Fal54]. Views [Bri53, Bur65, Gar70, Gra78, Jor52, Lau67, Lev54, Man58, Neu77, Osl83, OR13, Pat62, Pau81, Qui65, Sta74, Tur66, Wag67, Whi71, Zuc75]. VIII [May88]. Viktor [Cro71]. Villages [Coo14]. Villains [Cor77]. Vincent [Tob74, Zah12]. Vinci [Ran53]. Vindicated [Wes58]. Vindication [Jan78, Met60, Rot58, Sir18, Wes58]. Vineyard [Pog63]. Violence [Hoo73, Pes14, Rei21, Sch73a, Ste73]. Virgil [Kal95]. Virginia [Cai07, Col10, Woo94]. Virginie [Men18]. Virility [Hep55]. Virtue [And08, Boy06, Cot79, CW13, Gas00, Jur01, Mit87, Mul96, Per02, Woo19]. Virtues [Har98]. Virtuoso [Hou42a, Hou42b, Jan08, Pri95]. Virtuous [Car91, Pri95]. Vision [Alt08, Arm56, Gar12, Hed98, Nad91, Zal85, Zwi20, Thr77]. Visions [Ros24]. Visual [Cer05, Wil88]. Vitalism [Pac02]. vitis [Wei19b]. Vito [Goo84]. Vitoria [Tun14, vL97]. Vives [Col62, Cou04, Nau15]. Vivisection [Pre03]. Vnconfused [Pat62]. Vocabulary [Bes48, Kos04]. Voegelin [Ger15, Por02]. Voice [Fal20, Pod02, Stu04]. Vol [Rie61]. Volk [HL19]. Völkerpsychologie [Kal87]. Völkisch [HL19]. Volksgemeinschaft [Mit71b]. Volney [Gar54, Kim18, Val24]. Voltaire [For16, Hen77, Roe85, Ros55, Sch43b, Sho85]. Volterra [Goo84]. Volume [Ano40i, Ano40j, Ano41k, Ano42i, Ano42j, Ano43f, Ano44g, Ano44h, Ano45p, Ano45q, Ano46p, Ano46q, Ano47p, Ano48n, Ano48o, Ano49e, Ano49f, Ano50d, Ano51b, Ano52g, Ano53c, Ano53d, Ano54d, Ano55d, Ano56g, Ano57l, Ano57m, Ano58k, Ano58l, Ano59l, Ano59m, Ano59n, Ano60d, Ano60e, Ano61f, Ano61g, Ano62j, Ano62k, Ano62l, Ano62m, Ano63k, Ano63l, Ano63m, Ano63n, Ano64k, Ano64l, Ano64m, Ano64n, Ano65j, Ano65k, Ano65l, Ano65m, Ano66k, Ano66l, Ano66m, Ano66n, Ano67l, Ano67m, Ano67n, Ano67o, Ano68l, Ano68m, Ano68n, Ano68o, Ano69m, Ano69n, Ano69o, Ano69p, Ano70l, Ano70m, Ano70n, Ano70o, Ano71j, Ano71k, Ano71l, Ano71m, Ano72i, Ano72j, Ano72k, Ano72l, Ano73j, Ano73k, Ano73l, Ano73m, Ano74i, Ano74j, Ano74k, Ano74l].

Volume [Ano75l, Ano75m, Ano75n, Ano75o, Ano76j, Ano76k, Ano76l, Ano76m, Ano77j, Ano77k, Ano77l, Ano77m, Ano78i, Ano79g, Ano80n, Ano81k, Ano82i, Ano83k, Ano84k, Ano85j, Ano86o, Ano87l, Ano88o, Ano89l, Ano90n, Ano91m, Ano92n, Ano93m, Ano94m, Ano95l, Ano02e, Ano02f, Ano03e, Ano03g, Ano04e, Ano04f, Ano05c, Ano05d, Ano06e, Ano06f, Ano07e, Ano07f, Ano08e, Ano08f, Ano09e, Ano09f, Ano10c, Ano10d, Ano11b, Ano11c, Ano12e, Ano12f, Ano13d, Ano13e, Ano14d, Ano14f, Ano15c, Ano15d, Ano16e, Ano17e, Ano18d, Ano19d, Ano20e, Ano21d, Ano22d, Ano23e, Ano24a, Fer68]. Voluntarism [Gue83, Hei78, Hey79, Mal97]. Vossler [Gui08]. Vowel [Twi20]. Voyage [Bar73, McM66, Ser17]. Vry [Mie73]. vs [Ben82b, Con79, Kuh61, Spi44, Val04, Vie03, Wea80]. Vulcan [Lin74].

W [Gre15, Har85a, Rie61, Rog06, Tae50, Van93, Wil23]. W. [Not55, Pau79]. Wake [Hig91]. Wakefield [And08]. Waldo [Gre15]. Wallace [Bar58a, Hef78, Smi78]. Walled [Hed98]. Wallis [Pyc87]. Walpole [Hor80]. Walter [And11, Hal58, Lew01, Not16]. War [Ada79, Avi61, Bed21, Bon85, Bon56, Cam16, Cam22a, Cam22b, Cla23, Col82, Cra92, Fer73, Glo57, Har66, Jor52, Kap56, Par68a, Pur15, Ros12, Sch93, Sch22b, She44, Sin16, Smi65a, Zon14, vdL17, Bat24, Bon56, Lar10, Mac70, Moy23, Nad82, Pan21, Pas90, Por81, Rik54, Ter21]. Warburg [Gui08, Lev13, San20, Wen24]. Warburgs [Mar17]. Ward [Nel72, Nic68, Wal07]. Ware [Bri42]. Warfare [Mil70, Mor22b]. Wars [Cur42, Mor18, MS15, Tho94, Sal96]. Wartime [Pan21]. Was [Åsb21, Fal90, Lam09, Mon89, Mos41, Pfi97, Pre77, Rob62, Ros75, Ros83, Cat10, Gas59, Kas06]. Wasserman [Bin67]. Watchdogs [Sch73b]. Watts [Mac92]. Waves [Han60]. Way [Chi74, CH95, Far87, Kim14, Low61, Mye08, O’N12, Str51b, Dav10]. Ways [EST66]. Wealth [Hei73, Pac02, Whi76a]. Weaver [Mac80]. Webb [Ram01]. Weber [Cal22, Der10, Kae96, KE23, Mag16, Zha25]. Webster [Wol49, Byn84]. Weeds [Koc99]. Weight [TT24]. Weimar [Cre95, Lev13, Lie98]. Weinberg [Ano69l]. Welfare [Kri63]. Well [Cas81, SR81]. Well-Principled [SR81]. Wellek [Cre83, Wei69, Cul88]. Wells [Hyd56, Sol70, Sol73, Sti09]. Went [Bri53]. Wenz [McK82b]. Were [Ros77]. Werewolf [Shy14]. Werner [O’F92]. West [Bar58b, Bri23, Ger25, Goo90b, Jac68, Nee51a, Nee51b, Sch51, Sil01, Cas48, HE63]. Western [Boa47, Lac10, Lun77, Moo68, Lau44, Mer11, Mun79, Puz64]. Westernization [Lew58]. Weyer [Mor25]. Whately [McK81b, Ras77]. Wheel [Men20]. Wheeler [Söl16]. Where [Cha23b, Guc00]. Whewell [Cha24, Hef78, Rau88, Str55, Wil74b]. Which [Kja02, Lun77, Mye08, Her05, Isr06]. Whichcote [Gre81, Gre91b]. Whig [Con82]. Whiggery [Cre79]. Whiggish [May90]. Whigs [Pla13]. White [Guy02, Van59]. Whitehead [Ger15, Har59, Joh46, Low49, Mor90, Söl16]. Whitman [Blu63]. Who [AP92, Sir18]. Whole [Cla98, Dun69, Bui02]. Whose [Smi95, Wei72]. Whytt [Bas13]. Wiener [Sch50, Axi93, Boa72, EST66, Shi72]. Wilamowitz [Gro50]. Wilamowitz-Möllendorf [Gro50]. Wilberforce [Bak70]. Wild [You01]. Wilderness [Lee95]. Wildness [Lee95]. Wilhelm [Fro95, Hol50, Mas52, Sor83, Swe73]. Wilhelmine [Wer13]. Wilkins [Cla82, Gre81]. Will [Ban16, Dam79, Dan60, Gue95, Gun68, Men16, Mol97, Nad01, Ned99, Sin16, Sta08a, Tri49, Ver22, Wil96, Wil05, Wis97, Goo83, Gra65, Lun87, Osl83]. Willem [vE01]. William [Bal50, Ano88e, Aus52, Bak70, Bat00b, Bay49, BI60, Bux84, Cad95, Cha24, Cla85, Cop51, DeA87, Dot86, Fas97, Fei70, Fun94, Gib72, Guy20, Har02a, Hef78, Hen01, Hor80, Hut78, Kar65, Kau63, Keg58, Kin00, Kon97, Kos04, Kra40, Kre00, Lam09, Lev47b, McC20, Mud68, Nob65, Rau88, Sch93, Sho52, Sil78, Sil86, Str55, Str54, Thr11, Wei52, WB93, Wom86, Wor49, Zil41]. Willing [Har72]. Wilson [Jez71, Ste89]. Wilting [McG03]. Winckelmann [Kau01, Vic02]. Wing [Per18]. Wins [Ano47k]. Winstanley [Van65, Jur75]. Winzerus [Bur60]. Wisdom [Bar60a, Min98, Ter18]. Wise [LS23]. Wissenschaft [Bit17]. Wissenschaftsgeschichte [Ano98e]. Wit [Lun04, WA97, Wes77]. Witch [Mor25]. Witchcraft [Dav13, Ste65, Tea62, Tri69]. Withering [Ver84, Blo46]. Withering-Away [Ver84]. Within [Poi70, Ulr71, Wil51]. Without [Bar19, Lev72b, Hel19, Kin56, Low24, Wil51, Woo18]. Witness [Phi83]. Wittenberg [Bro05]. Wittgenstein [Ano88f, Joh88, Puc05, Roz02, War11]. Wolff [Bla61, Boa58, Cor75b, Frä75, Lac53, Van95]. Wolffian [Cal69]. Wolffianism [Fol17]. Wolfgang [Kan45]. Wolfson [Boa48b]. Wollaston [Fei77]. Wollstonecraft [Nic90, BB89, Con99, O’N07, Sir18]. Wollstonecraft’s [Jan78]. Woman [Ala77, AP92, Bro06, Fin14, Jan78, Moy75, Sei74, Sir18]. Women [Ban55, Cum73, CW13, Hil11, Nad82, Sie50, Sir18, Smi24a, Smi24b, Tob74, Wil75, Gre23a, Tor84]. Wonderful [Doa17, Ano88f, Joh88]. Wonderland [Pet85]. Wood [Eig09]. Woodrow [Ste89]. Woolf [Cai07, Woo94]. Word [Wil85]. Words [Cha23a, TT24]. Wordsworth [Bea40, Jos06, Wym49]. Work [Ano48a, Bly86, Bou87, Bra02, Bux84, Dyn81, Gar55, Gas59, Gui13, Had52, Isa10, Kin00, Mad54, Ros43a, Sam43]. Working [Ran65]. Working-Out [Ran65]. Works [Ano83f, Bak83, Dav94, Fer68, Flo95, Har02a, Kas77, Kli81, Mor09, Neg79, Tag86, Teg47, Tor84, Wil96, Wor49, Zea59]. Workshop [Met63]. World [Chr46, Cla98, Far91, Fit97, Goo84, Gre65, Gur00, Had43, Hef78, Kei04, Laz03, Mac88, Mil74, New95, Pan21, Pet01, Phi07, Pri74, Rus40, Sar85, TT24, Tuv50, Wes01, Wes76, Zwi20, dG84, Cla23, Kap56, Lau44, Rob23, Ros12, She44, Lac10, Smi65b]. World-Makers [Tuv50]. World-View [Sar85]. Worldliness [Bri00]. Worlds [Bar73, Sch83, Sil76]. Worms [Bac07]. Worship [MF06]. Wright [Ano45f, Ano45o, Gri12, Mad54, Wie45]. Writer [Bur60, Dra90]. Writers [Bar47, Sir18]. Writes [Fub96]. Writing [Ack73, Alt07, Bac07, Bre10, Bro05, Eng80, Fis66, Hel21a, Lar10, Mye93, Pod10, Rad93, Sko64, dS23, Ack73].

Writings [Ano85e, Fie96, Hil11, McD05, MT13, Min98, Per57, Per18, Ran64, Ril94, Shi72, Swe73, Wom86, Wri40, Ros85a]. Wrong [Bri53]. Wunderkammer [Wes01]. Wyck [Mil41]. Wyclif [Lah97, Mic03].

Xenophon [Wel76]. XI [Bak91]. XIII [Bak91, Dau77]. XIIth [Tat65]. XIV [Keo74, Llo69, Ruo66]. XIXth [Gra65]. XIXth-Century [Gra65]. Xomjakov [McN66]. XVIIe [Mor40]. XVIIIth [Kno65, Rou67]. XVIIth [Kno65, Kri64, Tat65]. XVth [Sig62]. XVth-Century [Sig62]. XXth [Sko64].

Yahoo [Fry54, Mur54]. Yahya [RS64]. Yang [Liu12]. Yannis [Ger25]. Yates [Spr76]. Year [And21]. Years [Dyk59, Gil86, Han95, Rob59, Sta09a, Wil87]. Yeoman [Gre64]. Yolton [Rog06]. York [Bar42]. Youmans [Haa48]. Young [Bia24, Bok17, Cro71, Gre63, Ham70, Kau63, Mil74, Swe73, Wet05]. Younger [Fas97]. Yugoslav [Rob20b]. Yunjidang [Kim14]. Yves [Wol07].

Zapolya [Coa77]. Zea [Gra75]. Zen [Mil99a]. Zeno [Gla01]. Zettelkasten [Hel19]. Zeus [War71]. Zhou [Wan05]. Zilsel [Kel50, LS17]. Zionism [Tay17]. Zionist [Loe70]. Znaniecki [Ano77e, McL76a]. Zumel [Wei19b]. Zunz [Bit17]. Zur [All74].

References

Amati:1985:PCH

[AA85] Frank Amati and Tony Aspromourgos. Petty contra Hobbes: a previously untranslated manuscript. Journal of the History of Ideas, 46 (1):127–132, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

        stable/2709780.

Achinstein:2020:HGM

[Ach20] Sharon Achinstein. Hugo Grotius and marriage’s global past: Conjugal thinking in early modern political thought. Journal of the History of Ideas, 81(2):195–215, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

        edu/pub/56/article/753534.

Ackerman:1973:BRW

[Ack73] Robert Ackerman. Book review: Writing about Writing about Myth. Journal of the History of Ideas, 34(1):147–155, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

        URL https://www.jstor.org/stable/2708952.

Ackerman:1975:FMR

[Ack75] Robert Ackerman. Frazer on myth and ritual. Journal of the History of Ideas, 36(1):115–134, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709014.

Acworth:1977:MHH

[Acw77] Richard Acworth. Malebranche and his heirs. Journal of the History of Ideas, 38(4):673–676, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708695.

Adams:1942:SAT

[Ada42] John Clarke Adams. Some antecedents of the theory of the corporative system. Journal of the History of Ideas, 3(2):182–189, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707176.

Adams:1949:SRS

[Ada49] Robert P. Adams. The social responsibilities of science in Utopia, New Atlantis and after. Journal of the History of Ideas, 10(3):374– 398, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707043.

Adamiak:1974:MED

[Ada74] Richard Adamiak. Marx, Engels, and Dühring. Journal of the History of Ideas, 35(1):98–112, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708745.

Adams:1979:TCW

[Ada79] Michael C. C. Adams. Tennyson’s crimean war poetry: a crosscultural approach. Journal of the History of Ideas, 40(3):405– 422, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709245.

Adamson:1980:GIF

[Ada80] Walter L. Adamson. Gramsci’s interpretation of Fascism. Journal of the History of Ideas, 41(4):615–633, October/December 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709277.

Adams:1989:ARP

[Ada89] John C. Adams. Alexander Richardson’s philosophy of art and the sources of the Puritan social ethic. Journal of the History of Ideas, 50(2):227–247, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709733.

Adams:1991:MMD

[Ada91] David Adams. Metaphors for mankind: The development of Hans Blumenberg’s anthropological metaphorology. Journal of the History of Ideas, 52(1):152–166, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709587.

Adams:1992:PFJ

[Ada92] James Eli Adams. Philosophical forgetfulness: John Stuart Mill’s “Nature”. Journal of the History of Ideas, 53(3):437–454, July/ September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709886.

Adamson:2001:ASA

[Ada01] Peter Adamson. Aristotelianism and the soul in the Arabic Plotinus. Journal of the History of Ideas, 62(2):211–232, April 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15108.

Adams:2016:FAS

[Ada16] Matthew S. Adams. Formulating an anarchist sociology: Peter Kropotkin’s reading of Herbert Spencer. Journal of the History of Ideas, 77(1):49–73, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/610560.

Adelman:2018:ENR

[Ade18] Jeremy Adelman. Empires, nations, and revolutions. Journal of the History of Ideas, 79(1):73–88, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/684833.

Adkins:2000:WIM

[Adk00] Gregory Matthew Adkins. When ideas matter: The moral philosophy of Fontenelle. Journal of the History of Ideas, 61(3):433–452, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15077.

Aerts:2019:HAE

[Aer19] Saskia Aerts. Historical approaches to epistemic authority: The case of neoplatonism. Journal of the History of Ideas, 80(3):343–363, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/729716.

Africa:1960:PTS

[Afr60] Thomas W. Africa. Phylarchus, Toynbee, and the Spartan myth. Journal of the History of Ideas, 21(2):266–272, April/June 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708199.

Africa:1961:OAM

[Afr61] Thomas W. Africa. The opium addiction of Marcus aurelius. Journal of the History of Ideas, 22(1):97–102, January/March 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707876.

Africa:1962:CGR

[Afr62] Thomas W. Africa. The city of God revisited: Toynbee’s reconsiderations. Journal of the History of Ideas, 23(2):282–292, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708163.

Africa:1993:ODT

[Afr93] Thomas W. Africa. The owl at dusk: Two centuries of classical scholarship. Journal of the History of Ideas, 54(1):143–163, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709865.

Ariew:1995:IBD

[AG95] Roger Ariew and Marjorie Grene. Ideas, in and before Descartes. Journal of the History of Ideas, 56(1):87–106, January 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710008.

Agassi:1969:LPH

[Aga69] Joseph Agassi. Leibniz’s place in the history of physics. Journal of the History of Ideas, 30(3):331–344, July/September 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708561.

Agassi:1973:CDH

[Aga73] Joseph Agassi. Continuity and discontinuity in the history of science. Journal of the History of Ideas, 34(4):609–626, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708892.

Appleby:1995:R

[AHJ95] Joyce Appleby, Lynn Hunt, and Margaret Jacob. Response. Journal of the History of Ideas, 56(4):675–680, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709999.

Allison:1981:LPV

[AJ81] Henry E. Allison and Nicholas Jolley. Locke’s pyrrhic victory. Journal of the History of Ideas, 42(4):672–674, October/December 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709126.

Abbamonte:2023:IBE

[AK23] Giancarlo Abbamonte and Craig Kallendorf. On indexing: The birth and early development of an idea. Journal of the History of Ideas, 84(3):465–486, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/901490.

Alaya:1967:ARP

[Ala67] Flavia M. Alaya. Arnold and renan on the popular uses of history. Journal of the History of Ideas, 28(4):551–574, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708529.

Alaya:1977:VSG

[Ala77] Flavia Alaya. Victorian science and the “Genius” of woman. Journal of the History of Ideas, 38(2):261–280, April/June 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708911.

Albaum:1955:MDP

[Alb55] Martin Albaum. The moral defenses of the physiocrats’ laissez-faire. Journal of the History of Ideas, 16(2):179–197, April 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707661.

Albury:1978:HOP

[Alb78] W. R. Albury. Halley’s ode on the Principia of Newton and the Epicurean revival in England. Journal of the History of Ideas, 39 (1):24–43, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709070.

Alborn:2017:GME

[Alb17] Timothy Alborn. The greatest metaphor ever mixed: Gold in the British Bible, 1750–1850. Journal of the History of Ideas, 78(3):427– 447, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665522.

Aldridge:1944:EMA

[Ald44] Alfred Owen Aldridge. The eclecticism of Mark Akenside’s “The Pleasures of Imagination”. Journal of the History of Ideas, 5(3):292–314, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707352.

Alexander:1966:JGH

[Ale66] W. M. Alexander. Johann Georg Hamann: Metacritic of Kant. Journal of the History of Ideas, 27(1):137–144, January/March 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708314.

Aljunied:2019:DSN

[Alj19] Khairudin Aljunied. Deformations of the secular: Naquib Al-Attas’s conception and critique of secularism. Journal of the History of Ideas, 80(4):643–663, October 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/735385.

Allen:1949:SSE

[All49] Phyllis Allen. Scientific studies in the English universities of the Seventeenth Century. Journal of the History of Ideas, 10(2):219–253, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707416.

Allen:1960:SP

[All60] R. E. Allen. The Socratic paradox. Journal of the History of Ideas, 21(2):256–265, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708198.

Allison:1966:LTP

[All66] Henry E. Allison. Locke’s theory of personal identity: a reexamination. Journal of the History of Ideas, 27(1):41–58, January/ March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708307.

Allentuck:1974:HFJ

[All74] Marcia Allentuck. Henry Fuseli and J. G. Herder’s Ideen Zur Philosophie Der Geschichte Der Menschheit in Britain: an unremarked connection. Journal of the History of Ideas, 35(1):113–120, January/ March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708746.

Allen:1975:AAF

[All75] Michael J. B. Allen. The absent angel in Ficino’s philosophy. Journal of the History of Ideas, 36(2):219–240, April/June 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708925.

Allen:1985:CCS

[All85] Peter Allen. S. T. Coleridge’s Church and State and the idea of an intellectual establishment. Journal of the History of Ideas, 46 (1):89–106, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709777.

Allen:1989:MCC

[All89] Peter Allen. Morrow on Coleridge’s Church and state. Journal of the History of Ideas, 50(3):485–489, July/September 1989. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709573.

Almasy:1978:PRH

[Alm78] Rudolph Almasy. The purpose of Richard Hooker’s polemic. Journal of the History of Ideas, 39(2):251–270, April/June 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708778.

Almond:1993:HMA

[Alm93] Philip C. Almond. Henry More and the apocalypse. Journal of the History of Ideas, 54(2):189–200, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709978.

Altman:2007:LSG

[Alt07] William H. F. Altman. Leo Strauss on “German nihilism”: Learning the art of writing. Journal of the History of Ideas, 68(4):587–612, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
223127.

Alter:2008:MST

[Alt08] Stephen G. Alter. Mandeville’s ship: Theistic design and philosophical history in Charles Darwin’s vision of natural selection. Journal of the History of Ideas, 69(3):441–465, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/242280.

Amidon:2007:CCC

[Ami07] Kevin Scott Amidon. Carrie Chapman Catt and the evolutionary politics of sex and race, 1885–1940. Journal of the History of Ideas, 68(2): 305–328, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
215050.

Anderson:1941:GRI

[And41] Eugene N. Anderson. German romanticism as an ideology of cultural crisis. Journal of the History of Ideas, 2(3):301–317, June 1941. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707133.

Anderson:1955:CGT

[And55] G. L. Anderson. Charles Gildon’s total academy. Journal of the History of Ideas, 16(2):247–251, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707666.

Anderson:1964:IJE

[And64] Wallace E. Anderson. Immaterialism in Jonathan Edwards’ early philosophical notes. Journal of the History of Ideas, 25(2):181–200, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708011.

Anderson:1976:CBT

[And76] Lorin Anderson. Charles Bonnet’s taxonomy and chain of being. Journal of the History of Ideas, 37(1):45–58, January/March 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708708.

Andrews:2002:MHF

[And02] Naomi J. Andrews. La mère humanité“: Femininity in the romantic socialism of Pierre Leroux and the Abbé A.-L. Constant. Journal of the History of Ideas, 63(4):697–716, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/39289.

Andrew:2004:SMP

[And04] Edward Andrew. The Senecan moment: Patronage and philosophy in the Eighteenth Century. Journal of the History of Ideas, 65(2):277– 299, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/173843.

Anderson:2008:SCV

[And08] Margaret Anderson. Stoic constructions of virtue in the vicar of wakefield. Journal of the History of Ideas, 69(3):419–439, July 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/242279.

Andrews:2011:WPO

[And11] Kit Andrews. Walter Pater as Oxford Hegeliano: Plato and Platonism and T. H. Green’s Prolegomena to Ethics. Journal of the History of Ideas, 72(3):437–459, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/445505.

Andersson:2021:PAF

[And21] Jenny Andersson. Planning the American future: Daniel Bell, future research, and the commission on the year 2000. Journal of the History of Ideas, 82(4):661–682, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/837676.

Angle:2000:SWA

[Ang00] Stephen C. Angle. Should we all be more English? Liang Qichao, Rudolf von Jhering, and rights. Journal of the History of Ideas, 61(2): 241–261, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15066.

Anonymous:1940:BRa

[Ano40a] Anonymous. Books received. Journal of the History of Ideas, 1(2): 252–254, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707337.

Anonymous:1940:BRb

[Ano40b] Anonymous. Books received. Journal of the History of Ideas, 1(3): 375–377, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707095.

Anonymous:1940:BRc

[Ano40c] Anonymous. Books received. Journal of the History of Ideas, 1 (4):509–512, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707129.

Anonymous:1940:FMa

[Ano40d] Anonymous. Front matter. Journal of the History of Ideas, 1(1):2, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707006.

Anonymous:1940:FMb

[Ano40e] Anonymous. Front matter. Journal of the History of Ideas, 1(2):129– 130, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707329.

Anonymous:1940:FMc

[Ano40f] Anonymous. Front matter. Journal of the History of Ideas, 1(3):257– 258, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707086.

Anonymous:1940:PRRa

[Ano40g] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 1(2):255, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707338.

Anonymous:1940:PRRb

[Ano40h] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 1(3):378–379, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707096.

Anonymous:1940:VIa

[Ano40i] Anonymous. Volume information. Journal of the History of Ideas, 1 (1):1, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707005.

Anonymous:1940:VIb

[Ano40j] Anonymous. Volume information. Journal of the History of Ideas, 1(4):514–526, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707121.

Anonymous:1941:A

[Ano41a] Anonymous. Announcement. Journal of the History of Ideas, 2(2): 255, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707119.

Anonymous:1941:BRa

[Ano41b] Anonymous. Books received. Journal of the History of Ideas, 2 (1):123–129, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707286.

Anonymous:1941:BRb

[Ano41c] Anonymous. Books received. Journal of the History of Ideas, 2(2): 250–254, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707117.

Anonymous:1941:BRc

[Ano41d] Anonymous. Books received. Journal of the History of Ideas, 2(3): 381–388, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707138.

Anonymous:1941:BRd

[Ano41e] Anonymous. Books received. Journal of the History of Ideas, 2 (4):508–514, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707025.

Anonymous:1941:FM

[Ano41f] Anonymous. Front matter. Journal of the History of Ideas, 2(1):??, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707278.

Anonymous:1941:PRRa

[Ano41g] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 2(1):130, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707287.

Anonymous:1941:PRRb

[Ano41h] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 2(2):255, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707118.

Anonymous:1941:PRRc

[Ano41i] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 2(3):389–390, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707139.

Anonymous:1941:PRRd

[Ano41j] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 2(4):515, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707026.

Anonymous:1941:VI

[Ano41k] Anonymous. Volume information. Journal of the History of Ideas, 2(4):516–527, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707017.

Anonymous:1942:A

[Ano42a] Anonymous. Announcement. Journal of the History of Ideas, 3(3): 357, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707310.

Anonymous:1942:BRa

[Ano42b] Anonymous. Books received. Journal of the History of Ideas, 3 (1):113–118, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707466.

Anonymous:1942:BRb

[Ano42c] Anonymous. Books received. Journal of the History of Ideas, 3(2): 232–236, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707180.

Anonymous:1942:BRc

[Ano42d] Anonymous. Books received. Journal of the History of Ideas, 3(3): 358–361, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707311.

Anonymous:1942:BRd

[Ano42e] Anonymous. Books received. Journal of the History of Ideas, 3 (4):469–471, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707319.

Anonymous:1942:CSJ

[Ano42f] Anonymous. Conservation of scholarly journals. Journal of the History of Ideas, 3(2):238, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707182.

Anonymous:1942:PRRa

[Ano42g] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 3(1):119, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707467.

Anonymous:1942:PRRb

[Ano42h] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 3(2):237, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707181.

Anonymous:1942:VIa

[Ano42i] Anonymous. Volume information. Journal of the History of Ideas, 3(1):1–2, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707457.

Anonymous:1942:VIb

[Ano42j] Anonymous. Volume information. Journal of the History of Ideas, 3(4):474–484, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707313.

Anonymous:1943:BRa

[Ano43a] Anonymous. Books received. Journal of the History of Ideas, 4 (1):119–124, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707240.

Anonymous:1943:BRb

[Ano43b] Anonymous. Books received. Journal of the History of Ideas, 4(2): 246–249, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707327.

Anonymous:1943:BRc

[Ano43c] Anonymous. Books received. Journal of the History of Ideas, 4(3): 374–376, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707261.

Anonymous:1943:BRd

[Ano43d] Anonymous. Books received. Journal of the History of Ideas, 4 (4):490–494, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707171.

Anonymous:1943:PRR

[Ano43e] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 4(3):377, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707262.

Anonymous:1943:VI

[Ano43f] Anonymous. Volume information. Journal of the History of Ideas, 4(4):497–506, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707164.

Anonymous:1944:BRa (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707106.

Anonymous:1944:BRb

[Ano44b] Anonymous. Books received. Journal of the History of Ideas, 5(2): 250–253, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707389.

Anonymous:1944:BRc

[Ano44c] Anonymous. Books received. Journal of the History of Ideas, 5(3): 377, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707358.

Anonymous:1944:BRd

[Ano44d] Anonymous. Books received. Journal of the History of Ideas, 5 (4):489–493, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707083.

Anonymous:1944:PARa

[Ano44e] Anonymous. Periodical articles and reprints received. Journal of the History of Ideas, 5(1):123–124, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707107.

Anonymous:1944:PARb

[Ano44f] Anonymous. Periodical articles and reprints received. Journal of the History of Ideas, 5(4):494, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707084.

Anonymous:1944:VIa

[Ano44g] Anonymous. Volume information. Journal of the History of Ideas, 5(1):1–2, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707098.

Anonymous:1944:VIb

[Ano44h] Anonymous. Volume information. Journal of the History of Ideas, 5(4):495–505, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707077.

Anonymous:1945:A

[Ano45a] Anonymous. Announcements. Journal of the History of Ideas, 6(1): 126, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707066.

Anonymous:1945:BRa

[Ano45b] Anonymous. Books received. Journal of the History of Ideas, 6 (1):122–124, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707064.

Anonymous:1945:BRb

[Ano45c] Anonymous. Books received. Journal of the History of Ideas, 6(2): 252–254, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707367.

Anonymous:1945:BRc

[Ano45d] Anonymous. Books received. Journal of the History of Ideas, 6(3): 378–379, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707301.

Anonymous:1945:BRd

[Ano45e] Anonymous. Books received. Journal of the History of Ideas, 6 (4):502–503, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707348.

Anonymous:1945:ECW

[Ano45f] Anonymous. Errata: Chauncey Wright’s defense of Darwin and the neutrality of science. Journal of the History of Ideas, 6(3):377, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707300. See

[Wie45].

Anonymous:1945:FMa

[Ano45g] Anonymous. Front matter. Journal of the History of Ideas, 6(1):1–2, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707052.

Anonymous:1945:FMb

[Ano45h] Anonymous. Front matter. Journal of the History of Ideas, 6(2):??, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707360.

Anonymous:1945:FMc

[Ano45i] Anonymous. Front matter. Journal of the History of Ideas, 6(3):??, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707289.

Anonymous:1945:FMd

[Ano45j] Anonymous. Front matter. Journal of the History of Ideas, 6(4):??, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707341.

Anonymous:1945:PARa

[Ano45k] Anonymous. Periodical articles and reprints received. Journal of the History of Ideas, 6(1):125–126, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707065.

Anonymous:1945:PARb

[Ano45l] Anonymous. Periodical articles and reprints received. Journal of the History of Ideas, 6(2):255, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707368.

Anonymous:1945:RPAa

[Ano45m] Anonymous. Reprints and periodical articles received. Journal of the History of Ideas, 6(3):380–381, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707302.

Anonymous:1945:RPAb

[Ano45n] Anonymous. Reprints and periodical articles received. Journal of the History of Ideas, 6(4):504, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707349.

Anonymous:1945:UEC

[Ano45o] Anonymous. An unpublished essay by Chauncey Wright (1830–1875). Journal of the History of Ideas, 6(1):89–90, January 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707058.

Anonymous:1945:VIa

[Ano45p] Anonymous. Volume information. Journal of the History of Ideas, 6(1):??, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707051.

Anonymous:1945:VIb

[Ano45q] Anonymous. Volume information. Journal of the History of Ideas, 6(4):505–508, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707340.

Anonymous:1946:Aa

[Ano46a] Anonymous. Announcement. Journal of the History of Ideas, 7(1): 127, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707276.

Anonymous:1946:Ab

[Ano46b] Anonymous. Announcement. Journal of the History of Ideas, 7(3): 379, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707411.

Anonymous:1946:BRa

[Ano46c] Anonymous. Books received. Journal of the History of Ideas, 7 (1):122–124, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707274.

Anonymous:1946:BRb

[Ano46d] Anonymous. Books received. Journal of the History of Ideas, 7(2): 250–251, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707074.

Anonymous:1946:BRc

[Ano46e] Anonymous. Books received. Journal of the History of Ideas, 7(3): 374–376, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707409.

Anonymous:1946:BRd

[Ano46f] Anonymous. Books received. Journal of the History of Ideas, 7 (4):499–501, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707036.

Anonymous:1946:CEP

[Ano46g] Anonymous. Correction: The evolutionism and pragmaticism of Peirce. Journal of the History of Ideas, 7(4):504, October 1946. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707038. See [Wie46a].

Anonymous:1946:FMa

[Ano46h] Anonymous. Front matter. Journal of the History of Ideas, 7(1):1–2, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707265.

Anonymous:1946:FMb

[Ano46i] Anonymous. Front matter. Journal of the History of Ideas, 7(2):??, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707068.

Anonymous:1946:FMc

[Ano46j] Anonymous. Front matter. Journal of the History of Ideas, 7(3):??, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707401.

Anonymous:1946:FMd

[Ano46k] Anonymous. Front matter. Journal of the History of Ideas, 7(4):??, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707029.

Anonymous:1946:PRRa

[Ano46l] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 7(1):125–126, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707275.

Anonymous:1946:PRRb

[Ano46m] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 7(2):252–254, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707075.

Anonymous:1946:PRRc

[Ano46n] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 7(3):377–378, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707410.

Anonymous:1946:PRRd

[Ano46o] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 7(4):502–504, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707037.

Anonymous:1946:VIa

[Ano46p] Anonymous. Volume information. Journal of the History of Ideas, 7(1):??, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707264.

Anonymous:1946:VIb

[Ano46q] Anonymous. Volume information. Journal of the History of Ideas, 7(4):505–506, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707028.

Anonymous:1947:A

[Ano47a] Anonymous. Announcements. Journal of the History of Ideas, 8(4): 499, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707435.

Anonymous:1947:BRa

[Ano47b] Anonymous. Books received. Journal of the History of Ideas, 8 (1):124–127, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707446.

Anonymous:1947:BRb

[Ano47c] Anonymous. Books received. Journal of the History of Ideas, 8(2): 249–252, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707161.

Anonymous:1947:BRc

[Ano47d] Anonymous. Books received. Journal of the History of Ideas, 8(3): 375–377, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707149.

Anonymous:1947:BRd

[Ano47e] Anonymous. Books received. Journal of the History of Ideas, 8 (4):492–496, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707433.

Anonymous:1947:CHC

[Ano47f] Anonymous. The Cohen–Holmes correspondence. Journal of the History of Ideas, 8(4):500, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707437.

Anonymous:1947:FMa

[Ano47g] Anonymous. Front matter. Journal of the History of Ideas, 8(1):??, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707439.

Anonymous:1947:FMb

[Ano47h] Anonymous. Front matter. Journal of the History of Ideas, 8(2):??, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707152.

Anonymous:1947:FMc

[Ano47i] Anonymous. Front matter. Journal of the History of Ideas, 8(3):??, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707141.

Anonymous:1947:FMd

[Ano47j] Anonymous. Front matter. Journal of the History of Ideas, 8(4):??, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707425.

Anonymous:1947:MNW

[Ano47k] Anonymous. Marjorie Nicolson wins Crawshay Prize awarded by the British Academy. Journal of the History of Ideas, 8(4):500, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707436.

Anonymous:1947:PRRa

[Ano47l] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 8(1):128, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707447.

Anonymous:1947:PRRb

[Ano47m] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 8(2):253–254, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707162.

Anonymous:1947:PRRc

[Ano47n] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 8(3):378–380, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707150.

Anonymous:1947:PRRd

[Ano47o] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 8(4):497–498, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707434.

Anonymous:1947:VI

[Ano47p] Anonymous. Volume information. Journal of the History of Ideas, 8(4):503–506, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707424.

Anonymous:1948:ALS

[Ano48a] Anonymous. Arthur O. Lovejoy at seventy-five: Reason at work. Journal of the History of Ideas, 9(4):403, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707212.

Anonymous:1948:BRa

[Ano48b] Anonymous. Books received. Journal of the History of Ideas, 9 (1):125–126, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707398.

Anonymous:1948:BRb

[Ano48c] Anonymous. Books received. Journal of the History of Ideas, 9(2): 245–250, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707231.

Anonymous:1948:BRc

[Ano48d] Anonymous. Books received. Journal of the History of Ideas, 9(3): 393–398, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707378.

Anonymous:1948:BRd

[Ano48e] Anonymous. Books received. Journal of the History of Ideas, 9 (4):519–521, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707221.

Anonymous:1948:FMa

[Ano48f] Anonymous. Front matter. Journal of the History of Ideas, 9(1):??, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707392.

Anonymous:1948:FMb

[Ano48g] Anonymous. Front matter. Journal of the History of Ideas, 9(2):??, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707224.

Anonymous:1948:FMc

[Ano48h] Anonymous. Front matter. Journal of the History of Ideas, 9(3):??, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707370.

Anonymous:1948:FMd

[Ano48i] Anonymous. Front matter. Journal of the History of Ideas, 9(4):??, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707211.

Anonymous:1948:PRRa

[Ano48j] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 9(1):127–128, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707399.

Anonymous:1948:PRRb

[Ano48k] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 9(2):251–254, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707232.

Anonymous:1948:PRRc

[Ano48l] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 9(3):399, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707379.

Anonymous:1948:PRRd

[Ano48m] Anonymous. Periodicals and reprints received. Journal of the History of Ideas, 9(4):522–523, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707222.

Anonymous:1948:VIa

[Ano48n] Anonymous. Volume information. Journal of the History of Ideas, 9(1):??, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707391.

Anonymous:1948:VIb

[Ano48o] Anonymous. Volume information. Journal of the History of Ideas, 9(4):525–526, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707210.

Anonymous:1949:BRa

[Ano49a] Anonymous. Books received. Journal of the History of Ideas, 10(1):150–154, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707208.

Anonymous:1949:BRb

[Ano49b] Anonymous. Books received. Journal of the History of Ideas, 10(2): 313–315, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707422.

Anonymous:1949:BRc

[Ano49c] Anonymous. Books received. Journal of the History of Ideas, 10(4):580–583, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707192.

Anonymous:1949:PAR

[Ano49d] Anonymous. Periodical articles received. Journal of the History of Ideas, 10(4):584, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707193.

Anonymous:1949:VIa

[Ano49e] Anonymous. Volume information. Journal of the History of Ideas, 10(1):1–2, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707195.

Anonymous:1949:VIb

[Ano49f] Anonymous. Volume information. Journal of the History of Ideas, 10(4):585–586, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707184.

Anonymous:1950:BRa

[Ano50a] Anonymous. Books received. Journal of the History of Ideas, 11(2): 248–254, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707251.

Anonymous:1950:BRb

[Ano50b] Anonymous. Books received. Journal of the History of Ideas, 11(4):511–513, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707596.

Anonymous:1950:PR

[Ano50c] Anonymous. Periodicals received. Journal of the History of Ideas, 11 (2):255, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707252.

Anonymous:1950:VI

[Ano50d] Anonymous. Volume information. Journal of the History of Ideas, 11(4):517–520, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707588.

Anonymous:1951:CDJ

[Ano51a] Anonymous. Corrigenda to Dr. Joseph Needham’s article, Part II. Journal of the History of Ideas, 12(4):628–629, October 1951. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707490. See [Nee51b].

Anonymous:1951:VI

[Ano51b] Anonymous. Volume information. Journal of the History of Ideas, 12(4):633–640, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707482.

Anonymous:1952:A

[Ano52a] Anonymous. Announcements. Journal of the History of Ideas, 13(4): 610, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707721.

Anonymous:1952:BRa

[Ano52b] Anonymous. Books received. Journal of the History of Ideas, 13(1):119–125, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707730.

Anonymous:1952:BRb

[Ano52c] Anonymous. Books received. Journal of the History of Ideas, 13(2): 281–284, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707617.

Anonymous:1952:BRc

[Ano52d] Anonymous. Books received. Journal of the History of Ideas, 13(4):605–609, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707720.

Anonymous:1952:FM

[Ano52e] Anonymous. Front matter. Journal of the History of Ideas, 13(1):??, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/3857202.

Anonymous:1952:PA

[Ano52f] Anonymous. Periodical articles. Journal of the History of Ideas, 13 (2):285, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707618.

Anonymous:1952:VI

[Ano52g] Anonymous. Volume information. Journal of the History of Ideas, 13(4):613–616, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707710.

Anonymous:1953:BR

[Ano53a] Anonymous. Books received. Journal of the History of Ideas, 14(4):634–636, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707708.

Anonymous:1953:FM

[Ano53b] Anonymous. Front matter. Journal of the History of Ideas, 14(1):??, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/3857200.

Anonymous:1953:VIa

[Ano53c] Anonymous. Volume information. Journal of the History of Ideas, 14(1):1–2, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707492.

Anonymous:1953:VIb

[Ano53d] Anonymous. Volume information. Journal of the History of Ideas, 14(4):641–642, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707696.

Anonymous:1954:BRa

[Ano54a] Anonymous. Books received. Journal of the History of Ideas, 15(1):193–197, January 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707657.

Anonymous:1954:BRb

[Ano54b] Anonymous. Books received. Journal of the History of Ideas, 15(4):602–606, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707680.

Anonymous:1954:FM

[Ano54c] Anonymous. Front matter. Journal of the History of Ideas, 15(1):??, January 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707646.

Anonymous:1954:VI

[Ano54d] Anonymous. Volume information. Journal of the History of Ideas, 15(4):608–610, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707671.

Anonymous:1955:BN

[Ano55a] Anonymous. Book notice. Journal of the History of Ideas, 16(3):420, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707642.

Anonymous:1955:BR

[Ano55b] Anonymous. Books received. Journal of the History of Ideas, 16 (4):563–565 + 568, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707512.

Anonymous:1955:FM

[Ano55c] Anonymous. Front matter. Journal of the History of Ideas, 16(1):1–2, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707523.

Anonymous:1955:VI

[Ano55d] Anonymous. Volume information. Journal of the History of Ideas, 16(4):569–570, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707502.

Anonymous:1956:A

[Ano56a] Anonymous. Announcement. Journal of the History of Ideas, 17(4): 566, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707791.

Anonymous:1956:BRa

[Ano56b] Anonymous. Books received. Journal of the History of Ideas, 17(1):140–144, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707694.

Anonymous:1956:BRb

[Ano56c] Anonymous. Books received. Journal of the History of Ideas, 17 (2):275–277 + 280, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707748.

Anonymous:1956:BRc

[Ano56d] Anonymous. Books received. Journal of the History of Ideas, 17(4):561–566, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707790.

Anonymous:1956:EHU

[Ano56e] Anonymous. Erratum: “History for Utopia: Saint-Simon and the Idea of Progress”. Journal of the History of Ideas, 17(4):568, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707792. See [Sim56].

Anonymous:1956:FM

[Ano56f] Anonymous. Front matter. Journal of the History of Ideas, 17(1):1–2, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707682.

Anonymous:1956:VI

[Ano56g] Anonymous. Volume information. Journal of the History of Ideas, 17(4):567–568, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707780.

Anonymous:1957:Ab

[Ano57a] Anonymous. Acknowledgements. Journal of the History of Ideas, 18(4):611, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707573.

Anonymous:1957:Aa

[Ano57b] Anonymous. Announcement. Journal of the History of Ideas, 18(4): 551, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707566.

Anonymous:1957:BRa

[Ano57c] Anonymous. Books received. Journal of the History of Ideas, 18(1):138–142, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707586.

Anonymous:1957:BRb

[Ano57d] Anonymous. Books received. Journal of the History of Ideas, 18(2): 285–289, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707631.

Anonymous:1957:BRc

[Ano57e] Anonymous. Books received. Journal of the History of Ideas, 18(3): 449–451, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707803.

Anonymous:1957:BRd

[Ano57f] Anonymous. Books received. Journal of the History of Ideas, 18(4):604–609, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707572.

Anonymous:1957:FMb

[Ano57g] Anonymous. Front matter. Journal of the History of Ideas, 18(1):2, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707576.

Anonymous:1957:FMa

[Ano57h] Anonymous. Front matter. Journal of the History of Ideas, 18(1):??, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/3857201.

Anonymous:1957:FMc

[Ano57i] Anonymous. Front matter. Journal of the History of Ideas, 18(2): ??, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707620.

Anonymous:1957:FMd

[Ano57j] Anonymous. Front matter. Journal of the History of Ideas, 18(3): ??, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707794.

Anonymous:1957:FMe

[Ano57k] Anonymous. Front matter. Journal of the History of Ideas, 18(4):??, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707558.

Anonymous:1957:VIa

[Ano57l] Anonymous. Volume information. Journal of the History of Ideas, 18 (1):1, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707575.

Anonymous:1957:VIb

[Ano57m] Anonymous. Volume information. Journal of the History of Ideas, 18(4):610–612, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707557.

Anonymous:1958:Ab

[Ano58a] Anonymous. Acknowledgements. Journal of the History of Ideas, 19(4):604, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707929.

Anonymous:1958:Aa

[Ano58b] Anonymous. Announcements. Journal of the History of Ideas, 19(2): 300, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707944.

Anonymous:1958:BRa

[Ano58c] Anonymous. Books received. Journal of the History of Ideas, 19(1):134–138, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707960.

Anonymous:1958:BRb

[Ano58d] Anonymous. Books received. Journal of the History of Ideas, 19(2): 294–299, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707943.

Anonymous:1958:BRc

[Ano58e] Anonymous. Books received. Journal of the History of Ideas, 19(3): 445–447, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708050.

Anonymous:1958:BRd

[Ano58f] Anonymous. Books received. Journal of the History of Ideas, 19(4):597–601, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707928.

Anonymous:1958:FMa

[Ano58g] Anonymous. Front matter. Journal of the History of Ideas, 19(1):2, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707947.

Anonymous:1958:FMb

[Ano58h] Anonymous. Front matter. Journal of the History of Ideas, 19(2): ??, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707931.

Anonymous:1958:FMc

[Ano58i] Anonymous. Front matter. Journal of the History of Ideas, 19(3): ??, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708036.

Anonymous:1958:FMd

[Ano58j] Anonymous. Front matter. Journal of the History of Ideas, 19(4):??, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707916.

Anonymous:1958:VIa

[Ano58k] Anonymous. Volume information. Journal of the History of Ideas, 19 (1):1, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707946.

Anonymous:1958:VIb

[Ano58l] Anonymous. Volume information. Journal of the History of Ideas, 19(4):602–604, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707915.

Anonymous:1959:Ac

[Ano59a] Anonymous. Acknowledgements. Journal of the History of Ideas, 20(4):615, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707898.

Anonymous:1959:Ab

[Ano59b] Anonymous. Advertisements. Journal of the History of Ideas, 20 (4):606, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707897.

Anonymous:1959:Aa

[Ano59c] Anonymous. Announcements. Journal of the History of Ideas, 20 (3):453–454, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708124.

Anonymous:1959:BRa

[Ano59d] Anonymous. Books received. Journal of the History of Ideas, 20(1):143–149, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707973.

Anonymous:1959:BRb

[Ano59e] Anonymous. Books received. Journal of the History of Ideas, 20(2): 283–289, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707827.

Anonymous:1959:BRc

[Ano59f] Anonymous. Books received. Journal of the History of Ideas, 20 (3):447–452, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708123.

Anonymous:1959:BRd

[Ano59g] Anonymous. Books received. Journal of the History of Ideas, 20(4): 601–605, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707896.

Anonymous:1959:FMa

[Ano59h] Anonymous. Front matter. Journal of the History of Ideas, 20(1):2, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707963.

Anonymous:1959:FMb

[Ano59i] Anonymous. Front matter. Journal of the History of Ideas, 20(2): ??, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707815.

Anonymous:1959:FMc

[Ano59j] Anonymous. Front matter. Journal of the History of Ideas, 20 (3):??, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708110.

Anonymous:1959:FMd

[Ano59k] Anonymous. Front matter. Journal of the History of Ideas, 20 (4):??, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707885.

Anonymous:1959:VIa

[Ano59l] Anonymous. Volume information. Journal of the History of Ideas, 20 (1):1, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707962.

Anonymous:1959:VIb

[Ano59m] Anonymous. Volume information. Journal of the History of Ideas, 20(3):??, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708109.

Anonymous:1959:VIc

[Ano59n] Anonymous. Volume information. Journal of the History of Ideas, 20 (4):613–616, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707884.

Anonymous:1960:BR

[Ano60a] Anonymous. Books received. Journal of the History of Ideas, 21(2):309–313, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708204.

Anonymous:1960:FM

[Ano60b] Anonymous. Front matter. Journal of the History of Ideas, 21 (1):??, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707995.

Anonymous:1960:ISH

[Ano60c] Anonymous. International Society for the History of Ideas. Journal of the History of Ideas, 21(2):314, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708205.

Anonymous:1960:VIa

[Ano60d] Anonymous. Volume information. Journal of the History of Ideas, 21(1):??, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707994.

Anonymous:1960:VIb

[Ano60e] Anonymous. Volume information. Journal of the History of Ideas, 21 (4):615–618, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708095.

Anonymous:1961:BRa

[Ano61a] Anonymous. Books received. Journal of the History of Ideas, 22 (1):136–141, January/March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707882.

Anonymous:1961:BRb

[Ano61b] Anonymous. Books received. Journal of the History of Ideas, 22 (3):425–431, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708138.

Anonymous:1961:BRc

[Ano61c] Anonymous. Books received. Journal of the History of Ideas, 22(4): 582–584, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708034.

Anonymous:1961:ERP

[Ano61d] Anonymous. Errata: The Renaissance perversion of pastoral. Journal of the History of Ideas, 22(3):431, July/September 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708139. See [Hen61].

Anonymous:1961:FM

[Ano61e] Anonymous. Front matter. Journal of the History of Ideas, 22 (1):2, January/March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707870.

Anonymous:1961:VIa

[Ano61f] Anonymous. Volume information. Journal of the History of Ideas, 22(1):1, January/March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707869.

Anonymous:1961:VIb

[Ano61g] Anonymous. Volume information. Journal of the History of Ideas, 22 (4):585–588, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708023.

Anonymous:1962:A

[Ano62a] Anonymous. Acknowledgments. Journal of the History of Ideas, 23 (4):572, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708177.

Anonymous:1962:BRa

[Ano62b] Anonymous. Books received. Journal of the History of Ideas, 23 (1):147–153, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708065.

Anonymous:1962:BRb

[Ano62c] Anonymous. Books received. Journal of the History of Ideas, 23(2):293–299, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708164.

Anonymous:1962:BRc

[Ano62d] Anonymous. Books received. Journal of the History of Ideas, 23 (3):424–429, July/September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708078.

Anonymous:1962:BRd

[Ano62e] Anonymous. Books received. Journal of the History of Ideas, 23(4): 565–569, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708176.

Anonymous:1962:FMa

[Ano62f] Anonymous. Front matter. Journal of the History of Ideas, 23 (1):2, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708053.

Anonymous:1962:FMb

[Ano62g] Anonymous. Front matter. Journal of the History of Ideas, 23(2):??, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708152.

Anonymous:1962:FMc

[Ano62h] Anonymous. Front matter. Journal of the History of Ideas, 23 (3):??, July/September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708068.

Anonymous:1962:FMd

[Ano62i] Anonymous. Front matter. Journal of the History of Ideas, 23 (4):??, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708167.

Anonymous:1962:VIa

[Ano62j] Anonymous. Volume information. Journal of the History of Ideas, 23(1):1, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708052.

Anonymous:1962:VIb

[Ano62k] Anonymous. Volume information. Journal of the History of Ideas, 23(2):??, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708151.

Anonymous:1962:VIc

[Ano62l] Anonymous. Volume information. Journal of the History of Ideas, 23(3):??, July/September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708067.

Anonymous:1962:VId

[Ano62m] Anonymous. Volume information. Journal of the History of Ideas, 23 (4):570–572, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708166.

Anonymous:1963:A

[Ano63a] Anonymous. Acknowledgments. Journal of the History of Ideas, 24 (4):601, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707992.

Anonymous:1963:BRa

[Ano63b] Anonymous. Books received. Journal of the History of Ideas, 24 (1):155–159, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707867.

Anonymous:1963:BRb

[Ano63c] Anonymous. Books received. Journal of the History of Ideas, 24(2):293–301, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707852.

Anonymous:1963:BRc

[Ano63d] Anonymous. Books received. Journal of the History of Ideas, 24 (3):440–443, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708221.

Anonymous:1963:BRd

[Ano63e] Anonymous. Books received. Journal of the History of Ideas, 24(4): 593–598, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707991.

Anonymous:1963:EN

[Ano63f] Anonymous. Editorial note. Journal of the History of Ideas, 24 (4):592, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707990.

Anonymous:1963:FMa

[Ano63g] Anonymous. Front matter. Journal of the History of Ideas, 24 (1):2, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707855.

Anonymous:1963:FMb

[Ano63h] Anonymous. Front matter. Journal of the History of Ideas, 24(2):??, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707842.

Anonymous:1963:FMc

[Ano63i] Anonymous. Front matter. Journal of the History of Ideas, 24 (3):??, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708208.

Anonymous:1963:FMd

[Ano63j] Anonymous. Front matter. Journal of the History of Ideas, 24 (4):??, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707976.

Anonymous:1963:VIa

[Ano63k] Anonymous. Volume information. Journal of the History of Ideas, 24(1):1, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707854.

Anonymous:1963:VIb

[Ano63l] Anonymous. Volume information. Journal of the History of Ideas, 24(2):??, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707841.

Anonymous:1963:VIc

[Ano63m] Anonymous. Volume information. Journal of the History of Ideas, 24(3):??, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708207.

Anonymous:1963:VId

[Ano63n] Anonymous. Volume information. Journal of the History of Ideas, 24 (4):599–602, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707975.

Anonymous:1964:BRd

[Ano64a] Anonymous. Book received. Journal of the History of Ideas, 25(4): 599–601, October/December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708190.

Anonymous:1964:BRa

[Ano64b] Anonymous. Books received. Journal of the History of Ideas, 25 (1):130–139, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708092.

Anonymous:1964:BRb

[Ano64c] Anonymous. Books received. Journal of the History of Ideas, 25(2):308–314, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708020.

Anonymous:1964:BRc

[Ano64d] Anonymous. Books received. Journal of the History of Ideas, 25 (3):458–462, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707913.

Anonymous:1964:EN

[Ano64e] Anonymous. Editorial note. Journal of the History of Ideas, 25 (1):140, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708093.

Anonymous:1964:FMa

[Ano64f] Anonymous. Front matter. Journal of the History of Ideas, 25 (1):2, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708081.

Anonymous:1964:FMb

[Ano64g] Anonymous. Front matter. Journal of the History of Ideas, 25(2):??, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708008.

Anonymous:1964:FMc

[Ano64h] Anonymous. Front matter. Journal of the History of Ideas, 25 (3):??, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707901.

Anonymous:1964:FMd

[Ano64i] Anonymous. Front matter. Journal of the History of Ideas, 25 (4):??, October/December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708180.

Anonymous:1964:HIF

[Ano64j] Anonymous. Historia integra: François Baudouin and his conception of history. Journal of the History of Ideas, 25(2):314, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708021.

Anonymous:1964:VIa

[Ano64k] Anonymous. Volume information. Journal of the History of Ideas, 25(1):1, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708080.

Anonymous:1964:VIb

[Ano64l] Anonymous. Volume information. Journal of the History of Ideas, 25(2):??, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708007.

Anonymous:1964:VIc

[Ano64m] Anonymous. Volume information. Journal of the History of Ideas, 25(3):??, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707900.

Anonymous:1964:VId

[Ano64n] Anonymous. Volume information. Journal of the History of Ideas, 25 (4):604–606, October/December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708179.

Anonymous:1965:A

[Ano65a] Anonymous. Acknowledgments. Journal of the History of Ideas, 26 (4):597, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708504.

Anonymous:1965:BRa

[Ano65b] Anonymous. Books received. Journal of the History of Ideas, 26 (1):155–157, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708408.

Anonymous:1965:BRb

[Ano65c] Anonymous. Books received. Journal of the History of Ideas, 26(2):295–301, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708237.

Anonymous:1965:BRc

[Ano65d] Anonymous. Books received. Journal of the History of Ideas, 26 (3):440–446, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708393.

Anonymous:1965:BRd

[Ano65e] Anonymous. Books received. Journal of the History of Ideas, 26(4): 598–601 + 604, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708505.

Anonymous:1965:FMa

[Ano65f] Anonymous. Front matter. Journal of the History of Ideas, 26 (1):2, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708396.

Anonymous:1965:FMb

[Ano65g] Anonymous. Front matter. Journal of the History of Ideas, 26(2):??, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708224.

Anonymous:1965:FMc

[Ano65h] Anonymous. Front matter. Journal of the History of Ideas, 26 (3):??, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708383.

Anonymous:1965:FMd

[Ano65i] Anonymous. Front matter. Journal of the History of Ideas, 26 (4):??, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708492.

Anonymous:1965:VIa

[Ano65j] Anonymous. Volume information. Journal of the History of Ideas, 26(1):1, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708395.

Anonymous:1965:VIb

[Ano65k] Anonymous. Volume information. Journal of the History of Ideas, 26(2):??, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708223.

Anonymous:1965:VIc

[Ano65l] Anonymous. Volume information. Journal of the History of Ideas, 26(3):??, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708382.

Anonymous:1965:VId

[Ano65m] Anonymous. Volume information. Journal of the History of Ideas, 26 (4):605–608, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708491.

Anonymous:1966:A

[Ano66a] Anonymous. Advertisements. Journal of the History of Ideas, 27(1):??, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708317.

Anonymous:1966:BRa

[Ano66b] Anonymous. Books received. Journal of the History of Ideas, 27 (1):150–156, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708316.

Anonymous:1966:BRb

[Ano66c] Anonymous. Books received. Journal of the History of Ideas, 27(2):310–314, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708648.

Anonymous:1966:BRc

[Ano66d] Anonymous. Books received. Journal of the History of Ideas, 27 (3):470–474, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708602.

Anonymous:1966:BRd

[Ano66e] Anonymous. Books received. Journal of the History of Ideas, 27(4): 626–632, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708346.

Anonymous:1966:EMI

[Ano66f] Anonymous. Erratum: Malthus and the idea of progress. Journal of the History of Ideas, 27(2):314, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708649. See [Lev66].

Anonymous:1966:FMa

[Ano66g] Anonymous. Front matter. Journal of the History of Ideas, 27 (1):2, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708304.

Anonymous:1966:FMb

[Ano66h] Anonymous. Front matter. Journal of the History of Ideas, 27(2):??, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708635.

Anonymous:1966:FMc

[Ano66i] Anonymous. Front matter. Journal of the History of Ideas, 27 (3):??, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708588.

Anonymous:1966:FMd

[Ano66j] Anonymous. Front matter. Journal of the History of Ideas, 27 (4):??, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708336.

Anonymous:1966:VIa

[Ano66k] Anonymous. Volume information. Journal of the History of Ideas, 27(1):1, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708303.

Anonymous:1966:VIb

[Ano66l] Anonymous. Volume information. Journal of the History of Ideas, 27(2):??, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708634.

Anonymous:1966:VIc

[Ano66m] Anonymous. Volume information. Journal of the History of Ideas, 27(3):??, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708587.

Anonymous:1966:VId

[Ano66n] Anonymous. Volume information. Journal of the History of Ideas, 27 (4):637–640, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708335.

Anonymous:1967:Aa

[Ano67a] Anonymous. Acknowledgments. Journal of the History of Ideas, 28 (4):620, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708538.

Anonymous:1967:Ab

[Ano67b] Anonymous. Announcements. Journal of the History of Ideas, 28 (4):620, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708539.

Anonymous:1967:BRa

[Ano67c] Anonymous. Books received. Journal of the History of Ideas, 28 (1):133–140, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708488.

Anonymous:1967:BRb

[Ano67d] Anonymous. Books received. Journal of the History of Ideas, 28(2):293–299, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708425.

Anonymous:1967:BRc

[Ano67e] Anonymous. Books received. Journal of the History of Ideas, 28 (3):455–456, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708632.

Anonymous:1967:BRd

[Ano67f] Anonymous. Books received. Journal of the History of Ideas, 28(4): 611–616, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708537.

Anonymous:1967:EPP

[Ano67g] Anonymous. Erratum: Perrenial philosophy: From Agostino Steuco to Leibniz. Journal of the History of Ideas, 28(1):140, January/ March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708489. See

[Sch66].

Anonymous:1967:FMa

[Ano67h] Anonymous. Front matter. Journal of the History of Ideas, 28 (1):2, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708476.

Anonymous:1967:FMb

[Ano67i] Anonymous. Front matter. Journal of the History of Ideas, 28(2):??, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708411.

Anonymous:1967:FMc

[Ano67j] Anonymous. Front matter. Journal of the History of Ideas, 28 (3):??, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708620.

Anonymous:1967:FMd

[Ano67k] Anonymous. Front matter. Journal of the History of Ideas, 28 (4):??, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708523.

Anonymous:1967:VIa

[Ano67l] Anonymous. Volume information. Journal of the History of Ideas, 28(1):1, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708475.

Anonymous:1967:VIb

[Ano67m] Anonymous. Volume information. Journal of the History of Ideas, 28(2):??, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708410.

Anonymous:1967:VIc

[Ano67n] Anonymous. Volume information. Journal of the History of Ideas, 28(3):??, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708619.

Anonymous:1967:VId

[Ano67o] Anonymous. Volume information. Journal of the History of Ideas, 28 (4):617–624, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708522.

Anonymous:1968:Ac

[Ano68a] Anonymous. Acknowledgments. Journal of the History of Ideas, 29 (4):633, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708301.

Anonymous:1968:Aa

[Ano68b] Anonymous. Advertisements. Journal of the History of Ideas, 29(2): ??, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708585.

Anonymous:1968:Ab

[Ano68c] Anonymous. Announcement. Journal of the History of Ideas, 29 (3):464, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708457.

Anonymous:1968:BRa

[Ano68d] Anonymous. Books received. Journal of the History of Ideas, 29 (1):152–154, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708473.

Anonymous:1968:BRb

[Ano68e] Anonymous. Books received. Journal of the History of Ideas, 29(2):311–314, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708584.

Anonymous:1968:BRc

[Ano68f] Anonymous. Books received. Journal of the History of Ideas, 29 (3):465–474, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708458.

Anonymous:1968:BRd

[Ano68g] Anonymous. Books received. Journal of the History of Ideas, 29(4): 623–630, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708300.

Anonymous:1968:FMa

[Ano68h] Anonymous. Front matter. Journal of the History of Ideas, 29 (1):2, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708461.

Anonymous:1968:FMb

[Ano68i] Anonymous. Front matter. Journal of the History of Ideas, 29(2):??, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708573.

Anonymous:1968:FMc

[Ano68j] Anonymous. Front matter. Journal of the History of Ideas, 29 (3):??, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708445.

Anonymous:1968:FMd

[Ano68k] Anonymous. Front matter. Journal of the History of Ideas, 29 (4):??, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708289.

Anonymous:1968:VIa

[Ano68l] Anonymous. Volume information. Journal of the History of Ideas, 29(1):1, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708460.

Anonymous:1968:VIb

[Ano68m] Anonymous. Volume information. Journal of the History of Ideas, 29(2):??, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708572.

Anonymous:1968:VIc

[Ano68n] Anonymous. Volume information. Journal of the History of Ideas, 29(3):??, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708444.

Anonymous:1968:VId

[Ano68o] Anonymous. Volume information. Journal of the History of Ideas, 29 (4):631–634, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708288.

Anonymous:1969:Ab

[Ano69a] Anonymous. Acknowledgments. Journal of the History of Ideas, 30 (4):619, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708617.

Anonymous:1969:Aa

[Ano69b] Anonymous. Advertisements. Journal of the History of Ideas, 30(1):138, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708252.

Anonymous:1969:BRa

[Ano69c] Anonymous. Books received. Journal of the History of Ideas, 30 (1):134–137, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708251.

Anonymous:1969:BRb

[Ano69d] Anonymous. Books received. Journal of the History of Ideas, 30(2):291–296, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708442.

Anonymous:1969:BRc

[Ano69e] Anonymous. Books received. Journal of the History of Ideas, 30 (3):455–457, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708570.

Anonymous:1969:BRd

[Ano69f] Anonymous. Books received. Journal of the History of Ideas, 30(4): 609–614, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708615.

Anonymous:1969:EOF

[Ano69g] Anonymous. Errata: Ophites and fathers. Journal of the History of Ideas, 30(4):619, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708616. See errata [Smi69].

Anonymous:1969:FMa

[Ano69h] Anonymous. Front matter. Journal of the History of Ideas, 30 (1):2, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708240.

Anonymous:1969:FMb

[Ano69i] Anonymous. Front matter. Journal of the History of Ideas, 30(2):??, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708428.

Anonymous:1969:FMc

[Ano69j] Anonymous. Front matter. Journal of the History of Ideas, 30 (3):??, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708558.

Anonymous:1969:FMd

[Ano69k] Anonymous. Front matter. Journal of the History of Ideas, 30 (4):??, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708605.

Anonymous:1969:RBW

[Ano69l] Anonymous. Reply to Bernard Weinberg’s review of my History of Modern Criticism. Journal of the History of Ideas, 30(2):281–282, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708440.

See [Wei69].

Anonymous:1969:VIa

[Ano69m] Anonymous. Volume information. Journal of the History of Ideas, 30(1):1, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708239.

Anonymous:1969:VIb

[Ano69n] Anonymous. Volume information. Journal of the History of Ideas, 30(2):??, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708427.

Anonymous:1969:VIc

[Ano69o] Anonymous. Volume information. Journal of the History of Ideas, 30(3):??, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708557.

Anonymous:1969:VId

[Ano69p] Anonymous. Volume information. Journal of the History of Ideas, 30 (4):615–620, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708604.

Anonymous:1970:Aa

[Ano70a] Anonymous. Acknowledgements. Journal of the History of Ideas, 31(4):637, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708269.

Anonymous:1970:Ab

[Ano70b] Anonymous. Announcement. Journal of the History of Ideas, 31 (4):640, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708271.

Anonymous:1970:BRa

[Ano70c] Anonymous. Books received. Journal of the History of Ideas, 31 (1):152–157, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708380.

Anonymous:1970:BRb

[Ano70d] Anonymous. Books received. Journal of the History of Ideas, 31(2):313–315, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708555.

Anonymous:1970:BRc

[Ano70e] Anonymous. Books received. Journal of the History of Ideas, 31 (3):475–480, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708520.

Anonymous:1970:BRd

[Ano70f] Anonymous. Books received. Journal of the History of Ideas, 31(4): 622–630, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708268.

Anonymous:1970:EJD

[Ano70g] Anonymous. Erratum: John Dewey’s ideas about the Great Depression. Journal of the History of Ideas, 31(4):637, October/ December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708270.

See [Cyw69].

Anonymous:1970:FMa

[Ano70h] Anonymous. Front matter. Journal of the History of Ideas, 31 (1):2, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708365.

Anonymous:1970:FMb

[Ano70i] Anonymous. Front matter. Journal of the History of Ideas, 31(2):??, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708542.

Anonymous:1970:FMc

[Ano70j] Anonymous. Front matter. Journal of the History of Ideas, 31 (3):??, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708508.

Anonymous:1970:FMd

[Ano70k] Anonymous. Front matter. Journal of the History of Ideas, 31 (4):??, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708255.

Anonymous:1970:VIa

[Ano70l] Anonymous. Volume information. Journal of the History of Ideas, 31(1):1, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708364.

Anonymous:1970:VIb

[Ano70m] Anonymous. Volume information. Journal of the History of Ideas, 31(2):??, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708541.

Anonymous:1970:VIc

[Ano70n] Anonymous. Volume information. Journal of the History of Ideas, 31(3):??, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708507.

Anonymous:1970:VId

[Ano70o] Anonymous. Volume information. Journal of the History of Ideas, 31 (4):631–637, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708254.

Anonymous:1971:A

[Ano71a] Anonymous. Acknowledgements. Journal of the History of Ideas, 32(4):638, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708986.

Anonymous:1971:BRa

[Ano71b] Anonymous. Books received. Journal of the History of Ideas, 32(1): 149–152 + 154+156+158+160, January/March 1971. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708333.

Anonymous:1971:BRb

[Ano71c] Anonymous. Books received. Journal of the History of Ideas, 32 (2):311–316 + 318+320, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708286.

Anonymous:1971:BRc

[Ano71d] Anonymous. Books received. Journal of the History of Ideas, 32(3): 471–474 + 476+480, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708362.

Anonymous:1971:BRd

[Ano71e] Anonymous. Books received. Journal of the History of Ideas, 32(4): 631–634, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708985.

Anonymous:1971:FMa

[Ano71f] Anonymous. Front matter. Journal of the History of Ideas, 32 (1):2, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708320.

Anonymous:1971:FMb

[Ano71g] Anonymous. Front matter. Journal of the History of Ideas, 32(2):??, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708274.

Anonymous:1971:FMc

[Ano71h] Anonymous. Front matter. Journal of the History of Ideas, 32 (3):??, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708349.

Anonymous:1971:FMd

[Ano71i] Anonymous. Front matter. Journal of the History of Ideas, 32 (4):??, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708972.

Anonymous:1971:VIa

[Ano71j] Anonymous. Volume information. Journal of the History of Ideas, 32(1):1, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708319.

Anonymous:1971:VIb

[Ano71k] Anonymous. Volume information. Journal of the History of Ideas, 32(2):??, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708273.

Anonymous:1971:VIc

[Ano71l] Anonymous. Volume information. Journal of the History of Ideas, 32(3):??, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708348.

Anonymous:1971:VId

[Ano71m] Anonymous. Volume information. Journal of the History of Ideas, 32 (4):636–640, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708971.

Anonymous:1972:BRa

[Ano72a] Anonymous. Books received. Journal of the History of Ideas, 33 (1):184–186 + 188, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709066.

Anonymous:1972:BRb

[Ano72b] Anonymous. Books received. Journal of the History of Ideas, 33 (2):346 + 348+350–352, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708880.

Anonymous:1972:BRc

[Ano72c] Anonymous. Books received. Journal of the History of Ideas, 33(3): 513 + 515+517+519–520, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709050.

Anonymous:1972:BRd

[Ano72d] Anonymous. Books received. Journal of the History of Ideas, 33 (4):650–656 + 658+660, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708865.

Anonymous:1972:FMa

[Ano72e] Anonymous. Front matter. Journal of the History of Ideas, 33 (1):2, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709053.

Anonymous:1972:FMb

[Ano72f] Anonymous. Front matter. Journal of the History of Ideas, 33(2):??, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708868.

Anonymous:1972:FMc

[Ano72g] Anonymous. Front matter. Journal of the History of Ideas, 33 (3):??, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709038.

Anonymous:1972:FMd

[Ano72h] Anonymous. Front matter. Journal of the History of Ideas, 33 (4):??, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708854.

Anonymous:1972:VIa

[Ano72i] Anonymous. Volume information. Journal of the History of Ideas, 33(1):1, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709052.

Anonymous:1972:VIb

[Ano72j] Anonymous. Volume information. Journal of the History of Ideas, 33(2):??, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708867.

Anonymous:1972:VIc

[Ano72k] Anonymous. Volume information. Journal of the History of Ideas, 33(3):??, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709037.

Anonymous:1972:VId

[Ano72l] Anonymous. Volume information. Journal of the History of Ideas, 33 (4):661–664, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708853.

Anonymous:1973:A

[Ano73a] Anonymous. Acknowledgments. Journal of the History of Ideas, 34 (4):687, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708902.

Anonymous:1973:BRa

[Ano73b] Anonymous. Books received. Journal of the History of Ideas, 34(2):315–320, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708735.

Anonymous:1973:BRb

[Ano73c] Anonymous. Books received. Journal of the History of Ideas, 34 (3):490–496, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708969.

Anonymous:1973:BRc

[Ano73d] Anonymous. Books received. Journal of the History of Ideas, 34(4): 674–680 + 682, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708901.

Anonymous:1973:EAM

[Ano73e] Anonymous. Erratum: Aristotle and modern biology. Journal of the History of Ideas, 34(4):687, October/December 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708903. See [Gre72b].

Anonymous:1973:FMa

[Ano73f] Anonymous. Front matter. Journal of the History of Ideas, 34 (1):2, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708940.

Anonymous:1973:FMb

[Ano73g] Anonymous. Front matter. Journal of the History of Ideas, 34(2):??, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708723.

Anonymous:1973:FMc

[Ano73h] Anonymous. Front matter. Journal of the History of Ideas, 34 (3):??, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708955.

Anonymous:1973:FMd

[Ano73i] Anonymous. Front matter. Journal of the History of Ideas, 34 (4):590, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708883.

Anonymous:1973:VIa

[Ano73j] Anonymous. Volume information. Journal of the History of Ideas, 34(1):1, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708939.

Anonymous:1973:VIb

[Ano73k] Anonymous. Volume information. Journal of the History of Ideas, 34(2):??, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708722.

Anonymous:1973:VIc

[Ano73l] Anonymous. Volume information. Journal of the History of Ideas, 34(3):??, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708954.

Anonymous:1973:VId

[Ano73m] Anonymous. Volume information. Journal of the History of Ideas, 34 (4):684–688, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708882.

Anonymous:1974:BRa

[Ano74a] Anonymous. Books received. Journal of the History of Ideas, 35 (1):170 + 172–176, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708753.

Anonymous:1974:BRb

[Ano74b] Anonymous. Books received. Journal of the History of Ideas, 35(2): 357–358 + 360+362+364–368, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708770.

Anonymous:1974:BRc

[Ano74c] Anonymous. Books received. Journal of the History of Ideas, 35 (3):538 + 540–544, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708801.

Anonymous:1974:BRd

[Ano74d] Anonymous. Books received. Journal of the History of Ideas, 35(4): 715–720, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709097.

Anonymous:1974:FMa

[Ano74e] Anonymous. Front matter. Journal of the History of Ideas, 35 (1):2–62, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708738.

Anonymous:1974:FMb

[Ano74f] Anonymous. Front matter. Journal of the History of Ideas, 35(2):??, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708756.

Anonymous:1974:FMc

[Ano74g] Anonymous. Front matter. Journal of the History of Ideas, 35 (3):??, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708788.

Anonymous:1974:FMd

[Ano74h] Anonymous. Front matter. Journal of the History of Ideas, 35 (4):578, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709084.

Anonymous:1974:VIa

[Ano74i] Anonymous. Volume information. Journal of the History of Ideas, 35(1):1, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708737.

Anonymous:1974:VIb

[Ano74j] Anonymous. Volume information. Journal of the History of Ideas, 35(2):??, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708755.

Anonymous:1974:VIc

[Ano74k] Anonymous. Volume information. Journal of the History of Ideas, 35(3):??, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708787.

Anonymous:1974:VId

[Ano74l] Anonymous. Volume information. Journal of the History of Ideas, 35 (4):721–728, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709083.

Anonymous:1975:A

[Ano75a] Anonymous. Acknowledgments. Journal of the History of Ideas, 36 (4):758, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709003.

Anonymous:1975:BRa

[Ano75b] Anonymous. Books received. Journal of the History of Ideas, 36 (1):185–186 + 188, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709021.

Anonymous:1975:BRb

[Ano75c] Anonymous. Books received. Journal of the History of Ideas, 36 (2):378 + 380+382–383, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708936.

Anonymous:1975:BRc

[Ano75d] Anonymous. Books received. Journal of the History of Ideas, 36 (3):573–576, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708666.

Anonymous:1975:BRd

[Ano75e] Anonymous. Books received. Journal of the History of Ideas, 36(4): 747 + 750–754, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709002.

Anonymous:1975:CLP

[Ano75f] Anonymous. Clark Library Professorship. Journal of the History of Ideas, 36(2):384, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708937.

Anonymous:1975:ENH

[Ano75g] Anonymous. Errata: The Nazi Holocaust as a persisting trauma for the non-Jewish mind. Journal of the History of Ideas, 36(4):758, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709004.

See [Fac75].

Anonymous:1975:FMa

[Ano75h] Anonymous. Front matter. Journal of the History of Ideas, 36 (1):2, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709007.

Anonymous:1975:FMb

[Ano75i] Anonymous. Front matter. Journal of the History of Ideas, 36(2):??, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708923.

Anonymous:1975:FMc

[Ano75j] Anonymous. Front matter. Journal of the History of Ideas, 36 (3):??, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708652.

Anonymous:1975:FMd

[Ano75k] Anonymous. Front matter. Journal of the History of Ideas, 36 (4):??, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708989.

Anonymous:1975:VIa

[Ano75l] Anonymous. Volume information. Journal of the History of Ideas, 36(1):1, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709006.

Anonymous:1975:VIb

[Ano75m] Anonymous. Volume information. Journal of the History of Ideas, 36(2):??, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708922.

Anonymous:1975:VIc

[Ano75n] Anonymous. Volume information. Journal of the History of Ideas, 36(3):??, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708651.

Anonymous:1975:VId

[Ano75o] Anonymous. Volume information. Journal of the History of Ideas, 36 (4):755–758, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708988.

Anonymous:1976:BRa

[Ano76a] Anonymous. Books received. Journal of the History of Ideas, 37 (1):185–191, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708720.

Anonymous:1976:BRb

[Ano76b] Anonymous. Books received. Journal of the History of Ideas, 37(2):377–382, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708833.

Anonymous:1976:BRc

[Ano76c] Anonymous. Books received. Journal of the History of Ideas, 37 (3):556–557 + 560+562+564+566–568, July/September 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708817.

Anonymous:1976:BRd

[Ano76d] Anonymous. Books received. Journal of the History of Ideas, 37 (4):721–722 + 724+726+728–730, October/December 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709035.

Anonymous:1976:FMa

[Ano76e] Anonymous. Front matter. Journal of the History of Ideas, 37 (1):2, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708705.

Anonymous:1976:FMc

[Ano76f] Anonymous. Front matter. Journal of the History of Ideas, 37(2):194, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708820.

Anonymous:1976:FMb

[Ano76g] Anonymous. Front matter. Journal of the History of Ideas, 37(2):??, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/3857203.

Anonymous:1976:FMd

[Ano76h] Anonymous. Front matter. Journal of the History of Ideas, 37 (3):??, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708804.

Anonymous:1976:FMe

[Ano76i] Anonymous. Front matter. Journal of the History of Ideas, 37 (4):??, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709024.

Anonymous:1976:VIa

[Ano76j] Anonymous. Volume information. Journal of the History of Ideas, 37(1):1, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708704.

Anonymous:1976:VIb

[Ano76k] Anonymous. Volume information. Journal of the History of Ideas, 37(2):193, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708819.

Anonymous:1976:VIc

[Ano76l] Anonymous. Volume information. Journal of the History of Ideas, 37(3):??, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708803.

Anonymous:1976:VId

[Ano76m] Anonymous. Volume information. Journal of the History of Ideas, 37 (4):731–736, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709023.

Anonymous:1977:BRa

[Ano77a] Anonymous. Books received. Journal of the History of Ideas, 38 (1):185 + 188–192, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708851.

Anonymous:1977:BRb

[Ano77b] Anonymous. Books received. Journal of the History of Ideas, 38 (2):365–367 + 370–376, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708920.

Anonymous:1977:BRc

[Ano77c] Anonymous. Books received. Journal of the History of Ideas, 38 (3):549 + 552–554, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708683.

Anonymous:1977:BRd

[Ano77d] Anonymous. Books received. Journal of the History of Ideas, 38(4): 728–730 + 732+734–742, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708702.

Anonymous:1977:ETZ

[Ano77e] Anonymous. Erratum: Thomas, Znaniecki, and Popper on falsification. Journal of the History of Ideas, 38(1):184, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708850. See [McL76a].

Anonymous:1977:FMa

[Ano77f] Anonymous. Front matter. Journal of the History of Ideas, 38 (1):2, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708836.

Anonymous:1977:FMb

[Ano77g] Anonymous. Front matter. Journal of the History of Ideas, 38(2):??, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708906.

Anonymous:1977:FMc

[Ano77h] Anonymous. Front matter. Journal of the History of Ideas, 38 (3):??, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708669.

Anonymous:1977:FMd

[Ano77i] Anonymous. Front matter. Journal of the History of Ideas, 38 (4):??, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708686.

Anonymous:1977:VIa

[Ano77j] Anonymous. Volume information. Journal of the History of Ideas, 38(1):1, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708835.

Anonymous:1977:VIb

[Ano77k] Anonymous. Volume information. Journal of the History of Ideas, 38(2):??, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708905.

Anonymous:1977:VIc

[Ano77l] Anonymous. Volume information. Journal of the History of Ideas, 38(3):??, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708668.

Anonymous:1977:VId

[Ano77m] Anonymous. Volume information. Journal of the History of Ideas, 38 (4):743–746, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708685.

Anonymous:1978:BRa

[Ano78a] Anonymous. Books received. Journal of the History of Ideas, 39(1): 158 + 160+162+164+166–168, January/March 1978. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709081.

Anonymous:1978:BRb

[Ano78b] Anonymous. Books received. Journal of the History of Ideas, 39(2): 335–338 + 340+342–344, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708785.

Anonymous:1978:BRc

[Ano78c] Anonymous. Books received. Journal of the History of Ideas, 39 (3):521–534, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709396.

Anonymous:1978:BRd

[Ano78d] Anonymous. Books received. Journal of the History of Ideas, 39(4): 686–694, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709452.

Anonymous:1978:FMa

[Ano78e] Anonymous. Front matter. Journal of the History of Ideas, 39 (1):1–165, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709068.

Anonymous:1978:FMb

[Ano78f] Anonymous. Front matter. Journal of the History of Ideas, 39 (2):198–341, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708772.

Anonymous:1978:FMc

[Ano78g] Anonymous. Front matter. Journal of the History of Ideas, 39 (3):434–535, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709381.

Anonymous:1978:FMd

[Ano78h] Anonymous. Front matter. Journal of the History of Ideas, 39(4): 578–699, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709440.

Anonymous:1978:VI

[Ano78i] Anonymous. Volume information. Journal of the History of Ideas, 39 (4):695–700, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709439.

Anonymous:1979:BRa

[Ano79a] Anonymous. Books received. Journal of the History of Ideas, 40 (3):491–504, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709254.

Anonymous:1979:BRb

[Ano79b] Anonymous. Books received. Journal of the History of Ideas, 40(4): 675–687, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709366.

Anonymous:1979:FMa

[Ano79c] Anonymous. Front matter. Journal of the History of Ideas, 40 (1):42–156, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709256.

Anonymous:1979:FMb

[Ano79d] Anonymous. Front matter. Journal of the History of Ideas, 40 (2):206–314, April/June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709146.

Anonymous:1979:FMc

[Ano79e] Anonymous. Front matter. Journal of the History of Ideas, 40 (3):444–503, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709240.

Anonymous:1979:FMd

[Ano79f] Anonymous. Front matter. Journal of the History of Ideas, 40(4): 562–691, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709355.

Anonymous:1979:VI

[Ano79g] Anonymous. Volume information. Journal of the History of Ideas, 40 (4):688–692, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709354.

Anonymous:1980:Aa

[Ano80a] Anonymous. Announcement. Journal of the History of Ideas, 41(1):174, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709113.

Anonymous:1980:Ab

[Ano80b] Anonymous. Announcement. Journal of the History of Ideas, 41(1):176, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709115.

Anonymous:1980:BRa

[Ano80c] Anonymous. Books received. Journal of the History of Ideas, 41 (1):160–171, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709112.

Anonymous:1980:BRb

[Ano80d] Anonymous. Books received. Journal of the History of Ideas, 41(2):344–352, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709467.

Anonymous:1980:BRc

[Ano80e] Anonymous. Books received. Journal of the History of Ideas, 41 (3):512–525, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709410.

Anonymous:1980:BRd

[Ano80f] Anonymous. Books received. Journal of the History of Ideas, 41(4): 699–710, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709282.

Anonymous:1980:FAa

[Ano80g] Anonymous. Forthcoming articles (1980–1981). Journal of the History of Ideas, 41(1):175–176, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709114.

Anonymous:1980:FAb

[Ano80h] Anonymous. Forthcoming articles (1980–1981). Journal of the History of Ideas, 41(3):526–527, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709411.

Anonymous:1980:FAc

[Ano80i] Anonymous. Forthcoming articles (1981–1982). Journal of the History of Ideas, 41(4):719–720, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709283.

Anonymous:1980:FMa

[Ano80j] Anonymous. Front matter. Journal of the History of Ideas, 41 (1):28–173, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709099.

Anonymous:1980:FMb

[Ano80k] Anonymous. Front matter. Journal of the History of Ideas, 41 (2):198–343, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709454.

Anonymous:1980:FMc

[Ano80l] Anonymous. Front matter. Journal of the History of Ideas, 41 (3):380–451, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709398.

Anonymous:1980:FMd

[Ano80m] Anonymous. Front matter. Journal of the History of Ideas, 41(4): 600–718, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709272.

Anonymous:1980:VI

[Ano80n] Anonymous. Volume information. Journal of the History of Ideas, 41 (4):712–717, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709271.

Anonymous:1981:BRa

[Ano81a] Anonymous. Books received. Journal of the History of Ideas, 42 (1):173–184, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709425.

Anonymous:1981:BRb

[Ano81b] Anonymous. Books received. Journal of the History of Ideas, 42(2):355–368, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709325.

Anonymous:1981:BRc

[Ano81c] Anonymous. Books received. Journal of the History of Ideas, 42 (3):552–568, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709193.

Anonymous:1981:BRd

[Ano81d] Anonymous. Books received. Journal of the History of Ideas, 42(4): 738–749, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709131.

Anonymous:1981:EMG

[Ano81e] Anonymous. Errata: In memory of George Boas. Journal of the History of Ideas, 42(3):568, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709195. See [Gom81].

Anonymous:1981:EMI

[Ano81f] Anonymous. Errata: Medieval irony. Journal of the History of Ideas, 42(3):568, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709194. See [Rei81].

Anonymous:1981:FMa

[Ano81g] Anonymous. Front matter. Journal of the History of Ideas, 42 (1):28–162, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709413.

Anonymous:1981:FMb

[Ano81h] Anonymous. Front matter. Journal of the History of Ideas, 42 (2):208–359, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709315.

Anonymous:1981:FMc

[Ano81i] Anonymous. Front matter. Journal of the History of Ideas, 42 (3):388–456, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709181.

Anonymous:1981:FMd

[Ano81j] Anonymous. Front matter. Journal of the History of Ideas, 42(4): 586–751, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709118.

Anonymous:1981:VI

[Ano81k] Anonymous. Volume information. Journal of the History of Ideas, 42 (4):752–755, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709117.

Anonymous:1982:BRa

[Ano82a] Anonymous. Books received. Journal of the History of Ideas, 43 (1):164–176, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709168.

Anonymous:1982:BRb

[Ano82b] Anonymous. Books received. Journal of the History of Ideas, 43(2):340–352, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709209.

Anonymous:1982:BRc

[Ano82c] Anonymous. Books received. Journal of the History of Ideas, 43 (3):519–528, July/September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709437.

Anonymous:1982:BRd

[Ano82d] Anonymous. Books received. Journal of the History of Ideas, 43(4): 690–698, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709352.

Anonymous:1982:FMa

[Ano82e] Anonymous. Front matter. Journal of the History of Ideas, 43 (1):152–163, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709157.

Anonymous:1982:FMb

[Ano82f] Anonymous. Front matter. Journal of the History of Ideas, 43 (2):194–296, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709197.

Anonymous:1982:FMc

[Ano82g] Anonymous. Front matter. Journal of the History of Ideas, 43 (3):378–516, July/September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709427.

Anonymous:1982:FMd

[Ano82h] Anonymous. Front matter. Journal of the History of Ideas, 43(4): 554–672, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709341.

Anonymous:1982:VI

[Ano82i] Anonymous. Volume information. Journal of the History of Ideas, 43 (4):699–704, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709340.

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. Journal of the History of Ideas, 44 (4):598, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709217.

Anonymous:1983:BRa

[Ano83b] Anonymous. Books received. Journal of the History of Ideas, 44 (1):165–176, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709313.

Anonymous:1983:BRb

[Ano83c] Anonymous. Books received. Journal of the History of Ideas, 44(2):337–352, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709144.

Anonymous:1983:BRc

[Ano83d] Anonymous. Books received. Journal of the History of Ideas, 44 (3):507–520, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709179.

Anonymous:1983:BRd

[Ano83e] Anonymous. Books received. Journal of the History of Ideas, 44(4): 694–708, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709225.

Anonymous:1983:EWR

[Ano83f] Anonymous. Erratum: The works of La Rochefoucauld in relation to Machiavellian ideas of morals an politics. Journal of the History of Ideas, 44(4):712, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709226. See [Bak83].

Anonymous:1983:FMa

[Ano83g] Anonymous. Front matter. Journal of the History of Ideas, 44 (1):18–167, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709300.

Anonymous:1983:FMb

[Ano83h] Anonymous. Front matter. Journal of the History of Ideas, 44 (2):230–310, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709133.

Anonymous:1983:FMc

[Ano83i] Anonymous. Front matter. Journal of the History of Ideas, 44 (3):368–454, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709170.

Anonymous:1983:FMd

[Ano83j] Anonymous. Front matter. Journal of the History of Ideas, 44(4): 578–693, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709212.

Anonymous:1983:VI

[Ano83k] Anonymous. Volume information. Journal of the History of Ideas, 44 (4):709–712, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709211.

Anonymous:1984:BRa

[Ano84a] Anonymous. Books received. Journal of the History of Ideas, 45 (1):162–176, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709338.

Anonymous:1984:BRb

[Ano84b] Anonymous. Books received. Journal of the History of Ideas, 45(2):328–335, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709297.

Anonymous:1984:BRc

[Ano84c] Anonymous. Books received. Journal of the History of Ideas, 45 (3):483–496, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709238.

Anonymous:1984:BRd

[Ano84d] Anonymous. Books received. Journal of the History of Ideas, 45(4): 625–627, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709378.

Anonymous:1984:ENB

[Ano84e] Anonymous. Errata: Nietzsche and Buddhism. Journal of the History of Ideas, 45(2):336, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709298. See [Elm83].

Anonymous:1984:ECC

[Ano84f] Anonymous. Errata: The concept of class in French culture: Prior to the revolution. Journal of the History of Ideas, 45(4):628, October/ December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709379.

See [Clo84].

Anonymous:1984:FMa

[Ano84g] Anonymous. Front matter. Journal of the History of Ideas, 45 (1):24–82, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709327.

Anonymous:1984:FMb

[Ano84h] Anonymous. Front matter. Journal of the History of Ideas, 45 (2):302–308, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709285.

Anonymous:1984:FMc

[Ano84i] Anonymous. Front matter. Journal of the History of Ideas, 45 (3):360–378, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709228.

Anonymous:1984:FMd

[Ano84j] Anonymous. Front matter. Journal of the History of Ideas, 45(4): 548–624, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709369.

Anonymous:1984:VI

[Ano84k] Anonymous. Volume information. Journal of the History of Ideas, 45 (4):629–632, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709368.

Anonymous:1985:BRa

[Ano85a] Anonymous. Books received. Journal of the History of Ideas, 46 (1):133–144, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709781.

Anonymous:1985:BRb

[Ano85b] Anonymous. Books received. Journal of the History of Ideas, 46(2): 297 + 300–304, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709641.

Anonymous:1985:BRc

[Ano85c] Anonymous. Books received. Journal of the History of Ideas, 46 (3):455–464, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709480.

Anonymous:1985:BRd

[Ano85d] Anonymous. Books received. Journal of the History of Ideas, 46(4): 618–626, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709549.

Anonymous:1985:EKE

[Ano85e] Anonymous. Erratum: Kepler’s early writings. Journal of the History of Ideas, 46(4):628, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709550. See [Ros85a].

Anonymous:1985:FMa

[Ano85f] Anonymous. Front matter. Journal of the History of Ideas, 46 (1):32–88, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709771.

Anonymous:1985:FMb

[Ano85g] Anonymous. Front matter. Journal of the History of Ideas, 46 (2):166–299, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709631.

Anonymous:1985:FMc

[Ano85h] Anonymous. Front matter. Journal of the History of Ideas, 46 (3):326–434, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709469.

Anonymous:1985:FMd

[Ano85i] Anonymous. Front matter. Journal of the History of Ideas, 46(4): 486–617, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709539.

Anonymous:1985:VI

[Ano85j] Anonymous. Volume information. Journal of the History of Ideas, 46 (4):627–630, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709538.

Anonymous:1986:BRa

[Ano86a] Anonymous. Books received. Journal of the History of Ideas, 47 (1):164–174, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709605.

Anonymous:1986:BRb

[Ano86b] Anonymous. Books received. Journal of the History of Ideas, 47(2):329–338, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709819.

Anonymous:1986:BRc

[Ano86c] Anonymous. Books received. Journal of the History of Ideas, 47 (3):513–524, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709669.

Anonymous:1986:BRd

[Ano86d] Anonymous. Books received. Journal of the History of Ideas, 47(4): 673–686, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709727.

Anonymous:1986:ERH

[Ano86e] Anonymous. Errata: Remembering Hulme. Journal of the History of Ideas, 47(1):161, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709603. See [Shu85].

Anonymous:1986:ETT

[Ano86f] Anonymous. Errata: Tocqueville’s two revolutions. Journal of the History of Ideas, 47(1):161, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709604. See [Kah85].

Anonymous:1986:EBP

[Ano86g] Anonymous. Erratum: Bringing philosophy down from the heavens. Journal of the History of Ideas, 47(4):686, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709728. See [Ber86a].

Anonymous:1986:FMa

[Ano86h] Anonymous. Front matter. Journal of the History of Ideas, 47 (1):36–163, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709591.

Anonymous:1986:FMb

[Ano86i] Anonymous. Front matter. Journal of the History of Ideas, 47 (2):214–328, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709808.

Anonymous:1986:FMc

[Ano86j] Anonymous. Front matter. Journal of the History of Ideas, 47 (3):360–496, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709656.

Anonymous:1986:FMd

[Ano86k] Anonymous. Front matter. Journal of the History of Ideas, 47(4): 546–610, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709715.

Anonymous:1986:Na

[Ano86l] Anonymous. Notices. Journal of the History of Ideas, 47(2):321–322, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709818.

Anonymous:1986:Nb

[Ano86m] Anonymous. Notices. Journal of the History of Ideas, 47(3): 511–512, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709668.

Anonymous:1986:Nc

[Ano86n] Anonymous. Notices. Journal of the History of Ideas, 47(4):672, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709726.

Anonymous:1986:VI

[Ano86o] Anonymous. Volume information. Journal of the History of Ideas, 47 (4):687–690, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709714.

Anonymous:1987:BRa

[Ano87a] Anonymous. Books received. Journal of the History of Ideas, 48 (1):175–184, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709617.

Anonymous:1987:BRb

[Ano87b] Anonymous. Books received. Journal of the History of Ideas, 48(2):361–368, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709563.

Anonymous:1987:BRc

[Ano87c] Anonymous. Books received. Journal of the History of Ideas, 48 (3):533–547, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709767.

Anonymous:1987:BRd

[Ano87d] Anonymous. Books received. Journal of the History of Ideas, 48(4): 725–736, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709698.

Anonymous:1987:ELU

[Ano87e] Anonymous. Erratum: Lovejoy’s unexplored option. Journal of the History of Ideas, 48(3):547, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709768. See [Oak87].

Anonymous:1987:FMa

[Ano87f] Anonymous. Front matter. Journal of the History of Ideas, 48 (1):22–174, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709607.

Anonymous:1987:FMb

[Ano87g] Anonymous. Front matter. Journal of the History of Ideas, 48 (2):246–360, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709552.

Anonymous:1987:FMc

[Ano87h] Anonymous. Front matter. Journal of the History of Ideas, 48 (3):452–520, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709757.

Anonymous:1987:FMd

[Ano87i] Anonymous. Front matter. Journal of the History of Ideas, 48(4): 594–718, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709686.

Anonymous:1987:Na

[Ano87j] Anonymous. Notices. Journal of the History of Ideas, 48(3):548, July/ September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709769.

Anonymous:1987:Nb

[Ano87k] Anonymous. Notices. Journal of the History of Ideas, 48(4): 723–724, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709697.

Anonymous:1987:VI

[Ano87l] Anonymous. Volume information. Journal of the History of Ideas, 48 (4):737–740, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709685.

Anonymous:1988:BRa

[Ano88a] Anonymous. Books received. Journal of the History of Ideas, 49 (1):173–180, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709712.

Anonymous:1988:BRb

[Ano88b] Anonymous. Books received. Journal of the History of Ideas, 49(2):353–362, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709506.

Anonymous:1988:BRc

[Ano88c] Anonymous. Books received. Journal of the History of Ideas, 49 (3):531–539, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709492.

Anonymous:1988:BRd

[Ano88d] Anonymous. Books received. Journal of the History of Ideas, 49(4): 692–702, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709681.

Anonymous:1988:ETD

[Ano88e] Anonymous. Errata: Thomas Davidson’s apeirotheism and its influence on William James and John Dewey. Journal of the History of Ideas, 49(2):363, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709508. See [DeA87].

Anonymous:1988:EWW

[Ano88f] Anonymous. Erratum: Wittgenstein’s “Wonderful Life”. Journal of the History of Ideas, 49(4):704, October/December 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709683. See [Joh88].

Anonymous:1988:FMa

[Ano88g] Anonymous. Front matter. Journal of the History of Ideas, 49 (1):46–172, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709700.

Anonymous:1988:FMb

[Ano88h] Anonymous. Front matter. Journal of the History of Ideas, 49 (2):228–352, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709495.

Anonymous:1988:FMc

[Ano88i] Anonymous. Front matter. Journal of the History of Ideas, 49 (3):450–530, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709482.

Anonymous:1988:FMd

[Ano88j] Anonymous. Front matter. Journal of the History of Ideas, 49(4): 586–691, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709672.

Anonymous:1988:Na

[Ano88k] Anonymous. Notices. Journal of the History of Ideas, 49(1):171, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709711.

Anonymous:1988:Nb

[Ano88l] Anonymous. Notices. Journal of the History of Ideas, 49(2):363, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709507.

Anonymous:1988:Nc

[Ano88m] Anonymous. Notices. Journal of the History of Ideas, 49(3): 540, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709493.

Anonymous:1988:Nd

[Ano88n] Anonymous. Notices. Journal of the History of Ideas, 49(4): 703–704, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709682.

Anonymous:1988:VI

[Ano88o] Anonymous. Volume information. Journal of the History of Ideas, 49 (4):705–708, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709671.

Anonymous:1989:BRa

[Ano89a] Anonymous. Books received. Journal of the History of Ideas, 50 (1):167–176, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709792.

Anonymous:1989:BRb

[Ano89b] Anonymous. Books received. Journal of the History of Ideas, 50(2):339–349, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709741.

Anonymous:1989:BRc

[Ano89c] Anonymous. Books received. Journal of the History of Ideas, 50 (3):513–523, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709576.

Anonymous:1989:BRd

[Ano89d] Anonymous. Books received. Journal of the History of Ideas, 50(4): 681–691, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709806.

Anonymous:1989:FMa

[Ano89e] Anonymous. Front matter. Journal of the History of Ideas, 50 (1):30–166, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709783.

Anonymous:1989:FMb

[Ano89f] Anonymous. Front matter. Journal of the History of Ideas, 50 (2):226–308, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709730.

Anonymous:1989:FMc

[Ano89g] Anonymous. Front matter. Journal of the History of Ideas, 50 (3):408–490, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709565.

Anonymous:1989:FMd

[Ano89h] Anonymous. Front matter. Journal of the History of Ideas, 50(4): 548–692, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709795.

Anonymous:1989:Na

[Ano89i] Anonymous. Notices. Journal of the History of Ideas, 50(2):337–338, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709740.

Anonymous:1989:Nb

[Ano89j] Anonymous. Notices. Journal of the History of Ideas, 50(3): 511–512, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709575.

Anonymous:1989:Nc

[Ano89k] Anonymous. Notices. Journal of the History of Ideas, 50(4): 679–680, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709805.

Anonymous:1989:VI

[Ano89l] Anonymous. Volume information. Journal of the History of Ideas, 50 (4):693–695, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709794.

Anonymous:1990:BRa

[Ano90a] Anonymous. Books received. Journal of the History of Ideas, 51 (1):164–174, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709754.

Anonymous:1990:BRb

[Ano90b] Anonymous. Books received. Journal of the History of Ideas, 51(2):341–352, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709521.

Anonymous:1990:BRc

[Ano90c] Anonymous. Books received. Journal of the History of Ideas, 51 (3):519–527, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709629.

Anonymous:1990:BRd

[Ano90d] Anonymous. Books received. Journal of the History of Ideas, 51(4): 689–697, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709654.

Anonymous:1990:FMa

[Ano90e] Anonymous. Front matter. Journal of the History of Ideas, 51 (1):26–161, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709743.

Anonymous:1990:FMb

[Ano90f] Anonymous. Front matter. Journal of the History of Ideas, 51 (2):198–340, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709510.

Anonymous:1990:FMc

[Ano90g] Anonymous. Front matter. Journal of the History of Ideas, 51 (3):378–516, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709619.

Anonymous:1990:FMd

[Ano90h] Anonymous. Front matter. Journal of the History of Ideas, 51(4): 598–700, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709644.

Anonymous:1990:Na

[Ano90i] Anonymous. Notices. Journal of the History of Ideas, 51(1): 162–163, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709753.

Anonymous:1990:Nb

[Ano90j] Anonymous. Notices. Journal of the History of Ideas, 51(2):336–337, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709520.

Anonymous:1990:Nc

[Ano90k] Anonymous. Notices. Journal of the History of Ideas, 51(3): 517–518, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709628.

Anonymous:1990:Nd

[Ano90l] Anonymous. Notices. Journal of the History of Ideas, 51(4): 685–688, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709653.

Anonymous:1990:OJF

[Ano90m] Anonymous. Obituary: John Fisher 1922–1989. Journal of the History of Ideas, 51(1):175, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709755.

Anonymous:1990:VI

[Ano90n] Anonymous. Volume information. Journal of the History of Ideas, 51 (4):701–703, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709643.

Anonymous:1991:BRa

[Ano91a] Anonymous. Books received. Journal of the History of Ideas, 52 (1):169–176, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709589.

Anonymous:1991:BRb

[Ano91b] Anonymous. Books received. Journal of the History of Ideas, 52(2):344–352, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709536.

Anonymous:1991:BRc

[Ano91c] Anonymous. Books received. Journal of the History of Ideas, 52 (3):525–532, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710053.

Anonymous:1991:BRd

[Ano91d] Anonymous. Books received. Journal of the History of Ideas, 52(4): 695–705, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709974.

Anonymous:1991:FMa

[Ano91e] Anonymous. Front matter. Journal of the History of Ideas, 52 (1):18–139, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709578.

Anonymous:1991:FMb

[Ano91f] Anonymous. Front matter. Journal of the History of Ideas, 52 (2):220–343, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709523.

Anonymous:1991:FMc

[Ano91g] Anonymous. Front matter. Journal of the History of Ideas, 52 (3):372–508, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710041.

Anonymous:1991:FMd

[Ano91h] Anonymous. Front matter. Journal of the History of Ideas, 52(4): 562–664, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709964.

Anonymous:1991:Na

[Ano91i] Anonymous. Notices. Journal of the History of Ideas, 52(1): 167–168, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709588.

Anonymous:1991:Nb

[Ano91j] Anonymous. Notices. Journal of the History of Ideas, 52(2):338–339, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709533.

Anonymous:1991:Nc

[Ano91k] Anonymous. Notices. Journal of the History of Ideas, 52(3): 521–523, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710051.

Anonymous:1991:Nd

[Ano91l] Anonymous. Notices. Journal of the History of Ideas, 52(4): 692–694, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709973.

Anonymous:1991:VI

[Ano91m] Anonymous. Volume information. Journal of the History of Ideas, 52 (4):706–708, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709963.

Anonymous:1992:A

[Ano92a] Anonymous. Announcement. Journal of the History of Ideas, 53 (4):714, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709947.

Anonymous:1992:BRa

[Ano92b] Anonymous. Books received. Journal of the History of Ideas, 53 (1):165–176, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709917.

Anonymous:1992:BRb

[Ano92c] Anonymous. Books received. Journal of the History of Ideas, 53(2):340–351, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709879.

Anonymous:1992:BRc

[Ano92d] Anonymous. Books received. Journal of the History of Ideas, 53 (3):521–528, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709893.

Anonymous:1992:BRd

[Ano92e] Anonymous. Books received. Journal of the History of Ideas, 53(4): 705–713, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709946.

Anonymous:1992:FMa

[Ano92f] Anonymous. Front matter. Journal of the History of Ideas, 53 (1):24–147, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709907.

Anonymous:1992:FMb

[Ano92g] Anonymous. Front matter. Journal of the History of Ideas, 53 (2):208–332, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709869.

Anonymous:1992:FMc

[Ano92h] Anonymous. Front matter. Journal of the History of Ideas, 53 (3):480, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709881.

Anonymous:1992:FMd

[Ano92i] Anonymous. Front matter. Journal of the History of Ideas, 53(4): 552–704, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709933.

Anonymous:1992:Na

[Ano92j] Anonymous. Notices. Journal of the History of Ideas, 53(1):164, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709916.

Anonymous:1992:Nb

[Ano92k] Anonymous. Notices. Journal of the History of Ideas, 53(2):338–339, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709878.

Anonymous:1992:Nc

[Ano92l] Anonymous. Notices. Journal of the History of Ideas, 53(3): 519–520, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709892.

Anonymous:1992:Nd

[Ano92m] Anonymous. Notices. Journal of the History of Ideas, 53(4): 703, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709945.

Anonymous:1992:VI

[Ano92n] Anonymous. Volume information. Journal of the History of Ideas, 53 (4):715–717, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709932.

Anonymous:1993:BRa

[Ano93a] Anonymous. Books received. Journal of the History of Ideas, 54(1):167–175, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709867.

Anonymous:1993:BRb

[Ano93b] Anonymous. Books received. Journal of the History of Ideas, 54(2): 333–341, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709987.

Anonymous:1993:BRc

[Ano93c] Anonymous. Books received. Journal of the History of Ideas, 54(3): 521–529, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710028.

Anonymous:1993:BRd

[Ano93d] Anonymous. Books received. Journal of the History of Ideas, 54(4):699–705, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709831.

Anonymous:1993:FMa

[Ano93e] Anonymous. Front matter. Journal of the History of Ideas, 54(1):118– 166, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709856.

Anonymous:1993:FMb

[Ano93f] Anonymous. Front matter. Journal of the History of Ideas, 54(2): 220, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709976.

Anonymous:1993:FMc

[Ano93g] Anonymous. Front matter. Journal of the History of Ideas, 54(3): 380, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710017.

Anonymous:1993:FMd

[Ano93h] Anonymous. Front matter. Journal of the History of Ideas, 54(4):560– 584, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709822.

Anonymous:1993:Na

[Ano93i] Anonymous. Notices. Journal of the History of Ideas, 54(1):165, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709866.

Anonymous:1993:Nb

[Ano93j] Anonymous. Notices. Journal of the History of Ideas, 54(2):331–332, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709986.

Anonymous:1993:Nc

[Ano93k] Anonymous. Notices. Journal of the History of Ideas, 54(3):519–520, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2710027.

Anonymous:1993:Nd

[Ano93l] Anonymous. Notices. Journal of the History of Ideas, 54(4):697, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709830.

Anonymous:1993:VI

[Ano93m] Anonymous. Volume information. Journal of the History of Ideas, 54(4):707–710, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709821.

Anonymous:1994:BRa

[Ano94a] Anonymous. Books received. Journal of the History of Ideas, 55(1):179–185, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709961.

Anonymous:1994:BRb

[Ano94b] Anonymous. Books received. Journal of the History of Ideas, 55(2): 349–358, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709905.

Anonymous:1994:BRc

[Ano94c] Anonymous. Books received. Journal of the History of Ideas, 55(3): 525–532, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709854.

Anonymous:1994:BRd

[Ano94d] Anonymous. Books received. Journal of the History of Ideas, 55(4):697–703, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709930.

Anonymous:1994:FMa

[Ano94e] Anonymous. Front matter. Journal of the History of Ideas, 55(1):16– 178, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709949.

Anonymous:1994:FMb

[Ano94f] Anonymous. Front matter. Journal of the History of Ideas, 55(2): ??, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709895.

Anonymous:1994:FMc

[Ano94g] Anonymous. Front matter. Journal of the History of Ideas, 55(3):384– 482, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709844.

Anonymous:1994:FMd

[Ano94h] Anonymous. Front matter. Journal of the History of Ideas, 55(4):552– 704, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709920.

Anonymous:1994:Na

[Ano94i] Anonymous. Notices. Journal of the History of Ideas, 55(1):175–177, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709960.

Anonymous:1994:Nb

[Ano94j] Anonymous. Notices. Journal of the History of Ideas, 55(2):347, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709904.

Anonymous:1994:Nc

[Ano94k] Anonymous. Notices. Journal of the History of Ideas, 55(3):523–524, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709853.

Anonymous:1994:Nd

[Ano94l] Anonymous. Notices. Journal of the History of Ideas, 55(4):695, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709929.

Anonymous:1994:VI

[Ano94m] Anonymous. Volume information. Journal of the History of Ideas, 55(4):705–708, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709919.

Anonymous:1995:BRa

[Ano95a] Anonymous. Books received. Journal of the History of Ideas, 56(1):165–175, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710015.

Anonymous:1995:BRb

[Ano95b] Anonymous. Books received. Journal of the History of Ideas, 56(2): 341–351, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709842.

Anonymous:1995:BRc

[Ano95c] Anonymous. Books received. Journal of the History of Ideas, 56(3): 523–529, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710039.

Anonymous:1995:BRd

[Ano95d] Anonymous. Books received. Journal of the History of Ideas, 56(4):685–693, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710001.

Anonymous:1995:FMa

[Ano95e] Anonymous. Front matter. Journal of the History of Ideas, 56(1):40– 164, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710003.

Anonymous:1995:FMb

[Ano95f] Anonymous. Front matter. Journal of the History of Ideas, 56(2):194– 308, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709833.

Anonymous:1995:FMc

[Ano95g] Anonymous. Front matter. Journal of the History of Ideas, 56(3):398– 500, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2710030.

Anonymous:1995:FMd

[Ano95h] Anonymous. Front matter. Journal of the History of Ideas, 56(4):554– 694, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709990.

Anonymous:1995:Na

[Ano95i] Anonymous. Notices. Journal of the History of Ideas, 56(1):163, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2710014.

Anonymous:1995:Nb

[Ano95j] Anonymous. Notices. Journal of the History of Ideas, 56(2):339–340, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709841.

Anonymous:1995:Nc

[Ano95k] Anonymous. Notices. Journal of the History of Ideas, 56(4):681–683, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2710000.

Anonymous:1995:VI

[Ano95l] Anonymous. Volume information. Journal of the History of Ideas, 56(4):695–698, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709989.

Anonymous:1996:BRa

[Ano96a] Anonymous. Books received. Journal of the History of Ideas, 57 (1):187–193, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14888.

Anonymous:1996:BRb

[Ano96b] Anonymous. Books received. Journal of the History of Ideas, 57(2): 375–382, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14901.

Anonymous:1996:BRc

[Ano96c] Anonymous. Books received. Journal of the History of Ideas, 57(3): 553–561, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14910.

Anonymous:1996:BRd

[Ano96d] Anonymous. Books received. Journal of the History of Ideas, 57 (4):745–752, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14922.

Anonymous:1997:BRa (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14929.

Anonymous:1997:BRb

[Ano97b] Anonymous. Books received. Journal of the History of Ideas, 58(2): 355–363, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14938.

Anonymous:1997:BRc

[Ano97c] Anonymous. Books received. Journal of the History of Ideas, 58(3): ??, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14949.

Anonymous:1997:BRd

[Ano97d] Anonymous. Books received. Journal of the History of Ideas, 58 (4):735–742, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14960.

Anonymous:1998:BRa

[Ano98a] Anonymous. Books received. Journal of the History of Ideas, 59 (1):185–194, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14978.

Anonymous:1998:BRb

[Ano98b] Anonymous. Books received. Journal of the History of Ideas, 59(2): 379–388, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14980.

Anonymous:1998:BRc

[Ano98c] Anonymous. Books received. Journal of the History of Ideas, 59(3): 551–555, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14990.

Anonymous:1998:BRd

[Ano98d] Anonymous. Books received. Journal of the History of Ideas, 59 (4):741–748, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15000.

Anonymous:1998:EBH

[Ano98e] Anonymous. Eklektik. Eine Begriffsgeschichte mit Hinweisen auf die Philosophie- und Wissenschaftsgeschichte (review). Journal of the History of Ideas, 59(1):173–182, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/207184.

Anonymous:1998:Na

[Ano98f] Anonymous. Notices. Journal of the History of Ideas, 59(2):377, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14984.

Anonymous:1998:Nb

[Ano98g] Anonymous. Notices. Journal of the History of Ideas, 59(3):549, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14994.

Anonymous:1998:Nc

[Ano98h] Anonymous. Notices. Journal of the History of Ideas, 59(4):739, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15008.

Anonymous:1999:BRa

[Ano99a] Anonymous. Books received. Journal of the History of Ideas, 60 (1):189–194, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15014.

Anonymous:1999:BRb

[Ano99b] Anonymous. Books received. Journal of the History of Ideas, 60(2): 375–384, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15023.

Anonymous:1999:BRc

[Ano99c] Anonymous. Books received. Journal of the History of Ideas, 60(3): 571–578, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15034.

Anonymous:1999:BRd

[Ano99d] Anonymous. Books received. Journal of the History of Ideas, 60 (4):761–769, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15045.

Anonymous:2000:BRa

[Ano00a] Anonymous. Books received. Journal of the History of Ideas, 61 (1):159–165, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15056.

Anonymous:2000:BRb

[Ano00b] Anonymous. Books received. Journal of the History of Ideas, 61(2): 349–360, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15067.

Anonymous:2000:BRc

[Ano00c] Anonymous. Books received. Journal of the History of Ideas, 61(3): 533–538, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15078.

Anonymous:2000:BRd

[Ano00d] Anonymous. Books received. Journal of the History of Ideas, 61 (4):703–711, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15090.

Anonymous:2000:EPO

[Ano00e] Anonymous. Errata: Paul Oskar Kristeller 1905–1999. Journal of the History of Ideas, 61(1):158, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15057. See [Mah99].

Anonymous:2000:ERP

[Ano00f] Anonymous. Errata: The rise of postmodernisms and the “End of Science”. Journal of the History of Ideas, 61(4):699, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15092. See [Hol00].

Anonymous:2001:BRa

[Ano01a] Anonymous. Books received. Journal of the History of Ideas, 62 (1):185–192, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15107.

Anonymous:2001:BRb

[Ano01b] Anonymous. Books received. Journal of the History of Ideas, 62(2): 367–379, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15109.

Anonymous:2001:BRc

[Ano01c] Anonymous. Books received. Journal of the History of Ideas, 62(3): 565–576, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15120.

Anonymous:2001:BRd

[Ano01d] Anonymous. Books received. Journal of the History of Ideas, 62 (4):747–752, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15129.

Anonymous:2002:BRa

[Ano02a] Anonymous. Books received. Journal of the History of Ideas, 63 (1):175–184, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15140.

Anonymous:2002:BRb

[Ano02b] Anonymous. Books received. Journal of the History of Ideas, 63(2): 365–370, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15149.

Anonymous:2002:BRc

[Ano02c] Anonymous. Books received. Journal of the History of Ideas, 63(3): 549–553, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15160.

Anonymous:2002:BRd

[Ano02d] Anonymous. Books received. Journal of the History of Ideas, 63 (4):735–744, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/39288.

Anonymous:2002:IVa

[Ano02e] Anonymous. Index to volume 63. Journal of the History of Ideas, 63(4):745–746, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/39291.

Anonymous:2002:IVb

[Ano02f] Anonymous. Index to volume 63. Journal of the History of Ideas, 63(4):747–748, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/39294.

Anonymous:2003:BRa

[Ano03a] Anonymous. Books received. Journal of the History of Ideas, 64 (1):137–147, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/42174.

Anonymous:2003:BRb

[Ano03b] Anonymous. Books received. Journal of the History of Ideas, 64(2): 337–341, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
43507.

Anonymous:2003:BRc

[Ano03c] Anonymous. Books received. Journal of the History of Ideas, 64(3): 513–519, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
47655.

Anonymous:2003:BRd

[Ano03d] Anonymous. Books received. Journal of the History of Ideas, 64 (4):661–668, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/52445.

Anonymous:2003:CV

[Ano03e] Anonymous. Contents of volume 64. Journal of the History of Ideas, 64(4):669–670, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/52446.

Anonymous:2003:DHI

[Ano03f] Anonymous. Dictionary of the history of ideas: Studies of selected pivotal ideas. Journal of the History of Ideas, 64(3):511, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/47658.

Anonymous:2003:IV

[Ano03g] Anonymous. Index to volume 64. Journal of the History of Ideas, 64(4):671–672, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/52449.

Anonymous:2003:MDF

[Ano03h] Anonymous. The Morris D. Forkosch Prize, the Selma V. Forkosch Prize. Journal of the History of Ideas, 64(3):iv, July 2003. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/47663.

Anonymous:2003:N

[Ano03i] Anonymous. Notices. Journal of the History of Ideas, 64(4):659, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/52453.

Anonymous:2004:BRa

[Ano04a] Anonymous. Books received. Journal of the History of Ideas, 65 (1):161–169, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/168330.

Anonymous:2004:BRb

[Ano04b] Anonymous. Books received. Journal of the History of Ideas, 65(2): 345–352, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
173844.

Anonymous:2004:BRc

[Ano04c] Anonymous. Books received. Journal of the History of Ideas, 65(3): 515–523, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
177662.

Anonymous:2004:BRd

[Ano04d] Anonymous. Books received. Journal of the History of Ideas, 65 (4):679–687, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/184064.

Anonymous:2004:CV

[Ano04e] Anonymous. Contents of volume 65. Journal of the History of Ideas, 65(4):689–690, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/184065.

Anonymous:2004:IV

[Ano04f] Anonymous. Index to volume 65. Journal of the History of Ideas, 65(4):691–692, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/184067.

Anonymous:2004:Na

[Ano04g] Anonymous. Notices. Journal of the History of Ideas, 65(1):159–160, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/168338.

Anonymous:2004:Nb

[Ano04h] Anonymous. Notices. Journal of the History of Ideas, 65(4):677–678, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/184072.

Anonymous:2005:BRa

[Ano05a] Anonymous. Books received. Journal of the History of Ideas, 66 (1):133–141, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/185255.

Anonymous:2005:BRb

[Ano05b] Anonymous. Books received. Journal of the History of Ideas, 66(2): 303–306, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
187184.

Anonymous:2005:CV

[Ano05c] Anonymous. Contents of volume 66. Journal of the History of Ideas, 66(4):651–652, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/191470.

Anonymous:2005:IV

[Ano05d] Anonymous. Index to volume 66. Journal of the History of Ideas, 66(4):653–654, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/191473.

Anonymous:2006:BRa

[Ano06a] Anonymous. Books received. Journal of the History of Ideas, 67 (1):209–218, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/194640.

Anonymous:2006:BRb

[Ano06b] Anonymous. Books received. Journal of the History of Ideas, 67(2): 409–415, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
198074.

Anonymous:2006:BRc

[Ano06c] Anonymous. Books received. Journal of the History of Ideas, 67(3): 547–553, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
202872.

Anonymous:2006:BRd

[Ano06d] Anonymous. Books received. Journal of the History of Ideas, 67 (4):735–742, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/204772.

Anonymous:2006:CV

[Ano06e] Anonymous. Contents of volume 67. Journal of the History of Ideas, 67(4):745–746, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/204773.

Anonymous:2006:IV

[Ano06f] Anonymous. Index to volume 67. Journal of the History of Ideas, 67(4):747–748, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/204774.

Anonymous:2006:Na

[Ano06g] Anonymous. Notice. Journal of the History of Ideas, 67(2):417, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/198079.

Anonymous:2006:Nb

[Ano06h] Anonymous. Notices. Journal of the History of Ideas, 67(3):555– 557, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
202878.

Anonymous:2006:Nc

[Ano06i] Anonymous. Notices. Journal of the History of Ideas, 67(4):743– 744, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
204779.

Anonymous:2007:BRa

[Ano07a] Anonymous. Books received. Journal of the History of Ideas, 68 (1):181–184, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/210685.

Anonymous:2007:BRb

[Ano07b] Anonymous. Books received. Journal of the History of Ideas, 68(2): 355–362, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
215052.

Anonymous:2007:BRc

[Ano07c] Anonymous. Books received. Journal of the History of Ideas, 68(3): 529–532, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
219815.

Anonymous:2007:BRd

[Ano07d] Anonymous. Books received. Journal of the History of Ideas, 68 (4):717–720, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/223129.

Anonymous:2007:CV

[Ano07e] Anonymous. Contents of volume 68. Journal of the History of Ideas, 68(4):727–728, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/223138.

Anonymous:2007:IV

[Ano07f] Anonymous. Index to volume 68. Journal of the History of Ideas, 68(4):712–714, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/223137.

Anonymous:2007:Na

[Ano07g] Anonymous. Notices. Journal of the History of Ideas, 68(1):185– 186, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
210690.

Anonymous:2007:Nb

[Ano07h] Anonymous. Notices. Journal of the History of Ideas, 68(3):533– 535, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
219820.

Anonymous:2007:Nc

[Ano07i] Anonymous. Notices. Journal of the History of Ideas, 68(4):721– 723, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
223126.

Anonymous:2008:BRa

[Ano08a] Anonymous. Books received. Journal of the History of Ideas, 69 (1):163–166, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/231006.

Anonymous:2008:BRb

[Ano08b] Anonymous. Books received. Journal of the History of Ideas, 69(2): 349–350, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
235866.

Anonymous:2008:BRc

[Ano08c] Anonymous. Books received. Journal of the History of Ideas, 69(3): 501–511, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
242284.

Anonymous:2008:BRd

[Ano08d] Anonymous. Books received. Journal of the History of Ideas, 69 (4):675–684, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/251820.

Anonymous:2008:CV

[Ano08e] Anonymous. Contents of volume 69. Journal of the History of Ideas, 69(4):687–688, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/251822/pdf.

Anonymous:2008:IV

[Ano08f] Anonymous. Index to volume 69. Journal of the History of Ideas, 69(4):689–690, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/251819/pdf.

Anonymous:2008:Na

[Ano08g] Anonymous. Notice. Journal of the History of Ideas, 69(1):167, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
231012.

Anonymous:2008:Nc

[Ano08h] Anonymous. Notice. Journal of the History of Ideas, 69(4):685, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/251821.

Anonymous:2008:Nb

[Ano08i] Anonymous. Notices. Journal of the History of Ideas, 69(3):513– 515, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
242285.

Anonymous:2009:BRa

[Ano09a] Anonymous. Books received. Journal of the History of Ideas, 70 (1):181–188, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/258200.

Anonymous:2009:BRb

[Ano09b] Anonymous. Books received. Journal of the History of Ideas, 70(2): 341–346, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
261982.

Anonymous:2009:BRc

[Ano09c] Anonymous. Books received. Journal of the History of Ideas, 70(3): 511–517, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
270109.

Anonymous:2009:BRd

[Ano09d] Anonymous. Books received. Journal of the History of Ideas, 70 (4):681–685, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/361793.

Anonymous:2009:CV

[Ano09e] Anonymous. Contents of volume 70. Journal of the History of Ideas, 70(4):691–692, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/364085/pdf.

Anonymous:2009:IV

[Ano09f] Anonymous. Index to volume 70. Journal of the History of Ideas, 70(4):693–694, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/361789/pdf.

Anonymous:2009:MDF

[Ano09g] Anonymous. Morris D. Forkosch Prize. Journal of the History of Ideas, 70(3):519, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/270110.

Anonymous:2009:Na

[Ano09h] Anonymous. Notice. Journal of the History of Ideas, 70(1):189, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
258201.

Anonymous:2009:Nb

[Ano09i] Anonymous. Notice. Journal of the History of Ideas, 70(4):687, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/364084.

Anonymous:2010:BRa

[Ano10a] Anonymous. Books received. Journal of the History of Ideas, 71(1): 163, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
370099.

Anonymous:2010:BRb

[Ano10b] Anonymous. Books received. Journal of the History of Ideas, 71(3): 505–515, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
386155.

Anonymous:2010:CV

[Ano10c] Anonymous. Contents of volume 71. Journal of the History of Ideas, 71(4):689–690, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/398304/pdf.

Anonymous:2010:IV

[Ano10d] Anonymous. Index to volume 71. Journal of the History of Ideas, 71(4):691–692, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/398305/pdf.

Anonymous:2010:Na

[Ano10e] Anonymous. Notice. Journal of the History of Ideas, 71(1):165, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
370096.

Anonymous:2010:Nb

[Ano10f] Anonymous. Notices. Journal of the History of Ideas, 71(2):335– 336, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
379776.

Anonymous:2010:Nc

[Ano10g] Anonymous. Notices. Journal of the History of Ideas, 71(3):517– 519, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
386156.

Anonymous:2010:Nd

[Ano10h] Anonymous. Notices. Journal of the History of Ideas, 71(4):685, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/398303.

Anonymous:2011:BR

[Ano11a] Anonymous. Books received. Journal of the History of Ideas, 72(3): 469–475, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
445509.

Anonymous:2011:CV

[Ano11b] Anonymous. Contents of volume 72. Journal of the History of Ideas, 72(4):641–642, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/453553.

Anonymous:2011:IV

[Ano11c] Anonymous. Index to volume 72. Journal of the History of Ideas, 72(4):643–644, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/453554.

Anonymous:2011:Na

[Ano11d] Anonymous. Notices. Journal of the History of Ideas, 72(2):329, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/429561.

Anonymous:2011:Nb

[Ano11e] Anonymous. Notices. Journal of the History of Ideas, 72(3):477– 478, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
445507.

Anonymous:2011:Nc

[Ano11f] Anonymous. Notices. Journal of the History of Ideas, 72(4):637, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/453552.

Anonymous:2012:BRa

[Ano12a] Anonymous. Books received. Journal of the History of Ideas, 73 (1):167–173, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/463145.

Anonymous:2012:BRb

[Ano12b] Anonymous. Books received. Journal of the History of Ideas, 73(2): 343–346, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
475334.

Anonymous:2012:BRc

[Ano12c] Anonymous. Books received. Journal of the History of Ideas, 73(3): 507–510, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
481422.

Anonymous:2012:BRd

[Ano12d] Anonymous. Books received. Journal of the History of Ideas, 73 (4):667–670, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488355.

Anonymous:2012:CV

[Ano12e] Anonymous. Contents of volume 73. Journal of the History of Ideas, 73(4):675–676, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488357.

Anonymous:2012:IV

[Ano12f] Anonymous. Index to volume 73. Journal of the History of Ideas, 73(4):677–678, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488358.

Anonymous:2012:Na

[Ano12g] Anonymous. Notices. Journal of the History of Ideas, 73(2):347, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/475338.

Anonymous:2012:Nb

[Ano12h] Anonymous. Notices. Journal of the History of Ideas, 73(3):511– 512, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
481423.

Anonymous:2012:Nc

[Ano12i] Anonymous. Notices. Journal of the History of Ideas, 73(4):671, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/488356.

Anonymous:2013:BRa

[Ano13a] Anonymous. Books received. Journal of the History of Ideas, 74(2): 331–334, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
504791.

Anonymous:2013:BRb

[Ano13b] Anonymous. Books received. Journal of the History of Ideas, 74(3): 497–501, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
511979.

Anonymous:2013:BRc

[Ano13c] Anonymous. Books received. Journal of the History of Ideas, 74 (4):677–681, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/523435.

Anonymous:2013:CV

[Ano13d] Anonymous. Contents of volume 74. Journal of the History of Ideas, 74(4):687–688, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/523437.

Anonymous:2013:IV

[Ano13e] Anonymous. Index to volume 74. Journal of the History of Ideas, 74(4):689–690, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/523438.

Anonymous:2013:Na

[Ano13f] Anonymous. Notices. Journal of the History of Ideas, 74(3):503, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/511978.

Anonymous:2013:Nb

[Ano13g] Anonymous. Notices. Journal of the History of Ideas, 74(4):683, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/523436.

Anonymous:2014:BRa

[Ano14a] Anonymous. Books received. Journal of the History of Ideas, 75 (1):161–162, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/535158.

Anonymous:2014:BRb

[Ano14b] Anonymous. Books received. Journal of the History of Ideas, 75(2): 339–346, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
544564.

Anonymous:2014:BRc

[Ano14c] Anonymous. Books received. Journal of the History of Ideas, 75(3): 513–516, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
549333.

Anonymous:2014:CV

[Ano14d] Anonymous. Contents of volume 75. Journal of the History of Ideas, 75(4):701–702, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/557243.

Anonymous:2014:ICI

[Ano14e] Anonymous. Ideas in context at 100: Introduction. Journal of the History of Ideas, 75(4):651–652, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/557237.

Anonymous:2014:IV

[Ano14f] Anonymous. Index to volume 75. Journal of the History of Ideas, 75(4):703–704, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/557244.

Anonymous:2014:Na

[Ano14g] Anonymous. Notices. Journal of the History of Ideas, 75(1):163– 164, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
535159.

Anonymous:2014:Nb

[Ano14h] Anonymous. Notices. Journal of the History of Ideas, 75(3):517– 518, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
549334.

Anonymous:2014:Nc

[Ano14i] Anonymous. Notices. Journal of the History of Ideas, 75(4):697, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/557242.

Anonymous:2015:BRa

[Ano15a] Anonymous. Books received. Journal of the History of Ideas, 76(2): 315–319, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
580302.

Anonymous:2015:BRb

[Ano15b] Anonymous. Books received. Journal of the History of Ideas, 76(3): 491–494, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
587275.

Anonymous:2015:CV

[Ano15c] Anonymous. Contents of volume 76. Journal of the History of Ideas, 76(4):661–662, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/597352.

Anonymous:2015:IV

[Ano15d] Anonymous. Index to volume 76. Journal of the History of Ideas, 76(4):663–664, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/597353.

Anonymous:2015:MDF

[Ano15e] Anonymous. Morris D. Forkosch Prize. Journal of the History of Ideas, 76(2):321, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/580303.

Anonymous:2015:Na

[Ano15f] Anonymous. Notices. Journal of the History of Ideas, 76(1):163– 164, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
566381.

Anonymous:2015:Nb

[Ano15g] Anonymous. Notices. Journal of the History of Ideas, 76(4):657– 658, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
597351.

Anonymous:2016:BRa

[Ano16a] Anonymous. Books received. Journal of the History of Ideas, 77 (1):173–177, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/610567.

Anonymous:2016:BRb

[Ano16b] Anonymous. Books received. Journal of the History of Ideas, 77(2): 347–353, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
620985.

Anonymous:2016:BRc

[Ano16c] Anonymous. Books received. Journal of the History of Ideas, 77(3): 509–515, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
626913.

Anonymous:2016:BRd

[Ano16d] Anonymous. Books received. Journal of the History of Ideas, 77 (4):679–682, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/635634.

Anonymous:2016:CV

[Ano16e] Anonymous. Contents of volume 77. Journal of the History of Ideas, 77(4):687–688, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/635636.

Anonymous:2016:Na

[Ano16f] Anonymous. Notices. Journal of the History of Ideas, 77(2):355– 356, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
620986.

Anonymous:2016:Nb

[Ano16g] Anonymous. Notices. Journal of the History of Ideas, 77(3):517– 518, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
626914.

Anonymous:2017:BRa

[Ano17a] Anonymous. Books received. Journal of the History of Ideas, 78 (1):157–161, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/648127.

Anonymous:2017:BRb

[Ano17b] Anonymous. Books received. Journal of the History of Ideas, 78(2): 309–312, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
652830.

Anonymous:2017:BRc

[Ano17c] Anonymous. Books received. Journal of the History of Ideas, 78(3): 497–501, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665525.

Anonymous:2017:BRd

[Ano17d] Anonymous. Books received. Journal of the History of Ideas, 78 (4):665–668, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/676151.

Anonymous:2017:CV

[Ano17e] Anonymous. Contents of volume 78. Journal of the History of Ideas, 78(4):675–676, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/676153.

Anonymous:2017:Na

[Ano17f] Anonymous. Notices. Journal of the History of Ideas, 78(1):163– 165, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
648128.

Anonymous:2017:Nb

[Ano17g] Anonymous. Notices. Journal of the History of Ideas, 78(2):313– 315, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
652831.

Anonymous:2017:Nc

[Ano17h] Anonymous. Notices. Journal of the History of Ideas, 78(3):503– 505, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665526.

Anonymous:2017:Nd

[Ano17i] Anonymous. Notices. Journal of the History of Ideas, 78(4):669– 671, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
676152.

Anonymous:2018:BRa

[Ano18a] Anonymous. Books received. Journal of the History of Ideas, 79 (1):169–171, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/684839.

Anonymous:2018:BRb

[Ano18b] Anonymous. Books received. Journal of the History of Ideas, 79(2): 331–336, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
693307.

Anonymous:2018:BRc

[Ano18c] Anonymous. Books received. Journal of the History of Ideas, 79(3): 497–499, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
703671.

Anonymous:2018:CV

[Ano18d] Anonymous. Contents of volume 79. Journal of the History of Ideas, 79(4):665–666, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/707778.

Anonymous:2018:Na

[Ano18e] Anonymous. Notices. Journal of the History of Ideas, 79(1):173– 175, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
684840.

Anonymous:2018:Nb

[Ano18f] Anonymous. Notices. Journal of the History of Ideas, 79(2):337– 339, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
693308.

Anonymous:2018:Nc

[Ano18g] Anonymous. Notices. Journal of the History of Ideas, 79(3):501– 503, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
703672.

Anonymous:2018:Nd

[Ano18h] Anonymous. Notices. Journal of the History of Ideas, 79(4):659– 661, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
707777.

Anonymous:2019:BRa

[Ano19a] Anonymous. Books received. Journal of the History of Ideas, 80 (1):171–175, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/715742.

Anonymous:2019:BRb

[Ano19b] Anonymous. Books received. Journal of the History of Ideas, 80(2): 333–336, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
722384.

Anonymous:2019:BRc

[Ano19c] Anonymous. Books received. Journal of the History of Ideas, 80(3): 501–503, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
729717.

Anonymous:2019:CV

[Ano19d] Anonymous. Contents of volume 80. Journal of the History of Ideas, 80(4):671–672, October 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/735387.

Anonymous:2019:Na

[Ano19e] Anonymous. Notices. Journal of the History of Ideas, 80(1):177– 179, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
715743.

Anonymous:2019:Nb

[Ano19f] Anonymous. Notices. Journal of the History of Ideas, 80(2):337– 339, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
722385.

Anonymous:2019:Nc

[Ano19g] Anonymous. Notices. Journal of the History of Ideas, 80(3):505– 507, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
729721.

Anonymous:2019:Nd

[Ano19h] Anonymous. Notices. Journal of the History of Ideas, 80(4):665– 667, October 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
735386.

Anonymous:2020:BRa

[Ano20a] Anonymous. Books received. Journal of the History of Ideas, 81 (1):157–162, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/747167.

Anonymous:2020:BRb

[Ano20b] Anonymous. Books received. Journal of the History of Ideas, 81(2): 327–330, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
753531.

Anonymous:2020:BRc

[Ano20c] Anonymous. Books received. Journal of the History of Ideas, 81(3): 499–501, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
761034.

Anonymous:2020:BRd

[Ano20d] Anonymous. Books received. Journal of the History of Ideas, 81 (4):663–665, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/772510.

Anonymous:2020:CV

[Ano20e] Anonymous. Contents of volume 81. Journal of the History of Ideas, 81(4):673–674, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/772512.

Anonymous:2020:Na

[Ano20f] Anonymous. Notices. Journal of the History of Ideas, 81(1):163– 165, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
747170.

Anonymous:2020:Nb

[Ano20g] Anonymous. Notices. Journal of the History of Ideas, 81(3):503– 505, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
761031.

Anonymous:2021:BRa

[Ano21a] Anonymous. Books received. Journal of the History of Ideas, 82 (1):175–178, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/781618.

Anonymous:2021:BRb

[Ano21b] Anonymous. Books received. Journal of the History of Ideas, 82(3): 539–540, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
798279.

Anonymous:2021:BRc

[Ano21c] Anonymous. Books received. Journal of the History of Ideas, 82 (4):703–704, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/837668.

Anonymous:2021:CV

[Ano21d] Anonymous. Contents of volume 82. Journal of the History of Ideas, 82(4):711–713, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/837670.

Anonymous:2021:Na

[Ano21e] Anonymous. Notices. Journal of the History of Ideas, 82(1):179– 181, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
781619.

Anonymous:2021:Nb

[Ano21f] Anonymous. Notices. Journal of the History of Ideas, 82(2):351– 353, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
790207.

Anonymous:2021:Nc

[Ano21g] Anonymous. Notices. Journal of the History of Ideas, 82(4):705– 707, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
837669.

Anonymous:2022:BRa

[Ano22a] Anonymous. Books received. Journal of the History of Ideas, 83 (1):171–173, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/847825.

Anonymous:2022:BRb

[Ano22b] Anonymous. Books received. Journal of the History of Ideas, 83(3): 523–525, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
858861.

Anonymous:2022:BRc

[Ano22c] Anonymous. Books received. Journal of the History of Ideas, 83 (4):685–688, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865888.

Anonymous:2022:CV

[Ano22d] Anonymous. Contents of volume 83. Journal of the History of Ideas, 83(4):695–697, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865890.

Anonymous:2022:Na

[Ano22e] Anonymous. Notices. Journal of the History of Ideas, 83(2):357– 359, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
855171.

Anonymous:2022:Nb

[Ano22f] Anonymous. Notices. Journal of the History of Ideas, 83(3):527– 529, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
858862.

Anonymous:2022:Nc

[Ano22g] Anonymous. Notices. Journal of the History of Ideas, 83(4):689– 691, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
865889.

Anonymous:2023:BRa

[Ano23a] Anonymous. Books received. Journal of the History of Ideas, 84 (1):197–199, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/879546.

Anonymous:2023:BRb

[Ano23b] Anonymous. Books received. Journal of the History of Ideas, 84(2): 409–410, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
886188.

Anonymous:2023:BRc

[Ano23c] Anonymous. Books received. Journal of the History of Ideas, 84(3): 585–587, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
901495.

Anonymous:2023:BRd

[Ano23d] Anonymous. Books received. Journal of the History of Ideas, 84 (4):803–805, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/909539.

Anonymous:2023:CV

[Ano23e] Anonymous. Contents of volume 83. Journal of the History of Ideas, 84(4):813–814, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/909541.

Anonymous:2023:Na

[Ano23f] Anonymous. Notices. Journal of the History of Ideas, 84(1):201– 203, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
879547.

Anonymous:2023:Nb

[Ano23g] Anonymous. Notices. Journal of the History of Ideas, 84(2):411– 413, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
886189.

Anonymous:2023:Nc

[Ano23h] Anonymous. Notices. Journal of the History of Ideas, 84(3):589– 591, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
901496.

Anonymous:2023:Nd

[Ano23i] Anonymous. Notices. Journal of the History of Ideas, 84(4):807– 809, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
909540.

Anonymous:2024:CV

[Ano24a] Anonymous. Contents of volume 85. Journal of the History of Ideas, 85(4):861–862, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/944589.

Anonymous:2024:Na

[Ano24b] Anonymous. Notices. Journal of the History of Ideas, 85(1):179– 181, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
917120.

Anonymous:2024:Nb

[Ano24c] Anonymous. Notices. Journal of the History of Ideas, 85(2):419– 421, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
926154.

Anonymous:2024:Nc

[Ano24d] Anonymous. Notices. Journal of the History of Ideas, 85(3):627– 629, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
933861.

Anonymous:2024:Nd

[Ano24e] Anonymous. Notices. Journal of the History of Ideas, 85(4):857, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/944588.

Anonymous:2025:N

[Ano25] Anonymous. Notices. Journal of the History of Ideas, 86(1):213, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
949933.

Anselment:1978:SCS

[Ans78] Raymond A. Anselment. Socrates and the clouds: Shaftesbury and a Socratic tradition. Journal of the History of Ideas, 39(2):171–182, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708773.

Anstey:1999:BOU

[Ans99] Peter Anstey. Boyle on occasionalism: an unexamined source. Journal of the History of Ideas, 60(1):57–81, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15013.

Ansart:2009:VMR

[Ans09] Guillaume Ansart. Variations on Montesquieu: Raynal and Diderot’s Histoire des deux Indes and the American Revolution. Journal of the History of Ideas, 70(3):399–420, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/270104.

Anthony:1966:MPU

[Ant66] H. Sylvia Anthony. Mercurius Politicus under Milton. Journal of the History of Ideas, 27(4):593–596 + 598–609, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708343.

Antin:1968:CGP

[Ant68] David Antin. Caxton’s the game and playe of the chesse. Journal of the History of Ideas, 29(2):269–278, April/June 1968. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708580.

Anthony:2021:MHP

[Ant21] Patrick Anthony. Making historicity: Paleontology and the proximity of the past in Germany, 1775–1825. Journal of the History of Ideas, 82(2):231–256, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/790202.

Antic:2024:PDH

[Ant24] Ana Antić. Psychiatry and decolonization: Histories of transcultural psychiatry in the Twentieth Century. Journal of the History of Ideas, 85(1):149–177, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/917119.

Ansell-Pearson:1992:WUT

[AP92] Keith Ansell-Pearson. Who is the Ubermensch? Time, truth, and woman in Nietzsche. Journal of the History of Ideas, 53(2):309–331, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709876.

Applebaum:1964:BHR

[App64] Wilbur Applebaum. Boyle and Hobbes: a reconsideration. Journal of the History of Ideas, 25(1):117–119, January/March 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708089.

Aquila:1988:CCP

[Aqu88] Richard E. Aquila. The Cartesian and a certain “Poetic” notion of consciousness. Journal of the History of Ideas, 49(4):543–562, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709673.

Archambault:1967:TFN

[Arc67] Paul J. Archambault. Thucydides in France: The notion of “Justice” in the memoires of Philippe De Commynes. Journal of the History of Ideas, 28(1):89–98, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708483.

Ardao:1963:ATP

[Ard63] Arturo Ardao. Assimilation and transformation of positivism in Latin America. Journal of the History of Ideas, 24(4):515–522, October/ December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707981.

Ariew:1992:TCP

[Ari92] Roger Ariew. Theory of comets at Paris during the Seventeenth Century. Journal of the History of Ideas, 53(3):355–372, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709882.

Armstrong:1956:DBN

[Arm56] David M. Armstrong. Discussion: Berkeley’s new theory of vision. Journal of the History of Ideas, 17(1):127–129, January 1956. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707690.

Armstrong:1965:JLD

[Arm65] Robert L. Armstrong. John Locke’s “Doctrine of Signs”: a new metaphysics. Journal of the History of Ideas, 26(3):369–382, July/ September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708387.

Armstrong:1969:CPL

[Arm69] Robert L. Armstrong. Cambridge platonists and Locke on innate ideas. Journal of the History of Ideas, 30(2):187–202, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708432.

Armitage:2000:EBR

[Arm00] David Armitage. Edmund Burke and reason of state. Journal of the History of Ideas, 61(4):617–634, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15087.

Armenteros:2007:HNN

[Arm07] Carolina Armenteros. From human nature to normal humanity: Joseph de Maistre, Rousseau, and the origins of moral statistics. Journal of the History of Ideas, 68(1):107–130, January 2007. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/210684.

Arnstein:1957:BCR

[Arn57] Walter L. Arnstein. The bradlaugh case: a reappraisal. Journal of the History of Ideas, 18(2):254–269, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707628.

Arneil:1994:TPP

[Arn94] Barbara Arneil. Trade, plantations, and property: John Locke and the economic defense of colonialism. Journal of the History of Ideas, 55(4):591–609, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709924.

Aronson:1987:SRO

[Aro87] Ronald Aronson. Sartre’s return to ontology: Critique, II, rethinks the basis of L’Être et le Néant. Journal of the History of Ideas, 48 (1):99–116, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709614.

Arriaza:1994:ANC

[Arr94] Armand Arriaza. Adam’s noble children: an Early Modern theorist’s concept of human nobility. Journal of the History of Ideas, 55(3):385– 404, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709846.

Aasbrink:2021:WRW

[Åsb21] Elisabeth Åsbrink. When race was removed from racism: Per Engdahl, the networks that saved Fascism and the making of the concept of ethnopluralism. Journal of the History of Ideas, 82(1):133–151, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
781616.

Asfour:1999:HPN

[Asf99] Amal Asfour. Hogarth’s post-Newtonian universe. Journal of the History of Ideas, 60(4):693–716, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15044.

Ashley:1978:DPC

[Ash78] Kathleen M. Ashley. Divine power in Chester cycle and late medieval thought. Journal of the History of Ideas, 39(3):387–404, July/ September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709384.

Ashley:1979:CCN

[Ash79] Kathleen Ashley. Chester Cycle and nominalist thought. Journal of the History of Ideas, 40(3):477, July/September 1979. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709251.

Atkinson:1974:NMF

[Atk74] James B. Atkinson. Naïveté and modernity: The French Renaissance battle for a literary vernacular. Journal of the History of Ideas, 35(2):179–196, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708757.

Atkins:2014:EOC

[Atk14] Jed W. Atkins. Euripides’s Orestes and the concept of conscience in Greek philosophy. Journal of the History of Ideas, 75(1):1–22, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/535151.

Atlas:1952:SMD

[Atl52] Samuel Atlas. Solomon Maimon’s doctrine of infinite reason and its historical relations. Journal of the History of Ideas, 13(2):168–187, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707609.

Attfield:1983:CAN

[Att83] Robin Attfield. Christian attitudes to nature. Journal of the History of Ideas, 44(3):369–386, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709172.

Atwell:1971:BC

[Atw71] John E. Atwell. A brief commentary. Journal of the History of Ideas, 32(3):433–436, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708357.

Abu-Uksa:2019:CCD

[AU19] Wael Abu-’Uksa. The construction of the concepts “Democracy” and “Republic” in Arabic in the Eastern and Southern Mediterranean, 1798–1878. Journal of the History of Ideas, 80(2):249–270, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/722380.

Augst:1965:DCM

[Aug65] Bertrand Augst. Descartes’s compendium on music. Journal of the History of Ideas, 26(1):119–132, January/March 1965. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708403.

Ausubel:1952:WCM

[Aus52] Herman Ausubel. William Cobbett and Malthusianism. Journal of the History of Ideas, 13(2):250–256, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707614.

Austin:1970:INS

[Aus70] William H. Austin. Isaac Newton on science and religion. Journal of the History of Ideas, 31(4):521–542, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708258.

Avineri:1961:PWH

[Avi61] Shlomo Avineri. The problem of war in Hegel’s thought. Journal of the History of Ideas, 22(4):463–474, October/December 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708025.

Avineri:1964:MJE

[Avi64] Shlomo Avineri. Marx and Jewish emancipation. Journal of the History of Ideas, 25(3):445–450, July/September 1964. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707911.

Avineri:1967:MI

[Avi67] Shlomo Avineri. Marx and the intellectuals. Journal of the History of Ideas, 28(2):269–278, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708422.

Avrich:1966:AAI

[Avr66] Paul Avrich. Anarchism and anti-intellectualism in Russia. Journal of the History of Ideas, 27(3):381–390, July/September 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708592.

Axinn:1971:KAF

[Axi71] Sidney Axinn. Kant, authority, and the French Revolution. Journal of the History of Ideas, 32(3):423–432, July/September 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708356. See comment [Dyk71].

Axinn:1993:OPP

[Axi93] Sidney Axinn. Obituary: Philip P. Wiener (1905–1992). Journal of the History of Ideas, 54(2):329, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709985.

Bachrach:2007:RHW

[Bac07] David Steward Bachrach. The rhetoric of historical writing: Documentary sources in histories of worms, c. 1300. Journal of the History of Ideas, 68(2):187–206, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/215051.

Bagley:1992:PE

[Bag92] Paul J. Bagley. On the practice of esotericism. Journal of the History of Ideas, 53(2):231–247, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709872.

Baker:1970:WWI

[Bak70] William Baker. William Wilberforce on the idea of Negro inferiority. Journal of the History of Ideas, 31(3):433–440, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708515.

Bakunin:1976:PLD

[Bak76] Jack S. Bakunin. Pierre Leroux on democracy, socialism, and the Enlightenment. Journal of the History of Ideas, 37(3):455–474, July/ September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708809.

Baker:1977:HSE

[Bak77] William Baker. Herbert Spencer and “Evolution” — a further note. Journal of the History of Ideas, 38(3):476, July/September 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708676.

Baker:1983:WRR

[Bak83] Susan Read Baker. The works of La Rochefoucauld in relation to Machiavellian ideas of morals and politics. Journal of the History of Ideas, 44(2):207–218, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2709136. See erratum [Ano83f].

Bakos:1991:QND

[Bak91] Adrianna E. Bakos. “Qui nescit dissimulare, nescit regnare”: Louis XI and raison d’ėtat during the reign of Louis XIII. Journal of the History of Ideas, 52(3):399–416, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2710044.

Ball:1942:LPN

[Bal42] M. Margaret Ball. The leadership principle in national socialism. Journal of the History of Ideas, 3(1):74–93, January 1942. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707462.

Baldensperger:1944:CTR

[Bal44] Fernand Baldensperger. 1793–1794: Climacteric times for “Romantic” tendencies in English ideology. Journal of the History of Ideas, 5(1):3–20, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707099.

Bald:1950:SWC

[Bal50] R. C. Bald. Sir William Chambers and the Chinese garden. Journal of the History of Ideas, 11(3):287–320, June 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707733.

Baldry:1956:HFA

[Bal56] H. C. Baldry. Hesiod’s five ages. Journal of the History of Ideas, 17(4):553–554, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707788.

Ballard:1960:LTS

[Bal60] Kaith Emerson Ballard. Leibniz’s theory of space and time. Journal of the History of Ideas, 21(1):49–65, January/March 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707998.

Bandel:1955:ECA

[Ban55] Betty Bandel. The English chroniclers’ attitude toward women. Journal of the History of Ideas, 16(1):113–118, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707530.

Bannister:1970:SFO

[Ban70] Robert C. Bannister. “The survival of the fittest is our doctrine”: History or histrionics? Journal of the History of Ideas, 31(3): 377–398, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708512.

Bantjes:2014:VPD

[Ban14] Rod Bantjes. “Vertical perspective does not exist”: The scandal of converging verticals and the final crisis of perspectiva artificialis. Journal of the History of Ideas, 75(2):307–338, April 2014. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/544563.

Bannon:2016:PES

[Ban16] Brad Bannon. President Edwards and the sage of highgate: Determinism, depravity, and the supernatural will. Journal of the History of Ideas, 77(1):27–47, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/610559.

Barzun:1941:RHP

[Bar41] Jacques Barzun. Romantic historiography as a political force in France. Journal of the History of Ideas, 2(3):318–329, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707134.

Barzun:1942:BRM

[Bar42] Jacques Barzun. Book review: Metapolitics. From the Romantics to Hitler. By Peter Viereck. New York: Alfred A. Knopf, 1941. 335 pp. $3.00. Journal of the History of Ideas, 3(1):107–110, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707464. See reply [Vie42].

Baron:1943:TMP

[Bar43] Hans Baron. Towards a more positive evaluation of the Fifteenth-Century Renaissance. Journal of the History of Ideas, 4(1):21–49, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707235.

Barzun:1945:HLA

[Bar45] Jacques Barzun. History as a liberal art. Journal of the History of Ideas, 6(1):81–88, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707057.

Baron:1947:BRN

[Bar47] Hans Baron. Book review: New Light on the Political Writers of the Florentine Renaissance. Journal of the History of Ideas, 8(2):241–248, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707160.

Baron:1950:BRF

[Bar50] Hans Baron. Book review: The First History of the Historical Concept of the Renaissance. Journal of the History of Ideas, 11(4):493–510, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707595.

Baron:1958:MPR

[Bar58a] Hans Baron. Moot problems of Renaissance interpretation: an answer to Wallace K. Ferguson. Journal of the History of Ideas, 19(1):26–34, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707950. See

[Fer58].

Baron:1958:PRI

[Bar58b] Samuel H. Baron. Plekhanov’s Russia: The impact of the West upon an “Oriental” society. Journal of the History of Ideas, 19(3):388–404, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708043.

Baron:1959:QAM

[Bar59a] Hans Baron. The querelle of the ancients and the moderns as a problem for Renaissance scholarship. Journal of the History of Ideas, 20(1): 3–22, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707964.

Barry:1959:HCP

[Bar59b] Robert M. Barry. History of Christian philosophy in the Middle Ages. Journal of the History of Ideas, 20(1):105–110, January 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707969.

Baron:1960:BRS

[Bar60a] Hans Baron. Book review: Secularization of Wisdom and Political Humanism in the Renaissance. Journal of the History of Ideas, 21 (1):131–150, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708005.

Baron:1960:MHO

[Bar60b] Hans Baron. Marvell’s “An Horatian Ode” and Machiavelli. Journal of the History of Ideas, 21(3):450–451, July/September 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708149.

Baron:1962:BRB

[Bar62] Hans Baron. Book review: A British Symposium on Italian Renaissance Civilization. Journal of the History of Ideas, 23(1): 143–146, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708064.

Barnard:1963:HTC

[Bar63] F. M. Barnard. Herder’s treatment of causation and continuity in history. Journal of the History of Ideas, 24(2):197–212, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707845.

Barnard:1971:PPC

[Bar71] F. M. Barnard. The “Practical Philosophy” of Christian Thomasius. Journal of the History of Ideas, 32(2):221–246, April/June 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708278.

Barlow:1973:IWR

[Bar73] Richard G. Barlow. Infinite worlds: Robert Burton’s cosmic voyage. Journal of the History of Ideas, 34(2):291–302, April/June 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708733.

Barasch:1975:CPB

[Bar75] Moshe Barasch. Character and physiognomy: Bocchi on Donatello’s St. George, a Renaissance text on expression in art. Journal of the History of Ideas, 36(3):413–430, July/September 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708654.

Barnouw:1981:SRF

[Bar81] Jeffrey Barnouw. The separation of reason and faith in Bacon and Hobbes, and Leibniz’s theodicy. Journal of the History of Ideas, 42 (4):607–628, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709121.

Barnard:1983:NCP

[Bar83] F. M. Barnard. National culture and political legitimacy: Herder and Rousseau. Journal of the History of Ideas, 44(2):231–253, April/ June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709138.

Barnouw:1988:ASP

[Bar88] Jeffrey Barnouw. “Aesthetic” for Schiller and Peirce: a neglected origin of pragmatism. Journal of the History of Ideas, 49(4):607– 632, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709676.

Barnouw:1992:PCP

[Bar92] Jeffrey Barnouw. Passion as “Confused” perception or thought in Descartes, Malebranche, and Hutcheson. Journal of the History of Ideas, 53(3):397–424, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709884.

Barash:1997:SM

[Bar97] Jeffrey Andrew Barash. The sources of memory. Journal of the History of Ideas, 58(4):707–717, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14959.

Baraz:1998:SEB

[Bar98] Daniel Baraz. Seneca, ethics, and the body: The treatment of cruelty in medieval thought. Journal of the History of Ideas, 59(2):195–215, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14979.

Baring:2016:IMC

[Bar16] Edward Baring. Ideas on the move: Context in transnational intellectual history. Journal of the History of Ideas, 77(4):567–587, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/635628.

Barr:2019:AOR

[Bar19] Timothy Barr. Without apparent occasion: Recent research on melancholy. Journal of the History of Ideas, 80(2):313–332, April 2019. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/722383.

Bassiri:2013:BUS

[Bas13] Nima Bassiri. The brain and the unconscious soul in Eighteenth-Century nervous physiology: Robert Whytt’s Sensorium Commune. Journal of the History of Ideas, 74(3):425–448, July 2013. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/511982.

Battenhouse:1948:DMC

[Bat48] Roy W. Battenhouse. The doctrine of man in Calvin and in Renaissance Platonism. Journal of the History of Ideas, 9(4):447–471, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707218.

Batz:1974:HAJ

[Bat74] William G. Batz. The historical anthropology of John Locke. Journal of the History of Ideas, 35(4):663–670, October/December 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709092.

Bates:1975:GTH

[Bat75] Thomas R. Bates. Gramsci and the theory of hegemony. Journal of the History of Ideas, 36(2):351–366, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708933.

Battisti:1977:DSU

[Bat77] Giuseppa S. Battisti. Democracy in Spinoza’s unfinished Tractatus Politicus. Journal of the History of Ideas, 38(4):623–634, October/ December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708691.

Bates:2000:MTL

[Bat00a] David Bates. The mystery of truth: Louis-Claude de Saint-Martin’s enlightened mysticism. Journal of the History of Ideas, 61(4):635–655, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15088.

Bates:2000:MED

[Bat00b] Donald G. Bates. Machina ex deo: William Harvey and the meaning of instrument. Journal of the History of Ideas, 61(4):577–593, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15089.

Bates:2019:IRE

[Bat19] Zach Bates. The idea of Royal Empire and the Imperial Crown of England, 1542–1698. Journal of the History of Ideas, 80(1):25–46, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/715734.

Batson:2024:SAP

[Bat24] Alexander D. Batson. Sinai and the Areopagus: Philip Melanchthon, natural law, and the beginnings of Athenian legal history in the shadow of the Schmalkaldic War. Journal of the History of Ideas, 85(4):713–748, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/944583.

Baumgardt:1945:BCM

[Bau45] David Baumgardt. Bentham’s “Censorial” method. Journal of the History of Ideas, 6(4):456–467, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707345.

Baumgardt:1946:LQL

[Bau46] David Baumgardt. Legendary quotations and lack of references. Journal of the History of Ideas, 7(1):99–102, January 1946. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707271. See note [Kah47].

Bauer:1970:PER

[Bau70] Robert J. Bauer. A phenomenon of epistemology in the Renaissance. Journal of the History of Ideas, 31(2):281–288, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708551.

Baum:2023:IDD

[Bau23] Jacob M. Baum. The idea of deafness as disability in Renaissance Germany. Journal of the History of Ideas, 84(4):621–652, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/909533.

Bayley:1949:PCP

[Bay49] Charles C. Bayley. Pivotal concepts in the political philosophy of William of Ockham. Journal of the History of Ideas, 10(2):199–218, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707415.

Baym:1965:TVS

[Bay65] Nina Baym. Thoreau’s view of science. Journal of the History of Ideas, 26(2):221–234, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708229.

Bloom:1951:JAE

[BB51] Edward A. Bloom and Lillian D. Bloom. Joseph Addison and Eighteenth-Century “Liberalism”. Journal of the History of Ideas, 12(4):560–583, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707486.

Bloom:1960:AMH

[BB60] Edward A. Bloom and Lillian D. Bloom. Addison on ‘moral habits of the mind’. Journal of the History of Ideas, 21(3):409–427, July/ September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708145.

Boas:1970:BRC

[BB70] George Boas and George Boas. Book review: Comment. Journal of the History of Ideas, 31(4):620–621, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708267.

Barker-Benfield:1989:MWE

[BB89] G. J. Barker-Benfield. Mary Wollstonecraft: Eighteenth-Century commonwealthwoman. Journal of the History of Ideas, 50(1):95– 115, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709788.

Breckman:2009:ERI

[BBGM09] Warren Breckman, Martin J. Burke, Anthony Grafton, and Ann E. Moyer. European reference index for the humanities. Journal of the History of Ideas, 70(2):349, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/261984.

Beach:1940:RNW

[Bea40] Joseph Warren Beach. Reason and nature in Wordsworth. Journal of the History of Ideas, 1(3):335–351, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707091.

Beardsley:1942:DMU

[Bea42] Monroe C. Beardsley. Dostoyevsky’s metaphor of the “Underground”. Journal of the History of Ideas, 3(3):265–290, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707305.

Beardsley:1971:BRM

[Bea71] Monroe C. Beardsley. Book review: Modes of Interpretation. Journal of the History of Ideas, 32(1):143–148, January/March 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708332.

Beardsley:1976:BRT

[Bea76] Monroe C. Beardsley. Book review: Tatarkiewicz’ History of Aesthetics. Journal of the History of Ideas, 37(3):549–555, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708816.

Beame:1982:UAM

[Bea82] Edmond M. Beame. The use and abuse of Machiavelli: The Sixteenth-Century French adaptation. Journal of the History of Ideas, 43 (1):33–54, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709159.

Beaujour:1983:ADE

[Bea83] Elizabeth Klosty Beaujour. Architectural discourse and Early Soviet literature. Journal of the History of Ideas, 44(3):477–495, July/ September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709177.

Beall:1991:HPD

[Bea91] E. F. Beall. Hesiod’s Prometheus and development in myth. Journal of the History of Ideas, 52(3):355–371, July/September 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710042.

Beame:1993:PH

[Bea93] Edmond M. Beame. The politiques and the historians. Journal of the History of Ideas, 54(3):355–379, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710018.

Beck:1947:PPP

[Bec47] Maximilian Beck. Plato’s problem in the Parmenides. Journal of the History of Ideas, 8(2):232–236, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707157.

Beck:1967:KS

[Bec67] Lewis White Beck. Kant’s strategy. Journal of the History of Ideas, 28(2):224–236, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708417.

Becker:1969:CSL

[Bec69] James F. Becker. The corporation spirit and its liberal analysis. Journal of the History of Ideas, 30(1):69–84, January/March 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708245.

Beck:1971:KRR

[Bec71] Lewis W. Beck. Kant and the right of revolution. Journal of the History of Ideas, 32(3):411–422, July/September 1971. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708355. See comment [Dyk71].

Beck:1972:LCK

[Bec72] Lewis White Beck. Lovejoy as a critic of Kant. Journal of the History of Ideas, 33(3):471–484, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709047.

Beck:1985:NBP

[Bec85] Hamilton Beck. The novel between 1740 and 1780: Parody and historiography. Journal of the History of Ideas, 46(3):405–416, July/ September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709475.

Beddeleem:2021:EBH

[Bed21] Martin Beddeleem. Epistemological battles on the home front: Early neoliberals at war against the social relations of science movement. Journal of the History of Ideas, 82(4):615–636, October 2021. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/837674.

Beenstock:2020:RUL

[Bee20] Zoe Beenstock. Reforming utilitarianism: Lyric poetry in J. S. Mill’s “Thoughts on Poetry and Its Varieties” and Autobiography. Journal of the History of Ideas, 81(4):599–620, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/772509.

Behrent:2008:MBS

[Beh08] Michael C. Behrent. The mystical body of society: Religion and association in Nineteenth-Century French political thought. Journal of the History of Ideas, 69(2):219–243, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/235865.

Bejczy:1997:TMC

[Bej97] Istvan Bejczy. Tolerantia: a medieval concept. Journal of the History of Ideas, 58(3):365–384, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/1/

article/14951.

Bellhouse:2008:BFM

[Bel08] David Bellhouse. Banishing fortuna: Montmort and De Moivre. Journal of the History of Ideas, 69(4):559–581, October 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/251815.

Bellucci:2015:LPA

[Bel15] Francesco Bellucci. Logic, psychology, and apperception: Charles S. Peirce and Johann F. Herbart. Journal of the History of Ideas, 76(1):69–91, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/566377.

Benjamin:1957:KBS

[Ben57] Edwin B. Benjamin. The King of Brobdingnag and secrets of state. Journal of the History of Ideas, 18(4):572–579, October 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707568.

Benson:1982:DHP

[Ben82a] Donald R. Benson. Dryden’s the hind and the panther: Transubstantiation and figurative language. Journal of the History of Ideas, 43(2):195–208, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709199.

Benton:1982:PEL

[Ben82b] Robert J. Benton. Political expediency and lying: Kant vs Benjamin constant. Journal of the History of Ideas, 43(1):135–144, January/ March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709165.

Benin:1984:CGD

[Ben84] Stephen D. Benin. The ’cunning of god’ and divine accommodation. Journal of the History of Ideas, 45(2):179–191, April/June 1984. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709286.

Bentley:2007:NN

[Ben07] Michael Bentley. News from nowhere. Journal of the History of Ideas, 68(3):375–379, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/219814.

Benite:2023:ANE

[Ben23] Zvi Ben-Dor Benite. The accountants of Nineveh: Exile Jews and capitalism in British Imperial thinking. Journal of the History of Ideas, 84(2):233–261, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/886181.

Beringause:1955:PPI

[Ber55] Arthur F. Beringause. The presentness of the past in Ireland. Journal of the History of Ideas, 16(2):240–246, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707665.

Beringause:1957:DMR

[Ber57] A. F. Beringause. The double martyrdom of Randolph Bourne. Journal of the History of Ideas, 18(4):594–603, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707571.

Berland:1962:HJA

[Ber62] Alwyn Berland. Henry James and the aesthetic tradition. Journal of the History of Ideas, 23(3):407–419, July/September 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708076.

Bercovitch:1967:PBF

[Ber67] Sacvan Bercovitch. The philosophical background to the fable of Emerson’s “American Scholar”. Journal of the History of Ideas, 28 (1):123–128, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708486.

Berlin:1969:FSO

[Ber69] Isaiah Berlin. [Feuerbach’s supposed objection to Hegel]: Reply to Orsini. Journal of the History of Ideas, 30(1):91–95, January/ March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708247. See

[Ors69].

Berman:1971:BCE

[Ber71] David Berman. Berkeley, Clayton, and an essay on spirit. Journal of the History of Ideas, 32(3):367–378, July/September 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708352.

Berry:1974:FEP

[Ber74a] Boyd M. Berry. The first English pediatricians and Tudor attitudes toward childhood. Journal of the History of Ideas, 35(4):561–577, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709086.

Berry:1974:ASC

[Ber74b] Christopher J. Berry. Adam Smith’s considerations on language. Journal of the History of Ideas, 35(1):130–138, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708748.

Berry:1977:HHC

[Ber77] Christopher J. Berry. From Hume to Hegel: The case of the Social Contract. Journal of the History of Ideas, 38(4):691–703, October/ December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708697.

Berlin:1978:MKD

[Ber78a] Isaiah Berlin. Marx’s kapital and Darwin. Journal of the History of Ideas, 39(3):519, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709395.

Berti:1978:AGD

[Ber78b] Enrico Berti. Ancient Greek dialectic as expression of freedom of thought and speech. Journal of the History of Ideas, 39(3): 347–370, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709382.

Berti:1980:RJS

[Ber80] Enrico Berti. Reply to James Seaton. Journal of the History of Ideas, 41(2):290–292, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709462. See [Sea80].

Berland:1986:BPH

[Ber86a] K. J. H. Berland. Bringing philosophy down from the heavens: Socrates and the new science. Journal of the History of Ideas, 47(2):299–308, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709816. See erratum [Ano86g].

Berman:1986:JBP

[Ber86b] David Berman. The jacobitism of Berkeley’s passive obedience. Journal of the History of Ideas, 47(2):309–319, April/June 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709817.

Bergman:1987:PHA

[Ber87] Jay Bergman. The perils of historical analogy: Leon trotsky on the French Revolution. Journal of the History of Ideas, 48(1):73–98, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709613.

Berti:1995:RU

[Ber95] Silvia Berti. At the roots of unbelief. Journal of the History of Ideas, 56(4):555–575, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709992.

Berry:2004:PRM

[Ber04] Jessica N. Berry. The Pyrrhonian revival in Montaigne and Nietzsche. Journal of the History of Ideas, 65(3):497–514, July 2004. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/177661.

Berns:2014:PPR

[Ber14] Andrew D. Berns. The place of paradise in Renaissance Jewish thought. Journal of the History of Ideas, 75(3):351–371, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/549326.

Berges:2018:FGP

[Ber18] Sandrine Bergès. Family, gender, and progress: Sophie de Grouchy and her exclusion in the publication of Condorcet’s sketch of human progress. Journal of the History of Ideas, 79(2):267–283, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/693304.

Bestor:1948:ESV

[Bes48] Arthur E. Bestor, Jr. The evolution of the socialist vocabulary. Journal of the History of Ideas, 9(3):259–302, June 1948. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707371.

Betts:1959:PBP

[Bet59] John Rickards Betts. P. T. Barnum and the popularization of natural history. Journal of the History of Ideas, 20(3):353–368, June/ September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708114.

Betz:2009:RSL

[Bet09] John R. Betz. Reading “Sibylline Leaves”: J. G. Hamann in the history of ideas. Journal of the History of Ideas, 70(1):93–118, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/258196.

Bevir:1993:EBB

[Bev93] Mark Bevir. Ernest Belfort Bax: Marxist, idealist, and positivist. Journal of the History of Ideas, 54(1):119–135, January 1993. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709863.

Bevir:2012:PAH

[Bev12] Mark Bevir. Post-analytic historicism. Journal of the History of Ideas, 73(4):657–665, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488354.

Bevir:2017:JRL

[Bev17] Mark Bevir. John Rawls in light of the archive: Introduction to the Symposium on the Rawls Papers. Journal of the History of Ideas, 78(2):255–263, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/652825.

Blair:1992:RHS

[BG92] Ann Blair and Anthony Grafton. Reassessing humanism and science. Journal of the History of Ideas, 53(4):535–540, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709934.

Boas:1959:NMP

[BH59] Marie Boas and Rupert Hall. Newton’s ‘mechanical principles’. Journal of the History of Ideas, 20(2):167–178, April 1959. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707817.

Ben-Israel:1960:WSH

[BI60] Hedva Ben-Israel. William Smyth, historian of the French Revolution. Journal of the History of Ideas, 21(4):571–585, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708105.

Bianco:2024:AYD

[Bia24] Giuseppe Bianco. The antihumanism of the young Deleuze: Sartre, Catholicism, and the perspective of the inhuman, 1945–48. Journal of the History of Ideas, 85(4):795–825, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/944586.

Bidney:1954:BRI

[Bid54] David Bidney. Book review: The Idea of the Savage in North American Ethnohistory. Journal of the History of Ideas, 15(2):322–327, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707777.

Biddle:1976:LCI

[Bid76] John C. Biddle. Locke’s critique of innate principles and Toland’s deism. Journal of the History of Ideas, 37(3):411–422, July/ September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708806.

Billings:2011:ETM

[Bil11] Joshua Billings. Epic and tragic music: The union of the arts in the Eighteenth Century. Journal of the History of Ideas, 72(1):99– 117, January 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
413476.

Bingham:1967:BRA

[Bin67] Alfred J. Bingham. Book review: Aspects of the Eighteenth Century by Earl R. Wasserman. Journal of the History of Ideas, 28 (1):129–132, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708487.

Bird:1962:TLT

[Bir62] Otto Bird. The tradition of the logical topics: Aristotle to Ockham. Journal of the History of Ideas, 23(3):307–323, July/September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708069.

Bird:1996:SCH

[Bir96] Alexander Bird. Squaring the circle: Hobbes on philosophy and geometry. Journal of the History of Ideas, 57(2):217–231, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14899.

Bishop:1952:KAI

[Bis52] Jonathan Bishop. Knowledge, action, and interpretation in Defoe’s novels. Journal of the History of Ideas, 13(1):3–16, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707723.

Bishop:1974:CJC

[Bis74] Donald H. Bishop. The carus-James controversy. Journal of the History of Ideas, 35(3):509–520, July/September 1974. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708798.

Bishop:1996:MMD

[Bis96] John D. Bishop. Moral motivation and the development of Francis Hutcheson’s philosophy. Journal of the History of Ideas, 57(2):277– 295, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14900.

Bisset:2015:JBT

[Bis15] Sophie Bisset. Jean Barbeyrac’s theory of permissive natural law and the foundation of property rights. Journal of the History of Ideas, 76(4):541–562, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/597346.

Bitzan:2017:LZM

[Bit17] Amos Bitzan. Leopold Zunz and the meanings of Wissenschaft. Journal of the History of Ideas, 78(2):233–254, April 2017. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/652824.

Bruno-Jofre:2019:LDN

[BJ19] Rosa Bruno-Jofré. Localizing Dewey’s notions of democracy and education: a journey across configurations in Latin America. Journal of the History of Ideas, 80(3):433–453, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/729718.

Blau:1949:JBE

[Bla49] Joseph L. Blau. Joel Barlow, enlightened religionist. Journal of the History of Ideas, 10(3):430–444, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707046.

Blau:1952:TLT

[Bla52] Joseph L. Blau. Tayler Lewis: True conservative. Journal of the History of Ideas, 13(2):218–233, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707612.

Blau:1957:RDT

[Bla57] Joseph L. Blau. Rosmini, Domodossola, and Thomas Davidson. Journal of the History of Ideas, 18(4):522–528, October 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707562.

Blackwell:1961:CWD

[Bla61] Richard J. Blackwell. Christian Wolff’s doctrine of the soul. Journal of the History of Ideas, 22(3):339–354, July/September 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708129.

Bland:1973:PL

[Bla73] D. E. Bland. Population and liberalism, 1770–1817. Journal of the History of Ideas, 34(1):113–122, January/March 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708947.

Black:1982:AMR

[Bla82] Robert Black. Ancients and moderns in the Renaissance: Rhetoric and history in Accolti’s dialogue on the preeminence of men of his own time. Journal of the History of Ideas, 43(1):3–32, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709158.

Blair:1990:TBC

[Bla90] Ann Blair. Tycho Brahe’s critique of Copernicus and the Copernican system. Journal of the History of Ideas, 51(3):355–377, July/ September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709620.

Black:1991:IRE

[Bla91a] Robert Black. Italian Renaissance education: Changing perspectives and continuing controversies. Journal of the History of Ideas, 52(2):315–334, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709531. See reply [Gre91c] and response [Bla91b].

Black:1991:RPG

[Bla91b] Robert Black. Reply to Paul Grendler. Journal of the History of Ideas, 52(3):519–520, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710050.

Blair:1992:HMN

[Bla92] Ann Blair. Humanist methods in natural philosophy: The commonplace book. Journal of the History of Ideas, 53(4):541–551, October/ December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709935.

Bland:1993:MJA

[Bla93] Kalman P. Bland. Medieval Jewish aesthetics: Maimonides, body, and scripture in Profiat Duran. Journal of the History of Ideas, 54(4):533– 559, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709823.

Blanchard:2001:PGA

[Bla01] W. Scott Blanchard. Petrarch and the genealogy of asceticism. Journal of the History of Ideas, 62(3):401–423, July 2001. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/15119.

Blair:2003:RSC

[Bla03] Ann Blair. Reading strategies for coping with information overload ca. 1550–1700. Journal of the History of Ideas, 64(1):11–28, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42175.

Blank:2013:JFU

[Bla13] Andreas Blank. Johannes von Felden on usucaption, justice, and the society of states. Journal of the History of Ideas, 74(3):403–423, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/511983.

Blank:2019:VJP

[Bla19] Andreas Blank. Value, justice, and presumption in the late scholastic controversy over price regulation. Journal of the History of Ideas, 80(2):183–202, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/722377.

Blank:2022:DDE

[Bla22] Daniel Blank. Debating drama in the Early Modern University: John Case, Aristotle’s Politics, and a previously unknown Oxford disputation. Journal of the History of Ideas, 83(3):387–406, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/858855.

Bleznick:1958:SRM

[Ble58] Donald W. Bleznick. Spanish reaction to Machiavelli in the Sixteenth and Seventeenth Centuries. Journal of the History of Ideas, 19(4):542– 550, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707922.

Blissett:1957:CS

[Bli57] William Blissett. Caesar and Satan. Journal of the History of Ideas, 18(2):221–232, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707625.

Blinderman:1982:HPP

[Bli82] Charles S. Blinderman. Huxley, Pater, and protoplasm. Journal of the History of Ideas, 43(3):477–486, July/September 1982. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709434.

Bloom:1946:WS

[Blo46] Solomon F. Bloom. The “Withering away” of the State. Journal of the History of Ideas, 7(1):113–121, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707273.

Bloch:1979:RHI

[Blo79] Jean H. Bloch. Rousseau and Helvétius on innate and acquired traits: The final stages of the Rousseau–Helvétius controversy. Journal of the History of Ideas, 40(1):21–41, January/March 1979. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709258.

Block:1984:EPA

[Blo84] Ed Block, Jr. Evolutionist psychology and aesthetics: The Cornhill Magazine, 1875–1880. Journal of the History of Ideas, 45(3): 465–475, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709236.

Blok:1996:PPB

[Blo96] Josine Blok. Proof and persuasion in Black Athena: The case of K. O. Muller. Journal of the History of Ideas, 57(4):705–724, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14921.

Bluestein:1963:ABH

[Blu63] Gene Bluestein. The advantages of barbarism: Herder and Whitman’s nationalism. Journal of the History of Ideas, 24(1):115–126, January/ March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707862.

Blum:2000:FPT

[Blu00] Paul Richard Blum. Francesco Patrizi in the “Time-Sack”: History and rhetorical philosophy. Journal of the History of Ideas, 61(1):59– 74, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15055.

Blythe:1986:MCD

[Bly86] James M. Blythe. The mixed constitution and the distinction between regal and political power in the work of Thomas Aquinas. Journal of the History of Ideas, 47(4):547–565, October/December 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709717.

Boas:1940:MLH

[Boa40] George Boas. The Mona Lisa in the history of taste. Journal of the History of Ideas, 1(2):207–224, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707333.

Boas:1947:BRH

[Boa47] George Boas. Book review: A History of Western Philosophy. by Bertrand Russell. Journal of the History of Ideas, 8(1):117–123, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707445.

Boas:1948:LHP

[Boa48a] George Boas. A. O. Lovejoy as historian of philosophy. Journal of the History of Ideas, 9(4):404–411, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707213.

Boas:1948:BRP

[Boa48b] George Boas. Book review: Professor Wolfson’s Philo. Journal of the History of Ideas, 9(3):385–392, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707377.

Boas:1951:BG

[Boa51] Marie Boas. Bacon and Gilbert. Journal of the History of Ideas, 12(3): 466–467, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707755.

Boas:1956:BRD

[Boa56] George Boas. Book review: La Disputa del Nuovo Mondo. Storia di una Polemica: 1750–1900 by Antonello Gerbi. Journal of the History of Ideas, 17(3):422–426, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707554.

Boas:1958:BRR

[Boa58] George Boas. Book review: Recent Books in the History of Philosophy by Hans M. Wolff. Journal of the History of Ideas, 19(4):571–592, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707926.

Boas:1959:BHP

[Boa59] George Boas. Bergson (1859–1941) and his predecessors. Journal of the History of Ideas, 20(4):503–514, October/December 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707889.

Boas:1963:BRC

[Boa63] George Boas. Book review: The Career of Philosophy from the Middle Ages to the Enlightenment, by John Herman Randall, Jr. Journal of the History of Ideas, 24(2):287–292, April/June 1963. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707851.

Boas:1964:BRB

[Boa64] George Boas. Book review: Bias and the History of Ideas. Journal of the History of Ideas, 25(3):451–457, July/September 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707912.

Boas:1967:BRT

[Boa67] George Boas. Book review: Tuzet: Cosmos and Imagination. Journal of the History of Ideas, 28(4):606–610, October/December 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708536.

Boas:1970:BRS

[Boa70] George Boas. Book review: The Study of Literature. Journal of the History of Ideas, 31(1):148–151, January/March 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708379.

Boas:1972:PPW

[Boa72] George Boas. To Philip P. Wiener. Journal of the History of Ideas, 33(3):355–357, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709039.

Bock:1951:HSM

[Boc51] Kenneth E. Bock. History and a science of man: an appreciation of George Cornewall Lewis. Journal of the History of Ideas, 12(4):599– 608, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707488.

Bochner:1965:WMG

[Boc65] Salomon Bochner. Why mathematics grows. Journal of the History of Ideas, 26(1):3–24, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708397.

Bock:1976:MPS

[Boc76] Kenneth E. Bock. The moral philosophy of Sir Henry Summer Maine. Journal of the History of Ideas, 37(1):147–154, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708716.

Boesche:1983:TCN

[Boe83] Roger Boesche. Tocqueville and Le Commerce: a newspaper expressing his unusual liberalism. Journal of the History of Ideas, 44(2):277–292, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709140.

Bok:2017:LIB

[Bok17] P. MacKenzie Bok. “The latest invasion from Britain”: Young Rawls and his community of American ethical theorists. Journal of the History of Ideas, 78(2):275–285, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/652827.

Bonner:1956:CWH

[Bon56] Thomas N. Bonner. Civil War historians and the “Needless War” doctrine. Journal of the History of Ideas, 17(2):193–216, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707742.

Bonadeo:1985:MW

[Bon85] Alfredo Bonadeo. Montaigne on war. Journal of the History of Ideas, 46(3):417–426, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709476.

Boodin:1943:DF

[Boo43] John Elof Boodin. The discovery of form. Journal of the History of Ideas, 4(2):177–192, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707323.

Boone:2000:CST

[Boo00] Rebecca Boone. Claude de Seyssel’s translations of ancient historians. Journal of the History of Ideas, 61(4):561–575, October 2000. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15091.

Borchardt:1968:ETN

[Bor68] Frank L. Borchardt. Etymology in tradition and in the Northern Renaissance. Journal of the History of Ideas, 29(3):415–429, July/ September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708452.

Bordeau:1971:JDI

[Bor71] Edward J. Bordeau. John Dewey’s ideas about the Great Depression. Journal of the History of Ideas, 32(1):67–84, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708325.

Bostetter:1970:CME

[Bos70] Edward E. Bostetter. Coleridge’s manuscript essay on the passions. Journal of the History of Ideas, 31(1):99–108, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708372.

Boss:1973:DSR

[Bos73] Ronald I. Boss. The development of social religion: a contradiction of French free thought. Journal of the History of Ideas, 34(4): 577–589, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708889.

Bos:1989:ELE

[Bos89] A. P. Bos. Exoterikoi Logoi and Enkyklioi Logoi in the Corpus Aristotelicum and the origin of the idea of the Enkyklios Paideia. Journal of the History of Ideas, 50(2):179–198, April/June 1989. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709731.

Botti:2017:RDB

[Bot17] Daniele Botti. Rawls on Dewey before the Dewey Lectures. Journal of the History of Ideas, 78(2):287–298, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/652828.

Bouwsma:1954:PSR

[Bou54] William J. Bouwsma. Postel and the significance of Renaissance cabalism. Journal of the History of Ideas, 15(2):218–232, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707768.

Bouwsma:1987:WB

[Bou87] William J. Bouwsma. Work on blumenberg. Journal of the History of Ideas, 48(2):347–354, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709562.

Boucher:1990:PHH

[Bou90] David Boucher. Practical Hegelianism: Henry Jones’s lecture tour of Australia. Journal of the History of Ideas, 51(3):423–452, July/ September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709623.

Boucher:1994:BIS

[Bou94] David Boucher. British idealism, the state, and international relations. Journal of the History of Ideas, 55(4):671–694, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709928.

Boucher:1997:SRG

[Bou97] David Boucher. The significance of R. G. Collingwood’s Principles of History. Journal of the History of Ideas, 58(2):309–330, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14940.

Bourke:2000:LAT

[Bou00] Richard Bourke. Liberty, authority, and trust in Burke’s idea of empire. Journal of the History of Ideas, 61(3):453–471, July 2000. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15079.

Bourg:2001:RME

[Bou01] Julian Bourg. The rhetoric of modal equivocacy in Cartesian transubstantiation. Journal of the History of Ideas, 62(1):121–140, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15098.

Boutcher:2003:ACR

[Bou03] Warren Boutcher. The analysis of culture revisited: Pure texts, applied texts, literary historicisms, cultural histories. Journal of the History of Ideas, 64(3):489–510, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/47656.

Bowers:1956:BSS

[Bow56] R. H. Bowers. Bacon’s spider simile. Journal of the History of Ideas, 17(1):133–135, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707692.

Bowler:1974:SFP

[Bow74] Peter J. Bowler. Sir Francis palgrave on natural theology. Journal of the History of Ideas, 35(1):144–147, January/March 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708750.

Bowler:1975:CME

[Bow75a] Peter J. Bowler. The changing meaning of “Evolution”. Journal of the History of Ideas, 36(1):95–114, January/March 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709013.

Bowler:1975:HSE

[Bow75b] Peter J. Bowler. Herbert Spencer and “Evolution” — an additional note. Journal of the History of Ideas, 36(2):367, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708934. See [Sim60].

Bowler:1976:MDC

[Bow76] Peter J. Bowler. Malthus, Darwin, and the concept of struggle. Journal of the History of Ideas, 37(4):631–650, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709028.

Bowles:1985:OPD

[Bow85] Paul Bowles. The origin of property and the development of Scottish historical science. Journal of the History of Ideas, 46(2):197–209, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709634.

Boyle:2002:PHA

[Boy02] Marjorie O’Rourke Boyle. Pure of heart: From ancient rites to Renaissance Plato. Journal of the History of Ideas, 63(1):41–62, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15141.

Boyle:2006:FVG

[Boy06] Deborah Boyle. Fame, virtue, and government: Margaret Cavendish on ethics and politics. Journal of the History of Ideas, 67(2):251– 289, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
198075.

Bracken:1956:BC

[Bra56] Harry M. Bracken. Berkeley and chambers. Journal of the History of Ideas, 17(1):120–126, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707689.

Bracken:1957:ABC

[Bra57] Harry M. Bracken. Andrew Baxter, critic of Berkeley. Journal of the History of Ideas, 18(2):183–204, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707623.

Bracken:1964:BS

[Bra64] Harry M. Bracken. Bayle not a sceptic? Journal of the History of Ideas, 25(2):169–180, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708010.

Braatz:1971:TNC

[Bra71] Werner E. Braatz. Two neo-conservative myths in Germany 1919– 32: The “Third Reich” and the “New State”. Journal of the History of Ideas, 32(4):569–584, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708978.

Branson:1979:JMS

[Bra79] Roy Branson. James Madison and the Scottish Enlightenment. Journal of the History of Ideas, 40(2):235–250, April/June 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709150.

Brauen:1982:AK

[Bra82] Fred Brauen. Athanasius Kircher (1602–1680). Journal of the History of Ideas, 43(1):129–134, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709164.

Brandist:2002:TRC

[Bra02] Craig Brandist. Two routes “to concreteness” in the work of the Bakhtin Circle. Journal of the History of Ideas, 63(3):521–537, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15161.

Braaksma:2024:TPS

[Bra24] Ymko Braaksma. Taking pragmatism seriously enough: Toward a deeper understanding of the British debate over pragmatism, ca. 1900–1910. Journal of the History of Ideas, 85(1):65–86, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/917116.

Breen:1952:GPD

[Bre52a] Quirinus Breen. Giovanni Pico Della Mirandola on the conflict of philosophy and rhetoric. Journal of the History of Ideas, 13(3):384– 412, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707604.

Breen:1952:MRG

[Bre52b] Quirinus Breen. Melancthon’s reply to G. Pico Della Mirandola. Journal of the History of Ideas, 13(3):413–426, June 1952. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707605.

Breslin:1961:PPA

[Bre61] Charles Breslin. Philosophy or philology: Auerbach and aesthetic historicism. Journal of the History of Ideas, 22(3):369–381, July/ September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708131.

Bredin:1972:ETS

[Bre72] Hugh Bredin. T. S. Eliot and Thomistic scholasticism. Journal of the History of Ideas, 33(2):299–306, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708875.

Breazeale:2001:FCP

[Bre01a] Daniel Breazeale. Fichte’s conception of philosophy as a “Pragmatic History of the Human Mind” and the contributions of Kant, Platner, and Maimon. Journal of the History of Ideas, 62(4):685–703, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15130.

Breckman:2001:EGC

[Bre01b] Warren Breckman. Eduard Gans and the crisis of Hegelianism. Journal of the History of Ideas, 62(3):543–564, July 2001. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/15122.

Breckman:2010:TTW

[Bre10] Warren Breckman. Times of theory: On writing the history of French theory. Journal of the History of Ideas, 71(3):339–361, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/386148.

Brinton:1940:NSU

[Bri40] Crane Brinton. The national socialists’ use of Nietzsche. Journal of the History of Ideas, 1(2):131–150, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707330.

Briefs:1941:EPR

[Bri41] Goetz A. Briefs. The economic philosophy of romanticism. Journal of the History of Ideas, 2(3):279–300, June 1941. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707132.

Brinton:1942:BRC

[Bri42] Crane Brinton. Book review: The Cultural Approach to History by Caroline F. Ware. Journal of the History of Ideas, 3(2):228–231, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707179.

Brinton:1953:SWW

[Bri53] Crane Brinton. Something went wrong: Three views of the heritage of the Early Nineteenth Century. Journal of the History of Ideas, 14(3): 457–462, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707813.

Bright:1979:ELR

[Bri79] Michael H. Bright. English literary romanticism and the Oxford movement. Journal of the History of Ideas, 40(3):385–404, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709244.

Bright:1985:PA

[Bri85] Michael Bright. The poetry of art. Journal of the History of Ideas, 46(2):259–277, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709638.

Brient:2000:HBH

[Bri00] Elizabeth Brient. Hans Blumenberg and Hannah Arendt on the “Unworldly Worldliness” of the Modern Age. Journal of the History of Ideas, 61(3):513–530, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15080.

Brinley:2023:LDR

[Bri23] Michael Brinley. Linguistic diplomacy: Roman Jakobson between East and West, 1956–68. Journal of the History of Ideas, 84(2):337– 363, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
886185.

Brower:1947:SA

[Bro47] Reuben A. Brower. Seven agamemnons. Journal of the History of Ideas, 8(4):383–405, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707426.

Brown:1953:MCS

[Bro53] Bernard E. Brown. Morris Cohen’s search for justice. Journal of the History of Ideas, 14(2):249–263, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707474.

Brown:1961:BRC

[Bro61] Harcourt Brown. Book review: Crocker’s Age of Crisis. Journal of the History of Ideas, 22(1):121–130, January/March 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707880.

Brown:1972:HLJ

[Bro72] Harcourt Brown. History and the learned journal. Journal of the History of Ideas, 33(3):365–378, July/September 1972. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709041.

Brodwin:1974:EVP

[Bro74] Stanley Brodwin. Emerson’s version of Plotinus: The flight to beauty. Journal of the History of Ideas, 35(3):465–483, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708794.

Brownley:1977:ARG

[Bro77] Martine Watson Brownley. Appearance and reality in Gibbon’s history. Journal of the History of Ideas, 38(4):651–666, October/ December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708693.

Brownley:1981:GAH

[Bro81] Martine Watson Brownley. Gibbon’s artistic and historical scope in the decline and fall. Journal of the History of Ideas, 42(4): 629–642, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709122.

Brown:1985:GTE

[Bro85] Harold I. Brown. Galileo on the telescope and the eye. Journal of the History of Ideas, 46(4):487–501, October/December 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709541.

Brown:1987:MSP

[Bro87] Theodore M. Brown. Medicine in the shadow of the Principia. Journal of the History of Ideas, 48(4):629–648, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709691.

Brown:1991:ETM

[Bro91] Gary I. Brown. The evolution of the term “Mixed Mathematics”. Journal of the History of Ideas, 52(1):81–102, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709583.

Brown:1992:DAS

[Bro92] Kevin L. Brown. Dating Adam Smith’s essay “Of the External Senses”. Journal of the History of Ideas, 53(2):333–337, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709877.

Brobjer:1997:NRP

[Bro97] Thomas H. Brobjer. Nietzsche’s reading and private library, 1885– 1889. Journal of the History of Ideas, 58(4):663–680, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14961.

Brown:1999:FGL

[Bro99] Vivienne Brown. The “Figure” of God and the limits to liberalism: a rereading of Locke’s essay and two treatises. Journal of the History of Ideas, 60(1):83–100, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15015.

Brobjer:2004:NVV

[Bro04a] Thomas H. Brobjer. Nietzsche’s view of the value of historical studies and methods. Journal of the History of Ideas, 65(2):301–322, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/173845.

Brown:2004:LEL

[Bro04b] Gregory Brown. Leibniz’s endgame and the ladies of the courts. Journal of the History of Ideas, 65(1):75–100, January 2004. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/168331.

Brosseder:2005:WWS

[Bro05] Claudia Brosseder. The writing in the Wittenberg sky: Astrology in Sixteenth-Century Germany. Journal of the History of Ideas, 66(4):557–576, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/191467.

Broad:2006:WIJ

[Bro06] Jacqueline Broad. A woman’s influence? John Locke and Damaris Masham on moral accountability. Journal of the History of Ideas, 67(3):489–510, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/202873.

Broad:2018:CCI

[Bro18] Jacqueline Broad. Conway and Charleton on the intimate presence of souls in bodies. Journal of the History of Ideas, 79(4):571–591, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/707773.

Brown:2024:SCJ

[Bro24] Angus Harwood Brown. Sieyès’s constitutional jury, the Pennsylvania Council of censors, and the debate on the conservative power in the French Revolution. Journal of the History of Ideas, 85(3):479– 508, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
933856.

Brush:1976:IIF

[Bru76] Stephen G. Brush. Irreversibility and indeterminism: Fourier to Heisenberg. Journal of the History of Ideas, 37(4):603–630, October/ December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709027.

Brickhouse:1983:OSM

[BS83] Thomas C. Brickhouse and Nicholas D. Smith. The origin of Socrates’ mission. Journal of the History of Ideas, 44(4):657–666, October/ December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709221.

Buchan:2021:NDB

[BS21] Bruce Buchan and Silvia Sebastiani. “No distinction of black or fair”: The natural history of race in Adam Ferguson’s lectures on moral philosophy. Journal of the History of Ideas, 82(2):207–229, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/790201.

Buickerood:2002:WER

[Bui02] James G. Buickerood. “The whole exercise of reason”: Charles Mein’s account of rationality. Journal of the History of Ideas, 63(4):639–658, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/39290.

Bullard:2005:RPK

[Bul05] Melissa Meriam Bullard. The Renaissance project of knowing: Lorenzo Valla and Salvatore Camporeale’s contributions to the Querelle between rhetoric and philosophy. Journal of the History of Ideas, 66(4):477–481, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/191468.

Bunzl:1995:PR

[Bun95] Martin Bunzl. Pragmatism to the rescue? Journal of the History of Ideas, 56(4):651–659, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709996.

Burroughs:1944:MFP

[Bur44] Josephine L. Burroughs. Marsilio Ficino, platonic theology. Journal of the History of Ideas, 5(2):227–242, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707387.

Burks:1949:CIH

[Bur49] Richard V. Burks. A conception of ideology for historians. Journal of the History of Ideas, 10(2):183–198, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707414.

Buranelli:1957:BRA

[Bur57a] Vincent Buranelli. Book review: Another Look at the Philosophes. Journal of the History of Ideas, 18(1):128–134, January 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707584.

Buranelli:1957:HPT

[Bur57b] Vincent Buranelli. The historical and political thought of boulainvilliers. Journal of the History of Ideas, 18(4):475–494, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707560.

Burns:1959:JMT

[Bur59] J. H. Burns. J. S. Mill and the term “Social Science”. Journal of the History of Ideas, 20(3):431–432, June/September 1959. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708120. See [Sen58, Igg59].

Burns:1960:WFP

[Bur60] J. H. Burns. “Winzerus”: a forgotten political writer. Journal of the History of Ideas, 21(1):124–130, January/March 1960. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708004.

Burns:1964:ETC

[Bur64] E. Bradford Burns. The Enlightenment in two colonial Brazilian libraries. Journal of the History of Ideas, 25(3):430–438, July/ September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707909.

Burton:1965:TRS

[Bur65] David H. Burton. Theodore Roosevelt’s social Darwinism and views on imperialism. Journal of the History of Ideas, 26(1):103–118, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708402.

Burnham:1974:MVC

[Bur74] Frederic B. Burnham. The more-Vaughan controversy: The revolt against philosophical enthusiasm. Journal of the History of Ideas, 35(1):33–49, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708741.

Burns:1981:ALF

[Bur81] Alfred Burns. Athenian literacy in the Fifth Century B.C. Journal of the History of Ideas, 42(3):371–387, July/September 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709182.

Burich:1987:HAS

[Bur87] Keith R. Burich. Henry Adams, the second law of thermodynamics, and the course of history. Journal of the History of Ideas, 48 (3):467–482, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709763.

Burnell:1993:PSU

[Bur93] Peter Burnell. The problem of service to unjust regimes in Augustine’s city of God. Journal of the History of Ideas, 54(2):177–188, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709977.

Burke:2003:IES

[Bur03] Peter Burke. Images as evidence in Seventeenth-Century Europe. Journal of the History of Ideas, 64(2):273–296, April 2003. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/43508.

Burgers:2011:MNM

[Bur11] Johannes Hendrikus Burgers. Max Nordau, Madison Grant, and racialized theories of ideology. Journal of the History of Ideas, 72 (1):119–140, January 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/413477.

Burke:2012:I

[Bur12] Martin J. Burke. Introduction. Journal of the History of Ideas, 73(1): 69, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
463137.

Butts:1959:HS

[But59] Robert E. Butts. Hume’s scepticism. Journal of the History of Ideas, 20(3):413–419, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708118.

Butterfield:1961:RRM

[But61] H. Butterfield. Reflections on religion and modern individualism. Journal of the History of Ideas, 22(1):33–46, January/March 1961.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707872.

Butterfield:1977:TEM

[But77] Herbert Butterfield. Toleration in early modern times. Journal of the History of Ideas, 38(4):573–584, October/December 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708688.

Butler:2006:IRP

[But06] Todd Wayne Butler. Image, rhetoric, and politics in the early Thomas Hobbes. Journal of the History of Ideas, 67(3):465–487, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/202874.

Butler:2015:MSP

[But15] Katherine Butler. Myth, science, and the power of music in the early decades of the Royal Society. Journal of the History of Ideas, 76(1):47– 68, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
566376.

Buxton:1984:IWJ

[Bux84] Michael Buxton. The influence of William James on John Dewey’s early work. Journal of the History of Ideas, 45(3):451–463, July/ September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709235.

Bensaude-Vincent:1988:WPT

[BV88] B. Bensaude-Vincent. When a physicist turns on philosophy Paul Langevin (1911–39). Journal of the History of Ideas, 49(2):319–338, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709503.

Bynack:1984:NWL

[Byn84] V. P. Bynack. Noah Webster’s linguistic thought and the idea of an American national culture. Journal of the History of Ideas, 45 (1):99–114, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709333.

Byrne:2006:HHM

[Byr06] James Steven Byrne. A humanist history of mathematics? Regiomontanus’s Padua oration in context. Journal of the History of Ideas, 67(1):41–61, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/194641.

Byrne:2023:PSI

[Byr23] Philippa Byrne. Portable scholasticism? The intellectual horizons of Gervase of Tilbury. Journal of the History of Ideas, 84(3):441– 464, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
901489.

Cadden:1995:SRM

[Cad95] Joan Cadden. Science and rhetoric in the Middle Ages: The natural philosophy of William of Conches. Journal of the History of Ideas, 56(1):1–24, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710004.

Cahoone:2017:MMR

[Cah17] Lawrence Cahoone. The metaphysics of Morris R. Cohen: From realism to objective relativism. Journal of the History of Ideas, 78(3): 449–471, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665523.

Caine:2007:SCV

[Cai07] Barbara Caine. Stefan Collini, Virginia Woolf, and the question of intellectuals in Britain. Journal of the History of Ideas, 68(3):369– 373, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
219816.

Caldwell:1962:ALT

[Cal62] R. L. Caldwell. Another look at Thomas Reid. Journal of the History of Ideas, 23(4):545–549, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708173.

Calinger:1969:NWC

[Cal69] Ronald S. Calinger. The Newtonian–Wolffian controversy: 1740–1759. Journal of the History of Ideas, 30(3):319–330, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708560.

Callahan:1986:RPM

[Cal86] John F. Callahan. Richard Peter McKeon 1900–1985. Journal of the History of Ideas, 47(4):653–662, October/December 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709724.

Calaresu:1997:IAR

[Cal97] Melissa Calaresu. Images of Ancient Rome in Late Eighteenth-Century Neapolitan historiography. Journal of the History of Ideas, 58(4):641–661, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14962.

Callisen:2012:GCI

[Cal12] Christian Thorsten Callisen. Georg Calixtus, Isaac Casaubon, and the consensus of antiquity. Journal of the History of Ideas, 73(1):1– 23, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
463134.

Callison:2022:PRE

[Cal22] William Callison. The politics of rationality in early neoliberalism: Max Weber, Ludwig von Mises, and the socialist calculation debate. Journal of the History of Ideas, 83(2):269–291, April 2022. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/855167.

Campbell:1970:MRA

[Cam70] Blair Campbell. La Mettrie: The robot and the automaton. Journal of the History of Ideas, 31(4):555–572, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708260.

Cameron:1984:HRP

[Cam84] David R. Cameron. The hero in Rousseau’s political thought. Journal of the History of Ideas, 45(3):397–419, July/September 1984. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709232.

Campbell:1986:FPR

[Cam86] Stuart L. Campbell. The four Paretos of Raymond Aron. Journal of the History of Ideas, 47(2):287–298, April/June 1986. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709815.

Camporeale:1996:LVO

[Cam96] Salvatore I. Camporeale. Lorenzo Valla’s oratio on the pseudodonation of Constantine: Dissent and innovation in Early Renaissance humanism. Journal of the History of Ideas, 57(1):9–26, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14889.

Campbell:2016:JPS

[Cam16] Ian W. S. Campbell. John Punch, Scotist holy war, and the Irish Catholic Revolutionary tradition in the Seventeenth Century. Journal of the History of Ideas, 77(3):401–421, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/626908.

Camp:2020:TMS

[Cam20] Pannill Camp. The theatre of moral sentiments: Neoclassical dramaturgy and Adam Smith’s impartial spectator. Journal of the History of Ideas, 81(4):555–576, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/772511.

Campbell:2022:JFF

[Cam22a] Ian Campbell. The Jewish family, forced baptism, and holy war in Early Modern Roman Scotism. Journal of the History of Ideas, 83(4):659–670, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865886.

Campbell:2022:RWN

[Cam22b] Ian Campbell. Rethinking war, nature, and supernature in early modern scholasticism: Introduction. Journal of the History of Ideas, 83(4):601–611, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865882.

Cannon:1961:JHI

[Can61] Walter F. Cannon. John Herschel and the idea of science. Journal of the History of Ideas, 22(2):215–239, April/June 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707834.

Cannon:1964:NRS

[Can64] Walter F. Cannon. The normative role of science in Early Victorian thought. Journal of the History of Ideas, 25(4):487–502, October/ December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708182.

Canavan:1973:RBJ

[Can73] Thomas L. Canavan. Robert Burton, Jonathan Swift, and the tradition of anti-Puritan invective. Journal of the History of Ideas, 34(2):227–242, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708727.

Cantor:1977:BRM

[Can77] G. N. Cantor. Berkeley, Reid, and the mathematization of mid-Eighteenth-Century optics. Journal of the History of Ideas, 38(3): 429–448, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708673.

Canovan:1984:BUJ

[Can84] Margaret Canovan. The un-Benthamite utilitarianism of Joseph Priestley. Journal of the History of Ideas, 45(3):435–450, July/ September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709234.

Capek:1954:JEC

[Cap54] Milic Capek. James’s early criticism of the automaton theory. Journal of the History of Ideas, 15(2):260–279, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707771.

Carr:1962:RTJ

[Car62] Robert Carr. The religious thought of John Stuart Mill: a study in reluctant scepticism. Journal of the History of Ideas, 23(4): 475–495, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708170.

Carvin:1972:LMC

[Car72] Walter P. Carvin. Leibniz on motion and creation. Journal of the History of Ideas, 33(3):425–438, July/September 1972. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709044.

Carlisle:1974:BTS

[Car74] Robert B. Carlisle. The birth of technocracy: Science, society, and Saint-Simonians. Journal of the History of Ideas, 35(3):445–464, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708793.

Carroll:1978:CMF

[Car78] P. Thomas Carroll. Correction to Margaret A. Fay’s article. Journal of the History of Ideas, 39(4):675–676, October/December 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709450. See [Fay78].

Cartwright:1984:KSN

[Car84] David E. Cartwright. Kant, Schopenhauer, and Nietzsche on the morality of pity. Journal of the History of Ideas, 45(1):83–98, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709332.

Carrithers:1986:MPH

[Car86] David Carrithers. Montesquieu’s philosophy of history. Journal of the History of Ideas, 47(1):61–80, January/March 1986. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709595.

Carrithers:1991:VRM

[Car91] David W. Carrithers. Not so virtuous republics: Montesquieu, Venice, and the theory of aristocratic republicanism. Journal of the History of Ideas, 52(2):245–268, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709527.

Carignan:2003:ARV

[Car03] Michael Carignan. Analogical reasoning in Victorian historical epistemology. Journal of the History of Ideas, 64(3):445–464, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/47657.

Carlin:2007:SPL

[Car07] Laurence Carlin. Selecting a phenomenalism: Leibniz, Berkeley, and the science of happiness. Journal of the History of Ideas, 68(1):57– 78, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
210686.

Carter:2008:HSV

[Car08] Stephen G. Carter. The “Historical Solution” versus the “Philosophical Solution”: The political commentary of Christopher Dawson and Jacques Maritain, 1927–1939. Journal of the History of Ideas, 69(1):93–115, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/231007.

Carhart:2019:PIL

[Car19] Michael C. Carhart. Practices of intellectual labor in the Republic of Letters: Leibniz and Edward Bernard on language and European origins. Journal of the History of Ideas, 80(3):365–386, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/729719.

Carhart:2023:DRK

[Car23] Michael C. Carhart. Donald R. Kelley (1931–2023). Journal of the History of Ideas, 84(4):v–vi, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/909531.

Cassirer:1942:GPDa

[Cas42a] Ernst Cassirer. Giovanni Pico Della Mirandola: a study in the history of Renaissance ideas. Journal of the History of Ideas, 3(2):123–144, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707173.

Cassirer:1942:GPDb

[Cas42b] Ernst Cassirer. Giovanni Pico Della Mirandola: a study in the history of Renaissance ideas. Journal of the History of Ideas, 3(3):319–346, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707307.

Cassirer:1945:BRF

[Cas45] Ernst Cassirer. Book review: Ficino’s Place in Intellectual History. Journal of the History of Ideas, 6(4):483–501, October 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707347.

Caspari:1947:ESF

[Cas47] Fritz Caspari. Erasmus on the social functions of Christian humanism. Journal of the History of Ideas, 8(1):78–106, January 1947. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707442.

Castell:1948:BRN

[Cas48] Alburey Castell. Book review: Northrop’s Meeting of East and West. Journal of the History of Ideas, 9(2):237–244, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707230.

Cassirer:1963:ACE

[Cas63] Thomas Cassirer. Awareness of the city in the encyclopedie. Journal of the History of Ideas, 24(3):387–396, July/September 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708214.

Cashdollar:1978:ACA

[Cas78] Charles D. Cashdollar. Auguste Comte and the American reformed theologians. Journal of the History of Ideas, 39(1):61–79, January/ March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709072.

Caspar:1981:ASW

[Cas81] Ruth Caspar. “All shall be well”: Prototypical symbols of hope. Journal of the History of Ideas, 42(1):139–150, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709421.

Caserta:1983:CM

[Cas83] Ernesto G. Caserta. Croce and Marxism. Journal of the History of Ideas, 44(1):141–149, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709310.

Cassedy:1988:MRR

[Cas88] Steven Cassedy. Mathematics, relationalism, and the rise of modern literary aesthetics. Journal of the History of Ideas, 49(1): 109–132, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709706.

Casillo:1992:LMO

[Cas92] Robert Casillo. Lewis Mumford and the organicist concept in social thought. Journal of the History of Ideas, 53(1):91–116, January/ March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709912.

Cassedy:2010:BRH

[Cas10] Steven Cassedy. Beethoven the Romantic: How E. T. A. Hoffmann got it right. Journal of the History of Ideas, 71(1):1–37, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/370100.

Caton:1973:KDE

[Cat73a] Hiram Caton. Kennington on Descartes’ evil genius. Journal of the History of Ideas, 34(4):639–641, October/December 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708894. See reply [Ken73].

Caton:1973:RCE

[Cat73b] Hiram Caton. Rejoinder: The cunning of the evil demon. Journal of the History of Ideas, 34(4):643–644, October/December 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708896.

Catana:2010:LRB

[Cat10] Leo Catana. Lovejoy’s readings of Bruno: Or how Nineteenth-Century history of philosophy was “Transformed” into the history of ideas. Journal of the History of Ideas, 71(1):91–112, January 2010. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/370097.

Caudill:1994:BEP

[Cau94] Edward Caudill. The bishop-eaters: The publicity campaign for Darwin and on the origin of species. Journal of the History of Ideas, 55(3): 441–460, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709849.

Cavarzere:2016:LNA

[Cav16] Marco Cavarzere. Land and nation: The ancient modernity of national geography (Piedmont, 1750–1800). Journal of the History of Ideas, 77(2):203–225, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/620979.

Celenza:2001:LAF

[Cel01] Christopher S. Celenza. Late antiquity and the Florentine Renaissance: Historiographical parallels. Journal of the History of Ideas, 62(1):17–35, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15099.

Celenza:2005:LVT

[Cel05] Christopher S. Celenza. Lorenzo Valla and the traditions and transmissions of philosophy. Journal of the History of Ideas, 66(4):483–506, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/191469.

Celenza:2014:ICI

[Cel14] Christopher S. Celenza. Ideas in context and the idea of Renaissance philosophy. Journal of the History of Ideas, 75(4):653–666, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/557238.

Censer:1989:BRR

[Cen89] Jack R. Censer. Book review: Revitalizing the Intellectual History of the French Revolution. Journal of the History of Ideas, 50(4): 652–666, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709803.

Cervantes:2005:CIC

[Cer05] Fernando Cervantes. Cervantes in Italy: Christian humanism and the visual impact of Renaissance Rome. Journal of the History of Ideas, 66(3):325–350, July 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/188682.

Cesar:2004:PAI

[Ces04] Floriano Jonas Cesar. Popular autonomy and Imperial power in bartolus of saxoferrato: an intrinsic connection. Journal of the History of Ideas, 65(3):369–381, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/177663.

Ceserani:2005:NIP

[Ces05] Giovanna Ceserani. Narrative, interpretation, and plagiarism in Mr. Roberton’s 1778 History of Ancient Greece. Journal of the History of Ideas, 66(3):413–436, July 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/188683.

Colp:1979:ISD

[CF79] Ralph Colp, Jr. and Margaret A. Fay. Independent scientific discoveries and the “Darwin-Marx” letter. Journal of the History of Ideas, 40(3):479, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709252.

Collins:1995:AGC

[CH95] Stephen L. Collins and James Hoopes. Anthony Giddens and Charles Sanders Peirce: History, theory, and a way out of the linguistic culde-sac. Journal of the History of Ideas, 56(4):625–650, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709995.

Colon:2015:ICN

[CH15] Gabriel Alejandro Torres Colón and Charles A. Hobbs. The intertwining of culture and nature: Franz Boas, John Dewey, and Deweyan strands of American anthropology. Journal of the History of Ideas, 76(1):139–162, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/566380.

Chalmers:1958:RSB

[Cha58a] David Chalmers. Ray Stannard Baker’s search for reform. Journal of the History of Ideas, 19(3):422–434, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708047.

Chambers:1958:BPT

[Cha58b] Clarke A. Chambers. The belief in progress in Twentieth-Century America. Journal of the History of Ideas, 19(2):197–224, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707935.

Charlton:1960:HAS

[Cha60] Kenneth Charlton. Holbein’s ambassadors’ and Sixteenth-Century education. Journal of the History of Ideas, 21(1):99–109, January/ March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708001.

Chambers:1963:CPL

[Cha63] A. B. Chambers. Chaos in Paradise Lost. Journal of the History of Ideas, 24(1):55–84, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707859.

Chapin:1966:CAC

[Cha66] Seymour L. Chapin. A case of arms control in the French Enlightenment. Journal of the History of Ideas, 27(2):285–295, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708645.

Chamberlin:1981:ACM

[Cha81] J. E. Chamberlin. An anatomy of cultural melancholy. Journal of the History of Ideas, 42(4):691–705, October/December 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709128.

Chaplin:2017:CNS

[Cha17] Joyce E. Chaplin. Can the nonhuman speak?: Breaking the chain of being in the Anthropocene. Journal of the History of Ideas, 78(4):509– 529, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
676144.

Chamedes:2023:HDT

[Cha23a] Giuliana Chamedes. How to do things with words: Antifascism as a differentially mobilizing ideology, from the popular front to the black power movement. Journal of the History of Ideas, 84(1):127–155, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
879543.

Chaplin:2023:HIR

[Cha23b] Joyce E. Chaplin. Historians of ideas rush in where stratigraphers fear to tread. Journal of the History of Ideas, 84(4):775–801, October 2023.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/909538.

Chakrabarti:2024:FIR

[Cha24] Upal Chakrabarti. “Facts” and “Ideas”: Richard Jones, William Whewell, and the entangled histories of science and political economy in Early Nineteenth-Century Britain. Journal of the History of Ideas, 85(3):509–537, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/933857.

Cherniss:1951:CEP

[Che51] Harold Cherniss. The characteristics and effects of presocratic philosophy. Journal of the History of Ideas, 12(3):319–345, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707750.

Cherry:1956:RCP

[Che56] George L. Cherry. The rôle of the Convention Parliament (1688–89) in parliamentary supremacy. Journal of the History of Ideas, 17(3):390– 406, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707551.

Cherniavsky:1959:KBA

[Che59] Michael Cherniavsky. Khan or Basileus: an aspect of Russian mediaeval political theory. Journal of the History of Ideas, 20(4): 459–476, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707886.

Cherno:1963:FMW

[Che63] Melvin Cherno. Feuerbach’s “Man is what He Eats”: a rectification. Journal of the History of Ideas, 24(3):397–406, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708215.

Chinard:1940:PAC

[Chi40] Gilbert Chinard. Polybius and the American Constitution. Journal of the History of Ideas, 1(1):38–58, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707009.

Ching:1974:TIC

[Chi74] Julia Ching. Truth and ideology: The Confucian way (tao) and its transmission (Tao-T’Ung). Journal of the History of Ideas, 35 (3):371–388, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708789.

Ching:1984:BRA

[Chi84] Julia Ching. Book review: Adventures in Cross-Cultural Sensibilities: Some Recent Studies of Chinese and Comparative Philosophy. Journal of the History of Ideas, 45(3):476–482, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709237.

Chouraqui:2015:PUS

[Cho15] Frank Chouraqui. “A principle of universal strife”: Ricoeur and Merleau-Ponty’s critiques of Marxist universalism, 1953–1956. Journal of the History of Ideas, 76(3):467–490, July 2015. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/587274.

Chroust:1945:CMD

[Chr45] Anton-Hermann Chroust. A contribution to the medieval discussion: Utrum Aristoteles Sit Salvatus. Journal of the History of Ideas, 6(2): 231–238, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707365.

Chroust:1946:FLJ

[Chr46] Anton-Hermann Chroust. The function of law and justice in the ancient world and the Middle Ages. Journal of the History of Ideas, 7(3): 298–320, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707403.

Chroust:1954:TPA

[Chr54] Anton-Hermann Chroust. Treason and patriotism in Ancient Greece. Journal of the History of Ideas, 15(2):280–288, April 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707772.

Christensen:1959:TNH

[Chr59] Merton A. Christensen. Taylor of Norwich and the higher criticism. Journal of the History of Ideas, 20(2):179–194, April 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707818.

Christensen:1964:NHP

[Chr64] Darrel E. Christensen. Nelson and Hegel on the philosophy of history. Journal of the History of Ideas, 25(3):439–444, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707910.

Christian:1971:MEF

[Chr71] Lynda Gregorian Christian. The metamorphoses of Erasmus’ “Folly”. Journal of the History of Ideas, 32(2):289–294, April/June 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708282.

Christian:1973:MET

[Chr73] William Christian. The moral economics of Tom Paine. Journal of the History of Ideas, 34(3):367–380, July/September 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708958.

Christensen:1989:MTS

[Chr89] Thomas Christensen. Music theory as scientific propaganda: The case of D’Alembert’s Élémens De Musique. Journal of the History of Ideas, 50(3):409–427, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709569.

Christl:2023:AOC

[Chr23] Robert Christl. Anarchism in one country: Diego Abad de Santillán and the invention of participatory national economic planning in interwar anarchism. Journal of the History of Ideas, 84(2):313–336, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/886184.

Cronin:1960:ETT

[CJ60] T. J. Cronin and S. J. Eternal truths in the thought of Descartes and of his adversary. Journal of the History of Ideas, 21(4):553–559, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708102.

Cassirer:1943:SRQ

[CJK^{+}43] Ernst Cassirer, Francis R. Johnson, Paul Oskar Kristeller, Dean P. Lockwood, and Lynn Thorndike. Some remarks on the question of the originality of the Renaissance. Journal of the History of Ideas, 4(1):49–74, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707236.

Clarke:1957:PDE

[Cla57] Jack Alden Clarke. The pastors of the desert on the eve of the French Revolution. Journal of the History of Ideas, 18(1):113–119, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707582.

Clauss:1982:JWE

[Cla82] Sidonie Clauss. John Wilkins’ essay toward a real character: Its place in the Seventeenth-Century episteme. Journal of the History of Ideas, 43(4):531–553, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709342.

Clarke:1985:QBW

[Cla85] Larry R. Clarke. The Quaker background of William Bartram’s view of nature. Journal of the History of Ideas, 46(3):435–448, July/ September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709478.

Claeys:1986:ISS

[Cla86] Gregory Claeys. “Individualism,” “Socialism,” and “Social Science”: Further notes on a process of conceptual formation, 1800–1850. Journal of the History of Ideas, 47(1):81–93, January/March 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709596.

Cladis:1992:DIS

[Cla92] Mark S. Cladis. Durkheim’s individual in society: a sacred marriage? Journal of the History of Ideas, 53(1):71–90, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709911.

Clark:1998:WIW

[Cla98] S. H. Clark. “The whole internal world his own”: Locke and metaphor reconsidered. Journal of the History of Ideas, 59(2):241–265, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14981.

Claeys:2000:SFO

[Cla00] Gregory Claeys. The “Survival of the Fittest” and the origins of social Darwinism. Journal of the History of Ideas, 61(2):223–240, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15068.

Clark:2011:AAA

[Cla11] Frederic Clark. Authenticity, antiquity, and authority: Dares Phrygius in Early Modern Europe. Journal of the History of Ideas, 72(2): 183–207, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
428851.

Claessens:2015:DBI

[Cla15] Guy A. J. Claessens. The drawing board of imagination: Federico Commandino and John Philoponus. Journal of the History of Ideas, 76(4):499–515, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/597344.

Clavey:2023:STC

[Cla23] Charles H. Clavey. “The stereotype takes care of everything”: Labor Antisemitism and critical theory during World War II. Journal of the History of Ideas, 84(4):711–742, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/909536.

Clemmer:1969:HTP

[Cle69] Robert Clemmer. Historical transcendentalism in Pennsylvania. Journal of the History of Ideas, 30(4):579–592, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708612.

Close:1969:CTA

[Clo69] A. J. Close. Commonplace theories of art and nature in classical antiquity and in the Renaissance. Journal of the History of Ideas, 30 (4):467–486, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708606.

Close:1971:PTA

[Clo71] Anthony J. Close. Philosophical theories of art and nature in classical antiquity. Journal of the History of Ideas, 32(2):163–184, April/ June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708275.

Cloeren:1975:NAH

[Clo75] Hermann J. Cloeren. The neglected analytical heritage. Journal of the History of Ideas, 36(3):513–529, July/September 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708660.

Clouatre:1984:CCF

[Clo84] Dallas L. Clouatre. The concept of class in French culture prior to the revolution. Journal of the History of Ideas, 45(2):219–244, April/ June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709289. See

errata [Ano84f].

Chen-Morris:2005:SIO

[CM05] Raz Chen-Morris. Shadows of instruction: Optics and classical authorities in Kepler’s Somnium. Journal of the History of Ideas, 66(2): 223–243, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
187185.

Ng:1993:TCT

[cN93] On cho Ng. A tension in Ch’ing thought: “Historicism” in Seventeenth- and Eighteenth-Century Chinese thought. Journal of the History of Ideas, 54(4):561–583, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709824.

Coates:1950:BLF

[Coa50a] Willson H. Coates. Benthamism, laissez faire, and collectivism. Journal of the History of Ideas, 11(3):357–363, June 1950. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707736.

Coates:1950:BRL

[Coa50b] Willson H. Coates. Book review: Liberalism and the Challenge of Fascism: Social Forces in England and France (1815–1870) by J. Salwyn Schapiro. Journal of the History of Ideas, 11(1):119–124, January 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707455.

Coats:1961:ASC

[Coa61] A. W. Coats. American scholarship comes of age: The Louisiana Purchase Exposition 1904. Journal of the History of Ideas, 22(3): 404–417, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708136.

Coates:1977:CDH

[Coa77] J. D. Coates. Coleridge’s debt to Harrington: a discussion of zapolya. Journal of the History of Ideas, 38(3):501–508, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708679.

Cochrane:1958:SM

[Coc58] Eric W. Cochrane. The settecento medievalists. Journal of the History of Ideas, 19(1):35–61, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707951.

Cochrane:1962:FLI

[Coc62] Eric W. Cochrane. French literature and the Italian tradition in Eighteenth-Century Tuscany. Journal of the History of Ideas, 23 (1):61–76, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708057.

Coen:2016:BTP

[Coe16] Deborah R. Coen. Big is a thing of the past: Climate change and methodology in the history of ideas. Journal of the History of Ideas, 77(2):305–321, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/620983.

Cogley:2007:SOK

[Cog07] Richard W. Cogley. “Some other kinde of being and condition”: The controversy in Mid-Seventeenth-Century England over the peopling of Ancient America. Journal of the History of Ideas, 68(1):35–56, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
210687.

Cohen:1940:BRF

[Coh40] Morris R. Cohen. Book review: Forces in American Criticism: a Study in the History of American Literary Thought by Bernard Smith. Journal of the History of Ideas, 1(2):241–251, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707336.

Cohen:1942:CAH

[Coh42] Morris R. Cohen. Causation and its application to history. Journal of the History of Ideas, 3(1):12–29, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707459.

Cohen:1948:HCC

[Coh48] Felix S. Cohen. The Holmes–Cohen correspondence. Journal of the History of Ideas, 9(1):3–52, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707393.

Cohen:1949:BRL

[Coh49] Felix S. Cohen. Book review: Legal Philosophy from Plato to Hegel. Journal of the History of Ideas, 10(4):575–579, October 1949. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707191.

Cohen:1954:SRB

[Coh54] I. Bernard Cohen. Some recent books on the history of science. Journal of the History of Ideas, 15(1):163–192, January 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707656.

Cohen:1962:APE

[Coh62] Maurice H. Cohen. The Aporias in Plato’s early dialogues. Journal of the History of Ideas, 23(2):163–174, April/June 1962. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708153.

Cohen:1968:BP

[Coh68] Gerald A. Cohen. Bourgeois and proletarians. Journal of the History of Ideas, 29(2):211–230, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708577.

Cohen:1976:ECO

[Coh76] I. Bernard Cohen. The Eighteenth-Century origins of the concept of scientific revolution. Journal of the History of Ideas, 37(2):257–288, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708824.

Cohen:1978:ARN

[Coh78] Lester H. Cohen. The American Revolution and natural law theory. Journal of the History of Ideas, 39(3):491–502, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709392.

Cohen:1980:SHR

[Coh80] L. Jonathan Cohen. Some historical remarks on the Baconian conception of probability. Journal of the History of Ideas, 41(2):219–231, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709457.

Cohen:1987:NTL

[Coh87] I. Bernard Cohen. Newton’s third law and universal gravity. Journal of the History of Ideas, 48(4):571–593, October/December 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709688.

Cohen:1994:HHT

[Coh94] I. Bernard Cohen. Harrington and Harvey: a theory of the state based on the new physiology. Journal of the History of Ideas, 55(2):187–210, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709896.

Colie:1959:SEE

[Col59] Rosalie L. Colie. Spinoza and the Early English deists. Journal of the History of Ideas, 20(1):23–46, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707965.

Colish:1962:MGV

[Col62] Marcia L. Colish. The mime of God: Vives on the nature of man. Journal of the History of Ideas, 23(1):3–20, January/March 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708054.

Colman:1966:NPM

[Col66] S. J. Colman. Nietzsche as politique et moraliste. Journal of the History of Ideas, 27(4):549–574, October/December 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708340.

Colish:1971:ILM

[Col71] Marcia L. Colish. The idea of liberty in Machiavelli. Journal of the History of Ideas, 32(3):323–350, July/September 1971. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708350.

Colp:1974:CBK

[Col74] Ralph Colp, Jr. The contacts between Karl Marx and Charles Darwin. Journal of the History of Ideas, 35(2):329–338, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708767.

Coleman:1976:PCE

[Col76] William Coleman. Providence, capitalism, and environmental degradation: English apologetics in an era of economic revolution. Journal of the History of Ideas, 37(1):27–44, January/March 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708707.

Coleman:1982:JJW

[Col82] Francis X. J. Coleman. John Jay on war. Journal of the History of Ideas, 43(1):145–151, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709166.

Colish:1999:RRM

[Col99] Marcia L. Colish. Republicanism, religion, and Machiavelli’s Savonarolan moment. Journal of the History of Ideas, 60(4):597–616, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15046.

Collini:2007:AR

[Col07] Stefan Collini. Author’s response. Journal of the History of Ideas, 68(3):395–405, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/219817.

Collins:2009:ITH

[Col09] Jeffrey R. Collins. Interpreting Thomas Hobbes in competing contexts. Journal of the History of Ideas, 70(1):165–180, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/258199.

Colish:2010:VB

[Col10] Marcia Colish. Virginia Brown 1940–2009. Journal of the History of Ideas, 71(1):161–162, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/370101.

Colish:2016:ALL

[Col16] Marcia L. Colish. 2015 Arthur O. Lovejoy Lecture the boys on the beach: Children’s games and baptismal grace in medieval thought. Journal of the History of Ideas, 77(3):359–378, July 2016. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/626906.

Comito:1971:RGD

[Com71] Terry Comito. Renaissance gardens and the discovery of paradise. Journal of the History of Ideas, 32(4):483–506, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708973.

Conroy:1960:BEA

[Con60] Graham P. Conroy. Berkeley and education in America. Journal of the History of Ideas, 21(2):211–221, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708195.

Conroy:1961:GBM

[Con61] Graham P. Conroy. George Berkeley on moral demostration. Journal of the History of Ideas, 22(2):205–214, April/June 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707833.

Constantin:1979:PED

[Con79] Charles Constantin. The Puritan ethic and the dignity of labor: Hierarchy vs. equality. Journal of the History of Ideas, 40(4): 543–561, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709358.

Conniff:1982:RHE

[Con82] James Conniff. Reason and history in early Whig thought: The case of Algernon Sidney. Journal of the History of Ideas, 43(3):397–416, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709430.

Conniff:1986:EBR

[Con86] James Conniff. Edmund Burke’s reflections on the coming revolution in Ireland. Journal of the History of Ideas, 47(1):37–59, January/ March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709594.

Connell:1996:LVO

[Con96] William J. Connell. Lorenzo Valla’s oratio on the pseudo-donation of Constantine: Dissent and innovation in Early Renaissance humanism. Journal of the History of Ideas, 57(1):1–7, January 1996. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14890.

Conniff:1999:EBH

[Con99] James Conniff. Edmund Burke and his critics: The case of Mary Wollstonecraft. Journal of the History of Ideas, 60(2):299–318, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15024.

Considine:2014:JJF

[Con14] John Considine. John Jamieson, Franz Passow, and the double invention of lexicography on historical principles. Journal of the History of Ideas, 75(2):261–281, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/544561.

Cooke:1973:JL

[Coo73] J. W. Cooke. Jefferson on liberty. Journal of the History of Ideas, 34 (4):563–576, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708888.

Cooper:1981:JBE

[Coo81] Robert Alan Cooper. Jeremy Bentham, Elizabeth Fry, and English prison reform. Journal of the History of Ideas, 42(4):675–690, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709127.

Cook:1999:BMT

[Coo99] Harold John Cook. Bernard Mandeville and the therapy of “The Clever Politician”. Journal of the History of Ideas, 60(1):101–124, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15016.

Cook:2014:MEE

[Coo14] Simon John Cook. The making of the English: English history, British identity, Aryan villages, 1870–1914. Journal of the History of Ideas, 75(4):629–649, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/557236.

Cope:1951:WJC

[Cop51] Jackson I. Cope. William James’s correspondence with Daniel Coit Gilman, 1877–1881. Journal of the History of Ideas, 12(4):609–627, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707489.

Copenhaver:1991:TTF

[Cop91] Brian P. Copenhaver. A tale of two fishes: Magical objects in natural history from antiquity through the scientific revolution. Journal of the History of Ideas, 52(3):373–398, July/September 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710043.

Copenhaver:2005:VOC

[Cop05] Brian P. Copenhaver. Valla our contemporary: Philosophy and philology. Journal of the History of Ideas, 66(4):507–525, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/191471.

Corbett:1975:CLL

[Cor75a] Theodore G. Corbett. The cult of lipsius: a leading source of Early Modern Spanish statecraft. Journal of the History of Ideas, 36 (1):139–152, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709016.

Corr:1975:CWL

[Cor75b] Charles A. Corr. Christian Wolff and Leibniz. Journal of the History of Ideas, 36(2):241–262, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708926.

Corcoran:1977:BOV

[Cor77] Paul E. Corcoran. The bourgeois and other villains. Journal of the History of Ideas, 38(3):477–485, July/September 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708677.

Corrigan:1980:BLL

[Cor80] Timothy J. Corrigan. Biographia literaria and the language of science. Journal of the History of Ideas, 41(3):399–419, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709401.

Cornilliat:1998:ERT

[Cor98] François Cornilliat. Exemplarities: a response to timothy hampton and karlheinz stierle. Journal of the History of Ideas, 59(4):613–624, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15001.

Corrigan:2010:MME

[Cor10] John Michael Corrigan. The metempsychotic mind: Emerson and consciousness. Journal of the History of Ideas, 71(3):433–455, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/386152.

Corrias:2018:WES

[Cor18] Anna Corrias. When the eyes are shut: The strange case of Girolamo Cardano’s Idolum in Somniorum Synesiorum Libri IIII (1562). Journal of the History of Ideas, 79(2):179–197, April 2018. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/693300.

Cottle:1979:JAV

[Cot79] Charles E. Cottle. Justice as artificial virtue in Hume’s treatise. Journal of the History of Ideas, 40(3):457–466, July/September 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709248.

Cotton:1981:JHT

[Cot81] James Cotton. James Harrington and Thomas Hobbes. Journal of the History of Ideas, 42(3):407–421, July/September 1981. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709184.

Cotkin:1994:MGP

[Cot94] George Cotkin. Middle-ground pragmatists: The popularization of philosophy in American culture. Journal of the History of Ideas, 55(2): 283–302, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709900.

Cotkin:2008:CAM

[Cot08a] George Cotkin. A conversation about morals and history. Journal of the History of Ideas, 69(3):493–497, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/242282.

Cotkin:2008:HMT

[Cot08b] George Cotkin. History’s moral turn. Journal of the History of Ideas, 69(2):293–315, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/235867.

Coulton:1944:HBM

[Cou44] G. G. Coulton. The historical background of Maritain’s humanism. Journal of the History of Ideas, 5(4):415–433, October 1944. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707079.

Coudert:1975:CPK

[Cou75] Allison Coudert. A Cambridge Platonist’s Kabbalist nightmare. Journal of the History of Ideas, 36(4):633–652, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708993.

Courtenay:1983:LMN

[Cou83] William J. Courtenay. Late medieval nominalism revisited: 1972– 1982. Journal of the History of Ideas, 44(1):159–164, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709312.

Courtenay:1987:AML

[Cou87] William J. Courtenay. Antiqui and moderni in late medieval thought. Journal of the History of Ideas, 48(1):3–10, January/March 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709608.

Cousins:2004:HFE

[Cou04] A. D. Cousins. Humanism, female education, and myth: Erasmus, Vives, and More’s To Candidus. Journal of the History of Ideas, 65(2):213–230, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/173846.

Campbell:1981:UAS

[CR81a] T. D. Campbell and I. S. Ross. The utilitarianism of Adam Smith’s policy advice. Journal of the History of Ideas, 42(1):73–92, January/ March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709418.

Cook:1981:PEH

[CR81b] Daniel J. Cook and Henry Rosemont, Jr. The pre-established harmony between Leibniz and Chinese thought. Journal of the History of Ideas, 42(2):253–267, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709319.

Cramer:1940:ICH

[Cra40] Frederick H. Cramer. Isolationism: a case-history. Journal of the History of Ideas, 1(4):459–493, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707125.

Craig:1944:LEP

[Cra44] Hardin Craig. Of the laws of ecclesiastical polity — first form. Journal of the History of Ideas, 5(1):91–104, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707103.

Cramer:1945:BCA

[Cra45] Frederick H. Cramer. Bookburning and censorship in Ancient Rome: a chapter from the history of freedom of speech. Journal of the History of Ideas, 6(2):157–196, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707362.

Craig:1992:NMM

[Cra92] Campbell Craig. The new meaning of modern war in the thought of Reinhold niebuhr. Journal of the History of Ideas, 53(4):687– 701, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709944.

Creel:1949:SC

[Cre49] Herrlee Glessner Creel. Sinism — a clarification. Journal of the History of Ideas, 10(1):135–140, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707205. See reply [Pri49].

Cress:1979:RWR

[Cre79] Lawrence Delbert Cress. Radical whiggery on the role of the military: Ideological roots of the American Revolutionary militia. Journal of the History of Ideas, 40(1):43–60, January/March 1979. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709259.

Creed:1983:RWK

[Cre83] Walter G. Creed. René Wellek and Karl Popper on the mode of existence of ideas in literature and science. Journal of the History of Ideas, 44(4):639–656, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709220.

Cremer:1995:PTE

[Cre95] Douglas J. Cremer. Protestant theology in Early Weimar Germany: Barth, Tillich, and Bultmann. Journal of the History of Ideas, 56(2): 289–307, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709839.

Crites:1966:BRC

[Cri66] Stephen D. Crites. Book review: A Critique of Kaufmann’s Hegel. Journal of the History of Ideas, 27(2):296–307, April/June 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708646.

Crimmins:1986:BRA

[Cri86] James E. Crimmins. Bentham on religion: Atheism and the secular society. Journal of the History of Ideas, 47(1):95–110, January/ March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709597.

Crimmins:1994:BPR

[Cri94] James E. Crimmins. Bentham’s political radicalism reexamined. Journal of the History of Ideas, 55(2):259–281, April 1994. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709899.

Crimmins:2002:BHI

[Cri02] James E. Crimmins. Bentham and Hobbes: an issue of influence. Journal of the History of Ideas, 63(4):677–696, October 2002. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/39292.

Crocker:1952:DSE

[Cro52] Lester G. Crocker. The discussion of suicide in the Eighteenth Century. Journal of the History of Ideas, 13(1):47–72, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707725.

Crocker:1953:PTF

[Cro53a] Lester Gilbert Crocker. The problem of truth and falsehood in the age of Enlightenment. Journal of the History of Ideas, 14(4):575–603, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707703.

Cross:1953:IPC

[Cro53b] Whitney R. Cross. Ideas in politics: The conservation policies of the two roosevelts. Journal of the History of Ideas, 14(3):421–438, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707810.

Crowe:1959:UUP

[Cro59] Charles R. Crowe. “This unnatural union of phalansteries and transcendentalists”. Journal of the History of Ideas, 20(4):495–502, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707888.

Crowe:1966:EPH

[Cro66] Charles Crowe. The emergence of progressive history. Journal of the History of Ideas, 27(1):109–124, January/March 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708311.

Cross:1971:YMM

[Cro71] Truman B. Cross. Young Marx, Marxism: Viktor Chernov’s use of the theses on Feuerbach. Journal of the History of Ideas, 32(4): 600–606, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708981.

Crocker:1985:IEP

[Cro85] Lester G. Crocker. Interpreting the Enlightenment: a political approach. Journal of the History of Ideas, 46(2):211–230, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709635.

Crowe:2005:DPR

[Cro05] Benjamin D. Crowe. Dilthey’s philosophy of religion in the “Critique of Historical Reason”: 1880–1910. Journal of the History of Ideas, 66(2):265–283, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/187186.

Crowe:2009:TGH

[Cro09] Benjamin D. Crowe. “Theismus des Gefühls”: Heydenreich, Fichte, and the transcendental philosophy of religion. Journal of the History of Ideas, 70(4):569–592, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/361798.

Crowe:2010:FVH

[Cro10] Benjamin Crowe. Faith and value: Heinrich Rickert’s theory of religion. Journal of the History of Ideas, 71(4):617–636, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/398300.

Cruttwell:1951:PPS

[Cru51] Patrick Cruttwell. Physiology and psychology in Shakespeare’s age. Journal of the History of Ideas, 12(1):75–89, January 1951. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707538.

Cuffel:1966:CGC

[Cuf66] Victoria Cuffel. The Classical Greek concept of slavery. Journal of the History of Ideas, 27(3):323–342, July/September 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708589.

Culotta:1970:TSH

[Cul70] M. C. Culotta. The temple, the synagogue, and Hebrew precedent. Journal of the History of Ideas, 31(2):273–276, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708549.

Culler:1988:BRW

[Cul88] Jonathan Culler. Book review: Wellek’s Modern Criticism. Journal of the History of Ideas, 49(2):347–351, April/June 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709505.

Cumming:1964:MHH

[Cum64] Robert D. Cumming. Mill’s history of his ideas. Journal of the History of Ideas, 25(2):235–256, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708014.

Cumming:1973:PCG

[Cum73] Alan Cumming. Pauline Christianity and Greek philosophy: a study of the status of women. Journal of the History of Ideas, 34(4): 517–528, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708885.

Cummings:1978:NTT

[Cum78] Philip W. Cummings. A note on the transmission of the title of Ramus’s Master’s thesis. Journal of the History of Ideas, 39(3):481, July/ September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709390.

Cummings:2010:MRB

[Cum10] Anthony M. Cummings. Musical references in Brucioli’s dialogi and their classical and medieval antecedents. Journal of the History of Ideas, 71(2):169–190, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/379773.

Cunningham:1967:AMN

[Cun67] Stanley B. Cunningham. Albertus Magnus on natural law. Journal of the History of Ideas, 28(4):479–502, October/December 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708525.

Curti:1942:AST

[Cur42] Merle Curti. The American scholar in three wars. Journal of the History of Ideas, 3(3):241–264, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707304.

Curti:1961:JAH

[Cur61] Merle Curti. Jane Addams on human nature. Journal of the History of Ideas, 22(2):240–253, April/June 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707835.

Curti:1966:AED

[Cur66] Merle Curti. The American exploration of dreams and dreamers. Journal of the History of Ideas, 27(3):391–416, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708593.

Curtis-Wendlandt:2013:SVW

[CW13] Lisa Curtis-Wendlandt. Staging virtue: Women, death, and liberty in Elise Reimarus’s Cato. Journal of the History of Ideas, 74(1):69– 92, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
495242.

Cywar:1969:JDT

[Cyw69] Alan Cywar. John Dewey: Toward domestic reconstruction, 1915– 1920. Journal of the History of Ideas, 30(3):385–400, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708564. See erra-

tum [Ano70g].

Dacome:2004:NMC

[Dac04] Lucia Dacome. Noting the mind: Commonplace books and the pursuit of the self in Eighteenth-Century Britain. Journal of the History of Ideas, 65(4):603–625, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/184066.

Dahlkvist:2015:EGU

[Dah15] Tobias Dahlkvist. The epileptic genius: The use of Dostoevsky as example in the medical debate over the pathology of genius. Journal of the History of Ideas, 76(4):587–608, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/597348.

Dales:1977:MVH

[Dal77] Richard C. Dales. A medieval view of human dignity. Journal of the History of Ideas, 38(4):557–572, October/December 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708687.

Dales:1980:AHM

[Dal80] Richard C. Dales. The de-animation of the heavens in the Middle Ages. Journal of the History of Ideas, 41(4):531–550, October/ December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709273.

Dales:1984:HHI

[Dal84] Richard C. Dales. Henry of Harclay on the infinite. Journal of the History of Ideas, 45(2):295–301, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709293.

Dales:1988:TET

[Dal88] Richard C. Dales. Time and eternity in the Thirteenth Century. Journal of the History of Ideas, 49(1):27–45, January/March 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709702.

Damrosch:1979:HRT

[Dam79] Leopold Damrosch, Jr. Hobbes as reformation theologian: Implications of the free-will controversy. Journal of the History of Ideas, 40 (3):339–352, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709241.

Dannenfeldt:1952:RPC

[Dan52] Karl H. Dannenfeldt. The Renaissance and the pre-classical civilizations. Journal of the History of Ideas, 13(4):435–449, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707711.

Daniels:1960:FWM

[Dan60] Robert Vincent Daniels. Fate and will in the Marxian philosophy of history. Journal of the History of Ideas, 21(4):538–552, October/ December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708101.

Daniel:1981:SCS

[Dan81] Stephen H. Daniel. Seventeenth-Century scholastic treatments of time. Journal of the History of Ideas, 42(4):587–606, October/ December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709120.

Daniel:1986:MRM

[Dan86] Stephen H. Daniel. Myth and rationality in Mandeville. Journal of the History of Ideas, 47(4):595–609, October/December 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709720.

Dauben:1977:GCP

[Dau77] Joseph W. Dauben. Georg Cantor and Pope Leo XIII: Mathematics, theology, and the infinite. Journal of the History of Ideas, 38 (1):85–108, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708842.

Davis:1943:MVE

[Dav43] Joe Lee Davis. Mystical versus enthusiastic sensibility. Journal of the History of Ideas, 4(3):301–319, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707256.

David:1951:CFM

[Dav51] Hans T. David. The cultural functions of music. Journal of the History of Ideas, 12(3):423–439, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707753.

Davis:1960:MP

[Dav60a] John W. Davis. The Molyneux problem. Journal of the History of Ideas, 21(3):392–408, July/September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708144.

Davis:1960:SCF

[Dav60b] Natalie Zemon Davis. Sixteenth-Century French arithmetics on the business life. Journal of the History of Ideas, 21(1):18–48, January/ March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707997.

Davidson:1963:LE

[Dav63] Edward H. Davidson. From Locke to Edwards. Journal of the History of Ideas, 24(3):355–372, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708212.

Davies:1966:CDS

[Dav66] Gordon L. Davies. The concept of denudation in Seventeenth-Century England. Journal of the History of Ideas, 27(2):278–284, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708644.

Davies:1969:PCB

[Dav69] Evan Davies. Psychological characteristics of beatle mania. Journal of the History of Ideas, 30(2):273–280, April/June 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708439.

Davis:1983:CCI

[Dav83] Walter W. Davis. China, the Confucian ideal, and the European age of Enlightenment. Journal of the History of Ideas, 44(4):523–548, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709213.

Davis:1994:AWR

[Dav94] Edward B. Davis. The anonymous works of Robert Boyle and the reasons why a Protestant should not turn papist (1687). Journal of the History of Ideas, 55(4):611–629, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709925.

Davis:1995:PPA

[Dav95] James A. Davis. Philosophical positivism and American atonal music theory. Journal of the History of Ideas, 56(3):501–522, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710038.

Davies:2010:TWC

[Dav10] Katherine Jane Davies. A “Third way” Catholic intellectual: Charles Du Bos, tragedy, and ethics in interwar Paris. Journal of the History of Ideas, 71(4):637–659, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/398301.

Davies:2013:RRS

[Dav13] S. F. Davies. The reception of Reginald Scot’s discovery of witchcraft: Witchcraft, magic, and radical religion. Journal of the History of Ideas, 74(3):381–401, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/511984.

Dawson:1993:BTT

[Daw93] Hugh J. Dawson. E. B. Tylor’s theory of survivals and Veblen’s social criticism. Journal of the History of Ideas, 54(3):489–504, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710025.

Day:2008:GSS

[Day08] Matthew Day. Godless savages and superstitious dogs: Charles Darwin, Imperial ethnography, and the problem of human uniqueness. Journal of the History of Ideas, 69(1):49–70, January 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/231008.

Duncan:1977:TRC

[DB77] Elmer H. Duncan and Robert M. Baird. Thomas Reid’s criticisms of Adam Smith’s theory of the moral sentiments. Journal of the History of Ideas, 38(3):509–522, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708680.

deBolla:2020:ILD

[dBJN^{+}20] Peter de Bolla, Ewan Jones, Paul Nulty, Gabriel Recchia, and John Regan. The idea of liberty, 1600–1800: a distributional concept analysis. Journal of the History of Ideas, 81(3):381–406, July 2020. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/761035.

deCeglia:2011:TBI

[dC11] Francesco Paolo de Ceglia. “It’s not true, but I believe it”: Discussions on jettatura in Naples between the end of the Eighteenth and beginning of the Nineteenth Centuries. Journal of the History of Ideas, 72(1):75–97, January 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/413475.

DeGre:1941:SKP

[De 41] Gerard De Gré. The sociology of knowledge and the problem of truth. Journal of the History of Ideas, 2(1):110–115, January 1941. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707284.

DeLevie:1953:PCO

[De 53] Dagobert De Levie. Patriotism and clerical office: Germany 1761– 1773. Journal of the History of Ideas, 14(4):622–627, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707706.

DeSeade:1979:SHH

[De 79] Esperanza Durán De Seade. State and history in Hegel’s concept of people. Journal of the History of Ideas, 40(3):369–384, July/ September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709243.

DeVivo:2003:HJV

[De 03] Filippo De Vivo. Historical justifications of Venetian power in the Adriatic. Journal of the History of Ideas, 64(2):159–176, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/43516.

DeWaal:2006:HIM

[De 06] Cornelis De Waal. Having an idea of matter: a Peircean refutation of Berkeleyan immaterialism. Journal of the History of Ideas, 67(2): 291–313, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
198083.

DeBont:2015:PPA

[De 15] Raf De Bont. “Primitives” and protected areas: International conservation and the “Naturalization” of indigenous people, ca. 1910– 1975. Journal of the History of Ideas, 76(2):215–236, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/580296.

DElia:1969:DBR

[D’E69] Donald J. D’Elia. Dr. Benjamin rush and the Negro. Journal of the History of Ideas, 30(3):413–422, July/September 1969. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708566.

DElia:2007:SPC

[D’E07] Anthony F. D’Elia. Stefano Porcari’s conspiracy against Pope Nicholas V in 1453 and republican culture in Papal Rome. Journal of the History of Ideas, 68(2):207–231, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/215054.

Deane:1972:LAE

[Dea72] Seamus F. Deane. Lord Acton and Edmund Burke. Journal of the History of Ideas, 33(2):325–335, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708878.

Dealy:1974:TMD

[Dea74] Glen Caudill Dealy. The tradition of monistic democracy in Latin America. Journal of the History of Ideas, 35(4):625–646, October/ December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709090.

DeArmey:1987:TDA

[DeA87] Michael H. DeArmey. Thomas Davidson’s apeirotheism and its influence on William James and John Dewey. Journal of the History of Ideas, 48(4):691–707, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709694. See errata [Ano88e].

Degen:2012:CFD

[Deg12] Andreas Degen. Concepts of fascination, from Democritus to Kant. Journal of the History of Ideas, 73(3):371–393, July 2012. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/481416.

Dege:2023:MML

[Deg23] Carmen Lea Dege. Myth, modernity, and the legacy of the axial age: Taylor, Habermas, Assmann, and Jaspers. Journal of the History of Ideas, 84(4):743–773, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/909537.

Deininger:1954:SHF

[Dei54] Whitaker T. Deininger. The skepticism and historical faith of Charles A. Beard. Journal of the History of Ideas, 15(4):573–588, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707676.

DelVecchio:1962:BR

[Del62] Giorgio Del Vecchio. Burlamaqui and Rousseau. Journal of the History of Ideas, 23(3):420–423, July/September 1962. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708077.

DellaTerza:1980:BRF

[Del80] Dante Della Terza. Book review: On Francesco de Sanctis (1817– 1883) and Realism. Journal of the History of Ideas, 41(2):335–341, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709466.

DelCaro:1989:DCN

[Del89] Adrian Del Caro. Dionysian classicism, or Nietzsche’s appropriation of an aesthetic norm. Journal of the History of Ideas, 50(4): 589–605, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709799.

Delph:1996:VGA

[Del96] Ronald K. Delph. Valla Grammaticus, Agostino Steuco, and the donation of Constantine. Journal of the History of Ideas, 57(1):55–77, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14898.

Delehanty:2007:MAM

[Del07] Ann T. Delehanty. Mapping the aesthetic mind: John Dennis and Nicolas Boileau. Journal of the History of Ideas, 68(2):233–253, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/215053.

DelSoldato:2017:SPS

[Del17] Eva Del Soldato. Saving the philosopher’s soul: The De pietate Aristotelis by Fortunio Liceti. Journal of the History of Ideas, 78(4):531– 547, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
676145.

Demos:1958:NHE

[Dem58] Raphael Demos. The Neo-Hellenic Enlightenment (1750–1821). Journal of the History of Ideas, 19(4):523–541, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707921.

Demaitre:1972:TCA

[Dem72] Ann Demaitre. The theater of cruelty and alchemy: Artaud and le grand oeuvre. Journal of the History of Ideas, 33(2):237–250, April/ June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708871.

DeMolen:1974:RMP

[DeM74] Richard L. DeMolen. Richard Mulcaster and the profession of teaching in Sixteenth-Century England. Journal of the History of Ideas, 35 (1):121–129, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708747.

Dendle:1999:HDP

[Den99] Peter Dendle. Hume’s dialogues and Paradise Lost. Journal of the History of Ideas, 60(2):257–276, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15025.

Derman:2010:SFM

[Der10] Joshua Derman. Skepticism and faith: Max Weber’s anti-utopianism in the eyes of his contemporaries. Journal of the History of Ideas, 71(3):481–503, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/386154.

DesJardins:1967:TOH

[Des67] Gregory DesJardins. Terms of De Officiis in Hume and Kant. Journal of the History of Ideas, 28(2):237–242, April/June 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708418.

Deutsch:1945:ASI

[Deu45] Karl W. Deutsch. Anti-Semitic ideas in the Middle Ages: International civilizations in expansion and conflict. Journal of the History of Ideas, 6(2):239–251, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707366.

Devine:1970:SBR

[Dev70] Francis Edward Devine. Stoicism on the best regime. Journal of the History of Ideas, 31(3):323–336, July/September 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708509.

DeVun:2008:JHS

[DeV08] Leah DeVun. The Jesus hermaphrodite: Science and sex difference in premodern Europe. Journal of the History of Ideas, 69(2):193– 218, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
235868.

Dewey:1941:JMA

[Dew41] John Dewey. James Marsh and American philosophy. Journal of the History of Ideas, 2(2):131–150, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707110.

Dewey:1959:LJD

[Dew59] John Dewey. Letters of John Dewey to Robert V. Daniels, 1946–1950. Journal of the History of Ideas, 20(4):569–576, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707892.

Dewulf:2018:RHC

[Dew18] Fons Dewulf. Revisiting Hempel’s 1942 contribution to the philosophy of history. Journal of the History of Ideas, 79(3):385–406, July 2018.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/703664.

deGrazia:1980:SLS

[dG80] Margreta de Grazia. The secularization of language in the Seventeenth Century. Journal of the History of Ideas, 41(2):319–329, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709464.

deGrazia:1984:CAW

[dG84] Sebastian de Grazia. Crossings to another world: Machiavelli and others. Journal of the History of Ideas, 45(1):145–151, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709336.

deHuszar:1945:NTD

[dH45] George de Huszar. Nietzsche’s theory of decadence and the transvaluation of all values. Journal of the History of Ideas, 6(3):259–272, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707290.

DeMas:1971:NMN

[DH71] Enrico De Mas and J. K. Houck. On the new method of a new science: a study of Giambattista Vico. Journal of the History of Ideas, 32(1):85–94, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708326.

Diamond:1982:SFH

[Dia82] Sigmund Diamond. Sigmund Freud, his Jewishness, and scientific method: The seen and the unseen as evidence. Journal of the History of Ideas, 43(4):613–634, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709346.

Dick:1980:OEL

[Dic80] Steven J. Dick. The origins of the extraterrestrial life debate and its relation to the scientific revolution. Journal of the History of Ideas, 41(1):3–27, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709100.

Dieckmann:1941:DCG

[Die41] Herbert Dieckmann. Diderot’s conception of genius. Journal of the History of Ideas, 2(2):151–182, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707111.

Dieckmann:1961:BRP

[Die61] Herbert Dieckmann. Book review: Pierre Bayle: ‘Philosopher of Rotterdam’. Journal of the History of Ideas, 22(1):131–136, January/ March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707881.

Diggins:2006:ALC

[Dig06] John P. Diggins. Arthur O. Lovejoy and the challenge of intellectual history. Journal of the History of Ideas, 67(1):181–208, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/194642.

Dillon:1974:CCC

[Dil74] George L. Dillon. Complexity and change of character in neo-classical criticism. Journal of the History of Ideas, 35(1):51–61, January/ March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708742.

Dinwiddy:1975:BTP

[Din75] J. R. Dinwiddy. Bentham’s transition to political radicalism, 1809–10. Journal of the History of Ideas, 36(4):683–700, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708996.

Devereux:1969:OLF

[DJ69] James A. Devereux and S. J. The object of love in Ficino’s philosophy. Journal of the History of Ideas, 30(2):161–170, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708430.

deMauriac:1949:AGP

[dM49] Henry M. de Mauriac. Alexander the great and the politics of “Homonoia”. Journal of the History of Ideas, 10(1):104–114, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707202.

Duffy:1988:NAT

[DM88] Michael F. Duffy and Willard Mittelman. Nietzsche’s attitudes toward the Jews. Journal of the History of Ideas, 49(2):301–317, April/ June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709502.

Doan:2017:MSW

[Doa17] Laura Doan. Marie Stopes’s wonderful rhythm charts: Normalizing the natural. Journal of the History of Ideas, 78(4):595–620, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/676148.

Dodds:1968:BRI

[Dod68] E. R. Dodds. Book review: The Idea of Progress in Classical Antiquity by Ludwig Edelstein. Journal of the History of Ideas, 29 (3):453–457, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708455.

Dodsworth:2008:IPE

[Dod08] F. M. Dodsworth. The idea of police in Eighteenth-Century England: Discipline, reformation, superintendence, c. 1780–1800. Journal of the History of Ideas, 69(4):583–604, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/251816.

Doering:2022:LTF

[Doe22] Jonathan Doering. The linguistic terror in France according to Jean Paulhan and Jean-Paul Sartre. Journal of the History of Ideas, 83(4):555–578, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865892.

Doja:2005:AHA

[Doj05] Albert Doja. The advent of heroic anthropology in the history of ideas. Journal of the History of Ideas, 66(4):633–650, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/191472.

Dombroski:1983:TMI

[Dom83] Robert S. Dombroski. Timpanaro’s materialism: an introduction. Journal of the History of Ideas, 44(2):311–326, April/June 1983. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709142.

Dombrowski:1984:VAM

[Dom84] Daniel A. Dombrowski. Vegetarianism and the argument from marginal cases in porphyry. Journal of the History of Ideas, 45 (1):141–143, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709335.

Dombrowski:1988:SAA

[Dom88] Daniel A. Dombrowski. St. Augustine, abortion, and libido crudelis. Journal of the History of Ideas, 49(1):151–156, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709708.

Doney:1955:CC

[Don55] Willis Doney. The Cartesian circle. Journal of the History of Ideas, 16(3):324–338, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707635.

Donnellan:1982:FNP

[Don82] Brendan Donnellan. Friedrich Nietzsche and Paul Rée: Cooperation and conflict. Journal of the History of Ideas, 43(4):595–612, October/ December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709345.

Dooley:1990:LCS

[Doo90] Brendan Dooley. From literary criticism to systems theory in early modern journalism history. Journal of the History of Ideas, 51(3): 461–486, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709625.

Dooley:1999:VFT

[Doo99] Brendan Maurice Dooley. Veritas filia temporis: Experience and belief in early modern culture. Journal of the History of Ideas, 60(3):487– 504, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15035.

Doody:2009:PNH

[Doo09] Aude Doody. Pliny’s natural history: Enkuklios Paideia and the ancient encyclopedia. Journal of the History of Ideas, 70(1):1–21, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
258192.

Doran:1940:ECS

[Dor40] Madeleine Doran. On Elizabethan “credulity”: With some questions concerning the use of the marvelous in literature. Journal of the History of Ideas, 1(2):151–176, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707331.

Doty:1986:CFW

[Dot86] Ralph Doty. Carneades, a forerunner of William James’s pragmatism. Journal of the History of Ideas, 47(1):133–138, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709599.

Douglass:2011:RDB

[Dou11] Robin Douglass. Rousseau’s debt to Burlamaqui: The ideal of nature and the nature of things. Journal of the History of Ideas, 72(2):209– 230, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
428852.

Downey:1940:RAI

[Dow40] Glanville Downey. Representations of abstract ideas in the Antioch mosaics. Journal of the History of Ideas, 1(1):112–113, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707013.

Drake:1970:RME

[Dra70] Stillman Drake. Renaissance music and experimental science. Journal of the History of Ideas, 31(4):483–500, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708256.

Drake:1975:ITR

[Dra75] Stillman Drake. Impetus theory reappraised. Journal of the History of Ideas, 36(1):27–46, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709009.

Drake:1977:GCP

[Dra77a] Stillman Drake. Galileo and the career of philosophy. Journal of the History of Ideas, 38(1):19–32, January/March 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708838.

Drake:1977:NPF

[Dra77b] Stillman Drake. Note on Professor Franklin’s paper. Journal of the History of Ideas, 38(2):315–316, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708915. See [Fra77].

Drake:1978:RER

[Dra78] Stillman Drake. Reply to Edward Rosen’s note. Journal of the History of Ideas, 39(1):148, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709079. See [Ros78].

Drake:1990:SAP

[Dra90] Richard Drake. Sibilla Aleramo and the peasants of the Agro Romano: a writer’s dilemma. Journal of the History of Ideas, 51(2):255–272, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709515.

Drescher:1964:TTD

[Dre64] Seymour Drescher. Tocqueville’s two democraties. Journal of the History of Ideas, 25(2):201–216, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708012.

Drochon:2019:RAM

[Dro19] Hugo Drochon. Raymond Aron’s “Machiavellian” liberalism. Journal of the History of Ideas, 80(4):621–642, October 2019. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/735384.

dosSantos:2023:HPF

[dS23] Pedro Cristovão dos Santos. Historians and programmers in the 1970s: Formal languages, the writing of history, and ideas of science. Journal of the History of Ideas, 84(1):157–177, January 2023. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/879544.

Ducheyne:2006:GIN

[Duc06] Steffen Ducheyne. Galileo’s interventionist notion of “Cause”. Journal of the History of Ideas, 67(3):443–464, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/202875.

Dudden:1957:MAM

[Dud57] Arthur P. Dudden. Men against monopoly: The prelude to trustbusting. Journal of the History of Ideas, 18(4):587–593, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707570.

Dudden:1961:NA

[Dud61] Arthur P. Dudden. Nostalgia and the American. Journal of the History of Ideas, 22(4):515–530, October/December 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708028.

Duffin:1980:ALE

[Duf80] Kathleen E. Duffin. Arthur O. Lovejoy and the emergence of novelty. Journal of the History of Ideas, 41(2):267–281, April/June 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709460.

Duffy:1995:MSR

[Duf95] Michael H. Duffy. Michelangelo and the sublime in romantic art criticism. Journal of the History of Ideas, 56(2):217–238, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709836.

Duguid:2015:AIP

[Dug15] Paul Duguid. The ageing of information: From particular to particulate. Journal of the History of Ideas, 76(3):347–368, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/587269.

Duhamel:1949:FRE

[Duh49] P. Albert Duhamel. The function of rhetoric as effective expression. Journal of the History of Ideas, 10(3):344–356, June 1949. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707041.

Duhamel:1953:OLJ

[Duh53] P. Albert Duhamel. The Oxford lectures of John Colet: an essay in defining the English Renaissance. Journal of the History of Ideas, 14(4):493–510, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707698.

Dumoulin:1981:BNC

[Dum81] Heinrich Dumoulin. Buddhism and Nineteenth-Century German philosophy. Journal of the History of Ideas, 42(3):457–470, July/ September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709187.

Dunn:1964:AGC

[Dun64] John Dunn. Authorship of Gregory’s critique of Hume. Journal of the History of Ideas, 25(1):128–129, January/March 1964. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708091.

Duncan:1969:PWE

[Dun69] Joseph E. Duncan. Paradise as the whole Earth. Journal of the History of Ideas, 30(2):171–186, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708431.

Dunning:2021:LAJ

[Dun21a] David E. Dunning. The logician in the archive: John Venn’s diagrams and Victorian historical thinking. Journal of the History of Ideas, 82(4):593–614, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/837673.

Dunstan:2021:CRC

[Dun21b] Sarah C. Dunstan. The capital of race capitals: Toward a connective cartography of black internationalisms. Journal of the History of Ideas, 82(4):637–660, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/837675.

Dupre:1974:BRR

[Dup74] Louis Dupré. Book review: Recent Literature on Marx and Marxism. Journal of the History of Ideas, 35(4):703–714, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709096.

Dupre:1982:BRS

[Dup82] Louis Dupré. Book review: Some Recent Philosophical Discussions of Religion. Journal of the History of Ideas, 43(3):505–518, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709436.

Durand:1943:TIF

[Dur43] Dana B. Durand. Tradition and innovation in Fifteenth Century Italy: “Il Primato Dell’ Italia” in the field of science. Journal of the History of Ideas, 4(1):1–20, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707234.

Dussinger:1981:LSL

[Dus81] John A. Dussinger. “The lovely system of Lord Shaftesbury”: an answer to Locke in the aftermath of 1688? Journal of the History of Ideas, 42(1):151–158, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709422.

Dwan:2011:EBE

[Dwa11] David Dwan. Edmund Burke and the emotions. Journal of the History of Ideas, 72(4):571–593, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/453549.

Dykhuizen:1952:ECL

[Dyk52] George Dykhuizen. An early chapter in the life of John Dewey. Journal of the History of Ideas, 13(4):563–572, October 1952. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707718.

Dykhuizen:1959:JDV

[Dyk59] George Dykhuizen. John Dewey: The Vermont years. Journal of the History of Ideas, 20(4):515–544, October/December 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707890.

Dykhuizen:1961:JDJ

[Dyk61] George Dykhuizen. John Dewey at Johns Hopkins (1882–1884). Journal of the History of Ideas, 22(1):103–116, January/March 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707877.

Dykhuizen:1962:JDU

[Dyk62] George Dykhuizen. John Dewey and the University of Michigan. Journal of the History of Ideas, 23(4):513–544, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708172.

Dyke:1971:C

[Dyk71] C. Dyke. Comments. Journal of the History of Ideas, 32(3): 437–440, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2708358. See [Bec71, Axi71].

Dymond:2021:HCF

[Dym21] Jeffrey Dymond. Human character and the formation of the state: Reconsidering Machiavelli and Polybius 6. Journal of the History of Ideas, 82(1):29–50, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/781611.

Dynes:1981:BRW

[Dyn81] Wayne Dynes. Book review: The Work of Meyer Schapiro: Distinction and Distance. Journal of the History of Ideas, 42(1):163–172, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709424.

Easton:1958:EED

[Eas58] Loyd D. Easton. Empiricism and ethics in dietzgen. Journal of the History of Ideas, 19(1):77–90, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707953.

Easton:1962:HNC

[Eas62] Loyd D. Easton. Hegelianism in Nineteenth-Century Ohio. Journal of the History of Ideas, 23(3):355–378, July/September 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708072.

Eastwood:1967:GQL

[Eas67] Bruce S. Eastwood. Grosseteste’s “Quantitative” law of refraction: a chapter in the history of non-experimental science. Journal of the History of Ideas, 28(3):403–414, July/September 1967. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708626.

Eaton:1980:MR

[Eat80] Henry Eaton. Marx and the Russians. Journal of the History of Ideas, 41(1):89–112, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709104.

Ebel:1969:TCN

[Ebe69] Julia G. Ebel. Translation and cultural nationalism in the reign of Elizabeth. Journal of the History of Ideas, 30(4):593–602, October/ December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708613.

Eberwein:1975:SJG

[Ebe75] Robert Eberwein. Samuel Johnson, George Cheyne, and the “Cone of Being”. Journal of the History of Ideas, 36(1):153–158, January/ March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709017.

Eccleshall:1976:RHS

[Ecc76] Robert Eccleshall. Richard Hooker’s synthesis and the problem of allegiance. Journal of the History of Ideas, 37(1):111–124, January/ March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708712.

Echeverria:1972:PRI

[Ech72] Durand Echeverria. The pre-revolutionary influence of Rousseau’s contrat social. Journal of the History of Ideas, 33(4):543–560, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708856.

Eckstein:1944:RST

[Eck44] Walter Eckstein. Rousseau and Spinoza: Their political theories and their conception of ethical freedom. Journal of the History of Ideas, 5(3):259–291, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707351.

Eckstein:1945:FNJ

[Eck45] Walter Eckstein. Friedrich Nietzsche in the judgment of posterity. Journal of the History of Ideas, 6(3):310–324, June 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707295.

Eckstein:1965:BRR

[Eck65] Walter Eckstein. Book review: Rousseau’s Theory of Liberty. Journal of the History of Ideas, 26(2):291–294, April/June 1965. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708236.

Edel:1946:SLM

[Ede46] Abraham Edel. A symposium: Levels of meaning and the history of ideas. Journal of the History of Ideas, 7(3):355–360, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707406.

Edelstein:1949:FMP

[Ede49] Ludwig Edelstein. The function of the myth in Plato’s philosophy. Journal of the History of Ideas, 10(4):463–481, October 1949. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707185.

Edelstein:1952:BRR

[Ede52] Ludwig Edelstein. Book review: Recent Trends in the Interpretation of Ancient Science. Journal of the History of Ideas, 13(4):573–604, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707719.

Edelstein:1963:ML

[Ede63] Ludwig Edelstein. In memory of A. O. Lovejoy. Journal of the History of Ideas, 24(4):451–456, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707977.

Eden:1998:BFA

[Ede98] Kathy Eden. “Between friends all is common”: The Erasmian adage and tradition. Journal of the History of Ideas, 59(3):405–419, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14989.

Edelstein:2018:CHR

[Ede18] Dan Edelstein. Christian human rights in the French Revolution. Journal of the History of Ideas, 79(3):411–426, July 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/703666.

Edelstein:2022:RPT

[Ede22] Dan Edelstein. A “Revolution” in political thought: Translations of Polybius Book 6 and the conceptual history of revolution. Journal of the History of Ideas, 83(1):17–40, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/847819.

Edwards:1970:BRR

[EE70] Paul Edwards and Paul Edwards. Book review: Reply to a Review. Journal of the History of Ideas, 31(1):143–145, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708377.

Efron:1997:JTS

[Efr97] Noah J. Efron. Jewish thought and scientific discovery in early modern Europe. Journal of the History of Ideas, 58(4):719–732, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14964.

Epstein:1984:DNR

[EG84] Julia L. Epstein and Mark L. Greenberg. Decomposing Newton’s rainbow. Journal of the History of Ideas, 45(1):115–140, January/ March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709334.

Egerton:1966:LPT

[Ege66] Frank N. Egerton III. The longevity of the patriarchs: a topic in the history of demography. Journal of the History of Ideas, 27(4): 575–584, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708341.

Egginton:1999:DHC

[Egg99] William Egginton. On Dante, hyperspheres, and the curvature of the Medieval Cosmos. Journal of the History of Ideas, 60(2):195–216, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15026.

Ehrenpreis:1952:SL

[Ehr52] Irvin Ehrenpreis. Swift on liberty. Journal of the History of Ideas, 13(2):131–146, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707607.

Eigen:2009:PHD

[Eig09] Edward A. Eigen. The plagiarism of the heathens detected: John Wood, the elder (1704–1754) on the translation of architecture and empire. Journal of the History of Ideas, 70(3):375–397, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/270103.

Eisele:1957:SPC

[Eis57] Carolyn Eisele. The scientist–philosopher C. S. Peirce at the Smithsonian. Journal of the History of Ideas, 18(4):537–547, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707564.

Ekirch:1950:FGA

[Eki50] Arthur A. Ekirch, Jr. Frederick Grimké: Advocate of free institutions. Journal of the History of Ideas, 11(1):75–92, January 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707453.

Elazar:2015:LCB

[Ela15] Yiftah Elazar. Liberty as a caricature: Bentham’s antidote to republicanism. Journal of the History of Ideas, 76(3):417–439, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/587272.

Elden:2024:FDA

[Eld24] Stuart Elden. Foucault and Dumézil on antiquity. Journal of the History of Ideas, 85(3):571–600, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/933859.

Elias:1967:BRB

[Eli67] Julius A. Elias. Book review: Boas on the Cult of Childhood. Journal of the History of Ideas, 28(3):451–454, July/September 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708631.

Elkins:1992:RP

[Elk92] James Elkins. Renaissance perspectives. Journal of the History of Ideas, 53(2):209–230, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709871.

Ellegaard:1957:DTN

[Ell57] Alvar Ellegård. The Darwinian theory and Nineteenth-Century philosophies of science. Journal of the History of Ideas, 18(3):362– 393, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707798.

Ellegard:1958:POP

[Ell58] Alvar Ellegard. Public opinion and the press: Reactions to Darwinism. Journal of the History of Ideas, 19(3):379–387, June 1958. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708042.

Ellis:1962:NCM

[Ell62] Brian D. Ellis. Newton’s concept of motive force. Journal of the History of Ideas, 23(2):273–278, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708161.

Elliott:1967:QCS

[Ell67] Charles F. Elliott. Quis custodiet sacra? Problems of Marxist revisionism. Journal of the History of Ideas, 28(1):71–86, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708481.

Ellenzweig:2003:LGR

[Ell03] Sarah Ellenzweig. The love of God and the radical Enlightenment: Mary Astell’s brush with Spinoza. Journal of the History of Ideas, 64(3):379–397, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/47659.

Elman:1983:NB

[Elm83] Benjamin A. Elman. Nietzsche and Buddhism. Journal of the History of Ideas, 44(4):671–686, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709223. See errata [Ano84e].

Elton:1997:ANE

[Elt97] William R. Elton. Aristotle’s Nicomachean ethics and Shakespeare’s Troilus and Cressida. Journal of the History of Ideas, 58(2):331–337, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14939.

Elukin:1992:JBH

[Elu92] Jonathan M. Elukin. Jacques Basnage and the history of the Jews: Anti-Catholic polemic and historical allegory in the Republic of Letters. Journal of the History of Ideas, 53(4):603–630, October/ December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709940.

Elukin:1998:NEH

[Elu98] Jonathan M. Elukin. A new essenism: Heinrich Graetz and mysticism. Journal of the History of Ideas, 59(1):135–148, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14970.

Elukin:2002:MRF

[Elu02] Jonathan M. Elukin. Maimonides and the rise and fall of the Sabians: Explaining Mosaic laws and the limits of scholarship. Journal of the History of Ideas, 63(4):619–637, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/39293.

Emerson:1967:PGH

[Eme67] Roger Emerson. Peter Gay and the heavenly city. Journal of the History of Ideas, 28(3):383–402, July/September 1967. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708625.

Emery:1984:MCO

[Eme84] Kent Emery, Jr. Mysticism and the coincidence of opposites in Sixteenth- and Seventeenth-Century France. Journal of the History of Ideas, 45(1):3–23, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709328.

Edelstein:2013:QQC

[EMR13] Dan Edelstein, Robert Morrissey, and Glenn Roe. To quote or not to quote: Citation strategies in the Encyclopédie. Journal of the History of Ideas, 74(2):213–236, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/504785.

Engel:1963:ENF

[Eng63] S. Morris Engel. An early Nietzsche fragment on language. Journal of the History of Ideas, 24(2):279–286, April/June 1963. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707850.

Eng:1980:LTR

[Eng80] Erling Eng. Locke’s Tabula Rasa and Freud’s “Mystic Writing Pad”. Journal of the History of Ideas, 41(1):133–140, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709107.

Engell:1995:OHL

[Eng95] James Engell. Obituary: Harry Levin 1912–1994. Journal of the History of Ideas, 56(1):161, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710013.

Engster:1998:MM

[Eng98] Dan Engster. The montaignian moment. Journal of the History of Ideas, 59(4):625–650, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15002.

English:2007:HHD

[Eng07] James F. English. Hazards of the higher debunkery. Journal of the History of Ideas, 68(3):363–368, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/219818.

English:2008:SCS

[Eng08] Adam C. English. “Science cannot stop with science”: Maurice Blondel and the sciences. Journal of the History of Ideas, 69(2):269– 292, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
235869.

Entin:2018:CRC

[Ent18] Gabriel Entin. Catholic Republicanism: The creation of the Spanish American republics during revolution. Journal of the History of Ideas, 79(1):105–123, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/684835.

Erlich:1973:RF

[Erl73] Victor Erlich. Russian formalism. Journal of the History of Ideas, 34 (4):627–638, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708893.

Erlin:2002:RMM

[Erl02] Matt Erlin. Reluctant modernism: Moses Mendelssohn’s philosophy of history. Journal of the History of Ideas, 63(1):83–104, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15142.

Eskildsen:2004:HGL

[Esk04] Kasper Risbjerg Eskildsen. How Germany left the Republic of Letters. Journal of the History of Ideas, 65(3):421–432, July 2004. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/177664.

Embree:1966:BRW

[EST66] Ainslie T. Embree, Conrad M. Schirokauer, and Arthur E. Tiedemann. Book review: Ways of Thinking of Eastern Peoples: India-China-Tibet-Japan by Hajime Nakamura, Philip P. Wiener. Journal of the History of Ideas, 27(1):145–149, January/March 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708315.

Evans:1955:ANT

[Eva55] Melbourne G. Evans. Aristotle, Newton, and the theory of continuous magnitude. Journal of the History of Ideas, 16(4):548–557, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707510.

Evans:1965:MDH

[Eva65] Mary Alice Evans. Mimicry and the Darwinian heritage. Journal of the History of Ideas, 26(2):211–220, April/June 1965. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708228.

Eyck:1957:EFI

[Eyc57] F. Gunther Eyck. English and French influences on German liberalism before 1848. Journal of the History of Ideas, 18(3):313–341, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707796.

Eze:2000:HRH

[Eze00] Emmanuel Chukwudi Eze. Hume, race, and human nature. Journal of the History of Ideas, 61(4):691–698, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15093.

Fackenheim:1975:BRN

[Fac75] Emil L. Fackenheim. Book review: The Nazi Holocaust as a Persisting Trauma for the Non-Jewish Mind. Journal of the History of Ideas, 36(2):369–376, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2708935. See errata [Ano75g].

Facius:2021:TGE

[Fac21] Michael Facius. Terms of government: Early Modern Japanese concepts of rulership and political geography in translation. Journal of the History of Ideas, 82(3):521–537, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/798278.

Fairchild:1941:RRR

[Fai41] Hoxie N. Fairchild. Romanticism and the religious revival in England. Journal of the History of Ideas, 2(3):330–338, June 1941. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707135.

Fakhry:1965:AFR

[Fak65] Majid Fakhry. Al-Farabi and the reconciliation of Plato and Aristotle. Journal of the History of Ideas, 26(4):469–478, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708494.

Falnes:1954:EPS

[Fal54] Oscar J. Falnes. European progress and the “Superior” races: As viewed by a fin-de-siècle liberal, Charles H. Pearson. Journal of the History of Ideas, 15(2):312–321, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707776.

Fallon:1988:AMC

[Fal88] Stephen M. Fallon. “To act or not”: Milton’s conception of divine freedom. Journal of the History of Ideas, 49(3):425–449, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709486.

Falkenstein:1990:WKN

[Fal90] Lorne Falkenstein. Was Kant a nativist? Journal of the History of Ideas, 51(4):573–597, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709647.

Falaky:2020:CCM

[Fal20] Fayçal Falaky. The cloche and its critics: Muting the Church’s voice in pre-revolutionary France. Journal of the History of Ideas, 81(2): 239–255, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
753536.

Farr:1987:WHL

[Far87] James Farr. The way of hypotheses: Locke on method. Journal of the History of Ideas, 48(1):51–72, January/March 1987. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709612.

Farrenkopf:1991:TSP

[Far91] John Farrenkopf. The transformation of Spengler’s philosophy of world history. Journal of the History of Ideas, 52(3):463–485, July/ September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710047.

Fasolt:1997:WDY

[Fas97] Constantin Fasolt. William Durant the Younger and conciliar theory. Journal of the History of Ideas, 58(3):385–402, July 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14952.

Fatovic:2005:ACR

[Fat05] Clement Fatovic. The anti-Catholic roots of liberal and republican conceptions of freedom in English political thought. Journal of the History of Ideas, 66(1):37–58, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/185256.

Faurot:1978:RAJ

[Fau78] J. H. Faurot. Reid’s answer to Joseph Priestley. Journal of the History of Ideas, 39(2):285–292, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708780.

Fay:1978:DMO

[Fay78] Margaret A. Fay. Did Marx offer to dedicate Capital to Darwin?: a reassessment of the evidence. Journal of the History of Ideas, 39 (1):133–146, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709077. See correction [Car78].

Feffer:1990:SSC

[Fef90] Andrew Feffer. Sociability and social conflict in George Herbert Mead’s interactionism, 1900–1919. Journal of the History of Ideas, 51(2):233–254, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709514.

Feffer:2005:PDD

[Fef05] Andrew Feffer. The presence of democracy: Deweyan exceptionalism and Communist teachers in the 1930s. Journal of the History of Ideas, 66(1):79–97, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/185257.

Feibleman:1944:ROM

[Fei44] James Feibleman. Reid and the origins of modern realism. Journal of the History of Ideas, 5(1):113–120, January 1944. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707105.

Feibleman:1946:TH

[Fei46] James Feibleman. The theory of Hamlet. Journal of the History of Ideas, 7(2):131–150, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707069.

Feinstein:1968:FQC

[Fei68] Blossom Feinstein. The Faerie Queene and cosmogonies of the Near East. Journal of the History of Ideas, 29(4):531–550, October/ December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708292.

Feinstein:1970:WJE

[Fei70] Howard M. Feinstein. William James on the emotions. Journal of the History of Ideas, 31(1):133–142, January/March 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708376.

Feinberg:1977:WHC

[Fei77] Joel Feinberg. Wollaston and his critics. Journal of the History of Ideas, 38(2):345–352, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708918.

Feld:1977:BRR

[Fel77] M. D. Feld. Book review: Revolution and Reaction in Early Modern Europe. Journal of the History of Ideas, 38(1):175–184, January/ March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708849.

Feldman:2004:PCT

[Fel04] Karen S. Feldman. Per canales Troporum: On tropes and performativity in Leibniz’s preface to Nizolius. Journal of the History of Ideas, 65(1):39–51, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/168332.

Fenves:1986:MDT

[Fen86] Peter Fenves. Marx’s doctoral thesis on two Greek atomists and the post-Kantian interpretations. Journal of the History of Ideas, 47 (3):433–452, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709662.

Ferguson:1951:IRS

[Fer51] Wallace K. Ferguson. The interpretation of the Renaissance: Suggestions for a synthesis. Journal of the History of Ideas, 12(4):483–495, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707483.

Ferguson:1954:RTR

[Fer54] Wallace K. Ferguson. Renaissance tendencies in the religious thought of Erasmus. Journal of the History of Ideas, 15(4):499–508, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707673.

Ferguson:1955:RRC

[Fer55] Arthur B. Ferguson. Renaissance realism in the “Commonwealth” literature of Early Tudor England. Journal of the History of Ideas, 16(3): 287–305, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707633.

Ferguson:1958:IIH

[Fer58] Wallace K. Ferguson. The interpretation of Italian humanism: The contribution of Hans Baron. Journal of the History of Ideas, 19(1):14– 25, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707949.

See reply [Bar58a].

Ferguson:1968:BRC

[Fer68] Arthur B. Ferguson. Book review: The Complete Works of St. Thomas More, Volume 4, Utopia by Edward Surtz, J. H. Hexter. Journal of the History of Ideas, 29(2):303–310, April/June 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708583.

Ferguson:1971:HTS

[Fer71] Arthur B. Ferguson. The historical thought of Samuel Daniel: a study in Renaissance ambivalence. Journal of the History of Ideas, 32(2):185–202, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708276.

Fernandez:1973:EJW

[Fer73] José A. Fernández. Erasmus on the just war. Journal of the History of Ideas, 34(2):209–226, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708726.

Feuer:1949:JSM

[Feu49] Lewis S. Feuer. John Stuart Mill and Marxian socialism. Journal of the History of Ideas, 10(2):297–303, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707419.

Feuer:1958:JDR

[Feu58] Lewis S. Feuer. John Dewey’s reading at College. Journal of the History of Ideas, 19(3):415–421, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708046.

Feuer:1959:JDB

[Feu59] Lewis S. Feuer. John Dewey and the Back to the People Movement in American thought. Journal of the History of Ideas, 20(4): 545–568, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707891.

Feuer:1979:FER

[Feu79] Lewis S. Feuer. The friendship of Edwin Ray Lankester and Karl Marx: The last episode in Marx’s intellectual evolution. Journal of the History of Ideas, 40(4):633–648, October/December 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709363.

Fidler:1994:JMG

[Fid94] Geoffrey C. Fidler. On Jean-Marie Guyau, immoraliste. Journal of the History of Ideas, 55(1):75–97, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709954.

Fiess:1955:BRT

[Fie55] Edward Fiess. Book review: Toynbee as Poet. Journal of the History of Ideas, 16(2):275–280, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2707669. See comment [Toy55].

Fiering:1976:ICA

[Fie76] Norman S. Fiering. Irresistible compassion: an aspect of Eighteenth-Century sympathy and humanitarianism. Journal of the History of Ideas, 37(2):195–218, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708821.

Field:1977:NR

[Fie77] Geoffrey G. Field. Nordic racism. Journal of the History of Ideas, 38(3):523–540, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708681.

Fieser:1996:ECB

[Fie96] James Fieser. The Eighteenth-Century British reviews of Hume’s writings. Journal of the History of Ideas, 57(4):645–657, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14920.

Finney:1947:EMS

[Fin47a] Gretchen L. Finney. Ecstasy and music in Seventeenth-Century England. Journal of the History of Ideas, 8(2):153–186, April 1947. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707154.

Finney:1947:OME

[Fin47b] Gretchen L. Finney. “Organical musick” and ecstasy. Journal of the History of Ideas, 8(3):273–292, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707143.

Fink:1972:BCE

[Fin72] Beatrice C. Fink. Benjamin Constant on equality. Journal of the History of Ideas, 33(2):307–314, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708876.

Finnegan:2014:EEC

[Fin14] Diarmid A. Finnegan. Eve and evolution: Christian responses to the first woman question, 1860–1900. Journal of the History of Ideas, 75(2):283–305, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/544562.

Firda:1971:GPH

[Fir71] Richard Arthur Firda. German philosophy of history and literature in the North American review: 1815–1860. Journal of the History of Ideas, 32(1):133–142, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708331.

Fisch:1947:BRP

[Fis47] Max H. Fisch. Book review: Professor Schneider’s History of American Philosophy. Journal of the History of Ideas, 8(4):484–491, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707432.

Fisch:1954:ABG

[Fis54] Max H. Fisch. Alexander Bain and the genealogy of pragmatism. Journal of the History of Ideas, 15(3):413–444, June 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707763.

Fisher:1956:MT

[Fis56] Peter F. Fisher. Milton’s theodicy. Journal of the History of Ideas, 17(1):28–53, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707684.

Fisher:1958:PCJ

[Fis58] Marvin Fisher. The pattern of conservatism in Johnson’s rasselas and Hawthorne’s tales. Journal of the History of Ideas, 19(2):173–196, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707934.

Fisher:1962:ML

[Fis62a] Peter F. Fisher. Milton’s logic. Journal of the History of Ideas, 23(1):37–60, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708056.

Fishman:1962:LHE

[Fis62b] Sterling Fishman. Lassalle on Heraclitus of Ephesus. Journal of the History of Ideas, 23(3):379–391, July/September 1962. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708073.

Fisher:1966:PWD

[Fis66] John Fisher. Plato on writing and doing philosophy. Journal of the History of Ideas, 27(2):163–172, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708636.

Fisher:1970:BRI

[Fis70] John Fisher. Book review: On Introducing the History of Ideas. Journal of the History of Ideas, 31(4):617–619, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708266.

Fisch:1972:PL

[Fis72] Max H. Fisch. Peirce and Leibniz. Journal of the History of Ideas, 33(3):485–496, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709048.

Fischer:1975:JLG

[Fis75] Klaus P. Fischer. John Locke in the German Enlightenment: an interpretation. Journal of the History of Ideas, 36(3):431–446, July/ September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708655.

Fisher:1982:LIR

[Fis82] Arthur L. Fisher. Lactantius’ ideas relating Christian truth and Christian society. Journal of the History of Ideas, 43(3):355–377, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709428.

Fitzmaurice:1997:CRP

[Fit97] Andrew Fitzmaurice. Classical rhetoric and the promotion of the new world. Journal of the History of Ideas, 58(2):221–243, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14942.

Fix:1989:ADE

[Fix89] Andrew Fix. Angels, devils, and evil spirits in Seventeenth-Century thought: Balthasar Bekker and the collegiants. Journal of the History of Ideas, 50(4):527–547, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709796.

Flechtheim:1947:HPP

[Fle47] Ossip K. Flechtheim. Hegel and the problem of punishment. Journal of the History of Ideas, 8(3):293–308, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707144.

Fleming:1959:COS

[Fle59] Donald Fleming. The centenary of the origin of species. Journal of the History of Ideas, 20(3):437–446, June/September 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708122.

Fleck:1963:ALN

[Fle63] George M. Fleck. Atomism in Late Nineteenth-Century physical chemistry. Journal of the History of Ideas, 24(1):106–114, January/ March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707861.

Fleisher:1966:TDM

[Fle66] Martin Fleisher. Trust and deceit in Machiavelli’s comedies. Journal of the History of Ideas, 27(3):365–380, July/September 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708591.

Flournoy:1946:BLT

[Flo46] F. R. Flournoy. British liberal theories of international relations (1848–1898). Journal of the History of Ideas, 7(2):195–217, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707071.

Flower:1949:MRA

[Flo49] Elizabeth Flower. The Mexican revolt against positivism. Journal of the History of Ideas, 10(1):115–129, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707203.

Floridi:1995:DSE

[Flo95] Luciano Floridi. The diffusion of Sextus Empiricus’s works in the Renaissance. Journal of the History of Ideas, 56(1):63–85, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2710007.

Fogel:2010:RBS

[Fog10] Joshua A. Fogel. The recent boom in Shanghai studies. Journal of the History of Ideas, 71(2):313–333, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/379777.

Fogel:2012:ALL

[Fog12] Joshua A. Fogel. 2011 Arthur O. Lovejoy Lecture: The gold seal of 57 CE and the afterlife of an inanimate object. Journal of the History of Ideas, 73(3):351–369, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/481415.

Foland:1958:ILN

[Fol58] Frances M. Foland. The impact of liberalism on nueva espana. Journal of the History of Ideas, 19(2):161–172, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707933.

Foley:2017:MSC

[Fol17] Adam Foley. Miltonic sublimity and the crisis of Wolffianism before Kant. Journal of the History of Ideas, 78(1):51–71, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/648122.

Fontana:1999:LCL

[Fon99] Benedetto Fontana. Love of country and love of God: The political uses of religion in Machiavelli. Journal of the History of Ideas, 60(4):639–658, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15047.

Fontana:2000:LKG

[Fon00] Benedetto Fontana. Logos and kratos: Gramsci and the ancients on hegemony. Journal of the History of Ideas, 61(2):305–326, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15069.

Forbes:1942:DMO

[For42] Elizabeth Livermore Forbes. Of the dignity of man: Oration of Giovanni Pico Della Mirandola, Count of Concordia. Journal of the History of Ideas, 3(3):347–354, June 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707308.

Forrest:1968:JD

[For68] Helen Forrest. Justinian I and the Descensus. Journal of the History of Ideas, 29(4):614–615, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708298.

Formigari:1974:LSL

[For74] Lia Formigari. Language and society in the Late Eighteenth Century. Journal of the History of Ideas, 35(2):275–292, April/June 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708762.

Force:1982:HJP

[For82] James E. Force. Hume and Johnson on prophecy and miracles: Historical context. Journal of the History of Ideas, 43(3):463–475, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709433.

Force:1984:HRS

[For84] James E. Force. Hume and the relation of science to religion among certain members of the Royal Society. Journal of the History of Ideas, 45(4):517–536, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709371.

Force:1989:BRO

[For89] James E. Force. Book review: The Origins of Modern Atheism. Journal of the History of Ideas, 50(1):153–162, January/March 1989. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709791.

Forth:1993:NDR

[For93] Christopher E. Forth. Nietzsche, decadence, and regeneration in France, 1891–95. Journal of the History of Ideas, 54(1):97–117, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709862.

Forsey:2003:DPA

[For03] Jane Forsey. The disenfranchisement of philosophical aesthetics. Journal of the History of Ideas, 64(4):581–597, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/52447.

Force:2005:ISE

[For05] Pierre Force. Innovation as spiritual exercise: Montaigne and Pascal. Journal of the History of Ideas, 66(1):17–35, January 2005. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/185258.

Force:2009:MCE

[For09] Pierre Force. Montaigne and the coherence of eclecticism. Journal of the History of Ideas, 70(4):523–544, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/361800.

Force:2016:EPP

[For16] Pierre Force. The “Exasperating Predecessor”: Pocock on Gibbon and Voltaire. Journal of the History of Ideas, 77(1):129–145, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/610564.

Forss:2023:MAO

[For23] Charlotta Forss. Mapping Atlantis: Olof Rudbeck and the use of maps in early modern scholarship. Journal of the History of Ideas, 84(2): 207–231, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
886180.

FernandezSebastian:2007:IHD

[FR07] Javier Fernández Sebastián and Pierre Rosanvallon. Intellectual history and democracy: an interview with Pierre Rosanvallon. Journal of the History of Ideas, 68(4):703–715, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/223135.

Frank:1942:CCL

[Fra42] Tenney Frank. Changing conceptions of literary and philological research. Journal of the History of Ideas, 3(4):401–414, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707315.

Frankena:1955:HMS

[Fra55] William Frankena. Hutcheson’s moral sense theory. Journal of the History of Ideas, 16(3):356–375, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707637.

Frangsmyr:1975:CWM

[Frä75] Tore Frängsmyr. Christian Wolff’s mathematical method and its impact on the Eighteenth Century. Journal of the History of Ideas, 36 (4):653–668, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708994.

Franklin:1977:SDI

[Fra77] Allan Franklin. Stillman Drake’s “Impetus Theory Reappraised”. Journal of the History of Ideas, 38(2):307–315, April/June 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708914. See note [Dra77b].

Francis:1978:HSM

[Fra78] Mark Francis. Herbert Spencer and the myth of laissez-faire. Journal of the History of Ideas, 39(2):317–328, April/June 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708783.

Franklin:1991:OHD

[Fra91] Julian H. Franklin. Obituary: Herbert A. Deane (1921–1991). Journal of the History of Ideas, 52(3):524, July/September 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710052.

Franco:2003:APM

[Fra03] Abel B. Franco. Avempace, projectile motion, and impetus theory. Journal of the History of Ideas, 64(4):521–546, October 2003. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/52448.

Fraenkel:2013:SMT

[Fra13] Carlos Fraenkel. Spinoza on miracles and the truth of the Bible. Journal of the History of Ideas, 74(4):643–658, October 2013. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/523433.

Frenkel:1964:SCS

[Fre64] F. E. Frenkel. Sex-crime and its socio-historical background. Journal of the History of Ideas, 25(3):333–352, July/September 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707903.

Freeden:1973:JHN

[Fre73] Michael Freeden. J. A. Hobson as a new liberal theorist: Some aspects of his social thought until 1914. Journal of the History of Ideas, 34 (3):421–443, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708962.

Friess:1940:BRD

[Fri40] Horace L. Friess. Book review: Dilthey on German History and Historians. Journal of the History of Ideas, 1(4):506–508, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707128.

Fritz:1954:BSO

[Fri54] Anita D. Fritz. Berkeley’s self — its origin in Malebranche. Journal of the History of Ideas, 15(4):554–572, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707675.

Fries:1973:SNA

[Fri73] Sylvia D. Fries. Staatstheorie and the new American science of politics. Journal of the History of Ideas, 34(3):391–404, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708960.

Frigo:2012:COS

[Fri12] Alberto Frigo. Can one speak of painting if one cannot hold a brush?: Giulio Mancini, medicine, and the birth of the connoisseur. Journal of the History of Ideas, 73(3):417–436, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/481418.

Frohman:1995:NST

[Fro95] Larry Frohman. Neo-stoicism and the transition to modernity in Wilhelm Dilthey’s philosophy of history. Journal of the History of Ideas, 56(2):263–287, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709838.

Frobert:2007:EHF

[Fro07] Ludovic Frobert. Elie Halévy’s first lectures on the history of European socialism. Journal of the History of Ideas, 68(2):329–353, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/215055.

Frye:1954:SYC

[Fry54] Roland Mushat Frye. Swift’s Yahoo and the Christian symbols for sin. Journal of the History of Ideas, 15(2):201–217, April 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707767. See comments [Fry54].

Fernandez-Santamaria:1980:RSS

[FS80] José A. Fernández-Santamaria. Reason of state and statecraft in Spain (1595–1640). Journal of the History of Ideas, 41(3):355–379, July/ September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709399.

Fubini:1996:HTV

[Fub96] Riccardo Fubini. Humanism and truth: Valla writes against the donation of Constantine. Journal of the History of Ideas, 57(1):79–86, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14891.

Fullinwider:1983:SFJ

[Ful83] S. P. Fullinwider. Sigmund Freud, John Hughlings Jackson, and speech. Journal of the History of Ideas, 44(1):151–158, January/ March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709311.

Fulton:1994:ASA

[Ful94] Ann Fulton. Apostles of Sartre: Advocates of early Sartreanism in American philosophy. Journal of the History of Ideas, 55(1):113–127, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709956.

Funkenstein:1994:PWJ

[Fun94] Amos Funkenstein. The polytheism of William James. Journal of the History of Ideas, 55(1):99–111, January 1994. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709955.

Furley:1981:GTI

[Fur81] David J. Furley. The Greek theory of the infinite universe. Journal of the History of Ideas, 42(4):571–585, October/December 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709119.

Gabriel:1955:EIC

[Gab55] Astrik L. Gabriel. The educational ideas of Christine De Pisan. Journal of the History of Ideas, 16(1):3–21, January 1955. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707524.

Gabriel:1956:BRH

[Gab56] Ralph H. Gabriel. Book review: Hartz on American Liberal Tradition. Journal of the History of Ideas, 17(1):136–138, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707693.

Gabbey:1998:MDR

[Gab98] Alan Gabbey. The melon and the dictionary: Reflections on Descartes’s dreams. Journal of the History of Ideas, 59(4):651–668, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15003.

Gaiger:2000:STL

[Gai00] Jason Gaiger. Schiller’s theory of landscape depiction. Journal of the History of Ideas, 61(1):115–132, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15058.

Gale:1974:CM

[Gal74] George Gale. Chew’s monadology. Journal of the History of Ideas, 35(2):339–348, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708768.

Galloway:1994:MSE

[Gal94] Andrew Galloway. The making of a social ethic in Late-Medieval England: From gratitudo to “Kyndenesse”. Journal of the History of Ideas, 55(3):365–383, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709845.

Galisanka:2017:JSF

[Gal17] Andrius Gališanka. Just society as a fair game: John Rawls and game theory in the 1950s. Journal of the History of Ideas, 78(2):299– 308, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
652829.

Gang:1956:HMC

[Gan56] T. M. Gang. Hobbes and the metaphysical conceit — a reply. Journal of the History of Ideas, 17(3):418–421, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707553. See [Wat55].

Garraty:1954:PSR

[Gar54] John A. Garraty. Preserved Smith, Ralph Volney Harlow, and psychology. Journal of the History of Ideas, 15(3):456–465, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707765.

Garfinkle:1955:SRE

[Gar55] Norton Garfinkle. Science and religion in England, 1790–1800: The critical response to the work of Erasmus Darwin. Journal of the History of Ideas, 16(3):376–388, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707638.

Garson:1970:LHR

[Gar70] Judith Garson. Literary history: Russian formalist views, 1916–1928. Journal of the History of Ideas, 31(3):399–412, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708513.

Garrett:1973:JPM

[Gar73] Clarke Garrett. Joseph Priestley, the Millennium, and the French Revolution. Journal of the History of Ideas, 34(1):51–66, January/ March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708943.

Garside:1975:SPH

[Gar75] Peter D. Garside. Scott and the “Philosophical” historians. Journal of the History of Ideas, 36(3):497–512, July/September 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708659.

Gardiner:1977:EPB

[Gar77] Judith Kegan Gardiner. Elizabethan psychology and Burton’s anatomy of melancholy. Journal of the History of Ideas, 38(3): 373–388, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708670.

Garrett:1984:SME

[Gar84] Clarke Garrett. Swedenborg and the Mystical Enlightenment in Late Eighteenth-Century England. Journal of the History of Ideas, 45 (1):67–81, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709331.

Garrison:1987:NRM

[Gar87] James W. Garrison. Newton and the relation of mathematics to natural philosophy. Journal of the History of Ideas, 48(4):609–627, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709690.

Garrard:1996:JMC

[Gar96] Graeme Garrard. Joseph de Maistre’s civilization and its discontents. Journal of the History of Ideas, 57(3):429–446, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14911.

Garsten:2012:LRV

[Gar12] Bryan Garsten. Liberalism and the rhetorical vision of politics. Journal of the History of Ideas, 73(1):83–93, January 2012. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/463139.

Garrod:2013:IAH

[Gar13] Raphaële Garrod. Idolatry and accommodation: “Histoires” and their natural-philosophical interpretations in Simon Goulart’s Commentaires et annotations sur la Sepmaine de Du Bartas (1583). Journal of the History of Ideas, 74(3):361–380, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/511985.

Gaster:1955:BRM

[Gas55] Theodor H. Gaster. Book review: Mythic Thought in the Ancient Near East. Journal of the History of Ideas, 16(3):422–426, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707644.

Gasking:1959:WWM

[Gas59] Elizabeth B. Gasking. Why was Mendel’s work ignored? Journal of the History of Ideas, 20(1):60–84, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707967.

Gass:2000:ETS

[Gas00] Michael Gass. Eudaimonism and theology in stoic accounts of virtue. Journal of the History of Ideas, 61(1):19–37, January 2000. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15059.

Gates:1967:SIK

[Gat67] Warren E. Gates. The spread of Ibn Khaldûn’s ideas on climate and culture. Journal of the History of Ideas, 28(3):415–422, July/ September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708627.

Gaukroger:1992:DED

[Gau92] Stephen Gaukroger. Descartes’s early doctrine of clear and distinct ideas. Journal of the History of Ideas, 53(4):585–602, October/ December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709939.

Gay:1963:MFT

[Gay63] John H. Gay. Matter and freedom in the thought of Samuel Clarke. Journal of the History of Ideas, 24(1):85–105, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707860.

Gordon-Bournique:1987:LHI

[GB87] Gladys Gordon-Bournique. A. O. Lovejoy and the “History of Ideas”. Journal of the History of Ideas, 48(2):207–210, April/June 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709554.

Gal:2010:BOD

[GCM10] Ofer Gal and Raz Chen-Morris. Baroque optics and the disappearance of the observer: From Kepler’s optics to Descartes’ doubt. Journal of the History of Ideas, 71(2):191–217, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/379779.

Geerken:1976:MSS

[Gee76] John H. Geerken. Machiavelli studies since 1969. Journal of the History of Ideas, 37(2):351–368, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708831.

Geerken:1999:MMR

[Gee99] John H. Geerken. Machiavelli’s Moses and Renaissance politics. Journal of the History of Ideas, 60(4):579–595, October 1999. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15048.

Geismann:1991:SBH

[Gei91] Georg Geismann. Spinoza — beyond Hobbes and Rousseau. Journal of the History of Ideas, 52(1):35–53, January/March 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709581.

Gelber:1952:ULS

[Gel52] Sidney Gelber. Universal language and the sciences of man in Berkeley’s philosophy. Journal of the History of Ideas, 13(4):482–513, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707714.

Gelfand:1989:CRF

[Gel89] Toby Gelfand. Charcot’s response to Freud’s rebellion. Journal of the History of Ideas, 50(2):293–307, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709736.

George:1957:ECO

[Geo57] Charles H. George. English Calvinist opinion on usury, 1600–1640. Journal of the History of Ideas, 18(4):455–474, October 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707559.

Gershoy:1951:TFH

[Ger51] Leo Gershoy. Three French historians and the Revolution of 1848. Journal of the History of Ideas, 12(1):131–146, January 1951. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707541.

Gert:1967:HPE

[Ger67] Bernard Gert. Hobbes and psychological egoism. Journal of the History of Ideas, 28(4):503–520, October/December 1967. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708526.

Gerber:1972:BRS

[Ger72] William Gerber. Book review: A Short-Title List of Subject Dictionaries of the Sixteenth, Seventeenth and Eighteenth Centuries as Aids to the History of Ideas by Giorgio Tonelli. Journal of the History of Ideas, 33(1):177–178, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709065.

Gerber:1973:BRT

[Ger73] William Gerber. Book review: Is there Progress in Philosophy? Journal of the History of Ideas, 34(4):669–673, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708900.

Gershkovich:2013:ISJ

[Ger13] Tatyana Gershkovich. Infecting, simulating, judging: Tolstoy’s search for an aesthetic standard. Journal of the History of Ideas, 74(1):115– 137, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
495244.

Gerolin:2015:IAN

[Ger15] Alessandra Gerolin. The influence of Alfred North Whitehead on Eric Voegelin. Journal of the History of Ideas, 76(4):633–655, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/597350.

Germana:2019:CHP

[Ger19] Nicholas A. Germana. The Creuzerstreit and Hegel’s philosophy of history. Journal of the History of Ideas, 80(2):271–288, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/722381.

Gere:2025:DBM

[Ger25] Cathy Gere. Dig beneath a mosque [review of Greeks, Romans, Germans: How the Nazis Usurped Europe’s Classical Past, by Johann Chapoutot, trans. Richard R. Nybakken, University of California Press, 2016, 505 pp.; The Indo-Europeans: Archaeology, Language, Race, and the Search for the Origins of the West, by Jean-Paul Demoule, trans. Rhoda Cronin, Allanic Oxford University Press, 2023, 563 pp.; Bureaucratic Archaeology, State, Science, and Past in Postcolonial India, by Ashish Avikunthak, Cambridge University Press, 2022, 328 pp.; Archaeology, Nation, and Race: Confronting the Past, Decolonizing the Future in Greece and Israel, by Raphael Greenberg and Yannis Hamilakis, Cambridge University Press, 2022, 218 pp.]. Journal of the History of Ideas, 86(1):193–211, January 2025. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/949932.

Gesemann:1965:HR

[Ges65] Wolfgang Gesemann. Herder’s Russia. Journal of the History of Ideas, 26(3):424–434, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708391.

Gewirtz:1941:ENM

[Gew41] Alan Gewirtz. Experience and the non-mathematical in the Cartesian method. Journal of the History of Ideas, 2(2):183–210, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707112.

Geyl:1948:BRT

[Gey48] P. Geyl. Book review: Toynbee’s System of Civilizations. Journal of the History of Ideas, 9(1):93–124, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707397.

Geyl:1955:BRT

[Gey55] Pieter Geyl. Book review: Toynbee the Prophet. Journal of the History of Ideas, 16(2):260–274, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2707668. See comment [Toy55].

Gibbs:1972:WAT

[Gib72] Lee W. Gibbs. William Ames’s technometry. Journal of the History of Ideas, 33(4):615–624, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708860.

Gilbert:1953:BRC

[Gil53] Felix Gilbert. Book review: The Composition and Structure of Machiavelli’s Discorsi. Journal of the History of Ideas, 14(1):136–156, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707500.

Gill:1969:TAS

[Gil69] James E. Gill. Theriophily in antiquity: a supplementary account. Journal of the History of Ideas, 30(3):401–412, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708565.

Gilbert:1983:NEJ

[Gil83] Felix Gilbert. The new edition of Johann Gustav Droysen’s Historik. Journal of the History of Ideas, 44(2):327–336, April/June 1983. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709143.

Gill:1985:AP

[Gil85] Christopher Gill. Ancient psychotherapy. Journal of the History of Ideas, 46(3):307–325, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709470.

Gilbert:1986:JBS

[Gil86] Felix Gilbert. Jacob Burckhardt’s student years: The road to cultural history. Journal of the History of Ideas, 47(2):249–274, April/ June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709813.

Gilbert:1987:C

[Gil87] Neal W. Gilbert. Comment. Journal of the History of Ideas, 48 (1):41–50, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709611. See [Tri87, Obe87].

Gilmore:2020:MCS

[Gil20] Nathaniel K. Gilmore. Montesquieu’s considerations on the state of Europe. Journal of the History of Ideas, 81(3):359–379, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/761036.

Gimmel:2008:RMC

[Gim08] Millie Gimmel. Reading medicine in the Codex de la Cruz Badiano. Journal of the History of Ideas, 69(2):169–192, April 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/235870.

Girill:1970:GPM

[Gir70] T. R. Girill. Galileo and platonistic methodology. Journal of the History of Ideas, 31(4):501–520, October/December 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708257.

Giustiniani:1985:HHM

[Giu85] Vito R. Giustiniani. Homo, humanus, and the meanings of ’humanism’. Journal of the History of Ideas, 46(2):167–195, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709633.

Givner:1962:SPL

[Giv62] David A. Givner. Scientific preconceptions in Locke’s philosophy of language. Journal of the History of Ideas, 23(3):340–354, July/ September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708071.

Guerlac:1969:BNP

[GJ69] Henry Guerlac and M. C. Jacob. Bentley, Newton, and Providence: The Boyle lectures once more. Journal of the History of Ideas, 30 (3):307–318, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708559.

Glausser:1990:TAL

[Gla90] Wayne Glausser. Three approaches to Locke and the slave trade. Journal of the History of Ideas, 51(2):199–216, April/June 1990. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709512.

Glazebrook:2001:ZAM

[Gla01] Trish Glazebrook. Zeno against mathematical physics. Journal of the History of Ideas, 62(2):193–210, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15111.

Glidden:1988:HBG

[Gli88] David K. Glidden. Hellenistic background for Gassendi’s theory of ideas. Journal of the History of Ideas, 49(3):405–424, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709485.

Glover:1957:WCH

[Glo57] Richard Glover. War and civilian historians. Journal of the History of Ideas, 18(1):84–100, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707580.

Gluck:2013:BCH

[Glu13] Mary Gluck. The Budapest coffee house and the making of “Jewish Modernity” at the Fin de Siècle. Journal of the History of Ideas, 74(2): 289–306, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
504789.

Gram:1980:PPH

[GM80] Moltke S. Gram and Richard M. Martin. The perils of plenitude: Hintikka contra Lovejoy. Journal of the History of Ideas, 41(3): 497–511, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709409.

Gocer:1999:HCC

[Goc99] Asli Gocer. A hypothesis concerning the character of Islamic art. Journal of the History of Ideas, 60(4):683–692, October 1999. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15049.

Godart:2008:PRC

[God08] Gerard Clinton Godart. “Philosophy” or “religion”? The confrontation with foreign categories in Late Nineteenth-Century Japan. Journal of the History of Ideas, 69(1):71–91, January 2008. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/231009.

Goldman:1952:OBE

[Gol52] Eric F. Goldman. The origins of Beard’s economic interpretation of the constitution. Journal of the History of Ideas, 13(2):234–249, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707613.

Goldberg:1972:BCD

[Gol72] Michael Goldberg. From Bentham to Carlyle: Dickens’ political development. Journal of the History of Ideas, 33(1):61–76, January/ March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709057.

Goldberg:1982:EPU

[Gol82a] Hillel Goldberg. An early psychologist of the unconscious. Journal of the History of Ideas, 43(2):269–284, April/June 1982. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709203.

Goldstein:1982:EFT

[Gol82b] Leslie F. Goldstein. Early feminist themes in French utopian socialism: The St.-Simonians and Fourier. Journal of the History of Ideas, 43(1):91–108, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709162.

Goldsmith:1988:RAM

[Gol88] M. M. Goldsmith. Regulating anew the moral and political sentiments of mankind: Bernard Mandeville and the Scottish Enlightenment. Journal of the History of Ideas, 49(4):587–606, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709675.

Gomperz:1943:PME

[Gom43] Heinrich Gomperz. Problems and methods of Early Greek science. Journal of the History of Ideas, 4(2):161–176, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707322.

Gombrich:1981:MGB

[Gom81] E. G. Gombrich. In memory of George Boas. Journal of the History of Ideas, 42(2):334–354, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2709324. See errata [Ano81e].

Goodwin:1964:ETA

[Goo64] Craufurd D. Goodwin. Evolution theory in Australian social thought. Journal of the History of Ideas, 25(3):393–416, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707906.

Goodman:1983:BAF

[Goo83] Lenn Evan Goodman. Bag.ya on the antinomy of free will and predestination. Journal of the History of Ideas, 44(1):115–130, January/ March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709308.

Goodstein:1984:RFV

[Goo84] Judith R. Goodstein. The rise and fall of Vito Volterra’s world. Journal of the History of Ideas, 45(4):607–617, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709376.

Goodman:1990:MGS

[Goo90a] Dena Goodman. The Martin Guerre story: a non-Persian source for Persian letter CXLI. Journal of the History of Ideas, 51(2):311–316, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709518.

Goodman:1990:EWP

[Goo90b] Russell B. Goodman. East–West philosophy in Nineteenth-Century America: Emerson and Hinduism. Journal of the History of Ideas, 51 (4):625–645, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709649.

Goodrum:2002:AAS

[Goo02] Matthew R. Goodrum. Atomism, atheism, and the spontaneous generation of human beings: The debate over a natural origin of the first humans in Seventeenth-Century Britain. Journal of the History of Ideas, 63(2):207–224, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15150.

Goodrum:2011:RVP

[Goo11] Matthew R. Goodrum. Recovering the vestiges of primeval Europe: Archaeology and the significance of stone implements, 1750–1800. Journal of the History of Ideas, 72(1):51–74, January 2011. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/413474.

Gootjes:2018:FOA

[Goo18] Albert Gootjes. The first orchestrated attack on Spinoza: Johannes Melchioris and the Cartesian network in Utrecht. Journal of the History of Ideas, 79(1):23–43, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/684830.

Gorham:1994:MBD

[Gor94] Geoffrey Gorham. Mind–body dualism and the Harvey–Descartes controversy. Journal of the History of Ideas, 55(2):211–234, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709897.

Gorton:2000:PT

[Gor00] Lisa Gorton. The paradox topos. Journal of the History of Ideas, 61(2):343–346, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15070.

Gordon:2008:PSA

[Gor08] Peter E. Gordon. The place of the sacred in the absence of God: Charles Taylor’s A Secular Age. Journal of the History of Ideas, 69(4):647–673, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/251818.

Gordon:2019:IRF

[Gor19a] Peter E. Gordon. Introduction: Reflections on the fiftieth anniversary of Hans Blumenberg’s The Legitimacy of the Modern Age. Journal of the History of Ideas, 80(1):67–73, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/715736.

Gordon:2019:SGL

[Gor19b] Peter E. Gordon. Secularization, genealogy, and the legitimacy of the modern age: Remarks on the Löwith–Blumenberg debate. Journal of the History of Ideas, 80(1):147–170, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/715741.

Gossman:1960:TUE

[Gos60] Lionel Gossman. Two unpublished essays on mathematics in the Hume papers. Journal of the History of Ideas, 21(3):442–449, July/ September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708148.

Gottschalk:1944:FPJ

[Got44] Louis Gottschalk. The French parlements and judicial review. Journal of the History of Ideas, 5(1):105–112, January 1944. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707104.

Gottfried:1975:ASC

[Got75] Paul Gottfried. Arthur Schopenhauer as a critic of history. Journal of the History of Ideas, 36(2):331–338, April/June 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708931.

Gottfried:1980:SIC

[Got80] Paul Gottfried. On the social implications and context of the Hegelian dialectic. Journal of the History of Ideas, 41(3):421–432, July/ September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709402.

Gould:1945:GMM

[Gou45] S. H. Gould. Gulliver and the moons of Mars. Journal of the History of Ideas, 6(1):91–101, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707059.

Gould:1961:CAS

[Gou61] James A. Gould. The concept of absolute space. Journal of the History of Ideas, 22(1):119–120, January/March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707879.

Gould:1974:SCF

[Gou74] Josiah B. Gould. The stoic conception of fate. Journal of the History of Ideas, 35(1):17–32, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708740.

Goulding:2006:HSE

[Gou06a] Robert Goulding. Histories of science in Early Modern Europe: Introduction. Journal of the History of Ideas, 67(1):33–40, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/194643.

Goulding:2006:MMP

[Gou06b] Robert Goulding. Method and mathematics: Peter Ramus’s histories of the sciences. Journal of the History of Ideas, 67(1):63–85, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/194644.

Gouwens:2010:EAC

[Gou10] Kenneth Gouwens. Erasmus, “Apes of Cicero,” and conceptual blending. Journal of the History of Ideas, 71(4):523–545, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/398296.

Gernand:1986:PKS

[GR86] Herbert W. Gernand and W. Jay Reedy. Planck, Kuhn, and scientific revolutions. Journal of the History of Ideas, 47(3):469–485, July/ September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709664.

Grace:1944:JMM

[Gra44] William J. Grace. Jacques Maritain and modern Catholic historical scholarship. Journal of the History of Ideas, 5(4):434–445, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707080.

Gray:1953:HEN

[Gra53] J. Glenn Gray. Heidegger “evaluates” Nietzsche. Journal of the History of Ideas, 14(2):304–309, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707479.

Graham:1961:LPE

[Gra61] John Graham. Lavater’s physiognomy in England. Journal of the History of Ideas, 22(4):561–572, October/December 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708032.

Grant:1962:LMT

[Gra62] Edward Grant. Late medieval thought, Copernicus, and the scientific revolution. Journal of the History of Ideas, 23(2):197–220, April/ June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708155.

Grace:1963:MSN

[Gra63a] William J. Grace. Milton, Salmasius, and the natural law. Journal of the History of Ideas, 24(3):323–336, July/September 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708210.

Gray:1963:RHP

[Gra63b] Hanna H. Gray. Renaissance humanism: The pursuit of eloquence. Journal of the History of Ideas, 24(4):497–514, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707980.

Grant:1965:CPF

[Gra65] A. Cameron Grant. Combe on phrenology and free will: a note on XIXth-Century secularism. Journal of the History of Ideas, 26 (1):141–147, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708406.

Grant:1968:NLO

[Gra68] A. Cameron Grant. New light on an old view. Journal of the History of Ideas, 29(2):293–301, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708582.

Gracia:1975:BRI

[Gra75] Jorge J. E. Gracia. Book review: Importance of the History of Ideas in Latin America: Zea’s Positivism in Mexico. Journal of the History of Ideas, 36(1):177–184, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709020.

Gray:1978:HSH

[Gra78] Robert Gray. Hobbes’ system and his early philosophical views. Journal of the History of Ideas, 39(2):199–215, April/June 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708775.

Gram:1981:IIC

[Gra81a] Moltke S. Gram. Intellectual intuition: The continuity thesis. Journal of the History of Ideas, 42(2):287–304, April/June 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709321.

Gras:1981:MRO

[Gra81b] Vernon W. Gras. Myth and the reconciliation of opposites: Jung and lévi-strauss. Journal of the History of Ideas, 42(3):471–488, July/ September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709188.

Graham:1986:SMA

[Gra86] Daniel W. Graham. Some myths about Aristotle’s biological motivation. Journal of the History of Ideas, 47(4):529–545, October/ December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709716.

Gravelle:1988:LVQ

[Gra88] Sarah Stever Gravelle. The Latin–vernacular question and humanist theory of language and culture. Journal of the History of Ideas, 49 (3):367–386, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709483.

Gravelle:1989:NTT

[Gra89] Sarah Stever Gravelle. A new theory of truth. Journal of the History of Ideas, 50(2):333–336, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709739.

Grafton:1992:KR

[Gra92] Anthony Grafton. Kepler as a reader. Journal of the History of Ideas, 53(4):561–572, October/December 1992. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709937.

Grafton:2003:SUE

[Gra03] Anthony Grafton. Some uses of eclipses in early modern chronology. Journal of the History of Ideas, 64(2):213–229, April 2003. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/43509.

Grafton:2006:HIP

[Gra06] Anthony Grafton. The history of ideas: Precept and practice, 1950– 2000 and beyond. Journal of the History of Ideas, 67(1):1–32, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/194645.

Grene:1943:HST

[Gre43] Marjorie Grene. Hume: Sceptic and tory? Journal of the History of Ideas, 4(3):333–348, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707258.

Greene:1953:SBS

[Gre53] D. J. Greene. Smart, Berkeley, the scientists and the poets: a note on Eighteenth-Century anti-Newtonianism. Journal of the History of Ideas, 14(3):327–352, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707805.

Greene:1954:ADN

[Gre54] John C. Greene. The American debate on the Negro’s place in nature, 1780–1815. Journal of the History of Ideas, 15(3):384–396, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707761.

Greene:1962:HMR

[Gre62] Robert A. Greene. Henry more and Robert Boyle on the spirit of nature. Journal of the History of Ideas, 23(4):451–474, October/ December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708169.

Gregor:1963:GGP

[Gre63] A. James Gregor. Giovanni Gentile and the philosophy of the young Karl Marx. Journal of the History of Ideas, 24(2):213–230, April/ June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707846.

Grennen:1964:CCC

[Gre64] Joseph E. Grennen. Chaucer’s characterization of the canon and his yeoman. Journal of the History of Ideas, 25(2):279–284, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708016.

Grendler:1965:URI

[Gre65] Paul F. Grendler. Utopia in Renaissance Italy: Doni’s “New World”. Journal of the History of Ideas, 26(4):479–494, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708495.

Green:1967:HIF

[Gre67] Louis Green. Historical interpretation in Fourteenth-Century Florentine chronicles. Journal of the History of Ideas, 28(2):161–178, April/ June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708413.

Greaves:1969:PSA

[Gre69a] Richard L. Greaves. Puritanism and science: The anatomy of a controversy. Journal of the History of Ideas, 30(3):345–368, July/ September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708562.

Greenlee:1969:BRP

[Gre69b] Douglas Greenlee. Book review: On Pragmatism. Journal of the History of Ideas, 30(4):603–608, October/December 1969. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708614.

Greenlee:1972:LCI

[Gre72a] Douglas Greenlee. Locke and the controversy over innate ideas. Journal of the History of Ideas, 33(2):251–264, April/June 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708872.

Grene:1972:AMB

[Gre72b] Marjorie Grene. Aristotle and modern biology. Journal of the History of Ideas, 33(3):395–424, July/September 1972. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709043. See erratum [Ano73e].

Green:1974:SMS

[Gre74] Paul D. Green. Spenser and the masses: Social commentary in The Faerie Queene. Journal of the History of Ideas, 35(3):389–406, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708790.

Green:1975:MCC

[Gre75] Jeffrey Martin Green. Montaigne’s critique of Cicero. Journal of the History of Ideas, 36(4):595–612, October/December 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708991.

Greetham:1980:CNB

[Gre80] D. C. Greetham. The concept of nature in Bartholomaeus Anglicus (FL. 1230). Journal of the History of Ideas, 41(4):663–677, October/ December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709279.

Greene:1981:WWI

[Gre81] Robert A. Greene. Whichcote, Wilkins, “Ingenuity,” and the reasonableness of Christianity. Journal of the History of Ideas, 42 (2):227–252, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709318.

Greene:1991:SSC

[Gre91a] Robert A. Greene. Synderesis, the spark of conscience, in the English Renaissance. Journal of the History of Ideas, 52(2):195–219, April/ June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709525.

Greene:1991:WCL

[Gre91b] Robert A. Greene. Whichcote, the candle of the Lord, and synderesis. Journal of the History of Ideas, 52(4):617–644, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709969.

Grendler:1991:RRB

[Gre91c] Paul F. Grendler. Reply to Robert Black. Journal of the History of Ideas, 52(2):335–337, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2709532. See [Bla91a, Bla91b].

Greene:1997:INN

[Gre97] Robert A. Greene. Instinct of nature: Natural law, synderesis, and the moral sense. Journal of the History of Ideas, 58(2):173–198, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14941.

Gregory:2014:CIS

[Gre14] Brad Gregory. Christian ideas in the Sixteenth and Seventeenth Centuries. Journal of the History of Ideas, 75(4):667–675, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/557239.

Greenham:2015:ABN

[Gre15] David Greenham. “Altars to the beautiful necessity”: The significance of F. W. J. Schelling’s “Philosophical Inquiries in the Nature of Human Freedom” in the development of Ralph Waldo Emerson’s concept of fate. Journal of the History of Ideas, 76(1):115–137, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/566379.

Green:2016:JAM

[Gre16] Jonathan Green. John Adams’s Montesquieuean moment: Enlightened historicism in the discourses on Davila. Journal of the History of Ideas, 77(2):227–251, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/620980.

Greenberg:2018:CPT

[Gre18] Udi Greenberg. Catholics, Protestants, and the tortured path to religious liberty. Journal of the History of Ideas, 79(3):461–479, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/703669.

Green:2023:WRK

[Gre23a] Karen Green. Women’s reception of Kant, 1790–1810. Journal of the History of Ideas, 84(2):263–285, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/886182.

Greenberg:2023:RGA

[Gre23b] Udi Greenberg. The remnants of giorgio agamben: The omnibus homo sacer upon its completion. Journal of the History of Ideas, 84(1):179–196, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/879545.

Grilli:1941:NPB

[Gri41] Marcel Grilli. The nationality of philosophy and bertrando spaventa. Journal of the History of Ideas, 2(3):339–371, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707136.

Griffiths:1956:AHF

[Gri56] J. Gwyn Griffiths. Archaeology and Hesiod’s five ages. Journal of the History of Ideas, 17(1):109–119, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707688.

Griffiths:1958:DHI

[Gri58] J. Gwyn Griffiths. Did Hesiod invent the “Golden Age”? Journal of the History of Ideas, 19(1):91–93, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707954.

Grimsley:1965:KL

[Gri65] Ronald Grimsley. Kierkegaard and Leibniz. Journal of the History of Ideas, 26(3):383–396, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708388.

Griffin:1974:HIH

[Gri74] C. M. Griffin. L. T. Hobhouse and the idea of harmony. Journal of the History of Ideas, 35(4):647–661, October/December 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709091.

Grigoriev:2012:CWT

[Gri12] Serge Grigoriev. Chauncey Wright: Theoretical reason in a naturalist account of human consciousness. Journal of the History of Ideas, 73(4):559–582, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488348.

Grinchpun:2021:SMN

[Gri21] Matías Grinchpun. Not six million nor thirty thousand: From “Holocaust Revisionism” to “State Terrorism” denial in Argentina, 1945– 2016. Journal of the History of Ideas, 82(1):153–174, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/781617.

Groth:1950:WMN

[Gro50] J. H. Groth. Wilamowitz-Möllendorf on Nietzsche’s birth of tragedy. Journal of the History of Ideas, 11(2):179–190, April 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707245.

Groenewegen:1967:ANP

[Gro67] P. D. Groenewegen. Authorship of the natural and political observations upon the bills of mortality. Journal of the History of Ideas, 28 (4):601–602, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708534.

Grossmann:1979:TER

[Gro79] Walter Grossmann. Toleration–exercitium religionis privatum. Journal of the History of Ideas, 40(1):129–134, January/March 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709265.

Grosholz:1996:PLA

[Gro96] Emily Grosholz. Plato and Leibniz against the materialists. Journal of the History of Ideas, 57(2):255–276, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14903.

Grosse:1999:RHJ

[Gro99] Jürgen Grosse. Reading history: On Jacob Burckhardt as sourcereader. Journal of the History of Ideas, 60(3):525–547, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15036.

Grosholz:2011:RAR

[Gro11] Emily Grosholz. Reference and analysis: The representation of time in Galileo, Newton, and Leibniz. Journal of the History of Ideas, 72(3):

333–350, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
445501.

Grote:2014:RER

[Gro14] Simon Grote. Review-essay: Religion and Enlightenment. Journal of the History of Ideas, 75(1):137–160, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/535157.

Grundy:1973:STE

[Gru73] Dominick Grundy. Skepticism in two essays by Montaigne and Sir Thomas Browne. Journal of the History of Ideas, 34(4):529–542, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708886.

Groarke:1991:SSH

[GS91] Leo Groarke and Graham Solomon. Some sources for Hume’s account of cause. Journal of the History of Ideas, 52(4):645–663, October/ December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709970.

Giovannetti-Singh:2023:ACJ

[GS23] Gianamar Giovannetti-Singh. Astronomical chronology, the Jesuit China mission, and Enlightenment history. Journal of the History of Ideas, 84(3):487–510, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/901491.

Greenberg:2018:ISF

[GSJ18] Udi Greenberg and Daniel Steinmetz-Jenkins. Introduction: Special forum on Christianity and human rights. Journal of the History of Ideas, 79(3):407–409, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/703665.

Guarneri:1982:IFA

[Gua82] Carl J. Guarneri. Importing fourierism to America. Journal of the History of Ideas, 43(4):581–594, October/December 1982. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709344.

Gubser:2005:THA

[Gub05] Mike Gubser. Time and history in Alois Riegl’s theory of perception. Journal of the History of Ideas, 66(3):451–474, July 2005. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/188684.

Gucer:2000:HEP

[Guc00] Kathryn Gucer. “Not heretofore extant in print”: Where the mad ranters are. Journal of the History of Ideas, 61(1):75–95, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15060.

Gudding:1996:PGD

[Gud96] Gabriel Gudding. The phenotype/genotype distinction and the disappearance of the body. Journal of the History of Ideas, 57(3):525–545, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14912.

Guerlac:1968:CAC

[Gue68] Henry Guerlac. Copernicus and Aristotle’s Cosmos. Journal of the History of Ideas, 29(1):109–113, January/March 1968. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708468.

Guerlac:1978:APO

[Gue78] Henry Guerlac. Amicus Plato and other friends. Journal of the History of Ideas, 39(4):627–633, October/December 1978. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709446.

Guerlac:1983:TVB

[Gue83] Henry Guerlac. Theological voluntarism and biological analogies in Newton’s physical thought. Journal of the History of Ideas, 44(2):219–229, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709137.

Guerlac:1986:CTC

[Gue86] Henry Guerlac. Can there be colors in the dark? Physical color theory before Newton. Journal of the History of Ideas, 47(1):3–20, January/ March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709592.

Guerrini:1989:EAE

[Gue89] Anita Guerrini. The ethics of animal experimentation in Seventeenth-Century England. Journal of the History of Ideas, 50(3):391– 407, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709568.

Guelzo:1995:CMR

[Gue95] Allen C. Guelzo. From Calvinist metaphysics to republican theory: Jonathan Edwards and James Dana on freedom of the will. Journal of the History of Ideas, 56(3):399–418, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2710033.

Guggisberg:1980:BRP

[Gug80] Hans R. Guggisberg. Book review: Protestant Ideals of Education in Historical Perspective: Two Approaches. Journal of the History of Ideas, 41(4):693–698, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709281.

Guilhamet:1985:SPS

[Gui85] Leon Guilhamet. Socrates and post-Socratic satire. Journal of the History of Ideas, 46(1):3–12, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709772.

Guillemin:2008:SLA

[Gui08] Anna Guillemin. The style of linguistics: Aby Warburg, Karl Vossler, and Hermann Osthoff. Journal of the History of Ideas, 69(4):605– 626, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
251814.

Guicciardini:2013:RMA

[Gui13] Niccolò Guicciardini. The role of musical analogies in Newton’s optical and cosmological work. Journal of the History of Ideas, 74(1):45– 67, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
495241.

Guilhot:2023:PKS

[Gui23] Nicolas Guilhot. “A primitive kind of superstition”: The idea of the paranoid style in art, psychiatry, and politics. Journal of the History of Ideas, 84(2):365–390, April 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/886186.

Gullace:1962:PMI

[Gul62] Giovanni Gullace. The pragmatist movement in Italy. Journal of the History of Ideas, 23(1):91–105, January/March 1962. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708059.

Gundersheimer:1962:LLH

[Gun62] Werner L. Gundersheimer. Louis Leroy’s humanistic optimism. Journal of the History of Ideas, 23(3):324–339, July/September 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708070.

Gunn:1968:IWL

[Gun68] J. A. W. Gunn. “Interest will not lie”: a Seventeenth-Century political Maxim. Journal of the History of Ideas, 29(4):551–564, October/ December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708293.

Gunter:1971:BTM

[Gun71] P. A. Y. Gunter. Bergson’s theory of matter and modern cosmology. Journal of the History of Ideas, 32(4):525–542, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708975.

Gunter:1982:BJ

[Gun82] Pete A. Y. Gunter. Bergson and Jung. Journal of the History of Ideas, 43(4):635–652, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709347.

Gupta:1975:FS

[Gup75] R. K. Gupta. Freud and Schopenhauer. Journal of the History of Ideas, 36(4):721–728, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708999.

Gurstein:2000:TCW

[Gur00] Rochelle Gurstein. Taste and “the conversible world” in the Eighteenth Century. Journal of the History of Ideas, 61(2):203–221, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15071.

Gustavson:1950:GLP

[Gus50] Carl G. Gustavson. German Lutheranism: a psychological study. Journal of the History of Ideas, 11(2):140–158, April 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707243.

Gutman:1941:MCR

[Gut41] Harry B. Gutman. The medieval content of Raphael’s “School of Athens”. Journal of the History of Ideas, 2(4):420–429, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707020.

Guthrie:1979:RRG

[Gut79] W. K. C. Guthrie. Reply to R. G. Mulgan. Journal of the History of Ideas, 40(1):128, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709264. See [Mul79].

Guyer:2002:LWB

[Guy02] Paul Guyer. Lewis White Beck on reasons and causes. Journal of the History of Ideas, 63(3):539–545, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15162.

Guy:2020:NER

[Guy20] Stéphane Guy. Negotiating an “Economic Revolution”: History, collectivism, and liberalism in William Clarke’s thought. Journal of the History of Ideas, 81(4):621–642, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/772513.

Ha:2020:RNM

[Ha20] Polly Ha. Revolutionizing the new model army: Ecclesiastical independence, social justice, and political legitimacy. Journal of the History of Ideas, 81(4):531–553, October 2020. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/772514.

Haar:1948:YCD

[Haa48] Charles M. Haar. E. L. Youmans: a chapter in the diffusion of science in America. Journal of the History of Ideas, 9(2):193–213, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707228.

Haara:2016:PPS

[Haa16] Heikki Haara. Pufendorf on passions and sociability. Journal of the History of Ideas, 77(3):423–444, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/626909.

Habib:1990:PCE

[Hab90] M. A. R. Habib. The prayers of childhood: T. S. Eliot’s manuscripts on Kant. Journal of the History of Ideas, 51(1):93–114, January/ March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709748.

Habib:1993:BRE

[Hab93] M. A. R. Habib. “Bergson Resartus” and T. S. Eliot’s manuscript. Journal of the History of Ideas, 54(2):255–276, April 1993. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709982.

Haber:1995:SES

[Hab95] Honi Haber. Schopenhauer as the embodiment of the Socratic and postmodern man: an examination of character. Journal of the History of Ideas, 56(3):483–499, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710037.

Hacking:1983:NCC

[Hac83] Ian Hacking. Nineteenth Century cracks in the concept of determinism. Journal of the History of Ideas, 44(3):455–475, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709176.

Hacohen:1998:KPV

[Hac98] Malachi H. Hacohen. Karl Popper, the Vienna Circle, and Red Vienna. Journal of the History of Ideas, 59(4):711–734, October 1998.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15004.

Hadas:1943:NCG

[Had43] Moses Hadas. From nationalism to cosmopolitanism in the greco-Roman world. Journal of the History of Ideas, 4(1):105–111, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707238.

Hadas:1950:ANS

[Had50] Moses Hadas. Aspects of nationalist survival under Hellenistic and Roman imperialism. Journal of the History of Ideas, 11(2):131–139, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707242.

Hadas:1952:BRR

[Had52] Moses Hadas. Book review: Recent Work in Ancient Politics. Journal of the History of Ideas, 13(1):110–118, January 1952. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707729.

Hadas:1958:BRO

[Had58a] Moses Hadas. Book review: Order and History. Journal of the History of Ideas, 19(3):442–444, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708049.

Hadas:1958:PHF

[Had58b] Moses Hadas. Plato in Hellenistic fusion. Journal of the History of Ideas, 19(1):3–13, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707948.

Haddock:1979:VDT

[Had79] B. A. Haddock. Vico’s “Discovery of the True Homer”: a case-study in historical reconstruction. Journal of the History of Ideas, 40(4): 583–602, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709360.

Hainds:1946:JSM

[Hai46] J. R. Hainds. John Stuart Mill and the Saint Simonians. Journal of the History of Ideas, 7(1):103–112, January 1946. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707272.

Hainds:1950:JME

[Hai50] J. R. Hainds. J. S. Mill’s examiner articles on art. Journal of the History of Ideas, 11(2):215–234, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707247.

Hall:1950:SOP

[Hal50] Thomas S. Hall. The scientific origins of the protoplasm problem. Journal of the History of Ideas, 11(3):339–356, June 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707735.

Hallie:1957:HBM

[Hal57] Philip P. Hallie. Hume, Biran and the méditatifs intérieurs. Journal of the History of Ideas, 18(3):295–312, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707795.

Halsted:1958:WBT

[Hal58] John B. Halsted. Walter Bagehot on toleration. Journal of the History of Ideas, 19(1):119–128, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707958.

Hall:1974:AGD

[Hal74] Thadd E. Hall. Abbé Gaudin’s development: From Enlightenment to revolution. Journal of the History of Ideas, 35(3):407–428, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708791.

Halverson:1992:HGO

[Hal92] John Halverson. Havelock on Greek orality and literacy. Journal of the History of Ideas, 53(1):148–163, January/March 1992. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709915.

Halevi:2002:TDN

[Hal02] Leor Halevi. The theologian’s doubts: Natural philosophy and the skeptical games of Ghazali. Journal of the History of Ideas, 63(1):19– 39, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15143.

Hammen:1953:SC

[Ham53] Oscar J. Hammen. The spectre of Communism in the 1840’s. Journal of the History of Ideas, 14(3):404–420, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707809.

Ham:1958:BRK

[Ham58] Edward B. Ham. Book review: Knox on Jansenist Enthusiasm. Journal of the History of Ideas, 19(2):283–293, April 1958. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707942.

Hamburger:1959:ACC

[Ham59] Max Hamburger. Aristotle and Confucius: a comparison. Journal of the History of Ideas, 20(2):236–249, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707821.

Hammen:1970:YMR

[Ham70] Oscar J. Hammen. The young Marx, reconsidered. Journal of the History of Ideas, 31(1):109–120, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708373.

Hammen:1972:ACR

[Ham72] Oscar J. Hammen. Alienation, Communism, and revolution in the Marx–Engels Briefwechsel. Journal of the History of Ideas, 33 (1):77–100, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709058.

Hamlin:1996:IAD

[Ham96] William M. Hamlin. Imagined apotheoses: Drake, Harriot, and Ralegh in the Americas. Journal of the History of Ideas, 57(3):405– 428, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14909.

Hampton:1998:ESS

[Ham98] Timothy Hampton. Examples, stories, and subjects in Don quixote and the heptameron. Journal of the History of Ideas, 59(4):597–611, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15005.

Hamerton:2008:MTS

[Ham08] Katharine J. Hamerton. Malebranche, taste, and sensibility: The origins of sensitive taste and a reconsideration of Cartesianism’s feminist potential. Journal of the History of Ideas, 69(4):533–558, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/251823.

Hanson:1959:CRK

[Han59] Norwood Russell Hanson. Copernicus’ role in Kant’s revolution. Journal of the History of Ideas, 20(2):274–281, April 1959. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707825.

Hanson:1960:WPN

[Han60] Norwood Russell Hanson. Waves, particles, and Newton’s ’fits’. Journal of the History of Ideas, 21(3):370–391, July/September 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708143.

Hanson:1961:CDK

[Han61] Norwood Russell Hanson. The Copernican disturbance and the Keplerian Revolution. Journal of the History of Ideas, 22(2):169–184, April/June 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707831.

Hankins:1967:IMS

[Han67] Thomas L. Hankins. The influence of Malebranche on the science of mechanics during the Eighteenth Century. Journal of the History of Ideas, 28(2):193–210, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708415.

Hankins:1991:BRF

[Han91] James Hankins. Book review: Forging Links with the Past. Journal of the History of Ideas, 52(3):509–518, July/September 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710049.

Hannaway:1992:GAH

[Han92] Owen Hannaway. Georgius Agricola as humanist. Journal of the History of Ideas, 53(4):553–560, October/December 1992. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709936.

Hankins:1995:BTA

[Han95] James Hankins. The “Baron Thesis” after forty years and some recent studies of Leonardo Bruni. Journal of the History of Ideas, 56(2):309– 338, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709840.

Handley:2017:DNS

[Han17] Sasha Handley. Deformities of nature: Sleepwalking and nonconscious states of mind in Late Eighteenth-Century Britain. Journal of the History of Ideas, 78(3):401–425, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/665521.

Hanebrink:2018:ATS

[Han18] Paul Hanebrink. An anti-totalitarian saint: The canonization of Edith Stein. Journal of the History of Ideas, 79(3):481–495, July 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/703670.

Hartshorne:1946:LGD

[Har46] Charles Hartshorne. Leibniz’s greatest discovery. Journal of the History of Ideas, 7(4):411–421, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707032.

Harrah:1959:ILM

[Har59] David Harrah. The influence of logic and mathematics on Whitehead. Journal of the History of Ideas, 20(3):420–430, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708119.

Hartigan:1966:SAW

[Har66] Richard Shelly Hartigan. Saint Augustine on war and killing: The problem of the innocent. Journal of the History of Ideas, 27(2):195–204, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708638.

Harrison:1971:EDV

[Har71] James Harrison. Erasmus Darwin’s view of evolution. Journal of the History of Ideas, 32(2):247–264, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708279.

Hare:1972:RGH

[Har72] Peter H. Hare. Rowland G. Hazard (1801–88) on freedom in willing. Journal of the History of Ideas, 33(1):155–164, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709063.

Harlow:1985:CCW

[Har85a] Barbara Harlow. Cairo curiosities: E. W. Lane’s account and Ahmad Amin’s dictionary. Journal of the History of Ideas, 46(2):279–286, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709639.

Harp:1985:TCD

[Har85b] Gillis J. Harp. Taylor, calhoun, and the decline of a theory of political disharmony. Journal of the History of Ideas, 46(1):107–120, January/ March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709778.

Hartman:1986:HCP

[Har86] Mark Hartman. Hobbes’s concept of political revolution. Journal of the History of Ideas, 47(3):487–495, July/September 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709665.

Hariman:1989:CMM

[Har89] Robert Hariman. Composing modernity in Machiavelli’s Prince. Journal of the History of Ideas, 50(1):3–29, January/March 1989. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709784.

Harris:1990:BRH

[Har90] H. S. Harris. Book review: History of the History of Philosophy. Journal of the History of Ideas, 51(1):115–120, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709749.

Harris:1993:RAR

[Har93] Marc L. Harris. Revelation and the American Republic: Timothy Dwight’s civic participation. Journal of the History of Ideas, 54(3): 449–468, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710023.

Harrison:1995:NSM

[Har95] Peter Harrison. Newtonian science, miracles, and the laws of nature. Journal of the History of Ideas, 56(4):531–553, October 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709991.

Harrison:1998:VAS

[Har98] Peter Harrison. The virtues of animals in Seventeenth-Century thought. Journal of the History of Ideas, 59(3):463–484, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14991.

Harrison:1999:PEM

[Har99] Peter Harrison. Prophecy, early modern apologetics, and Hume’s argument against miracles. Journal of the History of Ideas, 60(2):241– 256, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15028.

Hargraves:2002:PAC

[Har02a] Neil Hargraves. The “Progress of Ambition”: Character, narrative, and philosophy in the works of William Robertson. Journal of the History of Ideas, 63(2):261–282, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15151.

Harrison:2002:OSP

[Har02b] Peter Harrison. Original sin and the problem of knowledge in Early Modern Europe. Journal of the History of Ideas, 63(2):239–259, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15152.

Hark:2003:SST

[Har03] Michel Ter Hark. Searching for the searchlight theory: From Karl Popper to Otto Selz. Journal of the History of Ideas, 64(3):465–487, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/47660.

Harrison:2011:ASH

[Har11] Peter Harrison. Adam Smith and the history of the invisible hand. Journal of the History of Ideas, 72(1):29–49, January 2011. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/413473.

Harvey:2013:SBM

[Har13] Warren Zev Harvey. Spinoza on biblical miracles. Journal of the History of Ideas, 74(4):659–675, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/523434.

Hathaway:1968:CRP

[Hat68a] R. F. Hathaway. Cicero, De Re Publica II, and his Socratic view of history. Journal of the History of Ideas, 29(1):3–12, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708462.

Hattaway:1968:PSL

[Hat68b] Michael Hattaway. Paradoxes of Solomon: Learning in the English Renaissance. Journal of the History of Ideas, 29(4):499–530, October/ December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708291.

Hatch:1975:JPA

[Hat75] Ronald B. Hatch. Joseph Priestley: an addition to Hartley’s observations. Journal of the History of Ideas, 36(3):548–550, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708663.

Hattaway:1978:BKB

[Hat78] Michael Hattaway. Bacon and “Knowledge Broken”: Limits for scientific method. Journal of the History of Ideas, 39(2):183–197, April/ June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708774. See

reply [Hor82].

Haugen:1998:OIT

[Hau98] Kristine Louise Haugen. Ossian and the invention of textual history. Journal of the History of Ideas, 59(2):309–327, April 1998. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14982.

Haugen:2011:BTC

[Hau11] Kristine Louise Haugen. The birth of tragedy in the Cinquecento: Humanism and literary history. Journal of the History of Ideas, 72(3): 351–370, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
445500.

Havens:1953:SCC

[Hav53] Raymond D. Havens. Simplicity, a changing concept. Journal of the History of Ideas, 14(1):3–32, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707493.

Haven:1959:CHM

[Hav59] Richard Haven. Coleridge, Hartley, and the mystics. Journal of the History of Ideas, 20(4):477–494, October/December 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707887.

Hawkins:1994:DOC

[Haw94] M. J. Hawkins. Durkheim on occupational corporations: an exegesis and interpretation. Journal of the History of Ideas, 55(3):461–481, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709850.

Hawkes:2001:PCJ

[Haw01] David Hawkes. The politics of character in John Milton’s divorce tracts. Journal of the History of Ideas, 62(1):141–160, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15100.

Hay:1956:PCC

[Hay56] William H. Hay. Paul Carus: a case-study of philosophy on the frontier. Journal of the History of Ideas, 17(4):498–510, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707784.

Haywood:1959:IMS

[Hay59] C. Robert Haywood. The influence of mercantilism on social attitudes in the south, 1700–1763. Journal of the History of Ideas, 20 (4):577–586, October/December 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707893.

Hakeda:1963:BRP

[HE63] Yoshito Hakeda and Ainslie T. Embree. Book review: Philosophy and Culture, East and West. Journal of the History of Ideas, 24(4): 589–592, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707989.

Headley:1987:BRR

[Hea87] John M. Headley. Book review: The Reformation in Historical Thought. Journal of the History of Ideas, 48(3):521–532, July/ September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709766.

Headley:1988:RHM

[Hea88] John M. Headley. On the rearming of heaven: The machiavellism of tommaso campanella. Journal of the History of Ideas, 49 (3):387–404, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709484.

Hecht:2000:VLR

[Hec00] Jennifer Michael Hecht. Vacher de Lapouge and the rise of Nazi science. Journal of the History of Ideas, 61(2):285–304, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15072.

Hedges:1959:KBF

[Hed59] William L. Hedges. Knickerbocker, Bolingbroke, and the fiction of history. Journal of the History of Ideas, 20(3):317–328, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708112.

Hedley:1998:CII

[Hed98] Douglas Hedley. Coleridge’s intellectual intuition, the vision of God, and the walled garden of “Kubla Khan”. Journal of the History of Ideas, 59(1):115–134, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/14971.

Heffernan:1973:STA

[Hef73] William Heffernan. The slave trade and abolition in travel literature. Journal of the History of Ideas, 34(2):185–208, April/June 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708725.

Heffernan:1978:SOI

[Hef78] William C. Heffernan. The singularity of our inhabited world: William Whewell and A. R. Wallace in dissent. Journal of the History of Ideas, 39(1):81–100, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709073.

Heffernan:1981:PLD

[Hef81] William C. Heffernan. Percival Lowell and the debate over extraterrestrial life. Journal of the History of Ideas, 42(3):527–530, July/ September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709191.

Heilbroner:1973:PPD

[Hei73] Robert L. Heilbroner. The paradox of progress: Decline and decay in the wealth of nations. Journal of the History of Ideas, 34(2):243–262, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708728.

Heimann:1978:VIC

[Hei78] P. M. Heimann. Voluntarism and immanence: Conceptions of nature in Eighteenth-Century thought. Journal of the History of Ideas, 39(2):271–283, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708779.

Helm:1964:SRH

[Hel64] Bertrand Helm. Social roots of the heraclitean metaphysics. Journal of the History of Ideas, 25(4):565–571, October/December 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708186.

Hellman:1973:OLF

[Hel73] John Hellman. The opening to the left in French Catholicism: The role of the personalists. Journal of the History of Ideas, 34 (3):381–390, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708959.

Helm:1981:DLC

[Hel81] Paul Helm. Did Locke capitulate to Molyneux? Journal of the History of Ideas, 42(4):669–671, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709125.

Helbig:2019:LTH

[Hel19] Daniela K. Helbig. Life without toothache: Hans Blumenberg’s Zettelkasten and history of science as theoretical attitude. Journal of the History of Ideas, 80(1):91–112, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/715738.

Hellemans:2021:IMM

[Hel21a] Babette Hellemans. The immeasurability of the monastic mind: Writing about Peter abelard (1079–1142). Journal of the History of Ideas, 82(4):683–701, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/837677.

Hellman:2021:DLT

[Hel21b] Lisa Hellman. Drawing the lines: Translation and diplomacy in the Central Asian borderlands. Journal of the History of Ideas, 82(3): 485–501, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
798276.

Hendel:1944:A

[Hen44] C. W. Hendel. Announcements. Journal of the History of Ideas, 5(1):125–127, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707108.

Heninger:1961:RPP

[Hen61] S. K. Heninger, Jr. The Renaissance perversion of pastoral. Journal of the History of Ideas, 22(2):254–261, April/June 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707836. See errata [Ano61d].

Henze:1969:LAH

[Hen69] Donald F. Henze. The linguistic aspect of Hume’s method. Journal of the History of Ideas, 30(1):116–126, January/March 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708249.

Henry:1977:VM

[Hen77] Patrick Henry. Voltaire as moralist. Journal of the History of Ideas, 38(1):141–146, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708846.

Hendrix:1995:MPR

[Hen95] Scott Hendrix. Masculinity and patriarchy in Reformation Germany. Journal of the History of Ideas, 56(2):177–193, April 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709834.

Henry:2001:AEC

[Hen01] John Henry. Animism and empiricism: Copernican physics and the origins of William Gilbert’s experimental method. Journal of the History of Ideas, 62(1):99–119, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15101.

Hepburn:1955:GHV

[Hep55] Ronald W. Hepburn. George Hakewill: The virility of nature. Journal of the History of Ideas, 16(2):135–150, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707659.

Herrick:1948:SRM

[Her48] Francis H. Herrick. The second reform movement in Britain 1850– 1865. Journal of the History of Ideas, 9(2):174–192, April 1948. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707227.

Hershbell:1970:BPE

[Her70] Jackson P. Hershbell. Berkeley and the problem of evil. Journal of the History of Ideas, 31(4):543–554, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708259.

Herbert:1987:HPS

[Her87] Gary B. Herbert. Hobbes’s phenomenology of space. Journal of the History of Ideas, 48(4):709–717, October/December 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709695.

Herman:1992:HRA

[Her92] Arthur Herman. The Huguenot Republic and antirepublicanism in Seventeenth-Century France. Journal of the History of Ideas, 53(2):249–269, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709873.

Herrick:2005:BWN

[Her05] Tim Herrick. “A book which is no longer discussed today”: Tran Duc Thao, Jacques Derrida, and Maurice Merleau-Ponty. Journal of the History of Ideas, 66(1):113–131, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/185259.

Heron:2021:SOH

[Her21] Nicholas Heron. The superhuman origins of human dignity: Kantorowicz’s Dante. Journal of the History of Ideas, 82(3):427–452, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/798273.

Hess:2012:CFC

[Hes12] Jillian M. Hess. Coleridge’s fly-catchers: Adapting commonplacebook form. Journal of the History of Ideas, 73(3):463–483, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/481420.

Hesketh:2018:JRS

[Hes18] Ian Hesketh. John Robert Seeley, natural religion, and the Victorian conflict between science and religion. Journal of the History of Ideas, 79(2):309–329, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/693306.

Hetherington:1973:SKM

[Het73] N. S. Hetherington. Sources of Kant’s model of the stellar system. Journal of the History of Ideas, 34(3):461–462, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708965.

Hetherington:1981:PLP

[Het81] Norriss S. Hetherington. Percival Lowell: Professional scientist or interloper? Journal of the History of Ideas, 42(1):159–161, January/ March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709423.

Hetherington:1983:INI

[Het83] Norriss S. Hetherington. Isaac Newton’s influence on Adam Smith’s natural laws in economics. Journal of the History of Ideas, 44(3): 497–505, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709178.

Heyd:1979:RTC

[Hey79] Michael Heyd. From a rationalist theology to Cartesian voluntarism: David derodon and Jean-Robert chouet. Journal of the History of Ideas, 40(4):527–542, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709357.

Heyd:1982:PLH

[Hey82] David Heyd. The place of laughter in Hobbes’s theory of emotions. Journal of the History of Ideas, 43(2):285–295, April/June 1982. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709204.

Heydt:2017:PNR

[Hey17] Colin Heydt. The problem of natural religion in Smith’s moral thought. Journal of the History of Ideas, 78(1):73–94, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/648123.

Herzog:2021:FTA

[HG21] Dagmar Herzog and Stefanos Geroulanos. Fascisms and their afterli(v)es: an introduction. Journal of the History of Ideas, 82(1):73– 83, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
781613.

Havens:1957:BRR

[HH57] George R. Havens and George R. Havens. Book review: A Reply to Dr. Buranelli. Journal of the History of Ideas, 18(1):135–137, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707585.

Hein:1966:CNG

[HH66] Hilde Hein and George E. Hein. The chemistry of noble gases — a modern case history in experimental science. Journal of the History of Ideas, 27(3):417–428, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708594.

Hiatt:2009:DAH

[Hia09] Alfred Hiatt. Diplomatic arts: Hickes against Mabillon in the Republic of Letters. Journal of the History of Ideas, 70(3):351–373, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/270102.

Hickson:2010:MBL

[Hic10] Michael W. Hickson. The message of Bayle’s last title: Providence and toleration in the Entretiens de Maxime et de Thémiste. Journal of the History of Ideas, 71(4):547–567, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/398297.

Hickson:2018:PBS

[Hic18] Michael W. Hickson. Pierre Bayle and the secularization of conscience. Journal of the History of Ideas, 79(2):199–220, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/693301.

Higham:1954:IHN

[Hig54] John Higham. Intellectual history and its neighbors. Journal of the History of Ideas, 15(3):339–347, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707758.

Hightower:1961:ICL

[Hig61] James R. Hightower. Individualism in Chinese literature. Journal of the History of Ideas, 22(2):159–168, April/June 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707830.

Higgins:1986:NM

[Hig86] Kathleen Higgins. Nietzsche on music. Journal of the History of Ideas, 47(4):663–672, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709725.

Higonnet:1991:BRO

[Hig91] Patrice Higonnet. Book review: Orphans of the Enlightenment, or, in the Wake of François Furet. Journal of the History of Ideas, 52 (4):685–691, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709972.

Higham:1995:LRR

[Hig95] John Higham. The limits of relativism: Restatement and remembrance. Journal of the History of Ideas, 56(4):669–674, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709998.

Hill:1966:APP

[Hil66] James J. Hill. The aesthetic principles of the Peri Hupsous. Journal of the History of Ideas, 27(2):265–274, April/June 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708642.

Hilton:1972:BRI

[Hil72] Ronald Hilton. Book review: Is Intellectual History Irrelevant? The Case of the Aztecs. Journal of the History of Ideas, 33(2):337–344, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708879.

Hill:2001:ECA

[Hil01] Lisa Hill. Eighteenth-Century anticipations of the sociology of conflict: The case of Adam Ferguson. Journal of the History of Ideas, 62(2):281–299, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15112.

Hillman:2011:MMW

[Hil11] Susanne Hillman. Men with muskets, women with lyres: Nationality, citizenship, and gender in the writings of Germaine de Staël. Journal of the History of Ideas, 72(2):231–254, April 2011. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/428853.

Hinde:1992:DJB

[Hin92] John R. Hinde. The development of Jacob Burckhardt’s early political thought. Journal of the History of Ideas, 53(3):425–436, July/ September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709885.

Hipolito:2004:CGC

[Hip04] Jeffrey Hipolito. “Conscience the ground of consciousness”: The moral epistemology of Coleridge’s aids to reflection. Journal of the History of Ideas, 65(3):455–474, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/177665.

Hjermitslev:2011:PRD

[Hje11] Hans Henrik Hjermitslev. Protestant responses to Darwinism in Denmark, 1859–1914. Journal of the History of Ideas, 72(2):279– 303, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
428855.

Halewood:1965:HEA

[HL65] William H. Halewood and Marvin Levich. Houyhnhnm est animal rationale. Journal of the History of Ideas, 26(2):273–281, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708233.

Hare:2019:IVO

[HL19] J. Laurence Hare and Fabian Link. The idea of volk and the origins of Völkisch research, 1800–1930s. Journal of the History of Ideas, 80(4):575–596, October 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/735382.

Hampsher-Monk:1979:TCC

[HM79] Iain W. Hampsher-Monk. Tacit concept of consent in Locke’s two treatises of government: a note on citizens, travellers, and patriarchalism. Journal of the History of Ideas, 40(1):135–139, January/ March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709266.

Hobart:1971:HRT

[Hob71] Michael Hobart. History and religion in the thought of Herbert Butterfield. Journal of the History of Ideas, 32(4):543–554, October/ December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708976.

Hoberman:1980:HAP

[Hob80] Louisa S. Hoberman. Hispanic American political theory as a distinct tradition. Journal of the History of Ideas, 41(2):199–218, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709456.

Hoekstra:2015:CEH

[Hoe15] Kinch Hoekstra. The Clarendon edition of Hobbes’s Leviathan: Leviathan and its intellectual context. Journal of the History of Ideas, 76(2):237–257, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/ 56/article/580297. See response [Mal15].

Hofstadter:1941:PJT

[Hof41] Richard Hofstadter. Parrington and the Jeffersonian tradition. Journal of the History of Ideas, 2(4):391–400, October 1941. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707018.

Hofstadter:1954:C

[Hof54] Richard Hofstadter. Communication. Journal of the History of Ideas, 15(2):328–329, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707778.

Hoffheimer:1985:IST

[Hof85] Michael H. Hoffheimer. The influence of Schiller’s theory of nature on Hegel’s philosophical development. Journal of the History of Ideas, 46(2):231–244, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709636.

Holt:1940:ISH

[Hol40] W. Stull Holt. The idea of scientific history in America. Journal of the History of Ideas, 1(3):352–362, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707092.

Holborn:1948:HH

[Hol48] Hajo Holborn. History and the humanities. Journal of the History of Ideas, 9(1):65–69, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707395.

Holborn:1949:GMC

[Hol49] Hajo Holborn. Greek and modern concepts of history. Journal of the History of Ideas, 10(1):3–13, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707196.

Holborn:1950:WDC

[Hol50] Hajo Holborn. Wilhelm Dilthey and the critique of historical reason. Journal of the History of Ideas, 11(1):93–118, January 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707454.

Holborn:1953:HSC

[Hol53] Hajo Holborn. History and the study of the classics. Journal of the History of Ideas, 14(1):33–50, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707494.

Holborn:1960:BR

[Hol60] Hajo Holborn. Bismarck’s Realpolitik. Journal of the History of Ideas, 21(1):84–98, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708000.

Holt:1971:EHM

[Hol71] Niles R. Holt. Ernst Haeckel’s monistic religion. Journal of the History of Ideas, 32(2):265–280, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708280.

Holmes:1982:LUB

[Hol82] Stephen Holmes. Liberal uses of Bourbon legitimism. Journal of the History of Ideas, 43(2):229–248, April/June 1982. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709201.

Holton:2000:RPE

[Hol00] Gerald James Holton. The rise of postmodernisms and the “End of Science”. Journal of the History of Ideas, 61(2):327–341, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15073. See

errata [Ano00f].

Holley:2017:TPM

[Hol17] Jared Holley. Theory, practice, and modernity: Leo Strauss on Rousseau’s Epicureanism. Journal of the History of Ideas, 78(4):621– 644, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
676149.

Holm:2021:MFA

[Hol21] Sophie Holm. Multilingual foreign affairs: Translation and diplomatic agency in Eighteenth-Century Stockholm. Journal of the History of Ideas, 82(3):469–483, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/798275.

Hook:1968:EM

[Hoo68] Sidney Hook. The Enlightenment and Marxism. Journal of the History of Ideas, 29(1):93–108, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708467.

Hook:1973:MFM

[Hoo73] Sidney Hook. Myth and fact in the Marxist theory of revolution and violence. Journal of the History of Ideas, 34(2):271–280, April/ June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708730.

Horowitz:1974:SSI

[Hor74] Maryanne Cline Horowitz. The stoic synthesis of the idea of natural law in man: Four themes. Journal of the History of Ideas, 35(1):3–16, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708739.

Horsman:1976:ORA

[Hor76] Reginald Horsman. Origins of racial Anglo-Saxonism in Great Britain before 1850. Journal of the History of Ideas, 37(3):387–410, July/ September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708805.

Horowitz:1978:BRP

[Hor78] Maryanne Cline Horowitz. Book review: Paul Oskar Kristeller’s Impact on Renaissance Studies. Journal of the History of Ideas, 39(4): 677–683, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709451.

Horne:1980:PCS

[Hor80] Thomas Horne. Politics in a corrupt society: William Arnall’s defense of Robert walpole. Journal of the History of Ideas, 41(4): 601–614, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709276.

Horton:1982:BKB

[Hor82] Mary Horton. Bacon and “Knowledge Broken”: an answer to Michael Hattaway. Journal of the History of Ideas, 43(3):487–504, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709435.

See [Hat78].

Horner:1997:PPF

[Hor97] Richard Horner. A pragmatist in Paris: Frederic Rauh’s “Task of Dissolution”. Journal of the History of Ideas, 58(2):289–308, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14943.

Hoston:2024:NCD

[Hos24] Germaine A. Hoston. Neo-Confucianism and the development of German idealism. Journal of the History of Ideas, 85(2):257–287, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/926149.

Houghton:1941:HTR

[Hou41] Walter E. Houghton, Jr. The history of trades: Its relation to Seventeenth-Century thought: As seen in Bacon, Petty, Evelyn, and Boyle. Journal of the History of Ideas, 2(1):33–60, January 1941. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707280.

Houghton:1942:EVSa

[Hou42a] Walter E. Houghton, Jr. The English virtuoso in the Seventeenth Century: Part I. Journal of the History of Ideas, 3(1):51–73, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707461.

Houghton:1942:EVSb

[Hou42b] Walter E. Houghton, Jr. The English virtuoso in the Seventeenth Century: Part II. Journal of the History of Ideas, 3(2):190–219, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707177.

Houghton:1952:VAI

[Hou52] Walter E. Houghton. Victorian anti-intellectualism. Journal of the History of Ideas, 13(3):291–313, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707598.

Howe:1967:FOT

[How67] Nancy Howe. A further occurrence of theatrica: 1353. Journal of the History of Ideas, 28(4):603–605, October/December 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708535.

Howard:1999:JBR

[How99] Thomas Albert Howard. Jacob Burckhardt, religion, and the historiography of “Crisis” and “Transition”. Journal of the History of Ideas, 60(1):149–164, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15017.

Howland:2001:TLN

[How01] Douglas Howland. Translating liberty in Nineteenth-Century Japan. Journal of the History of Ideas, 62(1):161–181, January 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15102.

Hofer:2018:RAA

[HP18] O. P. Andrew Hofer and O. P. Alan Piper. Retracing the “Art of Arts and Science of Sciences” from Gregory the Great to Philo of Alexandria. Journal of the History of Ideas, 79(4):507–526, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/707770.

Hsueh:2002:GOT

[Hsu02] Vicki Hsueh. Giving orders: Theory and practice in the fundamental constitutions of Carolina. Journal of the History of Ideas, 63(3):425– 446, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15163.

Hellman:2021:TAG

[HTW21] Lisa Hellman and Birgit Tremml-Werner. Translation in action: Global intellectual history and early modern diplomacy. Journal of the History of Ideas, 82(3):453–467, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/798274.

Hudnut:1956:SSS

[Hud56] William H. Hudnut III. Samuel Stanhope Smith: Enlightened conservative. Journal of the History of Ideas, 17(4):540–552, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707787.

Hughes:1943:SAC

[Hug43] Merritt Y. Hughes. Spenser’s Acrasia and the Circe of the Renaissance. Journal of the History of Ideas, 4(4):381–399, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707165.

Huizinga:1943:HCF

[Hui43] J. Huizinga. History changing form. Journal of the History of Ideas, 4(2):217–223, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2707325. See reply [Kat44].

Hui:2017:MRP

[Hui17] Andrew Hui. The many returns of philology: a state of the field report. Journal of the History of Ideas, 78(1):137–156, January 2017.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/648126.

Humphrey:1972:ISK

[Hum72] Ted B. Humphrey. Internal structure in Kant’s thought. Journal of the History of Ideas, 33(1):43–60, January/March 1972. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709056.

Hunter:1952:MPM

[Hun52] William B. Hunter, Jr. Milton’s power of matter. Journal of the History of Ideas, 13(4):551–562, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707717.

Huntley:1953:STB

[Hun53] Frank Livingstone Huntley. Sir Thomas Browne and the metaphor of the circle. Journal of the History of Ideas, 14(3):353–364, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707806.

Hunter:1960:MIS

[Hun60] William B. Hunter. Milton on the incarnation: Some more heresies. Journal of the History of Ideas, 21(3):349–369, July/September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708142.

Hundert:1972:MHF

[Hun72a] E. J. Hundert. The making of Homo Faber: John Locke between ideology and history. Journal of the History of Ideas, 33(1):3–22, January/ March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709054.

Huntley:1972:DHC

[Hun72b] William B. Huntley. David Hume and Charles Darwin. Journal of the History of Ideas, 33(3):457–470, July/September 1972. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709046.

Hundert:1974:AMH

[Hun74] E. J. Hundert. The achievement motive in Hume’s political economy. Journal of the History of Ideas, 35(1):139–143, January/March 1974.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708749.

Hunting:1978:PBS

[Hun78] Claudine Hunting. The philosophes and black slavery: 1748–1765. Journal of the History of Ideas, 39(3):405–418, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709385.

Hundert:1986:SSD

[Hun86] E. J. Hundert. A satire of self-disclosure: From Hegel through Rameau to the Augustans. Journal of the History of Ideas, 47(2):235–248, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709812.

Hundert:1995:BME

[Hun95] E. J. Hundert. Bernard Mandeville and the Enlightenment’s maxims of modernity. Journal of the History of Ideas, 56(4):577–593, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709993.

Hunter:2000:CTD

[Hun00] Ian Hunter. Christian Thomasius and the desacralization of philosophy. Journal of the History of Ideas, 61(4):595–616, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15094.

Hurlbutt:1956:DHS

[Hur56] R. H. Hurlbutt III. David Hume and scientific theism. Journal of the History of Ideas, 17(4):486–497, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707783.

Hutton:1976:VTH

[Hut76] Patrick H. Hutton. Vico’s theory of history and the French Revolutionary tradition. Journal of the History of Ideas, 37(2):241–256, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708823.

Hutton:1978:TJO

[Hut78] Sarah Hutton. Thomas Jackson, Oxford Platonist, and William Twisse, Aristotelian. Journal of the History of Ideas, 39(4):635– 652, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709447.

Hutton:1987:AMR

[Hut87] Patrick H. Hutton. The art of memory reconceived: From rhetoric to psychoanalysis. Journal of the History of Ideas, 48(3):371–392, July/ September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709758.

Hutton:1992:POT

[Hut92] Patrick H. Hutton. The problem of oral tradition in Vico’s historical scholarship. Journal of the History of Ideas, 53(1):3–23, January/ March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709908.

Henninger-Voss:2002:HNS

[HV02] Mary J. Henninger-Voss. How the “New Science” of cannons shook up the Aristotelian Cosmos. Journal of the History of Ideas, 63(3):371– 397, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15168.

Hyde:1956:SHG

[Hyd56] William J. Hyde. The socialism of H. G. Wells in the Early Twentieth Century. Journal of the History of Ideas, 17(2):217–234, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707743.

Ianziti:2000:LPL

[Ian00] Gary Ianziti. A life in politics: Leonardo Bruni’s Cicero. Journal of the History of Ideas, 61(1):39–58, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15061.

Ianziti:2008:LBM

[Ian08] Gary Ianziti. Leonardo Bruni, the Medici, and the Florentine histories. Journal of the History of Ideas, 69(1):1–22, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/231010.

Iggers:1959:FRA

[Igg59] Georg G. Iggers. Further remarks about early uses of the term “Social Science”. Journal of the History of Ideas, 20(3):433–436, June/ September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708121.

See [Sen58, Bur59].

Iggers:1995:HHM

[Igg95] Georg G. Iggers. Historicism: The history and meaning of the term. Journal of the History of Ideas, 56(1):129–152, January 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710011.

Ilany:2012:DCP

[Ila12] Ofri Ilany. From divine commandment to political act: The Eighteenth-Century polemic on the extermination of the Canaanites. Journal of the History of Ideas, 73(3):437–461, July 2012. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/481419.

Immerwahr:1992:HRR

[Imm92] John Immerwahr. Hume’s revised racism. Journal of the History of Ideas, 53(3):481–486, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709889.

Immerwahr:2009:PUS

[Imm09] Daniel Immerwahr. Polanyi in the United States: Peter Drucker, Karl Polanyi, and the midcentury critique of economic society. Journal of the History of Ideas, 70(3):445–466, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/270106.

Inamura:2020:JML

[Ina20] Kazutaka Inamura. J. S. Mill on liberty, Socratic dialectic, and the logic behind political discourse. Journal of the History of Ideas, 81(2): 257–277, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
753537.

Insole:2011:KTI

[Ins11] Christopher Insole. Kant’s transcendental idealism and Newton’s divine sensorium. Journal of the History of Ideas, 72(3):413–436, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/445504.

Iozzia:2022:BSI

[Ioz22] Daniele Iozzia. A beginner’s success: The impact of Plotinus’s First Treatise among Christians. Journal of the History of Ideas, 83(1):1– 16, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
847818.

Irvine:1947:SFU

[Irv47] William Irvine. Shaw, the fabians, and the utilitarians. Journal of the History of Ideas, 8(2):218–231, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707156.

Isaac:2010:TWT

[Isa10] Joel Isaac. Theorist at work: Talcott Parsons and the Carnegie project on theory, 1949–1951. Journal of the History of Ideas, 71(2):287– 311, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
379775.

Israel:2006:EWE

[Isr06] Jonathan Irvine Israel. Enlightenment! Which Enlightenment? Journal of the History of Ideas, 67(3):523–545, July 2006. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/202876.

Israel:2016:JGP

[Isr16a] Jonathan Israel. J. G. A. Pocock and the “Language of Enlightenment” in his barbarism and religion. Journal of the History of Ideas, 77(1):107–127, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/610563.

Israel:2016:RDR

[Isr16b] Jonathan Israel. Rousseau, Diderot, and the “Radical Enlightenment”: a reply to Helena Rosenblatt and Joanna Stalnaker. Journal of the History of Ideas, 77(4):649–677, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/635633.

Israel:2019:PBC

[Isr19] Jonathan Israel. Pierre Bayle’s correspondence and its significance for the history of ideas. Journal of the History of Ideas, 80(3):479– 500, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
729720.

Itzkin:1978:BCU

[Itz78] Elissa S. Itzkin. Bentham’s Chrestomathia: Utilitarian legacy to English education. Journal of the History of Ideas, 39(2):303–316, April/ June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708782. See

note [Tay82].

Ivanov:2024:CTT

[Iva24] Sergey Ivanov. A calf from a tree-trunk: From a rustic proverb to a standard scholastic argument. Journal of the History of Ideas, 85(1):19–40, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/917114.

Jackson:1968:MEW

[Jac68] Carl T. Jackson. The meeting of East and West: The case of Paul Carus. Journal of the History of Ideas, 29(1):73–92, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708466.

Jacobitti:1975:LCI

[Jac75] Edmund E. Jacobitti. Labriola, Croce, and Italian Marxism (1895– 1910). Journal of the History of Ideas, 36(2):297–318, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708929.

Jack:1976:BRP

[Jac76a] Malcolm Jack. Book review: Progress and Corruption in the Eighteenth Century Mandeville’s “Private Vices, Public Benefits”. Journal of the History of Ideas, 37(2):369–376, April/June 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708832.

Jacob:1976:MSL

[Jac76b] Margaret C. Jacob. Millenarianism and science in the Late Seventeenth Century. Journal of the History of Ideas, 37(2):335–341, April/ June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708829.

Jacob:1977:BCP

[Jac77] J. R. Jacob. Boyle’s circle in the protectorate: Revelation, politics, and the Millennium. Journal of the History of Ideas, 38 (1):131–140, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708845.

Jack:1978:OSN

[Jac78] Malcolm Jack. One state of nature: Mandeville and Rousseau. Journal of the History of Ideas, 39(1):119–124, January/March 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709075.

Jacob:1982:BRS

[Jac82] Margaret Jacob. Book review: Science and Social Passion: The Case of Seventeenth-Century England. Journal of the History of Ideas, 43(2):331–339, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709208.

Jackson:1983:CEE

[Jac83] H. J. Jackson. Coleridge, etymology and etymologic. Journal of the History of Ideas, 44(1):75–88, January/March 1983. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709305.

Jacob:1985:HMP

[Jac85] Alexander Jacob. Henry More’s Psychodia Platonica and its relationship to Marsilio Ficino’s Theologia Platonica. Journal of the History of Ideas, 46(4):503–522, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709542.

Jacquette:1990:ANL

[Jac90] Dale Jacquette. Aesthetics and natural law in Newton’s methodology. Journal of the History of Ideas, 51(4):659–666, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709651.

Jackson:2012:FCG

[Jac12] Ben Jackson. Freedom, the common good, and the rule of law: Lippmann and Hayek on economic planning. Journal of the History of Ideas, 73(1):47–68, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/463136.

Jaffe:1980:CGC

[Jaf80] Kineret S. Jaffe. The concept of genius: Its changing role in Eighteenth-Century French aesthetics. Journal of the History of Ideas, 41(4):579–599, October/December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709275.

Janes:1978:RMW

[Jan78] R. M. Janes. On the reception of Mary Wollstonecraft’s: a vindication of the rights of woman. Journal of the History of Ideas, 39(2):293–302, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708781.

Janacek:2008:VHA

[Jan08] Bruce Janacek. A virtuoso’s history: Antiquarianism and the transmission of knowledge in the alchemical studies of Elias Ashmole. Journal of the History of Ideas, 69(3):395–417, July 2008. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/242278.

Jaynes:1970:PAM

[Jay70] Julian Jaynes. The problem of animate motion in the Seventeenth Century. Journal of the History of Ideas, 31(2):219–234, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708546.

Jenkins:1944:LFT

[Jen44] Iredell Jenkins. The laisser-faire theory of artistic censorship. Journal of the History of Ideas, 5(1):71–90, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707102.

Jennings:2007:DAL

[Jen07a] Jeremy Jennings. The debate about luxury in Eighteenth- and Nineteenth-Century French political thought. Journal of the History of Ideas, 68(1):79–105, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/210688.

Jennings:2007:VC

[Jen07b] Jeremy Jennings. The view from Calais. Journal of the History of Ideas, 68(3):381–387, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/219819.

Jenco:2013:RAV

[Jen13] Leigh Jenco. Revisiting Asian values. Journal of the History of Ideas, 74(2):237–258, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/504786.

Jenkins:2020:PHM

[Jen20] Bill Jenkins. Physiology of the haunted mind: Naturalistic theories of apparitions in Early Nineteenth-Century Scotland. Journal of the History of Ideas, 81(4):577–597, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/772515.

Jensen:2022:RMM

[Jen22] Jacob Jensen. Repurposing Mises: Murray Rothbard and the birth of anarchocapitalism. Journal of the History of Ideas, 83(2):315– 332, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
855169.

Jeremiah:2013:DLL

[Jer13] Edward Jeremiah. The development, logic, and legacy of reflexive concepts in Greek philosophy. Journal of the History of Ideas, 74(4):507– 529, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
523426.

Jessup:1954:ME

[Jes54] Bertram Jessup. The mind of Elia. Journal of the History of Ideas, 15(2):246–259, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707770.

Jeske:1986:CMP

[Jes86] Jeffrey Jeske. Cotton Mather: Physico-theologian. Journal of the History of Ideas, 47(4):583–594, October/December 1986. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709719.

Jezierski:1971:PPJ

[Jez71] John V. Jezierski. Parliament or people: James Wilson and blackstone on the nature and location of sovereignty. Journal of the History of Ideas, 32(1):95–106, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708327.

Jordan:2012:LHI

[JN12] Sara R. Jordan and Cary J. Nederman. The logic of the history of ideas and the study of comparative political theory. Journal of the History of Ideas, 73(4):627–641, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/488352.

Johnson:1940:GCP

[Joh40] Francis R. Johnson. Gresham College: Precursor of the Royal Society. Journal of the History of Ideas, 1(4):413–438, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707123.

Johnson:1942:THI

[Joh42] Francis R. Johnson. Thomas Hood’s inaugural address as mathematical lecturer of the city of London (1588). Journal of the History of Ideas, 3(1):94–106, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707463.

Johnstone:1944:RS

[Joh44] Paul H. Johnstone. The rural Socrates. Journal of the History of Ideas, 5(2):151–175, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707382.

Johnson:1946:WPH

[Joh46] A. H. Johnson. Whitehead’s philosophy of history. Journal of the History of Ideas, 7(2):234–249, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707073.

Johnson:1958:MA

[Joh58a] James William Johnson. The meaning of “Augustan”. Journal of the History of Ideas, 19(4):507–522, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707920.

Johnston:1958:LEM

[Joh58b] Charlotte Johnston. Locke’s examination of Malebranche and John Norris. Journal of the History of Ideas, 19(4):551–558, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707923.

Johnson:1959:BRL

[Joh59a] Francis R. Johnson. Book review: Lynn Thorndike’s History of Magic and Experimental Science. Journal of the History of Ideas, 20(2):282– 283, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707826.

Johnson:1959:SHR

[Joh59b] James William Johnson. The Scythian: His rise and fall. Journal of the History of Ideas, 20(2):250–257, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707822.

Johnson:1960:DAC

[Joh60] J. W. Johnson. “Of differing ages and climes”. Journal of the History of Ideas, 21(4):465–480, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708096.

Johnson:1961:NCB

[Joh61] James W. Johnson. That neo-classical bee. Journal of the History of Ideas, 22(2):262–266, April/June 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707837.

Johnson:1967:TAM

[Joh67a] Harold J. Johnson. Three ancient meanings of matter: Democritus, Plato, and Aristotle. Journal of the History of Ideas, 28(1):3–16, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708477.

Johnston:1967:KMV

[Joh67b] William M. Johnston. Karl Marx’s verse of 1836–1837 as a foreshadowing of his early philosophy. Journal of the History of Ideas, 28(2):259–268, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708421.

Johnson:1971:CIP

[Joh71a] James T. Johnson. The covenant idea and the Puritan view of marriage. Journal of the History of Ideas, 32(1):107–118, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708328.

Johnston:1971:SHP

[Joh71b] William M. Johnston. Syncretist historians of philosophy at Vienna 1860–1930. Journal of the History of Ideas, 32(2):299–305, April/ June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708284.

Johnson:1985:HCS

[Joh85] Curtis Johnson. The Hobbesian conception of sovereignty and Aristotle’s politics. Journal of the History of Ideas, 46(3):327–347, July/ September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709471.

John:1988:WWL

[Joh88] Peter C. John. Wittgenstein’s “Wonderful Life”. Journal of the History of Ideas, 49(3):495–510, July/September 1988. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709489. See erratum [Ano88f].

Johnson:2004:PHE

[Joh04] Christopher Johnson. “Periwigged heralds”: Epistemology and intertextuality in Early American cometography. Journal of the History of Ideas, 65(3):399–419, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/177666.

Johnson:2006:ICP

[Joh06] Carina L. Johnson. Idolatrous cultures and the practice of religion. Journal of the History of Ideas, 67(4):597–621, October 2006. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/204775.

Johnson:2022:FGI

[Joh22] Sarah Johnson. Farewell to the German ideology. Journal of the History of Ideas, 83(1):143–170, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/847824.

Jolley:1978:LLS

[Jol78] Nicholas Jolley. Leibniz on Locke and socinianism. Journal of the History of Ideas, 39(2):233–250, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708777.

Jones:1940:SCN

[Jon40] Richard F. Jones. Science and criticism in the neo-classical age of English literature. Journal of the History of Ideas, 1(4):381–412, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707122.

Jones:1967:NLH

[Jon67] Howard Mumford Jones. The nature of literary history. Journal of the History of Ideas, 28(2):147–160, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708412.

Jones:1973:LID

[Jon73] W. R. Jones. Lollards and images: The defense of religious art in Later Medieval England. Journal of the History of Ideas, 34(1):27–50, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708942.

Jones:1992:JSM

[Jon92a] H. S. Jones. John Stuart Mill as moralist. Journal of the History of Ideas, 53(2):287–308, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709875.

Jones:1992:TTT

[Jon92b] William David Jones. Toward a theory of totalitarianism: Franz Borkenau’s Pareto. Journal of the History of Ideas, 53(3):455– 466, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709887.

Jones:2013:PJH

[Jon13] Meirav Jones. Philo Judaeus and Hugo Grotius’s modern natural law. Journal of the History of Ideas, 74(3):339–359, July 2013. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/511986.

Jorgensen:1952:TVW

[Jor52] Paul A. Jorgensen. Theoretical views of war in Elizabethan England. Journal of the History of Ideas, 13(4):469–481, October 1952. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707713.

Joravsky:1959:SMB

[Jor59] David Joravsky. Soviet Marxism and biology before Lysenko. Journal of the History of Ideas, 20(1):85–104, January 1959. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707968.

Jordan:1985:TTA

[Jor85] Richard Douglas Jordan. Thomas Traherne and the art of meditation. Journal of the History of Ideas, 46(3):381–403, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709474.

Joshua:2006:WAA

[Jos06] Essaka Joshua. Wordsworth amongst the Aristotelians. Journal of the History of Ideas, 67(3):511–522, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/202877.

Joukovskaia:2023:BMB

[Jou23] Anna Joukovskaia. Bureaucracy: The making of a buzzword. Journal of the History of Ideas, 84(4):685–710, October 2023. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/909535.

Joy:1992:ERM

[Joy92] Lynn S. Joy. Epicureanism in Renaissance moral and natural philosophy. Journal of the History of Ideas, 53(4):573–583, October/ December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709938.

Jumonville:2008:CMH

[Jum08] Neil Jumonville. The complexity of moral history: Response to cotkin. Journal of the History of Ideas, 69(2):317–322, April 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/235871.

Jumonville:2012:REL

[Jum12] Neil Jumonville. Review-essay: Learn this forward but understand it backward. Journal of the History of Ideas, 73(1):147–162, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/463143.

Jung:1953:NES

[Jun53] Eva-Maria Jung. On the nature of evangelism in Sixteenth-Century Italy. Journal of the History of Ideas, 14(4):511–527, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707699.

Juretic:1975:DNM

[Jur75] George Juretic. Digger no millenarian: The revolutionizing of Gerrard Winstanley. Journal of the History of Ideas, 36(2):263–280, April/ June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708927.

Jurdjevic:2001:VCE

[Jur01] Mark Jurdjevic. Virtue, commerce, and the enduring Florentine Republican moment: Reintegrating Italy into the Atlantic Republican debate. Journal of the History of Ideas, 62(4):721–743, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15131.

Jurdjevic:2002:MSF

[Jur02] Mark Jurdjevic. Machiavelli’s sketches of Francesco Valori and the reconstruction of Florentine history. Journal of the History of Ideas, 63(2):185–206, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15153.

Jeanneret:1998:VED

[JW98] Michel Jeanneret and Caroline Warman. The vagaries of exemplarity: Distortion or dismissal? Journal of the History of Ideas, 59(4):565–579, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15006.

Kaelber:1996:WLM

[Kae96] Lutz Kaelber. Weber’s lacuna: Medieval religion and the roots of rationalization. Journal of the History of Ideas, 57(3):465–485, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14913.

Kahn:1947:NLQ

[Kah47] Ludwig W. Kahn. Note: Legendary quotations. Journal of the History of Ideas, 8(1):116, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707444. See [Bau46].

Kahn:1961:BRP

[Kah61] Charles H. Kahn. Book review: Plato’s Cretan City. Journal of the History of Ideas, 22(3):418–424, July/September 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708137.

Kahan:1985:TTR

[Kah85] Alan Kahan. Tocqueville’s two revolutions. Journal of the History of Ideas, 46(4):585–596, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709547. See errata [Ano86f].

Kalmar:1987:VLS

[Kal87] Ivan Kalmar. The Völkerpsychologie of Lazarus and Steinthal and the modern concept of culture. Journal of the History of Ideas, 48 (4):671–690, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709693.

Kallendorf:1995:VVP

[Kal95] Craig Kallendorf. From Virgil to Vida: The Poeta Theologus in Italian Renaissance commentary. Journal of the History of Ideas, 56(1):41– 62, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2710006.

Kaminsky:1952:EMG

[Kam52] Jack Kaminsky. The empirical metaphysics of George Henry lewes. Journal of the History of Ideas, 13(3):314–332, June 1952. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707599.

Kammen:1970:MCA

[Kam70] Michael Kammen. The meaning of colonization in American Revolutionary thought. Journal of the History of Ideas, 31(3):337–358, July/ September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708510.

Kann:1945:WMP

[Kan45] Robert A. Kann. Wolfgang Menzel: Pioneer of integral nationalism. Journal of the History of Ideas, 6(2):213–230, April 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707364.

Kang:2013:OEP

[Kan13] Taran Kang. Origin and essence: The problem of history in hannah arendt. Journal of the History of Ideas, 74(1):139–160, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/495245.

Kaplan:1956:SES

[Kap56] Sidney Kaplan. Social engineers as saviors: Effects of World War I on some American liberals. Journal of the History of Ideas, 17(3):347– 369, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707549.

Kaplan:1974:MME

[Kap74] Fred Kaplan. “The mesmeric mania”: The Early Victorians and animal magnetism. Journal of the History of Ideas, 35(4):691–702, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709095.

Kapust:2008:AUP

[Kap08] Daniel Kapust. On the ancient uses of political fear and its modern implications. Journal of the History of Ideas, 69(3):353–373, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/242276.

Kaplan:2023:OCM

[Kap23] Abram Kaplan. Occupy the commonplaces: Machiavelli and the Aristotelian tradition of the topics. Journal of the History of Ideas, 84(1):29–50, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/879539.

Kargon:1965:WRH

[Kar65] Robert Kargon. William Rowan Hamilton and Boscovichean atomism. Journal of the History of Ideas, 26(1):137–140, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708405.

Kargon:1966:THN

[Kar66] Robert Kargon. Thomas Hariot, the Northumberland Circle and early atomism in England. Journal of the History of Ideas, 27 (1):128–136, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708313.

Kargon:1969:MAV

[Kar69] Robert Kargon. Model and analogy in Victorian science: Maxwell’s critique of the French physicists. Journal of the History of Ideas, 30 (3):423–436, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708567.

Karier:1984:SSM

[Kar84] Clarence J. Karier. In search of self in a moral universe: Notes on George Herbert Mead’s functionalist theory of morality. Journal of the History of Ideas, 45(1):153–161, January/March 1984. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709337.

Kashap:1977:BRS

[Kas77] Paul Kashap. Book review: Some Recent Works on Spinoza’s Thought. Journal of the History of Ideas, 38(3):541–548, July/ September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708682.

Kassell:2006:AWL

[Kas06] Lauren Kassell. “All was this land full fill’d of faerie,” or magic and the past in Early Modern England. Journal of the History of Ideas, 67(1):107–122, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/194646.

Katz:1944:RJH

[Kat44] Joseph Katz. A reply to J. Huizinga on the form and function of history. Journal of the History of Ideas, 5(3):369–373, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707356. See [Hui43].

Katz:1954:PG

[Kat54] Joseph Katz. Plotinus and the gnostics. Journal of the History of Ideas, 15(2):289–298, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707773.

Katz:1984:CPS

[Kat84] Ruth Katz. Collective “Problem-Solving” in the history of music: The case of the Camerata. Journal of the History of Ideas, 45 (3):361–377, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709230.

Kates:1989:LRT

[Kat89] Gary Kates. From liberalism to radicalism: Tom Paine’s Rights of Man. Journal of the History of Ideas, 50(4):569–587, October/ December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709798.

Kaufmann:1948:NAS

[Kau48] Walter A. Kaufmann. Nietzsche’s admiration for Socrates. Journal of the History of Ideas, 9(4):472–491, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707219.

Kaufmann:1949:GHI

[Kau49] Walter A. Kaufmann. Goethe and the history of ideas. Journal of the History of Ideas, 10(4):503–516, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707187.

Kaufman:1963:WJL

[Kau63] Marjorie R. Kaufman. William James’s letters to a young pragmatist. Journal of the History of Ideas, 24(3):413–421, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708217.

Kaufmann:2001:AHO

[Kau01] Thomas DaCosta Kaufmann. Antiquarianism, the history of objects, and the history of art before Winckelmann. Journal of the History of Ideas, 62(3):523–541, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15123.

Koyre:1946:LB

[KCR46] Alexandre Koyré and Leonora Cohen-Rosenfield. Louis De bonald. Journal of the History of Ideas, 7(1):56–73, January 1946. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707268.

Konijnendijk:2023:MWR

[KE23] Roel Konijnendijk and Fernando Echeverría. Max Weber, the rise of the polis, and the “Hoplite Revolution” theory. Journal of the History of Ideas, 84(1):103–125, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/879542.

Kearns:1964:MFA

[Kea64] Francis E. Kearns. Margaret Fuller and the abolition movement. Journal of the History of Ideas, 25(1):120–127, January/March 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708090.

Keeton:1947:EMP

[Kee47] Morris T. Keeton. Edmund Montgomery — pioneer of organicism. Journal of the History of Ideas, 8(3):309–341, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707145.

Keevak:1992:DDT

[Kee92] Michael Keevak. Descartes’s dreams and their address for philosophy. Journal of the History of Ideas, 53(3):373–396, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709883.

Keedus:2012:LQP

[Kee12] Liisi Keedus. Liberalism and the question of “The Proud”: Hannah Arendt and Leo Strauss as readers of Hobbes. Journal of the History of Ideas, 73(2):319–341, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/475335/pdf.

Kegel:1958:WCM

[Keg58] Charles H. Kegel. William Cobbett and Malthusianism. Journal of the History of Ideas, 19(3):348–362, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708040.

Keil:1957:SQM

[Kei57] Harry Keil. Scabies and the Queen Mab passage in Romeo and Juliet. Journal of the History of Ideas, 18(3):394–410, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707799.

Keitt:2004:RES

[Kei04] Andrew W. Keitt. Religious enthusiasm, the Spanish Inquisition, and the disenchantment of the world. Journal of the History of Ideas, 65(2):231–250, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/173847.

Keller:1950:ZAI

[Kel50] A. C. Keller. Zilsel, the artisans, and the idea of progress in the Renaissance. Journal of the History of Ideas, 11(2):235–240, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707248.

Kelley:1964:HIF

[Kel64] Donald R. Kelley. Historia Integra: François Baudouin and his conception of history. Journal of the History of Ideas, 25(1):35–57, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708084.

Kelly:1965:LAF

[Kel65] George A. Kelly. Liberalism and aristocracy in the French restoration. Journal of the History of Ideas, 26(4):509–530, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708497.

Kelly:1968:RKH

[Kel68] George Armstrong Kelly. Rousseau, Kant, and history. Journal of the History of Ideas, 29(3):347–364, July/September 1968. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708447.

Kelley:1979:BRF

[Kel79] Donald R. Kelley. Book review: Foundations of Modern Political Thought. Journal of the History of Ideas, 40(4):663–673, October/ December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709365.

Kelly:1980:DIA

[Kel80] Aileen Kelly. The destruction of idols: Alexander Herzen and Francis Bacon. Journal of the History of Ideas, 41(4):635–662, October/ December 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709278.

Kelly:1981:LML

[Kel81] G. A. Kelly. From lèse-majesté to lèse-nation: Treason in Eighteenth-Century France. Journal of the History of Ideas, 42(2):269–286, April/ June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709320.

Kelley:1984:SAE

[Kel84] Donald R. Kelley. The science of anthropology: an essay on the very old Marx. Journal of the History of Ideas, 45(2):245–262, April/ June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709290.

Kelley:1987:HIH

[Kel87] Donald R. Kelley. Horizons of intellectual history: Retrospect, circumspect, prospect. Journal of the History of Ideas, 48(1): 143–169, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709616.

Kelley:1990:WHH

[Kel90] Donald R. Kelley. What is happening to the history of ideas? Journal of the History of Ideas, 51(1):3–25, January/March 1990. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709744.

Kelley:1995:BRP

[Kel95] Donald R. Kelley. Book review: Past Masters. Journal of the History of Ideas, 56(1):153–159, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710012.

Kelly:1996:VPJ

[Kel96] George Armstrong Kelly. Veils: The poetics of John Rawls. Journal of the History of Ideas, 57(2):343–364, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14904.

Kelley:1997:JHH

[Kel97] Donald R. Kelley. J. H. Hexter 1910–1996. Journal of the History of Ideas, 58(2):349–350, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14944.

Kelley:1998:RCP

[Kel98] Donald R. Kelley. Romanian cultural and political identity. Journal of the History of Ideas, 59(4):735–738, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15007.

Kelley:2000:J

[Kel00] Donald R. Kelley. JHI 2000. Journal of the History of Ideas, 61 (1):153–156, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15062.

Kelley:2001:EHI

[Kel01] Donald R. Kelley. Eclecticism and the history of ideas. Journal of the History of Ideas, 62(4):577–592, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15132.

Kelly:2004:CSP

[Kel04] Duncan Kelly. Carl Schmitt’s political theory of representation. Journal of the History of Ideas, 65(1):113–134, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/168333.

Kelley:2005:EE

[Kel05a] Donald R. Kelley. From the Executive Editor. Journal of the History of Ideas, 66(4):475–476, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/191474.

Kelley:2005:IHG

[Kel05b] Donald R. Kelley. Intellectual history in a global age. Journal of the History of Ideas, 66(2):155–167, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/187187.

Kelley:2008:JML

[Kel08] Donald R. Kelley. Joseph M. Levine 1933–2008. Journal of the History of Ideas, 69(3):499–500, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/242283.

Keller:2012:AIG

[Kel12] Vera Keller. Accounting for invention: Guido Pancirolli’s lost and found things and the development of Desiderata. Journal of the History of Ideas, 73(2):223–245, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/475336.

Kemp:2000:TMT

[Kem00] Catherine Kemp. Two meanings of the term “Idea”: Acts and contents in Hume’s treatise. Journal of the History of Ideas, 61(4):675– 690, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15095.

Kennington:1971:FDE

[Ken71] Richard Kennington. The finitude of Descartes’ evil genius. Journal of the History of Ideas, 32(3):441–446, July/September 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708359.

Kennington:1973:RC

[Ken73] Richard Kennington. Reply to Caton. Journal of the History of Ideas, 34(4):641–643, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708895. See [Cat73a].

Kennedy:1979:IDT

[Ken79] Emmet Kennedy. “Ideology” from Destutt De Tracy to Marx. Journal of the History of Ideas, 40(3):353–368, July/September 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709242.

Kennedy:1990:ABP

[Ken90] Rick Kennedy. The alliance between Puritanism and Cartesian logic at Harvard, 1687–1735. Journal of the History of Ideas, 51(4): 549–572, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709646.

Keohane:1974:NAL

[Keo74] Nannerl O. Keohane. Nonconformist absolutism in Louis XIV’s France: Pierre Nicole and Denis Veiras. Journal of the History of Ideas, 35(4):579–596, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709087.

Keohane:1977:RHE

[Keo77] Nannerl O. Keohane. The radical humanism of Étienne De la boétie. Journal of the History of Ideas, 38(1):119–130, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708844.

Kern:2024:AEH

[Ker24] Emily M. Kern. Alternate Edens: History, evolution, and origins in UNESCO’s Cultural and Scientific History of Mankind. Journal of the History of Ideas, 85(1):121–148, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/917118.

Kesselman:1968:FTG

[Kes68] Steven Kesselman. The frontier thesis and the Great Depression. Journal of the History of Ideas, 29(2):253–268, April/June 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708579.

Ketcham:1958:JMN

[Ket58] Ralph L. Ketcham. James Madison and the nature of man. Journal of the History of Ideas, 19(1):62–76, January 1958. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707952.

Kibre:1946:IIR

[Kib46] Pearl Kibre. The intellectual interests reflected in libraries of the Fourteenth and Fifteenth Centuries. Journal of the History of Ideas, 7(3): 257–297, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707402.

Kilgore:1961:NPE

[Kil61] W. J. Kilgore. Notes on the philosophy of education of Andres Bello. Journal of the History of Ideas, 22(4):555–560, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708031.

Killeen:2011:HKP

[Kil11] Kevin Killeen. Hanging up kings: The political Bible in Early Modern England. Journal of the History of Ideas, 72(4):549–570, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/453548.

Kim:2014:WBF

[Kim14] Sungmoon Kim. The way to become a female sage: Im Yunjidang’s Confucian feminism. Journal of the History of Ideas, 75(3):395– 416, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
549328.

Kim:2018:VFR

[Kim18] Minchul Kim. Volney and the French Revolution. Journal of the History of Ideas, 79(2):221–242, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/693302.

Kimmel:2021:EMI

[Kim21] Seth Kimmel. Early Modern Iberia, indexed: Hernando Colón’s Cosmography. Journal of the History of Ideas, 82(1):1–28, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/781610.

Kim:2024:SFT

[Kim24a] Agatha Seo-Hyun Kim. Seeing the future through a kaleidoscope: Uses of morphological type in Michelet’s national and natural histories. Journal of the History of Ideas, 85(3):539–569, July 2024. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/933858.

Kim:2024:SGP

[Kim24b] Minchul Kim. Sophie de Grouchy’s political thought in the Letters on Sympathy (1798). Journal of the History of Ideas, 85(2):237–255, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/926148.

King:1956:APM

[Kin56] Hugh R. King. Aristotle without prima materia. Journal of the History of Ideas, 17(3):370–389, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707550. See reply [Sol58].

Kinna:2000:WMA

[Kin00] Ruth Kinna. William Morris: Art, work, and leisure. Journal of the History of Ideas, 61(3):493–512, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15081.

Kirchner:1950:MMH

[Kir50] Walther Kirchner. Mind, mountain, and history. Journal of the History of Ideas, 11(4):412–447, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707590.

Kirk:1953:BPP

[Kir53] Russell Kirk. Burke and the philosophy of prescription. Journal of the History of Ideas, 14(3):365–380, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707807.

Kirchner:1994:MMH

[Kir94] Walther Kirchner. Mind, mountain, and history revisited. Journal of the History of Ideas, 55(2):303–305, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709901.

Kittrell:1966:LFE

[Kit66] Edward R. Kittrell. “Laissez faire” in English classical economics. Journal of the History of Ideas, 27(4):610–620, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708344.

Kitchen:1973:ATT

[Kit73] Martin Kitchen. August Thalheimer’s theory of Fascism. Journal of the History of Ideas, 34(1):67–78, January/March 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708944.

Kivy:1967:CMA

[Kiv67] Peter Kivy. Child Mozart as an aesthetic symbol. Journal of the History of Ideas, 28(2):249–258, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708420.

Kivy:1970:LFA

[Kiv70] Peter Kivy. “Lectures on the Fine Arts”: an unpublished manuscript of Thomas Reid’s. Journal of the History of Ideas, 31(1):17–32, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708367.

Kjaeldgaard:2002:PWA

[Kja02] Lasse Horne Kjaeldgaard. “The peak on which Abraham stands”: The pregnant moment of Søren Kierkegaard’s fear and trembling. Journal of the History of Ideas, 63(2):303–321, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15154.

Kaufmann:1966:BRB

[KK66] Walter Kaufmann and Walter Kaufmann. Book review: A Brief Reply. Journal of the History of Ideas, 27(2):308–310, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708647.

Klaiber:1971:NCL

[Kla71] Jeffrey L. Klaiber. The non-Communist Left in Latin America. Journal of the History of Ideas, 32(4):607–616, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708982.

Klaiber:1976:PCI

[Kla76] Jeffrey L. Klaiber. The posthumous christianization of the Inca Empire in colonial Peru. Journal of the History of Ideas, 37(3):

507–520, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708812.

Klauber:1993:BPO

[Kla93] Martin I. Klauber. Between Protestant orthodoxy and rationalism: Fundamental articles in the early career of Jean LeClerc. Journal of the History of Ideas, 54(4):611–636, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709826.

Kleingeld:1999:SVC

[Kle99] Pauline Kleingeld. Six varieties of cosmopolitanism in Late Eighteenth-Century Germany. Journal of the History of Ideas, 60(3): 505–524, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15037.

Kliger:1955:EUA

[Kli55] Samuel Kliger. Emerson and the usable Anglo-Saxon past. Journal of the History of Ideas, 16(4):476–493, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707505.

Klinck:1981:VTM

[Kli81] Dennis R. Klinck. Vestigia trinitatis in man and his works in the English Renaissance. Journal of the History of Ideas, 42(1):13–27, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709415.

Klinck:2006:LNC

[Kli06] Dennis R. Klinck. Lord Nottingham and the conscience of equity. Journal of the History of Ideas, 67(1):123–147, January 2006. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/194647.

Klink:2024:LAC

[Kli24] Maarten Klink. Legal analogies in Cicero’s political thought. Journal of the History of Ideas, 85(1):1–17, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/917113.

Kallet-Marx:1994:BRI

[KM94] Lisa Kallet-Marx. Book review: Institutions, Ideology, and Political Consciousness in Ancient Greece: Some Recent Books on Athenian Democracy. Journal of the History of Ideas, 55(2):307–335, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709902.

Kneller:1990:IFG

[Kne90] Jane Kneller. Imaginative freedom and the German Enlightenment. Journal of the History of Ideas, 51(2):217–232, April/June 1990. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709513.

Knight:1964:BRP

[Kni64] Frank H. Knight. Book review: Philosophy and History. Journal of the History of Ideas, 25(4):587–598, October/December 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708189.

Knowlson:1963:ILV

[Kno63] J. R. Knowlson. The ideal languages of Veiras, Foigny, and Tyssot De Patot. Journal of the History of Ideas, 24(2):269–278, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707849.

Knowlson:1965:IGU

[Kno65] James R. Knowlson. The idea of gesture as a universal language in the XVIIth and XVIIIth Centuries. Journal of the History of Ideas, 26 (4):495–508, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708496.

Kristeller:1981:JHR

[KNS^{+}81] Paul Oskar Kristeller, Ernest Nagel, Fritz Stern, James Gutmann, and Richard H. Popkin. John Herman Randall, Jr.: In memoriam. Journal of the History of Ideas, 42(3):489–501, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709189.

Kocher:1950:IGE

[Koc50] Paul H. Kocher. The idea of God in Elizabethan medicine. Journal of the History of Ideas, 11(1):3–29, January 1950. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707449.

Kocher:1957:FBS

[Koc57] Paul H. Kocher. Francis Bacon on the science of jurisprudence. Journal of the History of Ideas, 18(1):3–26, January 1957. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707577.

Kochin:1999:WCI

[Koc99] Michael Shalom Kochin. Weeds: Cultivating the imagination in Medieval Arabic political philosophy. Journal of the History of Ideas, 60(3):399–416, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15038.

Koch:2021:UEA

[Koc21] Julian Johannes Immanuel Koch. The “Urbild” of “Einbildung”: The archetype in the imagination in Eighteenth-Century German aesthetics. Journal of the History of Ideas, 82(4):569–591, October 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/837672.

Koertge:1977:GPA

[Koe77] Noretta Koertge. Galileo and the problem of accidents. Journal of the History of Ideas, 38(3):389–408, July/September 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708671.

Kohn:1940:GCE

[Koh40] Hans Kohn. The genesis and character of English nationalism. Journal of the History of Ideas, 1(1):69–94, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707011.

Kohn:1945:DN

[Koh45] Hans Kohn. Dostoevsky’s nationalism. Journal of the History of Ideas, 6(4):385–414, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707342.

Kohn:1949:PFN

[Koh49] Hans Kohn. The paradox of Fichte’s nationalism. Journal of the History of Ideas, 10(3):319–343, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707040.

Kohn:1951:EGN

[Koh51] Hans Kohn. The eve of German nationalism (1789–1812). Journal of the History of Ideas, 12(2):256–284, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707517.

Kohn:1956:FBB

[Koh56] Hans Kohn. France between Britain and Germany. Journal of the History of Ideas, 17(3):283–299, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707545.

Kohak:1964:GMR

[Koh64a] Erazim Kohak. T. G. Masaryk’s revision of Marxism. Journal of the History of Ideas, 25(4):519–542, October/December 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708184.

Kohn:1964:BRP

[Koh64b] Hans Kohn. Book review: Political Theory and the History of Ideas. Journal of the History of Ideas, 25(2):303–307, April/June 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708019.

Kohn:1969:RAT

[Koh69] Hans Kohn. Review-article: a turning point. Journal of the History of Ideas, 30(2):283–290, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708441.

Kohn:1970:BRM

[Koh70] Hans Kohn. Book review: The Multidimensional Enlightenment. Journal of the History of Ideas, 31(3):465–474, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708519.

Kohler:2017:SDJ

[Koh17] George Y. Kohler. “Scholasticism is a daughter of Judaism”: The discovery of Jewish influence on Medieval Christian thought. Journal of the History of Ideas, 78(3):319–340, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/665518.

Kolb:1991:HA

[Kol91] David Kolb. Heidegger at 100, in America. Journal of the History of Ideas, 52(1):140–151, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709586.

Kontler:1997:WRH

[Kon97] Laszlo Kontler. William Robertson’s history of manners in German, 1770–1795. Journal of the History of Ideas, 58(1):125–144, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14930.

Kooy:1999:RCI

[Koo99] Michael John Kooy. Romanticism and Coleridge’s idea of history. Journal of the History of Ideas, 60(4):717–735, October 1999. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15050.

Korshin:1974:JRD

[Kor74] Paul J. Korshin. Johnson and the Renaissance dictionary. Journal of the History of Ideas, 35(2):300–312, April/June 1974. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708764.

Korsmeyer:1975:RHA

[Kor75] Carolyn Wilker Korsmeyer. Relativism and Hutcheson’s aesthetic theory. Journal of the History of Ideas, 36(2):319–330, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708930.

Kostick:2004:WTL

[Kos04] Conor Kostick. William of Tyre, Livy, and the vocabulary of class. Journal of the History of Ideas, 65(3):353–368, July 2004. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/177667.

Kosicki:2024:CEC

[Kos24] Piotr H. Kosicki. Channeling Erasmus in Communist Poland: Leszek Kołakowski, Vatican II, and the reinvention of “Counter-Reformation”. Journal of the History of Ideas, 85(1):87–120, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/917117.

Kovesi:1976:MEB

[Kov76] Julius Kovesi. Marxist ecclesiology and biblical criticism. Journal of the History of Ideas, 37(1):93–110, January/March 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708711.

Koyre:1943:GP

[Koy43] Alexandre Koyré. Galileo and Plato. Journal of the History of Ideas, 4(4):400–428, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707166.

Koyre:1948:C

[Koy48] Alexandre Koyré. Condorcet. Journal of the History of Ideas, 9(2): 131–152, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707225.

Koyre:1998:PTF

[Koy98] Alexandre Koyre. Present trends of French philosophical thought. Journal of the History of Ideas, 59(3):531–548, July 1998. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14992.

Kristeller:1941:SPR

[KR41] Paul Oskar Kristeller and John Herman Randall, Jr. The study of the philosophies of the Renaissance. Journal of the History of Ideas, 2(4):449–496, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707022.

Koselleck:2006:C

[KR06] Reinhart Koselleck and Michaela Richter. Crisis. Journal of the History of Ideas, 67(2):357–400, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/198076.

Kraushaar:1940:LIP

[Kra40] Otto F. Kraushaar. Lotze’s influence on the pragmatism and practical philosophy of William James. Journal of the History of Ideas, 1(4):439–458, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707124.

Kramer:1997:HNM

[Kra97] Lloyd S. Kramer. Historical narratives and the meaning of nationalism. Journal of the History of Ideas, 58(3):525–545, July 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14953.

Kravitz:2018:MCV

[Kra18] Amit Kravitz. Moral comfort versus tragic downfall:: Kant’s concept of the dynamically sublime and Schelling’s tragic alternative. Journal of the History of Ideas, 79(4):613–634, October 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/707775.

Kress:2000:CMD

[Kre00] Jill M. Kress. Contesting metaphors and the discourse of consciousness in William James. Journal of the History of Ideas, 61(2):263–283, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15074.

Kristeller:1940:TIM

[Kri40] Paul Oskar Kristeller. The theory of immortality in Marsilio Ficino. Journal of the History of Ideas, 1(3):299–319, June 1940. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707089.

Kristeller:1944:FPP

[Kri44] Paul Oskar Kristeller. Ficino and Pomponazzi on the place of man in the universe. Journal of the History of Ideas, 5(2):220–226, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707386.

Kristeller:1946:PSH

[Kri46] Paul Oskar Kristeller. The philosophical significance of the history of thought. Journal of the History of Ideas, 7(3):360–366, June 1946.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707407.

Kristeller:1947:BRB

[Kri47] Paul O. Kristeller. Book review: Bruno’s Trial. Journal of the History of Ideas, 8(2):240, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707159.

Kristeller:1951:BRC

[Kri51a] Paul Oskar Kristeller. Book review: The Counter-Renaissance by Hiram Haydn. Journal of the History of Ideas, 12(3):468–472, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707756.

Kristeller:1951:MSA

[Kri51b] Paul Oskar Kristeller. The modern system of the arts: a study in the history of aesthetics Part I. Journal of the History of Ideas, 12(4):496– 527, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707484.

Kristeller:1952:MSA

[Kri52] Paul Oskar Kristeller. The modern system of the arts: a study in the history of aesthetics (II). Journal of the History of Ideas, 13(1):17–46, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707724.

Krieger:1953:MEH

[Kri53] Leonard Krieger. Marx and Engels as historians. Journal of the History of Ideas, 14(3):381–403, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707808.

Kristeller:1958:BRP

[Kri58] Paul Oskar Kristeller. Book review: From Platonism to Neoplatonism. Journal of the History of Ideas, 19(1):129–133, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707959.

Krieger:1960:HLS

[Kri60] Leonard Krieger. History and law in the Seventeenth Century: Pufendorf. Journal of the History of Ideas, 21(2):198–210, April/ June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708194.

Krieger:1963:IWS

[Kri63] Leonard Krieger. The idea of the welfare state in Europe and the United States. Journal of the History of Ideas, 24(4):553–568, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707984.

Krieger:1964:DPT

[Kri64] Leonard Krieger. The distortions of political theory: The XVIIth Century case. Journal of the History of Ideas, 25(3):323–332, July/ September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707902.

Krieger:1965:KCN

[Kri65] Leonard Krieger. Kant and the crisis of natural law. Journal of the History of Ideas, 26(2):191–210, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708227.

Krieger:1970:BRI

[Kri70] Leonard Krieger. Book review: Intellectual History as Art. Journal of the History of Ideas, 31(2):305–312, April/June 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708554.

Krieger:1973:AIH

[Kri73] Leonard Krieger. The autonomy of intellectual history. Journal of the History of Ideas, 34(4):499–516, October/December 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708884.

Kristeller:1983:CT

[Kri83] Paul Oskar Kristeller. “Creativity” and “Tradition”. Journal of the History of Ideas, 44(1):105–113, January/March 1983. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709307.

Kristeller:1995:CBA

[Kri95] Paul O. Kristeller. Comment on Black Athena. Journal of the History of Ideas, 56(1):125–127, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710010.

Kroeber:1953:DC

[Kro53] A. L. Kroeber. The delimitation of civilizations. Journal of the History of Ideas, 14(2):264–275, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707475.

Kroitor:1960:CDD

[Kro60] Harry P. Kroitor. Cowper, deism, and the divinization of nature. Journal of the History of Ideas, 21(4):511–526, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708099.

Kroll:1984:QLR

[Kro84] Richard W. F. Kroll. The question of Locke’s relation to Gassendi. Journal of the History of Ideas, 45(3):339–359, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709229.

Kruger:1955:HLS

[Kru55] Daniel H. Kruger. Hobson, Lenin, and Schumpeter on imperialism. Journal of the History of Ideas, 16(2):252–259, April 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707667.

Korpiola:2024:IPN

[KS24] Mia Korpiola and Jørn Øyrehagen Sunde. The influence of the principle “Necessitas Non Habet Legem” on Nordic medieval laws on theft. Journal of the History of Ideas, 85(4):681–711, October 2024. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/944582.

Kubrin:1967:NCC

[Kub67] David Kubrin. Newton and the cyclical Cosmos: Providence and the mechanical philosophy. Journal of the History of Ideas, 28(3): 325–346, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708622.

Kuehnemund:1942:GPD

[Kue42] Richard Kuehnemund. German prophets of doom and hope. Journal of the History of Ideas, 3(4):443–457, October 1942. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707317.

Kuhn:1949:DH

[Kuh49] Helmut Kuhn. Dialectic in history. Journal of the History of Ideas, 10(1):14–29, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707197.

Kuhn:1961:GGH

[Kuh61] Albert J. Kuhn. Glory or gravity: Hutchinson vs. Newton. Journal of the History of Ideas, 22(3):303–322, July/September 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708127.

Kukkonen:2000:PPL

[Kuk00] Taneli Kukkonen. Plenitude, possibility, and the limits of reason: a Medieval Arabic debate on the metaphysics of nature. Journal of the History of Ideas, 61(4):539–560, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15096.

Kusukawa:1997:LFI

[Kus97] Sachiko Kusukawa. Leonhart Fuchs on the importance of pictures. Journal of the History of Ideas, 58(3):403–427, July 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14954.

Kvastad:1977:SMH

[Kva77] Nils B. Kvastad. Semantics in the methodology of the history of ideas. Journal of the History of Ideas, 38(1):157–174, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708848.

LaVopa:1993:BRJ

[La 93] Anthony J. La Vopa. Book review: Jews and Germans: Old Quarrels, New Departures. Journal of the History of Ideas, 54(4):675–695, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709829.

Lach:1945:LC

[Lac45] Donald F. Lach. Leibniz and China. Journal of the History of Ideas, 6(4):436–455, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707344.

Lacy:1947:LJA

[Lac47] Edgar W. Lacy. Law and justice in antiquity and the Middle Ages. Journal of the History of Ideas, 8(3):373–374, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707148.

Lach:1953:SCW

[Lac53] Donald F. Lach. The Sinophilism of Christian Wolff (1679–1754). Journal of the History of Ideas, 14(4):561–574, October 1953. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707702.

Lacey:1965:EAS

[Lac65] A. R. Lacey. The eleatics and Aristotle on some problems of change. Journal of the History of Ideas, 26(4):451–468, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708493.

LaCasce:1970:SME

[LaC70] Steward LaCasce. Swift on medical extremism. Journal of the History of Ideas, 31(4):599–606, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708263.

LaCapra:1988:BRR

[LaC88] Dominick LaCapra. Book review: A Review of a Review. Journal of the History of Ideas, 49(4):677–687, October/December 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709680. See [Pag88].

Lackey:1999:KGL

[Lac99] Michael Lackey. Killing God, liberating the “Subject”: Nietzsche and post-God freedom. Journal of the History of Ideas, 60(4):737–754, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15051.

Lacy:2010:LRA

[Lac10] Tim Lacy. The Lovejovian roots of Adler’s philosophy of history: Authority, democracy, irony, and paradox in Britannica’s Great Books of the Western World. Journal of the History of Ideas, 71(1):113– 137, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
370102.

Ladd:1962:HDM

[Lad62] Everett C. Ladd, Jr. Helvetius and D’Holbach: “La Moralisation de la Politique”. Journal of the History of Ideas, 23(2):221–238, April/ June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708156.

Laerke:2014:LEN

[Lær14] Mogens Lærke. Leibniz, the encyclopedia, and the natural order of thinking. Journal of the History of Ideas, 75(2):237–259, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/544560.

Lahey:1997:WR

[Lah97] Stephen E. Lahey. Wyclif on rights. Journal of the History of Ideas, 58(1):1–20, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14931.

Lalande:1954:SHL

[Lal54] André Lalande. From science and hypothesis to last thoughts of H. Poincaré (1854–1912). Journal of the History of Ideas, 15(4):596–598, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707678.

Lamb:1956:RCC

[Lam56] Ursula Lamb. Religious conflicts in the conquest of Mexico. Journal of the History of Ideas, 17(4):526–539, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707786.

Lamb:1974:ASS

[Lam74] Robert Boyden Lamb. Adam Smith’s system: Sympathy not selfinterest. Journal of the History of Ideas, 35(4):671–682, October/ December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709093.

Lammi:1991:HGG

[Lam91] Walter Lammi. Hans-Georg Gadamer’s “Correction” of Heidegger. Journal of the History of Ideas, 52(3):487–507, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710048.

Lamb:2009:WWG

[Lam09] Robert Lamb. Was William Godwin a utilitarian? Journal of the History of Ideas, 70(1):119–141, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/258197.

Lang:1964:IB

[Lan64] Berel Lang. Intuition in Bloomsbury. Journal of the History of Ideas, 25(2):295–302, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708018.

Langmuir:1966:MHP

[Lan66] Gavin I. Langmuir. Majority history and post-biblical Jews. Journal of the History of Ideas, 27(3):343–364, July/September 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708590.

Land:1974:URP

[Lan74] Stephen K. Land. Universalism and relativism: a philosophical problem of translation in the Eighteenth Century. Journal of the History of Ideas, 35(4):597–610, October/December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709088.

Land:1976:LMT

[Lan76] Stephen K. Land. Lord Monboddo and the theory of syntax in the Late Eighteenth Century. Journal of the History of Ideas, 37 (3):423–440, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708807.

Land:1977:ASC

[Lan77] Stephen K. Land. Adam Smith’s “Considerations Concerning the First Formation of Languages”. Journal of the History of Ideas, 38 (4):677–690, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708696.

Lanyi:1980:DDR

[Lan80] Gabriel Lanyi. Debates on the definition of romanticism in literary France (1820–30). Journal of the History of Ideas, 41(1):141–150, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709108.

Lang:1995:API

[Lan95] Helen S. Lang. Aristotle’s Physics IV, 8: a vexed argument in the history of ideas. Journal of the History of Ideas, 56(3):353–376, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2710031.

Lang:2002:LAI

[Lan02] Timothy Lang. Lord Acton and “the insanity of nationality”. Journal of the History of Ideas, 63(1):129–149, January 2002. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/15144.

Lanzoni:2009:SM

[Lan09] Susan Lanzoni. Sympathy in mind (1876–1900). Journal of the History of Ideas, 70(2):265–287, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/261979.

Lane:2012:DOO

[Lan12a] Melissa Lane. Doing our own thinking for ourselves: On Quentin Skinner’s genealogical turn. Journal of the History of Ideas, 73(1):71– 82, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/

463138. See [Ski12].

Lane:2012:OSD

[Lan12b] Melissa Lane. The origins of the statesman–demagogue distinction in and after Ancient Athens. Journal of the History of Ideas, 73(2):179– 200, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
475337.

Larrabee:1941:HCG

[Lar41] Stephen A. Larrabee. Hazlitt’s criticism and Greek sculpture. Journal of the History of Ideas, 2(1):77–94, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707282.

Larsen:1959:MCP

[Lar59] Robert E. Larsen. Morris Cohen’s principle of polarity. Journal of the History of Ideas, 20(4):587–595, October/December 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707894.

Larsen:1962:ABN

[Lar62] Robert E. Larsen. The Aristotelianism of Bacon’s Novum Organum. Journal of the History of Ideas, 23(4):435–450, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708168.

Larson:1967:GL

[Lar67] James L. Larson. Goethe and linnæus. Journal of the History of Ideas, 28(4):590–596, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708532.

Large:1978:BRA

[Lar78] David C. Large. Book review: Art, Ideology, and Politics at Bayreuth, 1876–1976. Journal of the History of Ideas, 39(1):149–156, January/ March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709080.

Larabee:2010:BCG

[Lar10] Mark D. Larabee. Baedekers as casualty: Great War nationalism and the fate of travel writing. Journal of the History of Ideas, 71(3):457– 480, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
386153.

Laserson:1947:DRI

[Las47] Max M. Laserson. Democracy as a regulative idea and as an established regime: The democratic tradition in Russia and Germany. Journal of the History of Ideas, 8(3):342–362, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707146.

Lataner:1957:IDA

[Lat57] Albert Lataner. Introduction to Davidson’s autobiographical sketch. Journal of the History of Ideas, 18(4):529–536, October 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707563.

Latzer:1994:LCM

[Lat94] Michael Latzer. Leibniz’s conception of metaphysical evil. Journal of the History of Ideas, 55(1):1–15, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709950.

Latteur:2024:ALT

[Lat24] Olivier Latteur. Antiquarianism, local traditions, and urban identity in the Early Modern Netherlands: The controversy about the City of the Nervii. Journal of the History of Ideas, 85(3):455–477, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/933855.

Lauterbach:1944:MWW

[Lau44] Albert T. Lauterbach. Militarism in the Western World: a comparative study. Journal of the History of Ideas, 5(4):446–478, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707081.

Laudan:1967:NSL

[Lau67] Laurens Laudan. The nature and sources of Locke’s views on hypotheses. Journal of the History of Ideas, 28(2):211–223, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708416.

Lauxtermann:1990:HSP

[Lau90] P. F. H. Lauxtermann. Hegel and Schopenhauer as partisans of Goethe’s theory of color. Journal of the History of Ideas, 51(4): 599–624, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709648.

Laursen:1998:DHD

[Lau98] John Christian Laursen. David Hume and the Danish debate about freedom of the press in the 1770s. Journal of the History of Ideas, 59(1):167–172, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14972.

Laursen:2000:SDF

[Lau00] John Christian Laursen. Spinoza in Denmark and the fall of Struensee, 1770–1772. Journal of the History of Ideas, 61(2):189–202, April 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15075.

Lazier:2003:OGH

[Laz03] Benjamin Lazier. Overcoming gnosticism: Hans Jonas, Hans Blumenberg, and the legitimacy of the natural world. Journal of the History of Ideas, 64(4):619–637, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/52450.

LeVanBaumer:1945:CCR

[Le 45] Franklin Le Van Baumer. The conception of Christendom in Renaissance England. Journal of the History of Ideas, 6(2):131–156, April 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707361.

Leary:1977:BST

[Lea77] David E. Leary. Berkeley’s social theory: Context and development. Journal of the History of Ideas, 38(4):635–649, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708692.

Leach:1978:PCR

[Lea78] Eleanor Winsor Leach. Parthenian caverns: Remapping of an imaginative topography. Journal of the History of Ideas, 39(4): 539–560, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709441.

Leaman:1980:AAJ

[Lea80] Oliver Leaman. ’Abd Al-Jabbar and the concept of uselessness. Journal of the History of Ideas, 41(1):129–131, January/March 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709106.

Leary:1982:FIJ

[Lea82] David E. Leary. The fate and influence of John Stuart Mill’s proposed science of ethology. Journal of the History of Ideas, 43(1): 153–162, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709167.

Lebovics:1986:UAL

[Leb86] Herman Lebovics. The uses of America in Locke’s second treatise of government. Journal of the History of Ideas, 47(4):567–581, October/ December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709718.

Lees:1979:CUS

[Lee79] Andrew Lees. Critics of urban society in Germany, 1854–1914. Journal of the History of Ideas, 40(1):61–83, January/March 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709260.

Leerssen:1995:WWI

[Lee95] Joep Leerssen. Wildness, wilderness, and Ireland: Medieval and earlymodern patterns in the demarcation of civility. Journal of the History of Ideas, 56(1):25–39, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710005.

Lee:2009:RLH

[Lee09] Alexander Lee. Roman law and human liberty: Marsilius of Padua on property rights. Journal of the History of Ideas, 70(1):23–44, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/258193.

Lee:2011:PLP

[Lee11] Daniel Lee. Popular liberty, princely government, and the Roman law in Hugo Grotius’s De Jure Belli ac Pacis. Journal of the History of Ideas, 72(3):371–392, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/445503.

Leffler:1976:HRP

[Lef76] Phyllis K. Leffler. The “Histoire Raisonnée,” 1660–1720: a pre-Enlightenment genre. Journal of the History of Ideas, 37(2):219–240, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708822.

Legg:2025:SCI

[Leg25] Charlotte Ann Legg. The settler colonial ideal in Nineteenth-Century France: From revolutionary shipwreck to settler colonial shores. Journal of the History of Ideas, 86(1):109–139, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/949929.

LeMahieu:1979:MTS

[LeM79] D. L. LeMahieu. Malthus and the theology of scarcity. Journal of the History of Ideas, 40(3):467–474, July/September 1979. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709249.

Lenz:1958:HDC

[Len58] John W. Lenz. Hume’s defense of causal inference. Journal of the History of Ideas, 19(4):559–567, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707924.

Lennon:1977:JCP

[Len77] Thomas M. Lennon. Jansenism and the Crise Pyrrhonienne. Journal of the History of Ideas, 38(2):297–306, April/June 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708913.

Lennon:1997:BAP

[Len97] Thomas M. Lennon. Bayle’s anticipation of Popper. Journal of the History of Ideas, 58(4):695–705, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14965.

Lennon:2002:DBR

[Len02] Thomas M. Lennon. Did Bayle read Saint-Evremond? Journal of the History of Ideas, 63(2):225–237, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15155. See reply [Mor04a].

Lennon:2004:RM

[Len04] Thomas M. Lennon. A rejoinder to Mori. Journal of the History of Ideas, 65(2):335–341, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/ 1/article/173848. See [Len02, Len04].

Lester:1957:FNJ

[Les57] John A. Lester, Jr. Friedrich Nietzsche and John Davidson: a study in influence. Journal of the History of Ideas, 18(3):411–429, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707800.

Leslie:1972:MMD

[Les72] Margaret Leslie. Mysticism misunderstood: David Hartley and the idea of progress. Journal of the History of Ideas, 33(4):625–632, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708861.

Lestition:2007:CTN

[Les07] Steven O. Lestition. Countering, transposing, or negating the Enlightenment? A response to Robert Norton. Journal of the History of Ideas, 68(4):659–681, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/223128. See [Nor07].

Levinson:1947:NOJ

[Lev47a] Ronald B. Levinson. A note on one of James’s favorite metaphors. Journal of the History of Ideas, 8(2):237–239, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707158.

Levinson:1947:SLW

[Lev47b] Ronald B. Levinson. Sigwart’s logik and William James. Journal of the History of Ideas, 8(4):475–483, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707431.

Levenson:1950:BCL

[Lev50] Joseph R. Levenson. The breakdown of Confucianism: Liang Ch’I-Ch’Ao before exile — 1873–1898. Journal of the History of Ideas, 11(4):448–485, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707591.

Levy:1954:EVT

[Lev54] Aaron Levy. Economic views of Thomas Hobbes. Journal of the History of Ideas, 15(4):589–595, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707677.

Levi:1956:ISP

[Lev56] Albert William Levi. The idea of Socrates: The philosophic hero in the Nineteenth Century. Journal of the History of Ideas, 17(1):89–108, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707687.

Levin:1965:TSN

[Lev65] Harry Levin. Toward a sociology of the novel. Journal of the History of Ideas, 26(1):148–154, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708407.

Levin:1966:MIP

[Lev66] Samuel M. Levin. Malthus and the idea of progress. Journal of the History of Ideas, 27(1):92–108, January/March 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708310. See erratum [Ano66f].

Levin:1967:USC

[Lev67] Michael Levin. Uses of the social contract method: Vaughan’s interpretation of Rousseau. Journal of the History of Ideas, 28(4): 521–536, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708527.

Levin:1971:BRE

[Lev71] Harry Levin. Book review: Evangelizing Shakespeare. Journal of the History of Ideas, 32(2):306–310, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708285.

Levin:1972:BRL

[Lev72a] Harry Levin. Book review: A Literary Enormity: Sartre on Flaubert. Journal of the History of Ideas, 33(4):643–649, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708864.

Levine:1972:HE

[Lev72b] Norman Levine. Humanism without eschatology. Journal of the History of Ideas, 33(2):281–298, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708874.

Levy:1978:ASN

[Lev78] David Levy. Adam Smith’s “Natural law” and contractual society. Journal of the History of Ideas, 39(4):665–674, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709449.

Levy:1979:DLP

[Lev79] Donald Levy. The definition of love in Plato’s symposium. Journal of the History of Ideas, 40(2):285–291, April/June 1979. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709153.

Levin:1987:BRS

[Lev87a] Harry Levin. Book review: Santayana. Journal of the History of Ideas, 48(4):719–722, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709696.

Levine:1987:GHS

[Lev87b] Norman Levine. The German historical school of law and the origins of historical materialism. Journal of the History of Ideas, 48 (3):431–451, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709761.

Levine:1989:BMP

[Lev89] Joseph M. Levine. Bentley’s Milton: Philology and criticism in Eighteenth-Century England. Journal of the History of Ideas, 50(4): 549–568, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709797.

Levine:1991:GVQ

[Lev91] Joseph M. Levine. Giambattista Vico and the quarrel between the ancients and the moderns. Journal of the History of Ideas, 52 (1):55–79, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709582.

Levine:1997:EPJ

[Lev97] Joseph M. Levine. Erasmus and the problem of the Johannine comma. Journal of the History of Ideas, 58(4):573–596, October 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14966.

Levine:2003:MFE

[Lev03] Joseph M. Levine. Matter of fact in the English Revolution. Journal of the History of Ideas, 64(2):317–335, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/43510.

Levine:2005:IHH

[Lev05] Joseph M. Levine. Intellectual history as history. Journal of the History of Ideas, 66(2):189–200, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/187188.

Levine:2013:OWW

[Lev13] Emily J. Levine. The other Weimar: The Warburg Circle as Hamburg School. Journal of the History of Ideas, 74(2):307–330, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/504790.

Lewitter:1958:PGP

[Lew58] L. R. Lewitter. Peter the Great, Poland, and the Westernization of Russia. Journal of the History of Ideas, 19(4):493–506, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707919.

Lewis:1965:OME

[Lew65] Merrill Lewis. Organic metaphor and edenic myth in George Bancroft’s history of the United States. Journal of the History of Ideas, 26(4):587–592, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708502.

Lewisohn:1972:MCM

[Lew72] David Lewisohn. Mill and Comte on the methods of social science. Journal of the History of Ideas, 33(2):315–324, April/June 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708877.

Lewis:2001:WCE

[Lew01] Eric Lewis. Walter Charleton and early modern eclecticism. Journal of the History of Ideas, 62(4):651–664, October 2001. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/15133.

Lichtheim:1973:RPS

[Lic73] George Lichtheim. Reply to Professor Shue. Journal of the History of Ideas, 34(4):651–652, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708898. See [Shu73].

Lidtke:1966:ABG

[Lid66] Vernon L. Lidtke. August Bebel and German social democracy’s relation to the Christian churches. Journal of the History of Ideas, 27(2):245–264, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708641.

Lievestro:1956:TSA

[Lie56] Christiaan T. Lievestro. Tertullian and the sensus argument. Journal of the History of Ideas, 17(2):264–268, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707746.

Lieberman:1998:MFA

[Lie98] Benjamin David Lieberman. The meanings and function of antisystem ideology in the Weimar Republic. Journal of the History of Ideas, 59(2):355–375, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14983.

Lifschitz:2012:ALS

[Lif12] Avi Lifschitz. The arbitrariness of the linguistic sign: Variations on an Enlightenment theme. Journal of the History of Ideas, 73(4):537– 557, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
488347.

Link:1948:STC

[Lin48] Arthur S. Link. Samuel Taylor Coleridge and the economic and political crisis in Great Britain, 1816–1820. Journal of the History of Ideas, 9(3):323–338, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707373.

Linden:1974:FBA

[Lin74] Stanton J. Linden. Francis Bacon and alchemy: The reformation of Vulcan. Journal of the History of Ideas, 35(4):547–560, October/ December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709085.

Linden:1991:GSS

[Lin91] Stanton J. Linden. “By gradual scale sublim’d”: Jean d’Espagnet and the ontological tree in Paradise Lost, book V. Journal of the History of Ideas, 52(4):603–615, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709968.

Linstrum:2010:SLJ

[Lin10] Erik Linstrum. Strauss’s life of Jesus: Publication and the politics of the German public sphere. Journal of the History of Ideas, 71(4):593– 616, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
398299.

Liu:2004:THC

[Liu04] Haiming Liu. Transnational historiography: Chinese American studies reconsidered. Journal of the History of Ideas, 65(1):135–153, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/168334.

Liu:2012:RFC

[Liu12] Yu Liu. The religiosity of a former Confucian–Buddhist: The Catholic faith of Yang Tingyun. Journal of the History of Ideas, 73(1):25– 46, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
463135.

Livingston:2008:RSM

[Liv08] James Livingston. The return of the self-made man: Response to cotkin. Journal of the History of Ideas, 69(2):327–331, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/235872.

Lizzul:2022:LFP

[Liz22] Giorgio Lizzul. Liberality as a fiscal problem in medieval and Renaissance thought: a genealogy from Aristotle’s tyrant to Machiavelli’s Prince. Journal of the History of Ideas, 83(3):363–385, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/858854.

Lloyd:1969:APD

[Llo69] Charles E. Lloyd. The author of peace and Donne’s holy sonnet XIV. Journal of the History of Ideas, 30(2):251–252, April/June 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708436.

Lo:2017:PMF

[Lo17] Melissa Lo. The picture multiple: Figuring, thinking, and knowing in Descartes’s Essais (1637). Journal of the History of Ideas, 78(3): 369–399, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665520.

Lockey:2004:CEL

[Loc04] Brian Lockey. Conquest and English legal identity in Renaissance Ireland. Journal of the History of Ideas, 65(4):543–558, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/184068.

Lochman:2007:DDA

[Loc07] Daniel T. Lochman. Divus Dionysius: Authority, self, and society in John Colet’s reading of the ecclesiastical hierarchy2. Journal of the History of Ideas, 68(1):1–34, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/210689.

Loemker:1946:LJF

[Loe46] Leroy E. Loemker. Leibniz’s judgments of fact. Journal of the History of Ideas, 7(4):397–410, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707031.

Loemker:1947:NOP

[Loe47] Leroy E. Loemker. A note on the origin and problem of Leibniz’s discourse of 1686. Journal of the History of Ideas, 8(4):449–466, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707429.

Loemker:1955:BL

[Loe55] Leroy E. Loemker. Boyle and Leibniz. Journal of the History of Ideas, 16(1):22–43, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707525.

Loehr:1961:QIC

[Loe61a] Max Loehr. The question of individualism in Chinese art. Journal of the History of Ideas, 22(2):147–158, April/June 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707829.

Loemker:1961:LHE

[Loe61b] Leroy E. Loemker. Leibniz and the herborn encyclopedists. Journal of the History of Ideas, 22(3):323–338, July/September 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708128.

Loewenberg:1970:HZT

[Loe70] Peter Loewenberg. A hidden Zionist theme in Freud’s “My Son, the Myops … ” dream. Journal of the History of Ideas, 31(1): 129–132, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708375.

Logan:1975:RMR

[Log75] George M. Logan. The relation of Montaigne to Renaissance humanism. Journal of the History of Ideas, 36(4):613–632, October/ December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708992.

Lomonaco:2002:ASL

[Lom02] Jeffrey Lomonaco. Adam Smith’s “Letter to the Authors of the Edinburgh Review”. Journal of the History of Ideas, 63(4):659–676, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/39295.

Loubere:1956:LBP

[Lou56] Leo A. Loubère. Louis Blanc’s philosophy of history. Journal of the History of Ideas, 17(1):70–88, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707686.

Lovejoy:1940:RHI

[Lov40] Arthur O. Lovejoy. Reflections on the history of ideas. Journal of the History of Ideas, 1(1):3–23, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707007.

Lovejoy:1941:MRH

[Lov41] Arthur O. Lovejoy. The meaning of romanticism for the historian of ideas. Journal of the History of Ideas, 2(3):257–278, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707131.

Lovejoy:1942:CSA

[Lov42] Arthur O. Lovejoy. The Communism of Saint Ambrose. Journal of the History of Ideas, 3(4):458–468, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707318.

Lovejoy:1944:RPS

[Lov44] Arthur O. Lovejoy. Reply to Professor Spitzer. Journal of the History of Ideas, 5(2):204–219, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707385. See [Spi44].

Lovejoy:1946:GCB

[Lov46a] Arthur O. Lovejoy. Goldsmith and the chain of being. Journal of the History of Ideas, 7(1):91–98, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707270.

Lovejoy:1946:NPE

[Lov46b] Arthur O. Lovejoy. A note on Peirce’s evolutionism. Journal of the History of Ideas, 7(3):351–354, June 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707405.

Lovejoy:1949:HED

[Lov49] Arthur O. Lovejoy. Historiography and evaluation: a disclaimer. Journal of the History of Ideas, 10(1):141–142, January 1949. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707206.

Lovin:1967:BBI

[Lov67] Clifford R. Lovin. Blut und Boden: The ideological basis of the Nazi agricultural program. Journal of the History of Ideas, 28(2):279–288, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708423.

Lowith:1945:NDE

[Löw45] Karl Löwith. Nietzsche’s doctrine of eternal recurrence. Journal of the History of Ideas, 6(3):273–284, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707291.

Lowe:1949:IBJ

[Low49] Victor Lowe. The influence of Bergson, James and Alexander on Whitehead. Journal of the History of Ideas, 10(2):267–296, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707418.

Lowinsky:1954:MCR

[Low54] Edward E. Lowinsky. Music in the culture of the Renaissance. Journal of the History of Ideas, 15(4):509–553, October 1954. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707674.

Lowe:1961:JDM

[Low61] Irving Lowe. John Donne: The middle way. Journal of the History of Ideas, 22(3):389–397, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708134.

Lowry:1974:ACC

[Low74] S. Todd Lowry. The archaeology of the circulation concept in economic theory. Journal of the History of Ideas, 35(3):429–444, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708792.

Lowe:2024:SDG

[Low24] Ruby Lowe. The speech without doors: a genre, 1627–1769. Journal of the History of Ideas, 85(2):209–235, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/926147.

Luft:1991:DSI

[LS91] Eric V. D. Luft and Douglas G. Stenberg. Dostoevskii’s specific influence on Nietzsche’s preface to daybreak. Journal of the History of Ideas, 52(3):441–461, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710046.

Leitao:2017:ZTM

[LS17] Henrique Leitão and Antonio Sánchez. Zilsel’s thesis, maritime culture, and Iberian science in Early Modern Europe. Journal of the History of Ideas, 78(2):191–210, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/652822.

Lunbeck:2023:WAT

[LS23] Emma Lunbeck and Robert Stone. The wise adviser trap: Catastrophic decision-making in Herodotus and Thucydides. Journal of the History of Ideas, 84(3):417–439, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/901488.

Luban:2024:PSS

[Lub24] Daniel Luban. Prisoner, sailor, soldier, spy: Hobbes on coercion and consent. Journal of the History of Ideas, 85(2):185–208, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/926146.

Lukes:1971:MI

[Luk71] Steven Lukes. The meanings of “Individualism”. Journal of the History of Ideas, 32(1):45–66, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708324.

Lumiansky:1941:BEN

[Lum41] R. M. Lumiansky. Beginnings of English nationalism. Journal of the History of Ideas, 2(2):248–249, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707116.

Lunstead:1977:BRR

[Lun77] Jeffrey J. Lunstead. Book review: Recent Books Which Relate Indian and Western Philosophy. Journal of the History of Ideas, 38(4): 719–725, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708701.

Lundbaek:1983:INC

[Lun83] Knud Lundbaek. The image of neo-Confucianism in Confucius Sinarum Philosophus. Journal of the History of Ideas, 44(1):19– 30, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709302.

Lund:1987:TES

[Lun87] William R. Lund. Tragedy and education in the state of nature: Hobbes on time and the will. Journal of the History of Ideas, 48 (3):393–410, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709759.

Lund:2004:WJM

[Lun04] Roger D. Lund. Wit, judgment, and the misprisions of similitude. Journal of the History of Ideas, 65(1):53–75, January 2004. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/168335.

Lundberg:2021:MPC

[Lun21] Christa Lundberg. The making of a philosopher: The contemplative letters of Charles de bovelles. Journal of the History of Ideas, 82(2): 185–205, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
790200.

Lutz:1970:NLR

[Lut70] Rolland Ray Lutz, Jr. The “New Left” of Restoration Germany. Journal of the History of Ideas, 31(2):235–252, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708547.

Lynskey:1945:GCB

[Lyn45] Winifred Lynskey. Goldsmith and the chain of being. Journal of the History of Ideas, 6(3):363–374, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707298.

Lynch:1977:MEC

[Lyn77] Andrew J. Lynch. Montesquieu and the ecclesiastical critics of L’Esprit Des Lois. Journal of the History of Ideas, 38(3):487– 500, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708678.

Lynes:1982:DTE

[Lyn82] John W. Lynes. Descartes’ theory of elements: From Le Monde to the Principes. Journal of the History of Ideas, 43(1):55–72, January/ March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709160.

Lynch:2000:BTA

[Lyn00] Jack Lynch. Betwixt two ages cast: Milton, Johnson, and the English Renaissance. Journal of the History of Ideas, 61(3):397–413, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15082.

Lyon:1970:NHP

[Lyo70] Robert Lyon. Notes on Hume’s philosophy of political economy. Journal of the History of Ideas, 31(3):457–461, July/September 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708517.

Lyon:2003:BFP

[Lyo03] Gregory B. Lyon. Baudouin, Flacius, and the plan for the Magdeburg centuries. Journal of the History of Ideas, 64(2):253–272, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/43511.

MacLean:1949:ISS

[Mac49] Kenneth MacLean. Imagination and sympathy: Sterne and Adam Smith. Journal of the History of Ideas, 10(3):399–410, June 1949.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707044.

Mack:1955:FU

[Mac55] Mary Peter Mack. The fabians and utilitarianism. Journal of the History of Ideas, 16(1):76–88, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707528.

Maclear:1956:PAP

[Mac56] James Fulton Maclear. Popular anticlericalism in the Puritan Revolution. Journal of the History of Ideas, 17(4):443–470, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707781.

MacDonald:1967:FCO

[Mac67] Robert H. MacDonald. The frightful consequences of onanism: Notes on the history of a delusion. Journal of the History of Ideas, 28 (3):423–431, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708628.

MacGillivray:1970:THH

[Mac70] Royce MacGillivray. Thomas Hobbes’s history of the English Civil War: a study of Behemoth. Journal of the History of Ideas, 31 (2):179–198, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708544.

MacDonald:1971:DMS

[Mac71a] Robert H. MacDonald. A disputed maxim of state in “Forth Feasting” (1619). Journal of the History of Ideas, 32(2):295–298, April/ June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708283.

Macfie:1971:IHJ

[Mac71b] Alec Macfie. The invisible hand of Jupiter. Journal of the History of Ideas, 32(4):595–599, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708980.

Macpherson:1978:EPP

[Mac78] C. B. Macpherson. The economic penetration of political theory: Some hypotheses. Journal of the History of Ideas, 39(1):101–118, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709074.

Macpherson:1980:RPW

[Mac80] C. B. Macpherson. Reply to Professor Weaver. Journal of the History of Ideas, 41(1):154–156, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709110. See [Wea80].

MacLean:1988:HTC

[Mac88] Michael MacLean. History in a two-cultures world: The case of the German historians. Journal of the History of Ideas, 49(3): 473–494, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709488.

Maclear:1992:IWI

[Mac92] J. F. Maclear. Isaac Watts and the idea of public religion. Journal of the History of Ideas, 53(1):25–45, January/March 1992. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709909.

Maclean:1998:FRE

[Mac98] Ian Maclean. Foucault’s Renaissance episteme reassessed: an Aristotelian counterblast. Journal of the History of Ideas, 59(1):149–166, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14973.

MacPhail:2001:PHA

[Mac01] Eric MacPhail. The plot of history from antiquity to the Renaissance. Journal of the History of Ideas, 62(1):1–16, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15103.

MacCormack:2006:GDI

[Mac06] Sabine MacCormack. Gods, demons, and idols in the Andes. Journal of the History of Ideas, 67(4):623–647, October 2006. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/204776.

Madison:1945:AUS

[Mad45] Charles A. Madison. Anarchism in the United States. Journal of the History of Ideas, 6(1):46–66, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707055.

Madden:1954:CWL

[Mad54] Edward H. Madden. Chauncey Wright’s life and work: Some new material. Journal of the History of Ideas, 15(3):445–455, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707764.

Madden:1957:CEN

[Mad57] Edward H. Madden. Charles Eliot Norton on art and morals. Journal of the History of Ideas, 18(3):430–438, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707801.

Maestro:1973:PAC

[Mae73] Marcello Maestro. A pioneer for the abolition of capital punishment: Cesare beccaria. Journal of the History of Ideas, 34(3):463–468, July/ September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708966.

Maestro:1975:BFP

[Mae75] Marcello Maestro. Benjamin Franklin and the penal laws. Journal of the History of Ideas, 36(3):551–562, July/September 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708664.

Maestro:1978:LRP

[Mae78] Marcello Maestro. Lafayette as a reformer of penal laws. Journal of the History of Ideas, 39(3):503–510, July/September 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709393.

Maestro:1980:GMH

[Mae80] Marcello Maestro. Going metric: How it all started. Journal of the History of Ideas, 41(3):479–486, July/September 1980. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709407.

Maestro:1983:GFH

[Mae83] Marcello Maestro. Gaetano Filangieri and his laws of relative goodness. Journal of the History of Ideas, 44(4):687–691, October/ December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709224.

Magalhaes:2016:CAM

[Mag16] Pedro T. Magalhães. A contingent affinity: Max Weber, Carl Schmitt, and the challenge of modern politics. Journal of the History of Ideas, 77(2):283–304, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/620982.

Mahoney:1987:LHB

[Mah87] Edward P. Mahoney. Lovejoy and the hierarchy of being. Journal of the History of Ideas, 48(2):211–230, April/June 1987. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709555.

Mahoney:1999:POK

[Mah99] Edward P. Mahoney. Paul Oskar Kristeller 1905–1999. Journal of the History of Ideas, 60(4):758–760, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15052. See errata [Ano00e].

Mahlev:2015:PPA

[Mah15] Haim Mahlev. A philosophy of the patriarchs?: The agenda behind Christoph August Heumann’s Acta Philosophorum. Journal of the History of Ideas, 76(4):517–539, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/597345.

Maiorino:1977:BCG

[Mai77] Giancarlo Maiorino. The breaking of the circle: Giordano Bruno and the poetics of immeasurable abundance. Journal of the History of Ideas, 38(2):317–327, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708916.

Maik:1989:NDA

[Mai89] Linda L. Maik. Nordau’s degeneration: The American controversy. Journal of the History of Ideas, 50(4):607–623, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709800.

MaiaNeto:1997:ASE

[Mai97] José Raimundo Maia Neto. Academic skepticism in early modern philosophy. Journal of the History of Ideas, 58(2):199–220, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14945.

Mailander:2021:MSN

[Mai21] Elissa Mailänder. Masters of sex? Nazism, bigamy, and a university professor’s fight with society and the state (1930–1970). Journal of the History of Ideas, 82(1):109–131, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/781615.

Malone:1940:GPH

[Mal40a] Kemp Malone. Grundtvig’s philosophy of history. Journal of the History of Ideas, 1(3):281–298, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707088.

Malone:1940:REN

[Mal40b] Kemp Malone. The rise of English nationalism. Journal of the History of Ideas, 1(4):504–505, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707127.

Malone:1958:PSG

[Mal58] Kemp Malone. Primitivism in saxo grammaticus. Journal of the History of Ideas, 19(1):94–104, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707955.

Malet:1997:IBM

[Mal97] Antoni Malet. Isaac Barrow on the mathematization of nature: Theological voluntarism and the rise of geometrical optics. Journal of the History of Ideas, 58(2):265–287, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14946.

Mali:1999:RMB

[Mal99] Joseph Mali. The reconciliation of myth: Benjamin’s homage to Bachofen. Journal of the History of Ideas, 60(1):165–187, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15018.

Maloy:2009:ALP

[Mal09] J. S. Maloy. The Aristotelianism of Locke’s politics. Journal of the History of Ideas, 70(2):235–257, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/261977.

Malcolm:2015:CEH

[Mal15] Noel Malcolm. On the Clarendon edition of Hobbes’s Leviathan: a response. Journal of the History of Ideas, 76(2):301–314, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/580301. See

[Hoe15].

Maloy:2017:BPR

[Mal17] J. S. Maloy. Bodin’s Puritan readers and radical democracy in Early New England. Journal of the History of Ideas, 78(1):1–25, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/648120.

Mandelbaum:1942:CAH

[Man42] Maurice Mandelbaum. Causal analysis in history. Journal of the History of Ideas, 3(1):30–50, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707460.

Mandelbaum:1948:ALT

[Man48] Maurice Mandelbaum. Arthur O. Lovejoy and the theory of historiography. Journal of the History of Ideas, 9(4):412–423, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707214.

Mandelbaum:1955:CRT

[Man55] Maurice Mandelbaum. Concerning recent trends in the theory of historiography. Journal of the History of Ideas, 16(4):506–517, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707507.

Manuel:1956:EO

[Man56] Frank E. Manuel. From equality to organicism. Journal of the History of Ideas, 17(1):54–69, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707685.

Mandelbaum:1957:SBE

[Man57a] Maurice Mandelbaum. The scientific background of evolutionary theory in biology. Journal of the History of Ideas, 18(3):342–361, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707797.

Manning:1957:CDO

[Man57b] John Manning. Charles Dickens and the oswego system. Journal of the History of Ideas, 18(4):580–586, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707569.

Mandelbaum:1958:DRV

[Man58] Maurice Mandelbaum. Darwin’s religious views. Journal of the History of Ideas, 19(3):363–378, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708041.

Manns:1988:SIF

[Man88] James Manns. The Scottish influence on French aesthetic thought. Journal of the History of Ideas, 49(4):633–651, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709677.

Manns:1991:SIF

[Man91] James W. Manns. The Scottish influence on French aesthetic thought: Later developments. Journal of the History of Ideas, 52 (1):103–119, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709584.

Mannings:1994:JRT

[Man94] David Mannings. Jonathan Richardson, Thomas Gray, and the genealogy of art. Journal of the History of Ideas, 55(3):405–420, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709847.

Mancini:2008:TMT

[Man08] Matthew J. Mancini. Too many Tocquevilles: The fable of Tocqueville’s American reception. Journal of the History of Ideas, 69(2): 245–268, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
235873.

Manning:2013:HH

[Man13] Gideon Manning. The history of “Hylomorphism”. Journal of the History of Ideas, 74(2):173–187, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/504783.

Marks:1953:GUO

[Mar53] Harry J. Marks. Ground under our feet: Beard’s relativism. Journal of the History of Ideas, 14(4):628–633, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707707.

Marsak:1959:BFD

[Mar59a] Leonard M. Marsak. Bernard De Fontenelle: In defense of science. Journal of the History of Ideas, 20(1):111–122, January 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707970.

Marsh:1959:SPH

[Mar59b] Robert Marsh. The second part of Hartley’s system. Journal of the History of Ideas, 20(2):264–273, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707824.

Martin:1968:MNR

[Mar68] Neil A. Martin. Marxism, nationalism, and Russia. Journal of the History of Ideas, 29(2):231–252, April/June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708578.

Martin:1972:TCA

[Mar72] Rex Martin. The two cities in Augustine’s political philosophy. Journal of the History of Ideas, 33(2):195–216, April/June 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708869.

Marsh:1976:EAB

[Mar76] David Marsh. Erasmus on the antithesis of body and soul. Journal of the History of Ideas, 37(4):673–688, October/December 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709030.

Margolis:1986:CPP

[Mar86] Nadia Margolis. Christine De Pizan: The poetess as historian. Journal of the History of Ideas, 47(3):361–375, July/September 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709658.

Martin:1990:MHH

[Mar90] Dennis D. Martin. The via moderna, humanism, and the hermeneutics of late medieval monastic life. Journal of the History of Ideas, 51(2):179–197, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709511.

Marsh:2004:JCS

[Mar04] David Marsh. Julius Caesar Scaliger’s poetics. Journal of the History of Ideas, 65(4):667–676, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/184070.

Martinich:2009:IRT

[Mar09] A. P. Martinich. Interpreting the religion of Thomas Hobbes: an exchange: Hobbes’s erastianism and interpretation. Journal of the History of Ideas, 70(1):143–163, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/258198.

Margocsy:2010:RFN

[Mar10] Dániel Margócsy. “Refer to folio and number”: Encyclopedias, the exchange of curiosities, and practices of identification before Linnaeus. Journal of the History of Ideas, 71(1):63–89, January 2010. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/370098.

Marshall:2011:CSV

[Mar11] David L. Marshall. The current state of Vico scholarship. Journal of the History of Ideas, 72(1):141–160, January 2011. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/413478.

Martinich:2012:MLH

[Mar12] A. P. Martinich. A moderate logic of the history of ideas. Journal of the History of Ideas, 73(4):609–625, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/488351.

Marshall:2017:TW

[Mar17] David L. Marshall. A thousand Warburgs. Journal of the History of Ideas, 78(4):645–664, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/676150.

Martinez:2019:GSM

[Mar19] Raúl Martínez Martínez. Geoffrey Scott and modern architectural thought: The creation of a legacy throughout the Twentieth Century. Journal of the History of Ideas, 80(4):597–619, October 2019. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/735383.

Martins:2020:HNM

[Mar20] Pedro Martins. History, nation, and modernity: The idea of “Decadência” in Portuguese medievalist discourses (1842–1940). Journal of the History of Ideas, 81(3):451–471, July 2020. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/761038.

Marcocci:2022:ITE

[Mar22] Giuseppe Marcocci. Iberian theories of empire in the Sixteenth and Seventeenth Centuries. Journal of the History of Ideas, 83(4):671– 683, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
865887.

Marcus:2023:KOA

[Mar23] Hannah Marcus. Knowing old age in the Renaissance: Medicine, poetry, and spirituality in Ulisse Aldrovandi’s Encyclopedia of Old Age. Journal of the History of Ideas, 84(1):51–75, January 2023. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/879540.

Masur:1952:WDH

[Mas52] Gerhard Masur. Wilhelm Dilthey and the history of ideas. Journal of the History of Ideas, 13(1):94–107, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707727.

Massa:1977:GBI

[Mas77] Daniel Massa. Giordano Bruno’s ideas in Seventeenth-Century England. Journal of the History of Ideas, 38(2):227–242, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708909.

Masroori:2000:ETN

[Mas00] Cyrus Masroori. European thought in Nineteenth-Century Iran: David Hume and others. Journal of the History of Ideas, 61(4):657– 674, October 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15097.

Matthiessen:1943:BRC

[Mat43] F. O. Matthiessen. Book review: Criticism and Fiction. Journal of the History of Ideas, 4(3):368–373, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707260.

Mathur:1955:HKT

[Mat55] G. B. Mathur. Hume and Kant in their relation to the pragmatic movement. Journal of the History of Ideas, 16(2):198–208, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707662.

Matson:1957:HAP

[Mat57] Floyd W. Matson. History as art: The psychological-romantic view. Journal of the History of Ideas, 18(2):270–279, April 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707629.

Matheson:2006:EMP

[Mat06] Tamara Chaplin Matheson. Embodying the mind, producing the nation: Philosophy on French television. Journal of the History of Ideas, 67(2):315–341, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/198077.

Mautner:2005:GS

[Mau05] Thomas Mautner. Grotius and the skeptics. Journal of the History of Ideas, 66(4):577–601, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/191475.

Mayer:1988:TSU

[May88] Thomas F. Mayer. Thomas Starkey, an unknown conciliarist at the Court of Henry VIII. Journal of the History of Ideas, 49(2):207–227, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709497.

Mayr:1990:WHW

[May90] Ernst Mayr. When is historiography whiggish? Journal of the History of Ideas, 51(2):301–309, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709517.

Mayr:1992:IT

[May92] Ernst Mayr. The idea of teleology. Journal of the History of Ideas, 53(1):117–135, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709913.

Mayhew:2004:BGR

[May04] Robert Mayhew. British Geography’s Republic of Letters: Mapping an imagined community, 1600–1800. Journal of the History of Ideas, 65(2):251–276, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/173849.

Mazzeo:1953:MPP

[Maz53] Joseph Anthony Mazzeo. Metaphysical poetry and the poetic of correspondence. Journal of the History of Ideas, 14(2):221–234, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707472.

Mazzeo:1954:UAC

[Maz54] Joseph A. Mazzeo. Universal analogy and the culture of the Renaissance. Journal of the History of Ideas, 15(2):299–304, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707774.

Mazzeo:1957:DCL

[Maz57] Joseph Anthony Mazzeo. Dante’s conception of love. Journal of the History of Ideas, 18(2):147–160, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707621.

Mazzeo:1960:CMP

[Maz60] Joseph Anthony Mazzeo. Cromwell as Machiavellian Prince in Marvell’s “An Horatian Ode”. Journal of the History of Ideas, 21 (1):1–17, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707996.

Mazzeo:1962:SAR

[Maz62] Joseph Anthony Mazzeo. St. Augustine’s rhetoric of silence. Journal of the History of Ideas, 23(2):175–196, April/June 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708154.

Mazzeo:1972:SIH

[Maz72] Joseph Anthony Mazzeo. Some interpretations of the history of ideas. Journal of the History of Ideas, 33(3):379–394, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709042.

Mazur:2014:SHC

[Maz14] Peter A. Mazur. Searcher of hearts: Cesare Baronio’s history of conversion. Journal of the History of Ideas, 75(2):213–235, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/544559.

Malter:1981:MCG

[MB81] Rudolf Malter and Hamilton H. H. Beck. Main currents in the German interpretation of the critique of pure reason since the beginnings of neo-Kantianism. Journal of the History of Ideas, 42(3):531–551, July/ September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709192.

Martin:1995:HFS

[MB95] Raymond Martin and John Barresi. Hazlitt on the future of the self. Journal of the History of Ideas, 56(3):463–481, July 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710036.

McAleer:1999:GRP

[McA99] G. J. McAleer. Giles of Rome on political authority. Journal of the History of Ideas, 60(1):21–36, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15019.

McAlpin:2010:IER

[McA10] Mary K. McAlpin. Innocence of experience: Rousseau on puberty in the state of civilization. Journal of the History of Ideas, 71(2):241– 261, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
379778.

McColley:1942:EPA

[McC42] Grant McColley. An early poetic allusion to the Copernican theory. Journal of the History of Ideas, 3(3):355–357, June 1942. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707309.

McCutchan:1958:JEE

[McC58] J. Wilson McCutchan. Justice and equity in the English morality play. Journal of the History of Ideas, 19(3):405–410, June 1958. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708044.

McConnell:1963:HRP

[McC63] Allen McConnell. Helvetius’ Russian pupils. Journal of the History of Ideas, 24(3):373–386, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708213.

McClaughlin:1979:CDC

[McC79] Trevor McClaughlin. Censorship and defenders of the Cartesian faith in Mid-Seventeenth Century France. Journal of the History of Ideas, 40(4):563–581, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709359.

McCloskey:1990:RM

[McC90] Donald N. McCloskey. Reply to Munz. Journal of the History of Ideas, 51(1):143–147, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709751. See [Mun90].

McCalla:1994:PPP

[McC94] Arthur McCalla. Palingénésie Philosophique to Palingénésie Sociale: From a scientific ideology to a historical ideology. Journal of the History of Ideas, 55(3):421–439, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709848.

McCalla:2015:PRF

[McC15] Arthur McCalla. Paganism in Restoration France: Eckstein’s traditionalist orientalism. Journal of the History of Ideas, 76(4):563–585, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
597347.

McCray:2020:PMD

[McC20] J. Louise McCray. “Peril in the means of its diffusion”: William Godwin on truth and social media. Journal of the History of Ideas, 81(1):67–84, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/747168.

McDowell:2005:ICW

[McD05] Nicholas McDowell. Ideas of creation in the writings of Richard Overton the leveller and Paradise Lost. Journal of the History of Ideas, 66(1):59–78, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/185260.

McGrade:1963:CHP

[McG63] Arthur S. McGrade. The coherence of Hooker’s polity: The books on power. Journal of the History of Ideas, 24(2):163–182, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707843.

McGuire:1972:BCN

[McG72] J. E. McGuire. Boyle’s conception of nature. Journal of the History of Ideas, 33(4):523–542, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708855.

McGinn:1980:NT

[McG80] Robert E. McGinn. Nietzsche on technology. Journal of the History of Ideas, 41(4):679–691, October/December 1980. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709280.

McGetchin:2003:WFT

[McG03] Douglas T. McGetchin. Wilting florists: The turbulent early decades of the Société Asiatique, 1822–1860. Journal of the History of Ideas, 64(4):565–580, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/52451.

McGrath:2013:BCA

[McG13] Larry McGrath. Bergson comes to America. Journal of the History of Ideas, 74(4):599–620, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/523430.

McKeon:1947:ACD

[McK47] Richard McKeon. Aristotle’s conception of the development and the nature of scientific method. Journal of the History of Ideas, 8(1):3–44, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707440.

McKeon:1966:PDS

[McK66] Richard McKeon. Philosophy and the development of scientific methods. Journal of the History of Ideas, 27(1):3–22, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708305.

McKeon:1975:PTH

[McK75] Richard McKeon. Philosophy and theology, history and science in the thought of bonaventura and Thomas Aquinas. Journal of the History of Ideas, 36(3):387–412, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708653.

McKeon:1981:IPT

[McK81a] Richard McKeon. The interpretation of political theory and practice in Ancient Athens. Journal of the History of Ideas, 42(1):3–12, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709414.

McKerrow:1981:RWN

[McK81b] Ray E. McKerrow. Richard Whately on the nature of human knowledge in relation to ideas of his contemporaries. Journal of the History of Ideas, 42(3):439–455, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709186.

McKenzie:1982:RDM

[McK82a] Lionel A. McKenzie. Rousseau’s debate with Machiavelli in the social contract. Journal of the History of Ideas, 43(2):209–228, April/ June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709200.

McKim:1982:WAI

[McK82b] Robert McKim. Wenz on abstract ideas and Christian neo-Platonism in Berkeley. Journal of the History of Ideas, 43(4):665–671, October/ December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709349.

See reply [Wen82].

McKinney:1983:OMD

[McK83] Ronald H. McKinney. The origins of modern dialectics. Journal of the History of Ideas, 44(2):179–190, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709134.

McKim:1987:LAD

[McK87] Robert McKim. Luce’s account of the development of Berkeley’s immaterialism. Journal of the History of Ideas, 48(4):649–669, October/ December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709692.

McLachlan:1976:TZP

[McL76a] Hugh V. McLachlan. Thomas, Znaniecki, and Popper on falsification. Journal of the History of Ideas, 37(3):547, July/September 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708815. See erratum [Ano77e].

McLaren:1976:SSO

[McL76b] Angus McLaren. Sex and socialism: The opposition of the French Left to birth control in the Nineteenth Century. Journal of the History of Ideas, 37(3):475–492, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708810.

McLaverty:1986:DEL

[McL86] James McLaverty. From definition to explanation: Locke’s influence on Johnson’s dictionary. Journal of the History of Ideas, 47 (3):377–394, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709659.

McManmon:1966:PRI

[McM66] John J. McManmon. The problem of a religious interpretation of Gulliver’s fourth voyage. Journal of the History of Ideas, 27 (1):59–72, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708308.

McNeil:1945:CRF

[McN45] Gordon H. McNeil. The cult of Rousseau and the French Revolution. Journal of the History of Ideas, 6(2):197–212, April 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707363.

McNally:1966:CVX

[McN66] Raymond T. McNally. Chaadaev versus Xomjakov in the late 1830’s and the 1840’s. Journal of the History of Ideas, 27(1):73–91, January/ March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708309.

McNulty:1973:ASC

[McN73] Paul J. McNulty. Adam Smith’s concept of labor. Journal of the History of Ideas, 34(3):345–366, July/September 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708957.

McQuillan:2014:OPC

[McQ14] J. Colin McQuillan. Oaths, promises, and compulsory duties: Kant’s response to Mendelssohn’s Jerusalem. Journal of the History of Ideas, 75(4):581–604, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/557234.

McRae:1951:HPP

[McR51] Robert McRae. Hume as a political philosopher. Journal of the History of Ideas, 12(2):285–290, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707518.

McRae:1955:RTT

[McR55] Kenneth D. McRae. Ramist tendencies in the thought of Jean Bodin. Journal of the History of Ideas, 16(3):306–323, June 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707634.

McRae:1957:USB

[McR57] Robert McRae. The unity of the sciences: Bacon, Descartes, and Leibniz. Journal of the History of Ideas, 18(1):27–48, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707578.

McRae:1963:PBC

[McR63] Kenneth D. McRae. A postscript on Bodin’s connections with ramism. Journal of the History of Ideas, 24(4):569–571, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707985.

McRae:1965:IPT

[McR65] Robert McRae. “Idea” as a philosophical term in the Seventeenth Century. Journal of the History of Ideas, 26(2):175–190, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708226.

Morsel-Eisenberg:2021:ATR

[ME21] Tamara Morsel-Eisenberg. Anxieties of transmission: Rabbinic responsa and Early Modern “Print culture”. Journal of the History of Ideas, 82(3):377–404, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/798271.

Megill:1975:TEA

[Meg75] A. D. Megill. Theory and experience in Adam Smith. Journal of the History of Ideas, 36(1):79–94, January/March 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709012.

Megill:1987:RFH

[Meg87] Allan Megill. The reception of Foucault by historians. Journal of the History of Ideas, 48(1):117–141, January/March 1987. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709615.

Megill:2005:GHI

[Meg05] Allan Megill. Globalization and the history of ideas. Journal of the History of Ideas, 66(2):179–187, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/187189.

Melvin:1975:BTR

[Mel75] Peter H. Melvin. Burke on theatricality and revolution. Journal of the History of Ideas, 36(3):447–468, July/September 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708656.

Meli:1999:CLC

[Mel99] Domenico Bertoloni Meli. Caroline, Leibniz, and Clarke. Journal of the History of Ideas, 60(3):469–486, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15033.

Menzel:1956:SJH

[Men56] Johanna M. Menzel. The Sinophilism of J. H. G. Justi. Journal of the History of Ideas, 17(3):300–310, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707546.

Mengelberg:1961:LSH

[Men61] Kaethe Mengelberg. Lorenz Von Stein and his contribution to historical sociology. Journal of the History of Ideas, 22(2):267–274, April/ June 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707838.

Menchinger:2016:FWP

[Men16] Ethan L. Menchinger. Free will, predestination, and the fate of the Ottoman Empire. Journal of the History of Ideas, 77(3):445–466, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/626910.

Menin:2018:PVE

[Men18] Marco Menin. Paul et Virginie, or the enigma of evil: The double theodicy of Bernardin de Saint-Pierre. Journal of the History of Ideas, 79(4):593–612, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/707774.

Menon:2020:GSW

[Men20] Nikhil Menon. Gandhi’s spinning wheel: The charkha and its regenerative effects. Journal of the History of Ideas, 81(4):643–662, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/772516.

Merlan:1945:BF

[Mer45] Philip Merlan. Brentano and Freud. Journal of the History of Ideas, 6(3):375–377, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707299.

Merlan:1947:FCP

[Mer47] Philip Merlan. Form and content in Plato’s philosophy. Journal of the History of Ideas, 8(4):406–430, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707427.

Merlan:1948:PHJ

[Mer48] Philip Merlan. Parva Hamanniana: J. G. Hamann as a spokesman of the middle class. Journal of the History of Ideas, 9(3):380–384, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707376.

Merlan:1949:BFS

[Mer49a] Philip Merlan. Brentano and Freud — a sequel. Journal of the History of Ideas, 10(3):451, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707048.

Merlan:1949:EH

[Mer49b] Philip Merlan. Epicureanism and Horace. Journal of the History of Ideas, 10(3):445–451, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707047.

Merlan:1949:PHI

[Mer49c] Philip Merlan. Parva Hamanniana (II) Hamann and Schmohl. Journal of the History of Ideas, 10(4):567–574, October 1949. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707190.

Merlan:1950:LPP

[Mer50a] Philip Merlan. Lucretius — primitivist or progressivist. Journal of the History of Ideas, 11(3):364–368, June 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707737.

Merlan:1950:PHH

[Mer50b] Philip Merlan. Parva Hamanniana: Hamann and Galiani. Journal of the History of Ideas, 11(4):486–489, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707592.

Merlan:1953:AMS

[Mer53] Philip Merlan. Abstraction and metaphysics in St. Thomas’ Summa. Journal of the History of Ideas, 14(2):284–291, April 1953. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707477.

Merton:1966:MMS

[Mer66] Robert C. Merton. The “Motionless” motion of Swift’s flying island. Journal of the History of Ideas, 27(2):275–277, April/June 1966. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708643.

Merquior:1989:BCH

[Mer89] J. G. Merquior. Back to the context: Harry Levin, an appreciation. Journal of the History of Ideas, 50(4):667–678, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709804.

Merrick:1990:SCR

[Mer90] Jeffrey Merrick. Subjects and citizens in the remonstrances of the parlement of Paris in the Eighteenth Century. Journal of the History of Ideas, 51(3):453–460, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709624.

Merchant:2008:SNB

[Mer08] Carolyn Merchant. Secrets of nature: The Bacon debates revisited. Journal of the History of Ideas, 69(1):147–162, January 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/231011.

Mercer:2011:PRM

[Mer11] Ben Mercer. The paperback revolution: Mass-circulation books and the cultural origins of 1968 in Western Europe. Journal of the History of Ideas, 72(4):613–636, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/453551.

Metzger:1960:CVS

[Met60] Lore Metzger. Coleridge’s vindication of Spinoza: an unpublished note. Journal of the History of Ideas, 21(2):279–293, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708201.

Metzger:1963:BRW

[Met63] Lore Metzger. Book review: The Workshop of Productive Eclecticism. Journal of the History of Ideas, 24(1):143–149, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707865.

Methuen:1996:RHT

[Met96] Charlotte Methuen. The role of the heavens in the thought of Philip Melanchthon. Journal of the History of Ideas, 57(3):385–403, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14918.

Metcalf:2014:LE

[Met14] Thomas R. Metcalf. Liberalism and empire. Journal of the History of Ideas, 75(4):687–695, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/557241.

Mulsow:2006:ISA

[MF06] Martin Mulsow and Robert Folger. Idolatry and science: Against nature worship from Boyle to Rüdiger, 1680–1720. Journal of the History of Ideas, 67(4):697–711, October 2006. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/204778.

Mulsow:2004:APS

[MH04] Martin Mulsow and Janita Hamalainen. Ambiguities of the Prisca Sapientia in Late Renaissance humanism. Journal of the History of Ideas, 65(1):1–13, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/168336.

Michael:2003:JWB

[Mic03] Emily Michael. John Wyclif on body and mind. Journal of the History of Ideas, 64(3):343–360, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/1/

article/47661.

Middendorf:1960:DJM

[Mid60] John H. Middendorf. Dr. Johnson and mercantilism. Journal of the History of Ideas, 21(1):66–83, January/March 1960. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707999.

Miel:1969:PPR

[Mie69] Jan Miel. Pascal, Port-Royal, and Cartesian linguistics. Journal of the History of Ideas, 30(2):261–271, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708438.

Miedema:1973:BRK

[Mie73] Hessel Miedema. Book review: Karel Van Mander’s Grondt Der Edel Vry Schilder-Const: (“Foundations of the Noble and Free Art of Painting”). Journal of the History of Ideas, 34(4):653–668, October/ December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708899.

Mikulak:1970:DSG

[Mik70] Maxim W. Mikulak. Darwinism, Soviet genetics, and Marxism– Leninism. Journal of the History of Ideas, 31(3):359–376, July/ September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708511.

Miller:1941:BRN

[Mil41] Perry Miller. Book review: New England: Indian Summer by Van Wyck Brooks. Journal of the History of Ideas, 2(1):116–122, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707285.

Miller:1943:PSS

[Mil43] Perry Miller. “Preparation for salvation” in Seventeenth-Century New England. Journal of the History of Ideas, 4(3):253–286, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707254.

Miliband:1954:PRO

[Mil54] Ralph Miliband. The politics of Robert Owen. Journal of the History of Ideas, 15(2):233–245, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707769.

Miller:1956:BRB

[Mil56] J. Hillis Miller. Book review: Baird’s Ishmael. Journal of the History of Ideas, 17(4):555–560, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707789.

Miller:1961:JSM

[Mil61] Kenneth E. Miller. John Stuart Mill’s theory of international relations. Journal of the History of Ideas, 22(4):493–514, October/ December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708027.

Miliani:1963:UST

[Mil63] Domingo Miliani. Utopian socialism, transitional thread from romanticism to positivism in Spanish America. Journal of the History of Ideas, 24(4):523–538, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707982.

Miller:1964:CCN

[Mil64] Craig W. Miller. Coleridge’s concept of nature. Journal of the History of Ideas, 25(1):77–96, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708086.

Miles:1970:ICW

[Mil70] Wyndham D. Miles. The idea of chemical warfare in modern times. Journal of the History of Ideas, 31(2):297–304, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708553.

Miles:1974:YAN

[Mil74] Edwin A. Miles. The young American nation and the classical world. Journal of the History of Ideas, 35(2):259–274, April/June 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708761.

Miller:1976:AFC

[Mil76] B. Jaye Miller. Anarchism and French Catholicism in esprit. Journal of the History of Ideas, 37(1):163–174, January/March 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708718.

Capek:1987:CBA

[Mil87] Milič Čapek. The conflict between the absolutist and the relational theory of time before Newton. Journal of the History of Ideas, 48 (4):595–608, October/December 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709689.

Miller:1996:CCE

[Mil96] Peter N. Miller. Citizenship and culture in Early Modern Europe. Journal of the History of Ideas, 57(4):725–742, October 1996. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14923.

Milner:1997:FBT

[Mil97] Benjamin Milner. Francis Bacon: The theological foundations of Valerius Terminus. Journal of the History of Ideas, 58(2):245–264, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14947.

Milcinski:1999:ZAD

[Mil99a] Maja Milcinski. Zen and the art of death. Journal of the History of Ideas, 60(3):385–397, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15039.

Milnes:1999:EAM

[Mil99b] Tim Milnes. Eclipsing art: Method and metaphysics in Coleridge’s Biographia Literaria. Journal of the History of Ideas, 60(1):125–147, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15020.

Miller:2001:ABS

[Mil01] Peter N. Miller. The “Antiquarianization” of biblical scholarship and the London polyglot Bible (1653–57). Journal of the History of Ideas, 62(3):463–482, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15121.

Miller:2006:HRB

[Mil06] Peter N. Miller. History of religion becomes ethnology: Some evidence from Peiresc’s Africa. Journal of the History of Ideas, 67(4):675– 696, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
204777.

Mintz:1951:BRD

[Min51] Samuel I. Mintz. Book review: Douglas Bush’s Science and English Poetry. Journal of the History of Ideas, 12(1):155–157, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707543.

Mindel:1965:UMH

[Min65] Joseph Mindel. The uses of metaphor: Henry Adams and the symbols of science. Journal of the History of Ideas, 26(1):89–102, January/ March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708401.

Mintz:1968:HLH

[Min68] Samuel I. Mintz. Hobbes on the law of heresy: a new manuscript. Journal of the History of Ideas, 29(3):409–414, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708451.

Mintz:1970:HKL

[Min70] Samuel I. Mintz. Hobbes’s knowledge of the law: a reply. Journal of the History of Ideas, 31(4):614–615, October/December 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708265.

Miner:1998:VFP

[Min98] Robert C. Miner. Verum-factum and practical wisdom in the early writings of Giambattista Vico. Journal of the History of Ideas, 59(1):53–73, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14974.

Mirarchi:1977:FAM

[Mir77a] Lawrence A. Mirarchi. Force and absolute motion in Berkeley’s philosophy of physics. Journal of the History of Ideas, 38(4):705–713, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708698.

See reply [Sil77].

Mirarchi:1977:RPS

[Mir77b] Lawrence A. Mirarchi. A rejoinder to Professor Silver’s reply. Journal of the History of Ideas, 38(4):716–718, October/December 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708700.

Missner:1977:HML

[Mis77] Marshall Missner. Hobbes’s method in Leviathan. Journal of the History of Ideas, 38(4):607–621, October/December 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708690.

Missner:1983:SHP

[Mis83] Marshall Missner. Skepticism and Hobbes’s political philosophy. Journal of the History of Ideas, 44(3):407–427, July/September 1983. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709174.

Mitchell:1965:HR

[Mit65] Harvey Mitchell. Hobson revisited. Journal of the History of Ideas, 26(3):397–416, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708389.

Mittelstrass:1971:BRP

[Mit71a] Jürgen Mittelstrass. Book review: Philosophy as Artistic Achievement? Journal of the History of Ideas, 32(3):459–470, July/ September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708361.

Mitzman:1971:TGS

[Mit71b] Arthur Mitzman. Tönnies and German society, 1887–1914: From cultural pessimism to celebration of the Volksgemeinschaft. Journal of the History of Ideas, 32(4):507–524, October/December 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708974.

Mitzman:1987:MDC

[Mit87] Arthur Mitzman. Michelet, Danton, and the corruption of revolutionary virtue. Journal of the History of Ideas, 48(3):453–466, July/ September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709762.

Mitchell:1993:RSP

[Mit93] Harvey Mitchell. Reclaiming the self: The Pascal–Rousseau connection. Journal of the History of Ideas, 54(4):637–658, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709827.

Mitzman:1994:BRH

[Mit94] Arthur Mitzman. Book review: On the Historicity of High Culture. Journal of the History of Ideas, 55(1):159–173, January 1994. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709959.

Mitzman:1996:MSR

[Mit96] Arthur Mitzman. Michelet and social romanticism: Religion, revolution, nature. Journal of the History of Ideas, 57(4):659–682, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14924.

Michael:1989:CIS

[MM89] Emily Michael and Fred S. Michael. Corporeal ideas in Seventeenth-Century psychology. Journal of the History of Ideas, 50(1):31–48, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709785.

Michael:1990:TIG

[MM90] Fred S. Michael and Emily Michael. The theory of ideas in Gassendi and Locke. Journal of the History of Ideas, 51(3):379–399, July/ September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709621.

McKenzie-McHarg:2022:SPP

[MM22] Andrew McKenzie-McHarg. From status politics to the paranoid style: Richard Hofstadter and the pitfalls of psychologizing history. Journal of the History of Ideas, 83(3):451–475, July 2022. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/858858.

Modrak:1994:BRT

[Mod94] Deborah K. W. Modrak. Book review: Theophrastus and Recent Scholarship. Journal of the History of Ideas, 55(2):337–345, April 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709903.

Mohamed:2004:RTC

[Moh04] Feisal G. Mohamed. Renaissance thought on the celestial hierarchy: The decline of a tradition? Journal of the History of Ideas, 65(4):559–582, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/184071.

Molland:1978:MIS

[Mol78] A. G. Molland. Medieval ideas of scientific progress. Journal of the History of Ideas, 39(4):561–577, October/December 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709442.

Molivas:1997:RPD

[Mol97] Gregory I. Molivas. Richard Price, the debate on free will, and natural rights. Journal of the History of Ideas, 58(1):105–123, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14932.

Mommsen:1950:BRF

[Mom50] Theodor E. Mommsen. Book review: Frederick Antal’s Florentine Painting and its Social Background. Journal of the History of Ideas, 11(3):369–379, June 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707738.

Mommsen:1951:SAC

[Mom51] Theodor E. Mommsen. St. Augustine and the Christian idea of progress: The background of the city of God. Journal of the History of Ideas, 12(3):346–374, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707751.

Monk:1944:GBR

[Mon44] Samuel Holt Monk. “A grace beyond the reach of art”. Journal of the History of Ideas, 5(2):131–150, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707381.

Montague:1948:MFL

[Mon48] William Pepperell Montague. My friend Lovejoy. Journal of the History of Ideas, 9(4):424–427, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707215.

Monro:1974:GOM

[Mon74] D. H. Monro. Godwin, Oakeshott, and Mrs. Bloomer. Journal of the History of Ideas, 35(4):611–624, October/December 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709089.

Monfasani:1989:WLV

[Mon89] John Monfasani. Was Lorenzo Valla an ordinary language philosopher? Journal of the History of Ideas, 50(2):309–323, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709737.

Monheit:1997:GBA

[Mon97] Michael Leonard Monheit. Guillaume Bude, Andrea Alciato, Pierre de l’Estoile: Renaissance interpreters of Roman law. Journal of the History of Ideas, 58(1):21–40, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14933.

Monod:2019:ATD

[Mon19] Jean-Claude Monod. Archives, thresholds, discontinuities: Blumenberg and Foucault on historical substantialism and the phenomenology of history. Journal of the History of Ideas, 80(1):133–146, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/715740.

Moore:1945:SRC

[Moo45] B. Moore, Jr. Some readjustments in Communist theory: a note on the relation between ideas and social change. Journal of the History of Ideas, 6(4):468–482, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707346.

Moody:1951:GADa

[Moo51a] Ernest A. Moody. Galileo and Avempace: The dynamics of the Leaning Tower Experiment (I). Journal of the History of Ideas, 12(2):163– 193, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707514.

Moody:1951:GADb

[Moo51b] Ernest A. Moody. Galileo and Avempace: The dynamics of the Leaning Tower Experiment (II). Journal of the History of Ideas, 12(3):375– 422, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707752.

Moore:1968:BRO

[Moo68] John C. Moore. Book review: The Origins of Western Ideas. Journal of the History of Ideas, 29(1):141–151, January/March 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708472.

Moore:1976:LAL

[Moo76] J. T. Moore. Locke’s analysis of language and the assent to scripture. Journal of the History of Ideas, 37(4):707–714, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709033.

Moore:1979:CLP

[Moo79] John C. Moore. “Courtly love”: a problem of terminology. Journal of the History of Ideas, 40(4):621–632, October/December 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709362.

Mornet:1940:SEI

[Mor40] Daniel Mornet. La signification et l’Évolution de l’Idée de préciosité en France au XVIIe siècle. Journal of the History of Ideas, 1(2):225–231, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707334.

More:1941:BA

[Mor41] Louis Trenchard More. Boyle as alchemist. Journal of the History of Ideas, 2(1):61–76, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707281.

Morris:1945:NE

[Mor45] Charles Morris. Nietzsche — an evaluation. Journal of the History of Ideas, 6(3):285–293, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707292.

Mora:1953:SMP

[Mor53] José Ferrater Mora. Suárez and modern philosophy. Journal of the History of Ideas, 14(4):528–547, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707700.

Morse:1954:TTS

[Mor54] Richard M. Morse. Toward a theory of Spanish American government. Journal of the History of Ideas, 15(1):71–93, January 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707650.

Mora:1957:BRG

[Mor57] José Ferrater Mora. Book review: On the “Great Analogy”. Journal of the History of Ideas, 18(2):280–284, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707630.

Morawinska:1977:ECP

[Mor77a] Agnieszka Morawińska. Eighteenth-Century “paysages moralisés”. Journal of the History of Ideas, 38(3):461–475, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708675.

Morrison:1977:PHB

[Mor77b] James C. Morrison. Philosophy and history in Bacon. Journal of the History of Ideas, 38(4):585–606, October/December 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708689.

Moravia:1978:HMH

[Mor78a] Sergio Moravia. From homme machine to homme sensible: Changing Eighteenth-Century models of man’s image. Journal of the History of Ideas, 39(1):45–60, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709071.

Morrison:1978:BRT

[Mor78b] James C. Morrison. Book review: Three Interpretations of Vico. Journal of the History of Ideas, 39(3):511–518, July/September 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709394.

Morrison:1978:VPV

[Mor78c] James C. Morrison. Vico’s principle of verum is factum and the problem of historicism. Journal of the History of Ideas, 39(4): 579–595, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709443.

Morrison:1980:VS

[Mor80] James C. Morrison. Vico and Spinoza. Journal of the History of Ideas, 41(1):49–68, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709102.

Morrow:1986:NCC

[Mor86] John Morrow. The national church in Coleridge’s church and state: a response to Allen. Journal of the History of Ideas, 47(4):640– 652, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709723.

Moran:1989:PPE

[Mor89] Seán Farrell Moran. Patrick Pearse and the European revolt against reason. Journal of the History of Ideas, 50(4):625–643, October/ December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709801.

Morris:1990:WNL

[Mor90] Randall C. Morris. Whitehead and the new liberals on social progress. Journal of the History of Ideas, 51(1):75–92, January/March 1990.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709747.

Morgan:1992:CAF

[Mor92] David Morgan. Concepts of abstraction in French art theory from the Enlightenment to modernism. Journal of the History of Ideas, 53 (4):669–685, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709943.

Moran:1993:BPP

[Mor93] Francis Moran III. Between primates and primitives: Natural man as the missing link in Rousseau’s second discourse. Journal of the History of Ideas, 54(1):37–58, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709859.

Mortensen:1994:SMA

[Mor94] Preben Mortensen. Shaftesbury and the morality of art appreciation. Journal of the History of Ideas, 55(4):631–650, October 1994. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709926.

Morgan:1996:EAA

[Mor96] David Morgan. The enchantment of art: Abstraction and empathy from German romanticism to expressionism. Journal of the History of Ideas, 57(2):317–341, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14905.

Morley:1998:PEC

[Mor98] Neville Morley. Political economy and classical antiquity. Journal of the History of Ideas, 59(1):95–114, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14975.

Mori:2004:BSE

[Mor04a] Gianluca Mori. Bayle, Saint-Evremond, and fideism: a reply to Thomas M. Lennon. Journal of the History of Ideas, 65(2):323–334, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/173850.

See [Len02] and rejoinder [Len04].

Mori:2004:SR

[Mor04b] Gianluca Mori. A short reply. Journal of the History of Ideas, 65(2): 343–344, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
173851.

Mortimer:2009:HLH

[Mor09] Sarah Mortimer. Human liberty and human nature in the works of faustus socinus and his readers. Journal of the History of Ideas, 70(2): 191–211, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
261975.

Mori:2016:DJR

[Mor16] Giuliano Mori. Democritus Junior as reader of Auctoritates: Robert Burton’s method and the anatomy of melancholy. Journal of the History of Ideas, 77(3):379–399, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/626907.

Morelli:2018:RWC

[Mor18] Federica Morelli. Race, wars, and citizenship: Free people of color in the Spanish American independence. Journal of the History of Ideas, 79(1):143–156, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/684837.

Morefield:2020:CCH

[Mor20] Jeanne Morefield. Crashing the cathedral: Historical reassessments of Twentieth-Century international relations. Journal of the History of Ideas, 81(1):131–155, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/747169.

Morales:2022:NEC

[Mor22a] Edgardo Pérez Morales. The Neapolitan Enlightenment and the conceptual challenges of antislavery legislation in Colombia. Journal of the History of Ideas, 83(3):431–450, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/858857.

Mortimer:2022:WCL

[Mor22b] Sarah Mortimer. Warfare, Christianity, and the law of nature. Journal of the History of Ideas, 83(4):613–627, October 2022. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/865883.

Morton:2025:SMM

[Mor25] Peter A. Morton. Spiritual and medical melancholy in Lutheran responses to Johann Weyer’s criticism of the witch trials. Journal of the History of Ideas, 86(1):21–47, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/949926.

Mossner:1941:WHT

[Mos41] Ernest Campbell Mossner. Was Hume a Tory historian? Facts and reconsiderations. Journal of the History of Ideas, 2(2):225–236, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707114.

Mossner:1948:HEM

[Mos48] Ernest Campbell Mossner. Hume’s early memoranda, 1729–1740: The complete text. Journal of the History of Ideas, 9(4):492–518, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707220.

Mossner:1951:FAH

[Mos51] Ernest Campbell Mossner. The first answer to humes treatise: an unnoticed item of 1740. Journal of the History of Ideas, 12(2):291– 294, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707519.

Mossner:1960:PME

[Mos60] Ernest Campbell Mossner. “Of the principle of moral estimation: a discourse between David Hume, Robert Clerk, and Adam Smith”: an unpublished MS by Adam Ferguson. Journal of the History of Ideas, 21(2):222–232, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708196.

Mosse:1961:MON

[Mos61] G. L. Mosse. The mystical origins of national socialism. Journal of the History of Ideas, 22(1):81–96, January/March 1961. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707875.

Mosse:1966:BRN

[Mos66] George L. Mosse. Book review: E. Nolte on Three Faces of Fascism. Journal of the History of Ideas, 27(4):621–625, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708345.

Mossakowski:1973:SMC

[Mos73] Stanisław Mossakowski. The symbolic meaning of Copernicus’ seal. Journal of the History of Ideas, 34(3):451–460, July/September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708964.

Mossner:1978:RDH

[Mos78] Ernest Campbell Mossner. The religion of David Hume. Journal of the History of Ideas, 39(4):653–663, October/December 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709448.

Moss:1985:MEF

[Mos85] Bernard H. Moss. Marx and Engels on French social democracy: Historians or revolutionaries? Journal of the History of Ideas, 46(4): 539–557, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709544.

Moss:1998:PJL

[Mos98] Ann Moss. The Politica of Justus Lipsius and the commonplacebook. Journal of the History of Ideas, 59(3):421–436, July 1998. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14993.

Mott:2001:LCM

[Mot01] Margaret MacLeish Mott. Leonor de Caceres and the Mexican Inquisition. Journal of the History of Ideas, 62(1):81–98, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15104.

Mourant:1949:MNP

[Mou49] John A. Mourant. Mr. neill and physiocracy. Journal of the History of Ideas, 10(1):130–134, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707204.

Moynihan:1975:CEW

[Moy75] Robert D. Moynihan. Clarissa and the enlightened woman as literary heroine. Journal of the History of Ideas, 36(1):159–166, January/ March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709018.

Moyer:2003:HAS

[Moy03a] Ann E. Moyer. Historians and antiquarians in Sixteenth-Century Florence. Journal of the History of Ideas, 64(2):177–193, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/43512.

Moyn:2003:AFT

[Moy03b] Samuel Moyn. Amos Funkenstein on the theological origins of historicism. Journal of the History of Ideas, 64(4):639–657, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/52452.

Moyn:2007:IHD

[Moy07] Samuel Moyn. Intellectual history and democracy: an interview with Pierre rosanvallon: Introduction. Journal of the History of Ideas, 68(4):701–702, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/223130.

Moyn:2023:HAA

[Moy23] Samuel Moyn. Hannah Arendt among the Cold War liberals. Journal of the History of Ideas, 84(3):533–558, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/901493.

Makinen:2004:ICM

[MP04] Virpi Mäkinen and Heikki Pihlajamaki. The individualization of crime in medieval canon law. Journal of the History of Ideas, 65(4):525–542, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/184069.

Menudo:2020:RER

[MR20] José M. Menudo and Nicolas Rieucau. The rural economics of René de girardin: Landscapes at the service of L’Idéologie nobiliaire. Journal of the History of Ideas, 81(3):429–449, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/761030.

Morot-Sir:1961:SCD

[MS61] Edouard Morot-Sir. Sartre’s critique of dialectical reason. Journal of the History of Ideas, 22(4):573–581, October/December 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708033.

Mortimer:2015:LWT

[MS15] Sarah Mortimer and David Scott. Leviathan and the wars of the three kingdoms. Journal of the History of Ideas, 76(2):259–270, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/580298.

Meyer:2013:ECW

[MT13] Thomas Meyer and Bill Templer. Ernst Cassirer’s writings. Journal of the History of Ideas, 74(3):473–495, July 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/511980.

Mudford:1968:WLN

[Mud68] Peter G. Mudford. William Lawrence and the natural history of man. Journal of the History of Ideas, 29(3):430–436, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708453.

Mueller:1943:CIC

[Mue43] Gustav E. Mueller. Calvin’s institutes of the Christian religion as an illustration of Christian thinking. Journal of the History of Ideas, 4(3): 287–300, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707255.

Mueller:1945:PDP

[Mue45] Gustav E. Mueller. Pascal’s dialectical philosophy and his discovery of liberalism. Journal of the History of Ideas, 6(1):67–80, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707056.

Mueller:1958:HLT

[Mue58] Gustav E. Mueller. The Hegel legend of “Thesis-Antithesis-Synthesis”. Journal of the History of Ideas, 19(3):411–414, June 1958.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708045.

Mullett:1958:LCC

[Mul58] Charles F. Mullett. Lucretius in Clio’s chariot. Journal of the History of Ideas, 19(3):307–322, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708037.

Mullins:1963:SCE

[Mul63] Terence Y. Mullins. Sidgwick’s concept of ethical science. Journal of the History of Ideas, 24(4):584–588, October/December 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707988.

Mulhern:1968:NSS

[Mul68a] J. J. Mulhern. A note on stating “The Socratic Paradox”. Journal of the History of Ideas, 29(4):601–604, October/December 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708296.

Mulvaney:1968:EDL

[Mul68b] Robert J. Mulvaney. The early development of Leibniz’s concept of justice. Journal of the History of Ideas, 29(1):53–72, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708465.

Mulhern:1974:ASP

[Mul74] J. J. Mulhern. Aristotle and the Socratic paradoxes. Journal of the History of Ideas, 35(2):293–299, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708763.

Mulgan:1979:LGT

[Mul79] R. G. Mulgan. Lycophron and Greek theories of social contract. Journal of the History of Ideas, 40(1):121–128, January/March 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709263. See reply [Gut79].

Mulligan:1994:RBR

[Mul94] Lotte Mulligan. Robert Boyle, “Right Reason,” and the meaning of metaphor. Journal of the History of Ideas, 55(2):235–257, April 1994.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709898.

Mullin:1996:PPR

[Mul96] Amy Mullin. Purity and pollution: Resisting the rehabilitation of a virtue. Journal of the History of Ideas, 57(3):509–524, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14915.

Mulsow:1997:ESP

[Mul97] Martin Mulsow. Eclecticism or skepticism? A problem of the Early Enlightenment. Journal of the History of Ideas, 58(3):465–477, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14955.

Mulsow:2006:PUP

[Mul06] Martin Mulsow. Practices of unmasking: Polyhistors, correspondence, and the birth of dictionaries of pseudonymity in Seventeenth-Century Germany. Journal of the History of Ideas, 67(2):219–250, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/198078.

Mulsow:2009:CHA

[Mul09] Martin Mulsow. Christian humanism in the age of critical philology: Ralph Häfner’s Gods in Exile. Journal of the History of Ideas, 70(4):659–679, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/361794.

Mulder:2023:NTI

[Mul23] Nicholas Mulder. The neoliberal transition in intellectual and economic history. Journal of the History of Ideas, 84(3):559–583, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/901494.

Mumford:1981:CBP

[Mum81] Richard L. Mumford. Crane Brinton’s pattern and the Chinese Revolution of 1911. Journal of the History of Ideas, 42(4):707–722, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709129.

Munitz:1951:OUM

[Mun51] Milton K. Munitz. One universe or many? Journal of the History of Ideas, 12(2):231–255, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707516.

Mungello:1979:SRS

[Mun79] David E. Mungello. Some recent studies on the confluence of Chinese and Western intellectual history. Journal of the History of Ideas, 40 (4):649–661, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709364.

Mungello:1980:MCP

[Mun80a] David E. Mungello. Malebranche and Chinese philosophy. Journal of the History of Ideas, 41(4):551–578, October/December 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709274.

Munro:1980:CIC

[Mun80b] Donald J. Munro. The concept of “Interest” in Chinese thought. Journal of the History of Ideas, 41(2):179–197, April/June 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709455.

Munz:1990:RR

[Mun90] Peter Munz. The rhetoric of rhetoric. Journal of the History of Ideas, 51(1):121–142, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709750. See reply [McC90].

Murray:1954:MRM

[Mur54] W. A. Murray. Mr. Roland M. Frye’s article on Swift’s Yahoo. Journal of the History of Ideas, 15(4):599–601, October 1954. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707679. See [Fry54].

Murdoch:1958:NFM

[Mur58] Ruth T. Murdoch. Newton and the French muse. Journal of the History of Ideas, 19(3):323–334, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708038.

Murti:1968:BDG

[Mur68] V. V. Ramana Murti. Buber’s dialogue and Gandhi’s satyagraha. Journal of the History of Ideas, 29(4):605–613, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708297.

Myers:1975:ASC

[Mye75] M. L. Myers. Adam Smith as critic of ideas. Journal of the History of Ideas, 36(2):281–296, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708928.

Myers:1993:RCW

[Mye93] D. G. Myers. The rise of creative writing. Journal of the History of Ideas, 54(2):277–297, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709983.

Myers:2008:FWW

[Mye08] Benjamin Myers. “Following the way which is called heresy”: Milton and the heretical imperative. Journal of the History of Ideas, 69(3): 375–393, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
242277.

Nadel:1962:NLB

[Nad62] George H. Nadel. New light on Bolingbroke’s letters on history. Journal of the History of Ideas, 23(4):550–557, October/December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708174.

Nadelhaft:1982:ESC

[Nad82] Jerome Nadelhaft. The Englishwoman’s sexual Civil War: Feminist attitudes towards men, women, and marriage 1650–1740. Journal of the History of Ideas, 43(4):555–579, October/December 1982. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709343.

Nadler:1988:ADT

[Nad88] Steven M. Nadler. Arnauld, Descartes, and transubstantiation: Reconciling Cartesian metaphysics and real presence. Journal of the History of Ideas, 49(2):229–246, April/June 1988. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709498.

Nadler:1991:MVG

[Nad91] Steven Nadler. Malebranche and the vision in God: a note on the search after truth, III, 2, iii. Journal of the History of Ideas, 52(2):309–314, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709530.

Nadelman:1993:BSP

[Nad93] Heather L. Nadelman. Baconian science in Post-Bellum America: Charles Peirce’s “Neglected Argument for the Reality of God”. Journal of the History of Ideas, 54(1):79–96, January 1993. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709861.

Nadler:1994:CTL

[Nad94] Steven Nadler. Choosing a theodicy: The Leibniz–Malebranche– Arnauld connection. Journal of the History of Ideas, 55(4):573–589, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709923.

Nadler:1997:DDM

[Nad97] Steven M. Nadler. Descartes’s demon and the madness of Don Quixote. Journal of the History of Ideas, 58(1):41–55, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14934.

Naddaf:1998:OAC

[Nad98] Gerard Naddaf. On the origin of Anaximander’s cosmological model. Journal of the History of Ideas, 59(1):1–28, January 1998. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/712623.

Nadler:2001:GPJ

[Nad01] Steven M. Nadler. Gersonides on Providence: a Jewish chapter in the history of the general will. Journal of the History of Ideas, 62(1):37–57, January 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15105.

Nadler:2009:JS

[Nad09] Steven Nadler. The Jewish Spinoza. Journal of the History of Ideas, 70(3):491–510, July 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/270108.

Nadler:2013:STP

[Nad13a] Steven Nadler. Scripture and truth: a problem in Spinoza’s Tractatus Theologico-Politicus. Journal of the History of Ideas, 74(4):623– 642, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
523432.

Nadler:2013:SSC

[Nad13b] Steven Nadler. Spinoza and scripture: a colloquium introduction. Journal of the History of Ideas, 74(4):621–622, October 2013. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/523431.

Nadler:2019:SMB

[Nad19] Steven Nadler. Spinoza and menasseh ben Israel: Facts and fictions. Journal of the History of Ideas, 80(4):533–554, October 2019. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/735380.

Nadler:2022:MLR

[Nad22] Steven Nadler. The many lives of René Descartes. Journal of the History of Ideas, 83(3):501–522, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/858860.

Nagel:1957:MRC

[Nag57] Ernest Nagel. Morris R. Cohen in retrospect. Journal of the History of Ideas, 18(4):548–551, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707565.

Nagley:1968:BRS

[Nag68] Winfield E. Nagley. Book review: Søren Kierkegaard’s the Concept of Irony. Journal of the History of Ideas, 29(3):458–464, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708456.

Nahm:1947:TBT

[Nah47] Milton C. Nahm. The theological background of the theory of the artist as creator. Journal of the History of Ideas, 8(3):363–372, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707147.

Najemy:1999:PCM

[Naj99] John M. Najemy. Papirius and the chickens, or Machiavelli on the necessity of interpreting religion. Journal of the History of Ideas, 60(4):659–681, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15053.

Nakamura:1969:BRA

[Nak69] Hajime Nakamura. Book review: Asia in the Making of Europe. Journal of the History of Ideas, 30(3):451–455, July/September 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708569.

Nakhimovsky:2022:GLR

[Nak22] Isaac Nakhimovsky. Georg Lukács and Revolutionary Realpolitik, 1918–19: an essay on ethical action, historical judgment, and the history of political thought. Journal of the History of Ideas, 83(1):63– 85, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
847821.

Nannini:2021:BSP

[Nan21] Alessandro Nannini. At the bottom of the soul: The psychologization of the “Fundus Animae” between Leibniz and Sulzer. Journal of the History of Ideas, 82(1):51–72, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/781612.

Nartonis:2005:LAP

[Nar05] David K. Nartonis. Louis Agassiz and the Platonist story of creation at Harvard, 1795–1846. Journal of the History of Ideas, 66(3):437– 449, July 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/188685.

Nason:1946:LAC

[Nas46] John W. Nason. Leibniz’s attack on the Cartesian doctrine of extension. Journal of the History of Ideas, 7(4):447–483, October 1946.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707034.

Naumann:1952:GR

[Nau52] Walter Naumann. Goethe’s religion. Journal of the History of Ideas, 13(2):188–199, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707610.

Nauert:1957:MSA

[Nau57] Charles G. Nauert, Jr. Magic and skepticism in Agrippa’s thought. Journal of the History of Ideas, 18(2):161–182, April 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707622.

Nauta:2002:HRC

[Nau02] Lodi Nauta. Hobbes on religion and the Church between the elements of law and Leviathan: a dramatic change of direction? Journal of the History of Ideas, 63(4):577–598, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/39296.

Nauta:2011:PPG

[Nau11] Lodi Nauta. Philology as philosophy: Giovanni pontano on language, meaning, and grammar. Journal of the History of Ideas, 72(4):481– 502, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
453545.

Nauta:2015:OKJ

[Nau15] Lodi Nauta. The order of knowing: Juan Luis Vives on language, thought, and the topics. Journal of the History of Ideas, 76(3):325– 345, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
587268.

Nederman:1988:NSO

[Ned88] Cary J. Nederman. Nature, sin and the origins of society: The Ciceronian tradition in medieval political thought. Journal of the History of Ideas, 49(1):3–26, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709701.

Nederman:1991:AOP

[Ned91] Cary J. Nederman. Aristotelianism and the origins of “Political Science” in the Twelfth Centuryi. Journal of the History of Ideas, 52(2):179–194, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709524.

Nederman:1996:MAM

[Ned96] Cary J. Nederman. The meaning of “Aristotelianism” in medieval moral and political thought. Journal of the History of Ideas, 57 (4):563–585, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14925.

Nederman:1999:AGF

[Ned99] Cary J. Nederman. Amazing grace: Fortune, God, and free will in Machiavelli’s thought. Journal of the History of Ideas, 60(4):617–638, October 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15054.

Nederman:2005:EHE

[Ned05] Cary J. Nederman. Empire and the historiography of European political thought: Marsiglio of Padua, Nicholas of Cusa, and the medieval/ modern divide. Journal of the History of Ideas, 66(1):1–15, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/185261.

Nederman:2013:ALL

[Ned13] Cary J. Nederman. 2012 Arthur O. Lovejoy Lecture: Civil religion — metaphysical, not political: Nature, faith, and communal order in European thought, c. 1150–c. 1550. Journal of the History of Ideas, 74(1):1–22, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/495239.

Nederman:2024:MC

[Ned24] Cary J. Nederman. Marcia L. Colish (1937–2024). Journal of the History of Ideas, 85(3):iii–v, July 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/933853.

Needham:1951:HLLa

[Nee51a] Joseph Needham. Human laws and laws of nature in China and the West (i). Journal of the History of Ideas, 12(1):3–30, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707533.

Needham:1951:HLLb

[Nee51b] Joseph Needham. Human laws and laws of nature in China and the West (II): Chinese civilization and the laws of nature. Journal of the History of Ideas, 12(2):194–230, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2707515. See corrigenda [Ano51a].

Negley:1979:BRR

[Neg79] Glenn Negley. Book review: Recent Works on Utopian Literature. Journal of the History of Ideas, 40(2):315–320, April/June 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709155.

Neher:1977:CHL

[Neh77] André Neher. Copernicus in the Hebraic literature from the Sixteenth to the Eighteenth Century. Journal of the History of Ideas, 38(2):211–226, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708908.

Neill:1948:QP

[Nei48] Thomas P. Neill. Quesnay and physiocracy. Journal of the History of Ideas, 9(2):153–173, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707226.

Nelson:1972:LWC

[Nel72] Alvin F. Nelson. Lester Ward’s conception of the nature of science. Journal of the History of Ideas, 33(4):633–638, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708862.

Nelson:1976:AHP

[Nel76] Nicolas H. Nelson. Astrology, Hudibras, and the Puritans. Journal of the History of Ideas, 37(3):521–536, July/September 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708813.

Nelson:1987:CBR

[Nel87] Richard Nelson. Carl Becker revisited: Irony and progress in history. Journal of the History of Ideas, 48(2):307–323, April/June 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709560.

Nelles:2000:SBB

[Nel00] Paul Neave Nelles. Sainte-Beuve between Renaissance and Enlightenment. Journal of the History of Ideas, 61(3):473–492, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15083.

Neumeyer:1947:LGS

[Neu47] Eva Maria Neumeyer. The landscape garden as a symbol in Rousseau, Goethe and Flaubert. Journal of the History of Ideas, 8(2):187–217, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707155.

Neubauer:1967:DJB

[Neu67] John Neubauer. Dr. John Brown (1735–88) and Early German Romanticism. Journal of the History of Ideas, 28(3):367–382, July/ September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708624.

Neufeldt:1977:SPE

[Neu77] Leonard N. Neufeldt. The science of power: Emerson’s views on science and technology in America. Journal of the History of Ideas, 38(2):329–344, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708917.

Neuber:2025:CNP

[Neu25] Matthias Neuber. Consciousness in neorealism: Perry, Montague, and Holt. Journal of the History of Ideas, 86(1):141–168, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/949930.

Neville:2009:GEM

[Nev09] Kristoffer Neville. Gothicism and early modern historical ethnography. Journal of the History of Ideas, 70(2):213–234, April 2009. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/261976.

Newman:1995:HAR

[New95] Edgar Leon Newman. The historian as apostle: Romanticism, religion, and the first socialist history of the world. Journal of the History of Ideas, 56(2):239–261, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709837.

Newberry:2001:JBF

[New01] Paul A. Newberry. Joseph Butler on forgiveness: a presupposed theory of emotion. Journal of the History of Ideas, 62(2):233–244, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15113.

Nicolson:1940:KSJ

[Nic40] Marjorie Nicolson. Kepler, the Somnium, and John Donne. Journal of the History of Ideas, 1(3):259–280, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707087.

Nicolson:1948:LT

[Nic48] Marjorie Nicolson. A. O. Lovejoy as teacher. Journal of the History of Ideas, 9(4):428–438, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707216.

Nicolson:1968:WPD

[Nic68] Marjorie H. Nicolson. Ward’s “Pill and Drop” and men of letters. Journal of the History of Ideas, 29(2):177–196, April/June 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708575.

Nicholson:1990:ECS

[Nic90] Mervyn Nicholson. The eleventh commandment: Sex and spirit in wollstonecraft and Malthus. Journal of the History of Ideas, 51 (3):401–421, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709622.

Nicholls:2019:SGJ

[Nic19] Sophie Nicholls. Sovereignty and government in Jean Bodin’s Six Livres de la République (1576). Journal of the History of Ideas, 80(1):47–66, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/715735.

Nieli:1986:SIA

[Nie86] Russell Nieli. Spheres of intimacy and the Adam Smith problem. Journal of the History of Ideas, 47(4):611–624, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709721.

Nisbet:1944:BCS

[Nis44] Robert A. Nisbet. De Bonald and the concept of the social group. Journal of the History of Ideas, 5(3):315–331, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707353.

Niskanen:2021:APP

[Nis21] Samu Niskanen. Anselm’s predicament: The proslogion and antiintellectual rhetoric in the aftermath of the Berengarian controversy. Journal of the History of Ideas, 82(4):547–568, October 2021. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/837671.

Nordin:2020:NDF

[NL20] Jonas Nordin and John Christian Laursen. Northern declarations of freedom of the press: The relative importance of philosophical ideas and of local politics. Journal of the History of Ideas, 81(2):217– 237, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
753535.

Nobbs:1965:PIW

[Nob65] Douglas Nobbs. The political ideas of William Cleghorn, Hume’s academic rival. Journal of the History of Ideas, 26(4):575–586, October/ December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708501.

Noland:1961:IJJ

[Nol61] Aaron Noland. Individualism in Jean Jaures’ socialist thought. Journal of the History of Ideas, 22(1):63–80, January/March 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707874.

Noland:1967:PR

[Nol67] Aaron Noland. Proudhon and Rousseau. Journal of the History of Ideas, 28(1):33–54, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708479.

Nolte:1984:IDJ

[Nol84] Sharon H. Nolte. Industrial democracy for Japan: Tanaka Ōdō and John Dewey. Journal of the History of Ideas, 45(2):277–294, April/ June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709292.

Noone:1972:RTN

[Noo72] John B. Noone, Jr. Rousseau’s theory of natural law as conditional. Journal of the History of Ideas, 33(1):23–42, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709055.

Noon:1978:S

[Noo78] Georgia Noon. On suicide. Journal of the History of Ideas, 39 (3):371–386, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709383.

Northrop:1946:LTS

[Nor46] F. S. C. Northrop. Leibniz’s theory of space. Journal of the History of Ideas, 7(4):422–446, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707033.

Norwood:1947:AAN

[Nor47] Frederick A. Norwood. Attitude of the ante-Nicene fathers toward Greek artistic achievement. Journal of the History of Ideas, 8(4):431– 448, October 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707428.

North:1974:BRA

[Nor74] Helen F. North. Book review: Ancient Salt: The New Rhetoric and the Old. Journal of the History of Ideas, 35(2):349–356, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708769.

Norton:1975:GTF

[Nor75] David Fate Norton. George Turnbull and the furniture of the mind. Journal of the History of Ideas, 36(4):701–716, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708997.

Norford:1977:MMS

[Nor77] Don Parry Norford. Microcosm and macrocosm in Seventeenth-Century literature. Journal of the History of Ideas, 38(3):409– 428, July/September 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708672.

Norton:2001:JLD

[Nor01] David Fate Norton. From John Locke to Dugald Stewart. Journal of the History of Ideas, 62(2):359–365, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15114.

Norman:2002:NER

[Nor02] Judith Norman. Nietzsche and early romanticism. Journal of the History of Ideas, 63(3):501–519, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15164.

North:2004:MJF

[Nor04] Helen F. North. In memoriam: John F. Callahan. Journal of the History of Ideas, 65(1):155–157, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/168337.

Norton:2007:MCE

[Nor07] Robert Edward Norton. The myth of the counter-enlightenment. Journal of the History of Ideas, 68(4):635–658, October 2007. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/223131. See re-

sponse [Les07].

Norton:2008:IBE

[Nor08] Robert E. Norton. Isaiah Berlin’s “Expressionism,” or: “Ha! Du bist das Blökende!”. Journal of the History of Ideas, 69(2):339–347, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/235874.

Notestein:1955:MRW

[Not55] Robert B. Notestein. The moralist rigorism of W. G. Sumner. Journal of the History of Ideas, 16(3):389–400, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707639.

Nothaft:2011:SSA

[Not11] Carl Philipp Emanuel Nothaft. From Sukkot to Saturnalia: The attack on Christmas in Sixteenth-Century chronological scholarship. Journal of the History of Ideas, 72(4):503–522, October 2011. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/453546.

Nothaft:2016:WOE

[Not16] Carl Philipp Emanuel Nothaft. Walter Odington’s De etate mundi and the pursuit of a scientific chronology in Medieval England. Journal of the History of Ideas, 77(2):183–201, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/620978.

Nygaard:2016:FAN

[NS16] Stefan Nygård and Johan Strang. Facing asymmetry: Nordic intellectuals and center–periphery dynamics in European cultural space. Journal of the History of Ideas, 77(1):75–97, January 2016. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/610561.

Norton:1980:TRA

[NSR80] David Fate Norton and J. C. Stewart-Robertson. Thomas Reid on Adam Smith’s theory of morals. Journal of the History of Ideas, 41 (3):381–398, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709400.

Nurser:1961:RCL

[Nur61] John S. Nurser. The religious conscience in Lord Acton’s political thought. Journal of the History of Ideas, 22(1):47–62, January/March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707873.

Nye:1979:BCP

[Nye79] Mary Jo Nye. The Boutroux Circle and Poincaré’s conventionalism. Journal of the History of Ideas, 40(1):107–120, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709262.

Oake:1953:MRI

[Oak53] Roger B. Oake. Montesquieu’s religious ideas. Journal of the History of Ideas, 14(4):548–560, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707701.

Oake:1955:MAR

[Oak55] Roger B. Oake. Montesquieu’s analysis of Roman history. Journal of the History of Ideas, 16(1):44–59, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707526.

Oakley:1966:CR

[Oak66] Francis Oakley. From constance to 1688 revisited. Journal of the History of Ideas, 27(3):429–432, July/September 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708595.

Oakley:1968:JPT

[Oak68] Francis Oakley. Jacobean political theology: The absolute and ordinary powers of the king. Journal of the History of Ideas, 29 (3):323–346, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708446.

Oakley:1987:LUO

[Oak87] Francis Oakley. Lovejoy’s unexplored option. Journal of the History of Ideas, 48(2):231–245, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2709556. See erratum [Ano87e].

Oakley:1998:AOPb

[Oak98a] Francis Oakley. The absolute and ordained power of God and King in the Sixteenth and Seventeenth Centuries: Philosophy, science, politics, and law. Journal of the History of Ideas, 59(4):669–690, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15009.

Oakley:1998:AOPa

[Oak98b] Francis Oakley. The absolute and ordained power of God in Sixteenthand Seventeenth-Century theology. Journal of the History of Ideas, 59(3):437–461, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14995.

OBrien:1958:MPP

[O’B58] Michael O’Brien. Modern philosophy and platonic ethics. Journal of the History of Ideas, 19(4):451–472, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707917.

OBrien:1985:JCT

[O’B85] Charles H. O’Brien. The Jansenist campaign for toleration of Protestants in Late Eighteenth-Century France: Sacred or secular? Journal of the History of Ideas, 46(4):523–538, October/December 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709543.

OBrien:2008:ATR

[O’B08] Michael O’Brien. Amoralities not for turning: Response to cotkin. Journal of the History of Ideas, 69(2):323–326, April 2008. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/235875.

Oberg:1976:DHA

[Obe76] Barbara Bowen Oberg. David Hartley and the association of ideas. Journal of the History of Ideas, 37(3):441–454, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708808.

Oberman:1987:AML

[Obe87] Heiko A. Oberman. Via Antiqua and Via Moderna: Late medieval prolegomena to early reformation thought. Journal of the History of Ideas, 48(1):23–40, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2709610. See comment [Gil87].

OConnor:1964:PDF

[O’C64] Daniel D. O’Connor. Peirce’s debt to F. E. Abbot. Journal of the History of Ideas, 25(4):543–564, October/December 1964. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708185.

OCathasaigh:1984:SBP

[O’C84] Sean O’Cathasaigh. Skepticism and belief in Pierre Bayle’s Nouvelles Lettres Critiques. Journal of the History of Ideas, 45(3): 421–433, July/September 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709233.

Odegard:1965:LEV

[Ode65] Douglas Odegard. Locke’s epistemology and the value of experience. Journal of the History of Ideas, 26(3):417–423, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708390.

Odom:1966:ECM

[Odo66] Herbert H. Odom. The estrangement of celestial mechanics and religion. Journal of the History of Ideas, 27(4):533–548, October/ December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708339.

OFlaherty:1992:WHN

[O’F92] James C. O’Flaherty. Werner Heisenberg on the Nazi revolution: Three hitherto unpublished letters. Journal of the History of Ideas, 53(3):487–494, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709890.

Ogden:1949:PVC

[Ogd49] H. V. S. Ogden. The principles of variety and contrast in Seventeenth Century aesthetics, and Milton’s poetry. Journal of the History of Ideas, 10(2):159–182, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707413.

Ogden:1974:SAD

[Ogd74] John T. Ogden. From spatial to aesthetic distance in the Eighteenth Century. Journal of the History of Ideas, 35(1):63–78, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708743.

Ogilvie:2003:MBN

[Ogi03] Brian W. Ogilvie. The many books of nature: Renaissance naturalists and information overload. Journal of the History of Ideas, 64(1):29–40, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/42176.

Olsen:1988:CMD

[OH88] Mark Olsen and Louis-Georges Harvey. Contested methods: Daniel T. Rodgers’s contested truths. Journal of the History of Ideas, 49 (4):653–668, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709678.

Olivecrona:1974:ASN

[Oli74] Karl Olivecrona. Appropriation in the state of nature: Locke on the origin of property. Journal of the History of Ideas, 35(2):211–230, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708759.

Olssen:1968:MR

[Ols68] E. A. Olssen. Marx and the resurrection. Journal of the History of Ideas, 29(1):131–140, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708471.

Olson:1971:SPM

[Ols71] Richard Olson. Scottish philosophy and mathematics 1750–1830. Journal of the History of Ideas, 32(1):29–44, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708323.

Olshewsky:1991:CRH

[Ols91] Thomas M. Olshewsky. The classical roots of Hume’s skepticism. Journal of the History of Ideas, 52(2):269–287, April/June 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709528.

ONeill:2007:JAV

[O’N07] Daniel I. O’Neill. John Adams versus Mary Wollstonecraft on the French Revolution and democracy. Journal of the History of Ideas, 68(3):451–476, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/219821.

ONeill:2012:RMW

[O’N12] Daniel I. O’Neill. Revisiting the middle way: The logic of the history of ideas after more than a decade. Journal of the History of Ideas, 73(4):583–592, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/488349.

Ong:1953:PRN

[Ong53] Walter J. Ong. Peter Ramus and the naming of Methodism: Medieval science through Ramist homiletic. Journal of the History of Ideas, 14(2):235–248, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707473.

Oosterhoff:2013:PPP

[Oos13] Richard J. Oosterhoff. From pious to polite: Pythagoras in the Res publica litterarum of French Renaissance mathematics. Journal of the History of Ideas, 74(4):531–552, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/523427.

Ossa-Richardson:2013:PIT

[OR13] Anthony Ossa-Richardson. Possession or insanity?: Two views from the Victorian lunatic asylum. Journal of the History of Ideas, 74 (4):553–575, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/523428.

Orru:1986:ARE

[Orr86] Marco Orrù. Anomy and reason in the English Renaissance. Journal of the History of Ideas, 47(2):177–196, April/June 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709809.

Orr:2014:SJD

[Orr14] D. Alan Orr. Sir John Davies’s agrarian law for Ireland. Journal of the History of Ideas, 75(1):91–112, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/535155.

Orsini:1969:FSO

[Ors69] G. N. G. Orsini. Feuerbach’s supposed objection to Hegel. Journal of the History of Ideas, 30(1):85–90, January/March 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708246. See reply [Ber69].

Ortega:2018:CHI

[Ort18] Francisco A. Ortega. The conceptual history of independence and the colonial question in Spanish America. Journal of the History of Ideas, 79(1):89–103, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/684834.

OSullivan:2014:MOL

[O’S14] Luke O’Sullivan. Michael Oakeshott and the Left. Journal of the History of Ideas, 75(3):471–492, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/549331.

Olsen:2022:LLM

[OS22] Niklas Olsen and Quinn Slobodian. Locating Ludwig von Mises: Introduction. Journal of the History of Ideas, 83(2):257–267, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/855166.

Osborne:2002:FPN

[Osb02] Thomas M. Osborne. Faith, philosophy, and the nominalist background to Luther’s defense of the real presence. Journal of the History of Ideas, 63(1):63–82, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15145.

Osler:1970:JLC

[Osl70] Margaret J. Osler. John Locke and the changing ideal of scientific knowledge. Journal of the History of Ideas, 31(1):3–16, January/ March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708366.

Osler:1979:DCN

[Osl79] Margaret J. Osler. Descartes and Charleton on nature and God. Journal of the History of Ideas, 40(3):445–456, July/September 1979.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709247.

Osler:1983:PDW

[Osl83] Margaret J. Osler. Providence and divine will in Gassendi’s views on scientific knowledge. Journal of the History of Ideas, 44(4): 549–560, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709214.

Osler:1985:ETL

[Osl85] Margaret J. Osler. Eternal truths and the laws of nature: The theological foundations of Descartes’ philosophy of nature. Journal of the History of Ideas, 46(3):349–362, July/September 1985. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709472.

Osmond:1973:BSM

[Osm73] Rosalie E. Osmond. Body, soul, and the marriage relationship: The history of an analogy. Journal of the History of Ideas, 34(2):283–290, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708732.

Otis:2002:MCO

[Oti02] Laura Otis. The metaphoric circuit: Organic and technological communication in the Nineteenth Century. Journal of the History of Ideas, 63(1):105–128, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15146.

Owens:1975:MA

[Owe75] Robert R. Owens. The myth of anian. Journal of the History of Ideas, 36(1):135–138, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709015.

Oxtoby:2001:JC

[Oxt01] Willard Gurdon Oxtoby. Julia Ching, 1934–2001. Journal of the History of Ideas, 62(4):745–746, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15134.

Ozavci:2014:DIC

[Öza14] Hilmi Ozan Özavcı. Differing interpretations of la conscience collective and “the individual” in Turkey: Émile Durkheim and the intellectual origins of the Republic. Journal of the History of Ideas, 75(1):113– 136, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
535156.

Pace:1979:SCH

[Pac79] K. Claire Pace. “Strong contraries… happy discord”: Some Eighteenth-Century discussions about landscape. Journal of the History of Ideas, 40(1):141–155, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709267.

Packer:1989:KDM

[Pac89] Mark Packer. Kant on desire and moral pleasure. Journal of the History of Ideas, 50(3):429–442, July/September 1989. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709570.

Packham:2002:PPE

[Pac02] Catherine Packham. The physiology of political economy: Vitalism and Adam Smith’s wealth of nations. Journal of the History of Ideas, 63(3):465–481, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15159.

Pagden:1988:BRR

[Pag88] Anthony Pagden. Book review: Rethinking the Linguistic Turn: Current Anxieties in Intellectual History. Journal of the History of Ideas, 49(3):519–529, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709491. See review [LaC88].

Page:1996:SMI

[Pag96] Carl Page. Symbolic mathematics and the intellect militant: On modern philosophy’s revolutionary spirit. Journal of the History of Ideas, 57(2):233–253, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14906.

Palmer:1940:NIF

[Pal40] Robert R. Palmer. The national idea in France before the revolution. Journal of the History of Ideas, 1(1):95–111, January 1940. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707012.

Paludan:1972:LFR

[Pal72] Phillip S. Paludan. Law and the failure of reconstruction: The case of Thomas Cooley. Journal of the History of Ideas, 33(4):597–614, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708859.

Palti:1997:MHB

[Pal97] Elías José Palti. In memoriam: Hans Blumenberg (1920–1996), an unended quest. Journal of the History of Ideas, 58(3):503–524, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14956.

Palti:2006:PMI

[Pal06] Elías José Palti. The problem of “Misplaced ideas” revisited: Beyond the “History of ideas” in Latin America. Journal of the History of Ideas, 67(1):149–179, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/194648.

Palti:2009:BRB

[Pal09] Elías José Palti. Beyond revisionism: The bicentennial of independence, the early republican experience, and intellectual history in Latin America. Journal of the History of Ideas, 70(4):593–614, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/361797.

Palmer:2012:RLR

[Pal12] Ada Palmer. Reading Lucretius in the Renaissance. Journal of the History of Ideas, 73(3):395–416, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/481417.

Palti:2018:BHI

[Pal18a] Elías Palti. Beyond the “History of Ideas”: The issue of the “Ideological Origins of the Revolutions of Independence” revisited. Journal of the History of Ideas, 79(1):125–141, January 2018. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/684836.

Palti:2018:RHI

[Pal18b] Elías Palti. Revising history: Introduction to the symposium on the bicentennial of the Latin American Revolutions of independence. Journal of the History of Ideas, 79(1):65–71, January 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/684832.

Pankakoski:2021:WPA

[Pan21] Timo Pankakoski. Wartime pamphlets, anti-English metaphors, and the intensification of antidemocratic discourse in Germany after the First World War. Journal of the History of Ideas, 82(2):279–304, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/790204.

Pappe:1964:MT

[Pap64] H. O. Pappe. Mill and Tocqueville. Journal of the History of Ideas, 25(2):217–234, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708013.

Pappe:1979:SSL

[Pap79] H. O. Pappé. Sismondi’s system of liberty. Journal of the History of Ideas, 40(2):251–266, April/June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709151.

Parsons:1944:TDS

[Par44] Talcott Parsons. The theoretical development of the sociology of religion: a chapter in the history of modern social science. Journal of the History of Ideas, 5(2):176–190, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707383.

Parsons:1962:BRH

[Par62] Talcott Parsons. Book review: Hurst’s Law and Social Process in U. S. History. Journal of the History of Ideas, 23(4):558–564, October/ December 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708175.

Parry:1965:NOT

[Par65] Adam Parry. A note on the origins of teleology. Journal of the History of Ideas, 26(2):259–262, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708231.

Paret:1968:EPW

[Par68a] Peter Paret. Education, politics, and war in the life of Clausewitz. Journal of the History of Ideas, 29(3):394–408, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708450.

Parsons:1968:CHC

[Par68b] Coleman O. Parsons. The classical and humanist context of Paradise Lost, II, 496–505. Journal of the History of Ideas, 29(1):33–52, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708464.

Park:1969:LBM

[Par69a] Desiree Park. Locke and Berkeley on the Molyneux problem. Journal of the History of Ideas, 30(2):253–260, April/June 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708437.

Park:1969:HB

[Par69b] Roy Park. Hazlitt and Bentham. Journal of the History of Ideas, 30 (3):369–384, July/September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708563.

Parssinen:1973:TSO

[Par73] T. M. Parssinen. Thomas Spence and the origins of English land nationalization. Journal of the History of Ideas, 34(1):135–141, January/ March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708949.

Parkinson:1977:HPS

[Par77] G. H. R. Parkinson. Hegel, pantheism, and Spinoza. Journal of the History of Ideas, 38(3):449–459, July/September 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708674.

Paret:1988:CDS

[Par88a] Peter Paret. Continuity and discontinuity in some interpretations by Tocqueville and Clausewitz. Journal of the History of Ideas, 49 (1):161–169, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709710.

Partenheimer:1988:EHA

[Par88b] David Partenheimer. The education of Henry Adams in German philosophy. Journal of the History of Ideas, 49(2):339–345, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709504.

Paret:1991:OFG

[Par91] Peter Paret. Obituary: Felix Gilbert 1905–1991. Journal of the History of Ideas, 52(2):341, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709535.

Parsons:2001:MSE

[Par01] Jotham Parsons. Money and sovereignty in Early Modern France. Journal of the History of Ideas, 62(1):59–79, January 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15106.

Paret:2004:IAJ

[Par04] Peter Paret. From ideal to ambiguity: Johannes von Muller, Clausewitz, and the people in arms. Journal of the History of Ideas, 65 (1):101–111, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/168339.

Parsons:2007:DHI

[Par07] Jotham Parsons. Defining the history of ideas. Journal of the History of Ideas, 68(4):683–699, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/223132.

Park:2008:RBV

[Par08] Katharine Park. Response to Brian Vickers, “Francis Bacon, feminist historiography, and the dominion of nature”. Journal of the History of Ideas, 69(1):143–146, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/231013. See [Vic08].

Parkin:2015:HRL

[Par15] Jon Parkin. Hobbes and the reception of Leviathan. Journal of the History of Ideas, 76(2):289–300, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/580300.

Parasher:2024:MRP

[Par24] Tejas Parasher. M. N. Roy and the problem of parliamentary democracy. Journal of the History of Ideas, 85(3):601–626, July 2024. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/933860.

Passmore:1975:TA

[Pas75] John Passmore. The treatment of animals. Journal of the History of Ideas, 36(2):195–218, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708924.

Pasqualucci:1990:HMF

[Pas90] Paolo Pasqualucci. Hobbes and the myth of “Final War”. Journal of the History of Ideas, 51(4):647–657, October/December 1990. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709650.

Patrides:1958:RIM

[Pat58] C. A. Patrides. Renaissance ideas on Man’s upright form. Journal of the History of Ideas, 19(2):256–258, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707939.

Patrides:1959:RTC

[Pat59] C. A. Patrides. Renaissance thought on the celestial hierarchy: The decline of a tradition. Journal of the History of Ideas, 20(2):155–166, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707816.

Patrides:1962:RVV

[Pat62] C. A. Patrides. Renaissance views on the “Vnconfused Orders Angellick”. Journal of the History of Ideas, 23(2):265–267, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708159.

Patrides:1964:MA

[Pat64] C. A. Patrides. Milton and Arianism. Journal of the History of Ideas, 25(3):423–429, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707908.

Patrides:1967:SS

[Pat67] C. A. Patrides. The salvation of Satan. Journal of the History of Ideas, 28(4):467–478, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708524.

Patey:1986:JRB

[Pat86] Douglas Lane Patey. Johnson’s refutation of Berkeley: Kicking the stone again. Journal of the History of Ideas, 47(1):139–145, January/ March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709600.

Patellis:2001:HEU

[Pat01] Ioli Patellis. Hobbes on explanation and understanding. Journal of the History of Ideas, 62(3):445–462, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15124.

Paul:1971:MIR

[Pau71] Charles B. Paul. Music and ideology: Rameau, Rousseau, and 1789. Journal of the History of Ideas, 32(3):395–410, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708354.

Paul:1972:PDS

[Pau72] H. W. Paul. Pierre Duhem: Science and the Historian’s craft. Journal of the History of Ideas, 33(3):497–512, July/September 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709049.

Paul:1979:WSJ

[Pau79] Ellen Frankel Paul. W. Stanley Jevons: Economic revolutionary, political utilitarian. Journal of the History of Ideas, 40(2):267–283, April/June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709152.

Paul:1981:ICM

[Pau81] Diane Paul. “In the interests of civilization”: Marxist views of race and culture in the Nineteenth Century. Journal of the History of Ideas, 42(1):115–138, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709420.

Paul:1983:HSH

[Pau83] Ellen Frankel Paul. Herbert Spencer: The historicist as a failed prophet. Journal of the History of Ideas, 44(4):619–638, October/ December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709219.

Paul:1984:EL

[Pau84] Diane Paul. Eugenics and the Left. Journal of the History of Ideas, 45 (4):567–590, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709374.

Paxman:1993:LDP

[Pax93] David B. Paxman. Language and difference: The problem of abstraction in Eighteenth-Century language study. Journal of the History of Ideas, 54(1):19–36, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709858.

Patrizi:1943:PS

[PB43] Francesco Patrizi and Benjamin Brickman. On physical space. Journal of the History of Ideas, 4(2):224–245, April 1943. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707326.

Parekh:1973:HPI

[PB73] Bhikhu Parekh and R. N. Berki. The history of political ideas: a critique of Q. Skinner’s methodology. Journal of the History of Ideas, 34(2):163–184, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708724.

Parsons:1963:HRC

[PD63] R. Parsons and Giorgio Del Vecchio. On the history of the red cross. Journal of the History of Ideas, 24(4):577–583, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707987.

Pearce:1948:NMH

[Pea48] Roy Harvey Pearce. A note on method in the history of ideas. Journal of the History of Ideas, 9(3):372–379, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707375.

Pearce:1952:RMI

[Pea52] Roy Harvey Pearce. The “Ruines of Mankind”: The Indian and the Puritan mind. Journal of the History of Ideas, 13(2):200–217, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707611.

Pearson:1954:RCP

[Pea54] Lionel Pearson. Real and conventional personalities in Greek history. Journal of the History of Ideas, 15(1):136–145, January 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707653.

Peach:1980:HM

[Pea80] Bernard Peach. Hume’s mistake. Journal of the History of Ideas, 41(2):331–334, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709465.

Pearce:1994:LBN

[Pea94] Colin D. Pearce. Lord Brougham’s neo-paganism. Journal of the History of Ideas, 55(4):651–670, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709927.

Pearce:2014:BLR

[Pea14] Kenneth L. Pearce. Berkeley’s Lockean religious epistemology. Journal of the History of Ideas, 75(3):417–438, July 2014. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/549329.

Pearce:2015:SOP

[Pea15] Trevor Pearce. “Science organized”: Positivism and the metaphysical club, 1865–1875. Journal of the History of Ideas, 76(3):441–465, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/587273.

Peck:1996:HGT

[Pec96] Linda Levy Peck. Hobbes on the Grand Tour: Paris, Venice, or London? Journal of the History of Ideas, 57(1):177–183, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14892.

Pekacz:1999:SPO

[Pek99] Jolanta T. Pekacz. The Salonnières and the philosophies in Old Regime France: The authority of aesthetic judgment. Journal of the History of Ideas, 60(2):277–297, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15029.

Perkins:1957:CTW

[Per57] Merle L. Perkins. Civil theology in the writings of the abbé De Saint-Pierre. Journal of the History of Ideas, 18(2):242–253, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707627.

Perl:1966:NJL

[Per66] Margula R. Perl. Newton’s justification of the laws of motion. Journal of the History of Ideas, 27(4):585–592, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708342.

Perl:1969:PMN

[Per69] Margula R. Perl. Physics and metaphysics in Newton, Leibniz, and Clarke. Journal of the History of Ideas, 30(4):507–526, October/ December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708608.

Perkins:2002:VRC

[Per02] Franklin Perkins. Virtue, reason, and cultural exchange: Leibniz’s praise of Chinese morality. Journal of the History of Ideas, 63(3):447– 464, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15165.

Perry:2008:TTT

[Per08] Lewis Perry. Turn, turn, turn: Response to cotkin. Journal of the History of Ideas, 69(2):333–337, April 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/235876.

Person:2018:JRW

[Per18] John Person. Japanese right-wing discourse in international context: Minoda Muneki’s interwar writings on class and nation. Journal of the History of Ideas, 79(4):635–657, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/707776.

Pesic:2014:FBV

[Pes14] Peter Pesic. Francis Bacon, violence, and the motion of liberty: The Aristotelian background. Journal of the History of Ideas, 75(1):69– 90, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
535154.

Peterson:1955:BPM

[Pet55] Sven R. Peterson. Benjamin Peirce: Mathematician and philosopher. Journal of the History of Ideas, 16(1):89–112, January 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707529.

Peterson:1984:CRP

[Pet84] Susan Rae Peterson. The compatibility of Richard Price’s politics and ethics. Journal of the History of Ideas, 45(4):537–547, October/ December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709372.

Petersen:1985:TSA

[Pet85] Calvin R. Petersen. Time and stress: Alice in Wonderland. Journal of the History of Ideas, 46(3):427–433, July/September 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709477.

Peterson:1988:SCM

[Pet88] Willard J. Peterson. Squares and circles: Mapping the history of Chinese thought. Journal of the History of Ideas, 49(1):47–60, January/ March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709703.

Peterfreund:1994:SMO

[Pet94] Stuart Peterfreund. Scientific models in optics: From metaphor to metonymy and back. Journal of the History of Ideas, 55(1):59–73, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709953.

Peters:2001:DKS

[Pet01] Edward Peters. The desire to know the secrets of the world. Journal of the History of Ideas, 62(4):593–610, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15135.

Pettit:2012:FHO

[Pet12] Philip Pettit. Freedom in Hobbes’s ontology and semantics: a comment on Quentin Skinner. Journal of the History of Ideas, 73(1):111– 126, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/

463141. See [Ski12].

Petrescu:2015:CMS

[Pet15] Lucian Petrescu. Cartesian meteors and scholastic meteors: Descartes against the school in 1637. Journal of the History of Ideas, 76(1):25– 45, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
566375.

Peyre:1949:IEC

[Pey49] Henri Peyre. The influence of Eighteenth Century ideas on the French Revolution. Journal of the History of Ideas, 10(1):63–87, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707200.

Pfau:1990:ITF

[Pfa90] Thomas Pfau. Immediacy and the text: Friedrich Schleiermacher’s theory of style and interpretation. Journal of the History of Ideas, 51(1):51–73, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709746.

Pfizenmaier:1997:WIN

[Pfi97] Thomas C. Pfizenmaier. Was Isaac Newton an Arian? Journal of the History of Ideas, 58(1):57–80, January 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/118368;   https:/
/muse.jhu.edu/pub/1/article/14935.

Phillips:1952:BRF

[Phi52] Dayton Phillips. Book review: Ferguson’s History of the Periodic Conception of the Renaissance. Journal of the History of Ideas, 13(2): 266–280, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707616.

Phillips:1953:CBE

[Phi53] Harlan B. Phillips. Charles Beard: The English Lectures, 1899– 1901. Journal of the History of Ideas, 14(3):451–456, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707812.

Phillips:1970:OLN

[Phi70] D. C. Phillips. Organicism in the Late Nineteenth and Early Twentieth Centuries. Journal of the History of Ideas, 31(3):413–432, July/ September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708514.

Phillips:1971:JDR

[Phi71] D. C. Phillips. James, Dewey, and the reflex arc. Journal of the History of Ideas, 32(4):555–568, October/December 1971. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708977.

Phillips:1983:DWP

[Phi83] Mark Phillips. The disenchanted witness: Participation and alienation in Florentine historiography. Journal of the History of Ideas, 44(2):191–206, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709135.

Phillips:1989:MSL

[Phi89] Mark Phillips. Macaulay, Scott, and the literary challenge to historiography. Journal of the History of Ideas, 50(1):117–133, January/ March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709789.

Phillips:1996:RHA

[Phi96] Mark Phillips. Reconsiderations on history and antiquarianism: Arnaldo momigliano and the historiography of Eighteenth-Century Britain. Journal of the History of Ideas, 57(2):297–316, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14907.

Phillips:2007:OWO

[Phi07] Paul T. Phillips. One world, one faith: The quest for unity in Julian Huxley’s religion of evolutionary humanism. Journal of the History of Ideas, 68(4):613–633, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/223133.

Pickering:1989:NEL

[Pic89] Mary Pickering. New evidence of the link between Comte and German philosophy. Journal of the History of Ideas, 50(3):443–463, July/ September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709571.

Pietarinen:2009:SOA

[Pie09] Ahti-Veikko Pietarinen. Significs and the origins of analytic philosophy. Journal of the History of Ideas, 70(3):467–490, July 2009. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/270107.

Pierson:2010:RPG

[Pie10] Chris Pierson. The reluctant pirate: Godwin, justice, and property. Journal of the History of Ideas, 71(4):569–591, October 2010. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/398298.

Pimenta:2018:HCH

[Pim18] João Paulo Pimenta. History of concepts and the historiography of the independence of Brazil: a preliminary diagnosis. Journal of the History of Ideas, 79(1):157–168, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/684838.

Pine:1968:PPD

[Pin68] Martin Pine. Pomponazzi and the problem of “Double Truth”. Journal of the History of Ideas, 29(2):163–176, April/June 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708574.

Piper:1959:PSC

[Pip59] Herbert Piper. The pantheistic sources of Coleridge’s early poetry. Journal of the History of Ideas, 20(1):47–59, January 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707966.

Pirillo:2025:HKT

[Pir25] Diego Pirillo. How knowledge travels: Learned periodicals and the Atlantic Republic of Letters. Journal of the History of Ideas, 86(1):75– 107, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
949928.

Pitts:1986:SFE

[Pit86] Edward I. Pitts. Spinoza on freedom of expression. Journal of the History of Ideas, 47(1):21–35, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709593.

Plassart:2010:SJJ

[Pla10] Anna Plassart. A Scottish Jacobin: John Oswald on commerce and citizenship. Journal of the History of Ideas, 71(2):263–286, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/379772.

Plassart:2013:SWS

[Pla13] Anna Plassart. “Scientific whigs”?: Scottish historians on the French Revolution. Journal of the History of Ideas, 74(1):93–114, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/495243.

Plochmann:1953:NLT

[Plo53] George Kimball Plochmann. Nature and the living thing in Aristotle’s biology. Journal of the History of Ideas, 14(2):167–190, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707469.

Pocock:1978:BRC

[Poc78] J. G. A. Pocock. Book review: The Commons Debates of 1628. Journal of the History of Ideas, 39(2):329–334, April/June 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708784.

Pocock:1987:BGM

[Poc87] J. G. A. Pocock. Between Gog and Magog: The republican thesis and the Ideologia Americana. Journal of the History of Ideas, 48(2):325–346, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709561.

Pocock:2016:RC

[Poc16] J. G. A. Pocock. Response and commentary. Journal of the History of Ideas, 77(1):157–171, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/610566.

Podoksik:2002:VPT

[Pod02] Efraim Podoksik. The voice of poetry in the thought of Michael Oakeshott. Journal of the History of Ideas, 63(4):717–733, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/39297.

Podoksik:2010:OCL

[Pod10] Efraim Podoksik. One concept of liberty: Towards writing the history of a political concept. Journal of the History of Ideas, 71(2):219– 240, April 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
379774.

Poggioli:1963:NVP

[Pog63] Renato Poggioli. Naboth’s vineyard or the pastoral view of the social order. Journal of the History of Ideas, 24(1):3–24, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707856.

Pois:1970:TPW

[Poi70] Robert A. Pois. Two poles within historicism: Croce and Meinecke. Journal of the History of Ideas, 31(2):253–272, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708548.

Pojman:1983:CPK

[Poj83] Louis P. Pojman. Christianity and philosophy in Kierkegaard’s early papers. Journal of the History of Ideas, 44(1):131–140, January/ March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709309.

Polinger:1943:SSU

[Pol43] Elliot H. Polinger. Saint-Simon, the utopian precursor of the League of Nations. Journal of the History of Ideas, 4(4):475–483, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707169.

Pollin:1964:GLV

[Pol64] Burton R. Pollin. Godwin’s letters of verax. Journal of the History of Ideas, 25(3):353–373, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707904.

Pollin:1967:GLO

[Pol67] Burton R. Pollin. Godwin’s letter to Ogilvie, friend of Jefferson, and the Federalist propaganda. Journal of the History of Ideas, 28 (3):432–444, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708629.

Poole:1999:NIR

[Poo99] Randall Allen Poole. The neo-idealist reception of Kant in the Moscow Psychological Society. Journal of the History of Ideas, 60(2):319–343, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15030.

Poole:2005:FLC

[Poo05] William Poole. Francis Lodwick’s creation: Theology and natural philosophy in the Early Royal Society. Journal of the History of Ideas, 66(2):245–263, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/187190.

Popkin:1952:GTE

[Pop52] Richard H. Popkin. George Tucker, an Early American critic of Hume. Journal of the History of Ideas, 13(3):370–375, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707602.

Popkin:1953:JGP

[Pop53a] Richard H. Popkin. Joseph Glanvill: a precursor of David Hume. Journal of the History of Ideas, 14(2):292–303, April 1953. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707478.

Popkin:1953:SST

[Pop53b] Richard H. Popkin. Samuel Sorbière’s translation of Sextus Empiricus. Journal of the History of Ideas, 14(4):617–621, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707705.

Popkin:1954:DPR

[Pop54] Richard H. Popkin. The development of the philosophical reputation of Joseph glanvill. Journal of the History of Ideas, 15(2):305–311, April 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707775.

Popkin:1988:DBD

[Pop88] Richard H. Popkin. The dispersion of Bodin’s dialogues in England, Holland, and Germany. Journal of the History of Ideas, 49 (1):157–160, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709709.

Popkin:1993:SKS

[Pop93] Richard H. Popkin. Sources of knowledge of sextus empiricus in Hume’s time. Journal of the History of Ideas, 54(1):137–141, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709864.

Popper:2005:EPH

[Pop05] Nicholas Popper. The English polydaedali: How Gabriel Harvey read Late Tudor London. Journal of the History of Ideas, 66(3):351–381, July 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/188686.

Popper:2006:APM

[Pop06] Nicholas Popper. “Abraham, Planter of Mathematics”: Histories of mathematics and astrology in Early Modern Europe. Journal of the History of Ideas, 67(1):87–106, January 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/194649.

Porter:1972:AFL

[Por72] Dennis Porter. Anatole France and the liberalism of resignation. Journal of the History of Ideas, 33(1):101–114, January/March 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709059.

Porter:1978:PPG

[Por78] Roy S. Porter. Philosophy and politics of a geologist: G. H. Toulmin (1754–1817). Journal of the History of Ideas, 39(3):435–450, July/ September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709387.

Porter:1981:LPM

[Por81] J. M. Porter. Luther and political millenarianism: The case of the Peasants’ War. Journal of the History of Ideas, 42(3):389–406, July/ September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709183.

Porter:2002:EVN

[Por02] Clifford F. Porter. Eric Voegelin on Nazi political extremism. Journal of the History of Ideas, 63(1):151–171, January 2002. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/15147.

Post:1986:JRL

[Pos86] David M. Post. Jeffersonian revisions of Locke: Education, propertyrights, and liberty. Journal of the History of Ideas, 47(1):147– 157, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709601.

Poulet:1954:TR

[Pou54] Georges Poulet. Timelessness and romanticism. Journal of the History of Ideas, 15(1):3–22, January 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707647.

Power:1970:HMI

[Pow70] J. E. Power. Henry more and Isaac Newton on absolute space. Journal of the History of Ideas, 31(2):289–296, April/June 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708552.

Prize:2020:Nb

[PP20] Morris D. Forkosch Prize and Selma V. Forkosch Prize. Notices. Journal of the History of Ideas, 81(4):667–669, October 2020. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/772517.

Prakash:1955:HPH

[Pra55] Buddha Prakash. The Hindu philosophy of history. Journal of the History of Ideas, 16(4):494–505, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707506.

Pratt:1965:RE

[Pra65] Kenneth J. Pratt. Rome as eternal. Journal of the History of Ideas, 26(1):25–44, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708398.

Preu:1954:SIG

[Pre54] James Preu. Swift’s influence on Godwin’s doctrine of anarchism. Journal of the History of Ideas, 15(3):371–383, June 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707760.

Preston:1971:EEH

[Pre71] Joseph H. Preston. English ecclesiastical historians and the problem of bias, 1559–1742. Journal of the History of Ideas, 32(2):203–220, April/June 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708277.

Preston:1977:WTH

[Pre77] Joseph H. Preston. Was there an historical revolution? Journal of the History of Ideas, 38(2):353–364, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708919.

Preus:1979:MFA

[Pre79] J. Samuel Preus. Machiavelli’s functional analysis of religion: Context and object. Journal of the History of Ideas, 40(2):171–190, April/ June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709147.

Preus:1989:SVI

[Pre89] J. Samuel Preus. Spinoza, Vico, and the imagination of religion. Journal of the History of Ideas, 50(1):71–93, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709787.

Preece:2003:DCG

[Pre03] Rod Preece. Darwinism, Christianity, and the Great Vivisection Debate. Journal of the History of Ideas, 64(3):399–419, July 2003. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/47662.

Price:1948:SHC

[Pri48] Maurice T. Price. Sinism — a historical critique of H. G. Creel’s case for its pre-Confucian indigeneity. Journal of the History of Ideas, 9(2): 214–236, April 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707229.

Price:1949:RCC

[Pri49] Maurice T. Price. Reply to Creel’s clarification. Journal of the History of Ideas, 10(3):452–453, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707049. See [Cre49].

Prior:1954:BMS

[Pri54] Moody E. Prior. Bacon’s man of science. Journal of the History of Ideas, 15(3):348–370, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707759.

Price:1957:ECE

[Pri57] Kingsley Blake Price. Ernst Cassirer and the Enlightenment. Journal of the History of Ideas, 18(1):101–112, January 1957. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707581.

Priestley:1958:BRN

[Pri58] F. E. L. Priestley. Book review: The Notebooks of Samuel Taylor Coleridge by Kathleen Coburn. Journal of the History of Ideas, 19(4):593–596, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707927.

Priestley:1963:BRF

[Pri63] F. E. L. Priestley. Book review: Festschrift for Marjorie H. Nicolson. Journal of the History of Ideas, 24(3):433–439, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708220.

Price:1964:SCH

[Pri64a] John Valdimir Price. Sceptics in Cicero and Hume. Journal of the History of Ideas, 25(1):97–106, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708087.

Primer:1964:EDT

[Pri64b] Irwin Primer. Erasmus Darwin’s temple of nature: Progress, evolution, and the Eleusinian mysteries. Journal of the History of Ideas, 25(1):58–76, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708085.

Prinz:1969:BUM

[Pri69] Arthur M. Prinz. Background and ulterior motive of Marx’s “Preface” of 1859. Journal of the History of Ideas, 30(3):437–450, July/ September 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708568.

Priestley:1974:BRM

[Pri74] F. E. L. Priestley. Book review: Mapping the World of Ideas. Journal of the History of Ideas, 35(3):527–537, July/September 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708800.

Principe:1995:VRR

[Pri95] Lawrence M. Principe. Virtuous romance and romantic virtuoso: The shaping of Robert Boyle’s literary style. Journal of the History of Ideas, 56(3):377–397, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710032.

Prize:2012:N

[Pri12] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 73(1):175, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/463146.

Printy:2013:DMJ

[Pri13a] Michael Printy. The determination of man: Johann Joachim Spalding and the Protestant Enlightenment. Journal of the History of Ideas, 74(2):189–212, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/504784.

Prize:2013:N

[Pri13b] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 74 (2):355, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
504792.

Prize:2014:N

[Pri14] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 75(2):347–348, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/544565.

Prize:2015:N

[Pri15] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 76(3):495–496, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/587276.

Prize:2016:Na

[Pri16a] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 77(1):179–180, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/610568.

Prize:2016:Nb

[Pri16b] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 77(4):683–684, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/635635.

Priest:2017:CDT

[Pri17a] Greg Priest. Charles Darwin’s theory of moral sentiments: What Darwin’s ethics really owes to Adam Smith. Journal of the History of Ideas, 78(4):571–593, October 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/676147.

Prins:2017:GCJ

[Pri17b] Jacomien Prins. Girolamo Cardano and Julius Caesar Scaliger in debate about Nature’s musical secrets. Journal of the History of Ideas, 78(2):169–189, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/652821.

Prize:2020:Na

[Pri20] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 81(2):331–333, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/753532.

Prize:2021:N

[Pri21] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 82(3):541–543, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/798280.

Prize:2022:N

[Pri22] Morris D. Forkosch Prize. Notices. Journal of the History of Ideas, 83(1):175–177, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/847826.

Prochaska:1972:TPA

[Pro72] Franklyn K. Prochaska. Thomas Paine’s the age of reason revisited. Journal of the History of Ideas, 33(4):561–576, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708857.

Psalidas:1960:THB

[Psa60] Athanasios P. Psalidas. True happiness, or the basis of all religion. Journal of the History of Ideas, 21(4):481–496, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708097.

Puchner:2005:DLH

[Puc05] Martin Puchner. Doing logic with a hammer: Wittgenstein’s Tractatus and the polemics of logical positivism. Journal of the History of Ideas, 66(2):285–300, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/187191.

Purvis:2015:QWG

[Pur15] Zachary Purvis. Quiet war in Germany: Friedrich Schelling and Friedrich Schleiermacher. Journal of the History of Ideas, 76(3):369– 391, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
587270.

Puzzo:1959:GSH

[Puz59] Dante A. Puzzo. Gaetano Salvemini: an historiographical essay. Journal of the History of Ideas, 20(2):217–235, April 1959. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707820.

Puzzo:1964:RWT

[Puz64] Dante A. Puzzo. Racism and the Western tradition. Journal of the History of Ideas, 25(4):579–586, October/December 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708188.

Pycior:1987:MPW

[Pyc87] Helena M. Pycior. Mathematics and philosophy: Wallis, Hobbes, Barrow, and Berkeley. Journal of the History of Ideas, 48(2):265–286, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709558.

Quatrini:2019:ABC

[Qua19] Francesco Quatrini. Adam Boreel on collegiant freedom of speech. Journal of the History of Ideas, 80(4):511–531, October 2019. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/735379.

Quesada:1963:IML

[Que63] Francisco Miro Quesada. The impact of metaphysics on Latin-American ideology. Journal of the History of Ideas, 24(4):539– 552, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707983.

Quinones:1965:VTS

[Qui65] Ricardo J. Quinones. Views of time in Shakespeare. Journal of the History of Ideas, 26(3):327–352, July/September 1965. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708385.

Quillen:1992:TIP

[Qui92] Carol E. Quillen. A tradition invented: Petrarch, Augustine, and the language of humanism. Journal of the History of Ideas, 53(2):179–207, April/June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709870.

Quillen:2010:UPQ

[Qui10] Carol Quillen. The uses of the past in Quattrocento Florence: a reading of Leonardo Bruni’s dialogues. Journal of the History of Ideas, 71(3):363–385, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/386149.

Rabinowitz:1962:ASA

[Rab62] Margula Rabinowitz. Absolute space-again. Journal of the History of Ideas, 23(2):279–281, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708162.

Rachum:1995:MRE

[Rac95] Ilan Rachum. The meaning of “Revolution” in the English Revolution (1648–1660). Journal of the History of Ideas, 56(2):195–215, April 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709835.

Radcliffe:1993:RWM

[Rad93] Evan Radcliffe. Revolutionary writing, moral philosophy, and universal benevolence in the Eighteenth Century. Journal of the History of Ideas, 54(2):221–240, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709980.

Rahtz:2022:TTS

[Rah22] Joshua Rahtz. Two types of separation: Ludwig von Mises and German neoliberalism. Journal of the History of Ideas, 83(2):293– 313, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
855168.

Raju:1952:DIT

[Raj52] P. T. Raju. The development of Indian thought. Journal of the History of Ideas, 13(4):528–550, October 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707716.

Ramsey:2001:CIO

[Ram01] Rachel Ramsey. China and the ideal of order in John Webb’s An Historical Essay …. Journal of the History of Ideas, 62(3):483–503, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15125.

Ramelli:2014:DIO

[Ram14] Ilaria Ramelli. The divine as inaccessible object of knowledge in Ancient Platonism: a common philosophical pattern across religious traditions. Journal of the History of Ideas, 75(2):167–188, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/544557.

Ramberg:2019:TMG

[Ram19] Peter J. Ramberg. Transcendental materialism in the German Free Religious Movement: Science, nature, and theology in Kirchliche Reform, 1846–52. Journal of the History of Ideas, 80(3):409–431, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/729722.

Randall:1940:DSM

[Ran40] John Herman Randall, Jr. The development of scientific method in the school of Padua. Journal of the History of Ideas, 1(2):177–206, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707332.

Randall:1941:BRM

[Ran41] John Herman Randall, Jr. Book review: Man and Society in an Age of Reconstruction: Studies in Modern Social Structure by Karl Mannheim, Edward Shils. Journal of the History of Ideas, 2(3):372– 381, June 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707137.

Randall:1949:BRM

[Ran49] John Herman Randall, Jr. Book review: Morris R. Cohen’s the Meaning of Human History. Journal of the History of Ideas, 10(2):305–312, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707421.

Randall:1953:PLV

[Ran53] John Herman Randall, Jr. The place of Leonardo Da Vinci in the emergence of modern science. Journal of the History of Ideas, 14(2):

191–202, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707470.

Randall:1961:CID

[Ran61] John Herman Randall, Jr. The changing impact of Darwin on philosophy. Journal of the History of Ideas, 22(4):435–462, October/ December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708024.

Rand:1964:TMH

[Ran64] Calvin G. Rand. Two meanings of historicism in the writings of Dilthey, Troeltsch, and Meinecke. Journal of the History of Ideas, 25(4):503–518, October/December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708183.

Randall:1965:JSM

[Ran65] John Herman Randall, Jr. John Stuart Mill and the working-out of empiricism. Journal of the History of Ideas, 26(1):59–88, January/ March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708400.

Randall:1966:HGD

[Ran66a] John Herman Randall, Jr. T. H. Green: The development of English thought from J. S. Mill to F. H. Bradley. Journal of the History of Ideas, 27(2):217–244, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708640.

Rankin:1966:HPC

[Ran66b] Bayard Rankin. The history of probability and the changing concept of the individual. Journal of the History of Ideas, 27(4):483–504, October/December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708337.

Randall:1967:PTT

[Ran67] John Herman Randall, Jr. Plato’s treatment of the theme of the good life and his criticism of the Spartan ideal. Journal of the History of Ideas, 28(3):307–324, July/September 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708621.

Randall:1969:IUP

[Ran69] John Herman Randall, Jr. The intelligible universe of plotinos. Journal of the History of Ideas, 30(1):3–16, January/March 1969. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708241.

Randle:1970:NSD

[Ran70] Robert Randle. From national self-determination to national selfdevelopment. Journal of the History of Ideas, 31(1):49–68, January/ March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708369.

Randall:2023:RRC

[Ran23] Vicky Randall. The romance of the Republic: Class conflict and the problem of progress in Thomas Arnold’s History of Rome (1838– 42). Journal of the History of Ideas, 84(2):287–311, April 2023. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/886183.

Raphael:1969:ASI

[Rap69] D. D. Raphael. Adam Smith and ’the infection of David Hume’s society’: New light on an old controversy, together with the text of a hitherto unpublished manuscript. Journal of the History of Ideas, 30(2):225–248, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708434.

Rapaczynski:1981:LCP

[Rap81] Andrzej Rapaczynski. Locke’s conception of property and the principle of sufficient reason. Journal of the History of Ideas, 42(2):305–315, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709322.

Rappaport:1997:QEA

[Rap97] Rhoda Rappaport. Questions of evidence: an anonymous tract attributed to John toland. Journal of the History of Ideas, 58(2):339– 348, April 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14948.

Raphael:2016:GTN

[Rap16] Renée Raphael. Galileo’s two new sciences as a model of reading practice. Journal of the History of Ideas, 77(4):539–565, October 2016.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/635627.

Rapin:2021:FCT

[Rap21] Ami-Jacques Rapin. The first conceptualization of terrorism: Tallien, Roederer, and the “System of Terror” (August 1794). Journal of the History of Ideas, 82(3):405–426, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/798272.

Rashid:1977:RWC

[Ras77] Salim Rashid. Richard Whately and Christian political economy at Oxford and Dublin. Journal of the History of Ideas, 38(1): 147–155, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708847.

Rashid:1982:SIS

[Ras82] Salim Rashid. “He startled… as if he saw a spectre”: Tucker’s proposal for American independence. Journal of the History of Ideas, 43 (3):439–460, July/September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709432.

Rashid:1985:DSB

[Ras85] Salim Rashid. Dugald Stewart, “Baconian” methodology, and political economy. Journal of the History of Ideas, 46(2):245–257, April/ June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709637.

Rashid:2007:JNS

[Ras07] Salim Rashid. John von Neumann and scientific method. Journal of the History of Ideas, 68(3):501–527, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL http://www.

jstor.org/stable/30140935;   https://muse.jhu.edu/pub/56/
article/219822.

Ratner:1941:HAP

[Rat41] Sidney Ratner. The Historian’s approach to psychology. Journal of the History of Ideas, 2(1):95–109, January 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707283.

Ratner:1961:SCR

[Rat61] Joseph Ratner. Some comments on Rosen’s “Calvin’s Attitude Toward Copernicus”. Journal of the History of Ideas, 22(3):382– 385, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2708132. See [Ros60] and reply [Ros61].

Rather:1986:DFJ

[Rat86] L. J. Rather. Disraeli, Freud, and Jewish conspiracy theories. Journal of the History of Ideas, 47(1):111–131, January/March 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709598.

Ratner:1992:JDE

[Rat92] Sidney Ratner. John Dewey, empiricism, and experimentalism in the recent philosophy of mathematics. Journal of the History of Ideas, 53(3):467–479, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709888.

Raub:1988:RCW

[Rau88] Cymbre Quincy Raub. Robert Chambers and William Whewell: a Nineteenth-Century debate over the origin of language. Journal of the History of Ideas, 49(2):287–300, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709501.

Reaske:1972:DJE

[Rea72] Christopher R. Reaske. The Devil and Jonathan Edwards. Journal of the History of Ideas, 33(1):123–138, January/March 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709061.

Redwood:1974:CBD

[Red74] J. A. Redwood. Charles Blount (1654–93), deism, and English free thought. Journal of the History of Ideas, 35(3):490–498, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708796.

Redekop:1997:TAF

[Red97] Benjamin W. Redekop. Thomas Abbt and the formation of an enlightened German “Public”. Journal of the History of Ideas, 58(1):81–103, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14936.

Rees:1952:KPH

[Ree52] D. A. Rees. Kant’s “Physiology of the Human Understanding” and the classification of the sciences. Journal of the History of Ideas, 13(1):108–109, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707728.

Reedy:1983:LCR

[Ree83] W. Jay Reedy. Language, counter-revolution and the “Two Cultures”: Bonald’s traditionalist scientism. Journal of the History of Ideas, 44 (4):579–597, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709216.

Reeves:1991:AGR

[Ree91] Eileen Reeves. Augustine and Galileo on reading the heavens. Journal of the History of Ideas, 52(4):563–579, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709966.

Regoliosi:2005:SCC

[Reg05] Mariangela Regoliosi. Salvatore Camporeale’s contribution to theology and the history of the Church. Journal of the History of Ideas, 66(4):527–539, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/191476.

Rehding:2014:MHE

[Reh14] Alexander Rehding. Music-historical egyptomania, 1650–1950. Journal of the History of Ideas, 75(4):545–580, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/557233.

Reiss:1956:KRR

[Rei56] H. S. Reiss. Kant and the right of rebellion. Journal of the History of Ideas, 17(2):179–192, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707741.

Reif:1969:TTN

[Rei69] Patricia Reif. The textbook tradition in natural philosophy, 1600– 1650. Journal of the History of Ideas, 30(1):17–32, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708242.

Reiss:1981:MI

[Rei81] Edmond Reiss. Medieval irony. Journal of the History of Ideas, 42(2):209–226, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709317. See errata [Ano81f].

Reiss:1996:DBM

[Rei96] Timothy J. Reiss. Denying the body? Memory and the dilemmas of history in Descartes. Journal of the History of Ideas, 57(4):587–607, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14919.

Reidy:2017:RPD

[Rei17] David A. Reidy. Rawls on philosophy and democracy: Lessons from the archived papers. Journal of the History of Ideas, 78(2):265– 274, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
652826.

Reichardt:2021:FSI

[Rei21] Sven Reichardt. Fascism’s stages: Imperial violence, entanglement, and processualization. Journal of the History of Ideas, 82(1):85–107, January 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
781614.

Rempel:1979:IFT

[Rem79] Richard Rempel. From imperialism to free trade: Couturat, Halévy and Russell’s first crusade. Journal of the History of Ideas, 40 (3):423–443, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709246.

Renaldo:1976:BEB

[Ren76] John J. Renaldo. Bacon’s empiricism, Boyle’s science, and the Jesuit response in Italy. Journal of the History of Ideas, 37(4):

689–695, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709031.

Rescher:1960:NCA

[Res60] Nicholas Rescher. A Ninth-Century Arabic logician on: Is existence a predicate? Journal of the History of Ideas, 21(3):428–430, July/ September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708146.

Rescher:1963:AFL

[Res63] Nicholas Rescher. Al-Farabi on logical tradition. Journal of the History of Ideas, 24(1):127–132, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707863.

Reynolds:1953:SCH

[Rey53] Beatrice Reynolds. Shifting currents in historical criticism. Journal of the History of Ideas, 14(4):471–492, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707697.

Richards:1970:PIN

[RG70] Edward B. Richards and William R. Garner. The political implications of Nicholas Berdyaev’s philosophy. Journal of the History of Ideas, 31(1):121–128, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708374.

Riezler:1945:DKR

[RGE45] Kurt Riezler, James Gutmann, and Walter Eckstein. Discussion by Kurt Riezler. Journal of the History of Ideas, 6(3):294–306, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707293.

Reiss:1979:MVS

[RH79] Timothy J. Reiss and Roger H. Hinderliter. Money and value in the Sixteenth Century: The Monete Cudende ratio of Nicholas Copernicus. Journal of the History of Ideas, 40(2):293–313, April/June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709154.

Ribner:1952:SPS

[Rib52] Irving Ribner. Sir Philip Sidney on civil insurrection. Journal of the History of Ideas, 13(2):257–265, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707615.

Rice:1950:ERT

[Ric50] Eugene F. Rice, Jr. Erasmus and the religious tradition, 1495–1499. Journal of the History of Ideas, 11(4):387–411, October 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707589.

Richards:1979:IST

[Ric79] Robert J. Richards. Influence of sensationalist tradition on early theories of the evolution of behavior. Journal of the History of Ideas, 40(1):85–105, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709261.

Richter:1987:BHI

[Ric87] Melvin Richter. Begriffsgeschichte and the history of ideas. Journal of the History of Ideas, 48(2):247–263, April/June 1987. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709557.

Richards:2012:UBR

[Ric12] Jennifer Richards. Useful books: Reading vernacular regimens in Sixteenth-Century England. Journal of the History of Ideas, 73(2): 247–271, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
475339.

Ridgely:1956:DPN

[Rid56] Beverly S. Ridgely. Dalibray, Le Pailleur, and the “New Astronomy” in French Seventeenth-Century poetry. Journal of the History of Ideas, 17(1):3–27, January 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707683.

Ridgely:1962:MSG

[Rid62] Beverly S. Ridgely. Mellin De Saint-Gelais and the first vernacular reference to the Copernican system in France. Journal of the History of Ideas, 23(1):107–116, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708060.

Rie:1952:NA

[Rie52] Robert Rie. Nietzsche and after. Journal of the History of Ideas, 13(3): 349–369, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707601.

Rieff:1956:OFP

[Rie56] Philip Rieff. The origins of Freud’s political psychology. Journal of the History of Ideas, 17(2):235–249, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707744.

Rieser:1960:NMM

[Rie60] Max Rieser. The noetic models of mythology and metaphysics. Journal of the History of Ideas, 21(2):300–308, April/June 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708203.

Rieser:1961:BRH

[Rie61] Max Rieser. Book review: History of Philosophy, Vol. III by W. Tarkiewiez. Journal of the History of Ideas, 22(2):275–280, April/ June 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707839.

Rieser:1963:PPT

[Rie63] Max Rieser. Polish philosophy today. Journal of the History of Ideas, 24(3):423–432, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708219.

Riepe:1967:EIP

[Rie67] Dale Riepe. Emerson and Indian philosophy. Journal of the History of Ideas, 28(1):115–122, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708485.

Riegel:1968:CAA

[Rie68] Robert E. Riegel. Changing American attitudes toward prostitution (1800–1920). Journal of the History of Ideas, 29(3):437–452, July/ September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708454.

Rifkin:1953:AEC

[Rif53] Lester H. Rifkin. Aristotle on equality: a criticism of A. J. Carlyle’s theory. Journal of the History of Ideas, 14(2):276–283, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707476.

Rigotti:1986:BSA

[Rig86] Francesca Rigotti. Biology and society in the age of Enlightenment. Journal of the History of Ideas, 47(2):215–233, April/June 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709811.

Rigolot:1998:RCE

[Rig98] François Rigolot. The Renaissance crisis of exemplarity. Journal of the History of Ideas, 59(4):557–563, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15010.

Riker:1954:SGF

[Rik54] William H. Riker. Sidney George Fisher and the separation of powers during the Civil War. Journal of the History of Ideas, 15(3):397–412, June 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707762.

Riker:1957:DAF

[Rik57] William H. Riker. Dutch and American federalism. Journal of the History of Ideas, 18(4):495–521, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707561.

Riley:1973:HCS

[Ril73] Patrick Riley. How coherent is the social contract tradition? Journal of the History of Ideas, 34(4):543–562, October/December 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708887.

Riley:1994:NPW

[Ril94] Patrick Riley. “New” political writings of Leibniz. Journal of the History of Ideas, 55(1):147–158, January 1994. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709958.

Riley:1999:LPM

[Ril99] Patrick Riley. Leibniz’s political and moral philosophy in the Novissima Sinica, 1699–1999. Journal of the History of Ideas, 60(2):217– 239, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15031.

Ringler:1941:PNN

[Rin41] William Ringler. Poeta nascitur non fit: Some notes on the history of an aphorism. Journal of the History of Ideas, 2(4):497–504, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707023.

Ringer:1961:JJB

[Rin61] Alexander L. Ringer. J.-J. Barthelemy and musical utopia in Revolutionary France. Journal of the History of Ideas, 22(3):355–368, July/ September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708130.

Richards:1981:PPP

[RMG81] Judith Richards, Lotte Mulligan, and John K. Graham. “Property” and “People”: Political usages of Locke and some contemporaries. Journal of the History of Ideas, 42(1):29–51, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709416.

Robinson:1959:OHY

[Rob59] Elmo A. Robinson. One hundred years of philosophy teaching in California (1857–1957). Journal of the History of Ideas, 20(3): 369–384, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708115.

Roberts:1960:CCM

[Rob60] Louise Nisbet Roberts. A Chimera is a Chimera: a medieval tautology. Journal of the History of Ideas, 21(2):273–278, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708200.

Robbins:1962:MRW

[Rob62] Caroline Robbins. Marvell’s religion: Was He a new Methodist? Journal of the History of Ideas, 23(2):268–272, April/June 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708160.

Robbins:1975:FFC

[Rob75] Caroline Robbins. Faith and freedom (c. 1677–1729). Journal of the History of Ideas, 36(1):47–62, January/March 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709010.

Robinson:1980:ENT

[Rob80] David Robinson. Emerson’s natural theology and the Paris naturalists: Toward a theory of animated nature. Journal of the History of Ideas, 41(1):69–88, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709103.

Robertson:2006:MLP

[Rob06] David G. Robertson. Mind and language in Philo. Journal of the History of Ideas, 67(3):423–441, July 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/202879.

Robinson:2015:DPP

[Rob15] Emily Robinson. Defining progressive politics: Municipal socialism and anti-socialism in contestation, 1889–1939. Journal of the History of Ideas, 76(4):609–631, October 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/597349.

Robcis:2020:FFI

[Rob20a] Camille Robcis. Frantz Fanon, institutional psychotherapy, and the decolonization of psychiatry. Journal of the History of Ideas, 81(2): 303–325, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
753530.

Robertson:2020:CRP

[Rob20b] James Robertson. Communism as religious phenomenon: Phenomenology and Catholic socialism in Yugoslav Slovenia, 1927–42.

Journal of the History of Ideas, 81(2):279–301, April 2020. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/753538.

Robinson:2020:IUC

[Rob20c] Morgan Robinson. The idea of the Upelekwa: Constructing a transcontinental community in Eastern Africa, 1888–96. Journal of the History of Ideas, 81(1):85–106, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/747171.

Roberts:2023:TAC

[Rob23] Alexandre M. Roberts. Thinking about chemistry in Byzantium and the Islamic World. Journal of the History of Ideas, 84(4):595–619, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
909532.

Rodman:1968:BPP

[Rod68] Barbee-Sue Rodman. Bentham and the paradox of penal reform. Journal of the History of Ideas, 29(2):197–210, April/June 1968. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708576.

Rodgers:1988:KR

[Rod88] Daniel T. Rodgers. Keywords: a reply. Journal of the History of Ideas, 49(4):669–676, October/December 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709679.

Rodgers:2007:AC

[Rod07] Daniel T. Rodgers. Anatomy of a Cliché. Journal of the History of Ideas, 68(3):389–393, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/219823.

Roe:1985:VVN

[Roe85] Shirley A. Roe. Voltaire versus Needham: Atheism, materialism, and the generation of life. Journal of the History of Ideas, 46 (1):65–87, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709776.

Roehr:2003:FAS

[Roe03] Sabine Roehr. Freedom and autonomy in Schiller. Journal of the History of Ideas, 64(1):119–134, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/42177.

Rogerson:1953:APP

[Rog53] Brewster Rogerson. The art of painting the passions. Journal of the History of Ideas, 14(1):68–94, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707496.

Rogers:1966:BLR

[Rog66] G. A. J. Rogers. Boyle, Locke, and reason. Journal of the History of Ideas, 27(2):205–216, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708639.

Rogers:1972:DSD

[Rog72] James Allen Rogers. Darwinism and social Darwinism. Journal of the History of Ideas, 33(2):265–280, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708873.

Rogers:1978:LEN

[Rog78] G. A. J. Rogers. Locke’s essay and Newton’s Principia. Journal of the History of Ideas, 39(2):217–232, April/June 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708776.

Rogers:1979:LNC

[Rog79] G. A. J. Rogers. Locke, Newton, and the Cambridge Platonists on innate ideas. Journal of the History of Ideas, 40(2):191–205, April/ June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709148.

Rogers:1994:OEP

[Rog94] W. Kim Rogers. Ortega and ecological philosophy. Journal of the History of Ideas, 55(3):503–522, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709852.

Rogers:2006:MJW

[Rog06] G. A. J. Rogers. In memoriam: John W. Yolton 1921–2005. Journal of the History of Ideas, 67(2):419–421, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/198081.

Rollin:1971:HBP

[Rol71] Bernard E. Rollin. Hume’s blue patch and the Mind’s creativity. Journal of the History of Ideas, 32(1):119–128, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708329.

Rollmann:1978:DNV

[Rol78] Hans Rollmann. Deussen, Nietzsche, and Vedānta. Journal of the History of Ideas, 39(1):125–132, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709076.

Romein:1948:TH

[Rom48] Jan Romein. Theoretical history. Journal of the History of Ideas, 9(1):53–64, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707394.

Romanell:1960:OMT

[Rom60] Patrick Romanell. Ortega in Mexico: a tribute to Samuel Ramos. Journal of the History of Ideas, 21(4):600–608, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708107.

Romanell:1961:LAE

[Rom61] Patrick Romanell. Locke’s aphorisms on education and health. Journal of the History of Ideas, 22(4):549–554, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708030.

Romanell:1964:SMP

[Rom64] Patrick Romanell. Some medico-philosophical excerpts from the mellon collection of Locke papers. Journal of the History of Ideas, 25 (1):107–116, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708088.

Ron:2012:LHL

[Ron12] Amit Ron. The logic of the historian and the logic of the citizen. Journal of the History of Ideas, 73(4):643–655, October 2012. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/488353.

Rosen:1940:RRC

[Ros40] Edward Rosen. The Ramus–Rheticus correspondence. Journal of the History of Ideas, 1(3):363–368, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707093.

Rosen:1943:ATC

[Ros43a] Edward Rosen. The authentic title of Copernicus’s major work. Journal of the History of Ideas, 4(4):457–474, October 1943. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707168.

Rosenthal:1943:ASM

[Ros43b] Jerome Rosenthal. Attitudes of some modern rationalists to history. Journal of the History of Ideas, 4(4):429–456, October 1943. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707167.

Rosenthal:1955:VPH

[Ros55] Jerome Rosenthal. Voltaire’s philosophy of history. Journal of the History of Ideas, 16(2):151–178, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707660.

Rosen:1957:BRS

[Ros57a] Edward Rosen. Book review: Stillman Drake’s Discoveries and Opinions of Galileo. Journal of the History of Ideas, 18(3):439–448, June 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707802.

Rosenfield:1957:MRC

[Ros57b] Leonora Cohen Rosenfield. Morris R. Cohen, the teacher. Journal of the History of Ideas, 18(4):552–571, October 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707567.

Rosen:1960:CAT

[Ros60] Edward Rosen. Calvin’s attitude toward Copernicus. Journal of the History of Ideas, 21(3):431–441, July/September 1960. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708147. See [Rat61] and reply [Ros61].

Rosen:1961:RDR

[Ros61] Edward Rosen. A reply to Dr. Ratner. Journal of the History of Ideas, 22(3):386–388, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708133. See [Ros60, Rat61].

Rosenberg:1963:MLF

[Ros63] Nathan Rosenberg. Mandeville and laissez-faire. Journal of the History of Ideas, 24(2):183–196, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707844.

Ross:1968:SJH

[Ros68] Sydney Ross. Sir John Herschel’s marginal notes on Mill’s on liberty, 1859. Journal of the History of Ideas, 29(1):123–130, January/ March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708470.

Rosen:1970:PPP

[Ros70] Frederick Rosen. The principle of population as political theory: Godwin’s of population and the Malthusian controversy. Journal of the History of Ideas, 31(1):33–48, January/March 1970. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708368.

Rosen:1971:CAT

[Ros71a] Edward Rosen. Copernicus’ attitude toward the common people. Journal of the History of Ideas, 32(2):281–288, April/June 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708281.

Ross:1971:CTC

[Ros71b] Robert N. Ross. “To Charm Thy Curious Eye”: Erasmus Darwin’s poetry at the vestibule of knowledge. Journal of the History of Ideas, 32(3):379–394, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708353.

Rosen:1974:BRS

[Ros74] Edward Rosen. Book review: Studia Copernicana. Journal of the History of Ideas, 35(3):521–526, July/September 1974. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708799.

Rosen:1975:WCR

[Ros75] Edward Rosen. Was Copernicus’ Revolutions approved by the Pope? Journal of the History of Ideas, 36(3):531–542, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708661.

Rosen:1977:HSW

[Ros77] Edward Rosen. How the shackles were forged and later loosened. Journal of the History of Ideas, 38(1):109–117, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708843.

Rosen:1978:GP

[Ros78] Edward Rosen. Galileo and the philosophers. Journal of the History of Ideas, 39(1):147, January/March 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709078. See reply [Dra78].

Rosenblum:1982:RM

[Ros82] Nancy L. Rosenblum. Romantic militarism. Journal of the History of Ideas, 43(2):249–268, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709202.

Rosen:1983:WCN

[Ros83] Edward Rosen. Was Copernicus a neoplatonist? Journal of the History of Ideas, 44(4):667–669, October/December 1983. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709222.

Ross:1984:PPC

[Ros84] Stephanie Ross. Painting the passions: Charles LeBrun’s Conférence Sur l’Expression. Journal of the History of Ideas, 45(1):25–47, January/March 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709329.

Rosen:1985:BRK

[Ros85a] Edward Rosen. Book review: Kepler’s Early Writings. Journal of the History of Ideas, 46(3):449–454, July/September 1985. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709479. See erratum [Ano85e].

Rosen:1985:DSC

[Ros85b] Edward Rosen. The dissolution of the solid celestial spheres. Journal of the History of Ideas, 46(1):13–31, January/March 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709773.

Rose:1992:REC

[Ros92] Jonathan Rose. Rereading the English common reader: a preface to a history of audiences. Journal of the History of Ideas, 53(1):47–70, January/March 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709910.

Rosa:1996:SCC

[Ros96a] Susan Rosa. Seventeenth-Century Catholic polemic and the rise of cultural rationalism: an example from the empire. Journal of the History of Ideas, 57(1):87–107, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14893.

Rosand:1996:OMS

[Ros96b] David Rosand. Obituary: Meyer Schapiro (1904–1996). Journal of the History of Ideas, 57(3):547–549, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14916.

Rosenstein:1998:MTM

[Ros98] Roy Rosenstein. A medieval troubadour mobilized in the French resistance. Journal of the History of Ideas, 59(3):499–520, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14996.

Rosenberg:2003:EMI

[Ros03] Daniel Rosenberg. Early modern information overload. Journal of the History of Ideas, 64(1):1–9, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/42178.

Rose:2012:ASF

[Ros12] Anne C. Rose. An American science of feeling: Harvard’s psychology of emotion during the World War I era. Journal of the History of Ideas, 73(3):485–506, July 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/481421.

Rosenblatt:2016:CMP

[Ros16a] Helena Rosenblatt. On context and meaning in Pocock’s barbarism and religion, and on Gibbon’s “Protestantism” in his chapters on religion. Journal of the History of Ideas, 77(1):147–155, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/610565.

Rosenblatt:2016:RT

[Ros16b] Helena Rosenblatt. Rousseau, the “Traditionalist”. Journal of the History of Ideas, 77(4):627–635, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/635631.

Ross:2018:AAF

[Ros18] Tricia M. Ross. Anthropologia: an (almost) forgotten early modern history. Journal of the History of Ideas, 79(1):1–22, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/684829.

Rosenboim:2024:FUI

[Ros24] Or Rosenboim. A food utopia? Italian colonial visions of Tripolitania and Cyrenaica, 1911–13. Journal of the History of Ideas, 85(2):289– 320, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
926150.

Rothbard:1958:NBV

[Rot58] Murray N. Rothbard. A note on Burke’s vindication of natural society. Journal of the History of Ideas, 19(1):114–118, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707957.

Roth:1990:NEC

[Rot90] Michael S. Roth. Narrative as enclosure: The contextual histories of H. Stuart Hughes. Journal of the History of Ideas, 51(3): 505–515, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709627.

Roucek:1944:HCI

[Rou44] Joseph S. Roucek. A history of the concept of ideology. Journal of the History of Ideas, 5(4):479–488, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707082.

Rousseau:1967:MBS

[Rou67] G. S. Rousseau. Matt Bramble and the sulphur controversy in the XVIIIth Century: Medical background of Humphry Clinker. Journal of the History of Ideas, 28(4):577–589, October/December 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708531.

Rowbotham:1956:JFE

[Row56a] Arnold H. Rowbotham. The Jesuit figurists and Eighteenth-Century religious thought. Journal of the History of Ideas, 17(4):471–485, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707782.

Rowen:1956:STL

[Row56b] Herbert H. Rowen. A second thought on Locke’s first treatise. Journal of the History of Ideas, 17(1):130–132, January 1956. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707691.

Royse:1979:NDP

[Roy79] James R. Royse. Nominalism and divine power in the Chester cycle. Journal of the History of Ideas, 40(3):475–476, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709250.

Roy:2007:SHC

[Roy07] Ayon Roy. The specter of Hegel in Coleridge’s Biographia Literaria. Journal of the History of Ideas, 68(2):279–304, April 2007.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/215056.

Rozema:2002:TLP

[Roz02] David Rozema. Tractatus Logico-Philosophicus: a “Poem” by Ludwig Wittgenstein. Journal of the History of Ideas, 63(2):345–363, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15156.

Richter:2006:ITR

[RR06] Melvin Richter and Michaela Richter. Introduction: Translation of Reinhart Koselleck’s “Krise,” in Geschichtliche Grundbegriffe. Journal of the History of Ideas, 67(2):343–356, April 2006. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/198080.

Rescher:1964:YIA

[RS64] Nicholas Rescher and Fadlou Shehadi. Yahya Ibn ’Adi’s treatise “On the Four Scientific Questions Regarding the Art of Logic”. Journal of the History of Ideas, 25(4):572–578, October/December 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708187.

Russo:1968:HPO

[RS68] Joseph Russo and Bennett Simon. Homeric psychology and the oral epic tradition. Journal of the History of Ideas, 29(4):483–498, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708290.

Ruby:1986:OSL

[Rub86] Jane E. Ruby. The origins of scientific “Law”. Journal of the History of Ideas, 47(3):341–359, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709657.

Rubidge:1990:DMD

[Rub90] Bradley Rubidge. Descartes’s meditations and devotional meditations. Journal of the History of Ideas, 51(1):27–49, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709745.

Rubies:1991:HGD

[Rub91] Joan-Pau Rubiés. Hugo Grotius’s dissertation on the origin of the American peoples and the use of comparative methods. Journal of the History of Ideas, 52(2):221–244, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709526.

Rubies:2006:TEH

[Rub06] Joan Pau Rubiés. Theology, ethnography, and the historicization of idolatry. Journal of the History of Ideas, 67(4):571–596, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/204780.

Rudowski:1974:TSE

[Rud74] Victor Anthony Rudowski. The theory of signs in the Eighteenth Century. Journal of the History of Ideas, 35(4):683–690, October/ December 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709094.

Rudavsky:2001:GSH

[Rud01] Tamar Rudavsky. Galileo and Spinoza: Heroes, heretics, and hermeneutics. Journal of the History of Ideas, 62(4):611–631, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15136.

Rueger:1964:GCM

[Rue64] Zofia Rueger. Gerson, the Conciliar Movement and the right of resistance (1642–1644). Journal of the History of Ideas, 25(4): 467–486, October/December 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708181.

Rukgaber:2018:ISR

[Ruk18] Matthew Rukgaber. Immaterial spirits and the reform of first philosophy: The compatibility of Kant’s pre-critical metaphysics with the arguments in dreams of a spirit-seer. Journal of the History of Ideas, 79(3):363–383, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/703663.

Rumble:1965:LRS

[Rum65] Wilfrid E. Rumble, Jr. Legal realism, sociological jurisprudence and Mr. Justice Holmes. Journal of the History of Ideas, 26(4):

547–566, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708499.

Rumrich:1985:MDS

[Rum85] John Peter Rumrich. Milton, Duns Scotus, and the fall of Satan. Journal of the History of Ideas, 46(1):33–49, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709774.

Runia:2000:IRC

[Run00] David T. Runia. The idea and the reality of the city in the thought of Philo of Alexandria. Journal of the History of Ideas, 61(3):361–379, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15085.

Ruotolo:1966:TFD

[Ruo66] Lucio P. Ruotolo. The trinitarian framework of Donne’s holy sonnet XIV. Journal of the History of Ideas, 27(3):445–446, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708597.

Russell:1940:BMW

[Rus40] Bertrand Russell. Byron and the modern world. Journal of the History of Ideas, 1(1):24–37, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707008.

Ruse:1975:CDA

[Rus75] Michael Ruse. Charles Darwin and artificial selection. Journal of the History of Ideas, 36(2):339–350, April/June 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708932.

Russell:1985:HTH

[Rus85] Paul Russell. Hume’s treatise and Hobbes’s the elements of law. Journal of the History of Ideas, 46(1):51–63, January/March 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709775.

Russell:1988:SNR

[Rus88] Paul Russell. Skepticism and natural religion in Hume’s treatise. Journal of the History of Ideas, 49(2):247–265, April/June 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709499.

Russell:1993:EPF

[Rus93] Paul Russell. Epigram, pantheists, and freethought in Hume’s treatise: a study in esoteric communication. Journal of the History of Ideas, 54(4):659–673, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709828.

Rutherford:2014:ILV

[Rut14] Emily Rutherford. Impossible love and Victorian values: J. A. Symonds and the intellectual history of homosexuality. Journal of the History of Ideas, 75(4):605–627, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/557235.

Ryan:2001:PSB

[Rya01] Vanessa Lyndal Ryan. The physiological sublime: Burke’s critique of reason. Journal of the History of Ideas, 62(2):265–279, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15115.

Ryan:2009:RMG

[Rya09] Vanessa L. Ryan. Reading the mind: From George Eliot’s fiction to James Sully’s psychology. Journal of the History of Ideas, 70(4):615– 635, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
361796.

Saastamoinen:2010:PNE

[Saa10] Kari Saastamoinen. Pufendorf on natural equality, human dignity, and self-esteem. Journal of the History of Ideas, 71(1):39–62, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/370095.

Sacharoff:1972:SBP

[Sac72] Mark Sacharoff. Suicide and Brutus’ philosophy in Julius Caesar. Journal of the History of Ideas, 33(1):115–122, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709060.

Sachs:1976:MSF

[Sac76] Mendel Sachs. Maimonides, Spinoza, and the field concept in physics. Journal of the History of Ideas, 37(1):125–131, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708713.

Saccenti:2018:MNN

[Sac18] Riccardo Saccenti. The ministerium naturae:: Natural law in the Exegesis and theological discourse at Paris between 1160 and 1215. Journal of the History of Ideas, 79(4):527–545, October 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/707771.

Sailor:1962:CD

[Sai62] Danton B. Sailor. Cudworth and Descartes. Journal of the History of Ideas, 23(1):133–140, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708062.

Sailor:1964:MA

[Sai64] Danton B. Sailor. Moses and atomism. Journal of the History of Ideas, 25(1):3–16, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708082.

Sailor:1988:NDC

[Sai88] Danton B. Sailor. Newton’s debt to Cudworth. Journal of the History of Ideas, 49(3):511–518, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709490.

Sakamoto:2009:GHH

[Sak09] Kuni Sakamoto. The German Hercules’s heir: Pierre Gassendi’s reception of Keplerian ideas. Journal of the History of Ideas, 70(1):69– 91, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
258195.

Salmon:1989:SRE

[Sal89] J. H. M. Salmon. Stoicism and Roman example: Seneca and Tacitus in Jacobean England. Journal of the History of Ideas, 50(2):199–225, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709732.

Salas:1996:RPW

[Sal96] Charles G. Salas. Ralegh and the Punic Wars. Journal of the History of Ideas, 57(2):195–215, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/14908.

Sams:1943:SLD

[Sam43] Henry W. Sams. Self-love and the doctrine of work. Journal of the History of Ideas, 4(3):320–332, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707257.

Sanford:1944:SAH

[San44] Eva Matthews Sanford. The study of ancient history in the Middle Ages. Journal of the History of Ideas, 5(1):21–43, January 1944. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707100.

Sansom:1959:BRE

[San59] George Sansom. Book review: Early Sources of Japanese Tradition. Journal of the History of Ideas, 20(1):138–142, January 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707972.

Sandler:1960:LIT

[San60] S. Gerald Sandler. Lockean ideas in Thomas Jefferson’s bill for establishing religious freedom. Journal of the History of Ideas, 21 (1):110–116, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708002.

Sandin:1962:FUS

[San62] Robert T. Sandin. The founding of the Uppsala School. Journal of the History of Ideas, 23(4):496–512, October/December 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708171.

Sanford:1965:DD

[San65] William F. Sanford, Jr. Dana and Darwinism. Journal of the History of Ideas, 26(4):531–546, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708498.

Santurri:1982:TSP

[San82] Edmund N. Santurri. Theodicy and social policy in Malthus’ thought. Journal of the History of Ideas, 43(2):315–330, April/June 1982. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709207.

Santini:2020:SOH

[San20] Carlotta Santini. Searching for orientation in the history of culture: Aby Warburg and Leo Frobenius on the morphological study of the ifa-board. Journal of the History of Ideas, 81(3):473–497, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/761032.

Saperstein:1979:BRC

[Sap79] Marc Saperstein. Book review: Current Israeli Scholarship on Medieval Hebrew Literature. Journal of the History of Ideas, 40(1): 159–166, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709269.

Sargent:1982:BRT

[Sar82] Lyman Tower Sargent. Book review: Is There Only One Utopian Tradition? Journal of the History of Ideas, 43(4):681–689, October/ December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709351.

Sarf:1983:RKS

[Sar83] Harold Sarf. Reflections on Kierkegaard’s Socrates. Journal of the History of Ideas, 44(2):255–276, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709139.

Sarasohn:1985:MMP

[Sar85] Lisa T. Sarasohn. Motion and morality: Pierre Gassendi, Thomas Hobbes and the mechanical world-view. Journal of the History of Ideas, 46(3):363–379, July/September 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709473.

Sarafianos:2008:CBS

[Sar08] Aris Sarafianos. The contractility of Burke’s sublime and heterodoxies in medicine and art. Journal of the History of Ideas, 69(1):23– 48, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
231014.

Saresella:2014:DBC

[Sar14] Daniela Saresella. The dialogue between Catholics and Communists in Italy during the 1960s. Journal of the History of Ideas, 75(3):493– 512, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
549332.

Satris:1982:TVR

[Sat82] Stephen A. Satris. The theory of value and the rise of ethical emotivism. Journal of the History of Ideas, 43(1):109–128, January/ March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709163.

Savelle:1944:BRF

[Sav44] Max J. Savelle. Book review: Force and Freedom; Reflections on History by Jacob Burckhardt, James Hastings Nichols. Journal of the History of Ideas, 5(3):374–376, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707357.

Saveson:1959:DIJ

[Sav59] J. E. Saveson. Descartes’ influence on John Smith, Cambridge Platonist. Journal of the History of Ideas, 20(2):258–263, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707823.

Saveson:1960:DRD

[Sav60] J. E. Saveson. Differing reactions to Descartes among the Cambridge platonists. Journal of the History of Ideas, 21(4):560–567, October/ December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708103.

Sayre:1945:GC

[Say45] Farrand Sayre. Greek cynicism. Journal of the History of Ideas, 6(1):113–118, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707062.

Saccaro-Battisti:1983:CMP

[SB83] Giuseppa Saccaro-Battisti. Changing metaphors of political structures. Journal of the History of Ideas, 44(1):31–54, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709303.

Smith:1940:FAC

[SC40] Bernard Smith and Morris R. Cohen. On “Forces in American Criticism”. Journal of the History of Ideas, 1(3):369–374, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707094.

Scaglione:1997:PLP

[Sca97] Aldo D. Scaglione. Petrarchan love and the pleasures of frustration. Journal of the History of Ideas, 58(4):557–572, October 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14967.

Scanlan:1999:CAR

[Sca99] James P. Scanlan. The case against rational egoism in Dostoevsky’s notes from underground. Journal of the History of Ideas, 60(3):549– 567, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15040.

Schneider:1940:BRN

[Sch40] Herbert W. Schneider. Book review: The New England Mind. The Seventeenth Century by Perry Miller. Journal of the History of Ideas, 1(1):119–122, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707015.

Schapiro:1941:BRH

[Sch41] J. Salwyn Schapiro. Book review: History as the Story of Liberty by Benedetto Croce, Sylvia Sprigge. Journal of the History of Ideas, 2(4):505–508, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707024.

Schapiro:1943:JSM

[Sch43a] J. Salwyn Schapiro. John Stuart Mill, pioneer of democratic liberalism in England. Journal of the History of Ideas, 4(2):127–160, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707321.

Schilling:1943:ECA

[Sch43b] Bernard N. Schilling. The English case against Voltaire: 1789–1800. Journal of the History of Ideas, 4(2):193–216, April 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707324.

Schlatter:1943:PHC

[Sch43c] Richard B. Schlatter. The problem of historical causation in some recent studies of the English Revolution. Journal of the History of Ideas, 4(3):349–367, June 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707259.

Schneider:1944:BRG

[Sch44] Herbert W. Schneider. Book review: The Growth of American Thought by Merle Curti. Journal of the History of Ideas, 5(2):243– 249, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707388.

Schlatter:1945:THT

[Sch45a] Richard Schlatter. Thomas Hobbes and Thucydides. Journal of the History of Ideas, 6(3):350–362, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707297.

Schneider:1945:IDS

[Sch45b] Herbert W. Schneider. The influence of Darwin and Spencer on American philosophical theology. Journal of the History of Ideas, 6(1):3–18, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707053.

Schrade:1946:BCB

[Sch46a] Leo Schrade. Bach: The conflict between the sacred and the secular. Journal of the History of Ideas, 7(2):151–194, April 1946. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707070.

Schrecker:1946:LPI

[Sch46b] Paul Schrecker. Leibniz’s principles of international justice. Journal of the History of Ideas, 7(4):484–498, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707035.

Schrecker:1947:LAI

[Sch47] Paul Schrecker. Leibniz and the art of inventing algorisms. Journal of the History of Ideas, 8(1):107–116, January 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707443.

Schapiro:1949:C

[Sch49] J. Salwyn Schapiro. Comment. Journal of the History of Ideas, 10(2): 303–304, April 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707420.

Schneider:1950:BRE

[Sch50] Herbert W. Schneider. Book review: Evolution and the Founders of Pragmatism by Philip P. Wiener, John Dewey. Journal of the History of Ideas, 11(2):241–245, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707249.

Schwartz:1951:CTH

[Sch51] Benjamin Schwartz. Ch’ên Tu-Hsiu and the acceptance of the Modern West. Journal of the History of Ideas, 12(1):61–74, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707537.

Schueller:1952:UDM

[Sch52] Herbert M. Schueller. The use and decorum of music as described in British literature, 1700 to 1780. Journal of the History of Ideas, 13(1):73–93, January 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707726.

Schmidt:1953:ISP

[Sch53] H. D. Schmidt. The idea and slogan of “Perfidious Albion”. Journal of the History of Ideas, 14(4):604–616, October 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707704.

Schnitzler:1954:GVM

[Sch54] Henry Schnitzler. “Gay Vienna” — myth and reality. Journal of the History of Ideas, 15(1):94–118, January 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707651.

Schapiro:1956:LFA

[Sch56a] Meyer Schapiro. Leonardo and Freud: an art-historical study. Journal of the History of Ideas, 17(2):147–178, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707740.

Schmidt:1956:CNH

[Sch56b] Royal J. Schmidt. Cultural nationalism in Herder. Journal of the History of Ideas, 17(3):407–417, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707552.

Schmitter:1960:OMG

[Sch60] Dean Morgan Schmitter. The occasion for Marvell’s growth of popery. Journal of the History of Ideas, 21(4):568–570, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708104.

Schofield:1964:JPT

[Sch64] Robert E. Schofield. Joseph Priestley, the theory of oxidation and the nature of matter. Journal of the History of Ideas, 25(2):285–294, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708017.

Schirokauer:1965:BRB

[Sch65] Conrad Schirokauer. Book review: Balazs on China. Journal of the History of Ideas, 26(4):593–597, October/December 1965. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708503.

Schmitt:1966:PPA

[Sch66] Charles B. Schmitt. Perrenial philosophy: From Agostino Steuco to Leibniz. Journal of the History of Ideas, 27(4):505–532, October/ December 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708338.

See erratum [Ano67g].

Schaff:1973:MTR

[Sch73a] Adam Schaff. Marxist theory on revolution and violence. Journal of the History of Ideas, 34(2):263–270, April/June 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708729.

Schalk:1973:PWP

[Sch73b] David L. Schalk. Professors as watchdogs: Paul Nizan’s theory of the intellectual and politics. Journal of the History of Ideas, 34 (1):79–96, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708945.

Schankula:1980:LDS

[Sch80a] H. A. S. Schankula. Locke, Descartes, and the science of nature. Journal of the History of Ideas, 41(3):459–477, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709406.

Schuler:1980:SSA

[Sch80b] Robert M. Schuler. Some spiritual alchemies of Seventeenth-Century England. Journal of the History of Ideas, 41(2):293–318, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709463.

Schulz:1983:JAB

[Sch83] Constance B. Schulz. John Adams on “The Best of all Possible Worlds”. Journal of the History of Ideas, 44(4):561–577, October/ December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709215.

Schiffman:1984:MRS

[Sch84] Zachary S. Schiffman. Montaigne and the rise of skepticism in Early Modern Europe: a reappraisal. Journal of the History of Ideas, 45 (4):499–516, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709370.

Schmidt:1989:QEK

[Sch89a] James Schmidt. The question of Enlightenment: Kant, Mendelssohn, and the Mittwochsgesellschaft. Journal of the History of Ideas, 50(2):269–291, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709735.

Schochet:1989:BRR

[Sch89b] Gordon J. Schochet. Book review: Radical Politics and Ashcraft’s Treatise on Locke. Journal of the History of Ideas, 50(3):491– 510, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709574.

Schwoerer:1990:LLI

[Sch90] Lois G. Schwoerer. Locke, Lockean ideas, and the glorious revolution. Journal of the History of Ideas, 51(4):531–548, October/December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709645.

Schneewind:1991:NLS

[Sch91a] J. B. Schneewind. Natural law, skepticism, and methods of ethics. Journal of the History of Ideas, 52(2):289–308, April/June 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709529.

Schorske:1991:OLK

[Sch91b] Carl E. Schorske. Obituary: Leonard Krieger 1918–1990. Journal of the History of Ideas, 52(2):340, April/June 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709534.

Schott:1993:JAW

[Sch93] Linda Schott. Jane Addams and William James on alternatives to war. Journal of the History of Ideas, 54(2):241–254, April 1993. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709981.

Schmidt:1996:FTD

[Sch96] James Schmidt. The fool’s truth: Diderot, Goethe, and Hegel. Journal of the History of Ideas, 57(4):625–644, October 1996. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/14927.

Schneider:1998:ER

[Sch98] Ulrich Johannes Schneider. Eclecticism rediscovered. Journal of the History of Ideas, 59(1):173–182, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14969.

Schmaltz:1999:WCD

[Sch99] Tad M. Schmaltz. What has Cartesianism to do with Jansenism? Journal of the History of Ideas, 60(1):37–56, January 1999. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15021.

Schmitter:2002:RBP

[Sch02] Amy M. Schmitter. Representation and the body of power in French academic painting. Journal of the History of Ideas, 63(3):399–424, July 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15166.

Schmidt:2003:IEA

[Sch03] James Schmidt. Inventing the Enlightenment: Anti-Jacobins, British Hegelians, and the Oxford English Dictionary. Journal of the History of Ideas, 64(3):421–443, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/1/

article/47664.

Schmidt:2004:MTL

[Sch04] Jeremy Schmidt. Melancholy and the therapeutic language of moral philosophy in Seventeenth-Century thought. Journal of the History of Ideas, 65(4):583–601, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/184073.

Schneewind:2005:GHP

[Sch05a] J. B. Schneewind. Globalization and the history of philosophy. Journal of the History of Ideas, 66(2):169–178, April 2005. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/187192.

Schneider:2005:IDI

[Sch05b] Ulrich Johannes Schneider. The international dictionary of intellectual historians. Journal of the History of Ideas, 66(2):143–154, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/187193.

Schaefer:2007:IIF

[Sch07a] Richard Schaefer. Infallibility and intentionality: Franz Brentano’s diagnosis of German Catholicism. Journal of the History of Ideas, 68(3): 477–499, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
219824.

Schwartzberg:2007:JBF

[Sch07b] Melissa Schwartzberg. Jeremy Bentham on fallibility and infallibility. Journal of the History of Ideas, 68(4):562–586, October 2007. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/223134.

Schwegman:2010:SRT

[Sch10] Jeffrey Schwegman. The “System” as a reading technology: Pedagogy and philosophical criticism in Condillac’s Traité des systêmes. Journal of the History of Ideas, 71(3):387–409, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/386150.

Schmidt:2014:ECC

[Sch14a] James Schmidt. Enlightenment as concept and context. Journal of the History of Ideas, 75(4):677–685, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/557240.

Schwartz:2014:PRD

[Sch14b] Daniel Schwartz. Probabilism reconsidered: Deference to experts, types of uncertainty, and medicines. Journal of the History of Ideas, 75(3):373–393, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/549327.

Schatz:2015:LCD

[Sch15] Stephanie L. Schatz. Lewis Carroll’s dream-child and Victorian child psychopathology. Journal of the History of Ideas, 76(1):93–114, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
566378.

Schultz:2022:PIC

[Sch22a] Karie Schultz. Protestant intellectual culture and political ideas in the Scottish universities, ca. 1600–50. Journal of the History of Ideas, 83(1):41–62, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/847820.

Schwartz:2022:SCS

[Sch22b] Daniel Schwartz. Seventeenth-Century Scotism and the war just on both sides. Journal of the History of Ideas, 83(4):643–658, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/865885.

Schlichter:2023:EES

[Sch23a] Felix Schlichter. Euhemerus and Euhemerism in the Seventeenth and Eighteenth Centuries. Journal of the History of Ideas, 84(4):653– 683, October 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
909534.

Schwab:2023:RPH

[Sch23b] Maren Elisabeth Schwab. Rome as “Part of the Heavens”? Leon Battista Alberti’s Descriptio urbis Romae (ca. 1450) and Ptolemy’s Almagest. Journal of the History of Ideas, 84(1):1–27, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/879538.

Sclar:2020:PCO

[Scl20] David Sclar. Perfecting community as “One Man”: Moses H .ayim Luzzatto’s pietistic confraternity in Eighteenth-Century Padua. Journal of the History of Ideas, 81(1):45–66, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/747172.

Scott:1997:LTC

[Sco97] David Scott. Leibniz and the two clocks. Journal of the History of Ideas, 58(3):445–463, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14957.

Scott:1998:HBR

[Sco98] John T. Scott. The harmony between Rousseau’s musical theory and his philosophy. Journal of the History of Ideas, 59(2):287–308, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14985.

Scrivener:1978:GPR

[Scr78] Michael H. Scrivener. Godwin’s philosophy: a revaluation. Journal of the History of Ideas, 39(4):615–626, October/December 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709445.

Seaman:1968:MRA

[Sea68] Francis Seaman. Mach’s rejection of atomism. Journal of the History of Ideas, 29(3):381–393, July/September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708449.

Seaton:1980:DFS

[Sea80] James Seaton. Dialectics: Freedom of speech and thought. Journal of the History of Ideas, 41(2):283–289, April/June 1980. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709461. See reply [Ber80].

Seagrave:2011:HOM

[Sea11] S. Adam Seagrave. How old are modern rights?: On the Lockean roots of contemporary human rights discourse. Journal of the History of Ideas, 72(2):305–327, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/ 56/article/428856. See response [Tie11].

Seagrave:2012:IDH

[Sea12] S. Adam Seagrave. Identity and diversity in the history of ideas: a reply to brian tierney. Journal of the History of Ideas, 73(1):163– 166, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
463144.

Seidlin:1942:GB

[Sei42] Oskar Seidlin. Georg Brandes 1842–1927. Journal of the History of Ideas, 3(4):415–442, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707316.

Seigel:1965:IES

[Sei65] Jerrold E. Seigel. Ideals of eloquence and silence in Petrarch. Journal of the History of Ideas, 26(2):147–174, April/June 1965. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708225.

Seigel:1969:EEL

[Sei69] Jules Paul Seigel. The Enlightenment and the evolution of a language of signs in France and England. Journal of the History of Ideas, 30(1):96–115, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708248.

Seidel:1974:PBW

[Sei74] Michael A. Seidel. Poulain De La Barre’s the woman as good as the man. Journal of the History of Ideas, 35(3):499–508, July/September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708797.

Seidler:1983:KSS

[Sei83] Michael J. Seidler. Kant and the stoics on suicide. Journal of the History of Ideas, 44(3):429–453, July/September 1983. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709175.

Seigel:1987:APD

[Sei87] Jerrold Seigel. Autonomy and personality in Durkheim: an essay on content and method. Journal of the History of Ideas, 48 (3):483–507, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709764.

Seigel:1990:ASF

[Sei90] Jerrold Seigel. Avoiding the subject: a foucaultian itinerary. Journal of the History of Ideas, 51(2):273–299, April/June 1990. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709516.

Seliger:1958:RTD

[Sel58] M. Seliger. Race-thinking during the restoration. Journal of the History of Ideas, 19(2):273–282, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707941.

Seliger:1963:LNL

[Sel63] M. Seliger. Locke’s natural law and the foundation of politics. Journal of the History of Ideas, 24(3):337–354, July/September 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708211.

Selden:1974:HLM

[Sel74] Raman Selden. Hobbes and late metaphysical poetry. Journal of the History of Ideas, 35(2):197–210, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708758.

Sela:2013:TPP

[Sel13] Ori Sela. Through the prism of a plagiarism debate: Eighteenth through Twentieth-Century Sino–Japanese intellectual relations. Journal of the History of Ideas, 74(4):577–597, October 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/523429.

Senn:1958:EUT

[Sen58] Peter R. Senn. The earliest use of the term “Social Science”. Journal of the History of Ideas, 19(4):568–570, October 1958. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707925. See [Bur59, Igg59].

Serjeantson:2001:PAL

[Ser01] Richard Serjeantson. The passions and animal language, 1540–1700. Journal of the History of Ideas, 62(3):425–444, July 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15126.

Serjeantson:2017:FBV

[Ser17] Richard Serjeantson. Francis Bacon’s Valerius Terminus and the voyage to the “Great Instauration”. Journal of the History of Ideas, 78(3): 341–368, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
665519.

Sewell:2003:HBP

[Sew03] Keith C. Sewell. The “Herbert Butterfield Problem” and its resolution. Journal of the History of Ideas, 64(4):599–618, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/52454.

Selfridge-Field:1972:BGC

[SF72] Eleanor Selfridge-Field. Beethoven and Greek classicism. Journal of the History of Ideas, 33(4):577–595, October/December 1972. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708858.

Shrader-Frechette:1993:LLL

[SF93] Kristin Shrader-Frechette. Locke and limits on land ownership. Journal of the History of Ideas, 54(2):201–219, April 1993. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709979.

Sgarbi:2015:BVS

[Sga15] Marco Sgarbi. Benedetto Varchi on the soul: Vernacular Aristotelianism between reason and faith. Journal of the History of Ideas, 76(1): 1–23, January 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
566374.

Shapere:1963:DP

[Sha63] Dudley Shapere. Descartes and Plato. Journal of the History of Ideas, 24(4):572–576, October/December 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707986.

Shaffer:1970:MCK

[Sha70] Elinor S. Shaffer. Metaphysics of culture: Kant and Coleridge’s aids to reflection. Journal of the History of Ideas, 31(2):199–218, April/ June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708545.

Shaw:1988:BTI

[Sha88] Miranda Shaw. Buddhist and Taoist influences on Chinese landscape painting. Journal of the History of Ideas, 49(2):183–206, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709496.

Shalev:2002:MAT

[Sha02] Zur Shalev. Measurer of all things: John Greaves (1602=-1652), the Great Pyramid, and Early Modern metrology. Journal of the History of Ideas, 63(4):555–575, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/39298.

Sherwin:1944:DRW

[She44] Oscar Sherwin. Defoe reviews World War II. Journal of the History of Ideas, 5(3):359–368, June 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707355.

Sheldon:1946:LMU

[She46] Wilmon H. Sheldon. Leibniz’s message to US. Journal of the History of Ideas, 7(4):385–396, October 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707030.

Sherbo:1957:CSK

[She57] Arthur Sherbo. Christopher Smart’s knowledge of occult literature. Journal of the History of Ideas, 18(2):233–241, April 1957. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707626.

Shell:1977:REL

[She77] Marc Shell. Ruskin and the economy of literature. Journal of the History of Ideas, 38(1):65–84, January/March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708841.

Shelley:2000:IFM

[She00] Cameron Shelley. The influence of folk meteorology in the anaximander fragment. Journal of the History of Ideas, 61(1):1–17, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15063.

Shelford:2002:TGP

[She02] April Shelford. Thinking geometrically in Pierre-Daniel Huet’s Demonstratio evangelica (1679). Journal of the History of Ideas, 63(4):599–617, October 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/39299.

Sheehan:2003:PFB

[She03] Jonathan Sheehan. From philology to fossils: The biblical encyclopedia in Early Modern Europe. Journal of the History of Ideas, 64(1):41–60, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/42179.

Sheehan:2006:AIR

[She06a] Jonathan Sheehan. The altars of the idols: Religion, sacrifice, and the early modern polity. Journal of the History of Ideas, 67(4):649– 674, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
204782.

Sheehan:2006:ITA

[She06b] Jonathan Sheehan. Introduction: Thinking about idols in Early Modern Europe. Journal of the History of Ideas, 67(4):561–569, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/204781.

Shine:1945:JMO

[Shi45] Hill Shine. J. S. Mill and an open letter on the Saint-Simonian Society in 1832. Journal of the History of Ideas, 6(1):102–108, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707060.

Shin:1972:WPP

[Shi72] Kee Soo Shin. Writings of Professor Philip P. Wiener. Journal of the History of Ideas, 33(3):358–364, July/September 1972. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709040.

Shklar:1981:JAR

[Shk81] Judith N. Shklar. Jean D’Alembert and the rehabilitation of history. Journal of the History of Ideas, 42(4):643–664, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709123.

Shouse:1952:DHW

[Sho52] J. B. Shouse. David Hume and William James: a comparison. Journal of the History of Ideas, 13(4):514–527, October 1952. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707715.

Shoenberger:1979:LJR

[Sho79] Cynthia Grant Shoenberger. Luther and the justifiability of resistance to legitimate authority. Journal of the History of Ideas, 40(1):3–20, January/March 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709257.

Shoaf:1985:SSC

[Sho85] Richard Shoaf. Science, sect, and certainty in Voltaire’s dictionnaire philosophique. Journal of the History of Ideas, 46(1):121–126, January/March 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709779.

Shogimen:1999:RBT

[Sho99] Takashi Shogimen. The relationship between theology and canon law: Another context of political thought in the Early Fourteenth Century. Journal of the History of Ideas, 60(3):417–431, July 1999. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15041.

Shogimen:2010:APE

[Sho10] Takashi Shogimen. “Another” patriotism in Early Shōwa Japan (1930–1945). Journal of the History of Ideas, 71(1):139–160, January 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/370094.

Shogimen:2016:DEC

[Sho16] Takashi Shogimen. Dialogue, eurocentrism, and comparative political theory: a view from cross-cultural intellectual history. Journal of the History of Ideas, 77(2):323–345, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/620984.

Shortall:2018:TPC

[Sho18] Sarah Shortall. Theology and the politics of Christian human rights. Journal of the History of Ideas, 79(3):445–460, July 2018. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/703668.

Shortall:2024:BEP

[Sho24] Sarah Shortall. “Building the Earth”: Pierre Teilhard de Chardin, science, and the spirituality of the United Nations. Journal of the History of Ideas, 85(4):827–855, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/944587.

Shuchman:1962:ACC

[Shu62] Philip Shuchman. Aristotle’s conception of contract. Journal of the History of Ideas, 23(2):257–264, April/June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708158.

Shuchman:1963:CCC

[Shu63] Philip Shuchman. Comments on the criminal code of Plato’s Laws. Journal of the History of Ideas, 24(1):25–40, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707857.

Shugg:1968:CBM

[Shu68] Wallace Shugg. The Cartesian beast-machine in English literature (1663–1750). Journal of the History of Ideas, 29(2):279–292, April/ June 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708581.

Shue:1973:LNH

[Shu73] Henry G. Shue. Lukács: Notes on his originality. Journal of the History of Ideas, 34(4):645–650, October/December 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708897. See reply [Lic73].

Shute:1976:LGA

[Shu76] Michael N. Shute. A little great awakening: an episode in the American Enlightenment. Journal of the History of Ideas, 37(4): 589–602, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709026.

Shusterman:1985:RHN

[Shu85] Richard Shusterman. Remembering Hulme: a neglected philosopher– critic–poet. Journal of the History of Ideas, 46(4):559–576, October/ December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709545.

See errata [Ano86e].

Shurts:2022:III

[Shu22] Sarah Shurts. Identity, immigration, and Islam: Neo-reactionary and new-right perceptions and prescriptions. Journal of the History of Ideas, 83(3):477–499, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/858859.

Shyovitz:2014:CJT

[Shy14] David I. Shyovitz. Christians and Jews in the Twelfth-Century werewolf Renaissance. Journal of the History of Ideas, 75(4):521–543, October 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
557232.

Siegel:1950:MHA

[Sie50] Paul N. Siegel. Milton and the humanist attitude toward women. Journal of the History of Ideas, 11(1):42–53, January 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707451.

Siegel:1952:EHN

[Sie52] Paul N. Siegel. English humanism and the new Tudor aristocracy. Journal of the History of Ideas, 13(4):450–468, October 1952. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707712.

Siebert:1975:JHM

[Sie75] Donald T. Siebert, Jr. Johnson and Hume on miracles. Journal of the History of Ideas, 36(3):543–547, July/September 1975. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708662.

Siebert:1984:HII

[Sie84] Donald T. Siebert. Hume on idolatry and incarnation. Journal of the History of Ideas, 45(3):379–396, July/September 1984. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709231.

Sigmund:1962:IMP

[Sig62] Paul E. Sigmund, Jr. The influence of Marsilius of Padua on XVth-Century conciliarism. Journal of the History of Ideas, 23 (3):392–402, July/September 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708074.

Silver:1973:BPI

[Sil73] Bruce Silver. Berkeley and the principle of inertia. Journal of the History of Ideas, 34(4):599–608, October/December 1973. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708891.

Silver:1976:CMW

[Sil76] Bruce Silver. The conflicting microscopic worlds of Berkeley’s three dialogues. Journal of the History of Ideas, 37(2):343–349, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708830.

Silver:1977:RPM

[Sil77] Bruce Silver. Reply to Professor Mirarchi. Journal of the History of Ideas, 38(4):714–715, October/December 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708699. See [Mir77a].

Silver:1978:WBO

[Sil78] Bruce Silver. William Bartram’s and other Eighteenth-Century accounts of nature. Journal of the History of Ideas, 39(4):597–614, October/December 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709444.

Silver:1986:CQB

[Sil86] Bruce Silver. Clarke on the Quaker background of William Bartram’s approach to nature. Journal of the History of Ideas, 47 (3):507–510, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709667.

Silver:1993:BJR

[Sil93] Bruce Silver. Boswell on Johnson’s refutation of Berkeley: Revisiting the stone. Journal of the History of Ideas, 54(3):437–448, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710022.

Siljak:2001:BEW

[Sil01] Ana Siljak. Between East and West: Hegel and the origins of the Russian dilemma. Journal of the History of Ideas, 62(2):335–358, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15116.

Silver:2020:ET

[Sil20] Sean Silver. The emergence of texture. Journal of the History of Ideas, 81(2):169–194, April 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/753533.

Simonson:1947:IIH

[Sim47] Solomon Simonson. The idea of interpretation in Hebrew thought. Journal of the History of Ideas, 8(4):467–474, October 1947. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707430.

Simons:1955:RMB

[Sim55] Richard B. Simons. T. R. Malthus on British society. Journal of the History of Ideas, 16(1):60–75, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707527.

Simon:1956:HUS

[Sim56] Walter M. Simon. History for utopia: Saint-Simon and the idea of progress. Journal of the History of Ideas, 17(3):311–331, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707547. See erra-

tum [Ano56e].

Simon:1960:HSS

[Sim60] Walter M. Simon. Herbert Spencer and the “Social Organism”. Journal of the History of Ideas, 21(2):294–299, April/June 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708202. See comment [Bow75b].

Simon:1965:TCN

[Sim65] W. M. Simon. The ’two cultures’ in Nineteenth-Century France: Victor Cousin and Auguste Comte. Journal of the History of Ideas, 26(1):45–58, January/March 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708399.

Simon:1981:VMP

[Sim81] Lawrence H. Simon. Vico and Marx: Perspectives on historical development. Journal of the History of Ideas, 42(2):317–331, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709323.

Simons:1990:SPC

[Sim90] Margaret A. Simons. Sexism and the philosophical canon: On reading Beauvoir’s The Second Sex. Journal of the History of Ideas, 51 (3):487–504, July/September 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709626.

Simon:2004:SDM

[Sim04] Julia Simon. Singing democracy: Music and politics in Jean-Jacques Rousseau’s thought. Journal of the History of Ideas, 65(3):433–454, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/177668.

Simbirski:2016:CMH

[Sim16] Brian Simbirski. Cybernetic muse: Hannah arendt on automation, 1951–1958. Journal of the History of Ideas, 77(4):589–613, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/635629.

Singleton:1949:DM

[Sin49] Charles S. Singleton. Dante and myth. Journal of the History of Ideas, 10(4):482–502, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707186.

Singh:1963:HLN

[Sin63] Raghuveer Singh. Herakleitos and the law of nature. Journal of the History of Ideas, 24(4):457–472, October/December 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707978.

Singer:1989:HRC

[Sin89] Thomas C. Singer. Hieroglyphs, real characters, and the idea of natural language in English Seventeenth-Century thought. Journal of the History of Ideas, 50(1):49–70, January/March 1989. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709786.

Sinkoff:2000:BFJ

[Sin00] Nancy Sinkoff. Benjamin Franklin in Jewish Eastern Europe: Cultural appropriation in the Age of the Enlightenment. Journal of the History of Ideas, 61(1):133–152, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15064.

Sinclair:2016:BPW

[Sin16] Mark Sinclair. Bergson’s philosophy of will and the war of 1914– 1918. Journal of the History of Ideas, 77(3):467–487, July 2016. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/626911.

Siraisi:1991:GCA

[Sir91] Nancy G. Siraisi. Girolamo Cardano and the art of medical narrative. Journal of the History of Ideas, 52(4):581–602, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709967.

Siraisi:2003:HAM

[Sir03] Nancy G. Siraisi. History, antiquarianism, and medicine: The case of Girolamo Mercuriale. Journal of the History of Ideas, 64(2):231–251, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/43513.

Siraisi:2004:ORR

[Sir04] Nancy G. Siraisi. Oratory and rhetoric in Renaissance medicine. Journal of the History of Ideas, 65(2):191–211, April 2004. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/173852.

Sireci:2018:WWR

[Sir18] Fiore Sireci. “Writers who have rendered women objects of pity”: Mary Wollstonecraft’s literary criticism in the Analytical Review and A Vindication of the Rights of Woman. Journal of the History of Ideas, 79(2):243–265, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/693303.

Steinmetz-Jenkins:2018:MLS

[SJB18] Daniel Steinmetz-Jenkins and Kevin Brookes. The many liberalisms of serge audier. Journal of the History of Ideas, 79(1):45–63, January 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/684831.

Skagestad:2005:CBC

[Ska05] Peter Skagestad. Collingwood and Berlin: a comparison. Journal of the History of Ideas, 66(1):99–112, January 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/185262.

Skinner:1976:ASA

[Ski76] Andrew S. Skinner. Adam Smith and the American economic community an essay in applied economics. Journal of the History of Ideas, 37(1):59–78, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708709.

Skinner:2012:LAM

[Ski12] Quentin Skinner. On the liberty of the ancients and the moderns: a reply to my critics. Journal of the History of Ideas, 73(1):127– 146, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/

463142. See [Lan12a, Pet12].

Skinner:2025:JGP

[Ski25] Quentin Skinner. J. G. A. Pocock: a life in letters. Journal of the History of Ideas, 86(1):1–19, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/949925.

Skjonsberg:2022:CFS

[Skj22] Max Skjönsberg. Charles Francis Sheridan on the feudal origins and political science of the 1772 revolution in Sweden. Journal of the History of Ideas, 83(3):407–430, July 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/858856.

Skotheim:1964:WAH

[Sko64] Robert Allen Skotheim. The writing of American histories of ideas: Two traditions in the XXth Century. Journal of the History of Ideas, 25(2):257–278, April/June 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708015.

Skolimowski:1965:ALM

[Sko65] Henryk Skolimowski. Analytical–linguistic Marxism in Poland. Journal of the History of Ideas, 26(2):235–258, April/June 1965. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708230.

Skulsky:1958:APR

[Sku58] Harold Skulsky. Aristotle’s poetics revisited. Journal of the History of Ideas, 19(2):147–160, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707932.

Skulsky:1964:PLC

[Sku64] Harold Skulsky. Pain, law, and conscience in measure for measure. Journal of the History of Ideas, 25(2):147–168, April/June 1964. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708009.

Slive:1953:RHC

[Sli53] Seymour Slive. Rembrandt and his contemporary critics. Journal of the History of Ideas, 14(2):203–220, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707471.

Small:1983:NPT

[Sma83] Robin Small. Nietzsche and a platonist tradition of the Cosmos: Center everywhere and circumference nowhere. Journal of the History of Ideas, 44(1):89–104, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709306.

Small:1991:IRO

[Sma91] Robin Small. Incommensurability and recurrence: From Oresme to Simmel. Journal of the History of Ideas, 52(1):121–137, January/ March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709585.

Smeltzer:2020:TLA

[Sme20] Joshua Smeltzer. Technology, law, and annihilation: Carl Schmitt’s critique of utopianism. Journal of the History of Ideas, 81(1):107– 129, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
747173.

Smith:1941:KMS

[Smi41] Goldwin Smith. Karl Marx and St. George. Journal of the History of Ideas, 2(4):401–419, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707019.

Smith:1950:BMF

[Smi50] James Ward Smith. The British moralists and the fallacy of psychologism. Journal of the History of Ideas, 11(2):159–178, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707244.

Smith:1954:ETT

[Smi54] Lacey Baldwin Smith. English treason trials and confessions in the Sixteenth Century. Journal of the History of Ideas, 15(4):471–498, October 1954. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707672.

Smith:1961:RBI

[Smi61] Constance I. Smith. Richard Bentley and the innate idea of God: a correction. Journal of the History of Ideas, 22(1):117–118, January/ March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707878.

Smith:1962:LJA

[Smi62a] Constance I. Smith. Locke and John Austin on the idea of morality. Journal of the History of Ideas, 23(1):141–142, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708063.

Smith:1962:SIE

[Smi62b] Constance I. Smith. Some ideas on education before Locke. Journal of the History of Ideas, 23(3):403–406, July/September 1962. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708075.

Smith:1963:MCP

[Smi63] Brian Smith. Maine’s concept of progress. Journal of the History of Ideas, 24(3):407–412, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708216.

Smith:1964:JBC

[Smi64] Constance I. Smith. Jean Bodin and comparative law. Journal of the History of Ideas, 25(3):417–422, July/September 1964. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707907.

Smith:1965:HW

[Smi65a] Constance I. Smith. Hegel on war. Journal of the History of Ideas, 26(2):282–285, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708234.

Smith:1965:BRH

[Smi65b] Henry Nash Smith. Book review: Howard Mumford Jones: O Strange New World. Journal of the History of Ideas, 26(3):435–439, July/ September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708392.

Smith:1966:HF

[Smi66] Constance I. Smith. Heraclitus and fire. Journal of the History of Ideas, 27(1):125–127, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708312.

Smith:1967:DAI

[Smi67] Constance I. Smith. Descendit ad inferos — again. Journal of the History of Ideas, 28(1):87–88, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708482.

Smith:1969:OF

[Smi69] Constance I. Smith. Ophites and fathers. Journal of the History of Ideas, 30(2):249–250, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor. org/stable/2708435. See errata [Ano69g].

Smith:1970:BSR

[Smi70] Constance I. Smith. Bentham’s second rule. Journal of the History of Ideas, 31(3):462–463, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708518.

Smith:1973:SPG

[Smi73] Constance I. Smith. Saint Paul and a Greek idea. Journal of the History of Ideas, 34(2):281–282, April/June 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708731.

Smith:1976:GTI

[Smi76a] A. Mark Smith. Galileo’s theory of indivisibles: Revolution or compromise? Journal of the History of Ideas, 37(4):571–588, October/ December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709025.

Smith:1976:SHV

[Smi76b] Gordon Ross Smith. Shakespeare’s Henry v: Another part of the critical forest. Journal of the History of Ideas, 37(1):3–26, January/ March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708706.

Smith:1978:SME

[Smi78] Norah Smith. Sexual mores in the Eighteenth Century: Robert Wallace’s “Of Venery”. Journal of the History of Ideas, 39(3): 419–433, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709386.

Smith:1980:JME

[Smi80a] G. W. Smith. J. S. Mill on Edger and Réville: an episode in the development of Mill’s conception of freedom. Journal of the History of Ideas, 41(3):433–449, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709403.

Smith:1980:RPC

[Smi80b] Gordon Ross Smith. A rabble of princes: Considerations touching Shakespeare’s political orthodoxy in the second tetralogy. Journal of the History of Ideas, 41(1):29–48, January/March 1980. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709101.

Smithson:1983:MVA

[Smi83] Isaiah Smithson. The moral view of Aristotle’s poetics. Journal of the History of Ideas, 44(1):3–17, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709301.

Smith:1995:WTW

[Smi95] Bonnie G. Smith. Whose truth, whose history? Journal of the History of Ideas, 56(4):661–668, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709997.

Smith:2017:UKI

[Smi17] David Chan Smith. Useful knowledge, improvement, and the logic of capital in Richard Ligon’s True and Exact History of Barbados.

Journal of the History of Ideas, 78(4):549–570, October 2017. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/676146.

Smith:2019:LPS

[Smi19] Sophie Smith. The language of “Political Science” in Early Modern Europe. Journal of the History of Ideas, 80(2):203–226, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/722378.

Smith:2024:WIHa

[Smi24a] Sophie Smith. Women and intellectual history in the Twentieth Century, part one: Rethinking the “Origins” of US intellectual history. Journal of the History of Ideas, 85(3):425–454, July 2024. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/933854.

Smith:2024:WIHb

[Smi24b] Sophie Smith. Women and intellectual history in the Twentieth Century, part two: Activists, academics, and the future. Journal of the History of Ideas, 85(4):633–679, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/944581.

Snyder:1986:FRL

[Sny86] David C. Snyder. Faith and reason in Locke’s essay. Journal of the History of Ideas, 47(2):197–213, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709810.

Snyder:2011:MFC

[Sny11] James G. Snyder. Marsilio Ficino’s critique of the Lucretian alternative. Journal of the History of Ideas, 72(2):165–181, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/428850.

Socher:2006:FTO

[Soc06] Abraham P. Socher. Funkenstein on the theological origins of historicism: a critical note. Journal of the History of Ideas, 67(2):401– 408, April 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
198082.

Solmsen:1958:APM

[Sol58] Friedrich Solmsen. Aristotle and prime matter: a reply to Hugh R. King. Journal of the History of Ideas, 19(2):243–252, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707937. See [Kin56].

Solmsen:1963:NCG

[Sol63] Friedrich Solmsen. Nature as craftsman in Greek thought. Journal of the History of Ideas, 24(4):473–496, October/December 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707979.

Solmsen:1970:GWC

[Sol70] Friedrich Solmsen. George A. Wells on Christmas in Early New Testament criticism. Journal of the History of Ideas, 31(2):277–280, April/ June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708550.

Solmsen:1973:RW

[Sol73] Friedrich Solmsen. Reply to Wells. Journal of the History of Ideas, 34(1):144–145, January/March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708951. See [Wel73].

Solmsen:1978:PVA

[Sol78] Friedrich Solmsen. Platonic values in Aristotle’s science. Journal of the History of Ideas, 39(1):3–23, January/March 1978. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709069.

Solie:1982:MHM

[Sol82] Ruth A. Solie. Melody and the historiography of music. Journal of the History of Ideas, 43(2):297–308, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709205.

Solmsen:1983:PCS

[Sol83] Friedrich Solmsen. Plato and the concept of the soul (psyche): Some historical perspectives. Journal of the History of Ideas, 44 (3):355–367, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709171.

Soll:2000:AHA

[Sol00] Jacob Soll. Amelot de La Houssaye (1634=-1706) Annotates Tacitus. Journal of the History of Ideas, 61(2):167–187, April 2000. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15076.

Soll:2003:EHT

[Sol03a] Jacob Soll. Empirical history and the transformation of political criticism in France from Bodin to Bayle. Journal of the History of Ideas, 64(2):297–316, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/43515.

Soll:2003:IUH

[Sol03b] Jacob Soll. Introduction: The uses of historical evidence in Early Modern Europe. Journal of the History of Ideas, 64(2):149–157, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/43514.

Solch:2016:WWP

[Söl16] Dennis Sölch. Wheeler and Whitehead: Process biology and process philosophy in the Early Twentieth Century. Journal of the History of Ideas, 77(3):489–507, July 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/626912.

Soman:1967:AFD

[Som67] Alfred Soman. Arminianism in France: The D’Huisseau incident. Journal of the History of Ideas, 28(4):597–600, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708533.

Sommerville:1983:DBI

[Som83] C. John Sommerville. The distinction between indoctrination and education in England, 1549–1719. Journal of the History of Ideas, 44 (3):387–406, July/September 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709173.

Sommer:1998:GGU

[Som98] Elisabeth W. Sommer. Gambling with God: The use of the lot by the Moravian brethren in the Eighteenth Century. Journal of the History of Ideas, 59(2):267–286, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14986.

Sonntag:1974:AHF

[Son74] Otto Sonntag. Albrecht Von Haller on the future of science. Journal of the History of Ideas, 35(2):313–322, April/June 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708765.

Sorkin:1983:WHT

[Sor83] David Sorkin. Wilhelm Von Humboldt: The theory and practice of self-formation (Bildung), 1791–1810. Journal of the History of Ideas, 44(1):55–73, January/March 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709304.

Soss:1973:OTL

[Sos73] Neal M. Soss. Old Testament law and economic society. Journal of the History of Ideas, 34(3):323–344, July/September 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708956.

Soucy:1970:RRF

[Sou70] Robert Soucy. Romanticism and realism in the Fascism of drieu la rochelle. Journal of the History of Ideas, 31(1):69–90, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708370.

Southgate:1989:FLS

[Sou89] B. C. Southgate. “Forgotten and lost”: Some reactions to autonomous science in the Seventeenth Century. Journal of the History of Ideas, 50(2):249–268, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709734.

Southgate:1992:CTP

[Sou92] Beverley C. Southgate. “Cauterising the tumour of pyrrhonism”: Blackloism versus skepticism. Journal of the History of Ideas, 53 (4):631–645, October/December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709941.

Southgate:2000:BTT

[Sou00] Beverley C. Southgate. Blackloism and tradition: From theological certainty to historiographical doubt. Journal of the History of Ideas, 61(1):97–114, January 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15065.

Sparkes:1966:BQ

[Spa66] A. W. Sparkes. “Begging the question”. Journal of the History of Ideas, 27(3):462–463, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708600.

Spencer:1948:BRL

[Spe48] Theodore Spencer. Book review: Lovejoy’s Essays in the History of Ideas. Journal of the History of Ideas, 9(4):439–446, October 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707217.

Spencer:1955:BAF

[Spe55] Philip Spencer. “Barbarian assault”: The fortunes of a phrase. Journal of the History of Ideas, 16(2):232–239, April 1955. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707664.

Spector:1996:CFF

[Spe96] Jack J. Spector. Correspondence of Freud and Ferenczi. Journal of the History of Ideas, 57(2):365–372, April 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14902.

Spector:1998:BAG

[Spe98] Scott Spector. Beyond the aesthetic garden: Politics and culture on the margins of fin-de-siècle Vienna. Journal of the History of Ideas, 59(4):691–710, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15011.

Spektorowski:2002:MDC

[Spe02] Alberto Spektorowski. Maistre, Donoso Cortes, and the legacy of Catholic authoritarianism. Journal of the History of Ideas, 63(2):283– 302, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15157.

Spector:2013:FIC

[Spe13] Scott Spector. Forum: Ideas and the city introduction. Journal of the History of Ideas, 74(2):259–266, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/504787.

Spitzer:1944:GVH

[Spi44] Leo Spitzer. Geistesgeschichte vs. history of ideas as applied to Hitlerism. Journal of the History of Ideas, 5(2):191–203, April 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707384. See reply [Lov44].

Spitz:1952:SLH

[Spi52] Lewis W. Spitz. The significance of Leibniz for historiography. Journal of the History of Ideas, 13(3):333–348, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707600.

Spiegel:1955:TED

[Spi55a] Henry William Spiegel. Theories of economic development: History and classification. Journal of the History of Ideas, 16(4):518–539, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707508.

Spitz:1955:NLT

[Spi55b] Lewis W. Spitz. Natural law and the theory of history in Herder. Journal of the History of Ideas, 16(4):453–475, October 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707504.

Spiegelberg:1961:ABN

[Spi61] Herbert Spiegelberg. ‘Accident of birth’: a non-utilitarian motif in Mill’s philosophy. Journal of the History of Ideas, 22(4): 475–492, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708026.

Spivakovsky:1965:DHM

[Spi65] Erika Spivakovsky. Diego Hurtado de Mendoza and Averroism. Journal of the History of Ideas, 26(3):307–326, July/September 1965. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708384.

Spitz:1966:BRO

[Spi66] Lewis W. Spitz. Book review: Occultism and Despair of Reason in Renaissance Thought. Journal of the History of Ideas, 27(3): 464–469, July/September 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708601.

Spindler:1976:LME

[Spi76] Frank MacDonald Spindler. Lamennais and Montalvo: a European influence upon Latin American political thought. Journal of the History of Ideas, 37(1):137–146, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708715.

Spindler:1980:FBC

[Spi80] Frank MacDonald Spindler. Francisco Bilbao, Chilean disciple of Lamennais. Journal of the History of Ideas, 41(3):487–496, July/ September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709408.

Sprunger:1968:TPP

[Spr68] Keith L. Sprunger. Technometria: a prologue to Puritan theology. Journal of the History of Ideas, 29(1):115–122, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708469.

Sprunger:1976:JYN

[Spr76] Keith L. Sprunger. John Yates of Norfolk: The radical Puritan preacher as Ramist philosopher. Journal of the History of Ideas, 37 (4):697–706, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709032.

Sprinker:1980:GMH

[Spr80] Michael Sprinker. Gerard Manley Hopkins on the origin of language. Journal of the History of Ideas, 41(1):113–128, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709105.

Springborg:1994:HHH

[Spr94] Patricia Springborg. Hobbes, heresy, and the Historia Ecclesiastica. Journal of the History of Ideas, 55(4):553–571, October 1994. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709922.

Spurr:1988:RRR

[Spu88] John Spurr. “Rational religion” in Restoration England. Journal of the History of Ideas, 49(4):563–585, October/December 1988. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709674.

Stewart-Robertson:1981:WPS

[SR81] J. C. Stewart-Robertson. The well-principled savage, or the child of the Scottish Enlightenment. Journal of the History of Ideas, 42 (3):503–525, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709190.

Stewart-Robertson:1984:TRA

[SRN84] J. C. Stewart-Robertson and David Fate Norton. Thomas Reid on Adam Smith’s theory of morals. Journal of the History of Ideas, 45(2):309–321, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709295.

Stapleton:1941:HR

[Sta41] Laurence Stapleton. Halifax and Raleigh. Journal of the History of Ideas, 2(2):211–224, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707113.

Stanley:1957:HHT

[Sta57] Oma Stanley. T. H. Huxley’s treatment of ’nature’. Journal of the History of Ideas, 18(1):120–127, January 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707583.

Starkman:1959:QPJ

[Sta59] Miriam K. Starkman. Quakers, phrenologists, and Jonathan Swift. Journal of the History of Ideas, 20(3):403–412, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708117.

Starr:1967:CFI

[Sta67a] G. A. Starr. From casuistry to fiction: The importance of the Athenian Mercury. Journal of the History of Ideas, 28(1):17–32, January/ March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708478.

Stayer:1967:MRB

[Sta67b] James M. Stayer. The Münsterite rationalization of Bernhard Rothmann. Journal of the History of Ideas, 28(2):179–192, April/June 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708414.

Staudenbaur:1968:GFH

[Sta68] C. A. Staudenbaur. Galileo, Ficino, and Henry More’s psychathanasia. Journal of the History of Ideas, 29(4):565–578, October/ December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708294.

Stahlman:1973:BRR

[Sta73] William D. Stahlman. Book review: On Recent Copernicana. Journal of the History of Ideas, 34(3):483–489, July/September 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708968.

Staudenbaur:1974:BRP

[Sta74] C. A. Staudenbaur. Book review: Platonism, Theosophy, and Immaterialism: Recent Views of the Cambridge Platonists. Journal of the History of Ideas, 35(1):157–169, January/March 1974. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708752.

Starobinski:1977:EL

[Sta77] Jean Starobinski. Eloquence and liberty. Journal of the History of Ideas, 38(2):195–210, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708907.

Staum:1987:IRS

[Sta87] Martin S. Staum. Individual rights and social control: Political science in the French institute. Journal of the History of Ideas, 48 (3):411–430, July/September 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709760.

Staley:1989:AKC

[Sta89] Kevin Staley. Al-Kindi on creation: Aristotle’s challenge to Islam. Journal of the History of Ideas, 50(3):355–370, July/September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709566.

Stapleton:1991:EPC

[Sta91] Julia Stapleton. English pluralism as cultural definition: The social and political thought of George Unwin. Journal of the History of Ideas, 52(4):665–684, October/December 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709971.

Staum:1995:PPP

[Sta95] Martin Staum. Physiognomy and phrenology at the Paris athenee. Journal of the History of Ideas, 56(3):443–462, July 1995. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710035.

Staum:2004:NNF

[Sta04] Martin S. Staum. Nature and nurture in French ethnography and anthropology, 1859–1914. Journal of the History of Ideas, 65(3):475– 495, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/177669.

Stanley:2008:PMD

[Sta08a] Matthew Stanley. The pointsman: Maxwell’s Demon, Victorian free will, and the boundaries of science. Journal of the History of Ideas, 69(3):467–491, July 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/242281.

Stark:2008:PTA

[Sta08b] Ryan J. Stark. Protestant theology and apocalyptic rhetoric in Roger Ascham’s The Schoolmaster. Journal of the History of Ideas, 69(4):517–532, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/251817.

Staley:2009:JME

[Sta09a] Thomas W. Staley. The journal mind in its early years, 1876–1920: an introduction. Journal of the History of Ideas, 70(2):259–263, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/261978.

Staley:2009:KPM

[Sta09b] Thomas W. Staley. Keeping philosophy in mind: Shadworth H. Hodgson’s articulation of the boundaries of philosophy and science. Journal of the History of Ideas, 70(2):289–315, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/261980.

Stacey:2014:DDD

[Sta14] Peter Stacey. Definition, division, and difference in Machiavelli’s political philosophy. Journal of the History of Ideas, 75(2):189–212, April 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/544558.

Stalnaker:2016:JID

[Sta16] Joanna Stalnaker. Jonathan Israel in dialogue. Journal of the History of Ideas, 77(4):637–648, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/635632.

Steiner:1940:MSB

[Ste40] Arpad Steiner. A mirror for scholars of the Baroque. Journal of the History of Ideas, 1(3):320–334, June 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707090.

Stewart:1949:DS

[Ste49] H. L. Stewart. Dante and the schoolmen. Journal of the History of Ideas, 10(3):357–373, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707042.

Stedman:1951:DAC

[Ste51] Murray S. Stedman, Jr. “Democracy” in American communal and socialist literature. Journal of the History of Ideas, 12(1):147–154, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707542.

Stenerson:1953:EAT

[Ste53] Douglas C. Stenerson. Emerson and the agrarian tradition. Journal of the History of Ideas, 14(1):95–115, January 1953. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707497.

Stewart:1958:APH

[Ste58] Keith Stewart. Ancient poetry as history in the 18th Century. Journal of the History of Ideas, 19(3):335–347, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708039.

Steadman:1960:MFD

[Ste60] John M. Steadman. “Man’s first disobedience”: The causal structure of the Fall. Journal of the History of Ideas, 21(2):180–197, April/ June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708193.

Steadman:1962:FER

[Ste62] John M. Steadman. Felicity and end in Renaissance epic and ethics. Journal of the History of Ideas, 23(1):117–132, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708061.

Steadman:1965:EDC

[Ste65] John M. Steadman. Eve’s dream and the conventions of witchcraft. Journal of the History of Ideas, 26(4):567–574, October/December 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708500.

Steadman:1967:LRE

[Ste67a] John M. Steadman. Leviathan and Renaissance etymology. Journal of the History of Ideas, 28(4):575–576, October/December 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708530.

Stempel:1967:GMC

[Ste67b] Daniel Stempel. The garden: Marvell’s Cartesian ecstasy. Journal of the History of Ideas, 28(1):99–114, January/March 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708484.

Steinkraus:1973:MLK

[Ste73] Warren E. Steinkraus. Martin Luther King’s personalism and nonviolence. Journal of the History of Ideas, 34(1):97–111, January/ March 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708946.

Stepelevich:1974:RMS

[Ste74] Lawrence S. Stepelevich. The revival of max stirner. Journal of the History of Ideas, 35(2):323–328, April/June 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708766.

Stempel:1975:ARS

[Ste75] Daniel Stempel. Angels of reason: Science and myth in the Enlightenment. Journal of the History of Ideas, 36(1):63–78, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709011.

Stepelevich:1978:MSL

[Ste78a] Lawrence S. Stepelevich. Max Stirner and Ludwig Feuerbach. Journal of the History of Ideas, 39(3):451–463, July/September 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709388.

Stephens:1978:JLE

[Ste78b] Lester D. Stephens. Joseph Leconte’s evolutional idealism: a Lamarckian view of cultural history. Journal of the History of Ideas, 39 (3):465–480, July/September 1978. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709389.

Sternbach:1978:PDJ

[Ste78c] Robert Sternbach. Pascal and Dr. Johnson on immortality. Journal of the History of Ideas, 39(3):483–489, July/September 1978. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709391.

Stewart:1981:SCN

[Ste81] Larry Stewart. Samuel Clarke, Newtonianism, and the factions of Post-Revolutionary England. Journal of the History of Ideas, 42 (1):53–72, January/March 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709417.

Steinberg:1982:TTT

[Ste82] Michael Steinberg. The twelve tables and their origins: an Eighteenth-Century debate. Journal of the History of Ideas, 43(3):379–396, July/ September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709429.

Stepelevich:1985:MSH

[Ste85a] Lawrence S. Stepelevich. Max Stirner as Hegelian. Journal of the History of Ideas, 46(4):597–614, October/December 1985. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709548.

Stern:1985:HIH

[Ste85b] Laurent Stern. Hermeneutics and intellectual history. Journal of the History of Ideas, 46(2):287–296, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709640.

Stephanson:1988:RNP

[Ste88] Raymond Stephanson. Richard’s “Nerves”: The physiology of sensibility in clarissa. Journal of the History of Ideas, 49(2):267–285, April/June 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709500.

Steigerwald:1989:SPW

[Ste89] David Steigerwald. The synthetic politics of Woodrow Wilson. Journal of the History of Ideas, 50(3):465–484, July/September 1989. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709572.

Stern:2018:MKA

[Ste18] Eliyahu Stern. Marx and the Kabbalah: Aaron Shemuel Lieberman’s materialist interpretation of Jewish history. Journal of the History of Ideas, 79(2):285–307, April 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/693305.

Stierle:1998:TMC

[Sti98] Karlheinz Stierle. Three moments in the crisis of exemplarity: Boccaccio-Petrarch, Montaigne, and Cervantes. Journal of the History of Ideas, 59(4):581–595, October 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15012.

Stiles:2009:LMH

[Sti09] Anne Stiles. Literature in mind: H. G. Wells and the evolution of the mad scientist. Journal of the History of Ideas, 70(2):317–339, April 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/261981.

Stoll:1946:DOO

[Sto46] Elmer Edgar Stoll. The downfall of oratory: Our undemocratic arts. Journal of the History of Ideas, 7(1):3–34, January 1946. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707266.

Stoll:1951:SMD

[Sto51] Elmer E. Stoll. Symbolism in Moby-Dick. Journal of the History of Ideas, 12(3):440–465, June 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707754.

Stolnitz:1961:BSS

[Sto61] Jerome Stolnitz. ‘Beauty’: Some stages in the history of an idea. Journal of the History of Ideas, 22(2):185–204, April/June 1961. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707832.

Stocking:1962:LAS

[Sto62] George W. Stocking, Jr. Lamarckianism in American social science: 1890–1915. Journal of the History of Ideas, 23(2):239–256, April/ June 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708157.

Stone:1985:PTS

[Sto85] Harold Stone. The polemics of toleration: The scholars and publishers of Cassiodorus’ Variae. Journal of the History of Ideas, 46(2):147–165, April/June 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709632.

Strauss:1949:PPH

[Str49] Leo Strauss. Political philosophy and history. Journal of the History of Ideas, 10(1):30–50, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707198.

Stromberg:1951:HEC

[Str51a] R. N. Stromberg. History in the Eighteenth Century. Journal of the History of Ideas, 12(2):295–304, April 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707520.

Strong:1951:NMW

[Str51b] E. W. Strong. Newton’s “Mathematical Way”. Journal of the History of Ideas, 12(1):90–110, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707539.

Strong:1952:NG

[Str52] E. W. Strong. Newton and God. Journal of the History of Ideas, 13(2): 147–167, April 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707608.

Strout:1954:FHM

[Str54] Cushing Strout. Faith and history: The mind of William G. T. Shedd. Journal of the History of Ideas, 15(1):153–162, January 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707655.

Strong:1955:WWJ

[Str55] E. W. Strong. William Whewell and John Stuart Mill: Their controversy about scientific knowledge. Journal of the History of Ideas, 16(2):209–231, April 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707663.

Stromberg:1956:ICS

[Str56] Roland N. Stromberg. The idea of collective security. Journal of the History of Ideas, 17(2):250–263, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707745.

Strong:1957:NEN

[Str57] E. W. Strong. Newtonian explications of natural philosophy. Journal of the History of Ideas, 18(1):49–83, January 1957. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707579.

Stromberg:1972:UOA

[Str72] Roland N. Stromberg. Uncertainties and obscurities about the League of Nations. Journal of the History of Ideas, 33(1):139–154, January/ March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709062.

Stroud:1976:CTL

[Str76a] Michael Stroud. Chivalric terminology in late medieval literature. Journal of the History of Ideas, 37(2):323–334, April/June 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708828.

Struhl:1976:MNS

[Str76b] Paula Rothenberg Struhl. Mill’s notion of social responsibility. Journal of the History of Ideas, 37(1):155–162, January/March 1976. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708717.

Struever:2005:HP

[Str05] Nancy S. Struever. Historical priorities. Journal of the History of Ideas, 66(4):541–556, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/191477.

Struck:2016:ALL

[Str16] Peter T. Struck. 2013 Arthur O. Lovejoy Lecture: a cognitive history of divination in Ancient Greece. Journal of the History of Ideas, 77(1): 1–25, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
610558.

Streeter:2021:CTA

[Str21] Joseph Streeter. Conceptions of tolerance in antiquity and late antiquity. Journal of the History of Ideas, 82(3):357–376, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/798270.

Strassfeld:2022:ALM

[Str22] Jonathan Strassfeld. “I am aware that this letter may be offensive”: The unapologetic achievements of Ruth Barcan Marcus and Marjorie Glicksman Grene. Journal of the History of Ideas, 83(4):579–600, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
865893.

Stray:2023:CGL

[Str23] Christopher Stray. The Cambridge Greek lexicon: an essay-review. Journal of the History of Ideas, 84(2):391–408, April 2023. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/886187.

Stuurman:1989:SVH

[Stu89] Siep Stuurman. Samuel Van Houten and Dutch liberalism, 1860–90. Journal of the History of Ideas, 50(1):135–152, January/March 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709790.

Stuurman:1997:SCF

[Stu97] Siep Stuurman. Social Cartesianism: François Poulain de la Barre and the origins of the Enlightenment. Journal of the History of Ideas, 58(4):617–640, October 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/14968.

Stuurman:2004:VTR

[Stu04] Siep Stuurman. The voice of thersites: Reflections on the origins of the idea of equality. Journal of the History of Ideas, 65(2):171–189, April 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/173853.

Stuurman:2007:CEE

[Stu07] Siep Stuurman. Cosmopolitan egalitarianism in the Enlightenment: Anquetil duperron on India and America. Journal of the History of Ideas, 68(2):255–278, April 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/215057.

Styfhals:2015:EHK

[Sty15] Willem Styfhals. Evil in history: Karl Löwith and Jacob Taubes on modern eschatology. Journal of the History of Ideas, 76(2):191– 213, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
580295.

Styfhals:2019:MTO

[Sty19] Willem Styfhals. Modernity as theodicy: Odo Marquard reads Hans Blumenberg’s The Legitimacy of the Modern Age. Journal of the History of Ideas, 80(1):113–131, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/715739.

Sullivan:1983:LIJ

[Sul83] Eileen P. Sullivan. Liberalism and imperialism: J. S. Mill’s defense of the British Empire. Journal of the History of Ideas, 44(4): 599–617, October/December 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709218.

Sulek:1989:EPF

[Suł89] Antoni Sułek. The experiment of psammetichus: Fact, fiction, and model to follow. Journal of the History of Ideas, 50(4):645–651, October/December 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709802.

Summers:2017:AVP

[Sum17] Carol Summers. Adolescence versus politics: Metaphors in late colonial Uganda. Journal of the History of Ideas, 78(1):117–136, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/648125.

Suppes:1954:DPA

[Sup54] Patrick Suppes. Descartes and the problem of action at a distance. Journal of the History of Ideas, 15(1):146–152, January 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707654.

Susato:2012:TTP

[Sus12] Ryu Susato. Taming “The Tyranny of Priests”: Hume’s advocacy of religious establishments. Journal of the History of Ideas, 73(2):273– 293, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
475340.

Sutton:1953:PLP

[Sut53] Robert B. Sutton. The phrase libertas philosophandi. Journal of the History of Ideas, 14(2):310–316, April 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707480.

Sutcliffe:2003:JAR

[Sut03] Adam Sutcliffe. Judaism in the anti-religious thought of the clandestine French Early Enlightenment. Journal of the History of Ideas, 64(1):97–117, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/42180.

Swart:1962:IMN

[Swa62] Koenraad W. Swart. “Individualism” in the Mid-Nineteenth Century (1826–1860). Journal of the History of Ideas, 23(1):77–90, January/ March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708058.

Sweet:1973:YWH

[Swe73] Paul R. Sweet. Young Wilhelm Von Humboldt’s writings (1789–93) reconsidered. Journal of the History of Ideas, 34(3):469–482, July/ September 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708967.

Swerdlow:1993:MLH

[Swe93] N. M. Swerdlow. Montucla’s legacy: The history of the exact sciences. Journal of the History of Ideas, 54(2):299–328, April 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709984.

Sweet:1999:BBT

[Swe99] Dennis Sweet. The birth of the birth of tragedy. Journal of the History of Ideas, 60(2):345–359, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/15032.

Swinton:1975:HIG

[Swi75] W. E. Swinton. Historical interrelations of geology and other sciences. Journal of the History of Ideas, 36(4):729–738, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709000.

Sypher:1941:APL

[Syp41] Wylie Sypher. The African Prince in London. Journal of the History of Ideas, 2(2):237–247, April 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707115.

Sypher:1963:PHF

[Syp63] George Wylie Sypher. La Popeliniere’s Histoire de France: a case of historical objectivity and religious censorship. Journal of the History of Ideas, 24(1):41–54, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707858.

Sypher:1965:SBS

[Syp65] G. Wylie Sypher. Similarities between the scientific and the historical revolutions at the end of the Renaissance. Journal of the History of Ideas, 26(3):353–368, July/September 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708386.

Tzoref-Ashkenazi:2006:IIE

[TA06] Chen Tzoref-Ashkenazi. India and the identity of Europe: The case of Friedrich Schlegel. Journal of the History of Ideas, 67(4):713–734, October 2006. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
204771.

Tadd:2022:GLS

[Tad22] Misha Tadd. Global laozegetics: a study in globalized philosophy. Journal of the History of Ideas, 83(1):87–109, January 2022. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/847822.

Taeusch:1942:CUH

[Tae42] Carl F. Taeusch. The concept of “Usury” the history of an idea. Journal of the History of Ideas, 3(3):291–318, June 1942. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707306.

Taeusch:1950:AWM

[Tae50] Carl F. Taeusch. On Addison W. Moore. Journal of the History of Ideas, 11(4):490, October 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707593.

Taeusch:1953:RVD

[Tae53] Carl F. Taeusch. The religious view of the doctrine of natural rights. Journal of the History of Ideas, 14(1):51–67, January 1953. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707495.

Tager:1986:MPW

[Tag86] Michael Tager. Myth and politics in the works of Sorel and Barthes. Journal of the History of Ideas, 47(4):625–639, October/December 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709722.

Tanner:1977:LLP

[Tan77] Tony Tanner. Licence and licencing: To the presse or to the spunge. Journal of the History of Ideas, 38(1):3–18, January/March 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708837.

Tanay:1998:NFC

[Tan98] Dorit Esther Tanay. “Nos faysoms contre nature … ”: Fourteenth-Century Sophismata and the musical avant garde. Journal of the History of Ideas, 59(1):29–51, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14976.

Tang:1999:LTL

[Tan99] Yanfang Tang. Language, truth, and literary interpretation: a crosscultural examination. Journal of the History of Ideas, 60(1):1–20, January 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15022.

Tashiro:1969:EPE

[Tas69] Tom T. Tashiro. English poets, Egyptian onions, and the Protestant view of the Eucharist. Journal of the History of Ideas, 30(4): 563–578, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708611.

Tatarkiewicz:1963:CAA

[Tat63a] W. Tatarkiewicz. Classification of arts in antiquity. Journal of the History of Ideas, 24(2):231–240, April/June 1963. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707847.

Tatarkiewicz:1963:NMS

[Tat63b] W. Tatarkiewicz. A note on the modern system of the arts. Journal of the History of Ideas, 24(3):422, July/September 1963. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708218.

Tatarkiewicz:1965:TSE

[Tat65] W. Tatarkiewicz. Theatrica, the science of entertainment: From the XIIth to the XVIIth Century. Journal of the History of Ideas, 26(2):263–272, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708232.

Taylor:1955:CEI

[Tay55] J. Taylor. Copernicus on the evils of inflation and the establishment of a sound currency. Journal of the History of Ideas, 16(4):540–547, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707509.

Tayler:1981:MMH

[Tay81] Edward W. Tayler. In memoriam: Marjorie Hope Nicolson (1894– 1981). Journal of the History of Ideas, 42(4):665–667, October/ December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709124.

Taylor:1982:NRI

[Tay82] Brian Taylor. A note in response to Itzkin’s “Bentham’s Chrestomathia: Utilitarian Legacy to English Education”. Journal of the History of Ideas, 43(2):309–313, April/June 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709206. See [Itz78].

Taylor:1998:APA

[Tay98] Richard C. Taylor. Aquinas, the Plotiniana Arabica, and the metaphysics of being and actuality. Journal of the History of Ideas, 59(2): 217–239, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14987.

Taylor:2001:MAI

[Tay01] E. Derek Taylor. Mary Astell’s ironic assault on John Locke’s theory of thinking matter. Journal of the History of Ideas, 62(3):505–522, July 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15127.

Taylor:2017:BPJ

[Tay17] Simon W. Taylor. Between philosophy and Judaism: Leo Strauss’s skeptical engagement with Zionism. Journal of the History of Ideas, 78(1):95–116, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/648124.

Tung-Chi:1947:CMT

[TC47] Lin Tung-Chi. The Chinese mind: Its Taoist substratum. Journal of the History of Ideas, 8(3):259–272, June 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707142.

Teall:1962:WCE

[Tea62] John L. Teall. Witchcraft and Calvinism in Elizabethan England: Divine power and human agency. Journal of the History of Ideas, 23(1):21–36, January/March 1962. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708055.

Teggart:1940:PHI

[Teg40] Frederick J. Teggart. A problem in the history of ideas. Journal of the History of Ideas, 1(4):494–503, October 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707126.

Teggart:1942:CHE

[Teg42] Frederick J. Teggart. Causation in historical events. Journal of the History of Ideas, 3(1):3–11, January 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707458.

Teggart:1947:AHW

[Teg47] Frederick J. Teggart. The argument of Hesiod’s works and days. Journal of the History of Ideas, 8(1):45–77, January 1947. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707441.

Teichgraeber:1977:HPP

[Tei77] Richard Teichgraeber. Hegel on property and poverty. Journal of the History of Ideas, 38(1):47–64, January/March 1977. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708840.

Ten:1969:ML

[Ten69] C. L. Ten. Mill and liberty. Journal of the History of Ideas, 30 (1):47–68, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708244.

Tennant:1982:ARL

[Ten82] R. C. Tennant. The Anglican response to Locke’s theory of personal identity. Journal of the History of Ideas, 43(1):73–90, January/March 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709161.

Terracciano:2018:OPK

[Ter18] Pasquale Terracciano. The Origen of Pico’s Kabbalah: Esoteric wisdom and the dignity of man. Journal of the History of Ideas, 79(3): 343–361, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
703662.

Terzi:2021:CFU

[Ter21] Pietro Terzi. Contingency, freedom, and uchronic narratives: Charles Renouvier’s philosophy of history in the shadow of the Franco– Prussian War. Journal of the History of Ideas, 82(2):257–278, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/790203.

Thayer:1975:PQP

[Tha75] H. S. Thayer. Plato’s quarrel with poetry: Simonides. Journal of the History of Ideas, 36(1):3–26, January/March 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709008.

Thatcher:1983:NDL

[Tha83] David S. Thatcher. Nietzsche’s debt to Lubbock. Journal of the History of Ideas, 44(2):293–309, April/June 1983. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709141.

Theiss:2023:AAH

[The23] William Theiss. The Abbé d’Aubignac’s Homer and the culture of the street in Seventeenth-Century Paris. Journal of the History of Ideas, 84(1):77–102, January 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/879541.

Thiem:1979:GLA

[Thi79] Jon Thiem. The great library of Alexandria burnt: Towards the history of a symbol. Journal of the History of Ideas, 40(4):507–526, October/December 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709356.

Thijssen:1992:DHJ

[Thi92] J. M. M. H. Thijssen. David Hume and John Keill and the structure of continua. Journal of the History of Ideas, 53(2):271–286, April/ June 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709874.

Thorpe:1944:CAC

[Tho44] Clarence D. Thorpe. Coleridge as aesthetician and critic. Journal of the History of Ideas, 5(4):387–414, October 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707078.

Thorndike:1951:NCN

[Tho51] Lynn Thorndike. Newness and craving for novelty in Seventeenth-Century science and medicine. Journal of the History of Ideas, 12(4):584–598, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707487.

Thorndike:1955:CSS

[Tho55] Lynn Thorndike. Censorship by the Sorbonne of science and superstition in the first half of the Seventeenth Century. Journal of the History of Ideas, 16(1):119–125, January 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707531.

Thorndike:1958:JJG

[Tho58] Lynn Thorndike. Jean De jandun on gravitation. Journal of the History of Ideas, 19(2):253–255, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707938.

Thomas:1959:DS

[Tho59] Keith Thomas. The double standard. Journal of the History of Ideas, 20(2):195–216, April 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707819.

Thorslev:1975:RLC

[Tho75] Peter L. Thorslev, Jr. Romanticism and the literary consciousness. Journal of the History of Ideas, 36(3):563–572, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708665.

Thompson:1977:ICC

[Tho77] M. P. Thompson. The idea of conquest in controversies over the 1688 revolution. Journal of the History of Ideas, 38(1):33–46, January/ March 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2708839.

Thompson:1994:IED

[Tho94] Martyn P. Thompson. Ideas of Europe during the French Revolution and Napoleonic wars. Journal of the History of Ideas, 55(1):37–58, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709952.

Thomson:2007:CLE

[Tho07] Erik Thomson. Commerce, law, and erudite culture: The mechanics of Théodore Godefroy’s service to Cardinal Richelieu. Journal of the History of Ideas, 68(3):407–427, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/219825.

Thompson:2012:LHI

[Tho12] Martyn P. Thompson. The logic of the history of ideas: Mark Bevir and Michael Oakeshott. Journal of the History of Ideas, 73(4):593– 607, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
488350.

Thomas:2024:SEC

[Tho24] Tyler J. Thomas. The strategic emergence of Cartesianism: Descartes, public controversy, and the quarrel of Utrecht. Journal of the History of Ideas, 85(4):749–771, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/944584.

Thrane:1977:BPO

[Thr77] Gary Thrane. Berkeley’s “Proper Object of Vision”. Journal of the History of Ideas, 38(2):243–260, April/June 1977. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708910.

Throntveit:2011:WJE

[Thr11] Trygve Throntveit. William James’s ethical republic. Journal of the History of Ideas, 72(2):255–277, April 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/428854.

Thurston:1976:ATR

[Thu76] G. J. Thurston. Alexis De Tocqueville in Russia. Journal of the History of Ideas, 37(2):289–306, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708825.

Tierney:1954:OCT

[Tie54] Brian Tierney. Ockham, the conciliar theory, and the canonists. Journal of the History of Ideas, 15(1):40–70, January 1954. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707649.

Tierney:1963:NIE

[Tie63] Brian Tierney. Natura id est deus: a case of juristic pantheism? Journal of the History of Ideas, 24(3):307–322, July/September 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708209.

Tierney:1991:MR

[Tie91] Brian Tierney. Marsilius on rights. Journal of the History of Ideas, 52(1):3–17, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709579.

Tierney:2001:KPP

[Tie01a] Brian Tierney. Kant on property: The problem of permissive law. Journal of the History of Ideas, 62(2):301–312, April 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15117.

Tierney:2001:PNL

[Tie01b] Brian Tierney. Permissive natural law and property: Gratian to Kant. Journal of the History of Ideas, 62(3):381–399, July 2001. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15128.

Tierney:2011:RAS

[Tie11] Brian Tierney. Response to S. Adam Seagrave’s “How Old Are Modern Rights?: On the Lockean Roots of Contemporary Human Rights Discourse”. Journal of the History of Ideas, 72(3):461–468, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/445506. See [Sea11].

Tigerstedt:1970:FPP

[Tig70] E. N. Tigerstedt. Furor poeticus: Poetic inspiration in Greek literature before Democritus and Plato. Journal of the History of Ideas, 31(2):163–178, April/June 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708543.

Tillich:1944:EP

[Til44] Paul Tillich. Existential philosophy. Journal of the History of Ideas, 5(1):44–70, January 1944. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707101.

Tillich:1945:NBS

[Til45] Paul Tillich. Nietzsche and the bourgeois spirit. Journal of the History of Ideas, 6(3):307–309, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707294.

Timmermans:1999:ODC

[Tim99] B Timmermans. The originality of Descartes’s conception of analysis as discovery. Journal of the History of Ideas, 60(3):433–447, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/15042.

Tindall:1954:JJH

[Tin54] William York Tindall. James Joyce and the hermetic tradition. Journal of the History of Ideas, 15(1):23–39, January 1954. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707648.

Tinkler:1988:SHB

[Tin88] John F. Tinkler. The splitting of humanism: Bentley, Swift, and the English battle of the books. Journal of the History of Ideas, 49 (3):453–472, July/September 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709487.

Tinsley:1996:SGP

[Tin96] Barbara Sher Tinsley. Sozzini’s ghost: Pierre Bayle and Socinian toleration. Journal of the History of Ideas, 57(4):609–624, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14926.

Tipton:1969:TQB

[Tip69] Ian Tipton. Two questions on Bishop Berkeley’s panacea. Journal of the History of Ideas, 30(2):203–224, April/June 1969. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708433.

Tobin:1974:VBE

[Tob74] Rosemary Barton Tobin. Vincent of Beauvais on the education of women. Journal of the History of Ideas, 35(3):485–489, July/ September 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708795.

Todd:1974:ISG

[Tod74] Nigel Todd. Ideological superstructure in Gramsci and Mao Tse-Tung. Journal of the History of Ideas, 35(1):148–156, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708751.

Todd:1976:FCA

[Tod76] Robert B. Todd. The four causes: Aristotle’s exposition and the ancients. Journal of the History of Ideas, 37(2):319–322, April/June 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708827.

Tollebeek:1998:HRN

[Tol98] Jo Tollebeek. Historical representation and the nation-state in Romantic Belgium (1830–1850). Journal of the History of Ideas, 59(2):

329–353, April 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14988.

Tonsor:1959:LAD

[Ton59] Stephen J. Tonsor. Lord Acton on Dollinger’s historical theology. Journal of the History of Ideas, 20(3):329–352, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708113.

Tonsor:1964:HME

[Ton64] Stephen J. Tonsor. The historical morphology of Ernst von lasaulx. Journal of the History of Ideas, 25(3):374–392, July/September 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707905.

Tormey:1984:BRS

[Tor84] Judith Tormey. Book review: Some Recent Works on Historical Attitudes Toward Women. Journal of the History of Ideas, 45(4): 619–623, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709377.

Toulmin:1957:CEP

[Tou57] S. E. Toulmin. Crucial experiments: Priestley and Lavoisier. Journal of the History of Ideas, 18(2):205–220, April 1957. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707624.

Touey:1998:HNE

[Tou98] Daniel Touey. How Nietzsche explains and why. Journal of the History of Ideas, 59(3):485–498, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/1/

article/14997.

Touber:2013:SPS

[Tou13] Jetze Touber. Stones of passion: Stones in the internal organs as liminal phenomena between medical and religious knowledge in Renaissance Italy. Journal of the History of Ideas, 74(1):23–44, January 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/495240.

Townsend:1987:SKD

[Tow87] Dabney Townsend. From Shaftesbury to Kant: The development of the concept of aesthetic experience. Journal of the History of Ideas, 48(2):287–305, April/June 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709559.

Townsend:2019:GDP

[Tow19] Dabney Townsend. On genius: The development of a philosophical concept of genius in Eighteenth-Century Britain. Journal of the History of Ideas, 80(4):555–574, October 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/735381.

Toynbee:1955:C

[Toy55] Arnold J. Toynbee. Comment. Journal of the History of Ideas, 16 (3):421, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707643.

See [Fie55, Gey55].

Trattner:1964:GEE

[Tra64] Walter I. Trattner. God and expansion in Elizabethan England: John Dee, 1527–1583. Journal of the History of Ideas, 25(1):17–34, January/March 1964. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708083.

Trentman:1967:USN

[Tre67] John Trentman. The Uppsala School and the new logic. Journal of the History of Ideas, 28(3):445–450, July/September 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708630.

Tregenza:2021:SSU

[Tre21] Ian Tregenza. The “Servile State” down under: Hilaire Belloc and Australian political thought, 1912–53. Journal of the History of Ideas, 82(2):305–327, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/790205.

Trinkaus:1949:PFW

[Tri49] Charles Trinkaus. The problem of free will in the Renaissance and the reformation. Journal of the History of Ideas, 10(1):51–62, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707199.

Trimble:1950:ETH

[Tri50] William Raleigh Trimble. Early Tudor historiography, 1485–1548. Journal of the History of Ideas, 11(1):30–41, January 1950. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707450.

Trinkaus:1956:BRH

[Tri56] Charles Trinkaus. Book reviews: Humanistic and Political Literature in Florence and Venice at the Beginning of the Quattrocento: Studies in Criticism and Chronology by Hans Baron; The Crisis of the Early Italian Renaissance: Civic Humanism and Republican Liberty in an Age of Classicism and Tyranny by Hans Baron. Journal of the History of Ideas, 17(3):426–432, June 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707555.

Trimpi:1969:MUD

[Tri69] Helen P. Trimpi. Melville’s use of demonology and witchcraft in Moby-Dick. Journal of the History of Ideas, 30(4):543–562, October/ December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708610.

Trinkaus:1987:AVM

[Tri87] Charles Trinkaus. Antiquitas versus modernitas: an Italian humanist polemic and its resonance. Journal of the History of Ideas, 48 (1):11–21, January/March 1987. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709609. See comment [Gil87].

Trinkaus:1990:BRR

[Tri90] Charles Trinkaus. Book review: Renaissance Ideas and the Idea of the Renaissance. Journal of the History of Ideas, 51(4):667–684, October/ December 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709652.

Tribby:1991:CCC

[Tri91] Jay Tribby. Cooking (with) Clio and Cleo: Eloquence and experiment in Seventeenth-Century Florence. Journal of the History of Ideas, 52 (3):417–439, July/September 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710045.

Trinkaus:1996:LVP

[Tri96] Charles Edward Trinkaus. Lorenzo Valla on the problem of speaking about the Trinity. Journal of the History of Ideas, 57(1):27–53, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14894.

Triplett:2004:MFS

[Tri04] Hall Triplett. The misnomer of Freud’s “Seduction Theory”. Journal of the History of Ideas, 65(4):647–665, October 2004. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/184074.

Trompf:1973:CCR

[Tro73] G. W. Trompf. The concept of the Carolingian Renaissance. Journal of the History of Ideas, 34(1):3–26, January/March 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708941.

Truitt:2012:CDA

[Tru12] E. R. Truitt. Celestial divination and Arabic science in Twelfth-Century England: The history of Gerbert of Aurillac’s talking head. Journal of the History of Ideas, 73(2):201–222, April 2012. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/475341.

Tsao:2013:ECD

[Tsa13] Tiffany Tsao. Environmentalism and civilizational development in the colonial British histories of the Indian Archipelago (1783–1820). Journal of the History of Ideas, 74(3):449–471, July 2013. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/muse.jhu.edu/pub/56/article/511981.

Tsugawa:1961:DHL

[Tsu61] Albert Tsugawa. David Hume and Lord Kames on personal identity. Journal of the History of Ideas, 22(3):398–403, July/September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708135.

Tenorio-Trillo:2024:CCW

[TT24] Mauricio Tenorio-Trillo. Conceitos and conceptos: The weight of words in the Iberian world. Journal of the History of Ideas, 85(2): 389–417, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
926152.

Tuckness:2003:DRA

[Tuc03] Alex Scott Tuckness. Discourses of resistance in the American Revolution. Journal of the History of Ideas, 64(4):547–563, October 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/52455.

Tumbleson:1996:RRS

[Tum96] Raymond D. Tumbleson. “Reason and religion”: The science of Anglicanism. Journal of the History of Ideas, 57(1):131–156, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14895.

Tung:2014:VIS

[Tun14] Toy-Fung Tung. Vitoria’s ideas of supernatural and natural sovereignty: Adam and Eve’s marriage, the uncivil Amerindians, and the global Christian nation. Journal of the History of Ideas, 75(1):45– 68, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
535153.

Turbayne:1955:BMR

[Tur55] Colin M. Turbayne. Berkeley and Molyneux on retinal images. Journal of the History of Ideas, 16(3):339–355, June 1955. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707636.

Turner:1966:DAI

[Tur66] Ralph V. Turner. Descendit ad inferos: Medieval views on Christ’s descent into Hell and the salvation of the ancient just. Journal of the History of Ideas, 27(2):173–194, April/June 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708637.

Turck:1971:BRL

[Tur71] Dieter Turck. Book review: Leibniz’s Correspondence (1692). Journal of the History of Ideas, 32(4):627–630, October/December 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708984.

Turley:1975:PCS

[Tur75] Peter T. Turley. Peirce’s cosmic “Sheriff”. Journal of the History of Ideas, 36(4):717–720, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708998.

Tutino:2022:MMV

[Tut22] Stefania Tutino. The mystery of Mount Vesuvius’s crosses: Belief, credulity, and credibility in Post-Reformation Catholicism. Journal of the History of Ideas, 83(2):207–227, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/855164.

Tuve:1942:ILR

[Tuv42] Rosemond Tuve. Imagery and logic: Ramus and metaphysical poetics. Journal of the History of Ideas, 3(4):365–400, October 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707314.

Tuveson:1950:SWM

[Tuv50] Ernest Tuveson. Swift and the world-makers. Journal of the History of Ideas, 11(1):54–74, January 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707452.

Tuveson:1966:PGL

[Tuv66] Ernest Tuveson. “The pilot of the Galilean lake”. Journal of the History of Ideas, 27(3):447–458, July/September 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708598.

Tremml-Werner:2021:QPC

[TW21] Birgit Tremml-Werner. A question of political correctness: Translating friendship across time and space. Journal of the History of Ideas, 82(3):503–520, July 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/798277.

Twining:2020:EMD

[Twi20] Timothy Twining. The early modern debate over the age of the Hebrew vowel points: Biblical criticism and Hebrew scholarship in the Confessional Republic of Letters. Journal of the History of Ideas, 81(3):337–358, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/761033.

Tzamalikos:1991:OSV

[Tza91] Panayiotis Tzamalikos. Origen and the stoic view of time. Journal of the History of Ideas, 52(4):535–561, October/December 1991. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709965.

Ulreich:1971:PWF

[Ulr71] John C. Ulreich, Jr. A paradise within: The fortunate fall in Paradise Lost. Journal of the History of Ideas, 32(3):351–366, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708351.

Urbinati:2012:RAF

[Urb12] Nadia Urbinati. Republicanism after the French Revolution: The case of Sismonde de Sismondi. Journal of the History of Ideas, 73(1):95– 109, January 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
463140.

Urquhart:2014:CTH

[Urq14] Lela M. Urquhart. Competing traditions in the historiography of Ancient Greek colonization in Italy. Journal of the History of Ideas, 75(1):23–44, January 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/535152.

Vallins:1995:PEC

[Val95] David Vallins. Production and existence: Coleridge’s unification of nature. Journal of the History of Ideas, 56(1):107–124, January 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710009.

Valsania:2004:OOB

[Val04] Maurizio Valsania. “Our original barbarism”: Man vs. nature in Thomas Jefferson’s moral experience. Journal of the History of Ideas, 65(4):627–645, October 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/184075.

Valdez:2009:BRF

[Val09] Damian Valdez. Bachofen’s Rome and the fate of the feminine Orient. Journal of the History of Ideas, 70(3):421–443, July 2009. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/270105.

Valle:2022:ELA

[Val22] Tomás Antonio Valle. Eilhard Lubin, academic unorthodoxy, and the dynamics of confessional intellectual cultures. Journal of the History of Ideas, 83(2):181–206, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/855163.

Valmori:2024:BRD

[Val24] Niccolò Valmori. Breaking the revolutionary deadlock? Volney’s Leçons and the debate on the value of history. Journal of the History of Ideas, 85(4):773–794, October 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/944585.

Vann:1958:FAS

[Van58] Richard T. Vann. The free Anglo-Saxons: a historical myth. Journal of the History of Ideas, 19(2):259–272, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707940.

VanderZanden:1959:IWS

[Van59] James W. Vander Zanden. The ideology of white supremacy. Journal of the History of Ideas, 20(3):385–402, June/September 1959. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708116.

Vann:1965:LLG

[Van65] Richard T. Vann. The later life of gerrard winstanley. Journal of the History of Ideas, 26(1):133–136, January/March 1965. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708404.

VanDeWetering:1982:MPN

[Van82] Maxine Van De Wetering. Moralizing in Puritan natural science: Mysteriousness in earthquake sermons. Journal of the History of Ideas, 43(3):417–438, July/September 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709431.

VanPeursen:1993:WTA

[Van93] C. A. Van Peursen. E. W. Von Tschirnhaus and the Ars Inveniendi. Journal of the History of Ideas, 54(3):395–410, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2710020.

VanDerZande:1995:ICG

[Van95] Johan Van Der Zande. In the image of Cicero: German philosophy between Wolff and Kant. Journal of the History of Ideas, 56(3):419– 442, July 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2710034.

Vandenabeele:2001:NDK

[Van01] Bart Vandenabeele. On the notion of “Disinterestedness”: Kant, Lyotard, and Schopenhauer. Journal of the History of Ideas, 62 (4):705–720, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15137.

VanRooden:2015:RLA

[Van15] Aukje Van Rooden. Reconsidering literary autonomy: From an individual towards a relational paradigm. Journal of the History of Ideas, 76(2):167–190, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/580294.

Vanzo:2016:ERN

[Van16] Alberto Vanzo. Empiricism and rationalism in Nineteenth-Century histories of philosophy. Journal of the History of Ideas, 77(2):253– 282, April 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
620981.

Vanheeswijck:2017:PGT

[Van17] Guido M. Vanheeswijck. The philosophical genealogy of Taylor’s social imaginaries: a complex history of ideas and predecessors. Journal of the History of Ideas, 78(3):473–496, July 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/665524.

Vartanian:1950:TPM

[Var50] Aram Vartanian. Trembley’s polyp, La Mettrie, and Eighteenth-Century French materialism. Journal of the History of Ideas, 11(3): 259–286, June 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707732.

Varley:1953:IRK

[Var53] H. L. Varley. Imperialism and Rudyard Kipling. Journal of the History of Ideas, 14(1):124–135, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707499.

Vartanian:1973:BRD

[Var73] Aram Vartanian. Book review: Diderot and L’Encyclopédie. Journal of the History of Ideas, 34(2):303–311, April/June 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708734.

vanBunge:2017:SL

[vB17] Wiep van Bunge. Spinoza’s life: 1677–1802. Journal of the History of Ideas, 78(2):211–231, April 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/652823.

vanderLugt:2017:BMP

[vdL17] Mara van der Lugt. The body of Mahomet: Pierre Bayle on war, sex, and Islam. Journal of the History of Ideas, 78(1):27–50, January 2017. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/648121.

vanderZande:1998:MEC

[vdZ98] Johan van der Zande. The microscope of experience: Christian Garve’s translation of Cicero’s De Officiis (1783). Journal of the History of Ideas, 59(1):75–94, January 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14977.

vanderZande:2010:SHG

[vdZ10] Johan van der Zande. Statistik and history in the German Enlightenment. Journal of the History of Ideas, 71(3):411–432, July 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/386151.

vanEijnatten:2001:VTF

[vE01] Joris van Eijnatten. Vestige of the third force: Willem Bilderdijk, poet, anti-skeptic, millenarian. Journal of the History of Ideas, 62(2): 313–333, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
15110.

Velmet:2019:SAG

[Vel19] Aro Velmet. Sovereignty after gender trouble: Language, reproduction, and supranationalism in Estonia, 1980–2017. Journal of the History of Ideas, 80(3):455–478, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/729723.

Venturi:1963:OD

[Ven63] Franco Venturi. Oriental despotism. Journal of the History of Ideas, 24(1):133–142, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707864.

Venning:1976:HPC

[Ven76] Corey Venning. Hume on property, commerce, and empire in the good society: The role of historical necessity. Journal of the History of Ideas, 37(1):79–92, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708710.

Verene:1969:KHC

[Ver69] Donald Phillip Verene. Kant, Hegel, and Cassirer: The origins of the philosophy of symbolic forms. Journal of the History of Ideas, 30(1):33–46, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708243.

Vernon:1973:RCS

[Ver73] Richard Vernon. Rationalism and commitment in Sorel. Journal of the History of Ideas, 34(3):405–420, July/September 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708961.

Verdon:1982:LPH

[Ver82] Michel Verdon. On the laws of physical and human nature: Hobbes’ physical and social cosmologies. Journal of the History of Ideas, 43 (4):653–663, October/December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709348.

Vernon:1984:ACW

[Ver84] Richard Vernon. Auguste Comte and the withering-away of the state. Journal of the History of Ideas, 45(4):549–566, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709373.

Verbeek:1994:RFP

[Ver94] Theo Verbeek. Regius’s fundamenta physices. Journal of the History of Ideas, 55(4):533–551, October 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709921.

Verhaart:2022:ACD

[Ver22] Floris Verhaart. Across the confessional divide: Johannes Hoornbeeck, José de Acosta, and the role of force and free will in the development of a reformed missiology. Journal of the History of Ideas, 83(4):629–642, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/865884.

vanGorkom:2020:IKR

[vG20] Joris van Gorkom. Immanuel Kant on race mixing: The Gypsies, the Black Portuguese, and the Jews on St. Thomas. Journal of the History of Ideas, 81(3):407–427, July 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/761037.

vanIttersum:2011:KPD

[vI11] Martine J. van Ittersum. Knowledge production in the Dutch Republic: The household academy of Hugo Grotius. Journal of the History of Ideas, 72(4):523–548, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/453547.

Vickers:1990:DD

[Vic90] Brian Vickers. The dangers of dichotomy. Journal of the History of Ideas, 51(1):148–159, January/March 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709752.

Vickers:1992:FBP

[Vic92] Brian Vickers. Francis Bacon and the progress of knowledge. Journal of the History of Ideas, 53(3):495–518, July/September 1992. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709891.

Vick:2002:GOO

[Vic02] Brian E. Vick. Greek origins and organic metaphors: Ideals of cultural autonomy in neohumanist Germany from Winckelmann to Curtius. Journal of the History of Ideas, 63(3):483–500, July 2002. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15167.

Vickers:2008:FBF

[Vic08] Brian Vickers. Francis Bacon, feminist historiography, and the dominion of nature. Journal of the History of Ideas, 69(1):117–141, January 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/231015. See

response [Par08].

Vidal:1995:DAM

[Vid95] Mary Vidal. David among the moderns: Art, science, and the lavoisiers. Journal of the History of Ideas, 56(4):595–623, October 1995. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709994.

Viereck:1942:BRR

[Vie42] Peter Viereck. Book review: Reply by the author of Metapolitics. Journal of the History of Ideas, 3(1):110–112, January 1942. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707465. See [Bar42].

Viereck:1949:PMA

[Vie49] Peter Viereck. The poet in the machine age. Journal of the History of Ideas, 10(1):88–103, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707201.

Vieira:2003:MLV

[Vie03] Monica Brito Vieira. Mare Liberum vs. Mare Clausum: Grotius, Freitas, and Selden’s debate on dominion over the seas. Journal of the History of Ideas, 64(3):361–377, July 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/47665.

Vieira:2015:LCL

[Vie15] Mónica Brito Vieira. Leviathan contra Leviathan. Journal of the History of Ideas, 76(2):271–288, April 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/580299.

Vincent:2007:GPB

[Vin07] Andrew Vincent. German philosophy and British public policy: Richard Burdon Haldane in theory and practice. Journal of the History of Ideas, 68(1):157–179, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/210691.

Virtanen:1950:NAF

[Vir50] Reino Virtanen. Nietzsche and the action française: Nietzsche’s significance for French rightist thought. Journal of the History of Ideas, 11(2):191–214, April 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707246.

Virtanen:1986:CBP

[Vir86] Reino Virtanen. Claude Bernard’s prophecies and the historical relation of science to literature. Journal of the History of Ideas, 47(2):275–286, April/June 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709814.

Visser:1984:VPC

[Vis84] Margaret Visser. Vengeance and pollution in Classical Athens. Journal of the History of Ideas, 45(2):193–206, April/June 1984. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709287.

vanLiere:1997:VCC

[vL97] Katherine Elliot van Liere. Vitoria, Cajetan, and the conciliarists. Journal of the History of Ideas, 58(4):597–616, October 1997. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14963.

Vogt:1993:SFM

[Vog93] Philip Vogt. Seascape with fog: Metaphor in Locke’s essay. Journal of the History of Ideas, 54(1):1–18, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709857.

VonKuehnelt-Leddihn:1948:BBG

[Von48] Erik R. Von Kuehnelt-Leddihn. The Bohemian background of German National Socialism: The D.A.P., D.N.S.A.P. and N.S.D.A.P. Journal of the History of Ideas, 9(3):339–371, June 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707374.

VonLeyden:1958:AAP

[Von58] W. Von Leyden. Antiquity and authority: a paradox in the Renaissance theory of history. Journal of the History of Ideas, 19(4):473–492, October 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707918.

VonFritz:1981:AAE

[Von81] Kurt Von Fritz. Aristotle’s anthropological ethics and its relevance to modern problems. Journal of the History of Ideas, 42(2):187–207, April/June 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709316.

VonDerLuft:1984:SNG

[Von84] Eric Von Der Luft. Sources of Nietzsche’s “God is dead!” and its meaning for Heidegger. Journal of the History of Ideas, 45(2):263–276, April/June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709291.

Vorzimmer:1969:DMT

[Vor69] Peter Vorzimmer. Darwin, Malthus, and the theory of natural selection. Journal of the History of Ideas, 30(4):527–542, October/ December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708609.

Vovchenko:2012:MOR

[Vov12] Denis Vovchenko. Modernizing orthodoxy: Russia and the Christian East (1856–1914). Journal of the History of Ideas, 73(2):295– 317, April 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
475342.

Vozar:2025:AFE

[Voz25] Thomas Matthew Vozar. Academic freedom in the English Revolution: Libertas scholastica, libertas philosophandi, and the reformation of the universities. Journal of the History of Ideas, 86(1):49–73, January 2025. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
949927.

vanRuler:2000:MFS

[vR00] J. A. van Ruler. Minds, forms, and spirits: The nature of Cartesian disenchantment. Journal of the History of Ideas, 61(3):381–395, July 2000. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15084.

Vinci:2005:AMG

[VR05] Thomas C. Vinci and Jason Scott Robert. Aristotle and modern genetics. Journal of the History of Ideas, 66(2):201–221, April 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/187194.

Vucinich:1960:MRC

[Vuc60] Alexander Vucinich. Mathematics in Russian culture. Journal of the History of Ideas, 21(2):161–179, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708192.

W:1951:IN

[W.51] A. F. W. Introductory note. Journal of the History of Ideas, 12(1):31– 32, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707534.

Williamson:1997:GW

[WA97] Paul Williamson and Amal Asfour. Gainsborough’s wit. Journal of the History of Ideas, 58(3):479–501, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14950.

Wach:1946:RRS

[Wac46] Joachim Wach. The rôle of religion in the social philosophy of Alexis De Tocqueville. Journal of the History of Ideas, 7(1):74–90, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707269.

Wachbrit:1996:CSB

[Wac96] Robert Edward Wachbrit. Cartesian skepticism from bare possibility. Journal of the History of Ideas, 57(1):109–129, January 1996. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14896.

Wagner:1954:LLP

[Wag54] Vern Wagner. The lecture lyceum and the problem of controversy. Journal of the History of Ideas, 15(1):119–135, January 1954. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707652.

Wagar:1967:MVO

[Wag67] W. Warren Wagar. Modern views of the origins of the idea of progress. Journal of the History of Ideas, 28(1):55–70, January/March 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708480.

Waldman:1959:OLD

[Wal59] Theodore Waldman. Origins of the legal doctrine of reasonable doubt. Journal of the History of Ideas, 20(3):299–316, June/September 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708111.

Wallace:1963:BRC

[Wal63] John Wallace. Book review: The Cause Too Good. Journal of the History of Ideas, 24(1):150–154, January/March 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707866.

Walsh:1966:BS

[Wal66] James J. Walsh. Buridan and Seneca. Journal of the History of Ideas, 27(1):23–40, January/March 1966. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708306.

Waley:1970:PEM

[Wal70] Daniel Waley. The primitivist element in Machiavelli’s thought. Journal of the History of Ideas, 31(1):91–98, January/March 1970. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708371.

Wallace:1971:MBG

[Wal71] William A. Wallace. Mechanics from Bradwardine to Galileo. Journal of the History of Ideas, 32(1):15–28, January/March 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708322.

Walsh:1981:KCP

[Wal81] W. H. Walsh. Kant’s critique of pure reason: Commentators in English, 1875–1945. Journal of the History of Ideas, 42(4):723– 737, October/December 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709130.

Wallace:1984:SJF

[Wal84a] Dewey D. Wallace, Jr. Socinianism, justification by faith, and the sources of John Locke’s the reasonableness of Christianity. Journal of the History of Ideas, 45(1):49–66, January/March 1984. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709330.

Wallech:1984:ESS

[Wal84b] Steven Wallech. The elements of social status in Hume’s treatise. Journal of the History of Ideas, 45(2):207–218, April/June 1984. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709288.

Wallech:1986:CVR

[Wal86] Steven Wallech. “Class versus rank”: The transformation of Eighteenth-Century English social terms and theories of production. Journal of the History of Ideas, 47(3):409–431, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709661.

Wallace:1988:RRG

[Wal88] William A. Wallace. Randall redivivus: Galileo and the Paduan Aristotelians. Journal of the History of Ideas, 49(1):133–149, January/ March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709707.

Walmsley:1993:DCP

[Wal93] Peter Walmsley. Dispute and conversation: Probability and the rhetoric of natural philosophy in Locke’s essay. Journal of the History of Ideas, 54(3):381–394, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2710019.

Walhout:1996:HTA

[Wal96] Donald Walhout. The hermeneutical turn in American critical theory, 1830–1860. Journal of the History of Ideas, 57(4):683–703, October 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/14928.

Walbridge:1998:EGG

[Wal98] John Tuthill Walbridge. Explaining away the Greek gods in Islam. Journal of the History of Ideas, 59(3):389–403, July 1998. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/14998.

Walmsley:2004:LNP

[Wal04] Jonathan Walmsley. Locke’s natural philosophy in Draft A of the essay. Journal of the History of Ideas, 65(1):15–37, January 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/168340.

Wall:2007:JVJ

[Wal07] Byron Emerson Wall. John Venn, James Ward, and the chair of mental philosophy and logic at the University of Cambridge. Journal of the History of Ideas, 68(1):131–155, January 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/210692.

Wallace:2011:LTA

[Wal11] Joseph Wallace. Legal theories and ancient practices in John Selden’s Marmora Arundelliana. Journal of the History of Ideas, 72(3):393– 412, July 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
445502.

Walter:2019:DPT

[Wal19] Ryan Walter. Defending political theory after Burke: Stewart’s intellectual disciplines and the demotion of practice. Journal of the History of Ideas, 80(3):387–408, July 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/56/article/729724.

Wang:2005:ZDD

[Wan05] Robin Wang. Zhou Dunyi’s diagram of the supreme ultimate explained (taijitu shuo): a construction of the Confucian metaphysics. Journal of the History of Ideas, 66(3):307–323, July 2005. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/188687.

Warnke:1967:BRM

[War67] Frank J. Warnke. Book review: Mazzeo on the Renaissance. Journal of the History of Ideas, 28(2):289–292, April/June 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708424.

Warden:1971:MZ

[War71] J. R. Warden. The mind of Zeus. Journal of the History of Ideas, 32(1):3–14, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708321.

Ware:2011:ELW

[War11] Ben Ware. Ethics and the literary in Wittgenstein’s Tractatus Logico-Philosophicus. Journal of the History of Ideas, 72(4):595–611, October 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/453550.

Waserman:1972:BRG

[Was72] Manfred J. Waserman. Benjamin Rush on government and the harmony and derangement of the mind. Journal of the History of Ideas, 33(4):639–642, October/December 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708863.

Washell:1973:LLA

[Was73] Richard F. Washell. Logic, language, and Albert the great. Journal of the History of Ideas, 34(3):445–450, July/September 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708963.

Waszek:1984:TCM

[Was84] Norbert Waszek. Two concepts of morality: a distinction of Adam Smith’s ethics and its stoic origin. Journal of the History of Ideas, 45 (4):591–606, October/December 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709375.

Waswo:1989:MM

[Was89] Richard Waswo. Motives of misreading. Journal of the History of Ideas, 50(2):324–332, April/June 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709738.

Watson:1955:HMC

[Wat55] George Watson. Hobbes and the metaphysical conceit. Journal of the History of Ideas, 16(4):558–562, October 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707511. See reply [Gan56].

Watt:1971:LMC

[Wat71] E. D. Watt. “Locked In”: De Maistre’s critique of French Lockeanism. Journal of the History of Ideas, 32(1):129–132, January/March 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708330.

Watanabe:1972:ACC

[Wat72] Morimichi Watanabe. Authority and consent in Church government: Panormitanus, Aeneas Sylvius, Cusanus. Journal of the History of Ideas, 33(2):217–236, April/June 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708870.

Watkins:1992:MCL

[Wat92] Renee Neu Watkins. Mythology as code: Lapo Da Castiglionchio’s view of homosexuality and materialism at the Curia. Journal of the History of Ideas, 53(1):138–144, January/March 1992. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709914.

Wauters:2020:AIG

[Wau20] Bart Wauters. Aquinas, ius gentium, and the Decretists. Journal of the History of Ideas, 81(4):509–529, October 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/772518.

Wheeler-Barclay:1993:VES

[WB93] Marjorie Wheeler-Barclay. Victorian evangelicalism and the sociology of religion: The career of William Robertson Smith. Journal of the History of Ideas, 54(1):59–78, January 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709860.

Weaver:1980:EVI

[Wea80] Frederick Stirton Weaver. Exploitation vs. inequality in political theory: a critical note. Journal of the History of Ideas, 41(1): 151–154, January/March 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2709109. See reply [Mac80].

Weber:1979:BRC

[Web79] Eugen Weber. Book review: Cultures Apart. Journal of the History of Ideas, 40(3):481–490, July/September 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709253.

Weber:1981:FHF

[Web81] Eugen Weber. Fairies and hard facts: The reality of folktales. Journal of the History of Ideas, 42(1):93–113, January/March 1981. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709419.

Weber:2022:NET

[Web22] Isabella M. Weber. Neoliberal economic thinking and the quest for rational socialism in China: Ludwig von Mises and the Market Reform Debate. Journal of the History of Ideas, 83(2):333–356, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/855170.

Weinberg:1942:FLN

[Wei42] Julius R. Weinberg. The fifth letter of Nicholas of Autrecourt to Bernard of Arezzo. Journal of the History of Ideas, 3(2):220–227, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707178.

Weisinger:1945:IHD

[Wei45] Herbert Weisinger. Ideas of history during the Renaissance. Journal of the History of Ideas, 6(4):415–435, October 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707343.

Weisinger:1950:EOS

[Wei50] Herbert Weisinger. The English origins of the sociological interpretation of the Renaissance. Journal of the History of Ideas, 11(3):321– 338, June 1950. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707734.

Weigand:1952:PTF

[Wei52] Paul Weigand. Psychological types in Friedrich Schiller and William James. Journal of the History of Ideas, 13(3):376–383, June 1952. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707603.

Weinberg:1969:BRH

[Wei69] Bernard Weinberg. Book review: A History of Modern Criticism, 1750–1950 by René Wellek. Journal of the History of Ideas, 30 (1):127–133, January/March 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/ stable/2708250. See reply [Ano69l].

Weinstein:1972:BRW

[Wei72] Donald Weinstein. Book review: In Whose Image and Likeness? Interpretations of Renaissance Humanism. Journal of the History of Ideas, 33(1):165–176, January/March 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709064.

Weitzman:1975:ECL

[Wei75] Arthur J. Weitzman. Eighteenth-Century London: Urban paradise or fallen city? Journal of the History of Ideas, 36(3):469–480, July/ September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708657.

Weikart:1993:OSD

[Wei93] Richard Weikart. The origins of social Darwinism in Germany, 1859– 1895. Journal of the History of Ideas, 54(3):469–488, July 1993. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2710024.

Weinstein:1996:NLH

[Wei96] David Weinstein. The new liberalism of L. T. Hobhouse and the reenvisioning of Nineteenth-Century utilitarianism. Journal of the History of Ideas, 57(3):487–507, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14917.

Weikart:2002:DDD

[Wei02] Richard Weikart. Darwinism and death: Devaluing human life in Germany 1859–1920. Journal of the History of Ideas, 63(2):323–344, April 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15158.

Weinberg:2016:HKN

[Wei16] Joanna Weinberg. A humanist in the Kloyz: New perspectives on the Maharal of Prague and Jacques Bongars. Journal of the History of Ideas, 77(4):521–537, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/635626.

Weidner:2019:HDS

[Wei19a] Daniel Weidner. The history of dogma and the story of modernity: The modern age as “Second Overcoming of Gnosticism”. Journal of the History of Ideas, 80(1):75–90, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/715737.

Weinreich:2019:HBB

[Wei19b] Spencer J. Weinreich. Hagiography by the book: Bibliomancy and early modern cultures of compilation in Francisco Zumel’s De vitis patrum (1588). Journal of the History of Ideas, 80(1):1–23, January 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/715733.

Welter:1955:IPA

[Wel55] Rush Welter. The idea of progress in America: an essay in ideas and method. Journal of the History of Ideas, 16(3):401–415, June 1955.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707640.

Wells:1958:HD

[Wel58] G. A. Wells. Herder’s determinism. Journal of the History of Ideas, 19(1):105–113, January 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707956.

Wells:1960:HTP

[Wel60] G. A. Wells. Herder’s two philosophies of history. Journal of the History of Ideas, 21(4):527–537, October/December 1960. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708100.

Wells:1967:GE

[Wel67] George A. Wells. Goethe and evolution. Journal of the History of Ideas, 28(4):537–550, October/December 1967. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708528.

Wells:1969:SNT

[Wel69] George A. Wells. Stages of New Testament criticism. Journal of the History of Ideas, 30(2):147–160, April/June 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708429.

Wells:1971:GQO

[Wel71] George A. Wells. Goethe’s qualitative optics. Journal of the History of Ideas, 32(4):617–626, October/December 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708983.

Wells:1973:FSC

[Wel73] George A. Wells. Friedrich Solmsen on Christian origins. Journal of the History of Ideas, 34(1):143–144, January/March 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708950. See reply [Sol73].

Wellman:1976:SMX

[Wel76] Robert R. Wellman. Socratic method in xenophon. Journal of the History of Ideas, 37(2):307–318, April/June 1976. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708826.

Wellek:1984:BRB

[Wel84] René Wellek. Book review: Between Kant and Fichte: Karl Leonhard Reinhold. Journal of the History of Ideas, 45(2):323–327, April/ June 1984. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709296.

Wells:1994:PMO

[Wel94] Harwell Wells. The philosophical Michael Oakeshott. Journal of the History of Ideas, 55(1):129–145, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709957.

Wenz:1976:BCN

[Wen76] Peter S. Wenz. Berkeley’s Christian neo-Platonism. Journal of the History of Ideas, 37(3):537–546, July/September 1976. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708814.

Wenz:1982:BTC

[Wen82] Peter S. Wenz. Berkeley’s two concepts of impossibility: a reply to Mckim. Journal of the History of Ideas, 43(4):673–680, October/ December 1982. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709350.

See [McK82b].

Wen:2024:WJM

[Wen24] Xinyi Wen. When Jupiter meets Saturn: Aby Warburg, Karl Sudhoff, and astrological medicine in the Age of Disenchantment. Journal of the History of Ideas, 85(2):321–355, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/926151.

Werner:1972:DHA

[Wer72] John M. Werner. David Hume and America. Journal of the History of Ideas, 33(3):439–456, July/September 1972. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709045.

Wertz:1975:HHH

[Wer75] S. K. Wertz. Hume, history, and human nature. Journal of the History of Ideas, 36(3):481–496, July/September 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708658.

Wertz:1993:HHS

[Wer93] S. K. Wertz. Hume and the historiography of science. Journal of the History of Ideas, 54(3):411–436, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2710021.

Werner:2013:MJM

[Wer13] Meike G. Werner. Modern Jena as a model of cultural regeneration in Wilhelmine Germany. Journal of the History of Ideas, 74(2):267– 288, April 2013. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
504788.

West:1953:MAH

[Wes53] Robert H. West. Milton’s angelological heresies. Journal of the History of Ideas, 14(1):116–123, January 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707498.

Weston:1958:IPB

[Wes58] John C. Weston, Jr. The ironic purpose of Burke’s vindication vindicated. Journal of the History of Ideas, 19(3):435–441, June 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708048.

Westfall:1969:PLA

[Wes69] Carroll W. Westfall. Painting and the liberal arts: Alberti’s view. Journal of the History of Ideas, 30(4):487–506, October/December 1969. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708607.

Westfall:1976:BRC

[Wes76] Richard S. Westfall. Book review: The Changing World of the Newtonian Industry. Journal of the History of Ideas, 37(1):175–184, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708719.

Westgate:1977:GHW

[Wes77] Sam Westgate. George Herbert: “Wit’s an Unruly Engine”. Journal of the History of Ideas, 38(2):281–296, April/June 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708912.

Westfall:1987:NSP

[Wes87] Richard S. Westfall. Newton’s scientific personality. Journal of the History of Ideas, 48(4):551–570, October/December 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709687.

Westerhoff:1999:PCH

[Wes99] Jan C. Westerhoff. Poeta calculans: Harsdorffer, Leibniz, and the mathesis universalis. Journal of the History of Ideas, 60(3):449–467, July 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/15043.

Westerhoff:2001:WSB

[Wes01] Jan C. Westerhoff. A world of signs: Baroque pansemioticism, the polyhistor and the Early Modern Wunderkammer. Journal of the History of Ideas, 62(4):633–650, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/15138.

Weststeijn:2007:SSA

[Wes07] Thijs Weststeijn. Spinoza sinicus: an Asian paragraph in the history of the radical Enlightenment. Journal of the History of Ideas, 68(4):537–561, October 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/223136.

Weststeijn:2011:PPS

[Wes11] Arthur Weststeijn. The power of “Pliant Stuff”: Fables and frankness in Seventeenth-Century Dutch republicanism. Journal of the History of Ideas, 72(1):1–27, January 2011. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/413472.

Westler:2015:BTR

[Wes15] Brendon Westler. Between tradition and revolution: The curious case of Francisco Martínez Marina, the Cádiz Constitution, and Spanish liberalism. Journal of the History of Ideas, 76(3):393–416, July 2015. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/587271.

Wettersten:2005:NIY

[Wet05] John Wettersten. New insights on young Popper. Journal of the History of Ideas, 66(4):603–631, October 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/191478.

Wheeler:1955:SSG

[Whe55] Marcus Wheeler. Self-sufficiency and the Greek city. Journal of the History of Ideas, 16(3):416–420, June 1955. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707641.

White:1942:CMC

[Whi42] Lynn White, Jr. Christian myth and Christian history. Journal of the History of Ideas, 3(2):145–158, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707174.

White:1945:BRS

[Whi45a] Morton G. White. Book review: Social Darwinism in American Thought, 1860–1915 by Richard Hofstadter. Journal of the History of Ideas, 6(1):119–122, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707063.

Whitmore:1945:MMH

[Whi45b] Charles E. Whitmore. Mill and mathematics: an historical note. Journal of the History of Ideas, 6(1):109–112, January 1945. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707061.

White:1947:RAF

[Whi47] Morton G. White. The revolt against formalism in American social thought of the Twentieth Century. Journal of the History of Ideas, 8(2):131–152, April 1947. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707153.

White:1960:GIS

[Whi60] Hayden V. White. The Gregorian ideal and Saint Bernard of Clairvaux. Journal of the History of Ideas, 21(3):321–348, July/September 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708141.

White:1971:CVA

[Whi71] Donald A. White. Changing views of the Adventus Saxonum in Nineteenth and Twentieth Century English scholarship. Journal of the History of Ideas, 32(4):585–594, October/December 1971. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708979.

White:1976:BRA

[Whi76a] Donald White. Book review: Adam Smith’s Wealth of Nations. Journal of the History of Ideas, 37(4):715–720, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709034.

White:1976:PHI

[Whi76b] Lucia White. On a passage by Hume incorrectly attributed to Jefferson. Journal of the History of Ideas, 37(1):133–135, January/March 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708714.

Whitman:1986:NMT

[Whi86] James Whitman. Nietzsche in the magisterial tradition of German classical philology. Journal of the History of Ideas, 47(3): 453–468, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709663.

Whitney:1989:FBI

[Whi89] Charles Whitney. Francis Bacon’s instauratio: Dominion of and over humanity. Journal of the History of Ideas, 50(3):371–390, July/ September 1989. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709567.

Whitmer:2009:ETD

[Whi09] Kelly J. Whitmer. Eclecticism and the technologies of discernment in pietist pedagogy. Journal of the History of Ideas, 70(4):545–567, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/361799.

Wiener:1940:LPP

[Wie40] Philip P. Wiener. Leibniz’s project of a public exhibition of scientific inventions. Journal of the History of Ideas, 1(2):232–240, April 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707335.

Wiener:1943:BRP

[Wie43] Philip P. Wiener. Book review: Physics and Philosophy by James Jeans. Journal of the History of Ideas, 4(4):484–489, October 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707170.

Wiener:1945:CWD

[Wie45] Philip P. Wiener. Chauncey Wright’s defense of Darwin and the neutrality of science. Journal of the History of Ideas, 6(1):19–45, January 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707054. See errata

[Ano45f].

Wiener:1946:EPP

[Wie46a] Philip P. Wiener. The evolutionism and pragmaticism of Peirce. Journal of the History of Ideas, 7(3):321–350, June 1946. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707404. See correction [Ano46g].

Wiener:1946:LSH

[Wie46b] Philip P. Wiener. Logical significance of the history of thought. Journal of the History of Ideas, 7(3):366–373, June 1946. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707408.

Wiener:1946:PMC

[Wie46c] Philip P. Wiener. Peirce’s metaphysical club and the genesis of pragmatism. Journal of the History of Ideas, 7(2):218–233, April 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707072.

Wiener:1948:PLP

[Wie48] Philip P. Wiener. The pragmatic legal philosophy of N. St. John Green (1830–76). Journal of the History of Ideas, 9(1):70–92, January 1948. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707396.

Wiener:1951:BRC

[Wie51] Philip P. Wiener. Book review: Cassirer on the Problem of Knowledge. Journal of the History of Ideas, 12(2):305–309, April 1951. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707521.

Wiener:1956:GMB

[Wie56] Philip P. Wiener. G. M. Beard and Freud on ’American nervousness’. Journal of the History of Ideas, 17(2):269–274, April 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707747.

Wiener:1961:SPM

[Wie61] Philip P. Wiener. Some problems and methods in the history of ideas. Journal of the History of Ideas, 22(4):531–548, October/December 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708029.

Wiener:1963:JGP

[Wie63] Philip P. Wiener. James Gregory: On power. Journal of the History of Ideas, 24(2):241–268, April/June 1963. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707848.

Wiener:1968:BRE

[Wie68] Philip P. Wiener. Book review: The Encyclopedia of Philosophy by Paul Edwards. Journal of the History of Ideas, 29(4):616–622, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708299.

Wiener:1970:BRC

[Wie70] Philip P. Wiener. Book review: Comments on Professor Edwards’ Reply. Journal of the History of Ideas, 31(1):146–147, January/March 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708378.

Wiener:1973:TCC

[Wie73] Philip P. Wiener. Towards commemorating the centenary of Arthur O. Lovejoy’s birthday (October 10, 1873). Journal of the History of Ideas, 34(4):591–598, October/December 1973. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708890.

Wiener:1975:BRT

[Wie75] Philip P. Wiener. Book review: Three Recent Books on Pragmatism. Journal of the History of Ideas, 36(4):739–746, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709001.

Wiener:1980:MGB

[Wie80a] Philip P. Wiener. In memoriam: George Boas (1891–1980). Journal of the History of Ideas, 41(3):453–456, July/September 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709404.

Wiener:1980:MHM

[Wie80b] Philip P. Wiener. In memoriam: Howard Mumford Jones (1892– 1980). Journal of the History of Ideas, 41(3):457–458, July/September 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709405.

Wiener:1980:OHB

[Wie80c] Philip P. Wiener. Obituary: Herbert Butterfield (1900–1979). Journal of the History of Ideas, 41(1):157–158, January/March 1980. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709111.

Wiener:1986:RER

[Wie86] Philip P. Wiener. Remembering Edward Rosen. Journal of the History of Ideas, 47(1):159–161, January/March 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2709602.

Wigelsworth:2009:SCN

[Wig09] Jeffrey R. Wigelsworth. Samuel Clarke’s Newtonian soul. Journal of the History of Ideas, 70(1):45–68, January 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/258194.

Wilson:1941:SMN

[Wil41] Harold S. Wilson. Some meanings of “Nature” in Renaissance literary theory. Journal of the History of Ideas, 2(4):430–448, October 1941. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707021.

Wiley:1948:STB

[Wil48] Margaret L. Wiley. Sir Thomas Browne and the genesis of paradox. Journal of the History of Ideas, 9(3):303–322, June 1948. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707372.

Wilhelm:1951:PWC

[Wil51] Hellmut Wilhelm. The problem of within and without, a Confucian attempt in syncretism. Journal of the History of Ideas, 12(1):48–60, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707536.

Wilkins:1956:JDH

[Wil56] Burleigh Taylor Wilkins. James, Dewey, and Hegelian idealism. Journal of the History of Ideas, 17(3):332–346, June 1956. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707548.

Wilkins:1959:SNB

[Wil59] Burleigh Taylor Wilkins. Some notes on burckhardt. Journal of the History of Ideas, 20(1):123–137, January 1959. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707971.

Williams:1960:CET

[Wil60] Gwyn A. Williams. The concept of ’egemonia’ in the thought of Antonio Gramsci: Some notes on interpretation. Journal of the History of Ideas, 21(4):586–599, October/December 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708106.

Wilson:1965:DT

[Wil65] John B. Wilson. Darwin and the transcendentalists. Journal of the History of Ideas, 26(2):286–290, April/June 1965. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708235.

Willard:1966:MTL

[Wil66a] Charity Cannon Willard. The manuscript tradition of the Livre Des Trois Vertus and Christine de Pizan’s audience. Journal of the History of Ideas, 27(3):433–444, July/September 1966. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708596.

Wilson:1966:AT

[Wil66b] John B. Wilson. An analogue of transcendentalism. Journal of the History of Ideas, 27(3):459–461, July/September 1966. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2708599.

Wilson:1967:LLF

[Wil67] Margaret D. Wilson. Leibniz and Locke on “First Truths”. Journal of the History of Ideas, 28(3):347–366, July/September 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708623.

Wilson:1968:HTM

[Wil68] Stephen Wilson. History and traditionalism: Maurras and the action française. Journal of the History of Ideas, 29(3):365–380, July/ September 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708448.

Williams:1970:RSS

[Wil70a] Robert C. Williams. The Russian soul: a study in European thought and non-European nationalism. Journal of the History of Ideas, 31 (4):573–588, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708261.

Willman:1970:HLH

[Wil70b] Robert Willman. Hobbes on the law of heresy. Journal of the History of Ideas, 31(4):607–613, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708264.

Wilson:1971:FS

[Wil71] David S. Wilson. The flying spider. Journal of the History of Ideas, 32(3):447–458, July/September 1971. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2708360.

Wilson:1973:FCA

[Wil73] Stephen Wilson. Fustel De Coulanges and the Action Française. Journal of the History of Ideas, 34(1):123–134, January/March 1973. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708948.

Wilson:1974:NSP

[Wil74a] Curtis Wilson. Newton and some philosophers on Kepler’s “Laws”. Journal of the History of Ideas, 35(2):231–258, April/June 1974. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708760.

Wilson:1974:HWV

[Wil74b] David B. Wilson. Herschel and Whewell’s version of Newtonianism. Journal of the History of Ideas, 35(1):79–97, January/March 1974. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708744.

Williford:1975:BRW

[Wil75] Miriam Williford. Bentham on the rights of women. Journal of the History of Ideas, 36(1):167–176, January/March 1975. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709019.

Wilson:1980:ALM

[Wil80] Daniel J. Wilson. Arthur O. Lovejoy and the moral of the great chain of being. Journal of the History of Ideas, 41(2):249–265, April/ June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2709459.

Wilson:1985:HWV

[Wil85] Katharina M. Wilson. The history of the word “Vampire”. Journal of the History of Ideas, 46(4):577–583, October/December 1985. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709546.

Wilson:1987:LGC

[Wil87] Daniel J. Wilson. Lovejoy’s the great chain of being after fifty years. Journal of the History of Ideas, 48(2):187–206, April/June 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709553.

Wilson:1988:VSV

[Wil88] Catherine Wilson. Visual surface and visual symbol: The microscope and the occult in early modern science. Journal of the History of Ideas, 49(1):85–108, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709705.

Wilberding:1991:DDA

[Wil91] Erick Wilberding. A defense of Dionysius the Areopagite by Rubens. Journal of the History of Ideas, 52(1):19–34, January/March 1991. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709580.

Williams:1996:WPN

[Wil96] Linda L. Williams. Will to power in Nietzsche’s published works and the Nachlass. Journal of the History of Ideas, 57(3):447–463, July 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://muse.jhu.edu/pub/1/article/14914.

Williston:1997:DLE

[Wil97] Byron Williston. Descartes on love and/as error. Journal of the History of Ideas, 58(3):429–444, July 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/14958.

Williams:2005:JGW

[Wil05] David Lay Williams. Justice and the general will: Affirming Rousseau’s ancient orientation. Journal of the History of Ideas, 66(3): 383–411, July 2005. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
188688.

Wilson:2008:UPP

[Wil08] Andrew D. Wilson. The unity of physics and poetry: H. C. Ørsted and the aesthetics of force. Journal of the History of Ideas, 69(4):627– 646, October 2008. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
251813.

Wilkin:2019:FNR

[Wil19] Rebecca Wilkin. Feminism and natural right in François poulain de la barre and gabrielle suchon. Journal of the History of Ideas, 80(2): 227–248, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
722379.

Willson:2020:MLS

[Wil20] Justin Willson. A meadow that lifts the soul: Originality as anthologizing in the Byzantine Church interior. Journal of the History of Ideas, 81(1):1–21, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

56/article/747174.

Wilner:2022:BKPa

[Wil22a] Isaiah Lorado Wilner. Body knowledge, Part I: Dance, anthropology, and the erasure of history. Journal of the History of Ideas, 83(1):111– 142, January 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
847823.

Wilner:2022:BKPb

[Wil22b] Isaiah Lorado Wilner. Body knowledge, Part II: Motion, memory, and the mythology of modernity. Journal of the History of Ideas, 83(2): 229–255, April 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
855165.

Williamson:2023:ACE

[Wil23] Athanassia Williamson. Art as critical experience in Theodor W. Adorno and John Dewey. Journal of the History of Ideas, 84(3):511– 532, July 2023. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
901492.

Winnik:1970:SMT

[Win70] Herbert C. Winnik. Science and morality in Thomas C. Chamberlin. Journal of the History of Ideas, 31(3):441–456, July/September 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708516.

Winfrey:1981:CVJ

[Win81] John C. Winfrey. Charity versus justice in Locke’s theory of property. Journal of the History of Ideas, 42(3):423–438, July/September 1981. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709185.

Wish:1949:APM

[Wis49] Harvey Wish. Aristotle, Plato, and the Mason–Dixon Line. Journal of the History of Ideas, 10(2):254–266, April 1949. CODEN JHIDAT.

ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707417.

Wise:1968:IIP

[Wis68] Gene Wise. Implicit irony in Perry Miller’s New England mind. Journal of the History of Ideas, 29(4):579–600, October/December 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708295.

Wisner:1997:ECH

[Wis97] David A. Wisner. Ernst Cassirer, historian of the will. Journal of the History of Ideas, 58(1):145–161, January 1997. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/14937.

Wittkowsky:1943:SMP

[Wit43] George Wittkowsky. Swift’s modest proposal: The biography of an Early Georgian pamphlet. Journal of the History of Ideas, 4(1):75– 104, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707237.

Wittkower:1961:IAA

[Wit61] Rudolf Wittkower. Individualism in art and artists: a Renaissance problem. Journal of the History of Ideas, 22(3):291–302, July/ September 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708126.

Witt:1977:SCP

[Wit77] Ronald G. Witt. Salutati and contemporary physics. Journal of the History of Ideas, 38(4):667–672, October/December 1977. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708694.

Withers:2009:PST

[Wit09] Charles W. J. Withers. Place and the “Spatial Turn” in geography and in history. Journal of the History of Ideas, 70(4):637–658, October 2009. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/361795.

Wolfenstein:1944:SBT

[Wol44] Martha Wolfenstein. The social background of Taine’s philosophy of art. Journal of the History of Ideas, 5(3):332–358, June 1944.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707354.

Wolstein:1949:AWM

[Wol49] Benjamin Wolstein. Addison Webster Moore: Defender of instrumentalism. Journal of the History of Ideas, 10(4):539–566, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707189.

Wolstein:1950:NFT

[Wol50] Benjamin Wolstein. A note on the functional theory of habit. Journal of the History of Ideas, 11(4):490–492, October 1950. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707594.

Wolstein:1953:RSA

[Wol53] Benjamin Wolstein. The romantic Spinoza in America. Journal of the History of Ideas, 14(3):439–450, June 1953. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707811.

Wolfson:1956:FM

[Wol56] Philip J. Wolfson. Friedrich Meinecke (1862–1954). Journal of the History of Ideas, 17(4):511–525, October 1956. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707785.

Wollheim:1958:D

[Wol58] Richard Wollheim. Democracy. Journal of the History of Ideas, 19(2): 225–242, April 1958. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2707936.

Wolff:1960:KDH

[Wol60] Robert P. Wolff. Kant’s debt to Hume via Beattie. Journal of the History of Ideas, 21(1):117–123, January/March 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708003.

Wolfson:1961:EII

[Wol61] Harry A. Wolfson. Extradeical and intradeical interpretations of platonic ideas. Journal of the History of Ideas, 22(1):3–32, January/ March 1961. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707871.

Wolff:1968:LRP

[Wol68] Cynthia Griffin Wolff. Literary reflections of the Puritan character. Journal of the History of Ideas, 29(1):13–32, January/March 1968. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2708463.

Wolper:1970:RGI

[Wol70] Roy S. Wolper. The rhetoric of gunpowder and the idea of progress. Journal of the History of Ideas, 31(4):589–598, October/December 1970. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2708262.

Wolf:1976:HKL

[Wol76a] Ken Wolf. Hans Kohn’s liberal nationalism: The historian as prophet. Journal of the History of Ideas, 37(4):651–672, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709029.

Woll:1976:PHN

[Wol76b] Allen L. Woll. Positivism and history in Nineteenth-Century Chile: José victorino lastarria and valentín letelier. Journal of the History of Ideas, 37(3):493–506, July/September 1976. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708811.

Wolloch:2000:CHA

[Wol00] Nathaniel Wolloch. Christiaan Huygens’s attitude toward animals. Journal of the History of Ideas, 61(3):415–432, July 2000. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/1/article/15086.

Wolloch:2007:FCA

[Wol07] Nathaniel Wolloch. ”Facts, or conjectures”: Antoine-Yves Goguet’s historiography. Journal of the History of Ideas, 68(3):429–449, July 2007. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/219826.

Wolin:2016:ISJ

[Wol16a] Richard Wolin. Introduction to the symposium on Jonathan Israel’s democratic Enlightenment. Journal of the History of Ideas, 77(4):615– 626, October 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
635630.

Wolin:2016:SJG

[Wol16b] Richard Wolin. Symposium on J. G. A. Pocock’s barbarism and religion: Introduction. Journal of the History of Ideas, 77(1):99–106, January 2016. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
610562.

Womersley:1986:HWW

[Wom86] D. J. Womersley. The historical writings of William Robertson. Journal of the History of Ideas, 47(3):497–506, July/September 1986. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709666.

Woodman:1960:SCA

[Woo60] Ross G. Woodman. Shelley’s changing attitude to Plato. Journal of the History of Ideas, 21(4):497–510, October/December 1960. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708098.

Wood:1967:FPS

[Woo67] Neal Wood. Frontinus as a possible source for Machiavelli’s method. Journal of the History of Ideas, 28(2):243–248, April/June 1967. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2708419.

Woolf:1988:CLS

[Woo88] D. R. Woolf. Community, law and state: Samuel Daniel’s historical thought revisited. Journal of the History of Ideas, 49(1):61–83, January/March 1988. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709704.

Wood:1992:TRS

[Woo92] Neal Wood. Tabula Rasa, social environmentalism, and the “English Paradigm”. Journal of the History of Ideas, 53(4):647–668, October/ December 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709942.

Wood:1994:LBN

[Woo94] Joanne A. Wood. Lighthouse bodies: The neutral monism of Virginia Woolf and Bertrand Russell. Journal of the History of Ideas, 55(3):

483–502, July 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2709851.

Woodhouse:2018:SSI

[Woo18] Adam Woodhouse. Subjection without servitude:: The Imperial protectorate in Renaissance political thought. Journal of the History of Ideas, 79(4):547–569, October 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/56/article/707772.

Woolridge:2019:SVL

[Woo19] Paul Andrew Woolridge. Scrutiny’s virtue: Leavis, MacIntyre, and the case for tradition. Journal of the History of Ideas, 80(2):289– 311, April 2019. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
722382.

Woodford:2024:RWO

[Woo24] Benjamin Woodford. “The right we have to our owne bodies, goods, and liberties”: The freedom of the Ancient Constitution and Common Law in Milton’s early prose. Journal of the History of Ideas, 85(1):41– 63, January 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
917115.

Wormhoudt:1949:NNP

[Wor49] Arthur Wormhoudt. Newton’s natural philosophy in the behmenistic works of William law. Journal of the History of Ideas, 10(3):411–429, June 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222

(electronic). URL https://www.jstor.org/stable/2707045.

Wright:1940:SRW

[Wri40] Louis B. Wright. The significance of religious writings in the English Renaissance. Journal of the History of Ideas, 1(1):59–68, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707010.

Wright:1943:NSM

[Wri43] Louis B. Wright. The noble savage of Madagascar in 1640. Journal of the History of Ideas, 4(1):112–118, January 1943. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707239.

Wright:1949:BRP

[Wri49] Arthur F. Wright. Book review: Professor Northrop’s Chapter on the Traditional Culture of the Orient. Journal of the History of Ideas, 10(1):143–149, January 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707207.

Wright:1951:FRB

[Wri51] Arthur F. Wright. Fu I and the rejection of Buddhism. Journal of the History of Ideas, 12(1):33–47, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707535.

Wright:1960:SCC

[Wri60] Arthur F. Wright. The study of Chinese civilization. Journal of the History of Ideas, 21(2):233–255, April/June 1960. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2708197.

Wright:1980:HMM

[Wri80] John P. Wright. Hysteria and mechanical man. Journal of the History of Ideas, 41(2):233–247, April/June 1980. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2709458.

Wunder:2003:CCM

[Wun03] Amanda Jaye Wunder. Classical, Christian, and Muslim remains in the construction of Imperial Seville (1520–1635). Journal of the History of Ideas, 64(2):195–212, April 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.

edu/pub/1/article/43517.

Wiener:1950:BRC

[WW50] Philip P. Wiener and Philip P. Wiener. Book review: Comment. Journal of the History of Ideas, 11(2):246–247, April 1950. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707250.

Wyman:1949:CMW

[Wym49] Mary Wyman. Chinese mysticism and Wordsworth. Journal of the History of Ideas, 10(4):517–538, October 1949. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707188.

Yarrow:1942:FFD

[Yar42] Clarence H. Yarrow. The forging of fascist doctrine. Journal of the History of Ideas, 3(2):159–181, April 1942. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2707175.

Yarbrough:1986:JER

[Yar86] Stephen R. Yarbrough. Jonathan Edwards on rhetorical authority. Journal of the History of Ideas, 47(3):395–408, July/September 1986. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2709660.

Yeo:1996:ECC

[Yeo96] Richard R. Yeo. Ephraim Chambers’s Cyclopaedia (1728) and the tradition of commonplaces. Journal of the History of Ideas, 57(1):157– 175, January 1996. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/1/article/
14897.

Yeo:2003:SMB

[Yeo03] Richard R. Yeo. A solution to the multitude of books: Ephraim Chambers’s Cyclopaedia (1728) as “the best book in the universe”. Journal of the History of Ideas, 64(1):61–72, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/42181.

Yin:2022:RRA

[Yin22] Shoufu Yin. Redefining reciprocity: Appointment edicts and political thought in Medieval China. Journal of the History of Ideas, 83(4):533– 554, October 2022. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
865891.

Yoffie:2004:CJC

[Yof04] Adina M. Yoffie. Cocceius and the Jewish commentators. Journal of the History of Ideas, 65(3):393–398, July 2004. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/1/article/177670.

Yolton:1951:LUM

[Yol51] John W. Yolton. Locke’s unpublished marginal replies to John sergeant. Journal of the History of Ideas, 12(4):528–559, October 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2707485.

Yolton:1955:LSC

[Yol55] John W. Yolton. Locke and the Seventeenth-Century logic of ideas. Journal of the History of Ideas, 16(4):431–452, October 1955. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2707503.

Yolton:1979:LGP

[Yol79a] John W. Yolton. As in a looking-glass: Perceptual acquaintance in Eighteenth-Century Britain. Journal of the History of Ideas, 40(2):207–234, April/June 1979. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709149.

Yolton:1979:HAB

[Yol79b] John W. Yolton. Hume’s abstract in the bibliothéque raisonnée. Journal of the History of Ideas, 40(1):157–158, January/March 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709268.

Yolton:2001:LM

[Yol01] John W. Yolton. Locke’s man. Journal of the History of Ideas, 62(4):665–683, October 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/15139.

Yost:1951:LRH

[Yos51] R. M. Yost, Jr. Locke’s rejection of hypotheses about sub-microscopic events. Journal of the History of Ideas, 12(1):111–130, January 1951. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707540.

Young:1975:LDM

[You75] David B. Young. Libertarian demography: Montesquieu’s essay on depopulation in the lettres persanes. Journal of the History of Ideas, 36(4):669–682, October/December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.

org/stable/2708995.

Yousef:2001:SSW

[You01] Nancy Yousef. Savage or solitary?: The wild child and Rousseau’s man of nature. Journal of the History of Ideas, 62(2):245–263, April 2001. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/15118.

Yolton:1984:LST

[YY84] Jean S. Yolton and John W. Yolton. Locke’s suggestion of thinking matter and some Eighteenth-Century Portuguese reactions. Journal of the History of Ideas, 45(2):303–307, April/June 1984. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709294.

Zagorin:1990:HOM

[Zag90] Perez Zagorin. Hobbes on our mind. Journal of the History of Ideas, 51(2):317–335, April/June 1990. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709519.

Zagorin:1993:HEP

[Zag93] Perez Zagorin. Hobbes’s early philosophical development. Journal of the History of Ideas, 54(3):505–518, July 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2710026.

Zagorin:1999:TBT

[Zag99] Perez Zagorin. Two books on Thomas Hobbes. Journal of the History of Ideas, 60(2):361–371, April 1999. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15027.

Zagorin:2003:LPB

[Zag03] Perez Zagorin. Looking for pieter bruegel. Journal of the History of Ideas, 64(1):73–96, January 2003. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/pub/

1/article/42182.

Zahora:2012:TSP

[Zah12] Tomas Zahora. Thomist scholarship and plagiarism in the Early Enlightenment: Jacques Echard reads the Speculum morale, attributed to Vincent of Beauvais. Journal of the History of Ideas, 73(4):515– 536, October 2012. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
488346.

Zakariya:2024:HDC

[Zak24] Nasser Zakariya. Historicizing a dream of complete science. Journal of the History of Ideas, 85(2):357–388, April 2024. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.

jhu.edu/pub/56/article/926153.

Zaleski:1985:SPP

[Zal85] Carol G. Zaleski. St. Patrick’s purgatory: Pilgrimage motifs in a medieval otherworld vision. Journal of the History of Ideas, 46(4): 467–485, October/December 1985. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2709540.

Zaller:1993:FTE

[Zal93] Robert Zaller. The figure of the tyrant in English Revolutionary thought. Journal of the History of Ideas, 54(4):585–610, October 1993. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709825.

Zambelli:1998:PTF

[Zam98] Paola Zambelli. Present trends of French philosophical thought: Introduction. Journal of the History of Ideas, 59(3):521–530, July 1998. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/1/article/14999.

Zanoni:1979:DLP

[Zan79] C. P. Zanoni. Development of logical pragmatism in Italy. Journal of the History of Ideas, 40(4):603–619, October/December 1979. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709361.

Zanier:1987:PTR

[Zan87] Giancarlo Zanier. Platonic trends in Renaissance medicine. Journal of the History of Ideas, 48(3):509–519, July/September 1987. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://www.jstor.org/stable/2709765.

Zea:1959:HIL

[Zea59] Leopoldo Zea. History of ideas in Latin America: Recent works. Journal of the History of Ideas, 20(4):596–600, October/December 1959.

CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://www.jstor.org/stable/2707895.

Zebrowski:1994:RPB

[Zeb94] Martha K. Zebrowski. Richard Price: British platonist of the Eighteenth Century. Journal of the History of Ideas, 55(1):17–35, January 1994. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://www.jstor.org/stable/2709951.

Zeeveld:1946:SET

[Zee46] W. Gordon Zeeveld. Social equalitarianism in a Tudor crisis. Journal of the History of Ideas, 7(1):35–55, January 1946. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707267.

Zetterberg:1982:ENS

[Zet82] J. Peter Zetterberg. Echoes of nature in Salomon’s house. Journal of the History of Ideas, 43(2):179–193, April/June 1982. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2709198.

Zhang:2025:MWR

[Zha25] Chunjie Zhang. Max Weber and the re-enchantment of charisma. Journal of the History of Ideas, 86(1):169–192, January 2025. CO-DEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL

https://muse.jhu.edu/pub/56/article/949931.

Zhu:2002:WIF

[Zhu02] Rui Zhu. What if the father commits a crime? Journal of the History of Ideas, 63(1):1–17, January 2002. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://muse.jhu.edu/

pub/1/article/15148.

Zhulina:2021:TMR

[Zhu21] Alisa Zhulina. The tyrant and the martyr: Recent research on sovereignty and theater. Journal of the History of Ideas, 82(2):329– 349, April 2021. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
790206.

Zilsel:1940:CM

[Zil40] Edgar Zilsel. Copernicus and mechanics. Journal of the History of Ideas, 1(1):113–118, January 1940. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.jstor.org/

stable/2707014.

Zilsel:1941:OWG

[Zil41] Edgar Zilsel. The origins of William Gilbert’s scientific method. Journal of the History of Ideas, 2(1):1–32, January 1941. CODEN JHI-DAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https:/

/www.jstor.org/stable/2707279.

Zilsel:1945:GCS

[Zil45] Edgar Zilsel. The genesis of the concept of scientific progress. Journal of the History of Ideas, 6(3):325–349, June 1945. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://www.

jstor.org/stable/2707296.

Zammito:2010:JGH

[ZMM10] John H. Zammito, Karl Menges, and Ernest A. Menze. Johann Gottfried Herder revisited: The revolution in scholarship in the last quarter century. Journal of the History of Ideas, 71(4):661–684, October 2010. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (elec-

tronic). URL https://muse.jhu.edu/pub/56/article/398302.

Zon:2014:STS

[Zon14] Bennett Zon. Science, theology, and the simplicity of chant: Victorian musicology at war. Journal of the History of Ideas, 75(3):439– 469, July 2014. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
549330.

Zubovich:2018:APE

[Zub18] Gene Zubovich. American Protestants and the era of anti-racist human rights. Journal of the History of Ideas, 79(3):427–443, July 2018. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic).

URL https://muse.jhu.edu/pub/56/article/703667.

Zuckerman:1975:RTU

[Zuc75] Charles Zuckerman. The relationship of theories of universals to theories of Church government in the Middle Ages: a critique of previous views. Journal of the History of Ideas, 36(4):579–594, October/ December 1975. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://www.jstor.org/stable/2708990.

Zwierlein:2020:OCC

[Zwi20] Cornel Zwierlein. Orient contra China: Eusèbe Renaudot’s vision of world history (ca. 1700). Journal of the History of Ideas, 81(1):23– 44, January 2020. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-

3222 (electronic). URL https://muse.jhu.edu/pub/56/article/
747175.